A Complete Bibliography of the journal family The Two-Year College Mathematics Journal (1970–1983) and The College Mathematics

Journal (1984–date)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

26 February 2025 Version 1.11

Title word cross-reference

((p − 1)/2)^{2} [Lee24]. (0, π/2) [Li13a, Li13b]. (1 + 1/n)^{n} → e [Bad82]. (1 × 2) + (2 × 3) + … + n × (n + 1) = (1/3)[(n + 1)^{3} − (n + 1)] [AM87]. (fg)′ = f′g′ [MS76]. (n + 1/2) [Her12]. (x^{n})′ = nx^{n−1} [Hen90]. −1 [CMS20, Sch73]. −5 [LD15]. 1 [ASR19]. 1/(a + b) = 1/a + 1/b [BU02]. $11.95 [Hop18a]. 11/2 [Wat89]. $12.95 [Hop19]. 180^{∘} [Nak86]. 1^{2} [Lee24]. $20.40 [Gut19]. $24.95 [Gra18]. 25 [Ano95c]. 2^{2} [Lee24]. 2 × 2 [JK98]. 2 × n [Bro23]. 3 [HK19]. 4p [Ger00]. $50.00 [Kat18]. $54.00 [Ser17]. $56.95 [Gra18]. $7.99 [Hop18a]. [0, π] [FZ12]. ^{3} [SS83b]. ^{a} [Var76]. ^{b} [Var76]. ^{k} [Car87]. ^{x} [Gor84]. A [MM93]. A > B [Ros87]. A^{B} [Ros87]. aa^{T} + bb^{T} [TT03]. a · a′ + b · b′ = c · c′ [Gen89]. Arcsecx [Sch86c]. Arctan [Har87]. Arctanx [Eul93, Sch82b, Sch86c]. B^{A} [Ros87]. C [Ric15a, Kre10]. cos′ [Hes95]. cos(ax + b) [Eul97]. cos^{2n} X [Wie00]. Cosx [Sch84c, MS89]. d [Ric15a]. d/dx(x^{r}) = rx^{r−1} [Lin85]. D_{m} [Joh13]. D_{n} [Joh13]. δ [Ang77, Kin83, LeD73]. det(AB) = det(A) det(B) [Nat24]. e [And81, Bro04, Cha99, Chr83, Ćur14, KB99b, Sch72a, Seg82a, Ska97, SY93, Wie85]. e = ∑_{0}^{∞}(1/n!) [Sch94a]. e^{iθ} [Sim95]. e^{π} [Niv72]. e^{π} > π^{e} [Sch85b]. ϵ [Ang77, Kin83, LeD73]. ∃ [Wal04a]. f [Ayo87]. F (1) [Mil71]. F (d) [Mil71]. f(g(x)) = x [Zho12a]. f(t) = c/t [HSZ05]. F (x) = ∫_{x}^{2x} f(t) dt [HSZ05]. f^{−1} [Ayo87]. f^{−1}(x) = 1/f(x) [Mac03]. ∀ [Wal04a]. h [Woo75]. i [Str91]. i^{z} [PP21]. ∞/∞ [Car78a]. ∫_{a}^{b} x^{k} dx [Sch73]. ∫ f(x)e^{ax} dx [Kun70]. j [PU86]. k [Ger75, HS86, MT22, Sch73, WK73]. k + 1 [Ger75]. L [Cos17, Gar09]. L(2k − 1) [Che95]. lim_{d→0^{+}}(−d ln d) = 0 [Mat92]. lim_{n→∞} ^{n} \sqrt{n!}/n = e^{−1} [Sch89d, Wie87]. Ln [MW81]. ln 2 [Bur87a]. M [Ryd96, Daw94, Sch00]. mod(x^{2} + 1) [Sch72b]. n! [And80, Hsu80, Mos84, Sch80b]. N [CMS20, Dra00, Lon84, Put86, See98, Whi71, ASR19, Bai98, Buc83, Bum83, CC09, Coo13, EH19, Fra84, GV23, Han90, Hen94, Hsu82, KES09, Lar85, Mer17, Mey84, Pri99, SB84, Sch23, Smi87, SS21]. N + k [Cha09]. n − 1 [Boo79]. n^{2} [Sch23]. n × 2 [DR20]. ⊕ [BSF99]. p [DE77, Lee24]. P_{n}(D)Y = 0 [Kun73]. ϕ [LPP08]. ϕ(x) = 2(6n + 1) [DD95]. π [AT00, AMR13, Boa80, Col03, Sch86d, Sim09b]. Π(n) [Mar82]. π/4 [Bur87a]. π/8 [Gas23]. π^{e} [Niv72]. π_{1} ≈ Z ⊕ Z [KK09]. r [Dub92, Lin85]. R^{2} [DLM04]. R^{n} [DH03, Xu10]. sin′ [Hes95]. sin(ax + b) [Eul97]. Sin(θ_{1} + θ_{2}) = Sinθ_{1}Cosθ_{2} + Cosθ_{1}Sinθ_{2} [Gan81]. sin 1^{∘} [Ria95]. sin^{2n} X [Wie00]. Sin^{2}x [Sau90]. Sinθ [Sri99]. Sinφ_{1} + Sinφ_{2} = 1 [Mus79]. Sinx [Ros91, Sch84c]. sin x/x [GS90a, Li13a]. \sqrt{2} [Gen91]. \sqrt{n} [McC80a]. ∑_{k=1}^{∞} k^{j}/M^{k} [Gor89]. ∑_{n=1}^{∞} n^{m}x^{n} [Nea94]. ∑ a_{n}^{†} [Kem82]. t [Ekl00, Som99]. Tanx [Sch86c]. tan x/x [Li13b]. W [Moo23, PY04, Wan10]. w^{2} [Sch81a]. x [AB91]. x(a − x) [Lan74]. x^{1/2} [Mat94]. x^{1/x} [Sch84b, Sch89b]. X^{2} [NW98, Sch81a]. x^{n} − y^{n} [BJKL97]. ξ(2k) [Che95]. y = x [Ayo87]. y^{2} [Sch81a]. z^{2} [Sch81a]. | tan(A − B)| [Sch81b].

-Base [Woo75]. -by- [ASR19]. -Card [CC09]. -Children [EH19]. -Cube [Bai98]. -dimensional [Coo13]. -gons [GV23, See98]. -Handed [AKMMM23]. -Probabilities [Ekl00]. -purgated [AB91]. -Simplex [WK73]. -Site [Dra00]. -th [MT22, Pri99, Smi87]. -Times [Bum83]. -Tromino [Gar09]. -trominoes [Cos17]. -vectors [HK19]. -Year [Ano95c].

//sagemath.org [Den13]. /24 [Nel22b]. /4 [Mih02].

1/3 [Due93]. 1/e [KJ11]. 100th [Ano15h, Nac15a, Nac15b, Nac15c, Nac15d]. 101st [Nac16a, Nac16b, Nac16c, Nac16d, Nac16e]. 12th [BC22]. 13th [Baj23]. 14th [BT24]. 15th [BCT25]. 17th [Che18]. 17th-Century [Che18]. 19 [BR24, RS21]. 1980s [Pag20]. 1981 [Ald81]. 1990s [WW20].

2 [Cun08]. 2000s [Bei20]. 2009 [Nun10, RC20]. 2019 [Ric19a]. 28S [SS89]. 2nd [Ols13].

322 [Sim19].

’88 [KGSP88].

978-0-19-872349-3 [Hop18a]. 978-0-19-875523-4 [Hop18a]. 978-0-465-09425-7 [Gut19]. 978-0-486-43278-6 [Hop19]. 978-0-801-88709-3 [Gra18]. 978-0-8218-3777-1 [Ser17]. 978-0-88385-793-9 [Kat18]. 978-1-421-42515-3 [Gra18]. 978-1-93951-213-0 [Kat18].

= [AP92, Koh12, NW98, Sch81a].

Aaron [BBHS02]. ab=c [Pen73]. Abacists [Rey93]. Abbott [Spo16]. Abbott-and-Costello [Spo16]. Abel [Gre11]. Abelian [Koh12, Lea76]. Ability [Joh93b]. Absolute [Lew73, OB77, SP79]. Absolutely [KS12]. Absorbing [CH17]. Abstract [Boy81, EB99, Gre00, Mac96]. Abstraction [Hal82, Smo85]. Academic [Joh72a, Per73]. Accurate [Rob85]. Achievement [Beh73, Dou73]. Acknowledgement [Ano89a]. Acknowledgements [Ano84a, Ano85a, Ano95a]. Acknowledgment [Ano95b]. Acknowledgments [Ano96a, Ano97a]. Action [KM13]. Active [Pen20, WWW83]. Activities [Coh77b]. Activity [Sha18]. Actuarial [Byr11]. Acute [Van14]. ad [Ber80]. Adding [San19]. Addition [cH94]. Additional [Man84]. Additions [Ano09a, Ano10a, Ano11a, Ano12a, Ano13a]. Additive [Cal98]. Additivity [BSF99]. Address [Ten83]. Adjusting [JW15]. Admirable [ST19]. Admissions [Gia73]. Adult [Mac78]. Advanced [CS17, Cog99]. Advances [Mos88]. Advantage [LP16]. Adventure [CCH^{+}21, Vod97]. Aerodynamic [CSTM10]. Affected [Chr96]. Affine [GRSS19, Por93, Won00]. After [Gin04]. Aftermath [O’B82]. Again [CE23, Ric12]. Against [Gar01, Knu18, Hen91]. Age [Phr90, Col03]. Agnesi [Gra18, GM99]. Ago [Dun23]. Ahlfors [Alb98]. Aided [Hal93, Leh90]. Aids [Ess82]. Aimee [Kat18]. Airplanes [NS17]. Airy [HMNW94]. Albert [Mau83a]. Album [Ale83b]. Algebra [AB91, BS02, Bid86, Boy81, BB76, BM93, CJLP92, CJLP93, Car93, Che18, DK01, Din13, Dun91, EB99, Gol16, Gre00, Gre85, Hav78, HP99a, HP99b, Hir98, HM23, Isa95, LaT93, Lax76, Lia70, LM90, Mac92b, Mac96, Mer74, MS23, Mue20, PH77, Per71, Pom77, PL09, Por93, Sha25, Sha80, She78, SHL86, Sta79, Tuc93, Wal96, WC77, YD07, Col13, Sri10]. Algebraic [BB16, DPV21, KK13c, Rou12, Sch83c, Wie86]. Algorists [Rey93]. Algorithm [Bai98, Ben94, CL85, Cus83, EG75, FLP18, Fla87a, Fla87b, Fla87c, Fla87d, FF87, Fla88a, FT88, Fla88b, Gue01, Kir88, Kun74, KC90, Lea76, Mat88, McC23, MW83, Pet70, Prc82, Sza05, Whi85b, Luc98]. Algorithmic [Dou85, Ham85, Kor85, Mau85]. Algorithms [Arg84, AG87, Hil85b, KP85, Pri90, RS96, Rob78]. Alice [Wil02b]. Alive [Alb81a, Ped80b]. All-Stars [Som03]. Allocating [JSW22]. Allowed [Hyd88]. Alluring [Eck83]. Almost [FN04, Hoo00, MS84, Nel16e, TT12, Ano70e, Ano71f]. Alphametics [Bro80]. also [Kos17, Swe22]. Alternate [BQ08, Bom82, Fri88, PS79, Sch72a, Sch84a, Sch86f]. Alternating [AE17, AE19, BQ08, BCEW85, DeM13, FN04, Har88, Kat12, MM21, Mor86, Nel12, Nel14a, Pla18, Wu24]. Alternation [Tam87]. Alternative [Bar81b, Bur84, Che81, Coc76, Gur19, Hil90, Joh89, Kun74, Nym83, Pro76, Sha80, Smi80, SY93, Tei23, Wei76b]. Alternatives [Tal84]. Altitudes [Ber80]. Always [Car93, LS12]. Am [Wal11, Hof16, Mah20, Sch89b]. AM-GM [Hof16, Mah20]. Amalgamation [Men88]. Amateur [AC00]. Amatyc [Bal82a, Bal82b]. Amazing [CD93, McC02, Swa05]. Ambassador [Cun08]. Ambikeshwar [Hop19]. America [Wag99]. American [Ser17, Ald79a, BS14a, Her10, Sch23, Zit05]. Among [PM82a, PM83, SL84, Ste76b]. Amortization [KD99, WD90]. AMS [Fle79]. Analog [Sim02]. Analogies [Hof82]. Analysis [Cas99, EW70, Gao97, Jon04, KD92, MC24, MS19, Mil74, TYM88, Vel23, Wil02a]. Analytic [Leh90, Lew84, VK80, Vie78]. Analyze [Mas91, PL94]. Analyzing [Lau20b, McC01]. Anatomy [Moo78a]. Ancient [CFLR12, Dor05, Rey93]. Andrews [Ano96n, Mat96]. Angas [Ken87]. Angle [Alb17, BC18a, BL22b, Fra78, Fra12b, Gau84, JV15, Kob13a, Lut84, MT23, MR95, Mue07, Nel89a, Pet83, Sub15, Wan21]. Angle-Sum [MT23]. Angles [Dia17, Kos20, Kow08, Nak86, SY12, Wil24a]. Angling [Kal86]. Animal [Lor00]. Animating [MP10]. Anne [Ano11b]. Annealing [Alb07b]. Annihilation [Kal87]. Anniversary [Ano15h, Nac15a, Nac15b, Nac15c, Nac15d, Nac16a, Nac16b, Nac16c, Nac16d, Nac16e]. Announcing [Ano95c]. Annual [BC22, Baj23, BT24, BCT25]. Annuities [Rei87]. Anomalous [Boa72]. Anonymity [Hop14b]. Answer [TT21]. Anthony [Ano84b]. Anticommuting [Zho17a]. Antidifferentiation [Sch78b]. Antisubmarine [WWW83]. Antoine [BM88]. Anxiety [Hil80a, Hil80b]. Any [Blå16b, Hor74, LJ22, Lyc23, Man75, PS12, Sch73]. Anymore [Hat07]. Anyway [Boa81b, BU02, Hue92, SH81]. AP [Ken99b]. Apart [BM88]. Aperiodic [CS95]. APL [LeC86, Nei89]. Apologies [Hau05]. Apostol [Alb97]. Appearance [MS88]. Applause [Alb81b]. Apple [Bai86]. Applicable [Man78]. Applicant [Gla12]. Application [Bia98, BL22a, BR83a, Coo84, For92, Fre95, Gan81, Gao97, Heg96, Her93, IM11, KD99, Kun11, NM96, Pug80, Rog07, Slo76, Vie78, Wan10]. Applications [AS88, BHB04, BHS^{+}21, BM93, Cla73, Din22, Fis75, Fon96, GRSS19, GW18, Hah98, Hel73, Hug00, Liu09, Nel18, Nel23, Sch74b, Sch85c, Sha80, Wag84, Wai78, Wal87]. Applied [Cho82, Lat83, Sch89e, Smi11]. Applies [Gem89, Reb79]. Applying [Hol81a]. Apportionment [McC23, NZ03, Swe22]. Approach [Alb02, Ang77, BD78, Bar10, BQ08, Ben94, BR83b, Cal95, Cha07, Cha92, Che81, CCH76, Daw80, DG80, Dou85, DF96, Dub92, Eck98, FN07, GKR19, Gra95, Gre76, Gür22, Hav78, Hec75, Hoe09, Hol76, Hol81b, Hub81, JB92, Jon20b, Kac19, KL11, Kri25, Kun80a, Kun88, Lew84, Lib81, MdMH19, Moo78b, MR98, Pac78, PS79, Pag87b, Sch76, Sch81a, Sch82a, SS82, Sch86b, Sch72c, Sch83c, SB84, Sem74, SR16, SA79, SK01, Suz14, Swe22, Tou95, Tre94, VK80, Wie86, WW88]. Approaches [Hen94, Sch84a]. Approximate [FN16, Gau84, Kre97, LS01, MP10, Mur80a]. Approximates [Ett74]. Approximating [And80, Col92, Sch85a, Ven76]. Approximation [Bro04, Dia17, LL01, McC80a, Nam24, Pet83, Ros76, Sch80b, Wej83]. Approximations [ANANFT01, AM01, Bou24, Gor93, KB99b, TL00]. Apps [Qui16, Qui17]. Apropos [Fek88]. Aptitude [Beh73]. Arbitrary [FKY13, SS22, Tan15, dA95]. Arc [Bar97, CCHR24, Chi73, Een97, Gil98, NM96, Rid98, Sim11b]. Archaeology [Wil82b]. Archimedean [Ara11, CF17, DeT94, KN15]. Archimedes [Bur87b, GW18, Lan95, McG18, Mer18, Osl06, Sch11b, Sny24, Swa13, VV24]. Archives [Ald79a]. Archytas [Hug89]. Arclength [DS99]. Arcsine [Nam24]. Arctangent [Bra01b, Ins89, Una14b]. Area [Alp87, BZ04, Bel15, BC16b, BC17, Bra86, CW90, Cha13, CDL^{+}01, DH03, Gol75, Gor03, Her04, Kwo16, Lew86, Mey86, Mil78, Nel15e, Par95, Par84, Pla17c, RM71, RS85, San24, Sch95c, SS83a, Str90b, The23, Ton97, WCP88, Zam77]. Areas [AB17, FM00, GV23, MT19, Rei87, Sal85]. Argument [Abb24, Ric93]. Arguments [Nic93]. Arising [RS23a]. Arithmetic [ALS12, Bai85, BS00, Bea06, Bri91, Cas99, Cia17, CD93, Cra12, Dom11, FH00, Her93, Hol81a, Kat03, Kin74, Lon84, Mer17, Nel89b, Nel15a, PM82b, Pla78, Pla14, Pla16, RS23a, Sch78a, Sch82c, Sch88b, Sch70, Som80, Edd85, MP88]. Arm [BM98a, Con17]. Aroma [LN99]. Arrangements [DR06]. Arranging [GG22a]. Arrays [Edg22]. Arrivals [BW19]. Art [Lin19]. Artemas [AC00]. Articulation [GW70]. Arts [Fei76, Gra99a, Gra99b, Len74, SW76]. Arts-II [Gra99b]. Asking [Dav94]. Aspect [Len74]. Assessing [BM95]. Asset [Din13]. Assignment [GS90b]. Assisted [BW91]. Associated [Ano93i, Ger00]. Association [Ald78]. Associativity [Dob72, Mal22b]. Assume [SH81]. Assumed [Alb81a]. Astray [Gut19]. Astroid [Sch95c]. Astronomy [Mot75]. Asylum [Smu82a, Smu82b]. Asymmetric [Gar99, SJC00]. Asymptote [BH93]. Asymptotic [Kal85a]. Atmosphere [Str04]. Attainable [PP84]. Attempt [EB99]. Attic [MP12]. Attitude [Beh73]. Attitudes [Mas84]. Attractive [RS23b]. Attractor [RS23b]. Atypical [Pio21]. Auctions [JF88]. Audio [Mor73, Wil72]. Audio-Tutorial [Mor73, Wil72]. Augustus [Sim11a]. Aureate [Spo23b]. Authority [Alb96]. Authors [Ano79e, Ano84i, Ano85g, Ano86f, Ano87f, Ano89l, Ano97n]. Autobiography [Rei86]. Automatic [Nei89]. Automobile [FN09]. Auxiliary [MM85, Sil89]. Average [CSWW01, Dee99, Dob85, Dun97a, Zac18]. Averages [Bou24, Roz94].

Averaging [Fal85, FG19, HT84, Pfa11]. Averse [MZ24]. Avoid [Ful05]. Avoiding [CL90]. Avoids [Nie94]. Awards [Ano08p, Ano08q, Ano11h, Ano14h, Ano14i, Ano15g, Ano16f, Ano17f, Ano18f, Ano21h, Ano24g, Ano12g, Ano19g, Ano20f, Ano22d]. Awareness [Dav85]. Axes [Sch92b]. Axes-Not [Sch92b]. Axiom [Oma09]. Axis [Hwa12, Knu18, Wan21]. Ayse [Kat18].

B [AR86, Sri10, Luc98, Var76, AP92]. b-algorithm [Luc98]. Babbage [Ait16, FF24]. Babylonian [Bid86]. Bach [KC87]. Back [Bon21, BJ24, DM10]. Backward [Ste81b]. Bag [May24, Mon21]. Balance [And96, Bai86]. Balanced [BB76, LM06, SS17]. Balancing [Edg23, Jon12a]. Ball [Sim06, Tan15]. Balloon [Gru94]. Ballot [KK13a]. Balls [Jep79]. Baltimore [Gra18]. Band [ACJ^{+}06, Kuc16, Sek01, SC03, Swa11]. Bank [Bol83]. Bankoff [Ale92]. Bargaining [Kir96]. Barge [GS03]. Barn [Sch97a]. Barometer [Pag79b]. Barrels [Sin90]. Barycenter [FG19]. Base [Fre86, Pla17c, Woo75, smbV16]. Baseball [AS22, Som03, Tho98a]. Based [PH83, WJ73, KB99b]. Basel [Ben12b, BM13, Mor16]. Bases [CJ23, HB92, Tou96]. Basic [BB76, Bru83, Gut19, Jon71, Kal84a, Led77, Sch74c, Sch76, Sew18, Tan17]. Basics [All03]. Basin [CMS20]. Basis [ALT10b, Hil92a, Smi80]. Basketball [SH12]. Bat [Bai87, Log79]. Batting [Roz94]. Battle [Nas85]. Baye [Fra87, MB82, Ros12, Sha25]. Be [Bon14, BB15, Dor77b, Eck81, Fla11, Gei84, HSL77, Mil85, PS15, Ren81, SS93, Sim02, Wei76a, TYM88]. Bear [CKM86]. Beat [PV16]. Beauties [Ber06]. Beautiful [PP21]. Beauty [Gut19]. Bed [WS05]. Bedroom [Sto07]. Been [BW13]. Beetles [GG22a]. before [Cur05a]. Begin [But81]. Beginning [Boy81, Mer74]. Behavior [HS00]. Behaviour [Hig23]. Behind [HJ17, San06, Sim11a]. Behold [AM87, Ayo87, Buc85b, Edd85, Har86, Har87, MP88, Nak86]. Being [DPV21, LL71, vR01]. Ben [Ric19a, DS12]. Ben-Hur [DS12]. Bernoulli [Mer22b, Coo82, Mor15, Par13, Per94, Rod80, Wie85]. Bernoullis [Dun87]. Bernstein [BMV06]. Bers [AR87]. Bertrand [Lo 97]. Besides [Luc84]. Best [Ash97, BL22a, Mau85, Roe93, Ska96, VK80, WJ73]. Bet [Joh17]. Bets [PV16]. Better [Bet12, Buc85b, Daw97, Dun84, FE01, Gar24, Hal85, Kep85, Mas84, Wat89, You70]. Between [BPR15, CH17, CAS20, Dun97a, FBH83, Mer22a, Nel88, Pet14, Sch88a, Vel23, CW90, CGSO21, CG24, Nel92, OHC07]. Beyond [Alb05a, Hal92, Moo86, Niv79, Wal95a]. Bézier [Pen19]. Bézout [BL22a]. Bhaskara [BS00]. Bi [Fla87a, Fla87b, Fla87c, Fla87d, FF87, Fla88a, FT88, Fla88b, Mat88]. Bi-Month [Fla87a, Fla87b, Fla87c, Fla87d, FF87, Fla88a, FT88, Fla88b, Mat88]. Biangular [NW10]. Bibliography [CS74]. Bicentennial [Zit05]. Bicycle [Fin02, NR16]. Bicylinder [DeT94]. Bifurcation [Per83, RS00]. Bifurcations [MP07, SR10]. Big [Cos17, Eck07, Ric19a, Tru83]. Bigger [Ros01]. Biggest [BW98, Lan93b]. Bike [Sch21a]. Billiard [Jep79]. Binary [Hof78a, Hos18b, Kin74, KKS90b, Mar90b]. Bingo [AS02]. Binomial [Abe15, Fri89, Gui95, Hat07, HP99a, Kat09, LM23, Mal22a, Mat25a, Str77, Tha18]. Binomials [Osl00]. Binominal [Fra87]. Biographer [WK94]. Biology [KS03]. Biotech [BHB04]. Bird [Daw90]. Birkhoff [AW83]. Birth [Wie90]. Birthday [Ale79d, HS86, LM16, Sch16, Tat02, Tre19]. Birthplace [Som03]. Births [FE01]. Bisecting [BC16b, BC17, Coo24, Kos20]. Bisection [CL85]. Bit [Reb89]. Bivariate [AFL^{+}21, Fon09]. Bizarre [Gri17]. Black [Ric19a, Sew18]. Blackjack [Buc83]. Blast [Wil24b]. Block [Dr24a, GL92, Hua25]. Blokus [AC22b]. Board [Ano70c, Ano70d, Ano71d, Ano71e, Ano72c, Ano72d, Ano73d, Ano73e, Ano73f, Ano74c, Ano74d, Ano74e, Ano74f, Ano75a, Ano75b, Ano75c, Ano75d, Ano76a, Ano76b, Ano76c, Ano76d, Ano77a, Ano77b, Ano77c, Ano77d, Ano77e, Ano78a, Ano78b, Ano78c, Ano78d, Ano78e, Ano79a, Ano79b, Ano79c, Ano79d, Ano80d, Ano80e, Ano80f, Ano80g, Ano80h, Ano81e, Ano81f, Ano81g, Ano81h, Ano81i, Ano82a, Ano82b, Ano82c, Ano82d, Ano82e, Ano83a, Ano83b, Ano83c, Ano83d, Ano83e, Ano84c, Ano84d, Ano84e, Ano84f, Ano84g, Ano85b, Ano85c, Ano85d, Ano85e, Ano85f, Ano86a, Ano86b, Ano86c, Ano86d, Ano86e, Ano87a, Ano87b, Ano87c, Ano87d, Ano87e, Ano88b, Ano88c, Ano88d, Ano88e, Ano88f, Ano89b, Ano89c, Ano89d, Ano89e, Ano89f, Ano90a, Ano90b, Ano90c]. Board [Ano90d, Ano90e, Ano91a, Ano91b, Ano91c, Ano91d, Ano91e, Ano92a, Ano92b, Ano92c, Ano92d, Ano92e, Ano93a, Ano93b, Ano93c, Ano93d, Ano93e, Ano94c, Ano94d, Ano94e, Ano94f, Ano94g, Ano95e, Ano95f, Ano95g, Ano95h, Ano95i, Ano96d, Ano96e, Ano96f, Ano96g, Ano96h, Ano97d, Ano97e, Ano97f, Ano97g, Ano97h, Ano98b, Ano98c, Ano98d, Ano98e, Ano98f, Ano99a, Ano99b, Ano99c, Ano99d, Ano99e, Ano00a, Ano00b, Ano00c, Ano00d, Ano00e, Ano01f, Ano01g, Ano01h, Ano01i, Ano01j, Ano02f, Ano02g, Ano02h, Ano02i, Ano02j, Ano03f, Ano03g, Ano03h, Ano03i, Ano03j, Ano04f, Ano04g, Ano04h, Ano04i, Ano04j, Ano05f, Ano05g, Ano05h, Ano05i, Ano05j, Ano06f, Ano06g, Ano06h, Ano06i, Ano07e, Ano07f, Ano07g, Ano07h, Ano07i, Ano08f, Ano08g]. Board [Ano08h, Ano08i, Ano08j, Ano09f, Ano09g, Ano09h, Ano09i, Ano09j, Ano10b, Ano10c, Ano10d, Ano10e, Ano10f, Ano11c, Ano11d, Ano11e, Ano11f, Ano11g, Ano12b, Ano12c, Ano12d, Ano12e, Ano12f, Ano13b, Ano13c, Ano13d, Ano13e, Ano13f, Ano14b, Ano14c, Ano14d, Ano14e, Ano14f, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, AKMMM23]. Boardwalk [AR97]. Boas [Boa83]. Bobo [CK14]. Body [Swa11]. Boggle [Nee13]. Bolzano [EO17]. Bombelli [AB91]. Bond [Heg96]. Bonus [RS16]. Book [AB83, AH82, Ale83a, Ale07, Ano15a, Bac09, Bar11, Bla01, Boa19, Bou81, Bra13, Bur16, Cam82, Car17, Che15, CB09, Col10, Col13, Cox12, Cra11a, Dev05, Dic17, Ewi99, Fus13, Gar07, Gar15, Gol17, Gol10, Gol11b, Gol13, Gra03, Gra18, Gre16, Gut19, Hen14, Her09, Hop14a, Hop16, Hop18a, Hop19, Hug06, Kap17, Kar13, Kat18, KB99a, Ken01, Kho14, Kin10, Kly15, Kra01, Lan83, Lin00a, LMK79, LF79, LTM79, LRM79, LM79, LG80, LB80, LC80, LH80, LGA81, Mal81, McD00, Mor11, Mur10, Nun10, O’R15, Ols13, Ric19a, Ros17, Ros00, Ros02, Ros06, Ros16, Rou09, RC20, Sau99, Sch11a, Ser17, Sol99, Sri10, Swe02, Swe03, Tro82, Wag11, Wim02]. Books [Edg18, Gut19]. Boolean [She78]. Bordering [CM98a]. Borrower [Mil85]. Boston [Ric19a]. Both [Fra75, Fre08, Mil85]. Bottles [Boa80]. Bound [LM10]. Boundaries [Bai91, CMS20]. Boundary [All16, The23, Whi11]. Bounding [HM97, Sha15]. Bounds [PM82b, Pfi88, SP79, Und84]. Bowl [Som00]. Box [McG18, Pin74a, Chu09, CDH^{+}14, Daw80, Dun84, HGGB08, JSSW06, KK13a]. Boxes [Nel16c, smbV16]. Brachistochrone [Joh04, Zen96]. Bracing [Fre12]. Brahmagupta [BS98]. Brainteasers [Ano09w]. Brams [Jon10]. Break [Hau08]. Breakage [GS11]. Breaking [AOW05, Ano79f, Ano79g, Ano79h, Ano80i, BBW79]. Breakthroughs [Rob84a]. Bressoud [Ric19a]. Brewery [Col94]. Bridge [DM03, Gro10, HM05, JF88, Zee13]. Brief [Eve02, Kle89, Pie77]. Bringing [Cat19]. British [Joh76]. Broken [Gro05a, Kri25]. Brown [Hen22]. Bryn [Dun23]. Bubbles [AP74]. Buckled [Man98]. Buffon [Arn94, Kac19]. Build [Dun84]. Building [Ald78]. Bulgarian [Hop12]. Bunch [Mor04]. Bureau [Ano71c]. Burns [Ano11b]. Bus [Wil99b]. Business [Man75, Yin87]. Buy [Gar24].

C [Nac15b, Nac16b]. Caesar [LP87]. Cake [Fle99, TT12]. Cakes [HM10]. CAL [Whi88]. Calculate [Sul09]. Calculations [Bar22, Kal84a, LLH21, NRT20]. Calculator [MW83, Seg82a, Wai78]. Calculators [Ano80a, Ano80b, Ano80c, Ano81c, Ano81d, BB82, Bro82, Cha83, CM83, Cus83, Gil82, Gor82, KP85, Kir88, Lev81, Lic81, LN80, McC80b, MDC80, Rey93, Roq79, SW79, Sto83, Sul09]. Calculus [Aar00, Bai89, Bel84, Bel15, BM98a, Bit70, Boa71b, Bod74, BI10, BW13, BE95, Bre02, Bri92, Bro99a, But81, BW91, CS17, CdSS17, Cat19, Cha07, Cib93, Cla73, Cog99, Col94, Col03, CM14b, CS93, Daw19, Dic06, DM03, DM10, FG92, Far02, Fra86b, Gen73, Gor92, Gri74, Hah98, Ham84, Heg87, Heg96, Hof16, HJ17, Kle84, KD99, KK13b, Kuc16, LJ73, Lar75, Lax84, LeD73, Lea16b, Lea18, Lei07, LS22d, LRSZ14, Lin00b, LB97, Mac84, Mac95, McG18, MS78, Nie94, Niv79, Nun10, Pal98, Pec70, Pen03, Per83, Qui22, Ral84, Ric19a, Ric18, Rid95, Roj16, San77, SE96, Sch78a, Sch21b, Sil90, Sta79, Sta71, Thu00, Ton15, Tre92, Var76, Var77, Vel23, Wag84, Wey23, Xu14, YD07, Zia91, Zob13, Boy70, Ric19a]. Calendar [Ano73a, Ano73b, Ano73c, Ano74a, Ano74b, Ano81a, Ano81b]. Calendars [CAS20]. Camel [Cla21]. Can [Bon08, Bon14, CS86a, Dor77b, Ewi95, FPS13, Fal85, Fin02, Guy84, HSL77, Ken91, Kou95, Kre17, Lan93b, Lyc23, PV16, PS12, Roe93, Ruc87, Thu00, THMM24, Wei76a, ZV07]. Canadian [Wil02a]. Cancellation [Boa72]. Candies [Adn98]. Candy [Row15]. Canine [Mal16]. Cannonball [DS99]. Canonical [Cos94, SH18]. Can’t [Har96a, Sch97b, Sim02]. Cantor [Iga08, KR98, Kno85, MR96, MDL11, Mih02, Mih04, San19, Smu85]. Capacity [Nie94]. Capped [SH14]. Caps [ZV07]. Capsule [BW91, KKS90a, KKS91, Nie91, PJS81, Bai98]. Capsules [Ano01a, Ano01b, Ano01c, Ano01d, Ano01e, Ano02a, Ano02b, Ano02c, Ano02d, Ano02e, Ano03a, Ano03b, Ano03c, Ano03d, Ano03e, Ano04a, Ano04b, Ano04c, Ano04d, Ano04e, Ano05a, Ano05b, Ano05c, Ano05d, Ano05e, Ano06a, Ano06b, Ano06c, Ano06d, Ano06e, Ano07a, Ano07b, Ano07c, Ano07d, Ano08a, Ano08b, Ano08c, Ano08d, Ano08e, Ano09b, Ano09c, Ano09d, Ano09e, Dan92, Gor92, HB92, Hil92a, KKS^{+}89, KKS^{+}90c, KKS90b, KKS^{+}90d, Kra91, Pag79a, PZA^{+}79, PKL^{+}79, YMH90]. CAPTCHA [AOW05]. Captured [Ste73]. Capturing [HS96]. Car [WS96]. Carbon [Zob13]. Carcassonne [CK17]. Card [Bos00, CC09, CD93, CC95, Ehr21, Gol16, MS20, Nag21, PDR21, San06, Swa05]. Card-Collecting [Bos00]. Cardinal [SR12b]. Cardinality [BW15, Fra86a, LM21]. Cards [Gar00a, Sim02]. Carefully [HM10]. Cares [NW98]. Carlo [Cea84, Nea93]. Carolina [Joh72a]. Carpenter [Win91]. Carroll [Ait16, McC02, Rat10, RT07]. Carryless [ALS12]. Cars [BJ24, Kir96]. Cartwright [TM01]. CAS [Hen91]. Case [And96, Hen91, Kai01, KM15, Liu01, Sch74c, Tan86b, Van14, Vel23, WD90, CG19]. Cases [Biv97, Ser15]. Casino [Con99]. Catalan [IMW12, Kos09, KG12]. Catan [AMM15, AKMMM23]. Catenaries [Cox25]. Catenary [Cel99, SR12a, SR12b]. CATIA [BM77]. Cauchy [De 10, KKS90a, MM85, Par16, Pow80, SS79, Ste75, Zho17b]. Cauchy-Euler [MM85]. Caught [AKMMM23]. Causes [Hil80a, Hil80b]. Caustic [LB97]. Cautionary [Kru00]. Cavalieri [Eve91]. Celebration [Hop15]. Centaurs [DI08]. Center [BC17, GW18, Hu21, SJ11, Sim00, The23]. Centered [Cag20a, Nel15b]. Centering [ST02]. Centers [Mar08, Ric05]. Central [Bia98, PM82a, PM83]. Centroid [BC17, Lo 91, Mer18, Sch77]. Century [Che18, Dun23, Lea18, AC00, Joh76]. Certain [Bur84, CL90, For17, Lin70, SS22]. Cevians [Moo83]. ch [Joh17]. Chain [BR24, Buk08, Jon04, Ken99c, Loe21, Thu82]. Chains [CH17, LM92, Rid90]. Chair [Sch95c]. Challenge [CS17, Hol79, Ped80a, Rob73]. Challenges [BS15, Fra87]. Challenging [BB14, Fra86a, Woo70]. Chance [Hor74, Jon20a, Mer74, Sim02, WJ73]. Change [Fre86, Gra11b, Hop14c, IM11, Kha13, Lu22, Ric19a]. Changing [Joh89, Ric12]. Channel [Ros04]. Chaos [LLH21, Mar24, Par91, Nat91]. Chaotic [MS91, Mer97, PP21, Str91]. Characterization [BB16, Mar98a, OS16, Xu10]. Characterizing [RR98]. Charity [EG75]. Charles [Gol13]. Charlotte [Ken87]. Chart [Kem82, Shi78]. Charts [FM77]. Chasing [BBHS02]. Chebysheff [Sch89c]. Chebyshev [BMV06, Cas03, Tou95]. Check [Moo80]. Checkerboards [DL73]. Chen [RC20, Gor24]. Chern [CL83]. Chess [MST21, Wag14]. Chessboard [ASR19, BLW89, DT13, Pau19]. Chessboards [Gar09]. Chestnuts [Cib93]. Chicago [Fei76]. Chicken [CGSO21]. Child [JW15, KO00]. Children [EH19]. China [Bra90, Che18]. Chinese [How02, Per17, Sch87b]. Chiò [FL75]. Choice [Eis98b, EW70]. Choking [MS19]. Chomp [MM13]. Choose [CCH^{+}21]. Choosing [Led16]. Chord [Mas11]. Chordal [Een97].

Chords [FM00, Hol82]. Christiaan [Buk08]. Christmas [KB21]. Chu [Spi16]. Chung [Cal95]. Chutes [CHJ11, CG14, FGL^{+}18]. Cida [Cun08]. Cida-2 [Cun08]. Ciphers [LP87]. Circle [Bel15, BB15, CW09, FT18, Hoe87, LJ22, Mar08, Spo23b, Fei76]. Circle-Squaring [CW09]. Circles [Ber14a, Cha80, DK08, Due96, Fri24, Jur94, Kar14, Lon83, WP90, Oku22]. Circular [Gor08, Hal12, Par83, Ric15a, Ric93, SM13]. Circularity [BG04]. Circumference [LJ22]. Circumscribe [JW22]. Circus [Far15]. City [Shu02]. Clamped [Uet88]. Clapping [Haa91]. Clarify [Ker83]. Clarifying [NF03]. Class [Fer98, Gro96, Kha03, Lei07, LW10, Pie71, Sch86b, Zah72, Ziz13]. Classes [FM77, Heg87, Sch72b, SB84, Shu02]. Classic [Nas23, Roj16, Tea06]. Classical [Hug00, ST20]. Classics [Ale79a, Ale79b, Ale79c]. Classification [Fra86a, GG22b]. Classifying [Loe14, Ped80a, VB86, smbH02]. Classroom [AD78, Ano01a, Ano01b, Ano01c, Ano01d, Ano01e, Ano02a, Ano02b, Ano02c, Ano02d, Ano02e, Ano03a, Ano03b, Ano03c, Ano03d, Ano03e, Ano04a, Ano04b, Ano04c, Ano04d, Ano04e, Ano05a, Ano05b, Ano05c, Ano05d, Ano05e, Ano06a, Ano06b, Ano06c, Ano06d, Ano06e, Ano07a, Ano07b, Ano07c, Ano07d, Ano08a, Ano08b, Ano08c, Ano08d, Ano08e, Ano09b, Ano09c, Ano09d, Ano09e, AMM15, Bai98, Bru83, BW91, CK17, Cla21, Cog99, Dan92, Fag20, FG92, GK77, Gor92, Gru94, HB92, Hil92a, KK13b, KS88, KKS^{+}89, KKS^{+}90c, KKS90b, KKS90a, KKS^{+}90d, KKS91, Kra91, Lom94, MdMH19, MC78, Nat91, Nie91, Pag79a, PZA^{+}79, PKL^{+}79, Rob73, Sch72a, Sch74a, Sch81a, Sch82a, Sni77, Wai78, WW88, YMH90]. Clay [AL00]. Climate [Kha13, KK13b, WM13]. Climbing [Cur05b]. Climbs [DS12]. Clipped [Ren14a]. Clique [Hol23]. Clock [JK98, Per17]. Clocks [Ste81a]. Close [Mur80a]. Closed [Seg83, Tan15]. Closed-Form [Seg83]. Closer [CK14, Egg20]. Closing [Ben76a, BL75, Zir76]. Clothing [Sau90]. Club [AOW05, RS96]. CMJ [Ano15h, Hop20, Kly20, Nac15a, Nac15b, Nac15c, Nac15d, Nac16a, Nac16b, Nac16c, Nac16d, Nac16e, Ano95c]. Cobb [Goe11]. Cobwebs [NF03]. Coconuts [Kun98, Mor04, SB97]. Code [Kin74, Liu09]. Codes [Liu01, Nag21, Sal88]. Coefficient [Jon74, Mat25a, CG19]. Coefficients [Abe15, AFL^{+}21, Bra19, De 10, Gol94, Gri71, Gui95, HP99a, Kat09, Lau20b, Mel22, Mun04]. Coffee [Ban06, LB97]. Cognitive [BE03]. Coherence [Hil84b]. Coin [Ber02, Blo98, Cha79, Eid93, Hof78b, LP16, Lip03, MP03, Mor04, Spe86, MP03]. Coin-Tossing [Cha79, MP03]. Coincidences [Hat07]. Coincident [BL22a]. Coins [MG97]. Colin [Hed00]. Collaborative [GECW13]. Collapsed [Hoo00]. Collatz [Sol23]. Collecting [Bos00]. Collector [DLPS16]. College [Ano70l, Bra90, Col71, Coo74, Hav78, Joh72a, Spa71, Ano76g, Ano77k, Ano77l, Ano77m, Ano77n, Ano77o, Ano78k, Ano78l, Ano78m, Ano78n, BM95, Bot78, BEN78, DN70, DP71, Dou73, Dun23, Fle79, Fle80, Goo70, Has71b, HA70, Hoo97, Led77, Lei83, Lin75, Lin76a, Lin76b, Pag20, Per73, Pri73b, RM71, Sch70, Sem74, Tox70, Tre92, Bei20, WW20]. Colleges [CS74, Fis75, Jon71, Lai70, LL71, Man70, Man78, McK79, Ste76b]. Collegiate [SS93]. Collide [Ful00]. Collisions [Hil89]. Colonial [San02]. Color [Mit81]. Colorado [DN70]. Colored [DeT98]. Coloring [ASR19, BLW89, Jep86, Reb79]. Colors [GV24]. Columbia [Spa71]. Column [Bou23]. Columns [MM93]. Combinations [Hyd88]. Combinatorial [Ada80, Alb02, BO95, DR06, Haa91, KM13, Mat25a, Mic14]. Combinatorics [AC22b, BM93, MR03, Pug80, Row15, Spe86, WW95, Hop18a]. Combining [CMP15]. Come [Atk02, Gla93, Ler21]. Comets [BBB^{+}13, MS20]. Coming [Ano70a, Ano70b, Ano71a, Ano71b, Ano72a, Ano72b, Wag99]. Commensurable [Par07]. Commentary [Ano83j, Ano83k, Dav82a, Dav82b, Dav82c, Fus13, Kin79a, Kin79b, Kin79c, Kin79d, Kin79e, Kin80a, Kin80d, Kin80e, Kin80b, Kin80c, Kin81a, Kin81b, Kin81c]. Comments [Che96, LM23, Sch74c]. Committed [Kip70]. Committing [Law71]. Common [Bes14, Ćur14, Jor77, Lei83, PH83]. Communication [GS96]. Community [CS74, Coo74, HA70, Joh72a, Led77, Tox70]. Commutativity [AB92]. Commute [Lyc23, PS12]. Commuting [Sou14]. Compactly [Ash20]. Compactness [Tan15]. Comparative [Mil74]. Comparing [Not71, Ros87, Var76]. Comparison [Mor86]. Compass [Hec03]. Compensating [Egg20]. Competition [Bra80, Pag86, Wie90]. Compleat [CS88]. Complementary [Ban06, Tal84]. Complete [Nel77, RR89]. Completing [LC23, MM14, Moo78b]. Complex [Bra85, CMS20, Due96, Ham07, Hol81a, KK93a, KM89, LH89, McA83, Moo86, Nac17e, Nat91, NL84, San21, Sch72b, Spo17, Ves85, Wei76a]. Complexes [BMO^{+}00]. Complexity [Hay92]. Complicated [Mas91]. Componendo [Kob14]. Composite [GV24, Mur80c, Sch90c]. Composition [MDC80]. Compositions [Biv97, NF03]. Compound [Rob78]. Compounding [Cho81]. Computation [Bla91, Bro74, DW88, Ekl01, Gan84, McK76a, Rob85, TW98, TB20]. Computational [BR83b, Cla21, ES85, Smi80]. Compute [Lam24]. Computed [CM90, Kob17]. Computer [AA78, All77, Ano94a, Ano94b, Ano95d, Ano96b, Ano96c, Ano97b, Ano97c, Ano98a, AG87, Aus76, Bai98, Ban90, Bit70, Ble75, BE95, Bri92, Bro74, Bro80, BW91, Col75, CCH76, CS93, Dan92, Den77, Dor77a, Gaz78, Gem89, GR74, Gil82, GG84, Gor92, Gre76, Hal93, Hic70, HB92, Hil92a, JJ78, Ker83, KS88, KKS^{+}89, KKS^{+}90c, KKS90b, KKS90a, KKS^{+}90d, KKS91, Kra91, Lan77, Lar75, Leh90, LM90, Luc84, Mas91, Nie91, Pen86, Pet70, Reb89, Rid95, San77, SP76, SHL86, SK77, Sta71, Tro78, Tym81, YMH90, You74]. Computer-Aided [Hal93, Leh90]. Computer-Graphical [Gre76]. Computer-Oriented [Pet70, BE95]. Computerized [BM77]. Computers [Ano80a, Ano80b, Ano80c, Ano81c, Ano81d, BB82, BC94, Bro82, Cha83, Cof09, Cog99, CM83, Cus83, Gor82, Har96a, Lev81, Lic81, LL71, LN80, McC80b, Rid90, Roq79, SW79, Sto83]. Computing [Cos94, DS87, Har10, Hir24, IB22, Kie85, LPT11, Mac96, MMOM14, Ste83, Tal84, Tee17].

Computo [Sup84]. Concave [Buc85b]. Concavity [You93]. Concentric [DK08]. Concept [Cur05b, Kle89, MdMH19]. Conceptions [CDL^{+}01]. Concepts [GP76]. Concerning [Abb93, Mar82, NRT20, Ven76]. Conchospirals [Boy99]. Concise [GI18]. Concluding [Ano88g]. Conclusion [ASW94]. Concurrency [Hoe97]. Condensation [LLH21, RT07]. Condition [Dob72, GEJ95, PS15]. Conditional [Huf89]. Conditionally [Smo24]. Conditioning [Roz94, ES85]. Conditions [DF96, Mat14, Sch99a, Whi11, You93]. Cone [Cul97]. Conference [Mit72]. Confidence [AD78, SD93]. Conformal [MT23]. Conformality [Fee01]. Confuse [Cof09]. Confusion [Ros70]. Congruence [Eve91]. Congruences [Lef83]. Congruent [ANU14, GV23, Wet10]. Congruent-like [GV23]. Conic [Ayo03, Ber14a, Cul90, Ser17, Ven80, Wil03]. Conical [Kwo16, SH18]. Conics [Cam19, Gas94, Sch92b]. Conjecture [BEN78, KS06, Ser15, Sol23, Zac18, SS81a]. Conjectures [Alb81b, LD15, SY87, SS81a]. Connect [HR84]. Connect-it [HR84]. Connected [DR20, Wag20]. Connecting [Hil90]. Connection [CG24, EG75, Kin83, Lea76, Mey07, PW21]. Connections [Mer22a, OHC07]. Connector [Sel70]. Conquer [MG97]. Consecutive [ANU14, Ber94, HMF20, LM20, Nel14a, Nel22a, Pon07, Pri18]. Consequences [Bur86b, Nel24d, Pen71, PV05, Sch77]. Consider [Hal12]. Consistency [Sch99a]. Constance [Ale80]. Constancy [Hop14c]. Constant [AMR13, Bur85, ES85, HR24, HSZ05, LJ22, Pin01, Ric15a, Ric19a, Sha15]. Constituents [Luc84]. Constrained [Bai91, Ziz98]. Constructed [Ano11b, Smi11]. Constructing [AKMMM23, Boy81, Bra80, Hec03, Kar22, Moo85, Pul88]. Construction [Arm74, BR99, BM77, EB99, Fra78, Fra14, Ger00, Grü84a, Jor77, MM73, Moo83, Seg00]. Constructions [LR24, Woo00]. Consultants [Ano71c]. Consumer [Bai85]. Contain [Hah88, Han82, Mar08]. contained [Lin85]. Content [Lei83, Tho84]. Contest [Rid98]. Contests [PM74]. Continued [Par05, Den12]. Continuing [smb12]. Continuity [BK85, Cro95, Hat11, McM03, MdMH19, MS17b]. Continuous [BR24, Bri78, Cal00, Gra94, Het97, Kar22, Kre17, Lar74, SS99, Tal84]. Continuous-Time [BR24]. Continuously [Nan14]. Contour [Ska91]. Contract [Mil74]. Contractions [AL15]. Contrapositive [Hel73]. Contrary [Fal05]. Contribution [AC00, GM99, Lea16b]. Control [LR20]. Controversy [Rey93]. Contumacious [GY00]. Conundrum [Mal16]. Converge [Ćur14, KS12]. Convergence [AMP87, Bro04, KG12, LM92]. Convergent [Ash97, CL85, Smo24]. Conversation [AS82a, AS82b, AT84, AA90, AA91, AA93, AW83, Ale89, Ale92, Bar79]. Converting [CAS20]. Convex [ACJ^{+}06, BR83a, BR87, BR89, Gre74a, PS82a, Sek01, SC03, Sun21, Swa11]. Convexity [Bel84, Bra01c, HT84, HMP10, You93]. Convolving [CMP15]. Conway [Ben17, Guy82]. Cookie [BK14, KX15]. Cooperative [DN70]. Coordinate [Sch81b]. Coordinates [BR83a, BR87, BR89, Gor93, Hil92a, Lew86, McM03, Mur95, NW10]. COOTIE} [Hir00, BF00, DW98]. Cops [STW17]. Copyright [Gem89]. CORDIC [Ekl01, Sul09]. Core [Rou13]. Corgi [BPP15]. Corner [AA78, All77, Ano70g, Ano94a, Ano94b, Ano95d, Ano96b, Ano96c, Ano97b, Ano97c, Ano98a, Aus76, Ble75, Col75, Den77, Dor77a, Gaz78, GR74, JJ78, Lan77, Mor02, SP76, SK77, Tro78, You74]. Corners [ASW94]. Correct [MS84, Sal88]. Correcting [Liu01, Var11]. Correction [Ano21f, Ano21g, Ano22c, Ano23f]. Corrections [Ano09a, Ano10a, Ano11a, Ano12a, Ano13a, Kir97, PJS81]. Correlation [AFL^{+}21, Fon09, Hub81, Huf89, Kun80a, She22]. Correspondence [Pen71, Pet14]. Corresponding [Nun10, Rou12]. Corset [AB17]. Coset [Kle91]. Cosine [Cip87, Kin13, Nic93, PK83, Pet14, Una10, Wan14a]. Cosines [LS22d]. Costello [Spo16]. Cotangent [Sub15]. Couch [Mor02]. Couldn’t [RB21]. Countability [GP75, Tou96]. Countable [CS03]. Counterexample [PW21]. Counterexamples [Mor86, Pom77, Str97b]. Counterfeit [MG97]. Counting [Bro23, DeM12, DeT10, DL73, Grü84b, Hof78a, JB92, KB21, Lar89, LR13, Nel14d, Pet11, Sch91a, Wag20]. Counts [Sta98]. Course [All03, BE95, BEN78, Bus71, But81, CJLP93, EB99, GS90b, Hah88, Led77, Len74, Mar91, Pec70, Sac22, Ste77, Wil72]. Courses [Car70, Hoo97, Luc84]. Cousin [Nel24a]. Covariances [CH17]. Cover [Ano18b, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, Ano22a, Ano22b, Ano23e, Ano23a, Ano23b, Ano23c, Ano23d, Ano24e, Ano24a, Ano24b, Ano24c, Ano24d, Ano25, Tho98a, Ano18a, Ano18c, Ano18d, Ano18e, Ano19a, Ano19b, Ano19c, Ano19d, Ano19e, Ano20a, Ano20b, Ano20c, Ano20d, Ano20e]. Covering [Cag23]. COVID [BR24, RS21]. COVID-19 [BR24, RS21]. Cow [Hal12]. Coxeter [Log80, Ped80b]. Cozenage [Whi85a]. Crackers [Som99]. Cramer [Bal78, Bom09, Bur83a, Fri88, Hua25, Kal87, Kun88, Var11]. Create [Fin02]. Created [Bon14]. Creative [Bra13, DP80]. Credit [WN94]. CRICISAM [Sta71]. Crime [BH22]. Criteria [Per73]. Criterion [CM14a, Tei23]. Critical [DRW96, GS10, Nor88]. Cross [FS15, HK19, Knu18]. Cross-Sections [Knu18]. Crossing [LPT11]. Crossover [RM71]. Crossword [Ano14a, Ano24f, KK09, Sul24]. Crotchets [Mey91]. Crowd [BNUW23]. Crush [Row15]. Cryptography [Fag20]. Cryptology [LP87]. Cryptosystems [LP87]. Cube [Bai98, Ben12a, GV19, Hug89, Dav82d, Rok14]. Cubes [ANU14, DL73, GRV21, Nel24b, Par11, Tur81]. Cubic [BT07, Che22, Fak99, GS10, Hil90, KW98, KR96, Lar75, Lut80, Nel16c, Pen19, San21, Wal95b]. Cubics [SS22, Sue94]. Cubing [Ben12a]. Cubits [GP02]. Cuboctahedron [HP99a, HP99b, Wal00b]. Cultural [Man74]. Cup [LB97, Sul16]. CUPM [Ano71c, Goo70, Jon71, Per73]. Cups [Ban06, Ste12]. Cures [Hil80a, Hil80b]. Curiosa [Bri75a]. Curious [Due93, Lut17, PV05, RT07, Sch89e, TR23]. Curl [TPB99]. Curling [LR20]. Currency [PS82b]. Curricular [Law71]. Curriculum [CJLP93, Gol83a, Len76, Rob84b, Sch70]. Curse [Sup84]. Curtains [Ric18]. Curvature [GSS15, Hoe09, Smi00]. Curve [Bra01a, BG04, Cav93, DC95, Ken20, SS00, Sch94b, She77, Ska96, Ste80]. Curves [BL22a, Boy10, Boy15, Bro00, Bro01, Cam19, Che96, Gol75, Kal85a, Kro82, Mat17, MHS92, Nan14, ST19, Sim07, SS94, Ste80, Swa13, Wil98b]. Customer [Yin87]. Cuts [BB11]. Cutter [Har95]. Cutting [ASW94, BB11, Cul97, HM10, Knu18, You24]. Cyanobacteria [GECW13]. Cycles [Dal19, LS12]. Cyclic [Bea06, BPS12, Eck83, OS16, PS22, Pet11, Sar23, Wal86]. Cyclically [Kin74]. Cycling [Kal96a]. Cycloid [DeT88]. Cyclotomic [Lef83]. Cylinder [Cul97, GE09, Siu84, Zam77]. Cylindrical [LS01, McM03, Mur95]. Czédli [HMT11]. Czédli-type [HMT11].

D [Nat91]. D.B.E [TM01]. D.C [Coo70]. Dame [TM01]. Dan [Bro99b]. Dance [Job11, SH10]. Dangers [Ste74]. Dantzig [AR86]. Darboux [OP76]. Data [Cav93, Fon09, GK88, Hei78, Led16]. Database [CS86b, CS87, CS88]. Date [Guy84]. Dates [CAS20]. David [Ano88a, Mey06]. Day [AW90, Sch04]. Days [Mue20]. DC [Kat18]. Dead [OS81]. Deal [Dr24b]. Decade [Bei20, Hop20, Kly20]. Decagon [Nel16g]. Decimal [Lon81]. Decimals [AT08, CS06, Gan84, Lea84, McK76a, Pla78]. Decision [Mar98b, Zee13]. Deck [Sim02]. Decks [Buc83]. Decomposition [Coa11, Kal96b, MR98, Sob97]. Decompositions [Edg20]. defective [CG19]. Defended [Kle84]. Defending [MR79]. Defense [FH94, WK94]. Define [Car72, CEO88]. Defined [Bus71]. Defining [Lew86]. Definite [Bra16, DD10, Gür22, Har10, Sin12, Und84, Win95]. Definition [CM14b, Har10, Sni77, SY93]. Definitions [Pet14, Ric22]. Deflection [Kal06]. Deformation [Lar74]. Degenerate [Ven80]. Degree [BD96a, BD96b, Boy10, Coo74, Fei76, Kow16, Lai70, Ros76, SS22]. Degrees [Jav17]. Delay [Hop15]. Delta [Ett74, Hun98, Ven76, Yan80]. Deltas [Ben76b, Lin74b]. Delusions [Hal93]. Demise [BR17]. Demonstrating [Pal98]. Demonstration [AD78]. Demonstrations [Ban90, FG92]. Density [Che05, KC84, Stu95]. Dependence [FBH83]. Dependent [ESS08]. Depletion [Gro08]. Depreciation [Heg87]. Deranged [Spi10]. Derby [Dru92]. Derivation [And80, Grz24, Hes95, Lin85, OAHS21, Qui22, Sch80a, Sew18]. Derivative [Cat19, Cla11, DH03, Eis98a, Kou95, Mur80c, Sch82b, Sch13, Sri99, Str97b, Tha18]. Derivatives [AM02, BF77, Biv97, Bri78, CL15, Cip87, Fla20, Jos16, KO00, Lar74, Mau02, McM03, Mor91, PS79, PE82, Pin72, Pro76, Sch84c, Sch86c, Sed80, Smi87, Spi16, Str90a]. Derive [HM95]. Deriving [HW84, MM85]. Descarte [Fre95]. Descartes [BW13]. Descent [Dru95]. Description [Som80]. Desert [Bai09]. Design [Zah72]. Designer [CS06]. Designing [Far02, Fre08, Har95, Mat17, Ros12, Tho98a]. Designs [Fer98, Ger00]. Detect [Kre17, MG97, Nan14, Sal88]. Detecting [Moo80]. Determinant [CM98a, LLH21, Spi06]. Determinantal [Mar92a]. Determinants [Buc85a, CDNN02, DS87, Fek88, FL75, Gob88, Hua25, LPT11, LW18, Mar90a, RT07, Sue94, TW90]. Determination [Ria95].

Determine [Kou95]. Determined [Bra19]. Determining [BG04, Nea93]. Deuce [KRSW18]. Developing [CC19, Hav78]. Development [Joh72b, WM76]. Developmental [BS72, Mil74, Mit72, Ros70, Tox70]. Deviation [Boo79, Fal81, Mur84]. Device [Ros77]. Devours [BK14]. Diacu [Ols13]. Diagnostics [Hoa88]. Diagonal [BM98b]. Diagram [Dev91a, KH84, RS00]. Diagrams [Bri78, Daw80, Grü84a, Grü84b, KR98, Loe14, McC23, NO92]. Dialectic [Kor85]. Dialogue [KS03]. Diameter [LJ22]. Diameters [Sch85d]. Diametric [Bea09]. Dianne [O’H85]. Dice [BMRS13, CC19, FS99, FGL^{+}18, Gri17, MR03, MS12, Mur81, Nef82, Pom84, Pug80, SS17]. Dice-Sum [Mur81, Pug80]. Did [CdSS17, Lan95, Par13, Sto04, Zho17b]. Diet [Seg87]. Difference [ANU14, BY97, Jon20b, Lon84, MM14, Shi10, Wan14a, Zia91, Yiu12]. Differences [All16, Jon20b, Str90a]. Different [Boy81, Cha07, Fra12b, GV23, Jon24, MHS92, Nic93]. Differentiability [AL05, BK85, McM03, Ros91]. Differentiable [Bum83, Cal00, FM94, Nan14]. Differential [Ban90, Bla94, Boy94, BJ24, Bus71, Cho82, Dav94, Ehr94, Gug92, HM95, Hub94, HS00, KK13b, Lau20b, LS22d, Lom94, MF94, MT97, Mun04, Nas85, Oli73, Par13, Pet14, Sac22, SS77, Str04, UK96, Wol23, Ziz98, CG19]. Differentials [Bai89, DM03, DM10, Dra13, FK94]. Differentiate [Dra13, KES09, THMM24]. Differentiating [Mil78]. Differentiation [Eul86, Gar93, Hen91, MP14, Nei89, Sub04, Tro04, Wie83, Wie01]. Difficult [Pen20]. Digit [Kal96a, Smi89]. Digital [Asa84, KD92]. Digits [Fri84, PU86, Tre94]. Digression [Man81]. Dihedoku [Nac17d, Nac17a, Nac17b, Nac17c]. Dihedral [JB92, LW18, Som80]. Dilemma [Bar09]. Dimension [Han90, Tan15, Leh22]. Dimensional [BZ04, JW19, Kat03, Coo13, HMT11]. Dimensions [Fra84, Hal13, HP99b, Mey84, Nog17, Put86, San21, Sch13]. Dimple [GG82]. Dinner [DK08]. Diophantine [AA04, Joh22, SB97]. Dipsticks [LS01]. Dirac [Bar81b]. Direct [EH16, Ins89, Pet11]. Directions [Boy94]. Directrices [Moo85]. Directrix [BR99]. Dirichlet [BR89, Pin74a, SS99]. Dirk [Tat02]. Disappearing [Kno85]. Discipline [Knu91, Mac84]. Discontinuity [Van14]. Discontinuous [Ash20, Har96b]. Discover [Ald79b, Ska97]. Discover- [Ska97]. Discovered [Rou13]. Discovering [Dor05, MT97, NW10, Var77, Kat18]. Discovery [BEN78, Cho80a, Daw80, GG84, Gre00, OB79, PH77, SA79]. Discrete [BR89, Cat19, CC95, FG19, Ham84, Han20, Hua88, JD98, Kat18, Lar85, Lop93, Luc84, NR16, Ral84, San91, Tal84, Tha18, Wag84, Woo84]. Discretization [Bor11]. Discussion [Mor73, Sni77]. Disease [GO12, Mic12]. Disguise [MM13, Pri99]. Disjoint [ES23b, EH24]. Disk [Jur94]. Disks [Key94]. Dispersion [PM82a, PM83]. Dissection [Dr24a, Wu23a]. Dissections [HL12]. Distance [Dee99, Dun97a, Gaz84, Gaz92, GI18, PG79, Sas81, Sch94b]. Distances [CGSO21, Pul88]. Distinctly [Woo84]. Distinguished [Pro95]. Distributed [Alb02]. Distributes [Ste15]. Distribution [Alb12, BH22, Fri84, PU86, SM00, Som03]. Distributions [AFL^{+}21, MS91, MR03]. Distributive [Cas87]. Distributivity [Gre75]. Divergence [Eck97, LM06]. Divergent [Ash97]. Diverging [Sch86b]. Diversions [Wal86]. Diversity [FGL^{+}18]. Divide [MG97]. Divided [Ric15a]. Dividendo [Kob14]. Divides [MS17a]. Dividing [FT18, Rei73b, SB97, Suz99]. Divisibility [CM14a, FH21, Gan17, Hut83, Prc82, Whi85b]. Division [Bar10, Bur99, Coc76, Daw97, Fre95, PC81, Ruc87]. Divisions [Bro23]. Divisor [PH83]. Divisors [AS10, Mer22a, OHC07]. Do [Atk02, Bes14, BC04, Bro99a, Dan98, Dav84b, Dav85, Dic06, Fak99, Har96a, Mal16, MS16, MP07, Par13, Pen03, PG06, Ren82, Sch11b, WN94, Wey23, Wil72, smbV16]. Doctor [Coo74, Fei76, Smu82a, Smu82b]. Doctorate [Pri73b]. Dodecagon [Dr24a, Nel15e]. Does [BH93, Boa86, Bra00, Gla93, Joh93b, MS76, SS83a, VC10, Zho12a]. Doesn’t [TYM88]. Dog [Ric19a, CdSS17]. Dogs [BI10, Dic06, MP07, Pen03, PG06, Sch11b, Wey23]. Doing [GS90b]. Dolciani [Hil92b]. Dollars [Adn98]. Domain [Cur10]. Domination [DT13, Die81]. Dominoes [Men04]. Don [AS82a, AS82b]. Done [HSL77, Ken00]. Don’t [BI10, Bro99a, KS12]. Donuts [NK22]. Doodling [Gre73]. Doom [Nie21]. Doors [KK93b]. Dot [Jon11]. Dots [Hil90]. Double [Bom09, Cid21, GCM18, Hug89, Kob13a, LPT11, Lu22, MMOM14, Nel89a, Sch80a, Sub15, Una10]. Double-Angle [Nel89a]. Double-Crossing [LPT11]. Double-Sidedness [GCM18]. Doubling [Moo86]. Doughnut [vR97]. Douglas [Goe11]. Dover [Hop19]. Down [CE23, Gla93, HP14, KC87]. Downs [Ste12]. Draft [Daw97]. Draining [GE09]. Dramatic [Wil02b]. Draw [Jon04, Ken91, Sul16]. Drawer [SS99]. Drawing [Bai98, McC23]. Drawings [Ber94]. Dream [HMN^{+}11]. Dreidel [Tra96]. Driver [Iga08, Wil99b]. Drivers [IK06]. Drug [Bra00, HJ17]. Duals [Ara11]. Dublin [Ser17]. Dud [Eck81]. Dull [Swa05]. Dunnit [GN74]. Duration [Heg96, OAHS21]. During [EB99, Bei20, Pag20, WW20]. Dust [KR98]. dx} [SS83b]. Dyed [CBL19]. Dynamic [Bri96]. Dynamical [Bla91, Bla94, Dev91b, Kat18, San91, Yer97]. Dynamically [WM13]. Dynamics [DW09, Kie20, KR11, PP21, Wal95b]. Dyson [Alb94].

Each [Nel15c, Sch99b]. Earlier [MT22, PJS81]. Early [Ria95, Wey23]. Earth [Dee99, Gar24, Rou13]. Easiest [Tor11]. Easy [Bum86, For17, Hat83, Jos16]. Eccentrically [Gas94]. Eccentricity [Ayo03]. Echelon [Lay93, Lot23, VB86]. Eclectic [GRSS19]. Eclipse [MT19]. Ecosystems [GECW13]. Ed [Len74]. Eddy [Par95]. Edge [KR98, Poo12]. edition [Ols13, Ser17]. Editor [Lak24a, Lak25a, ACF76, Ano70f, Ano88g, Ano20g, BEL75, Bra94, Coh77a, Fos99, FBF78, GKB^{+}75, GGF78, Has71a, Hop14b, Hop14c, Hop15, Hop18b, Kly19, LMC^{+}70, Lak24d, Lak24b, (Ed24, Lak24e, Lak24c, Lak25b, Lof78, PSSP76, She76, SP71, TD77, TP74, Wen74, Wes94, Woo72a]. Editorial [Alb79, Ano70c, Ano70d, Ano71d, Ano71e, Ano72c, Ano72d, Ano73d, Ano73e, Ano73f, Ano74c, Ano74d, Ano74e, Ano74f, Ano75a, Ano75b, Ano75c, Ano75d, Ano76a, Ano76b, Ano76c, Ano76d, Ano77a, Ano77b, Ano77c, Ano77d, Ano77e, Ano78a, Ano78b, Ano78c, Ano78d, Ano78e, Ano79a, Ano79b, Ano79c, Ano79d, Ano80d, Ano80e, Ano80f, Ano80g, Ano80h, Ano81e, Ano81f, Ano81g, Ano81h, Ano81i, Ano82a, Ano82b, Ano82c, Ano82d, Ano82e, Ano83a, Ano83b, Ano83c, Ano83d, Ano83e, Ano84c, Ano84d, Ano84e, Ano84f, Ano84g, Ano85b, Ano85c, Ano85d, Ano85e, Ano85f, Ano86a, Ano86b, Ano86c, Ano86d, Ano86e, Ano87a, Ano87b, Ano87c, Ano87d, Ano87e, Ano88b, Ano88c, Ano88d, Ano88e, Ano88f, Ano89b, Ano89c, Ano89d, Ano89e, Ano89f, Ano90a, Ano90b]. Editorial [Ano90c, Ano90d, Ano90e, Ano91a, Ano91b, Ano91c, Ano91d, Ano91e, Ano92a, Ano92b, Ano92c, Ano92d, Ano92e, Ano93a, Ano93b, Ano93c, Ano93d, Ano93e, Ano94c, Ano94d, Ano94e, Ano94f, Ano94g, Ano95e, Ano95f, Ano95g, Ano95h, Ano95i, Ano96d, Ano96e, Ano96f, Ano96g, Ano96h, Ano97d, Ano97e, Ano97f, Ano97g, Ano97h, Ano98b, Ano98c, Ano98d, Ano98e, Ano98f, Ano99a, Ano99b, Ano99c, Ano99d, Ano99e, Ano00a, Ano00b, Ano00c, Ano00d, Ano00e, Ano01f, Ano01g, Ano01h, Ano01i, Ano01j, Ano02f, Ano02g, Ano02h, Ano02i, Ano02j, Ano03f, Ano03g, Ano03h, Ano03i, Ano03j, Ano04f, Ano04g, Ano04h, Ano04i, Ano04j, Ano05f, Ano05g, Ano05h, Ano05i, Ano05j, Ano06f, Ano06g, Ano06h, Ano06i, Ano07e, Ano07f, Ano07g, Ano07h, Ano07i, Ano08f]. Editorial [Ano08g, Ano08h, Ano08i, Ano08j, Ano09f, Ano09g, Ano09h, Ano09i, Ano09j, Ano10b, Ano10c, Ano10d, Ano10e, Ano10f, Ano11c, Ano11d, Ano11e, Ano11f, Ano11g, Ano12b, Ano12c, Ano12d, Ano12e, Ano12f, Ano13b, Ano13c, Ano13d, Ano13e, Ano13f, Ano14b, Ano14c, Ano14d, Ano14e, Ano14f, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Lat75a, Lat75b, Lat76, Lat77, Pag88]. Editorials [Ano74g, Ano75e]. Education [And96, CS74, Fri78, HSL77, Man70, Mat96, Mit74, SS93, You70]. Edwards [O’H85]. Effect [Kie20]. Effective [Pag79b]. Effects [Bur77, CH99, Dou73]. Efficient [Jos83, KP85, Kir88, Wie87]. Efficiently [Gra11b]. Egyptian [Ann98, BW15, NRT20]. Eigenargument [Law21]. Eigenpictures [Sch95a, ZF97]. Eigenstuff [Hoo00]. Eigenvalues [Due96, HM23, LM92, Tho00, VK00, Zho17a]. Eigenvector [Sch95a]. Eigenvectors [HB92, Tol22]. eight [Nor09]. Eighteen [Nor09]. Eighteenth [Joh76]. Eighty [Nor09]. Eighty-eight [Nor09]. Eisenstein [LP94]. Election [NZ03]. Elections [Bir10]. Elective [Set74]. Electronic [Arg84, Hop18a, Kat18, Hop18a, Kat18]. Elementary [And80, Bai89, Bel84, Bia98, BF77, Boy94, BM93, CJLP92, Cla73, CM14b, Dub92, Ekl01, FM77, Gao97, Gen73, Gra11a, Hof16, Joh22, Jor77, Ken99a, KP86, Lia70, Lot23, Mar82, Mey91, Moo23, Mor16, Mor15, Mue20, MS22, Mye77, OAHS21, Pap17, Per71, Sal18, Ste77, Var76, Wal04b, Wat89]. Elements [Lyc23, PS12, Ano81j]. Elephants [Boa80]. Eliminating [Ziz98]. Elimination [Her93, Roz88, Yer97]. Ellen [AR88]. Ellipse [Bro84, Cap17, HW84, Kal06, Kin13, Moo85, Moo89, Niv79, Pfi88, Seg00, Usi86, Win91]. Ellipses [AN16, Jon93, Par83, Pfi91]. Ellipses-Revisited [Jon93]. Elliptic [Bro00, Bro01]. Elusive [Kre10]. Elvis [BPP15, Mal16, Zho17b, Zho18]. Emendations [Ano09a, Ano10a, Ano11a, Ano12a, Ano13a]. Emergency [BW19, CHS87]. Emmy [Sri10]. Empirical [KL11, Led16]. Enclosing [FM00]. Encounter [MR96]. Encounters [Jon20a, Lan76]. Encouraging [Mai70]. End [MS16]. Endless [Ald23]. Endnotes [Lak24a, Lak25a]. Ends [Lie97]. Energy [Cho81]. English [Sch95d]. Enhancing [Wag87]. Enjoyment [EK22]. Enough [DW88, Hil85b, KH84, Shu02]. Enrique [Bra13]. Ensphering [SH14]. Entrance [Bra90]. Entries [IB22]. Enumeration [MW03]. Envelope [Sim97]. Envelopes [Cam19, SS22]. Environment [Wat85, Whi88]. Environmental [Lam77]. Eötvös [Wie90]. Epicycloids [GG84]. Epidemic [Roz07, Wan10]. Epidemiology [GO12, Mar13]. Episodes [Bra13]. Epitrochoid [SA99]. Epsilon [Yan80]. Epsilons [Ben76b, Lin74b]. Equal [AS96, Bea09, BN03, Hig23, PS82a, Par84]. Equalities [BN03]. Equality [Jus89, Pro76]. Equalizing [Tam87]. Equally [Bon14]. Equals [Law24, NS87]. Equation [AA04, Ban90, BS00, Cho80c, CM20, D’A23, DeT84, DD95, HMNW94, Hub94, HS00, Joh22, Jon20b, Lau20b, Lut80, MM85, Par13, Par16, ST19, Sew18, Ves85]. Equations [Ali17, Bla94, Bom82, Boy94, Boy15, BJ24, Bus71, Che96, Che22, Chr96, Dav94, Ehr94, Ett74, Far15, Fla86, Gre76, Gug92, HM95, Hor90a, HW84, Kal01, Ken91, KK13b, Lom94, MF94, MS78, Mis23, MT97, Mun04, MM85, Nas85, Nie03, Oli73, Sac22, SS77, Sch85a, SS94, Str04, UK96, Wol23, CG19]. Equiangular [Boy07, Sco87]. Equilateral [AN12, Ber20, Oku22]. Equivalence [HJ17, Liu11, Pin74a, Ric22, SB84]. Equivalences [Bel84]. Equivalent [Dob72, HH88]. Eradicating [GO12]. Erdos [Ale81, Ser15]. ergo [Sup84]. Erratum [Ano92f, Ano19f]. Error [Bor11, Eis97, Liu01, Mer05, Min91, Moo78a, Moo80]. Error-Correcting [Liu01]. Error-Detecting [Moo80]. Errors [Sal88]. Escher [Ber14b]. Escribed [Lon83]. Establishment [Ald79a]. esteem [Wil90]. Estimate [And81, Bou24, GEJ95, Ham20]. Estimates [Mer05]. Estimating [Har88, Ros01, Woo25]. etc [Mat94]. Eternal [Lar89]. Ethics [Shu02]. Euclid [Ano81j, Ben94, CS03, Ste81b]. Euclidean [BMO^{+}00, DE77, EG75, Hob13, Kun74]. Euler [Ada80, AMR13, Bil18, Bor11, BS04, Bur85, CC23, De 10, Dun91, Dun20, FK94, GG23, Ham20, Jon12b, Kat99, KS06, Lew84, LPP08, May13, Moo91, MM85, Par16, Sal18, SS82, ST19, Sha15]. Eulerian [Kri95]. Europe [Man70]. Evaluate [Str79, Wil24b]. Evaluating [BGR86, Gob88, Gor84, Gor89, RT07, Sch73, SS83b, Wie83]. Evaluation [Bit70, FL75, Max74, Muk24, Pri90, Sch81b, Smi91, WE73]. Eve [Tat02].

Even [Asa84, Bén05, Ful05, Her12, PS22, Thu15, Wu23b]. Events [Ano70a, Ano70b, Ano71a, Ano71b, Ano72a, Ano72b, Ano73a, Ano73b, Ano73c, Ano74a, Ano74b, BB16, FBH83, Mon82]. Ever [BL22a]. Every [Nel15d, Sch00, Sou14]. Everything [Ano71f, Tho92, Ano70e]. Everywhere [DI08]. Evolute [SS00]. Evolution [Kle89, Sim02]. Evolutionary [Bar09]. Exact [IL24, LS01, Mac92b, Ros76]. Exactly [Bum83, HS86, Mic12]. Examination [Bra90]. Examinations [EW70]. Examine [Ive88]. Examining [McM03]. Example [CM17, GP76, MS23, OB79, Pal98]. Examples [Boy81, GV23, Min94, Sch90b]. Exams [Dun23, Spi10]. Exceptional [Ćur06, Sch74c]. Excerpts [Ano15a]. Exchange [PS82b]. Exclusion [Mel22, Tre18]. Excursion [Cid21]. Exercise [Arm74]. Exercises [Ham85]. Exhibit [AG14]. Existence [MS17b]. Exists [Ash97]. Exotic [Car98]. Exotic-Looking [Car98]. Expansion [Bri91, Eul93]. Expansions [ANANFT01]. Expect [FPS13, ZV07]. Expectations [Huf89, LM16]. Expected [Gre20, Hal80, LM10, Shu79]. Experience [GK88, Ska91]. Experiment [Bru83, BEN78, Gru94, Lia70, Per71, SW76]. Experimental [Boa71b, Tre94]. Experiments [Lom94]. Experts [Han88]. Explain [Hof82]. Explains [Ano11b, FN09]. Explicit [Abb24, Cer15, Coo82, Wan21]. Exploiting [BJKL97]. Exploration [BNUW23]. Exploring [Ald23, AC22b, PDR21, Ryd96, Ste76a, Str90b]. Exponential [AL05, Bil93, Ćur06, DW09, Dra13, EI89, Fee01, FM05, GKR19, Gor94, HM95, Kal01, Kha03, KS24, Nel88, Nel92, Sch85a, TB20]. exponentials [Den12]. Exponentiation [KP85, Ste15]. Exponents [Mor10a]. Exposition [CJLP92]. Expression [Wan21]. Expressions [IL24]. Extended [HP99a]. Extending [Kat12, Mer22b, Per94, PDR21, Rod80, Sch87a]. Extension [Abb10, Arn94, Die23, FF24, Gin04, Gre00, HS86, Put86]. Extensions [Adl96, CS06, Kre17]. Extracting [Mar24]. Extrapolating [Sim19]. Extrema [BM97]. Extremal [AFL^{+}21]. Extremely [Kru00]. Extremum [MP88]. Eyeball [Gar22]. Eyes [Yin87].

F [Saa70, SS83b]. F.R.S. [TM01]. FA [Sul16]. Fable [FH95]. Face [KN15]. Facilities [CHS87]. Facility [Buc87]. Fack [Cra11b]. Factor [Cur05b]. Factored [Dor77b]. Factorial [Sch84d]. Factorials [Fri81, Kie85, Tre94]. Factoring [Ben17, BB01, Bod71, Bri75b, Fri81, Luc98, Pra09, Rat81]. Factorization [BJKL97, Her93, SS77, UK96]. Factorizations [CGSO21, Fag20, Wal87]. Factors [Gre77]. Faculty [DP71]. Fair [Bar10, Daw97, Fra87, Guy84, Lip03, PS82b, Tra96]. Fairer [Cra74]. Fairest [BB11]. Fakin [Ros18]. Fall [Cur05b, Oss81, Ros01]. Fall-Factor [Cur05b]. Fallacies [Amm00, Ano89g, Ano89h, Ano89i, Ano89j, Ano89k, Ano90j, Ano90f, Ano90g, Ano90h, Ano90i, Ano91f, Ano91g, Ano91h, Ano91i, Ano91j, Ano92g, Ano92h, Ano92i, Ano92j, Ano93f, Ano93g, Ano93h, Ano94h, Ano94i, Ano94j, Ano94k, Ano95j, Ano95k, Ano95l, Ano95m, Ano95n, Ano96i, Ano96j, Ano96k, Ano96l, Ano96m, Ano97i, Ano97j, Ano97k, Ano97l, Ano97m, Ano98g, Ano98h, Ano98i, Ano98j, Ano99f, Ano99g, Ano99h, Ano99i, Ano99j, Ano00f, Ano01k, Ano01o, Ano01l, Ano01m, Ano01n, Ano02k, Ano02l, Ano02m, Ano02n, Ano02o, Ano03k, Ano03l, Ano03m, Ano03n, Ano03o, Ano04k, Ano04l, Ano04m, Ano04n, Ano04o, Ano05k, Ano05l, Ano05m, Ano05n, Ano05o, Ano06j, Ano06k, Ano06l, Ano06m, Ano06n, Ano07j, Ano07k]. Fallacies [Ano07l, Ano07m, Ano07n, Ano08k, Ano08l, Ano08m, Ano08n, Ano08o, Bur94, Cla92, FG93, Gue00a, Gue00b, Lax98, Mil93, Tee00]. Falling [BM88, BJK17, SS96]. False [Bom09]. Familiar [BGR86, Fon96, JS75, Sch72a, Sch84a, Sch86d, Sch91b, Shp14]. Families [BC18b, Cam19, DRW96, FH11, FST23]. Family [BM97, FKY13, FH21, Mih02, Sou14]. Fano [Ehr21]. Fans [Pri95]. Fantasy [AS22]. Far [CS86a, Edg18, Lee99, Ruc87]. Far-reaching [Lee99]. Farewell [Pag88]. Faro [Ano91q]. Fascinating [HB97]. Fashion [Guy84]. Fast [Du20, Sim00]. Fast-Food-Frusta [Sim00]. Fast-Growing [Du20]. Faster [Gol83b]. Fastest [BPR15, BMV14]. Fate [Sch89e]. Father [AR86]. Faulhaber [Cer15, Das11, Her12, Kri95]. Favorite [Cha91]. Feature [Lut17]. FedEx [Mor10b]. Feeder [Daw90]. Feigenbaum [KR98]. Feller [Cal95]. Fences [HMNW94]. Fermat [BZ04, BC18a, Bro00, Kat99, Kau20, Lea73, Mor10a, Rat81, SS81a, Whi85a, Zho12b]. Fether [Smu82a, Smu82b]. Fiat [Dav80]. Fib [BF23]. Fibonacci [AG87, BQ99, BB08b, CDNN02, Car78b, CK12, Coo84, HB97, Hay92, Hen94, Hon15, Mar09, Pon17, Pri18, Ryd96, Spi06]. Field [AW90, Isa96, Swa74]. Fields [Jus89, Nag21, Olm73]. Fifth [Hop20, Ros76, Sin12, SS21]. Fifth-Degree [Ros76]. Figures [Cas03, Dun97a, Nie01, Wet10]. Fill [Nor09]. Filled [Gob88, TW90]. Filling [WS81]. Film [Ano14g, Wil16]. Finally [Bei89]. Financial [Din13]. Find [Blå16b, BL22a, Cra73a, HM05]. Finding [AB17, Bén05, Boy10, CM98a, HS00, KH08, Par91, Red89, Rob78, Und84, Win95]. Finest [CFLR12]. Finite [All16, BL78, Bro01, CH17, Dob72, DE77, Ekl00, Jon20b, Lib81, MZ94, Mey07, Nag21, Pet11, Sch21b, Sie12, Swa74, Tou13, Zia91]. Fire [Cib93]. Fireflies [ZGN06]. Fireworks [Dan92]. First [BSC17, CJLP93, Fri21, Fri84, GM99, Hsu82, Kly20, Kou95, Lea18, MC20, Mer17, OB77, PV13, PU86, Per83, Ric15a, Sac22, Sch23, SS21, Str04, Woo70, Wri95]. First-Order [Fri21]. First-Year [Woo70]. Fish [RC20]. Fishing [BC94]. Fit [Ska96, VK80]. Fits [Wil02b]. Fitted [Rog07]. Fitting [Cav93, Led16, Pen19]. Five [BC18b, Mor04, MW83]. Five-Function [MW83]. Fixed [AL15, But81, MR95, TT21]. Flag [Sch23]. Flagpole [DR06]. Flatland [Sch06]. Flattening [Ken20]. Flaws [Amm00, Ano89g, Ano89h, Ano89i, Ano89j, Ano89k, Ano90j, Ano90f, Ano90g, Ano90h, Ano90i, Ano91f, Ano91g, Ano91h, Ano91i, Ano91j, Ano92g, Ano92h, Ano92i, Ano92j, Ano93f, Ano93g, Ano93h, Ano94h, Ano94i, Ano94j, Ano94k, Ano95j, Ano95k, Ano95l, Ano95m, Ano95n, Ano96i, Ano96j, Ano96k, Ano96l, Ano96m, Ano97i, Ano97j, Ano97k, Ano97l, Ano97m, Ano98g, Ano98h, Ano98i, Ano98j, Ano99f, Ano99g, Ano99h, Ano99i, Ano99j, Ano00f, Ano01k, Ano01o, Ano01l, Ano01m, Ano01n, Ano02k, Ano02l, Ano02m, Ano02n, Ano02o, Ano03k, Ano03l, Ano03m, Ano03n, Ano03o, Ano04k, Ano04l, Ano04m, Ano04n, Ano04o, Ano05k, Ano05l, Ano05m, Ano05n, Ano05o, Ano06j, Ano06k, Ano06l, Ano06m, Ano06n, Ano07j, Ano07k]. Flaws [Ano07l, Ano07m, Ano07n, Ano08k, Ano08l, Ano08m, Ano08n, Ano08o, Bur94, Cla92, FG93, Gue00a, Gue00b, Lax98, Mil93, Tee00]. flex [IMW12]. Flexagons [Poo12]. Flexible [Gia73, Lon76]. Flight [Boy99]. Flimflam [Amm00, Ano89g, Ano89h, Ano89i, Ano89j, Ano89k, Ano90j, Ano90f, Ano90g, Ano90h, Ano90i, Ano91f, Ano91g, Ano91h, Ano91i, Ano91j, Ano92g, Ano92h, Ano92i, Ano92j, Ano93f, Ano93g, Ano93h, Ano94h, Ano94i, Ano94j, Ano94k, Ano95j, Ano95k, Ano95l, Ano95m, Ano95n, Ano96i, Ano96j, Ano96k, Ano96l, Ano96m, Ano97i, Ano97j, Ano97k, Ano97l, Ano97m, Ano98g, Ano98h, Ano98i, Ano98j, Ano99f, Ano99g, Ano99h, Ano99i, Ano99j, Ano00f, Ano01k, Ano01l, Ano01m, Ano01n, Ano02k, Ano02l, Ano02m, Ano02n, Ano02o, Ano03k, Ano03l, Ano03m, Ano03n, Ano03o, Ano04k, Ano04l, Ano04m, Ano04n, Ano04o, Ano05k, Ano05l, Ano05m, Ano05n, Ano05o, Ano06j, Ano06k, Ano06l, Ano06m, Ano06n, Ano07j, Ano07k, Ano07l]. Flimflam [Ano07m, Ano07n, Ano08k, Ano08l, Ano08m, Ano08n, Ano08o, Bur94, Cla92, FG93, Gue00a, Gue00b, Lax98, Mil93, Tee00]. Flimplam [Ano01o]. Flipping [Cra74, Spe86, Zuc09]. Flips [Ros18]. Flitting [Nat19]. Floating [Con17]. Floating-Arm [Con17]. Floats [Bro05]. Florin [Ols13]. Flourishing [Lin21]. Flow [FM77, Kem82, Shi78]. Flowering [Bur92]. Flowers [Far15]. Fly [MV12, Nat19]. Flying [RC20]. Foci [Job11, Moo85, SH10]. Focus [AM10, BR99]. Fog [Car93]. Fold [CG24, FN09, MC20]. Fold-over [MC20]. Folding [Ber06, Fre03, Fre11, Pri95, WC99]. Folds [KR11]. Follow [Hat83, Ler21, Tur85a]. Follow-Up [Ler21, Tur85a]. Food [Bra00, Daw90, Sim00]. Football [Gal81b, Jon04, LP16, Wil02a]. Forced [Pro95]. Forcing [Fay97]. Foresight [FS79]. Forest [Zob13]. Forget [Mun04]. Forgotten [Ali17, BB08b, Mar09]. Form [Bla16a, Car78a, Cer15, Hos18a, Sch88b, Seg83, ST19, Ven80, Lot23]. Formal [Gal81a, Ste74]. Format [BB76]. Forms [Cos94, FS13, Lay93, Ziz98]. Formula [Ada80, Alp87, And80, Boo79, Bra86, Coc76, Coo82, Dob84, Fon96, FS15, Fre86, Fri89, Hen94, Her04, cH94, Ins89, Jan80, Kal84b, KKS90a, Lee99, Lin85, Mac95, Moo91, Nel21, PL81, Sal18, San24, Sch74c, Sch94a, Sch86f, SY78, TB20, TT21, Una14a, Whi71, Wil24b]. Formulae [EH16, MP14, Men88]. Formulas [Ash20, Car87, Che95, Gug96, Hoe09, MT23, Nel89a, PK83, Sch86d, Seg83, SM91]. Formulations [Man98]. Forringer [Spo23a]. Fortresses [Mic11]. Fortune [RS16]. Forum [Pag84a, KGSP88]. Forward [McG90, Ste81b]. Foster [EB99]. Found [Joy14]. Fountains [Sim09a]. Four [ASW94, BMV14, Cus83, DD10, Gre75, Mey91, Mit81, Nor09, Reb89, Rei73a, ST98, Tea06]. Four-Bit [Reb89]. Four-Color [Mit81]. Four-Function [Cus83]. Four-gon [ASW94]. Four-Letter [ST98]. Four-Minute [BMV14]. Fourbar [CG24]. Fourier [Hug00, KD92, Kos17]. Fourth [Lea18, Som00, SS21]. Fractal [Bar84]. Fractals [CMS20, CMP15, Man84, Nat91, Nie91, Wal96, Wil91, Nat91, Nat91]. Fraction [Par05]. Fractional [KO00, MC24, Mor10a]. Fractions [Ann98, BW15, Bur84, Dal19, FG10, Gur19, Hen91, Hua91, Jos83, Kal96a, Nym83, Pac78, Pla78, Sch83c, Stu13, Wie86, YD07, Den12]. Fred [Ano84h]. Fredonia [Col71]. Free [Abb93, Dav92, SH12]. Freeman [Alb94]. Frénet [Hoe09]. Frenzy [Smi14]. Frequencies [Mur81, Pug80, Tre94]. Frequency [Gre20]. Frequent [Gre20]. Fresh [MR98]. Freshman [Dou73, Wil72]. Freshmen [BEN78]. Fresnel [Che09]. Friend [BK11]. Friends [Mih02, Sch21a]. Friendship [Nun10]. Frieze [smbH02]. Fringes [Cul90]. Fritsch [Par95]. Frusta [Sim00]. Frustum [CSTM10, Kwo16]. Fubini [AM02]. Fun [Val15]. Function [Ano92f, AS96, BF13, Bri96, Cha92, CC80, Ćur06, Cus83, Ekl01, Ell83, Fee01, GKR19, GS90a, Goe11, Gor94, Har96b, HM95, Hon15, Huf89, Hun98, Kle89, KS24, Kra03, LeD73, LPP08, Mal17, MR96, MP10, Mer22a, MW83, Mur80c, Ni16, Ovc11, PY04, Pen19, PL94, Ric22, SH18, Sil89, Sim95, SS81b, Tro04, Wan10, Xu10, Xu14]. Functional [Gao97, Hol76, MS78]. Functions [AB16, Ald79b, AL05, Ash20, Atk02, BKSU19, Ben75, Boa85, Bod71, Bra85, Bri75b, Bri78, Buc85b, Bum83, Bus71, BW91, CCH^{+}21, CP89, Cha07, Cip87, Cra77, DC95, DW09, Din22, Dob85, Dra13, DL75, EH16, EI89, FM05, FM94, GS03, Gra95, GL00, Ham07, Har72, Hig23, IL24, Jun74, Kar22, Key94, Mac03, Mas91, MDC80, MR03, Mor91, MS22, Nan14, Nel88, Nel92, O’L98, OT72, PS79, Pin75, Pio21, Pon17, Rib97, RR98, Sal85, San21, Sha95, Sno89, Sur11, Wat87]. Fundamental [Bel15, CP89, Dun91, Hir98, Mac95, Pal98, PL09, Ton15, Zob13, BSC17]. Funnels [HMNW94]. Further [HP99a, Jan80, Nel18, Sch74d]. Future [Ano81a, Ano81b, Fre12, Kle71].

G [Ste83]. Gabriel [Fle99]. Gaetana [Gra18]. Gaie [Mey06]. Gaitpath [Lor00]. Galileo [ST20]. Galleries [Mic11]. Gallian [Hau08]. Galois [BSC17, Ehr21]. Gambler [BS14a, MZ24, OAHS21]. Gambling [Con99]. Game [CCHR24, CC19, Cra73b, Ehr24, FG07, Gol16, Han70, KRSW18, LM92, MP03, Mar98b, MS20, McC02, NRT20, Pri76, Ruc87, Tam87, Tra96, TYM88, NS17]. Games [Alb05a, Alb05b, Gri17, HR84, KX15, McC01, Pie71, Str85, Tor11, Vod97, Wil02a]. Gamow [Lin00b]. Gap [BB08a, DM03]. Garden [KO00]. Gardner [Alb05a, Alb05b, Ano09k, Bar79, Kho12, MR79]. Garrett [AW83]. Gas [Shu02]. Gateway [HP93]. Gathering [Bos00]. Gaussian [Ben94, Her93, Roz88, Ste76a, Sza05, Wat12, Yer97]. Gavin [Hen22]. gcd [KC90, Sza05, Law24]. Geek [Mul19]. Gehman [Ald81]. Geloxia [Suz99]. Gem [Mar82]. Gems [CJLP92, Hon79e, Hon79a, Hon79b, Hon79c, Hon79d, Hon80a, Hon80b, Hon80c, Hon81a, Hon81b, Hon81c, Hon82a, Hon82b, Hon83a, Hon83b, Hon95]. General [BS00, CC80, DeT84, Fri78, Gra11a, HSL77, Mit74, MM85, Sch88b, Whi71]. Generalization [BZ04, Bur83a, CC23, Gor03, Gre00, Kos09, Mas11, OB79, Say95, Sch89d]. Generalizations [HS96, Kat03, KM89, KS96]. Generalized [Blo01, CF17, Coo84, DeT88, Hil85a, How02, JHG11, Kal01, Meh22, Sch16, Sob97, Tru19, Woo75, Woo25, Kos17, Pow80]. Generalizing [Gen73, PS03]. Generate [Rib97]. Generated [For17, Mat25b, Sur11, Tre20, Nor88]. Generating [Ald79b, Car98, Ell83, EI89, Hon15, MR03, Pin75, Pon17, SS81b, Sno89, Wat87]. Generation [Lan77]. Generator [Pin74b]. Generators [Reb89]. Generic [HJ17]. Genes [Wri95]. Genesis [Joh72b]. Genocchi [CM17]. Genocchi-Peano [CM17]. Genomics [Han20]. Genuine [Fre95]. Geometer [CdSS17, Hoe97, Jur94]. Geometic [MP88]. Geometric [Adl96, Bar10, Bel84, Ber14a, Ber20, Blo01, Bou99, CS06, Cia17, DeT84, DG80, DF96, Dun97a, DL73, Eck98, Edd85, GKR19, Gra95, Ham07, cH94, Hol81b, Kat03, Kob98, LP94, Mar98a, MM93, Mat92, Muk24, Mus79, Nel89b, Nel15a, Nie78, PM82b, Pet14, Pla18, Pow80, Rat81, Sar23, Sch78a, Sch82c, Sch88b, Sch92a, Sch72c, SY87, Seg83, ST20, Sem74, SR16, SK01, Tan86a, Tir86, Tou95, Wu23a, Wu24]. Geometrical [Bal78, Bid86, Hoe84, Hor90a, Sch76]. Geometrically [Kal85a]. Geometrico [LB94]. Geometries [DE77, FS15]. GEOMETRY [Grü81, Alb81a, Bar84, Boy07, BL78, Can86, Die81, Dil14, ES23b, Far00, Gao97, GRSS19, Gol16, Gro96, Gug96, Her02, HP93, HP99a, HP99b, Kle72, Log80, Mar10, McC98a, MW97, NW10, OS81, Oss81, Ped80b, Ped98b, Sch81b, Sch86d, Swa74, VK80, Vie78, Yiu12]. George [Ser17, AR86, Ale79d, Ano96n, Ano06o, Ano08p, Ano08q, Ano11h, Ano12g, Ano14h, Ano14i, Ano15g, Ano16f, Ano17f, Ano18f, Ano19g, Ano20f, Ano21h, Ano22d, Ano24g]. Gergonne [Moo20]. Germain [MGC19]. Gerrymander [Suz14]. Gerrymandering [HMP10]. Gestures [LR20]. Get [Boa81a, Hen22, Lan95]. Gets [Wri95]. Getting [Ste77, TL00]. Giant [Sim11b]. Gini [LZ18]. Girard [Wil24b]. Girls [Col13]. Give [SS83a]. Given [Lan93b, Moo85, Pla17c, Woo00]. Glass [Che22, DHR10]. Glasses [IK06]. Gleanings [Coo70]. Glimpse [Ale83b]. Global [Kac17]. GM [Hof16, Mah20, Sch89b]. Go [Tee07]. Goal [Isa96]. Goat [Fra84, Mey84]. God [Alb04, JSW14, Joy14, Rok14, Ruc82, Gra18]. Gödel [Hof82, Wie75a, Wie75b]. Goes [Gla93]. Going [Due96, SH81, Wal11]. Goldbach [EKP15]. Golden [Gra23, Nel25, Spo23b, Edw14, Fal05, FH11, MS88, Mar92b, Nel16f, SW95]. Gompertz [Byr11]. gon [ASW94]. gons [GV23, See98]. Gonzalez [Col13, Bra13]. González-Velasco [Bra13]. Good [Ald88, AM01, Car77, Dav94, Fla11, Kro79, KH84, Ren82, SH81]. Gould [Bro06]. GPS [Kal02]. Graceful [KR91]. Grades [She77]. Gradient [HM05, Mah20]. Graduate [Lai70, McK79]. Graeco [KS06]. Graeco-Latin [KS06]. Graham [Kol81]. Grain [Knu18]. Grandma [BG15]. Granola [BG15]. Graph [Ano93i, AC22b, BF13, BDCH85, CCKL25, CEO88, DW88, Kin13, Ols96, Qui16, RS96, Sha18, Wag14]. Graphical [DW88, FN07, Gre76, Hor90a]. Graphics [Fra86b, GG84, Mas91, Pen86, Wag87]. Graphing [DeT84, Lar75, LH89, SS89, Ves85]. Graphs [AS10, Ayo87, Bei89, Bic17, BD96a, BD96b, Bre83, BW91, Ess82, Fak99, GK13, KR91, Mor91, PL94, Sha18, Str93, STW17, Wal87]. Grashof [CG24]. Gravitation [Lin00b]. Gravitational [Jus89]. Gravity [Sim00, Str72]. Grayscale [LZ18]. Grazing [Fra84, Mey84]. Great [Alb81b, LS94]. Greatest [PH83]. Greatness [Som02]. Greedy [Pau19]. Green [Ber14b]. Greenhouse [Kie20]. Gregory [Grz24]. Grid [Bro23, CHS87, RS23a]. Gridiron [Isa95, Nie78]. Grids [Abb93, Dav92, Hof78a]. Grinstead [Gol13]. Grizzly [CKM86]. Gröbner [ALT10b]. Ground [CF17]. Group [Bia98, Bre08, Cri11, IM11, Joh00, Joh13, Moo95, Ped72, Som80, Tea06, WJ73, Win95, CJLP93]. Group-Based [WJ73]. Groupoid [BW15]. Groups [Dob72, GK77, Gro05b, JB92, LN14, Lea76, LW18, MDL11, OS16, Pet11, Sie12, Tou13, Wei76b, smbH02]. Growing [Du20, Tuc93]. Growth [App01, Mos88]. Guards [Mic11]. Gudermann [Rob97]. Guessing [Ahu79, Mar98b, McC02, Pól78, Pól79]. Guest [Wes94]. Guide [Col13, MZ24, RC20]. Guided [Han20, Sie12]. Guidelines [Ano79e, Ano84i, Ano85g, Ano86f, Ano87f, Ano89l, Ano97n, Pri89]. Guitar [Hug00]. Gunnery [Gro97]. Guy [AA93].

H [Nat91]. H. [Bro06, Log80]. Hadamard [Nel23]. Half [Den07, Sol23]. Half-Row [Den07]. Hall [DY04]. Halley [Gro97]. Halmos [Alb82, Alb04]. Hamilton [RB21]. Hammer [Ahn09]. Hamming [Ehr21, Liu09, Nag21]. Hand [BDLS18, Gao13, Sul09, Thu15, Wai78]. Hand-Held [Wai78]. Handbook [Hen22]. Handed [AKMMM23]. Hands [JK98]. Hanging [Buk08, Daw90]. Hank [BBHS02]. Hanoi [Bon18, IB13]. Hanoi-ing [Bon18]. Harcourt [Ric19a]. hardback [Gra18, Ser17]. Harder [Bon16, Ros01]. Hardest [Mab10]. Harmonic [AE17, AE19, Bén05, Bob95, BCEW85, CK12, Dun87, Eck97, Fay97, FN04, Gor24, Les04, LM06, Nel89b, Pro95, Zuc16]. Harmonic-like [LM06]. Harmonic-Type [Les04]. Harmony [DK08]. Harris [Sha18]. Harry [Ald81]. Harvesting [CKM86]. Haste [Eri22]. Hats [Hat04, Hat07]. Hausdorff [Koh12]. Having [ST19]. Haystack [Cra73a]. Haystacks [Fra88]. Headquarters [Ald78]. Heads [Sch90a]. Heating [SE96]. Hebrew [CAS20]. Hedging [PV16]. Height [Mar90b, Som02]. Held [Wai78]. Helen [Mar10]. Helical [Far08]. Hell [Ren85]. Help [Mas91]. Helpful [Ros77]. Helsinki [Lan80]. Henry [AT84]. Heptagon [LR24]. Heptagonal [Cag24a]. Herbert [CL90, Pag84b]. Here [DI08]. Heresy [Law71]. Hermann [Una14a]. Hermit [HGGB08]. Hermite [Nel23]. Heron [Alp87, CV14, Dob84, Her04, Pra09, San24]. Heuristic [Hug89]. Hexaflexagons [Gar12, IMW12, Poo12]. Hexagon [Aug12, HKM22, HP99a, HP99b, Nel16g]. Hexagonal [Cag24b, Wag14]. Hexagons [AN16, Wu23a]. Hexagram [KH08, RR89]. Hexagrams [Gar00b]. Hidden [Bai91, WD90]. High [Hal13, LS94]. Higher [HP99b, RC20]. Highlights [Ano86g, Ano88h, Ano88i, Ano88j, Ano89m, Ano93j, Ano93k, Ano94l, Ano94m, Ano95o, Ano96o, Ano96p, Ano97o, Ano97p, Ano97q, Ano98k, Ano99k, Ano99l, Ano99m, Ano99n, Ano99o, Ano00g, Ano00h, Ano00i, Ano00j, Ano00k, Ano01p, Ano01q, Ano01r, Ano01s, Ano01t, Ano02p, Ano02q, Ano02r, Ano02s, Ano02t, Ano03p, Ano03q, Ano03r, Ano03s, Ano03t, Ano04p, Ano04q, Ano04r, Ano04s, Ano04t, Ano05p, Ano05q, Ano05r, Ano05s, Ano05t, Ano06p, Ano06q, Ano06r, Ano06s, Ano06t, Ano07o, Ano07p, Ano07q, Ano07r, Ano07s, Ano08r, Ano08s, Ano08t, Ano08u, Ano08v, Ano09l, Ano09m, Ano09n, Ano09o, Ano09p, Ano10g, Ano10h, Ano10i, Ano10j, Ano10k, Ano11i, Ano11j, Ano11k, Ano11l, Ano11m, Ano12h, Ano12i, Ano12j, Ano12k, Ano12l, Ano13g, Ano13h]. Highlights [Ano13i, Ano13j, Ano13k, Ano14j, Ano14k, Ano14l, Ano14m, Ano14n, Ano15i, Ano15j, Ano15k, Ano15l, Ano15m, Ano16g, Ano16h, Ano16i, Ano16j, Ano16k, Ano17g, Ano17h, Ano17i, Ano17j, Ano17k, Ano17l, Ano17m, Ano18g, Ano18h, Ano19h, Ano19i, Ano22e, Ask92, BB88, Bar91, BS23, BBS24a, BBS24b, BBS25, Ber98, Ber87, Bix95, BC85, Bus96, Cha84, CB85, Dav86, Dev89, Dev95, Dos96, Dud94, Dun97b, ES23a, EBS24a, EBS24b, EBS24c, FA85, Fla90a, Fla90b, Fla90c, Fla90d, Fla90e, Fle93, Her98, Hil84a, Hil96, Hog84, Hon87, Jon85, Kap85, Kap92, Kat92, Kid93, Lan93a, LS19a, LS19b, LS19c, LS20a, LS20b, LS20c, LS20d, LS20e, LS21a, LS21b, Lei21, LS21c, LS21d, LE22, LS22a, LS22b, LS22c, TLSES23, LS23a, LS23b, Mau84, McL87, MS87, NN91, Niv86, Noe88, Ote98]. Highlights [Pag87a, PLS18, Ped98a, Ren89, Ros92, Sau94, Sca91, Sch89a, SSG89, Sch86e, Sel91, Sen95, Str97a, Swe92, Swe95, Tat91, Tat94, Wat84, Whe86]. Highway [CSWW01, DR04]. Hill [CE23, Pag72]. Hindsight [FS79, Sch83a]. Hippocrates [Hug89]. Hiroshi [Col13]. Historian [Wil94]. Historical [Hah98, Tur85b]. History [BSC17, Boy70, CDL^{+}01, Eve02, Gal81a, KP86, Lar89, Len76, Mit81, Ols13, Pie77, Ric19a, Bra13]. Hockey [Hir24]. Hockey-Stick [Hir24]. Hodges [Ser17]. Hold [Boa86, Roj16]. Hölder [Goe11]. Holding [Kir92, Sch92a]. holds [Boa86]. Holes [Cos17]. Holiday [AOW05]. Home [BBHS02, MP12, SE96]. Homework [McG90]. Homicide [Smi78]. Homogeneity [BSF99]. Homogeneous [SS77]. Homogenized [BHB04]. Homographic [Cap17]. Homomorphism [Moo95]. Homomorphisms [Arp74, Joh13]. Honeycomb [DT13]. Honor [MR79]. Hope [Min03]. Hopkins [Gra18]. Horn [Kos20]. Horse [PV16]. Horton [Guy82]. Hossenfelder [Gut19]. Hostile [Bai09]. Houghton [Ric19a]. Hour [Hol79]. House [Pau85]. Housework [VC10]. HP [SS89]. HP-28S [SS89]. http [Den13]. Hudde [Cur05a]. Hulls [Sun21]. Human [FIRW14, Ler21, Lin21, MR79]. Hundred [Nor09]. Hundredth [Tat02]. Hungarian [CL90]. Hur [DS12]. Huygens [Buk08]. Hyperbola [Bro84, Cap17, Fra14, HW84]. Hyperbolic [BS96, Ben75, Bur86b, Dra23, EH16, FS15, GS03, Hob13, JW22, Lam24, Lan76, Nel88, Nel92, Sal85, SR12b, Sha95, Sob95]. Hyperbolicity [BMO^{+}00]. Hypersphere [Pet15]. Hypocycloids [GG84, Sch85d]. Hypotenuses [Fos16].

I.C.M. [Lan80]. Ian [Hop18a]. Ideals [Sur11]. Ideas [Ker83, Ric19a]. Idempotent [Fag20]. Idempotents [Sib12]. Identification [Gal91, KK93a]. Identifying [HP99b]. Identities [Ald79b, Arp74, BQ99, Edg21, FKY13, Gre76, Mol16, Pin74b, Ros89, Sch81d, Sil90, Spi06, Tre20]. Identity [Abe15, AC22a, Bal22, Bal23, BL22a, Hir24, Jon12b, KM89, Kob98, Kob13a, Mal22a, Mat25a, MGC19, Nel14c, Nel16g, Nel16a, Spi16, Sub15, Una14b, Val15, Wan14a, Wan14b]. If [Ger75, HSZ05, MS17a, NW98, Sel70, SH81]. II [Alb05a, Bal23, BD96b, Cha07, Gra99b, HP99b, RY13, Wie75b]. III [ES85]. III-Conditioning [ES85]. Ill [Roz94]. Ill-Conditioning [Roz94]. Illinois [Fei76, GW70]. Illogical [Lin19]. Illuminism [GRSS19]. Illustrate [Nie91]. illustrated [Col13]. Illustrating [Sch90b]. Image [Bri91, KK93a, Tru23]. Images [Hig23, LZ18]. Imaginary [BBB03, Wag77]. Imbalance [Sal14]. Imbedding [Nef83]. Immigrant [Wag99]. Impacts [Gol11a]. Impaired [Wit17]. Impatient [CHJ11]. Implement [Lu22]. Implementing [FLP18]. Implication [Ber75, Sni77]. Implicit [Boy15]. Importance [Ral84, Tuc93]. Impossible [PDR21]. Improper [Fly95, Ste23, Tei23]. Improve [Ken99a, Thu00]. Improvement [Sch90b]. Improving [Bro04, Mit74]. Inclusion [SM13, Tre18]. Inclusion-Exclusion [Tre18]. Income [Fal85, HT84]. Inconvenient [AB16]. Increase [Fal85]. Increasing [FM94, Ler21]. Independence [Bis80, Che05, DT13, Mur77, Ter96]. Independent [BB16, Bis80, Eis17, Her03, Hol23, Mon82, Oma09, SW95]. Indeterminate [Car78a, FS13]. Index [Hoo97, LZ18, MS19, Ano95c]. India [Bre02]. Indices [OS16, PS15]. Indiscretion [Woo84]. Individualized [Gor92, Gre77, MC78, Rid95]. Induction [BS78, Coc76, Dav84a, EH24, Ger75, Hat11, Hay92, Ruc87, Sch95b, Sol23, SH81, Val15]. Inductive [CS95, Tan15]. Inegrals [Har10]. Inequalities [BR00, Bur83b, Dub92, Gre76, HH88, Lew73, Nel23, Raz15, Sch74b]. Inequality [BB97, Bur86b, FPS13, Gre73, GI18, Hof16, JS75, Kat03, Kun11, KK13c, Mah20, Mer22b, Nel89b, Nel15a, Nel25, Saw17, Sch78a, Sch80a, Sch82c, Sch88a, Sch88b, Sch89b, Sch91b, Siu81, SS79, Sun21, Wat89, Edd85, Goe11, MP88, Mor15, Per94, Rod80, Sch89c, SK89, Ste75, Wag77, Wie85]. Inescapable [Nie21]. Inexpensive [Wai78]. Inference [HSL77, RC99]. Infinite [Cic86, Edw14, FH21, FST23, Gri78, Kem82, Kre97, Mal22b, MM21, Mar90b, MDC80, Nam24, Pau19, Shi78, Tho00, Una09, Ric19a]. Infinitely [MHS92]. Infinitude [Abb24, Meh22, Nor17, Sad17]. Infinitum [Ber80]. Infinity [Fra86a, Sch72c, Sch74d, Smu85, Wag11, Hop18a]. Inflection [Bro01, Eck81, Kou95, OP76]. Influence [Hoa88]. Informal [Gal81a, Wil02b]. Information [Ano70m, Ano72j, Ano72k, Ano73n, Ano73o, Ano73p, Ano74p, Ano74q, Ano74r, Lan72, McC01]. ing [Bon18]. Inge [Rou13]. Inheritance [AN16, Ann98]. Inheritances [Rei73b]. Initial [Cha07, Mat14, Nie21]. Inn [AOW05]. Inner [Cha13, Gra11a, Rou13]. Innovative [Max74]. Inoue [Col13]. Inquisitiveness [Mai70]. Inradii [Ric15b]. Insanity [RY13]. Inscribed [Lon83, Mae19, Mey86]. Inscribed-Triangle [Mae19]. Insects [Boy99]. Inservice [TA74]. Insight [FS79, Gre74a, Hol81b, Hua25]. Insights [Dra00]. Insignia [ST19]. Inspired [BPP15, Du20]. Inspires [Sch23]. Instances [Kna85]. Instant [RY13, Sch83a, Sch85b]. Institute [BE95]. Institutes [Chi82]. Institutional [Ano72e, Ano72f, Ano73g, Ano74h, Ano74i]. Instruction [BJ24, BS72, Hic70, Lia70, Mor73, Not71, SHL86, WJ73, WE73]. Instructional [Coh77b, DW74, Wai74]. Instructor [MC78, Pri73b]. Instructors [Sha18]. Insulation [MP12]. Integer [ANU14, BL22b, Bri91, CDH^{+}14, Fje95, Fos16, FST23, Han82, Han85, Her93, Hon15, KES09, Lut84, Mac92b, Nel16b, Sch73, Sin90, WN96]. Integer-Sided [Lut84]. Integers [CS06, Edg22, Han82, Hir24, HMF20, Hsu82, Kri95, LM20, Lon81, Mer17, Pfa11, Pon07, Sch23, Sch90c, Sno89, Ste76a, Sza05, Wat12, Zia91]. Integrability [Tho12]. Integrable [Pio21]. Integral [BC75, Bra16, CS17, DD10, EH16, Gro19, Gür22, Ins89, KKS90a, Lin78, Lut11, MS89, Mat94, Moo83, Mur80a, Nel18, PL81, Pon17, Ros76, Sch74c, Sin12, Tho23, Win95]. Integrals [Bro74, Bur00, Che09, Fly95, Key94, Lu22, MMOM14, Nic93, Rog07, Ros01, Say95, Ste23, Str79, Str90a, Tei23, Tru19, Und84, Wie83, Wie00, Wit17]. Integrated [Pec70]. Integrating [Gol83a, Hei78]. Integration [Ahn09, AS88, BHS^{+}21, Bre92, Bur86a, Cha99, Cho82, Cra77, DH12b, Gil91, Gri71, Gur19, HLL98, Hol81b, Hor90b, Hud11, Joh89, Ken99a, Kri25, Kro16, MZ94, Mer05, Mey91, Mur80b, Nea93, Nel95, Olm73, SA79, Ste74, Vel23, Ven74, Yan85]. Intellectual [SY87]. Inter [GP76]. Inter-related [GP76]. Interaction [Bra00]. Interactive [Fra86b]. Interchange [Eck11]. Interchanging [Ven74]. Intercity [Pul88]. Interest [Hol78, Rob78, Tan86b, Way80]. Interesting [Ang77, Bur86b, But81, Cra73b, Pen71, Pin75, Vie78]. Interfractile [Pag83]. Intermediate [JW22, JD98, Sha80, Wal11]. International [Pag85, RP85, Tur85b]. Interpolating [Hil92a, Lin74a]. Interpolation [LJ73, LL01, Lu22, McK76b, Sch87b]. Interpretation [Hoe84, Hug89, MM93, Mon82, RS96, Tan86a]. Interpretations [GJJ23, Mus79]. Intersections [FM05]. Interval [AD78, Cas99]. Intervals [SD93, Tan17]. Interview [AR86, AR87, AR88, Alb91, Alb96, Alb97, Alb98, Alb04, Alb05a, Alb05b, Ale79e, Ale79f, Ale80, Ale81, Ano09k, Bar81a, Bro06, Hau08, Jon10, Log80, Mau83a, Pag84b, Pag85, TM01, Tat02]. Interviewed [Ale79d]. Interviews [Ano15a]. Intrinsic [FS15]. Introduce [Bre92, Chr83, Hoa88]. Introducing [BMO^{+}00]. Introduction [Alb07b, Kat18, KS24, MS18, Nat91, Nel24c, Pag09, Sch97b, STW17, Hop18a, Hop18a]. Introductory [Hah88, Hei78, LaT93, Mac78, Rob84b]. Intuition [Mur77, PL82]. Intuitive [Bis80, Coa11, GJJ23, Hol76, Mis23, WC77]. Invalid [FG10]. Invented [Bre02]. Inventor [PS81]. Inverse [AS96, Boa85, CM98a, Cid21, Cra77, DR13, EH16, Ful05, GSS15, Gro93, Han85, Key94, MM93, MP14, Mor91, Nau94, Ovc11, Pap17, Sal85, Tro04]. Inverse-Square [Nau94]. Inverses [GCM18, Han82, Zem71]. Inversion [Hat07, Siu81]. Investigating [Biv97, FC91, Jur94]. Investigation [BH22, Sch83b, TV94]. Investigations [Cla21, DL75, Gor92, Rid95]. Investment [DG80, Thu82]. Invisible [Sch90c]. Invitation [MZ94]. Involutions [WW88, WCP88]. Involved [Gre77, Kip70, Ste77]. Involving [Abe15, Arm74, BHS^{+}21, Bru83, Kat09, Lew73, WCP88, Oku22]. Iroha [Col13]. Irrational [Cho80a, Law21, Mis13, Mor10a]. Irrationalities [Sch74a]. Irrationality [Bum86, Gen91, Jon12b, Lin70]. Irrationals [SL84]. Irregular [CEO88]. Irving [Alb91]. Isaac [Ric87, WN94]. ISBN [Gra18, Gut19, Hop18a, Hop19, Kat18, Ser17]. Islands [HMT11]. Isn’t [Cra74, DW88, Biv86]. Isomorphisms [AM83a]. Isoperimetric [Cha91, Wal95a]. Isosceles [Bas16, Nel16e]. Issue [Dev91b]. Issues [DK01, Tho84]. Iterated [Ano11b]. Iterating [Ni16]. Iteration [But81, MR95]. Iterations [RZ99]. Iterative [EK22, LaT93, Mue07, Ria95]. Itself [Ful05]. Ivan [AA91].

J [Gol13, Jon74, Jon10, Nac15c, Nac16c]. Jaccard [GI18]. Jacobsthal [Nel24c]. Jan [Cur05a]. January [KGSP88]. Jazz [Boa79b]. Jean [Alb96]. Jeeps [Bai09]. Jello [Ber14b]. Jensen [SK89, Sun21]. Jerusalem [Sch07]. Jews [Her10]. Job [Boa81a]. Joe [Hau08]. John [Wag11, Alb07a, Guy82, Ste83]. Johns [Gra18]. Johnson [Kat18, Kar14]. Join [RS96]. Jordan [CJ23, Cos94, HB92]. Josephine [O’H85]. Josephus [MP23]. Journal [BS15, Hoo97, Pag20, Ano70l, Bei20, WW20]. Journals [Zit05]. Journey [Wag99, Hen22, Bra13]. Joy [Hil92b]. Judge [FH95, Han88]. Julia [Rei86]. Julie [RC20]. Junior [BC22, Baj23, BT24, BCT25, CS74, DN70]. Juniors [BM95]. Jürgens [Nat91]. Jury [Fra87]. Just [Hat07, Sch92b, Sch97b]. Justification [Sni77].

Kalamazoo [Tox70]. Kaplansky [Alb91]. Kathleen [Kat18]. Katsuura [Mic14]. Keep [BM88]. Kemeny [Ste83]. KenKen [Nac17e, Wat12, Nac22]. Kennedy [Gor24]. Keno [Smi72]. Kepler [LB94, Ler21, Wal95a]. Kernel [BC75, Tru23]. Key [Cav96, Ker83, LP87]. Kick [Isa96]. Kill [Shu02]. Kind [Jon24, NW10, OB77, Tho12, Wri95]. Kind-Absolute [OB77]. King [Log80]. Klein [BCK^{+}21]. Kleinian [Nac22]. Kline [Ale79e, Ale79f]. Knaves [CM20, Fal86, LR13, Ros14]. Knight [Bea12]. Knightdoku [Nac18e, Nac18a, Nac18b, Nac18c, Nac18d]. Knights [CM20, Fal86, LR13, Ros14]. Knights-and-Knaves [CM20]. Knot [HMN^{+}11, MS23]. Knots [GV24, MW97, Sul96]. Know [CdSS17, Dav85, Dic06, Har96a, MP07, Pen03, PG06, Sch11b, Wey23, Zho17b]. Knowledge [EB99]. Knuth [AS82a, AS82b, FLP18, Knu91]. Koan [Rek16]. Koch [Mat17]. Koch-Like [Mat17]. Königsberg [HW04]. Kuhn [DF96]. Kunusu [Knu91].

L [Kol81, TM01, Her93]. L-U} [Her93]. Lab [CS93]. Labeling [Bre83]. Laboratory [BS72, Dou73, Moo78b, Wil73, WM76, Spa71]. Labyrinths [Sch11a]. Ladder [BJK17, KK08, LM92, SS96]. Ladders [CHJ11, CG14, FGL^{+}18]. Lagrange [Bal22, Bal23, Smi86, Ziz98]. Lagrange-Another [Smi86]. Lagrangian [Con17]. Lake [FN16, GECW13, MS12]. Lakes [CW90]. Lambert [Moo23, PY04, Wan10]. Lambert- [Moo23]. Lamplighter [Sie12]. Lane [Ale89]. Language [Tan17]. Laplace [Arn94, Den07, Gug92, NO92]. Large [BMV06, CG14, Hol23, Jon20a, Kie85, Ros01]. Larger [Niv72]. Largest [Sar23]. Larkin [Rek16]. Lars [Alb98]. Last [Kau20, Mor10a, Mue20, SS81a, Whi85a]. Lasting [GM99]. Late [Ano79f, Ano79g, Ano79h, Ano80i, BBW79]. Late-Breaking [Ano79f, Ano79g, Ano79h, Ano80i, BBW79]. Lateral [Kwo16]. Latin [PV09, KS06]. Lattice [Ben94, Pag87b, Par11, Sas91, Sch90c, Sco87, Zuc16]. Lattices [Ros89]. Laurie [Gol13]. Law [Gem89, Hel73, LS22d, MB82, Pow80, Qui22, Byr11, Gro93, Nel24d]. Laws [Ler21]. Lazy [STW17]. LCM [Law24]. Lead [Guy84]. Leads [Gut19, Mac95, Val15]. League [Jon04, LP16]. Learn [GS90b, Wil72]. Learned [Ken99b, Nun10]. Learning [Doa89, Pen20, Pri89, Wat85, Whi88, RC20]. Least [She77, Ste75, Tee93, VK80]. Lecture [BB76, Hil92b, Mor73, Not71, Pen20, Wei76b]. Lecture-Discussion [Mor73]. Lecture-Program [BB76]. Lecture-Text [Not71]. Lefschetz [Tuc83]. Left [CRS10, Sim02]. Leftist [Ric08]. Legal [AKMMM23]. Legislative [Swe22]. Legs [Kau20]. Lehmann [Rou13]. Leibniz [Cur05a, DC95, Gra94, Grz24, Jon12b, Spi16, Sub04]. Lemma [FG10]. Lender [Mil85]. Length [Bar97, Chi73, Daw76, Een97, Gil98, Jac97, NM96, Rid98, Sim11b]. Lengths [GS11, Kri25]. Leon [Ale92]. Leonardo [Bri75a, Lem16]. Less [Din22]. Lessons [GO12, Hat83, Hol79]. Let [Boa81a, Han88, Her90]. Letter [Ano20g, Fos99, Kly19, Lak24d, Lak24b, (Ed24, Lak24e, Lak24c, Lak25b, Pri72, ST98]. Letters [ACF76, Ano70f, Ano79f, Ano79g, Ano79h, Ano80i, BEL75, BBW79, Coh77a, FBF78, GKB^{+}75, GGF78, LMC^{+}70, Lof78, PSSP76, She76, SP71, TD77, Thu82, TP74, Wen74, Woo72a]. Letting [Len15]. Level [Fre11, Kac17, MP12, San77]. Leventhal [Ric19a]. Leverage [Hoa88]. Lewis [McC02, Rat10, RT07]. L’Hôpital [Bum84, Car78a, Eck11, Hua88, Gil97]. L’Hôspital [Eul93]. l’Hospital [Bou99]. Liability [Fal85]. Liar [Wal04b]. Liberal [Gra99a, Gra99b, Len74, SW76]. License [Hut83]. Lie [Bar22]. Lies [BC94]. Life [Mau85, Nun10]. Lifetimes [AMR13]. Light [AW85, Cib93, HS90]. Lighter [MF94, Thi79a, Thi79b, Thi79c, Thi79d, Thi79e, Thi80a, Thi80b, Thi80c, Thi80d]. Lights [Tor11]. Like [GS90b, Mat17, Pla17a, SS93, Wil72, AN10, GV23, LM06, Nea94, RT07]. Likelihood [JW19]. Limit [Adl96, Aus82, BKSU19, Bet12, Biv86, BGR86, CM90, CC23, Eck11, Ham20, Muk24, Pla14, Sch72a, Sch76, Tha18]. Limiting [BHS^{+}21, Hig23, Suz14]. Limits [AH81, Ash20, CM14b, Daw19, Day13, Gor84, Hos18a, Kau20, LW10, MS17b, Min92, Nan14, Sch97b, Sed80].

Lincoln [Sto04]. Lindelöf [PW21]. Line [Ayo87, Biv86, BS04, But03, DL75, ES23b, EH24, Gaz84, Hec03, Lew84, MS18, PG79, Sas81, SS82, Ste75, Tro89, Tur81, VK80, FG19]. Line-to-Line [DL75]. Linear [AR86, Ano93i, BC75, Bou23, BM93, CJLP92, CJLP93, Car93, Che18, Col13, DK01, Din13, Dob85, DL75, Fra92, Fri21, Gra95, GL00, Gre85, Gug92, GEJ95, Gug96, HM23, Hsi89, Isa95, Kal01, Kal02, LaT93, Lax76, Lu22, Mac92b, MS23, Mue20, Mun04, Nel77, Oli73, Por93, SS77, Sch99a, Seg87, Sem74, Sha25, SB97, Sob97, Tuc93, UK96, Wal96, CJLP93]. Linearizing [AAC04]. Linearly [CL85]. Lines [Alf15, BC16b, CW23, FM05, NM96, SS00, SH85, Smi00, Ste01, Tro04, Wil03]. Link [Liu01]. Linkage [CG24]. Linking [BCK^{+}21]. Links [BCK^{+}21]. Lipman [AR87]. Lissajous [Cas03, MW97]. List [RC99]. Literacy [Hec75]. Literature [Boa79b, CS74]. Little [Bon18]. Live [Gei84]. Lives [BPP15, Tre92]. Living [Alb81a]. Lizard [Bir15]. Lobb [Kos09]. Local [You93]. Locating [BS14b, Can89]. Location [Buc87, PDR21]. Locations [CW23, EK92]. Loci [AM83b, Mar92a, Nor88]. Locks [Cav96]. Locus [AM10, Mau83b]. Log [Jor77]. Logarithm [BD78, Blå16b, Cus83, Fra24, Kal01, Kir88, MW83, Ric22]. Logarithmic [Ni16]. Logarithms [AM01, Bur99, Chr83, Fre86, Pie77, Smo24]. Logic [Lin19, Mar91, Nee13, Set74, She75]. Logical [BR83b, Set74, Sch11a]. Logistic [Bra01a, Cav93, Wil02a]. Lognormal [SM00]. LOGO [FC91, Wat85]. Long [Bas16, BM98a, Bur99, Sch04, Som00]. Longest [Sch90a]. Look [AS02, Bar22, BB01, BC18b, Cha91, CK14, Egg20, Fal81, Fra88, Fri78, HR24, HMNW94, Lan72, Lar85, Lon83, Mal17, Sch84b, Sim95, Tre92, WS73]. Looking [Bri88, Car98, Che22, HLL98, Hsu80, JD90, Mos84, NS17]. Loopholes [Ben76a, BL75, Zir76]. Loose [Lie97]. Lord [FE01]. Lore [Eck83]. Lose [Jon04]. Losers [Min03]. Loss [LN99]. Losses [KM15]. Lost [Ols13, Gut19]. Lottery [Ber94, Gar24, Ilt00, JSW22, Kai01]. Love [Ped98b]. Low [VK00]. Lower [Coc76, Spa71]. Lucas [BQ99, Pon17]. Luckiest [Gar24]. Luddhar [Lut80]. Lunda [Ger00]. Lunda-designs [Ger00]. Lunes [Dra23]. Luoshu [TT12].

M [Gol13, Kat18, Log80, Sch11a, Sri10, Nac15d, Nac16d]. M&m [SS05]. MAA [Ano15a, Kat18, Ano70l, Ano15h, Has71b, Lan72, Nac15a, Nac15b, Nac15c, Nac15d, Nac16a, Nac16b, Nac16c, Nac16d, Nac16e]. Mac [Ale89]. Machin [Nel21]. Machine [Nel77, Ruc82, SA99, Col03]. Machine-Oriented [Nel77]. Machines [SS22]. Maclaurin [Eul93, Hed00]. Mad [BPS12, Sch95b]. Madcap [Ric19a]. Madden [Kat18]. Made [Bum86, Jos16, Man74, NL84, Tra96, Bos95]. Magic [AM83a, Bea12, BB14, Bos00, Bro01, Gar00b, Ger00, KH08, Lau20a, Nag21, RR89, Val15, Wag94, WR89]. Magician [Bon08]. Magicians [BB14]. Magus [Guy82]. Major [Boa71a]. Majors [Ken00]. Make [Bol83, Dal19]. Makers [Smi81]. Makes [BG15, Eri22, Hat83]. Making [Bur95, Gra11b, JY98, SY87, WS05]. Man [CL83, Joy14, Man74, Ric87, Sim07]. Man-Made [Man74]. Managing [CHS86]. Mancala [FLS14, TTW^{+}13]. Mandelbrot [Nat91, Bar84, Bri88, Dev91a, Ewi95]. Manga [Col13]. Manipulate [Fos12]. Manuscript [Ano70k]. Many [DW19, JF88, Lon81, MHS92]. Map [Fee01, Mit81, Pul88, Reb79]. Mapping [Bri78]. Mappings [MS91]. Maps [Loe14, MT23, Ska91]. Marbles [DHR10]. Marching [PR18]. Margin [Eis97]. Marginal [AFL^{+}21]. Maria [Gra18]. Marked [KR91]. Markets [Din13]. Markov [BR24, CH17, Jon04, LM92, Rid90]. Marriage [BCD^{+}24]. Marry [Tru23]. Martin [Alb05a, Alb05b, AC00, Ano09k, Bar79, Kho12, MR79]. Mary [TM01, AR88, Hil92b]. Mascheroni [Bil18, Sha15]. Mass [GW18, Hu21, SJ11]. Massimo [Gra18]. Master [Cav96]. Mastering [FLP18]. Mastermind [FLP18]. Mastery [Hav78, Hec75]. Match [BP12]. Matches [HS86]. Matchstick [Edg16]. Materials [Coh77b, Zah72]. Math [Car70, Cat19, FSW04, Gut19, Hil80a, Hil80b, Len74, PM74, Smi14, ST98, Col13, Nun10]. Math-Ed [Len74]. Mathematical [Alb05b, AW90, And96, Arg84, BPP15, BC22, Baj23, BT24, BCT25, BN03, BF13, BS78, Coc76, Dav85, Dru92, Fra88, FH95, Ger75, GO12, GG84, GM99, Guy82, Hec75, Hon79e, Hon79a, Hon79b, Hon79c, Hon79d, Hon80a, Hon80b, Hon80c, Hon81a, Hon81b, Hon81c, Hon82a, Hon82b, Hon83a, Hon83b, Hon95, Jon11, Kac17, KS03, KS96, Ler21, Mai70, Mot75, Nie91, Pag85, Pag86, Ren81, RP85, Ruc87, Sch83b, Ser17, Smi81, SH81, Wag94, Alb05a]. Mathematician [Alb94, AC00, Atk05, Bai88, CL90, Gra18]. Mathematicians [Cha77, God88, Har96a, KGSP88, Mor04, PM88, Wil90]. Mathematics [Alb81b, Alb07a, Ald79a, AC00, All03, Ano70l, Ano71f, Beh73, Bei20, Boa71a, Boa79b, Bra13, Bru83, BEN78, BS72, BM77, CS74, Car75, Chi82, Cog99, CC95, CCH76, CDK^{+}90, CS86b, CS87, CS88, Cuo00, DP71, Dav80, Dav82d, DW98, Den13, DK08, Dou73, Dud97, DW74, ES85, EW70, Ell84, Fei76, FIRW14, Fis75, FA71, FMA72, FM77, Fri78, Gal91, Gar00a, GK77, Gol11a, Gol83a, GS90b, Gra99a, Gra99b, Gre77, Had13, Ham84, Han70, Has71b, Hat83, Hau08, HA70, Her90, Her10, Hic70, Hil92b, Hol79, Hoo97, Ill17, Joh76, Jon71, Ken00, Kor85, Lai70, Lat83, LS94, LeC86, Led77, Lei83, Len76, Lin75, Lin21, Lop93, Luc84, Mac78, Man74, Man75, Man78, Mat96, May13, Mey07, Mil74, Mit72, Mit74, Pag20, Pri89, Ral84, Ric87]. Mathematics [RM71, Rob84b, San06, San02, SS93, Shu02, Slo76, Sne88, Spa71, SW76, Ste73, Ste76b, Sto07, Tal84, Tee07, Tee12, Tox70, Tre92, TS83, Tuc93, Wag84, Wag11, WE73, Wai74, WW20, Whi88, Wil73, WM76, Wil82b, Wil72, Woo70, You70, Zah72, Zas70, Zee13, Zit05, Ano70e]. Mathologist [Alb82]. Matrices [BM98b, CD93, Han82, Han85, Hoo00, Hua25, Ins93, JK98, Kha03, Kra91, LP14, MS23, Str93, Str77, Swa73, TTW^{+}13, VB86, VK00, Zem71, Zho17a]. Matriculation [Dun23]. Matrix [Alp89, BS02, Bur98, CM98a, Cid21, Coa11, DR13, Fla86, FH11, GCM18, Gol94, Han90, HM95, HM23, Kal96b, LL01, LW18, Lot23, Mal22b, Moo20, Pap17, TB20, Tho00, Wan21, ZF97, CG19]. Matroids [LN14]. Maturity [Rob78]. Maupertuis [Sim11b]. Maurer [Dan85]. Maverick [Alb82]. Mawr [Dun23]. Max [BAW87, Bai91, HP14, Ziz98]. Max-Min [BAW87, Bai91, Ziz98]. Maxim [Smi81]. Maxim-Makers [Smi81]. Maxima [Ano92f, CL15, Cha92, LM90, Sch74b]. Maximal [Kal06, Lat83, Tan17]. Maximization [CC19]. Maximize [Lan74, Mey86]. Maximizing [DS99, Eck07, Sim09a]. Maximum [Car72, Kal84b, Pla17c, Ric15b, SS83a]. Maybe [Moo86, TYM88]. Mazzotti [Gra18]. McCarthy [Str85]. McNugget [CGSO21]. MD [Gra18]. Mean [Boa81b, Bur83b, Edd85, Fal81, Har86, Hoe84, Kac17, Kat03, Mer17, MP88, Mur84, Mye77, Nel89b, Nel15a, PM82b, Pla14, Pla16, Say95, Sch78a, Sch82c, Sch85c, Sch88b, Sil89, SS81b, Smi86, Sto04, Pow80]. Mean-Harmonic [Nel89b]. Mean-Value [Boa81b]. Meaning [DC95]. Means [Dub92, Hub94, Pla16, Sch88a, Zuc16]. Measles [Roz07]. Measure [Kow08]. Measures [Che05, Nie94, PM82a, PM83]. Measuring [TPB99]. Mechanical [CW09, Osl06]. Mechanisms [vD09]. Media [Ano86g, Ano88h, Ano88i, Ano88j, Ano89m, Ano93j, Ano93k, Ano94l, Ano94m, Ano95o, Ano96o, Ano96p, Ano97o, Ano97p, Ano97q, Ano98k, Ano99k, Ano99l, Ano99m, Ano99n, Ano99o, Ano00g, Ano00h, Ano00i, Ano00j, Ano00k, Ano01p, Ano01q, Ano01r, Ano01s, Ano01t, Ano02p, Ano02q, Ano02r, Ano02s, Ano02t, Ano03p, Ano03q, Ano03r, Ano03s, Ano03t, Ano04p, Ano04q, Ano04r, Ano04s, Ano04t, Ano05p, Ano05q, Ano05r, Ano05s, Ano05t, Ano06p, Ano06q, Ano06r, Ano06s, Ano06t, Ano07o, Ano07p, Ano07q, Ano07r, Ano07s, Ano08r, Ano08s, Ano08t, Ano08u, Ano08v, Ano09l, Ano09m, Ano09n, Ano09o, Ano09p, Ano10g, Ano10h, Ano10i, Ano10j, Ano10k, Ano11i, Ano11j, Ano11k, Ano11l, Ano11m, Ano12h, Ano12i, Ano12j, Ano12k, Ano12l, Ano13g, Ano13h]. Media [Ano13i, Ano13j, Ano13k, Ano14j, Ano14k, Ano14l, Ano14m, Ano14n, Ano15i, Ano15j, Ano15k, Ano15l, Ano15m, Ano16g, Ano16h, Ano16i, Ano16j, Ano16k, Ano17g, Ano17h, Ano17i, Ano17j, Ano17k, Ano17l, Ano17m, Ano18g, Ano18h, Ano19h, Ano19i, Ano22e, Ask92, BB88, Bar91, BS23, BBS24a, BBS24b, BBS25, Ber98, Ber87, Bix95, BC85, Bus96, Cha84, CB85, Dav86, Dev89, Dev95, Dos96, Dud94, Dun97b, ES23a, EBS24a, EBS24b, EBS24c, FA85, Fla90a, Fla90b, Fla90c, Fla90d, Fla90e, Fle93, Her98, Hil84a, Hil96, Hog84, Hon87, Jon85, Kap85, Kap92, Kat92, Kid93, Lan93a, LS19a, LS19b, LS19c, LS20a, LS20b, LS20c, LS20d, LS20e, LS21a, LS21b, Lei21, LS21c, LS21d, LE22, LS22a, LS22b, LS22c, TLSES23, LS23a, LS23b, Mau84, McL87, MS87, NN91, Niv86, Noe88, Ote98]. Media [Pag87a, PLS18, Ped98a, Ren89, Ros92, Sau94, Sca91, Sch89a, SSG89, Sch86e, Sel91, Sen95, Str97a, Swe92, Swe95, Tat91, Tat94, Wat84, Whe86]. Median [Doa89, Fal81, VU21]. Medians [Sch95b]. Medical [Ros12, Whi85a]. Medieval [Dor05]. Meeting [Lew95, Coo70]. Meetings [Ano81a, Ano81b]. Meets [Bil18, CK12, Log79]. MegaMenger [Bic17]. Memoir [Her10]. Memoriam [Ano81k]. Memories [Kir97]. Memoryless [AMR13]. Merca [Zac18]. Mercator [NM96]. Merrill [Ald81]. Mersenne [Asa84]. Meta [Cuo00]. Meta-Problems [Cuo00]. Metaphors [Hof82]. Method [Bea17, BQ08, Ben17, Bom82, Bra19, Bro07, Ett74, GW18, Hsi89, KM13, LC23, Mil74, Mis23, Mor73, MM85, PH77, Rat81, Ria95, RT07, Smi87, Spo23a, Tho84, Ven76, Wei76b, LS12, Bor11, Chr96, FN07, FL75, Het97, Lut80, Pri99, Wal95b]. Methodology [Rob84a]. Methods [Fly95, LaT93, Lan77]. Metric [Dom11, Jon24, Nef83, Rat10, Rok14, Wal95a]. Metropolitan [BH22]. Michigan [JW15]. Microcomputer [Hsi89]. Microcomputers [DW87]. Midpoint [Buc85b]. Midsummer [HMN^{+}11]. Midy [Gin04]. Mifflin [Ric19a]. Might [BW13]. Migration [JMW24]. Mile [AAC04, BMV14]. Milk [BHB04]. Millennium [Ols13]. Milwaukee [RM71]. Min [BAW87, Bai91, HP14, Ziz98]. Min-Max [HP14]. Mind [BB08a, Hos18b, MR79, Wal04a]. Mind-Reading [Hos18b]. Mineola [Hop19]. Mini [Bod74, Car70]. Mini-Calculus [Bod74]. Mini-Math [Car70]. Minima [Ano92f, CL15, Cha92, LM90, Sch74b]. Minimal [Chu09, Dru95, HLL98, McC98b, Nee13, Par95, Pra17]. Minimization [PH83]. Minimizing [LN99, Rou12]. Minimum [AT00, Arm74, Gor03, Han90, Jac97, Kal84b, Kri25]. Minitab [Gre85]. Minkowski [Pfi88]. Minority [Tre92]. Minute [BMV14, BF13, Jon11]. Miscellanea [Ano99p, Ano99q, Ano00l, Ano00m, Ano01u, Ano01v, Ano01w, Ano02u, Ano02v, Ano02w, Ano03u, Ano03v, Cup00]. Misconceptions [Mar92b]. Miscopied [PV05]. Misguided [Smi81]. Misleading [Rob84b]. Miss [MST21]. Missing [BN03, Kai01, Liu01]. Mistake [Gri74, Kho12]. Mistaken [KS06]. Mistakes [DP80]. Misunderstood [LP94]. Mixed [AM02, Pro76]. Mobile [Far02]. Möbius [You24, Ano11b]. Mod [ALS12, Ryd96, Sch00, Sib12]. Model [BR24, BO95, DW74, Heg87, KKS91, LR20, Lon76, Mic12, SE96, Seg87, TA74, Wan10]. Modeling [BW19, BQ13, Bur98, Fis75, Gar00a, KS03, KK13b, Led16, Lor00, Pag86, San91, WM13]. Models [App01, BHB04, DY04, HM05, Kac17, Lau20b, Led16, Mas84, Roz07]. Modern [BS15, Mer18, Sny24, Sri10]. Modifying [CJ23]. Modularizing [SW76]. Modulo [ANU14, Daw94, KES09, Lee24, LD15, Nel13, Nel17b, Nel24a, SB84]. Modulus [LH89]. Moiré [Cul90]. Moise [Boo96]. Molehills [Bur95]. Moments [Gra90]. Money [Dan98]. Monkey [Mor04]. Monoid [CGSO21]. Monotone [Che13]. Monotonicity [Li13a, Li13b]. Monster [BK14, KX15]. Montana [Rei73a]. Monte [Cea84, CD93, Nea93]. Month [Fla87a, Fla87b, Fla87c, Fla87d, FF87, Fla88a, FT88, Fla88b, Mat88]. Moon [Bra01c]. Moone [Sim07]. Moore [Ano88a]. Morgan [Gro16, Sim11a]. Morphs [Kin13]. Morrie [Nel24d]. Morris [Ale79e, Ale79f]. Morsel [Dun20]. Mortality [Byr11]. Most [Din22, Fak99, Ste23]. Mother [Sri10]. Motion [BL78, PY04, dA95]. motions [GG22b]. Motivate [Doa89, MDC80]. Motivating [Seg82a]. Motivation [Rat81]. Motivational [McG90]. Mountains [Bur95]. Move [Hop18b]. Moving [Mor02]. Much [Dan98, Eis98a]. Multi [IB13]. Multi-Peg [IB13]. Multiple [Biv97, BL22a, Eis98b, EW70, Hat07, HM05, Kri25]. Multiple-Choice [Eis98b, EW70]. Multiples [Can89, Nel22b]. Multiplication [Kun80b, Mal22b, Pet70, WR89]. Multiplicity [HKM22]. Multipliers [Fla20, Ziz98]. Multiply [RB21]. Multiplying [BC16a, Suz99]. Multivariable [CS93, Fra86b, Hof16, LW10, NM96, Rid95]. Multivariate [MP23]. Mundane [Din22]. Music [DK08, Haa91]. Musical [McC98a, Shi88]. Must [Car93, Eck81, HSZ05]. Mutilated [Gar09]. Mutually [PV09]. My [Cur10, Joh17]. Mystic [Aug12]. Myth [Ric87].

n [Car87, JD90, Jon74, Luc98, Sch99b]. Naccis [BK14]. Name [Alb81a, Par16]. National [Jon04, LP16, Bra90]. Natural [Bur99, Chr83, Fra24, GKR19, Gre20, KC84, Ken99c, Kro83, Law24, Ric22, Sad17, ST19, Smo24, SS21]. Nature [AH81]. Navigating [Sch07]. NCTM [Coo70]. Nearly [Gin04, Nel16c]. Nearness [PM82a, PM83]. Necessarily [Eck81]. Necessary [BMRS13, Dud97]. Neck [CS86a]. Necklace [BM88]. Need [Ald88, BI10, Dan98, GK88, Hil84b, Mas84, Ren82, Smo85]. Needle [Arn94, Cra73a, Kac19]. Needs [Boa81b, Dou85, Hue92]. Negative [Sch73, SR16, WD90]. Neglected [BD78, Cha92, Jan80, Len74, SY78]. Neighbors [Elr12, Zas70]. Neile [Lea16b]. Neither [Ash97]. Nelsen [Edg18]. Nested [KP85, MP10, Nel17a]. Nests [KS03]. Networks [GS96]. Neumann [Gol83b]. Neurons [Kir97]. Neusis [LR24]. Neutrals [Ros14]. Nevertheless [Boa81a]. Newly [For17]. News [Ano72g, Ano73h, Ano73i, Ano73j, Ano74j, Ano74k, Ano74l, Ano74m, Ano75g, Ano75f, Ano75h, Ano76g, Ano77k, Ano77l, Ano77m, Ano77n, Ano77o, Ano78k, Ano78l, Ano78m, Ano78n, Ano79f, Ano79g, Ano79h, Ano80i, Ano83j, Ano83k, Bal82a, Bal82b, Bot78, BBW79, Chi75, Chi76, Dav82a, Dav82b, Dav82c, Kin79a, Kin79b, Kin79c, Kin79d, Kin79e, Kin80a, Kin80d, Kin80e, Kin80b, Kin80c, Kin81a, Kin81b, Kin81c, Lin76a, Lin76b]. Newton [LS12, Wil24b, Bro04, Chr96, CMS20, CSTM10, Cur05a, FN07, FH94, Het97, McG18, Nau94, Pri99, Ric87, Sim11b, Smi87, Wag77, Wal95b, WN94, WK94, Wil94]. NFL [IB22, Joh93b]. Nigeria [Zas70]. Nilpotent [Loe14]. Nim [Rou14, FLS14]. Nine [BCR^{+}11, CD93, GP02]. Nine-Card [CD93]. Nines [Lon81]. Nineteenth [AC00]. Ninetieth [Ale79d]. Niven [AA91, KC84]. No [Bea06, Cal00, GG82, Ilt00, Tan88, Hau05, Knu91]. Nod [Lo 97]. Noether [Sri10]. Non [ACJ^{+}06, Bea17, BG04, SC03]. Non-Circularity [BG04]. Non-Convex [ACJ^{+}06]. Non-Proof [Bea17]. Non-Smooth [ACJ^{+}06]. Nonagon [LR24]. Nonagonal [Cag20a, Cag24a]. Nonconsecutive [Mat14]. None [Fra86a]. Nonfield [Pom77]. Nonlab [Hor74]. Nonlecture [Hor74]. Nonlinear [MC20]. Nonlogarithmic [Bad82]. Nonnegative [Mel22]. Nonperiodic [JM97]. Nonprogrammed [Hor74]. Nonstandard [Gor86, Kun88]. Nontransitive [Gri17, SS17]. Nontrivial [BC17]. nor [Ash97]. Normal [Eis17, PS15, SS00, Smi00, Stu95, Tha18, TL00]. Normalizers [Joh00]. Normals [BT07, Ros14]. Norms [DLM04]. North [Joh72a]. Notation [Gar93, Hol76]. Notational [Hil89]. Note [Ben76a, Ben76b, Ber75, Bra94, BS78, Car78a, Een97, Eul86, Eul97, Fra78, Gri78, Kro82, Lin74b, Lin78, Loe21, McC80c, Mur80c, Pin72, Sch74d, Seg82b, SS79, Tho23, Ton15, Wes94, Wri89, Yan85, Zen96]. Notebook [Bar22]. Notes [Ano72g, Ano73h, Ano73i, Ano73j, Ano74j, Ano74k, Ano15n, Ano16l, Ano17n, Ano18i, Lan80, Shi88]. Nothing [Blå16b, Sta98]. Notices [Ano74l, Ano74m, Ano75g, Ano75f, Ano75h, Chi75, Chi76]. Notion [DC95]. Novel [Eck98, Grz24, KB99b, LC23]. Null [Kal84a]. Number [Blå16b, BB08b, Cag23, Cho80a, Edg21, FG07, Gre74b, GEJ95, Hen94, IB22, JB92, Jep79, Joh13, Jon12a, JSW14, Joy14, Kai01, KM89, Lan77, Lau25, Mac92b, Mar90b, Mar09, McC02, Nel15d, PS22, Pau85, Pla17a, Reb89, Rib94, Rok14, Sch95d, Sob95, Sta88, SY93, Wie85, YD07]. Number-Guessing [McC02]. Number-Pad [FG07]. Number-Theoretic [JB92]. Numberland [Wil02b]. Numbers [Asa84, AG87, Ben17, BCK^{+}21, Ber94, Bon14, Bro99a, BDDY13, Cag15, Cag20a, Cag20b, Cag23, Cag24b, Cag24a, Cag24c, Car78b, Coo82, DPV21, Eck83, Edg20, Fra88, Fra87, Gal91, HKM22, Hay92, Hir24, Hon15, Hut83, IMW12, Kal84b, KC84, KC87, Kro83, Law24, Lea16a, Lea73, Lin70, Luc83, LM24, MM73, McA83, Nel15b, Nel15c, Nel16d, Nel16e, Nel17b, Nel19, Nel22a, Nel24c, Nel24b, Oma09, Pon17, Pri18, Rib97, Ric08, Ric19b, Ric19c, Ryd96, Sad17, Sch72b, Shi10, Smo24, Spo16, Spo17, SS21, THMM24, Wat12, Wei76a, Wie84, Wil82a, Wri95, Yiu12]. Numerical [Alb07a, And81, AS88, Bur86a, Fly95, Gue01, Ken99a, Mer05, TT12]. Numerically [Wil98b]. NY [Hop19].

O [Nat91]. Oakland [RC20]. Objects [ST20]. Oblique [Dun77]. Observation [Fri21]. Observations [Jan80]. Obstacle [BPR15, BE03]. Obstructed [CHS87]. Obtaining [And81, KC90]. Obvious [Wal11]. Occupational [Mit72]. Octoplex [Mic11]. Odd [Dav84a, Her12, Sch23]. Oddity [EF94]. Oddly [FN09]. Odds [Bra80, Gar01, PV16]. ODEs [Fri21]. Off [BD96a, BD96b, Gar24, MS16, Roz94]. Often [Ber94]. Oil [Nie94]. Old [Bon08, BC18b, Cha91, Cib93, Fre03, Gan17, Hal13, Hat04, Sim95, WS73]. Oldest [Boa19]. Olympiad [BC22, Baj23, BT24, BCT25, KS96, Pag85, RP85]. Olympiads [Tur85b]. On-Line [Tro89]. Once [MZ24]. One [Alb81b, BCR^{+}11, Boy10, CM15, CM20, Dou85, Dr24b, Eid93, FPS13, Fre11, HMT11, Kir97, Lut84, Mat91, Nat91, PP84, RS23a, Ros77, Sne88, Swe22, Tat02, Tol22, Wal87, Wol71, Zem71, Roj16]. One-dimensional [HMT11]. One-Factorizations [Wal87]. One-Round [CM15]. One-Sided [Wol71, Zem71]. One-Way [RS23a]. Ones [Wat89]. Online [All03]. Only [Han82, Ric19a, Sne88]. Open [Den13, Gia73, Sin76]. Operation [Joh00]. Operational [WC77]. Operations [Gre75, KKS90b]. Operator [BY97, Sob97, UK96]. Operators [Cho82, SS77]. Opportunities [AH81]. Optics [MT97]. Optimal [Bet12, Bos00, CC19, EK92, FG19, Gal81b, MP12, MS18, RS21, SH12].

Optimization [BL22b, CDH^{+}14, EF94, Eri22, PG06, Suz14, Tea06]. Optimizing [Nog17]. Orbit [Dev91a, Tee17, Wil94]. Orbits [LB94, Nau94, TW98, Won00]. Order [Bén05, Dub92, DE77, Fri21, Ger00, Gil97, Gre11, Han90, HLL98, Joh89, Lon84, MC20, Mat14, RZ99, SS77, Str04, Ven74]. Ordered [GV23, Olm73, Tro89, Wei76a]. Ordering [Coc76, Pin74a]. Orders [Sch72c, Sch74d]. Ordinary [HS00, SS77]. Orientability [Wol71]. Orientation [HK19, IMW12]. Oriented [Nel77, Pet70, BE95]. Origin [Ban15, Her02, HS96]. Original [DC95]. Origins [Mey07]. Orlin [Ric19a]. Orthogonal [Ban04, Joh93a, Kra91, PV09, Smi80]. Oscillations [HM05, Pro95]. Oscillator [Pro95]. Oscillators [Fay97]. Other [Bri75a, FS13, MW81, Ros12, ST98, Sch73, Sch99b, Tou96]. Ought [Ren81]. Our [BL22a, Coc76]. Outdated [Rob84b]. Over-represented [Her10]. Overlapping [Jus89]. Overtime [Jon04, LP16]. Own [CCH^{+}21]. Owner [Yin87]. Oxford [Hop18a].

P [Gol13, Hil92b]. Pace [Kha13]. Packing [WP90]. Packings [MS18]. Pad [FG07]. Painting [Sch06]. Pair [BD96a, BD96b, Cal95, HKP15]. Pairing [Sul16]. Pairs [GV23, Hig23]. Pan [Bur72]. Paper [Bir15, Sza05, TPB99]. Paper-and-Pencil [Sza05]. paperback [Gra18, Hop18a, Hop19, Kat18]. Paperclip [Hoe09]. Parabola [AM10, BT07, BH93, DC95, Hol82, Osl06, Sas91, Sny24, Spo22, Ste80, Ste01, Woo00]. Parabolas [Cow76, JHG11]. Parabolic [FM00, Mas11, RS85]. Paradigm [She78]. Paradox [BJK17, Fje95, MW18, Nel79, PS81, SS96, Wal04b, Zam77, Cea81, Cea84, JW19, KL11, Kna85, MZ24, Min03, Ric24, Tan86a]. Paradoxical [Sto07]. Parallel [BS04, Knu18, Kro82, Ste80]. Parallelism [Liu11]. Parameter [Nor88, Wie01]. Parameter-generated [Nor88]. Parameters [Kit19]. Parametric [Boy10, Che96, Far15, Pen19, SS94]. Parametrically [Hil90]. Parametrization [Tir86]. Parametrizing [Wil98b]. Parenthesization [Kos09]. Parents [Dan98]. Paris [Alb81a]. Parity [BDCH85, LM24, Mer22a]. Parking [CCH^{+}21]. Parrondo [Min03]. Part [AS82a, AS82b, Ale79e, Ale79f, BD96b, Hil80b, HP99b, Nat91, PM82a, PM83, vD09, Nym83, Alb05a, Alb05b, Bal23, BD96a, HP99a, Wie75a, Wie75b]. Partial [AM02, Bur84, Cra12, Gur19, Hen91, Hua91, Hub81, Jos83, McM03, Nym83, Pro76, Sch83c, Ser15, Wie86, YD07]. Partially [Cal00, Tro89]. Participants [Bit70]. Participation [DP71]. Particular [Kow24, Kun72, Kun73]. Partisan [Suz14]. Partition [Ald79b, Loe14, LM24, Mer22a]. Partitioning [Sad17, Sta88, SY12, Tro89]. Partitions [BPS12, HKP15, Hof74, Mer22a, OHC07]. Parts [Bur86a, DH12b, Ehr24, FN92, Gil91, Hor90b, Kro16, Mur80b, SA79, Yan85]. Party [BPS12]. Pascal [Aug12, Can89, Den07, HP99a, HP99b, Hyd88, Jon83, LM20, Lon84, Luc83, Nea94, Nef82, Osl03, Pla17a, Pla17b, Put86, Ros77, Smi73, Wal00a, WK73, Wri89, Wri95]. Pascal-Like [Pla17a, Nea94]. Pass [Lon76]. Passing [Eis98b, Joh93b]. Passive [WWW83]. Past [San02, Wil24b]. Path [BPR15, Mar98a, Ren85, SA99, Ter96]. Paths [FG19, GRV21, Jac97, LR24, Luc83, Mar90b, NR16, Zuc16]. Patrick [Kat18]. Pattern [Dr24a, Ols89]. Patterns [Cla11, DL73, Gol94, Kal96a, Sch13, smbH02]. Paul [Alb82, Alb04, Ale81]. Pawns [Elr12]. Pay [Eis98a]. Payments [JW15]. PDEs [MC20]. Peak [Lon76]. Peaks [BL22a]. Peano [CM17]. Pedagogically [AB16]. Pedestal [Ami18]. Peg [IB13]. Peitgen [Nat91]. Pell [Pla17a]. Pen [Nas23, Sch97a]. Pencil [Sza05]. Pencils [Cap17]. Pendulum [Rob97]. Penetrating [Bai09]. Penner [Ano81k]. Penny [Cra73b, Cra74]. Penrose [CS95]. Pentagon [Nel16g]. Pentagonal [Cag23, Cag24b, Cag24c]. Pentagons [Wil24a]. Pentagram [Cho80a]. People [Zas70]. Peremptory [Fra87]. Perfect [Asa84, Fra75, Nel16d, Nel17b, She77]. Perfecting [Sim02]. Perfectly [Hot02]. Performance [MS19]. Perimeter [BC16b, Coo24, Par84, Pfi88, Pla17c, Ton97, WCP88]. Periodic [BY97, CP89, Fay97, Fer98, HM05, OT72]. Periods [CP89]. Permudoku [Nac14]. Permutable [Wil24a]. Permutahedron [Cri11]. Permutation [Wil93]. Permutations [HP14, LM92, TT21, Vod97]. Permuted [Kin74]. Perplex [Spo17]. Perplexes [PL09]. Perplexing [Die23, Ric10]. Persistence [Sch83b]. Person [BB11, Str85, Tam87]. Personal [Lan80]. Perspective [BM98b, Fra12b, JW19, LRSZ14, McK79, Moo89]. Perspectives [Man84]. Petal [Gra90]. Peters [Sch11a, Wag11]. Petersburg [Cea81, Cea84, KL11, MZ24, Tre19]. Peterson [Gol13]. PGA [MS19]. PhD [Hen22]. Phenomenal [ABCS01]. Phenomenon [JMW24]. Philip [Rek16]. Philosophy [Gol11a, Her90]. Phoebe [Bro05]. Phorecaster [ABCS01]. Photographs [Tee17]. Phyllotaxy [Car78b]. Physical [FG92]. Physicist [Alb94, FH94]. Physics [Boa71a, DK08, Gut19, Pec70]. Physics-Calculus [Pec70]. Pi [AT00, DLM04, Grz24, Jon12b, Mar24, Nam24, Nel25, Woo25]. Picard [PW21]. Picking [Mae19]. Picture [Ale83b, FH00, Ren82, Tho98b]. Pictures [JK98, Ken91, Ken99a, Nel79]. Picturing [Bra85, Cic86, Sch95a]. Pie [BB11]. Pie-Cutting [BB11]. Piece [Blå16b, Kri25, TT12]. Pieces [JM97]. Piecewise [BMO^{+}00]. Pigeonhole [EKP15, Reb79]. Pile [Moo20]. Pills [DW74]. Pineapple [Car78b]. Pinpoint [Nat19]. Pioneer [Alb07a, Mey06, Ste83]. Pisa [Dea71]. Pitfalls [DW88]. Pivot [Roz88]. Pizza [WW95]. Placement [AMM15, ES23b, EH24]. Plaited [Ped73]. Plan [Col71]. Planar [Por93, SS94]. Plane [Alb17, Boo96, BL78, BG04, Cam19, Can86, Che96, CS95, Gaz92, GPV24, Hof74, Ins93, KR11, Mie83, Moo91, San24, Sob95, VU21, WP90, GG22b]. Planes [Bro01, Kra03]. Planetary [TW98]. Planets [BBB^{+}13]. Planimeter [Egg20, Lei07]. Plants [CCKL25]. Platonic [Gro05b, Ped73, Sch12]. Play [Bra00, KK12, OAHS21]. Players [Gri17]. Playing [Gar00a, Sim06]. Playoffs [Joh17]. Plays [KX15]. Plea [Sch21b]. Pleasant [HF93]. Please [RS00, Ska96]. Pleasing [Fer98]. Plimpton [Sim19]. Plows [Lew95]. Plug [Bos95]. Plutarch [Lan95]. Poem [Alb17, CSTM10]. Poems [Emm09, Hal10]. Poi [Far15]. Poincaré [Hec03, Jur94]. Point [But81, CW23, Dea71, Eck81, Gaz84, Gaz92, Gla12, Mac95, MR95, PG79, Par83, Poo12, Rou12, Sas81, Sch94b, Smi91, Smi11, Tru83, Tur81, BC18a]. Point-and-Line [Tur81]. Point-Slope [Mac95]. Points [AL15, BPR15, BY97, Bro01, DRW96, Dun97a, Fak99, GS10, HS90, HGGB08, HS96, Jep86, Kou95, Nor88, OP76, Sch90c, TT21, VU21, Zho12b]. Poison [Shu02]. Poisson [BW19]. Pol [D’A23]. Polar [Boy15, Egg20, Ess82, Gor93, Lew86, McC90, Pri95]. Polishing [Lee98]. Polite [Alb81b]. Poll [Eis97]. Pollak [AT84]. PóLya [Ano08q, Ale79d, Ano06o, Ano08p, Ano11h, Ano12g, Ano14h, Ano14i, Ano15g, Ano16f, Ano17f, Ano18f, Ano19g, Ano20f, Ano22d, Ale83b, Ano21h, Ano24g]. Polygon [DeT98, JV15, PS22, smbV16]. Polygon-Base [smbV16]. Polygonal [AB17, Cag20b, DeT88, Edg20, Edg22, Hug79]. Polygons [Fra75, Fre12, Hoe87, Lau20a, PS82a, Pag87b, Sco87]. Polyhedra [Ped80a, Sco87]. Polyhedrons [Car72]. Polynomial [Cho80b, DRW96, Die23, Dor77b, Ett74, Fla86, GK13, Gor93, Hig23, Lar74, Lau25, Lin74a, Lon72, MHS92, PE82, Ric10, San21, Sch87b, Sch84d, Sch00]. Polynomials [AM83b, AB92, BB01, BMV06, Cas03, Cer15, FKY13, GCW96, GS10, Her12, Hil92a, HM97, Jon20a, Jon20b, KP85, Kri95, KC90, Lef83, LH89, May97a, MP10, Mel22, Nor88, O’L98, Pin72, Pom77, Pri90, Red89, Rog07, Ros76, Sch72b, SS22, Suz99, Wal95b, Yan80, BSC17]. Polyomino [Cof09, HL12]. Polyominoes [Fre11]. Polytechnic [BE95]. Polytope [Put86]. Polytopes [Zho12b]. Pompeiu [BlC09]. Pool [Bon16, Daw97, IB22, Mab10]. Pooling [RS21]. Poor [Thu82]. Pop [Boa80, Wal00b]. Pop-up [Wal00b]. Pope [Atk05]. Popular [TYM88]. Population [CKM86]. Porisms [Mac92a]. Port [Sie24]. Port-and-Sweep [Sie24]. Portfolio [Bar23]. Position [Bom09]. Positional [Edg21]. Positioning [CHS87]. Positive [Che13, Lon81, Lyc23, Mer17, PS12, Sch23, Sou14, Wis99]. Positivity [Smi91]. Possible [BL22a, FG19, Sha80]. Possible-Paths [FG19]. Post [Man70]. Post-Secondary [Man70]. Potatoes [May24]. Potent [Lax76]. Pound [May24]. Power [ANANFT01, Alf15, Boa82, CC80, Day13, EI89, Fri89, HP14, RR98, Sch88a, Sch81c, Sno89, Str79]. Powerful [Bea17, Nel22a]. Powers [Abe15, Aie87, Daw94, HB97, Her12, Hir24, HMF20, Hsu82, Kri95, Kro83, Lea16a, LM20, MT22, Pfa11, Ric19a, SS21, Wis99, Wu23b, Zia91]. pp [Gra18, Gut19, Hop18a, Hop19, Kat18, Ser17]. Practical [Boa82, Vie78]. Practice [Hud11]. Praise [Lax84]. Pre [Sta79]. Pre-Calculus [Sta79]. Precalculus [Gol75, Sch80b]. Precise [Ric24]. Precision [BN03]. Predator [KKS91]. Predator-Prey [KKS91]. Predetermined [Buc85a, Fek88, Kri25]. Predicting [KM15]. Prehistory [Rou14]. Prelude [McC98a]. Preparation [HP99a, Per73]. Preparatory [Lei83]. Preparing [Col71]. Presentation [Wil02b]. Presenting [DF96, Sel70]. Preserving [BKSU19]. President [Kha13]. Presidential [Som02]. Press [Gra18, Hop18a, Kat18, Nun10, Ric19a, RC20]. Pressure [FIRW14, Ler21, MS19]. Pretrigonometry [Usi86]. Prey [KKS91]. Price [Eid93, Tol22, TYM88]. Pricing [Din13]. Prime [Ell83, Gre74b, Lee24, MS17a, Nel24a, PS15, Rib94, Rib97]. Prime-Generating [Ell83]. Primer [Wal95a]. Primes [Abb24, Asa84, Can89, Che05, LD15, Meh22, Nel19, Nor17, Sad17]. Primitive [Eck92]. Princeton [Nun10, Ric19a]. Principal [Hwa12]. Principia [Nau94]. Principle [BS78, Coc76, Pin74a, Reb79, SS99, Pin74a]. Principles [Sew18]. Priority [AOW05]. Prismoid [Ami18]. Prisms [Nog17, SH14]. Probabilistic [FBH83, Fal86, Mal22a, MP23, SL84]. Probabilities [AS02, BR87, Bra80, BDLS18, Ekl00, Fra87, Sul16]. Probability [AP92, AFL^{+}21, Aus82, AMM15, AKMMM23, Eck07, ESS08, Eis98b, ES23b, EH24, Han20, Jan80, Kep85, KRSW18, KJ11, KP86, Lie97, Lyc23, MS91, MW18, MST21, NS87, Nel79, NRT20, PS12, Ric24, Sha25, SY78, Sou14, Sto07, Swa05, TL00, Gol13]. Probes [Lom94]. Problem [Adl96, Ann98, Aus82, BLW89, BCR^{+}11, Bon18, BCD^{+}24, Bou99, BC18b, Buk08, Cha79, Cha09, Che81, Chu09, CHS86, DW87, DHR10, DLPS16, DS99, EH19, EG75, Eri22, Fla11, Fre03, GL92, GSS15, Gor86, Gor03, Haa91, Hal80, Han20, Hat04, Hat07, HS86, Hof78a, Hof78b, HS96, Hud11, Joh04, KK08, KS96, Kos09, Kri25, Lam77, Lar89, LM16, Log79, MP23, Mit81, Moo20, Mor10b, Nac17e, Nas23, Niv79, OAHS21, OB79, PV13, Pau85, Pau21, Per83, Pom84, RR89, Roj16, Sch16, Sch89e, Sch95a, SB84, ST20, Shu79, SB97, Smi78, Swe22, Tea06, Ten83, WS73, WS96, Wil99b, Wil94, pX91, Zen96, Oku22, Arn94, Ben12b, BM13, Gla12, Iga08, Mor16, Sch87b]. Problem-Solving [SB84]. Problem-Then [Pom84]. Problems [All16, Ano73k, Ano73l, Ano74n, Ano74o, Ano75i, Ano75j, Ano75k, Ano75l, Ano76e, Ano76f, Ano77f, Ano77g, Ano77h, Ano77i, Ano77j, Ano78f, Ano78g, Ano78h, Ano78i, Ano78j, Ano79i, Ano83f, Ano83g, Ano83h, Ano83i, Ano84j, Ano84k, Ano84l, Ano84m, Ano84n, Ano85h, Ano85i, Ano85j, Ano85k, Ano85l, Ano86h, Ano86i, Ano86j, Ano86k, Ano86l, Ano87g, Ano87h, Ano87i, Ano87j, Ano87k, Ano88k, Ano88l, Ano88m, Ano88n, Ano88o, Ano89n, Ano89o, Ano89p, Ano89q, Ano89r, Ano90k, Ano90l, Ano90m, Ano90n, Ano90o, Ano91k, Ano91l, Ano91m, Ano91n, Ano91o, Ano92k, Ano92l, Ano92m, Ano92n, Ano92o, Ano93l, Ano93m, Ano93n, Ano93o, Ano93p, Ano94n, Ano94o, Ano94p, Ano95p, Ano95q, Ano95r, Ano95s, Ano95t, Ano96q, Ano96r, Ano96s, Ano96t, Ano96u]. Problems [Ano97r, Ano97s, Ano97t, Ano97u, Ano97v, Ano98l, Ano98m, Ano98n, Ano98o, Ano98p, Ano99r, Ano99s, Ano99t, Ano99u, Ano99v, Ano00n, Ano00o, Ano00p, Ano00q, Ano00r, Ano01x, Ano01y, Ano01z, Ano01-27, Ano01-28, Ano02x, Ano02y, Ano02z, Ano02-27, Ano02-28, Ano03w, Ano03x, Ano03y, Ano03z, Ano03-27, Ano04u, Ano04v, Ano04w, Ano04x, Ano04y, Ano05u, Ano05v, Ano05w, Ano05x, Ano05y, Ano06u, Ano06v, Ano06w, Ano06x, Ano06y, Ano07t, Ano07u, Ano07v, Ano07w, Ano07x, Ano08w, Ano08x, Ano08y, Ano08z, Ano08-27, Ano09q, Ano09r, Ano09s, Ano09t, Ano09u, Ano10l, Ano10m, Ano10n, Ano10o, Ano10p, Ano11n, Ano11o, Ano11p, Ano11q, Ano11r, Ano12m, Ano12n, Ano12o, Ano12p, Ano12q, Ano13l, Ano13m, Ano13n, Ano13o, Ano13p, Ano14o, Ano14p, Ano14q]. Problems [Ano14r, Ano14s, Ano15o, Ano15p, Ano15q, Ano15r, Ano15s, Ano16m, Ano16n, Ano16o, Ano16p, Ano16q, Ano17o, Ano17p, Ano17q, Ano17r, Ano17s, Ano17t, Ano18j, Ano18k, Ano18l, Bai91, BM97, BS15, BR89, BL22b, Buc87, Cha07, Cof09, CDH^{+}14, Cuo00, Daw76, Fra86a, Gro93, Hed00, Joh94, Jus73, Jus74a, Jus74b, Jus76a, Jus76b, Jus79a, Jus79b, Jus79c, Jus79d, Jus80a, Jus80b, Jus80c, Jus80d, Jus80e, Jus81a, Jus81b, Jus81c, Jus81d, Jus81e, Jus82a, Jus82b, Jus82c, Jus82d, Jus82e, LS94, Mau83b, MP88, Moo91, Nel94, Nie21, OC18, OC19a, OC19b, OC19c, OC19d, OC19e, OC20a, OC20b, OC20c, OC20d, OC20e, OC21a, OC21b, OC21c, OC21d, OC21e, OC22a, OC22b, OC22c, OC22d, OC22e, OC23e, OC23a, OC23b, OC23c, OC23d, OST24d, OST24a, OST24b, OST24c, EST24, OST25, PP84, Rob78, Sie24, Ste85, Str72, Vau08]. Problems [Wag14, WCP88, Ziz98]. Procedures [DeT84]. Proceedings [Mit72]. Process [BW19]. Processes [BHS^{+}21]. Processing [KD92]. Produce [FG19, WW00]. Produced [BCR^{+}11]. Product [CFLR12, Daw80, EK11, HK19, Law24, McC80c, MS17a, Mol16, Nel25, Pet11, Sal18, Sch86f]. Product-to-Sum [Mol16]. Productive [Min91]. Products [Gra11a, Kle91, McA83, Nel24a, Nic93, Usi79]. Professional [DP71]. Professor [Knu91, Smu82a, Smu82b]. Professors [Joh72a]. Profile [Kol81]. Profiles [Gro10]. Profit [vR01]. Program [BB76, Car70, DN70, Knu91, RM71, Smi72, Tox70]. Program-Ron [Knu91]. Programmed [Lia70, Not71]. Programming [AR86, Hsi89, Knu91, Sem74, WN96]. Programs [Car75, Fri78, Gre77, HB92, Mar91, Ste76b, Wag87]. Progress [GW70]. Progression [Min94]. Progressions [Bai85, SK01]. Project [Far08, Gre00, PM74, Stu95, Win95]. Projectile [PY04, dA95]. Projectiles [Min94]. Projections [Fee01, Joh93a]. Projective [Dil14]. Projectors [Ban04]. Projects [Ano01-30, Ano04z, Ano06z, Ano07z, Ano09x]. PROLOG [Lop93]. Proof [Ada80, Alb81b, AN10, AN12, ANU14, AN22, AMP87, Ami18, AE17, AH81, Bad82, Bal78, Bal22, Bal23, Bar81b, Bas16, Bea17, Ber14a, Ber18, Bil18, Bos95, Cag15, Cag20a, Cag20b, Cag24b, Cag24a, Cha13, Cia17, Coa11, Coo24, Cow76, Cra12, CS03, Dav84a, Day13, DeM13, DH12a, Dr24a, Edg16, Edw14, EO17, Eis17, EH16, FZ12, Fos16, Fri24, Fri88, Gao13, GCM18, Gas23, Gen91, Gil97, Gor24, GI18, Heo15, cH94, Hu21, Hwa12, Ins89, JV15, Jav17, Jon12a, Ken99c, Kob13a, Kob13b, Kob14, Kob17, KK13c, Kwo16, LP94, Lau20a, Lea16a, Leh90, Lem16, Li13a, Li13b, Lot23, MM14, Mal22a, Man81, MM21, Mat25a, Mat92, May97a, Meh22, Mic14, Mol16, Mol17, Mor16, MGC19, Mor15, Muk24, Nat24, Nel12, Nel13, Nel14a, Nel14b, Nel14d]. Proof [Nel14c, Nel15e, Nel15b, Nel15c, Nel15d, Nel15a, Nel16f, Nel16c, Nel16g, Nel16d, Nel16a, Nel16b, Nel16e, Nel17a, Nel17b, Nel19, Nel21, Nel22a, Nel22b, Nel24a, Nel25, Pal17, PS22, Pap17, Par95, PV13, Per03, Pla14, Pla16, Pla17a, Pla17b, Pla17c, Pla18, Pow80, Pri18, Pro76, Raz15, Ren14b, Ren14a, Ren81, Ric15b, Sal18, Sar23, Saw17, Sch72a, Sch78a, Sch82c, Sch85b, Sch88a, Sch89c, SK89, Sch91b, Sch94a, Sch23, Sch86f, Sch95b, SH14, Ser15, Sha15, She78, Sol23, Spo22, Sub15, Sun21, Tan15, Thu15, Tre18, Tur81, Tym81, Una09, Una10, Una14b, Una14a, Usi86, Wan14a, Wan14b, Wie87, Woo25, Wu23a, Wu23b, Wu24, Xu14, Zac18, Zho18, Kat18, Wag11]. Proofs [Ang77, Ber20, BCR^{+}11, Coc76, DR06, Ehr24, Gal81a, Hay92, Hen90, JS75, Lei07, MM21, Nor17, Rob90, SS21, Wat89]. Propeller [Gar99, SJC00]. Proper [Sha95]. Properties [Hoe87, Shp14]. Property [AMR13, AT08, Bro84, Cas87, Cow76, DeT94, Fri24, KN15, Mas11, PS82a, PS22, Par95, Sch77, Sch95b, Spi06, Spo22, TR23, Usi86, FN07]. Proportion [MS20]. Proportional [FM00]. Proportionality [Lau20b]. Proportions [Har86, Kob14]. Proposition [Bas16]. Prospective [Ano79e]. Prosthesis [SY87]. Prove [Ald79b, Sad17, Wal11]. Proved [CFLR12, Ric15a]. Providence [Ser17]. Providing [You70]. Proving [Dob84, Len15, Pól78, Pól79]. Pseudo [MM93]. Pseudorandom [Reb89]. Ptolemy [DH12a]. Public [LP87]. Public-key [LP87]. Publisher [Ano70g, Pri72]. Pull [BJ24]. Pull-Back [BJ24]. Pulley [Bou99]. Pump [WRR98]. Pumping [FG10]. Punxsutawney [ABCS01]. Pure [Hal85]. purgated [AB91]. Purposes [Boa82]. Pursuit [See98, Sim07]. Push [Tee12]. Push-To [Tee12]. Put [Lan93b]. Putnam [HS96]. Putting [DM10, JM97]. Puzzle [Ano14a, Ano24f, Die23, Kat03, KK09, Lam96, Nac14, Nac15a, Nac15b, Nac15c, Nac15d, Nac16a, Nac16b, Nac16c, Nac16d, Nac17a, Nac17b, Nac17c, Nac18a, Nac18b, Nac18c, Nac18d, Par05, Ric10, Sul24]. Puzzles [Ano09w, Ano15h, Lev14, LM21, Nac16e, Nac17d, Nac18e, Nee13, Ped73, Sal14, Wil93]. Puzzling [How18, vD09]. Pyramid [Lo 91, Wal00a]. Pyramids [McC98b]. Pythagoras [Fos16, NK22, Ren14b, Ren14a]. Pythagorean [AA19, AN10, AN12, AN22, Bas16, BS96, CG24, Coo13, DW19, Dun77, Eck92, For92, For17, Gaz92, Gen89, Har86, Heo15, Hil85a, JY98, Kal86, Kau20, Kle72, Lem16, Len15, MW03, Mol17, Nel14b, Nel16c, Nel16e, Nel19, PS03, Sch82a, Spo17, Spo23a, SY12, Tir86, WW00, Wie84, Hop19]. Pythagorean-like [AN10].

Quadratic [Cho80c, DeT84, Hor90a, Kal84b, Kra03, Lon72, Mis23, Nie03, PV05, San21, SD93, Ves85, Xu10, Xu14, LP94]. Quadratics [For17, SS22]. Quadrature [Bur87b, Osl06, Sny24]. Quadric [Tee93]. Quadrilateral [Coo24, Gre74a, Mey86, Sar23]. Quadrilaterals [AC22a, Bea06, Bea09]. Quadrinomials [Bri20]. Quadruples [Nel14b]. Quaint [Hed00]. Qualifications [Joh72a]. Quantitative [BM95]. Quarter [Rok14]. Quarter-Turn [Rok14]. Quarterbacks [Joh93b]. Quartic [MHS92]. Quasi [Dun77, Seg83]. Quasi-Geometric [Seg83]. Quasi-Pythagorean [Dun77]. Quasigroup [Lyc23]. Quaternion [Moo86]. Quaternions [Ban15, Sch81d]. Queens [Pau19, Cha09]. Quenching [Ehr94]. Queries [Cho79]. Question [EH20, Hol78, LJ73, TT21, Way80]. Questioning [Alb96]. Questions [Dav94, Kun89, Pag79b]. Queueing [BO95]. Quick [Cho80b, Seg00]. Quiz [Ziz13]. Quotient [Cha91, Cur05a, DH12b, EK11]. Quotient-Rule-Integration-by-Parts [DH12b]. Quotients [Pla78]. Qwirkle [BDLS18].

Ra [Sup84]. Rabbits [Bri75a]. Race [Fre12]. Racing [PV16]. Racquetball [BP12, KRSW18]. Radial [WP90]. Radical [Bom82]. Radii [Lon83, Sim11b]. Rail [Man98]. Rain [Bai02, Bai13]. Ralston [Ano84b]. Ramanujan [BS02, Par05]. Ramifications [AA04]. RAMS [Sup84]. Random [BR89, Eis17, EG75, GS11, Gre74b, HS96, JR88, Lan77, LM92, LM10, Mur84, Wri89]. Randomized [Kun89]. Randomly [AKMMM23, Tre20]. Randomness [Ber02]. Range [LM10, TT12]. Ranges [Pag83]. Rank [Cal98, Kir97, Kun11, LP14, VK00]. Rank-One [Kir97]. Rankings [Coo86, JWW16]. Ranks [Alb12]. Raphson [CMS20]. Rascal [ALT10a]. Rate [Joh93b, PS82b]. Rates [AW85, Ful00, PG06]. Rather [Bas16, Fre88, PG06]. Ratio [Che13, Een97, Fal05, FS15, Gro08, LJ22, MS88, Mar92b, Nel16f, Nel25, PS82a, PM82b, SR16, SW95]. Rational [ANANFT01, AM01, BBB03, BF00, Boy10, Boy15, Bro99a, BW91, Can86, Cho80b, DPV21, Den12, Dia17, Dr24b, Gre75, GS10, Hot02, Lau25, Lin85, Lut80, Lyc23, LM24, McC80a, O’L98, Pac78, Pag87b, PS12, Red89, Sou14, Tym81]. Rationals [CS03, GP75, Mih04]. Rationing [Bur77]. Ratwyt [Fra12a]. Rault [Kat18]. Raymond [Sch11a, Gar82]. Reach [Niv79]. reaching [Lee99]. Reader [KGSP88]. Reading [AB91, Hos18b, Ken99b]. Real [BT07, FH00, JK98, MM73, Moo86, Oma09, RS00, Ska96, Sta88, Hwa12]. Realities [Lat83]. Reality [Flu88]. Realization [BDCH85]. Really [Bro15, Pen20, Sto04]. Rearranged [AE17, Sch86a]. Rearrangement [BM13, Eck97]. Rearrangements [AE19, BCEW85, Smo24]. Reasoning [Ric15a, RC99]. Rebalance [Bar23]. Reciprocal [AS96, BDDY13, DL75]. Reciprocals [Cag15]. Reciprocity [LP94]. Recirculation [BHB04]. Recommendations [CJLP93, Jon71, Per73]. Reconstructed [HM23]. Reconstruction [Tym81]. Record [BBHS02]. Records [Rib94]. Recounting [BQ99]. Recruiting [PM74]. Recruitment [CDK^{+}90]. Rectangle [Ber18, DR20, Lan93b, Ren14a]. Rectangles [Lat83]. Rectangular [Bro23, Ess82, Pri95]. Rectangular-to-Polar [Pri95]. Recurrence [Lam96, LP14, TV94]. Recurrences [Mat14]. Recursion [Aie87, Hay92, MR03]. Recursions [WW00]. Recursive [Car87, Che95, Coc76, Ekl01, MW03, Pla14, TT21]. Recursively [Bur87a, CM90, Mat25b]. Red [AKMMM23]. Rediscovering [Aug12, Kal85b]. Reduced [Lot23, VB86]. Reduction [Bou23, Gre11, KC90, SM91]. Reductions [FG10]. Redundancy [GS96]. Redundant [Cas87]. Redux [GW18, NW10]. Reencountered [Dav88]. Reexamining [Cel99]. Referees [Ano07y, Ano08-28, Ano09v, Ano10q, Ano11s, Ano12r, Ano13q, Ano14t, Ano15t, Ano16r, Ano17u, Ano18m, Ano19j, Ano20h, Ano21i, Ano22f, Ano23g, Ano24h]. Reflection [Bro84, Cow76, Usi86]. Reflections [Ayo87, Bri84, Cha09, Mal15, MC24]. Reflective [Spo22]. Reframing [AA19]. Regarding [DD95, TT21]. Region [ACJ^{+}06, RS85, The23]. Regions [Grü84b, MC20]. Regression [Fra92, Hoa88, Lut17, Wil02a]. Regular [Cai73, Dr24a, Fra75, Fre12, Jac97, JV15, Nel15e, See98, Wu23a]. Reid [Ale80]. Reinventing [CV14]. Rejection [Mil71, Saa70]. Relabeling [FS99]. Related [AW85, Bob95, Cea81, Cea84, Dub92, Ful00, Hof78a, Mer22a, Mye77, PG06, Ric24, ST20, Som02, Swe22, GP76]. Relating [BK85, HP99a, HP99b]. Relation [Lam96, Liu11, Moo23, TV94]. Relations [AN16, Bri75b, CW90, Gil97, PM82a, PM83, Slo77]. Relationship [Cag24b, Cag24a, Cag24c, Nel88, Nel92, Str90b, Vel23]. Relative [Ano92f, Boy10, Cha92, Hil92a]. Relativity [DR04, Sha95]. Relaxing [CS17]. Relevant [Cla73]. Reliability [GS96]. Remainder [BSL09, Bre08, How02, Per17, Sch87b]. Remark [Kit19, Mar82]. Remarks [Abb93, Ano88g, Ven76]. Remedial [Man74, Ros70]. Remember [Hat83]. Reminiscence [Tuc83]. Removing [Jon11]. Rencontres [Dav88]. Rensselaer [BE95]. Rental [WS96]. Reordered [Ric19a]. Repeatability [SL84]. Repeating [AT08, Fje95, Gan84, Lea84, McK76a, Pla78]. Repetitions [Hyd88]. Replace [EH20]. Replies [Ano96n, WK94]. Report [BC22, Baj23, BT24, BCT25, Cun08, Fle79, GW70, Kha13]. Representations [ALT10b, Lon81, Pio21, Sch81c]. represented [Her10]. Reprint [Hop19, Ser17]. Repunit [Fra88]. Rescue [Ler21]. Research [Ano01-30, Ano04z, Ano06z, Ano07z, Ano09x, Chi82, Mer74, Not71, SS93, WM76]. Reserve [CHS86]. Residue [Sch72b]. Resistance [PY04, dA95]. Resolution [Wal04b]. Resolved [LJ73]. Resolving [Chr96, MZ24]. Resources [DK01, JSW22]. Respect [Gre75, Wie01]. Response [Ano84b, Ano84h, Ano88a, Dan85, Gia73, Kun89, Mat96, Wil94]. Responses [KGSP88]. Restaurant [Yin87]. Restless [Mey06]. Result [BAW87, BS02, BF77, Jep79, Kri95, Nel16b]. Results [Gar00b, Mye77, Not71, Sch84a]. Retention [CDK^{+}90]. Rethinking [Bro99a]. Retirement [Dan98]. Retreat [HA70]. Retrolife [Elr12]. Return [Mey84]. Revealed [Ber02]. Reveals [Ric19a]. Revenge [Str85, VV24]. Review [Ale07, Ale08, Ano14g, Ano15a, AG14, Bac09, Bar11, Bla01, Bra13, Bur16, Car17, Che15, Col13, Cox12, Cra11a, CO92, Dev05, Dic17, Ewi99, Fus13, Gar07, Gar15, Gol17, Gol10, Gol11b, Gol13, Goo08, Gra03, Gra18, Gre16, Gut19, Hen14, Hop14a, Hop16, Hop18a, Hop19, Hug06, HN91, Kap17, Kar13, Kat18, KB99a, Ken01, Kho14, Kin10, Kly15, Kra01, Lin00a, McD00, Mor11, Mur10, Nat91, Nun10, O’R15, Ols13, Qui16, Qui17, Ric19a, Ros17, Ros00, Ros02, Ros06, Ros16, RC20, Sau99, Sch11a, Ser17, Sol99, Sri10, Swe02, Swe03, TB91, Wag11, Wil16, Wim02, dF89]. Reviews [AB83, AH82, Ale83a, And84, Ano70h, Ano70i, Ano71g, Ano71h, Ano72h, Ano72i, Ano73m, Ano80j, Ano80k, Ano80l, Ano81l, Ano81m, Ano84o, Ano84p, Ano92p, Ano00s, Ano01-29, Arg93, AA95, Bau98, Bau93, BBM87, BSR85, Bou81, BR94, BB86, Cam82, Cam91, Cho91, CSL^{+}74, CM98b, CB09, Col10, DeT93, DF85, DL84, Fin94, Fin97, Fis96, FW94, Fra97, Gil95a, Gil95b, GSH87, Gul93, Har97, Hen95, Her09, Hoe90, Hos86, HAC88, JBC89, Kal99, KK90, KLB^{+}87, Lan83, Lat73, LH86, LEHE87, LP88, LHL88, Lei90, LH92, Lei93, Lei95, Lei96, Lei97, Lew02, LMK79, LF79, LTM79, LRM79, LM79, LG80, LB80, LC80, LH80, LGA81, Lin81a, Lin81b, Lin81c, Lin82a, Lin82b, Lin82c, Lin83a, Lin83b, Lin83c, Mal81, May97b, Moo84, MS94, MH09]. Reviews [MSFA85, Pag89, PGNH90, Pal92, Pen89, RCS86, Rou09, Sag92, SZ89, SDC85, Sol94, Tan85, Tha01, Tho96, Tro82, VW73, Wei91, Wil96, Wil98a, Wil99a, Wil98c, WHC^{+}74, WHD^{+}74, WHS^{+}75b, WH75, WHS^{+}75a, WHT^{+}75, WHH^{+}76, WHPT76, WHB^{+}76, WHL^{+}76, WHLD77, WHGR77, WHMB77, WHFM77, WHMS77, WHN78, WHLB78, WHS78, WHDS78, WHH^{+}74, Zeh96, dFMN^{+}90]. Revisions [Ano09a, Ano10a, Ano11a, Ano12a, Ano13a]. Revisited [Ale79a, Ale79b, Ale79c, BlC09, BZ05, BJK17, Cha79, Che09, Chi73, EK11, GP75, Hal80, Hat04, Jon93, Kar14, Mal22b, Mar09, Nel92, Pau21, SJC00, Smi78, SM91, pX91]. Revisiting [Ban04, Gil98]. Revolution [AK75, Knu18, Lam24, MMOM14, RR98]. Reward [Ano91p]. Rewards [Pen20]. Rewriting [Bur92]. Rhombi [CS95]. Rhumb [NM96]. RI [Ser17]. Rich [Ban90]. Richard [AA93]. Richardson [Sch90b]. Ridge [Daw76, Lon76]. Ridge-Length [Daw76]. Riemann [BHS^{+}21, Bil18, BGR86, Gor94, Gro19, Kac19, MS89, Mur80a, Olm73, Pio21, Tho12, Tho23, Tru19]. Riemann-Integrable [Pio21]. Right [DS08, Fos16, Gao13, KS88, Lan95, Lon83, Nel16a, Par84, Ren14b, Thu15, TYM88]. Right-Hand [Gao13, Thu15]. rigid [GG22b]. Rigor [Bro99a, Gal81a]. Ringing [IM11]. Rings [Ano91p, Arp74, Kle91]. Rise [Kac17, Oss81]. Risk [MZ24]. Risk-Averse [MZ24]. Roads [Wag11]. Roast [Cib93]. Robbers [STW17, ZV07]. Robbins [Pag84b]. Roberts [Ano84h]. Robin [Hop18a, AA90]. Robinson [JM97, Rei86]. Rock [Bir15, Cur05b]. Rocket [FSW04]. Rod [GS11]. Roger [Edg18]. Rogers [JMW24]. Role [Ano70l, Bai88, Bra00, Mas84, MC78]. Roll [Car93]. Rolle [Smi86, Gen73]. Roller [Dru92]. Rolling [AM10, Eid93]. Roman [CAS20]. Ron [Kat18, Knu91]. Ronald [Kol81]. Rooks [MST21]. Room [BW19]. Root [BBB03, Edg23, Lut80, Nel89b, Par91, Sch00, Smi87, TT03]. Root-Finding [Par91]. Rooted [Mar90b]. Rooting [Ben12a]. Roots [Aie87, AM83b, Asa84, BT07, BBB03, BS14b, Bri20, Bro99b, Cho80b, CMS20, Dor05, Ett74, GCW96, GS10, HM97, Hug79, Lau25, Law21, May97a, McK76b, Mer17, Mis13, Nel17a, NL84, Pri99, Red89, San21, Sue94, Ven76, Ves85, Wag77, Wej83, Zac18]. Rose [Gra90, Har95, She97]. Roses [Hal92]. Rotating [BF13]. Rotation [Alp89, GG23, Gro05b, Ins93, Knu18, Pen88, Sch92b, Wan21]. Rotations [Bri84, Due96, Kra91, Mal15]. Roulette [BS14a, SR12a]. Round [CM15, NS87, Pag79b, Roz94, RS16]. Round-off [Roz94]. Rounded [Nie01]. Rounding [Bor11, GL00]. Rounds [NS87]. Route [Par91]. Routine [Ste85]. Row [Den07, KC90, Lot23, Pla17a, Pla17b, VB86]. Row-reduced [VB86]. Rubber [Kuc16]. Rubberbanding [Kir92]. Rubik [Dav82d, GPV24, JSW14, Rok14]. Rudin [AR88]. Rug [Sch95c]. Ruin [BS14a, OAHS21]. Rule [Bil93, Buc85b, Cur05a, Daw80, Day13, DH12b, Fre95, Fri89, Gao13, Gil97, Gro97, Hua25, Ken99c, Kro79, Loe21, Mar98b, Thu15, Tro04, Bal78, Bom09, Bum84, Bur83a, Bur87b, Car78a, Eck11, Eul93, Fri88, Hua88, Kal87, Kre97, KR96, Kun88, MB82, Pin01, Rog07, Str85, Uet88, Var11]. Ruler [Kow08]. Rulers [KR91]. Rules [EK11, Gar93, Hut83, LW18, Mer05, Ric12]. Run [AAC04, BBHS02, BMV14, Sch90a]. Running [Bai02, Lor00, Zho18]. Runs [Ber02, Ilt00]. Rupert [JW08]. Russell [Lo 97].

S [Kat18, Log80, JSW14]. Saari [Bir10, Hau05]. Sabine [Gut19]. Saga [smb12]. SAGE [Den13]. sagemath.org [Den13]. Sahin [Kat18]. Saidak [Abb24]. Saint [Cea81]. Saints [Bro15]. Salinon [CF17]. Salmon [Ser17]. Same [Bro15, Gla93, Hoo00]. Sample [Boo79, LM10, Mur84, Nea93, Woo72b]. Sampler [TS83]. Samples [RS21]. Sampling [Wri89]. sangaku [Oku22]. Sanity [Wil99b]. Sarrus [LW18]. SARS [Roz07]. Satan [Smu85]. Satellite [Tee17]. Saunders [Ale89]. Saupe [Nat91, Nat91]. Savings [Cho81]. Saws [Vau08]. Scarce [JSW22]. Scenes [Sim11a]. Scholar [Wag99]. Scholes [Sew18]. School [FM77, LS94]. Schools [McK79]. Schwarz [SS79, Zho17b]. Science [Boa71b, Byr11, Mey07, Wag83]. Scientists [Luc84]. Scissors [Bir15]. Scores [Fos12, SP79]. Scott [Ken87]. Scoundrels [Bro15]. Scratch [Gar24]. Scratch-off [Gar24]. Sea [Kac17, PL82, Sim09b]. Seam [AS22]. Search [Liu01, Ste76b, Str91]. Searching [CW23]. Seashells [Boy07]. Season [BC94]. Seasonal [Mar13]. sec [SS83b]. Secant [Ful05]. Secants [Mal15]. Second [Ano70j, Bel15, Lon83, Mal17, Mat14, Mer74, Nel24d, RZ99, SS77, Str97b, Swe22]. Second-Order [Mat14]. Second-Year [Ano70j]. Secondary [Man70]. Secret [Gin04]. Secretary [Gla12]. Secrets [Ano91q, JLR11, Ric19a]. Section [Ayo03, Ven80, Wil03]. Sections [BS15, Ber14a, Cul90, Knu18, Ser17]. Sector [Mel22]. See [Ewi95, Pól76a, Pól76b, Pól76c]. Segment [Gor08]. Segments [ES23b, EH24, FM00, FG19]. Selected [Gui95]. Selection [PS82b]. Selective [Kal87]. Self [Boy07, BS72, Dav85, Han85, Kob17, Lin85, Suz14, Wil90]. Self-Awareness [Dav85]. Self-contained [Lin85]. Self-esteem [Wil90]. Self-Instruction [BS72]. Self-Inverse [Han85]. Self-Limiting [Suz14]. Self-Similar [Boy07]. Self-Similarity [Kob17]. Selling [Bar23]. Selves [BL22a]. Semester [EB99, Per83]. Semigroup [PS12]. Semigroups [Sou14]. Semipartial [Hub81]. Semiregular [Sco87]. Senet [NRT20]. Sensei [Knu91]. Sensitive [Kun89]. Separation [Moo91]. Sequel [Mos84]. Sequence [Blo98, Bob95, CK12, CK14, Coo84, Cra12, Du20, Ols96, San77, Whi71]. Sequences [Alf15, BNUW23, Blo01, Ćur14, Ekl01, FM00, Jon83, Kut81, Lon84, Men88, Pri73a, Sch84d, SS05]. Sequencing [Coh77b]. Sequential [JMW24]. Serendipitous [Dor05, For92, MR96, Per03]. Series [AMP87, AE17, AE19, Ash97, Ben12b, BM13, Bén05, Ber20, Bes14, Boa82, Bob95, Bro04, BCEW85, BDDY13, Bur00, Cag15, Cea81, Cha99, CK12, Che13, Cia17, Col92, DG80, Dun87, Eck97, Eck98, Edw14, EI89, Eul97, FN04, FST23, Gil88, Gor24, Gor89, Gri78, Gro16, Grz24, Har88, HP14, Hug00, Jon83, Kal93, KS12, Kat12, Kem82, Kep85, KB99b, KG12, Kos17, Kre97, Les04, Lib81, LM06, MM21, Mor86, MS84, Nam24, Nea94, Nel12, Nie78, OW02, Ovc11, Pio21, Pla18, Sch86a, Sch86b, Sch87a, Sch92a, Sch81c, Seg83, Seg89, SR16, Shi78, Smo24, Sno89, Una09, Wu23a, Wu24]. Series-based [KB99b]. Serret [Hoe09]. Services [Lan72]. Session [RP85]. Set [Bri88, Cro95, Dev91a, Ewi95, Hol23, Joh72b, Mih04, San19, Sch74b, Sek01, SC03, VU21, Gol16, MS20, NS17, Nat91]. Sets [Abb93, BBB^{+}13, BE03, Dav92, Fon09, Iga08, Mar08, McA83, PV09, Sch99b, SW95, Tro89, Wag20]. Settlement [AMM15]. Settlers [AMM15]. Seventeen [Hol79]. Several [Lan76, Sch99b]. Sextet [Sch12]. Shad [Cra11b]. Shadows [Hob13]. Shanks [Bro99b]. Shape [Roe93]. Shaped [The23]. Shapes [Kle71, Nog17, Sch99b, Shp14]. Shaping [Ser15]. Share [Par16]. Sharing [Fre13, Sin90]. Sharma [Hop19]. Sheep [Sau90]. Shell [McG18]. Shells [Key94]. shen [CL83]. Shiing [CL83]. Shiing-shen [CL83]. Shin [Col13]. Shirts [CBL19]. Shoe [MW81]. Short [Bus71, Car70, EO17, Eis17, Hop18a, Mal22a, Meh22, MS18, Mor16, Nat24, Nor17, Pap17, Raz15]. Shortcut [Hua91]. Shortened [PC81]. Shorter [Wie87]. Shortest [GRV21, Mar98a, SA99]. Shot [Bol83, Mab10, Smi86]. Should [Bre92, Eis98a, God88, Hah88, KGSP88, PM88, Roz88, Tan88, Ziz13]. Shouldn’t [Grü81]. Show [TYM88]. Shrinking [Pet15]. Shuffling [CC95]. Shutting [RT07]. Side [BS04, MF94, Ros12, Thi79a, Thi79b, Thi79c, Thi79d, Thi79e, Thi80a, Thi80b, Thi80c, Thi80d]. Sided [Lut84, Wol71, Zem71]. Sidedness [Cid21, GCM18, Pap17]. Sides [Bea09, DW19, Len15, Sin90]. Sidney [Ano81k]. Siege [Ols13]. Sierpiński [Hop19]. Sieving [EG75, Gre74b]. Sign [Cla11, MM21, PL94, Sch13]. Signal [KD92]. Signed [Alb12]. Significance [DK97, Ols89, VU21]. Significant [Smi89]. Signpost [Ano83j, Ano83k, Dav82a, Dav82b, Dav82c, Kin79a, Kin79b, Kin79c, Kin79d, Kin79e, Kin80a, Kin80d, Kin80e, Kin80b, Kin80c, Kin81a, Kin81b, Kin81c]. Signs [Fre95]. Silent [Ait16]. Similar [Boy07, Roj23]. Similarities [Sac22]. Similarity [Kob17]. Similitude [Sch99b]. Simmon [Ros04]. Simple [AMP87, AC22b, Bla91, BC75, Bou99, Car87, Coo82, Cow76, Gao13, GEJ95, GP02, cH94, HS00, KW98, KS24, Mar98b, McC23, MW18, Nat24, Prc82, Rob97, Sch78b, ST19, Sil89, Ven80, Wri89]. Simpler [Thu15]. Simplex [Hsi89, WK73]. Simplified [Daw19, LM16]. Simplifies [Gre11]. Simpson [Bur87b, JW19, Kna85, Kre97, KR96, Pin01, Ric24, Rog07, Tan86a, Uet88]. Simulated [Alb07b]. Simulation [Cea84, Con17, Flu88, Fra92, Lew95]. Simulations [Ker83]. Since [Ken00]. Sine [BL22a, Cip87, DRC98, FZ12, cH94, Hue92, Kow16, MP14, Nel14c, Nic93, PK83, Pet14, Sal18, SR12b, Una10, Wan14b]. Sines [Usi79]. Single [DW88, Ken91, Kow16, McG18, Smi91, Sou14, Wil73]. Single-Variable [McG18]. Singles [Blo98]. Singular [Coa11, JLR11, JB13, MR98, ZF97]. Singularly [Kal96b]. SIR [Wan10]. Site [Dra00]. Six [Hen90, Kal93, Zer02]. sixth [Ser17]. Sizes [Nea93]. Sketch [Tur85b]. Sketching [Wit17]. Sketching-Impaired [Wit17]. Sketchpad [Hoe97, Jur94]. Skew [SH85]. Skills [BM95, Hav78]. Skin [Bra16, DD10, Sin12]. Skinned [Ste23]. Skinning [Gür22, Nel18]. Skip [Ziz13]. Sledge [Ahn09]. Sledge-Hammer [Ahn09]. Slicing [vR97]. Slide [JSW14, WM76]. Slide-Tape [WM76]. Sliding [GL92, KK08]. Slippery [Bea11]. Slope [Bea11, Mac95]. Slouching [Bai13]. Slower [BK11]. Slowly [Sch86b]. Small [BMV06, BF23, Dia17, GK77, Ste76b, Wei76b]. Smart [Ruc82]. Smartphone [Qui16]. Smith [Ser17, Wil82a]. Smooth [ACJ^{+}06, NR16]. Smullyan [Gar82, Sch11a]. Snakes [GPV24]. Snarks [smb12]. Sneaking [Ped72]. Snell [Qui22, Gol13]. Snowfalls [Boa80]. Snowplow [pX91]. Social [Mey07]. Socially [Cla73]. Society [Ser17]. Sock [Pau21]. Soft [SJ11]. Software [And84, Ano84o, Ano84p, Ano92p, Ano00s, Ano01-29, Arg93, AA95, Bau98, Bau93, BBM87, BSR85, BR94, BB86, Cam91, Cho91, CM98b, CO92, CS86b, CS87, CS88, DeT93, DF85, Den13, DL84, Fin94, Fin97, Fis96, FW94, Fra97, Gem89, Gil95a, Gil95b, GSH87, Gul93, Har97, Hen95, Hoe90, Hos86, HAC88, HN91, JBC89, Kal99, KK90, KLB^{+}87, LH86, LEHE87, LP88, LHL88, Lei90, LH92, Lei93, Lei95, Lei96, Lei97, Lew02, May97b, Moo84, MS94, MSFA85, Pag89, PGNH90, Pal92, Pen89, RCS86, Sag92, SZ89, SDC85, Sol94, Tan85, Tha01, Tho96, TB91, Wei91, Wil96, Wil98a, Wil99a, Wil98c, Zeh96, dF89, dFMN^{+}90]. Solar [Phr90]. Solid [Vie78]. Solids [Ara11, Cai73, DH03, Gro05b, Knu18, Lam24, MMOM14]. Solitaire [Bia98, Hop12, Sie24]. Solitary [Han70]. Solomon [Sim09b, Tuc83]. Solution [Ano14a, Aus82, BF00, Bon18, Bro80, Gor86, Joh22, KW98, Kow24, Mit81, Moo20, Nel77, NW98, Nie21, RR89, Ric19c, SA99, Sul24]. Solutions [Ano73k, Ano73l, Ano74n, Ano74o, Ano75i, Ano75j, Ano75k, Ano75l, Ano76e, Ano76f, Ano77f, Ano77g, Ano77h, Ano77i, Ano77j, Ano78f, Ano78g, Ano78h, Ano78i, Ano78j, Ano90k, Ano90l, Ano90m, Ano90n, Ano90o, Ano91k, Ano91l, Ano91m, Ano91n, Ano91o, Ano92k, Ano92l, Ano92m, Ano92n, Ano92o, Ano93l, Ano93m, Ano93n, Ano93o, Ano93p, Ano94n, Ano94o, Ano94p, Ano95p, Ano95q, Ano95r, Ano95s, Ano95t, Ano96q, Ano96r, Ano96s, Ano96t, Ano96u, Ano97r, Ano97s, Ano97t, Ano97u, Ano97v, Ano98l, Ano98m, Ano98n, Ano98o, Ano98p, Ano99r, Ano99s, Ano99t, Ano99u, Ano99v, Ano00n, Ano00o, Ano00p, Ano00q, Ano00r, Ano01x, Ano01y, Ano01z, Ano01-27, Ano01-28, Ano02x, Ano02y, Ano02z, Ano02-27, Ano02-28, Ano03w, Ano03x, Ano03y, Ano03z, Ano03-27]. Solutions [Ano04u, Ano04v, Ano04w, Ano04x, Ano04y, Ano05u, Ano05v, Ano05w, Ano05x, Ano05y, Ano06u, Ano06v, Ano06w, Ano06x, Ano06y, Ano07t, Ano07u, Ano07v, Ano07w, Ano07x, Ano08w, Ano08x, Ano08y, Ano08z, Ano08-27, Ano09q, Ano09r, Ano09s, Ano09t, Ano09u, Ano09w, Ano10l, Ano10m, Ano10n, Ano10o, Ano10p, Ano11n, Ano11o, Ano11p, Ano11q, Ano11r, Ano12m, Ano12n, Ano12o, Ano12p, Ano12q, Ano13l, Ano13m, Ano13n, Ano13o, Ano13p, Ano14o, Ano14p, Ano14q, Ano14r, Ano14s, Ano15h, Ano15o, Ano15p, Ano15q, Ano15r, Ano15s, Ano16m, Ano16n, Ano16o, Ano16p, Ano16q, Ano17o, Ano17p, Ano17q, Ano17r, Ano17s, Ano17t, Ano18j, Ano18k, Ano18l, Bou99, BL22b, Cho80c, Chu09, CDH^{+}14, D’A23, For17, Hor90a, Joh94, Jus73, Jus74a, Jus74b, Jus76a, Jus76b]. Solutions [KH08, Kun72, Kun73, Mac92b, Nac16e, Nac17d, Nac18e, Nee13, Nel94, OC18, OC19a, OC19b, OC19c, OC19d, OC19e, OC20a, OC20b, OC20c, OC20d, OC20e, OC21a, OC21b, OC21c, OC21d, OC21e, OC22a, OC22b, OC22c, OC22d, OC22e, OC23e, OC23a, OC23b, OC23c, OC23d, OST24d, OST24a, OST24b, OST24c, EST24, OST25, PP84, Sch85a, Smu82b, Tea06]. Solvable [Mic12]. Solve [LM90, Mis23, Moo91, Nie03, Sch99a]. Solved [Par13]. Solving [Bom82, Bou23, CG19, CM20, DW87, Hof78b, Lut80, Oli73, OB79, SS22, SB84, She75, UK96, Wol23]. Soma [GP02]. Some [Adl96, AK75, Bel84, BD96a, BD96b, Biv97, Bla91, Boy15, BM93, Bum84, Bur86b, Bur77, Cea81, Che96, Cla73, DK97, Daw76, Den07, Fis75, Gar00b, Gre74a, Hel73, Hil80a, Hil80b, Hoe87, Hof78b, Lan80, Lat83, Len81, Mau83b, Mer74, MS78, Mun04, Mye77, NRT20, Pri89, RS96, Rob84a, Sch74c, Sch77, SL84, Sch90b, ST19, ST20, Sel70, Ser15, SS81a, Sta79, Str72, Val15]. Something [Bes14, Sta98]. Sometimes [LS12]. Sonar [WWW83]. Sophie [MGC19]. Sophomores [BEN78]. Source [Den13]. Southern [Zas70]. Space [Ald23, Kal84a, Kra91, Nor09, Sim06, WS81]. Space-Filling [WS81]. Spaces [Car98, ESS08]. Spaceships [NS17]. Spacewalk [Hes95]. Spanish [San02]. Sparsely [KR91]. Special [Biv97, Dev91b, Hos18a]. Specified [AFL^{+}21]. Spectacle [Sim09a]. Spectral [Sob97]. Speed [AW85, CSWW01]. Speeding [Gue01]. Spell [Gra94]. Sphere [Seg82b, Siu84, DS08]. Spheres [Cha80, GY00]. Spherical [Ara11, Ban04, FS15, Hob13, Mur95, Par83]. Spheroid [Sch07]. Spider [Hes95, MV12]. Spiders [Rid90]. Spills [Nie94]. Spin [Sny24]. Spinners [CG14, FGL^{+}18]. Spinout [Lam96]. Spiral [Gra23]. Spirals [Boy99, CBL19, Mat25b]. Spirograph [Hal92, Sim97]. Spirolaterals [FC91]. Splines [Hil90, KR96]. Splits [Doa89]. Spock [Bir15]. Sports [Hei78, Min94]. Spread [BR24, Mic12]. Spreadsheet [Arg84]. Spreadsheets [Hsi89, Sno89, Vod97]. Spring [CH99, HA70, SJ11]. Sprinkler [SR10]. Sprouts [Pri76]. Square [Abb93, Bro99b, Cag24c, Cha13, Dav92, Dr24a, Edg23, HKP15, Han90, Jep86, Jon12a, Jun74, LC23, MM14, McK76b, Mer17, Mis13, Moo78b, Nau94, Nel89b, Nel16e, Nel17a, Pra17, Pri99, SS83a, TT03, Wej83, Zac18]. Square-Arithmetic [Nel89b]. Square-Free [Abb93, Dav92]. Square-Sum [HKP15]. Squares [AM83a, BB14, Ber18, Bro01, Cos17, DeM12, DeM13, DeT10, Ger00, JY98, KS06, Lee24, LS79, MM14, Nel13, Nel14a, Nel14d, Nel24b, PV09, PR18, Pri18, Ste75, Tee93, VK80, Wag20, WR89, Wis99]. Squaring [Ben12a, CW09, Gor08]. Squeeze [BHS^{+}21, Ćur14]. SSD [Sch83b]. St. [Cea84, KL11, MZ24, Tre19]. Stability [Bar09, SR10]. Stable [BCD^{+}24, Fre08]. Stack [BC04]. Stacking [Jon93, Pfi91]. Staircase [DS12].

Stairway [Far08]. Stan [Bar81a]. Stand [RS00, Ska96]. Standard [Boo79, Fal81, Gor86, Mur84, PL81, SP79]. Star [Cag23, JV15, Lee98, Nak86, The23]. Star-Shaped [The23]. Stars [Som03, WC99]. Statement [Ano70l]. Statements [BR83b]. States [BR87, CH17, Ric05]. Station [Sim06]. Statistical [Che05, EW70, HSL77, Rob84a, RC99, VU21]. Statistician [Ive88]. Statisticians [Sne88]. Statistics [Ald88, Bai88, Bra00, Eve02, Far00, GK88, God88, Hah88, Hei78, KGSP88, Ker83, LRSZ14, Mos88, PM88, Qui17, RS16, Ste77, Swe22, Tan88, Wag83]. Statistics-Related [Swe22]. Steep [Pag72]. Steiner [Rou12]. Step [Lo 91]. Stephen [Dan85]. Sterling [HR24]. Steven [Nun10, Ric19a, Jon10]. Stewart [Hop18a]. Stick [CS86a, Hir24, Kri25]. Sticker [DLPS16]. Stieltjes [Gro19]. Stiffening [CH99]. Still [Alb96, Sch82c]. Stillwell [Wag11]. Sting [KS03]. Stirling [Ham20, HR24, Hir24, OB77, Wri95]. Stochastic [Bis80, Bur98, DR13, EG75]. Stoke [Sul96]. Stoppers [JW08]. Storage [LS01]. Story [HM23, Lev14]. Straight [FG19, Mab10]. Straight-In [Mab10]. Straight-line [FG19]. Straightedge [Hec03, Woo00]. Strange [FGL^{+}18]. Strategic [CHS86]. Strategies [AMM15, Bos00]. Strategy [CC19, Con99, Gal81b, Gar24, Pie71]. Straus [Ser15]. Straw [IK06, Iga08, JSSW06]. Street [RS23a]. Stretch [Ruc87]. Stretched [Ber14a]. Strictly [FM94]. String [Blå16b, Bur92, Che81, Pen86]. Strings [Sch12, Sch90c]. Strip [You24]. Strogatz [Nun10, Ric19a]. Strong [Tho12]. Stronger [BB97, FPS13]. Strongly [Sch90c]. Strophoid [Job11]. Structure [Oss81]. Struik [Tat02]. Student [Ano01-30, Ano04z, Ano06z, Ano07z, Ano09x, BCR^{+}11, CS17, EB99, Ger75, Hol81b, HSL77, Mac78, MS78, Nun10, Sch83b, Ste73, Tea06, WE73, Wil73, Som99]. Student-It [HSL77]. Student-Produced [BCR^{+}11]. Students [Boy81, CDK^{+}90, CDL^{+}01, Due96, LS94, LB97, Mer74, Mit72, Pól76a, Pól76b, Pól76c, SE96, Sha18, Ste81a, Ste77, Tre92, Wil72, Woo72b]. Study [CJLP93, Dav85, Fra92, Jon74, KM15, Liu01, Per73, Vel23]. Studying [KR98, Tre92]. Stuff [Fre95, SH81]. Stupid [BMRS13, Moo78a]. Style [CFLR12, Kat18, OB77]. Subadditive [FZ12]. Subgroup [OS16]. Subgroups [JB92, Pet11, PS15]. Subharmonic [Seg89]. Subliminal [Ros04]. Subnormals [BW13]. Subprime [BNUW23]. Subrings [Arp74]. Subscribers [Ano72e, Ano72f, Ano73g, Ano74h, Ano74i]. Subscription [Ano70m, Ano72j, Ano72k, Ano73n, Ano73o, Ano73p, Ano74p, Ano74q, Ano74r]. Subsequence [CL90]. Subset [Kre17]. Subsets [DR20]. Subsidiary [Bai88]. Subspaces [Lay93, Str93]. Substance [Oss81]. Substitution [Moo80, Vel23]. Substitutions [BR17]. Subtle [JLR11]. Subtleties [Bum84]. Subtracting [San19]. Subtraction [She22]. Success [WJ73]. Successful [Gre77]. Succession [MB82]. Successive [Sub04]. Sudoku [PV09, ALT10b, BCD^{+}24]. Sufficient [PS15]. Sugar [DL73]. Suggest [Ken99a]. Suggested [Hil80a, Hil80b, PM74]. Suggestion [LeD73, Sch70]. Suggestions [Sel70, Ste74]. Sum [Ahu79, Alb12, Bur95, DeM12, DeM13, DeT10, Eis17, Gor89, HKP15, Har88, HMF20, JV15, Kac19, Kal93, Kob17, Len81, MT23, Mat91, MT22, Mol16, Mur81, Nak86, NS87, Nel15c, OT72, PK83, Pla17b, Pri18, Pug80, Ric15b, Sch23, SP79, Shi10, Spi06, Tur81, Una09, Una10, Wan14b, Wu23b, Zia91, Sup84, Yiu12, AE19]. Sum-Difference [Shi10, Yiu12]. Sum-Guessing [Ahu79]. Summation [FN92, Lib81, McC90, Pri73a]. Summations [Kat09]. Summer [LS94, Tox70]. Summing [Bur00, Car87, LPP08, Nel14d, Nel24b, Sch92a, Smo24]. Sums [Alb02, BQ08, Bén05, Ber18, Bra01b, BGR86, Cag20b, CP89, CL90, Cra12, DK97, Den07, Edg22, Ekl00, FST23, Gor94, Gui95, Her12, Hir24, Hsu82, Kob13b, Kre97, Kri95, Kro83, Len81, Les04, LM20, Mar90a, MT22, Mur80a, Nel14a, Nel15d, PE82, Pfa11, Pfi88, Pla17a, Pon07, Pri73a, Sch86a, SS21, Str90a, Tru19, Wil24b, Wis99, Har87]. Sun [Bra01c, Dee99]. Sunshades [FN09]. Super [Log79, Som00]. Superbat [Bai87]. Superior [BKSU19]. Superposition [BR87]. Support [JW15, Mer74, Zee13]. Supported [Ash20]. Supposer [SY87]. Surface [CW90, HLL98, Kwo16, LL01, Seg82b, Str90b, The23, Ton97, Zam77]. Surfaces [Boy07, LH89, Tee93, Wol71]. Surfacing [Bei89]. Surpass [Ral84]. Surprise [DD95, ES85]. Surprises [BPP15, Boa79a]. Surprising [AK75, BAW87, Eve91, Nel16b, OHC07]. Survey [All03, Car75, Fle79, Fle80, Kle89, Lan72]. Surveyor [Bra86]. Survival [Lin75]. Surviving [GM99]. Survivor [IB22]. Suspense [KK13a]. Suspension [Gro10, HM05]. Sustainability [Zee13]. SVD [GJJ23, Kal96b, LL01]. Sweep [Sie24]. Sweeping [LR20]. Sweepstakes [Eck07]. Swimming [Zho18]. Swing [WRR98]. Swinging [Fre08]. Sylvester [Che22, Kun11]. Symbolic [HB92, She75]. Symmedian [Smi11]. Symmetric [BPR15, BSC17, Raz15]. Symmetry [Cri11, GK13, Gro05a, Nel95]. Symposium [Ped80b]. Synchronized [WM76]. Synchronizing [ZGN06]. Synthetic [Fre95, Kun80b, PC81]. System [AMR13, Den13, GEJ95, Kal02, LP16, LM90, Oma09, Phr90, Reb89, Rou12, Sch83b]. Systematic [Che18]. Systems [Bla91, Bla94, Bou23, Bur92, Dev91b, Edg21, Gug92, HM95, Mac92b, Nat91, Nel77, Oli73, Sch99a, SHL86, Wie84, Yer97, CG19, Kat18].

T [CBL19]. T-Shirts [CBL19]. Table [Fre08, Kno85, Pul88, Sch97b, Vau08]. Tables [DK08, Joh00, Jon12b, Jor77, Lon84, TL00]. Tabular [Gil91, Hor90b]. Taco [BW98]. Tactile [FLP18]. Tails [Hal13]. Takahashi [Col13]. Take [AR97, Gla93, Nas23, Str85]. Taking [Bet12, Cur05b, Hop10, KS03]. Tale [Kru00, OW02]. Talent [Daw97]. Tales [Gol13]. Talking [Kun72]. Tangent [Biv86, But03, FM05, Fri24, Gas23, Joh93a, Kob13a, Kra03, MS16, MP14, Ste01, Tro04, Wil03]. Tangential [Lau20a]. Tangentially [Dob84]. Tangents [Aar00, Gas94, Jav17, Mal15, Nel22b, Seg00, TR23]. Tank [Nie94]. Tanks [LS01]. Tanks-Exact [LS01]. Tape [WM76]. Tarr [Smu82a, Smu82b]. Tax [Fal85, HT84]. Taxes [Els82]. Taxicab [Jon24, Rat10, Wal95a]. Taylor [Kat18, Alb96, BSL09, Bes14, Bos95, Eul97, Gor93, Kal85b, Kos17, Kre10, MP10, O’L98]. Taylor-made [Bos95]. Tea [BPS12]. Teach [Bon08, God88, Grü81, Han88, Her90, KGSP88, Kun80b, PM88, Tan88]. Teacher [Ano71f, Fla11, Has71b, LeD73, Lin75, McK79, Nun10, Ano70e]. Teachers [Col71, Coo74, FM77, GK88, HA70, Lai70, Per73, Ste81a]. Teaches [Ree88]. Teaching [AB16, Ald88, BS15, Bea11, Ber75, Bla94, Boa79b, BN03, Byr11, Car93, Dav82d, DK01, DR13, DP80, Ell84, FLP18, Fos12, Gol11a, Gor24, Hal12, Hud11, KP86, LeC86, Lew73, Lut11, Mil74, Mit72, NO92, Per71, Ric12, RC99, Sch70, Shu02, Tho23, Thu00, Var11, Whi88, Wil91, Zho12a]. Team [Joh17, Kep85, Pag85]. Technical [RM71]. Technique [DL75, Gur19, Hug00, Nel77, Sch78b, She09]. Techniques [Jos83, Kun89]. Technology [BN03, Qui16, Qui17]. Telegraphic [Ano80j, Ano80k, Ano80l, Ano81l, Ano81m, Lin81a, Lin81b, Lin81c, Lin82a, Lin82b, Lin82c, Lin83a, Lin83b, Lin83c]. Telescope [Tee12, RT07]. Telescopes [Tee07]. Telescoping [Ben12b, Pri73a]. Temperature [DY04]. Ten [Cav96, May24, RC99]. Ten-Pound [May24]. Tendency [PM82a, PM83]. Tennis [Coo86]. Tension [Blo01]. Tent [Sri10]. Tenzi [BP20]. Term [Whi71]. Terminal [BQ13]. Terminally [Har96b]. Terms [MZ94, MT22, Ovc11, Ros70]. Ternary [MR96]. Test [CS17, Che13, Cho80b, Eis98b, Fos12, Gro16, Kat12, Lin78, Mor86, Str97b, Ven80, Wag77]. Testbank [BM77]. Testing [PR85, Tea06, Zah72]. Tests [FH21, Gan17, Ros12, Woo72b]. Tetrahedra [SH85]. Tetrahedral [Cag15]. Tetrahedron [Her04, HP99a, HP99b, Jac97, Mar98a]. Tetration [EK22]. TEXas [Kat18]. TEXas-Style [Kat18]. Text [Min91, Not71]. Textbook [Zah72]. Texts [BN03]. th [Hen94, MT22, Pri99, Smi87, CMS20, Whi71]. Thébault [PV13]. Their [Ara11, CS06, Hil92a, Lon81, Mar08, Pól76a, Pól76b, Pól76c, Sha18, Zas70, Gro05b, Law24, Sch85d]. Them [Cal95, Led16, Pól76a, Pól76b, Pól76c]. Theme [Mer18, Osl03]. Theorem [AN22, AB17, AL00, BBB03, BSL09, BHS^{+}21, Bel15, BSC17, DW19, Dun91, FF24, FG19, Gar22, Gin04, GG23, Gre74b, HF93, Hir98, Hoe97, Hwa12, JW22, JD98, Kar14, Kob14, Lea18, LM23, Mac95, Mic14, Moo95, Mye77, NP22, Pal98, Saa70, Say95, Sch74a, Sch85c, Sch88b, Sil89, Smi86, Ton15, Var77, Zob13, Abb10, AA19, AM02, AN12, Bar81b, Bas16, BZ04, BlC09, Cal95, Coo13, DH12a, EO17, Fra87, Gaz92, Gen89, Gen73, Gre11, Har86, Heo15, Hof82, How02, Kal85b, Kau20, Kre10, LP94, Lem16, Len15, McG18, Mol17, Mor10a, OP76, Pal17, PW21, PS03, Ros12, Sha25, Sub04, Sul96, Tou95, Tre18, Wal11, Whi85a, Wie75a, Wie75b]. Theorems [AN10, Boa81b, Bro15, Eve91, Leh22, Mil71, PL09, Rei73a, Tan17, WR89]. Theoretic [JB92]. Theorist [CCKL25]. Theory [AC22b, Bia98, Bla91, BSC17, Bre08, Flu88, IM11, Jep79, Joh72b, Kir96, LR20, Lin00b, Mac92b, MS23, Qui16, Sha18, Som00, Wag14, YD07]. There [DI08, Ols89, Rib97, Shu02]. Thin [Ber18]. Thing [KK12]. Things [Hop18b, Ken99b, Mer97]. Thinking [Bri84, Her03, Nie91]. Third [Nel15d].

Thirst [Ehr94]. Thorns [Hal92]. Thorns-Beyond [Hal92]. Those [Boa81b]. Thought [Bon16, Daw90, Hal85]. Thousand [HS90, Nor09, Tho98b]. Thread [Dav83]. Three [BNUW23, BZ04, Ber20, Bro00, Emm09, Fon96, Hal10, Ham85, Hol79, Kar14, Kat03, Lea16a, Man98, Mey86, Nel15c, San21, Sch86d, SH14, Str85, Wu23b]. Three-Dimensional [BZ04]. Three-Hour [Hol79]. Thrills [Hal82]. Throw [SH12]. Thyself [Ive88]. Tickets [Gar24]. Tie [CBL19]. Tie-Dyed [CBL19]. Ties [Pen88]. Tiling [CS95, Cos17, Gar09, Men04, Mie83]. Tilings [JM97, Pra17]. Time [Arm74, BR24, Dru95, Gla93, JWW16, KS88, May13, Sha95, Nym83]. Times [AAC04, Bum83, CL83]. Timetables [Ols13]. Timing [Tho92]. Tin [Roe93]. Tion [Rob90]. Tip [Bea11, Byr11, DR13, Fos12, Hal12, Hud11, Lut11, Ric12, Var11, Zho12a]. Tippy [You06]. Today [Alb81b, Ziz13]. Todd [Alb07a]. ToGa [MP03]. Together [JM97]. Tom [Alb97]. Tony [Col13]. Too [THMM24, Ell84]. Tool [Gor24, Lax76, McG90, Sne88]. Tools [BS15, FLP18]. Top [RC99]. Topic [Ano95c, Hoo97]. Topics [Hei78]. Topological [Cai73, CM14b]. Topology [FN09, Kre17]. Toronto [Ped80b]. Torricelli [BC18a, Gro93]. Torsion [Hoe09]. Torus [Mur95, WS81]. Toss [LP16]. Tosses [Blo98]. Tossing [Ber02, Cha79, Hof78b, Lip03, MP03, Nef82]. Total [MT19]. Touch [Alb04]. Tour [BMV06, Han20, Sie12, MS19]. Tournament [Sch92a, Wal87]. Tours [Bea12]. Tower [Bon21, DHR10, IB13]. Towers [Daw94]. Trace [KM13]. Track [Fin02, PV16]. Trades [Zho18]. Traditional [Mil74, Mor73, Wag14]. Trafalgar [Nas85]. Traffic [CHS87]. Tragedy [Mul19]. Trails [Bro00, Hal13]. Training [RP85, TA74]. Trammel [Sch11b]. trans [Hop19]. Transcendental [Cha07]. Transcendentals [Wey23]. Transfer [San77]. Transform [BC75, Gug92, NO92]. Transformation [Ano93i, Won00]. Transformations [AB16, AS88, But03, Por93, Swa73]. Transforms [Den07]. Transient [CH17]. Transitions [AC87]. translated [Col13]. Transom [Cra11b]. Trapezoid [Ren14b]. Trapezoidal [Bou24, Buc85b]. Traveler [Boa79a, Bar09]. Traveling [Bes14]. Treatise [Ser17]. Treatment [Che18, Gra11a, Moo23]. Trebuchet [Con17]. Tree [CCKL25, For17, Mar90b, Du20]. Trees [Bur98, Coo86, KB21, Spo23a]. Trends [Ste76b]. Trials [Fra87]. Triangle [ALT10a, BB97, BR00, BC16b, BC17, Ber20, BS04, BR89, CO99, CG24, Das11, DS08, Dun77, FPS13, Fos16, GI18, HB97, HP99a, HP99b, IMW12, Lan93b, Lar89, LM20, Mae19, Nea94, Nel16a, OB77, Pla17a, Pla17c, RS23a, San24, Saw17, Sch77, Sch84d, Oku22, Can89, Den07, Jon83, Lon84, Nef82, Osl03, Pla17b, Put86, Ros77, Smi73, Wri89, Wri95]. Triangles [AN12, Bas16, BS98, BS00, Bir10, CG24, CBL19, DeM12, Dr24b, Edg16, Edw14, Har08, Hyd88, JY98, JW22, Lon83, Lut84, Mie83, Nel14d, Nel14c, Par07, Par84, Roj23, Sas91, Sin90, Swa74, SY12, Zuc09, Hop19]. Triangular [BDDY13, Cag20a, Cag24c, DT13, Edg21, Jon12a, Kob13b, Lea16a, LS79, Nel15b, Nel15c, Nel15d, Nel16d, Nel16e, Nel19, Nel24b, Wat12].

Triangulations [DeT98]. Tribonacci [BNUW23]. Trick [Ehr21, Hos18b, PDR21, San06, She09, Swa05, Val15, Wag94]. Tricks [CC09, Nag21, Bon08]. Tricky [vD09]. Trig [BR17, Boa19, Mor91]. Trigonometric [Bal22, Bal23, Cho80c, Dra13, Ham07, IL24, Jos16, MT23, Nel15a, PS79, Pin74b, Ros89, Sch74a, Sch74b, ST19, Sil90, Wil24b]. Trigonometry [Ban04, Bus71, Ins93, Min91]. Trinomial [CO99]. Trip [MS23]. Triple [McC80c, Sch86f, Wit17]. Triples [BS96, Dun77, Eck92, For17, Hil85a, Kal86, Kau20, Kle72, MW03, Nel16e, RB21, Sch82a, Spo17, Spo23a, Tir86, WW00]. Triplet [vD09]. Triplets [For92, Hud11]. Triquetras [Mac92a]. Trisection [Bro07, Gau84, MR95, Mue07, Pet83]. Trisectrices [Eid93]. Trochoids [Far15, Hal92]. Tromino [Gar09]. trominoes [Cos17]. Trough [You06]. Truck [FIRW14, IK06, Ler21, Iga08]. True [SH81, Tan86b]. Trust [Sch97b]. Truth [HW04, Wag11]. Tucker [DF96, Mau83a, TS83]. Tugging [GS03]. Tuning [Har08]. Turn [CRS10, Rok14]. Turning [Fak99]. Turns [Hop10, Lei07]. Tutorial [Mor73, Wil72]. TV [TYM88]. Twenty [Lea18]. Twenty-First [Lea18]. Twice [Lut84, Sch91a]. Twist [smbV16]. Two [Ano70l, Ano76g, Ano77k, Ano77l, Ano77m, Ano77n, Ano77o, Ano78k, Ano78l, Ano78m, Ano78n, Ano92f, BB16, BB11, Bea09, BPR15, Bot78, Boy10, Bro15, Bur72, Car87, Cha92, Chi82, Col71, Ćur14, DP71, Dr24b, Edg23, Eid93, Eve91, Fak99, Fis75, Fle79, Fle80, Fon09, Fre11, Goo70, Gre77, Gri17, Hah98, Has71b, IK06, JW19, JS75, Kal84b, Kut81, Lai70, Law24, Leh22, LL71, Lin75, Lin76a, Lin76b, LM23, Liu09, Lyc23, Man70, Man78, McK79, Mue20, Nan14, Nat91, Nel15c, Nor17, OT72, OW02, Pag20, Pap17, Per73, Pla16, PS12, Pri73b, Sch84a, Sch13, Sch21a, Sem74, Str97b, Tam87, Tol22, Vau08, Wat89, Wet10, Oku22, MP88]. Two-Dimensional [JW19]. Two-for-One [Dr24b]. Two-Pan [Bur72]. Two-Person [BB11, Tam87]. Two-Sidedness [Pap17]. Two-Variable [Str97b]. Two-Year [Ano70l, Ano76g, Ano77k, Ano77l, Ano77m, Ano77n, Ano77o, Ano78k, Ano78l, Ano78m, Ano78n, Bot78, Col71, Fis75, Fle80, Goo70, Has71b, Lai70, LL71, Lin75, Lin76a, Lin76b, Man70, McK79, Pag20, Per73, Pri73b, Sem74, DP71, Fle79]. TYC [FA71, FMA72]. TYCMJ [Ano79e]. Tying [Lie97]. Type [Hos18a, Les04, HMT11]. Types [Slo77]. Typology [Tou13].

U} [Her93]. U.S.A. [Tur85b]. Ubiquitous [Adl96]. UFOs [NS17]. Ulam [Bar81a]. Ultimate [Con99]. Unbalanced [vR01]. Uncertainty [Wag83]. Unchained [Cox25]. Uncountably [Sur11]. Undecidable [Fla11]. Under-represented [Her10]. Underachieves [Mil74]. Underdetermined [Kal02]. Undergraduate [Boa71a, CDK^{+}90, Hoo97, SHL86, Tuc93]. Underground [Had13]. Undersampled [DRC98]. Understand [Hat83, Hub94]. Understanding [GECW13, HP93, JB13, JM97, Lei83, Pag79b, PR85]. Understands [Ger75]. Undetermined [De 10, Gol94, Gri71]. Unexpected [LD15, MS88]. Unfair [Rob84b]. Unfamiliar [Shp14]. Unfoldings [GV19]. Unicycle [Fin02, NR16]. Unified [Daw19, Lib81]. Uniform [BK85, MdMH19, Wis99]. Uniformly [Alb02, Gob88, TW90]. Unifying [BM97]. Unimodular [BS14b, Bri20, Won00]. Union [Fon09, Sch99b]. Unique [Lot23]. Uniqueness [Bir15, Sch81c]. Unit [GS11, Gol75, Jep86, Pra17, Tan15]. United [Ric05]. Units [WM76]. Unity [Hil84b]. Universal [Die81, JW08]. Universe [Ric19a]. University [Fei76, Gra18, Hop18a, Hop19, Nun10, Ric19a, PH77, Sin76]. Unknotting [CCHR24]. Unlawful [Bar22]. Unless [Thu82]. Unorthodox [Gür22]. Unpredictable [HS00]. Unusual [Kut81, Mau83b, Saw17, TT21]. Up-Down [HP14]. Up-to-Date [Guy84]. Update [CS87]. Upper [LM10]. Upside [KC87]. Upside-Down [KC87]. Uptake [Zob13]. Upward [You93]. USA [BC22, Baj23, BT24, BCT25, Pag85]. USAMO [Tur85a]. Use [Ald79b, CM14b, GK77, Har72, Hei78, Hic70, Kou95, Pin75, Rey93, Ros77, Sne88, Tym81]. Used [Kir96, LL71]. Useful [Gar93]. Uses [BM77]. Using [AMM15, BW19, Blå16b, Bou24, BGR86, Car72, CW23, Cra77, De 10, DW87, Doa89, DM03, Dra13, DF96, Edw14, Ekl01, Fra92, Gor84, Gre85, Hat11, HB92, HM95, Hoa88, Hoo97, HM05, JW22, Jur94, Kal84b, KD92, Kre97, Lar75, LeC86, LH89, LM90, Lop93, Lu22, Mas91, MR03, Moo91, Nam24, Nel24b, NO92, Per73, PL94, Sch99a, SB84, Str79, Suz99, SY12, TL00, Var76, Wil24b, Xu14, Zia91, smbH02, Bir10, Har10, JSW22]. Utmost [Str72].

V [Alb98, IMW12]. V-flex [IMW12]. Validation [EM17]. Valley [Tox70]. Valuable [Kal96b]. Value [AT00, All16, Boa81b, Cha07, Coa11, Hal80, JW22, JD98, MR98, Mye77, Nie21, OB77, Rob78, Say95, Sch85c, Sch88b, Shu79, Sil89, Smi86, Wag87, Wal11]. Values [BDLS18, Cic86, Dob85, DLM04, Lew73, Pon17, ZF97]. Vandermonde [Spi16]. Variable [Bra85, McG18, Mun04, Str97b, Woo70]. Variables [Alb02, Ano92f, Cha92, Eis17, Lu22, Nan14]. Variance [Rob85]. Variant [Gar22]. Variation [Gaz92, Kin74, Kit19, Mar13, PV13]. Variations [CF17, Dic06, Hos18b, KK08, Mer18, NP22, Osl03, Pau85]. Variety [BR00]. Varignon [NP22, Pal17]. Vary [Len15]. Vector [AC22a, Car98, DM03, Hub81, HM05, JLR11, Kob98, Kun80a, McC80c, SS82, Sch81d, Sch86f]. Vectors [Dea71, Ful00, GJJ23, JB13, Joh93a, KK93a, Wat85, HK19]. Vehicles [BJ24]. Velasco [Bra13]. Velocity [BQ13]. Venn [Grü84a, Grü84b, KH84]. Verhulst [Bra01a]. Verification [EM17]. Versed [Boa83]. Version [Abb24, Het97, Str97b]. Versions [SS99]. Versus [Bis80, FIRW14, Ler21, Kor85, Tho84, Wag84]. Vertex [JV15, Nak86]. Vertices [BD96a, BD96b, PS22]. Very [Bla91, Eis17, Eve02, HF93, Hop18a, Mor15]. Via [Nel21, Bel15, BMO^{+}00, Bur86a, Bur99, Bur00, Cat19, Cha99, DR06, Den07, ES23b, EH24, Eul93, FKY13, GG84, Har86, Hen91, Hof16, Hua25, Kal87, Knu18, Loe14, LL01, MM14, Mah20, McG18, Pfi88, Sch86d, Sch88b, Sch89b, Spi06, Spi16, Sue94, Tha18, Tou96]. Vibrating [Che81, Pen86].

Videocassettes [Wai74]. Viète [CFLR12, EM17]. View [BS96, FH94, Gla12, Ham07, Kob98, Moo89, Osl06, Par83, Sch06]. Viewed [BS02]. Viewpoint [Bla94, Fal05]. Vigor [Gal81a]. viii [Hop19]. Vinci [Lem16]. Visible [NL84]. Visibles [BZ05]. Vision [Bri92]. Visit [Zer02]. Visits [BDCH85]. Visual [AE19, Bri84, Cal95, EM17, Edg20, Edg21, Edg22, Gan81, MM21, Mon82, Par95, SR16, Ska91, SS21]. Visualization [Bri96, Gro19, MC24, Min92]. Visualizing [Bir10, MW97, Moo95, San21]. Viviani [Abb10, Zho12b]. Vocational [Ano70j]. Volleyball [KM15]. Voltaire [Sim11b]. Volume [Ami18, Ano15n, Ano16l, Ano17n, Ano18i, CW90, DH03, FN16, Lo 91, Mil78, Pet15, SH85, Str90b, Ton97]. Volume-Surface [Str90b]. Volumes [AK75, Car72, Knu18, Lam24, MMOM14, RR98, Win95]. Voting [Rat10, Str79]. vs [Bla91, Bor11, Hol23, Man74, Mil74, Mor73, Oss81, Rey93, Str90a, WWW83].

W [Bro06, Mau83a, Sri10]. Waclaw [Hop19]. Wada [FN07]. Waiter [Roj16]. Waiting [CRS10]. Walk [AR97, BR89]. Walking [BK11, BDCH85]. Walks [JR88]. Wander [MS16]. War [CM15]. Ware [DY04]. Warfare [WWW83]. Waring [Wil24b]. Was [Ano11b, Bre02, Sch04, Atk05]. Washington [Kat18, Coo70, HA70, PH77]. Wasp [KS03]. Waste [Eri22]. Water [IK06, Ill17, Sim09a, Boa86]. Wavefronts [Daw80]. Waves [Bes14, DRC98]. Way [Alb05b, Bas16, BMV14, But81, Chr83, Gar24, Hsu80, Mau85, MdMH19, Mos84, Nie91, Ols96, PP84, RS23a, Sch07, Sin12]. Ways [Bra16, DD10, Hof78b, JD90, Kal93, Mey86, Nor09]. Weak [AL15]. Weakened [Str97b]. Web [Qui17]. Wedding [Fle99]. Weierstrass [EO17]. Weighing [MG97]. Weighings [Bur72]. Weight [May24]. Weighted [Dub92, Hoe84, Hof16, JSW22]. Weights [Pin01]. Weird [Stu13]. Well [Alb81a, BC18b, Coc76, Moo86, Nie01, Ped80b, Pin74a]. Well-Ordering [Coc76, Pin74a]. Well-Rounded [Nie01]. Welsh [BPP15]. Western [Man70]. Wheel [Lei07, RS16]. Wheels [Bre08]. Where [Atk02, Hyd88, MS16, Ols89, Sch73, WN94]. Which [Niv72, Sch07]. Whiff [BSC17]. While [Nun10]. Whipper [Cur05b]. Whipper-The [Cur05b]. Who [Ano71f, Boa81b, CdSS17, Gar82, GN74, Hue92, Joy14, Kip70, NW98, Par13, Ree88, Ric15a, Tan88, VC10, Ano70e, Kip70]. Whodunits [She75]. Whom [Fre88, Ree88]. Whose [BD96a, BD96b, DH03, Han82]. Wiles [Bos95]. Will [Ger75, Ral84, RS00, Ska96, JMW24]. William [Gol13, Lea16b]. Wilson [Hop18a, AA90, Tre18]. Win [BP20, CM15, Eck07, Jon04]. Window [San02]. Winner [Ilt00, Str85]. Winner-Take-All [Str85]. Winners [Ano06o, Tur85a]. Winning [Bon16, BP12, IB22, KRSW18, Wil02a]. Wins [Joh17, Kep85, KM15]. Wisdom [Ric19a]. Wishes [Lan95]. Witch [GM99]. Without [AN10, AN12, ANU14, AN22, Ami18, AE17, Bal22, Bal23, Bar23, Ber18, Ber20, Bil18, CL15, Cag15, Cag20a, Cag20b, Cag24b, Cag24a, Cha13, Cia17, Coo24, Cra12, DeM13, DH12a, Dr24a, Edg16, Edw14, Ehr24, FZ12, Fos16, Fri24, Gas23, Heo15, Hu21, JV15, Jav17, Jon12a, Kob13a, Kob14, Kob17, KK13c, Kwo16, Lau20a, Lea16a, LS22d, Li13a, Li13b, MM14, MM21, Mol16, Mol17, MGC19, Muk24, Nel12, Nel13, Nel14a, Nel14b, Nel14d, Nel14c, Nel15e, Nel15b, Nel15c, Nel15d, Nel15a, Nel16f, Nel16c, Nel16g, Nel16d, Nel16a, Nel16b, Nel16e, Nel17a, Nel17b, Nel19, Nel21, Nel22a, Nel22b, Nel24a, Nel25, Pal17, PS22, PV13, Pla78, Pla14, Pla16, Pla17a, Pla17b, Pla17c, Pla18, Pri18, Qui22, Ren14b, Ren14a, Ric15b, Sch23, Sed80, SH14, Ser15]. Without [Sha15, Smi87, Sub15, Sun21, Una09, Una10, Una14b, Una14a, Wan14a, Wan14b, Woo25, Wu23a, Wu23b, Wu24, Zho18, smbH02, Aar00, BR99, Day13, Ins93, Kob13b, TL00, Xu14]. Wobegon [MS12]. Wolf [Sau90]. Woman [GM99, Mac78]. Women [Bru83, Cha77, May13]. Won’t [WN94]. Word [Hed00, Log79]. Words [AN10, AN12, ANU14, AN22, Ami18, AE17, Bal22, Bal23, BR99, Ber18, Ber20, Bil18, Cag15, Cag20a, Cag20b, Cag24b, Cag24a, Cha13, Cia17, Coo24, Cra12, DeM13, DH12a, Dr24a, Edg16, Edw14, Ehr24, FZ12, Fos16, Fri24, Gas23, Heo15, Hu21, JV15, Jav17, Jon12a, Kob13a, Kob13b, Kob14, Kob17, KK13c, Kwo16, Lau20a, Lea16a, Li13a, Li13b, MM14, MM21, Mol16, Mol17, MGC19, Muk24, Nel12, Nel13, Nel14a, Nel14b, Nel14d, Nel14c, Nel15e, Nel15b, Nel15c, Nel15d, Nel15a, Nel16f, Nel16c, Nel16g, Nel16d, Nel16a, Nel16b, Nel16e, Nel17a, Nel17b, Nel19, Nel21, Nel22a, Nel22b, Nel24a, Nel25, Pal17, PS22, PV13, Pla14, Pla16, Pla17a, Pla17b, Pla17c, Pla18, Pri18, Ren14b, Ren14a, Ric15b, ST98, Sch23, Sch95d, SH14, Ser15, Sha15, Sub15]. Words [Sun21, Tho98b, Una09, Una10, Una14b, Una14a, Wan14a, Wan14b, Woo25, Wu23a, Wu23b, Wu24, Zho18]. Work [Fre13, GM99, HP99a, Mer97, WS73]. Working [TA74]. works [Reb79]. Workshop [LS94, Ait16]. World [Ber14b, BW98, Fee01, Gra18, Gri17, Kep85, Lin19, Man74, Ric19a]. Worse [FE01, Yin87]. Worst [Ash97]. Worth [Mal17, Tho98b]. Would [Mal16, SS93]. Wrinkle [Fre03]. Write [Kha13]. Writer [Alb94]. Writing [Gol83a, GS90b, Pri89]. Written [Coh77b, GM99]. Wrong [Kro16].

X [Kat18, SS83b]. xiii [Kat18]. xiv [Kat18]. xv [Ser17]. xvi [Hop18a]. xviii [Hop18a]. xx [Gra18].

Yardstick [Bai86]. Year [Ano70l, Ano70j, Ano76g, Ano77k, Ano77l, Ano77m, Ano77n, Ano77o, Ano78k, Ano78l, Ano78m, Ano78n, Ano95c, Bot78, Col71, Fis75, Fle80, Goo70, Has71b, Lai70, LL71, Lin75, Lin76a, Lin76b, Man70, Man78, McK79, Pag20, Per73, Pri73b, Sem74, Woo70, DP71, Fle79]. Years [Gin04, Hop12, Kly20]. Yes [Cal00]. Yeshiva [Hop19]. York [Gut19, Hop19, Ric19a]. Yoruba [Zas70]. Youth [Hop14b]. Yuki [Col13].

Z [JR88]. Zag [LR24]. Zbikowski [CM14a]. Zeno [Kir96]. Zero [AS10, Hos18a, KC90, Lut11]. Zero-Divisors [AS10]. Zero-Over-Zero [Hos18a]. Zero-Row [KC90]. Zeroing [Hun98]. Zeros [Lon72, LH89, Mel22, PE82].

Zig [LR24]. Zig-Zag [LR24]. Zork [Sup84]. Zoser [Lo 91].

References

Acker:1978:CC

[AA78] Charles John Acker and Joe Frank Allison. Computer corner. Two-Year College Mathematics Journal, 9(5):303–308, November 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1978.11974597.

Albers:1990:CRW

[AA90] D. J. Albers and G. L. Alexanderson. A conversation with Robin Wilson. College Mathematics Journal, 21(3):178–195, May 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1990.11973305.

Albers:1991:CIN

[AA91] Donald J. Albers and G. L. Alexanderson. A conversation with Ivan Niven. College Mathematics Journal, 22(5):370–402, November 1991. CODEN ???? ISSN 0746-8342 (print), 1931-

1346 (electronic). URL http://www.tandfonline.com/doi/
abs/10.1080/07468342.1991.11973414.

Albers:1993:CRK

[AA93] Donald J. Albers and Gerald L. Alexanderson. A conversation with Richard K. Guy. College Mathematics Journal, 24(2): 122–148, March 1993. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1993.11973518.

Arney:1995:SR

[AA95] David C. Arney and Daniel J. Arney. Software reviews. College Mathematics Journal, 26(5):401–403, November 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1995.11973739.

Andreescu:2004:DER

[AA04] Titu Andreescu and Dorin Andrica. On a Diophantine equation and its ramifications. College Mathematics Journal, 35(1):15–21, January 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2004.11922045.

Adelstein:2019:RPT

[AA19] Ian M. Adelstein and George L. Ashline. Reframing the Pythagorean Theorem. College Mathematics Journal, 50(1):28– 35, January 2019. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2019.1557983.

Ash:2004:LMR

[AAC04] Garrett I. Ash, J. Marshall Ash, and Stefan Catoiu. Linearizing mile run times. College Mathematics Journal, 35(5):370– 374, November 2004. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2004.11922098.

Aarao:2000:TC

[Aar00] Jorge Aarão. Tangents without calculus. College Mathematics Journal, 31(5):406–407, November 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2000.
11974183.

Albers:1983:BR

[AB83] Donald J. Albers and R. P. Boas. Book reviews. Two-Year College Mathematics Journal, 14(3):266–267, June 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1983.11972696.

Arcavi:1991:RBP

[AB91] Abraham Arcavi and Maxim Bruckheimer. Reading Bombelli’s xpurgated algebra. College Mathematics Journal, 22(3):212–219, May 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1991.11973383.

Avital:1992:CP

[AB92] Shmuel Avital and Edward Barbeau. Commutativity of polynomials. College Mathematics Journal, 23(5):386–395, November 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1992.11973488.

Abel:2016:PIF

[AB16] Todd Abel and Jeremy Brazas. Pedagogically inconvenient functions for teaching transformations. College Mathematics Journal, 47(3):200–206, May 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.47.3.200.

Anderson:2017:FPA

[AB17] Stuart M. Anderson and Owen D. Byer. Finding polygonal areas with the corset theorem. College Mathematics Journal, 48 (3):171–178, May 2017. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.48.3.171.

Abbott:1993:RCS

[Abb93] H. L. Abbott. Remarks concerning “square-free sets on square grids”. College Mathematics Journal, 24(4):353–356, September 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1993.11973555.

Abboud:2010:VTE

[Abb10] Elias Abboud. Viviani’s Theorem and its extension. College Mathematics Journal, 41(3):203–211, May 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://

www.tandfonline.com/doi/abs/10.4169/074683410X488683.

Abbasi:2024:EVS

[Abb24] Sarmad Abbasi. An explicit version of Saidak’s argument for the infinitude of primes. College Mathematics Journal, 55(4): 349, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2023.2282928.

Aaron:2001:PPP

[ABCS01] Michael A. Aaron, Brewster B. Boyd, Jr., Melanie J. Curtis, and Paul M. Sommers. Punxsutawney’s phenomenal phorecaster. College Mathematics Journal, 32(1):26–29, January 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2001.11921847.

Abel:2015:IIP

[Abe15] Ulrich Abel. On an identity involving powers of binomial coefficients. College Mathematics Journal, 46(2):138, March 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.46.2.138.

Agnew:1987:T

[AC87] Jeanne L. Agnew and James R. Choike. Transitions. College Mathematics Journal, 18(2):124–133, March 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1987.
11973020.

Allaire:2000:AMA

[AC00] Patricia R. Allaire and Antonella Cupillari. Artemas Martin: an amateur mathematician of the Nineteenth Century and his contribution to mathematics. College Mathematics Journal, 31(1): 22–34, January 2000. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2000.11974104.

Aebi:2022:VIQ

[AC22a] Christian Aebi and Grant Cairns. A vector identity for quadrilaterals. College Mathematics Journal, 53(5):392–393, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2022.2118995.

Austin:2022:ECG

[AC22b] Jathan Austin and Emelie Curl. Exploring combinatorics and graph theory with simple Blokus. College Mathematics Journal, 53(4):273–281, 2022. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2022.2100147.

Angel:1976:LE

[ACF76] Allen Angel, Orrin Cocks, and Leonard Feldman. Letters to the Editor. Two-Year College Mathematics Journal, 7(2):1–2, May 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1976.11974422.

Aarao:2006:NSB

[ACJ^{+}06] J. Aarão, A. Cox, C. Jones, M. Martelli, and A. Westfahl. A non-smooth band around a non-convex region. College Mathematics Journal, 37(4):269–278, September 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2006.11922194.

Andrepont:1978:CDC

[AD78] Wayne Andrepont and Peter Dickinson. Classroom demonstration of a confidence interval. Two-Year College Mathematics Journal, 9(1):34–36, January 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1978.
11974543.

Adamson:1980:CPE

[Ada80] Iain T. Adamson. A combinatorial proof of Euler’s formula. Two-Year College Mathematics Journal, 11(4):272–273, September 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1980.11972457.

Adler:1996:SEU

[Adl96] David N. Adler. Some extensions of a ubiquitous geometric limit problem. College Mathematics Journal, 27(4):290–291, September 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1996.11973796.

Adnan:1998:CD

[Adn98] Saad M. Adnan. Candies and dollars. College Mathematics Journal, 29(5):414–415, November 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1998.
11973979.

An:2017:PWR

[AE17] Yajun An and Tom Edgar. Proof without words: Rearranged alternating harmonic series. College Mathematics Journal, 49 (1):35, January 2017. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2017.1389564.

An:2019:SVR

[AE19] Yajun An and Tom Edgar. “Sum” visual rearrangements of the alternating harmonic series. College Mathematics Journal, 50(4):280–285, 2019. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2019.1655375.

Agarwal:2021:ECC

[AFL^{+}21] Anurag Agarwal, David L. Farnsworth, Carl V. Lutzer, James E. Marengo, and J. A. Stephen Viggiano. Extremal correlation coefficients for bivariate probability distributions with specified marginal distributions. College Mathematics Journal, 52(1):45– 53, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2021.1845539.

Atkins:1987:FNC

[AG87] John Atkins and Robert Geist. Fibonacci numbers and computer algorithms. College Mathematics Journal, 18(4):328–336, September 1987. CODEN ???? ISSN 0746-8342 (print), 1931-

1346 (electronic). URL http://www.tandfonline.com/doi/
abs/10.1080/07468342.1987.11973055.

Armstrong:2014:ER

[AG14] Calvin Armstrong and Susan Goldstine. Exhibit review. College Mathematics Journal, 45(4):254–256, September 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
45.4.254.

Appel:1981:NPL

[AH81] Kenneth Appel and Wolfgang Haken. The nature of proof: Limits and opportunities. Two-Year College Mathematics Journal, 12(2):118–119, March 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1981.11972508.

Alexanderson:1982:BR

[AH82] G. L. Alexanderson and J. E. Householder. Book reviews. Two-Year College Mathematics Journal, 13(4):286–288, September 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1982.11972634.

Ahner:2009:SHI

[Ahn09] Henry F. Ahner. Sledge-hammer integration. College Mathematics Journal, 40(1):6–9, January 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2009.
11922328.

Ahuja:1979:SG

[Ahu79] Mangho Ahuja. On sum-guessing. Two-Year College Mathematics Journal, 10(2):95–99, March 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1979.
11972331.

Aieta:1987:PRR

[Aie87] Joseph F. Aieta. Powers and roots by recursion. College Mathematics Journal, 18(5):411–416, November 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1987.
11973067.

Aitken:2016:BCS

[Ait16] Neil Aitken. Babbage and Carroll in the Silent Workshop, 1867. College Mathematics Journal, 47(3):215, May 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
47.3.215.

Alexanderson:1975:SSV

[AK75] G. L. Alexanderson and L. F. Klosinski. Some surprising volumes of revolution. Two-Year College Mathematics Journal, 6 (3):13–15, September 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1975.11974385.

Austin:2023:CRH

[AKMMM23] Jathan Austin, Brian G. Kronenthal, Jonathon Miller, and Susanna Molitoris-Miller. Caught “Red”-Handed? The probability of randomly constructing a legal Catan board. College Mathematics Journal, 54(4):355–366, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2023.2237850.

Azad:2000:TC

[AL00] H. Azad and A. Laradji. On a theorem of Clay. College Mathematics Journal, 31(5):405–406, November 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2000.
11974182.

Anselone:2005:DEF

[AL05] Philip M. Anselone and John W. Lee. Differentiability of exponential functions. College Mathematics Journal, 36(5):388– 393, November 2005. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2005.11922154.

Acosta:2015:WCF

[AL15] Daniel Acosta and Terry Lawson. Weak contractions and fixed points. College Mathematics Journal, 46(1):35–41, January 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.46.1.35.

Albers:1979:E

[Alb79] Don Albers. Editorial. Two-Year College Mathematics Journal, 10(1):2, January 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1979.11972314.

Albers:1981:GAW

[Alb81a] Donald J. Albers. “geometry is alive and well and living in Paris under an assumed name”. Two-Year College Mathematics Journal, 12(4):226, September 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1981.11972535.

Albers:1981:PAP

[Alb81b] Donald J. Albers. Polite applause for a proof of one of the great conjectures of mathematics: What is a proof today? Two-Year College Mathematics Journal, 12(2):82, March 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972503.

Albers:1982:PHM

[Alb82] Donald J. Albers. Paul Halmos: Maverick mathologist. Two-Year College Mathematics Journal, 13(4):226–242, September 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1982.11972621.

Albers:1991:IIK

[Alb91] Donald J. Albers. Interview with Irving Kaplansky. College Mathematics Journal, 22(2):98–117, March 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1991.
11973366.

Albers:1994:FDM

[Alb94] Donald J. Albers. Freeman Dyson: Mathematician, physicist, and writer. College Mathematics Journal, 25(1):2–21, January 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.jstor.org/stable/2687079;
 http://www.tandfonline.com/doi/abs/10.1080/07468342.
1994.11973574.

Albers:1996:SQA

[Alb96] Donald J. Albers. Still questioning authority: an interview with Jean Taylor. College Mathematics Journal, 27(4):250– 266, September 1996. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1996.11973789.

Albers:1997:ITA

[Alb97] Donald J. Albers. An interview with Tom Apostol. College Mathematics Journal, 28(4):250–270, September 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1997.11973873.

Albers:1998:ILV

[Alb98] Donald J. Albers. An interview with Lars V. Ahlfors. College Mathematics Journal, 29(2):82–92, March 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/07468342.
1998.11973921.

Albert:2002:SUD

[Alb02] Jeanne Albert. Sums of uniformly distributed variables: a combinatorial approach. College Mathematics Journal, 33(3):201–206, May 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2002.11921941.

Albers:2004:TGI

[Alb04] Don Albers. In touch with God: an interview with Paul Halmos. College Mathematics Journal, 35(1):2–14, January 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2004.11922044.

Albers:2005:MGB

[Alb05a] Don Albers. “Mathematical games” and beyond: Part II of an interview with Martin Gardner. College Mathematics Journal, 36(4):301–314, September 2005. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2005.11922143.

Albers:2005:WMG

[Alb05b] Don Albers. On the way to “mathematical games”: Part I of an interview with Martin Gardner. College Mathematics Journal, 36(3):178–190, May 2005. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2005.11922126.

Albers:2007:JTN

[Alb07a] Don Albers. John Todd — numerical mathematics pioneer. College Mathematics Journal, 38(1):2–23, January 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2007.11922213.

Albright:2007:ISA

[Alb07b] Brian Albright. An introduction to simulated annealing. College Mathematics Journal, 38(1):37–42, January 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2007.
11922216.

Albright:2012:DSS

[Alb12] Brian Albright. The distribution of the sum of signed ranks. College Mathematics Journal, 43(3):232–236, May 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
43.3.232.

Albiston:2017:PAP

[Alb17] Jordie Albiston. A plane angle poem. College Mathematics Journal, 48(1):30, January 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.48.1.30.

Alder:1978:HBA

[Ald78] Henry L. Alder. A headquarters building for the association. Two-Year College Mathematics Journal, 9(4):195–202, September 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1978.11974577.

Alder:1979:EAA

[Ald79a] Henry L. Alder. Establishment of the archives of American mathematics. Two-Year College Mathematics Journal, 10(1):51–52, January 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1979.11972324.

Alder:1979:UGF

[Ald79b] Henry L. Alder. The use of generating functions to discover and prove partition identities. Two-Year College Mathematics Journal, 10(5):318–329, November 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1979.
11972377.

Alder:1981:HMG

[Ald81] Henry L. Alder. Harry Merrill Gehman 1898–1981. Two-Year College Mathematics Journal, 12(4):283–284, September 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972549.

Alder:1988:NGT

[Ald88] Henry L. Alder. The need for good teaching of statistics …. College Mathematics Journal, 19(1):20–21, January 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1988.11973081.

Aldous:2023:EES

[Ald23] David Aldous. Exploring endless space. College Mathematics Journal, 54(3):179–185, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2023.2201150.

Alexanderson:1979:CRa

[Ale79a] G. L. Alexanderson. Classics revisited. Two-Year College Mathematics Journal, 10(1):40–42, January 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1979.
11972321.

Alexanderson:1979:CRb

[Ale79b] G. L. Alexanderson. Classics revisited. Two-Year College Mathematics Journal, 10(2):119–122, March 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1979.
11972336.

Alexanderson:1979:CRc

[Ale79c] G. L. Alexanderson. Classics revisited. Two-Year College Mathematics Journal, 10(4):283–286, September 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1979.
11972370.

Alexanderson:1979:GPI

[Ale79d] G. L. Alexanderson. George pólya interviewed on his ninetieth birthday. Two-Year College Mathematics Journal, 10(1):13–19, January 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1979.11972316.

Alexanderson:1979:IMKa

[Ale79e] G. L. Alexanderson. An interview with Morris Kline: Part 1. Two-Year College Mathematics Journal, 10(3):172–178, June 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1979.11972346.

Alexanderson:1979:IMKb

[Ale79f] G. L. Alexanderson. An interview with Morris Kline: Part 2. Two-Year College Mathematics Journal, 10(4):259–264, September 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1979.11972361.

Alexanderson:1980:ICR

[Ale80] G. L. Alexanderson. An interview with Constance Reid. Two-Year College Mathematics Journal, 11(4):226–238, September 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1980.11972447.

Alexanderson:1981:IPE

[Ale81] G. L. Alexanderson. An interview with Paul Erdős. Two-Year College Mathematics Journal, 12(4):249–259, September 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972540.

Alexanderson:1983:BR

[Ale83a] G. L. Alexanderson. Book reviews. Two-Year College Mathematics Journal, 14(4):365–366, September 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1983.
11972719.

Alexanderson:1983:GPP

[Ale83b] G. L. Alexanderson. A glimpse at the Pólya picture album. Two-Year College Mathematics Journal, 14(4):274–294, September 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1983.11972700.

Alexanderson:1989:CSM

[Ale89] G. L. Alexanderson. A conversation with Saunders Mac Lane. College Mathematics Journal, 20(1):2–26, January 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1989.11973198.

Alexanderson:1992:CLB

[Ale92] G. L. Alexanderson. A conversation with Leon Bankoff. College Mathematics Journal, 23(2):98–117, March 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1992.
11973442.

Alexanderson:2007:BR

[Ale07] Gerald L. Alexanderson. Book review. College Mathematics Journal, 38(5):405–408, November 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2007.
11922266.

Alexanderson:2008:R

[Ale08] Gerald L. Alexanderson. Review. College Mathematics Journal, 39(5):419–421, November 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2008.11922323.

Alfaro:2015:SPL

[Alf15] Ricardo Alfaro. Sequences of power lines. College Mathematics Journal, 46(2):113–120, March 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
46.2.113.

Ali:2017:FE

[Ali17] Kazim Ali. Forgotten equations. College Mathematics Journal, 48(2):111, March 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.48.2.111.

Allinger:1977:CC

[All77] Glenn D. Allinger. Computer corner. Two-Year College Mathematics Journal, 8(1):47–50, January 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1977.
11974481.

Allen:2003:SOM

[All03] G. Donald Allen. A survey of online mathematics course basics. College Mathematics Journal, 34(4):270–279, September 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2003.11922017.

Allen:2016:BVP

[All16] Paul T. Allen. Boundary value problems and finite differences. College Mathematics Journal, 47(1):34–41, January 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.47.1.34.

Alperin:1987:HAF

[Alp87] Roger C. Alperin. Heron’s area formula. College Mathematics Journal, 18(2):137–138, March 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1987.
11973023.

Alperin:1989:MR

[Alp89] Roger C. Alperin. The matrix of a rotation. College Mathematics Journal, 20(3):230, May 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1989.11973235.

Applegate:2012:CAM

[ALS12] David Applegate, Marc LeBrun, and N. J. A. Sloane. Carryless arithmetic mod 10. College Mathematics Journal, 43(1):43–50, January 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.43.1.043.

Anggoro:2010:RT

[ALT10a] Alif Anggoro, Eddy Liu, and Angus Tulloch. The Rascal triangle. College Mathematics Journal, 41(5):393–395, November 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/074683410X521991.

Arnold:2010:GBR

[ALT10b] Elizabeth Arnold, Stephen Lucas, and Laura Taalman. Gröbner basis representations of Sudoku. College Mathematics Journal, 41(2):101–112, March 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/074683410X480203.

Amir-Moez:1983:IMS

[AM83a] Ali R. Amir-Moéz. Isomorphisms on magic squares. Two-Year College Mathematics Journal, 14(1):48–51, January 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1983.11972651.

Amir-Moez:1983:RPL

[AM83b] Ali R. Amir-Moéz. Roots of polynomials and loci. Two-Year College Mathematics Journal, 14(4):313–317, September 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1983.11972704.

Amir-Moez:1987:B

[AM87] Ali. R. Amir-Moéz. Behold! (1 × 2) + (2 × 3) + … + n × (n + 1) = (1/3)[(n + 1)^{3} − (n + 1)]. College Mathematics Journal, 18(4): 318, September 1987. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1987.11973051.

Apostol:2001:GRA

[AM01] Tom M. Apostol and Mamikon A. Mnatsakanian. Good rational approximations to logarithms. College Mathematics Journal, 32(3):172–179, May 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2001.11921872.

Aksoy:2002:MPD

[AM02] Asuman Aksoy and Mario Martelli. Mixed partial derivatives and Fubini’s Theorem. College Mathematics Journal, 33(2):126–130, March 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2002.11921930.

Agarwal:2010:LFR

[AM10] Anurag Agarwal and James Marengo. The locus of the focus of a rolling parabola. College Mathematics Journal, 41(2):129–133, March 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/074683410X480230.

Amiras:2018:PWV

[Ami18] Lucas Amiras. Proof without words: Volume of a pedestal prismoid. College Mathematics Journal, 49(2):92, March 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2018.1424423.

Ammon:2000:FFF

[Amm00] Theodore G. Ammon. Fallacies, flaws, and flimflam. College Mathematics Journal, 31(1):48–49, January 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2000.
11974107.

Austin:2015:SCU

[AMM15] Jathan Austin and Susanna Molitoris-Miller. The settlers of catan: Using settlement placement strategies in the probability classroom. College Mathematics Journal, 46(4):275–282, September 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.46.4.275.

Amir-Moez:1987:SPS

[AMP87] A. R. Amir-Moéz and Jean Pedersen. A simple proof of series convergence. College Mathematics Journal, 18(5):410, November 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1987.11973066.

Agarwal:2013:SLM

[AMR13] Anurag Agarwal, James E. Marengo, and Likin C. Simon Romero. System lifetimes, the memoryless property, Euler’s constant, and π. College Mathematics Journal, 44(3):210–219, May 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.44.3.210.

Alsina:2010:PWN

[AN10] Claudi Alsina and Roger B. Nelsen. Proof without words: New Pythagorean-like theorems. College Mathematics Journal, 41(5): 370, November 2010. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2010.11922451.

Alsina:2012:PWP

[AN12] Claudi Alsina and Roger B. Nelsen. Proof without words: The Pythagorean Theorem with equilateral triangles. College Mathematics Journal, 43(3):226, May 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
43.3.226.

Agarwal:2016:IRH

[AN16] Mahesh Agarwal and Narasimhamurthi Natarajan. Inheritance relations of hexagons and ellipses. College Mathematics Journal, 47(3):208–214, May 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.47.3.208.

Alsina:2022:PWA

[AN22] Claudi Alsina and Roger Nelsen. Proof without words: Another Pythagorean theorem. College Mathematics Journal, 53 (1):38, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2022.1997306.

Aguilera-Navarro:2001:RAP

[ANANFT01] Maria Cecília K. Aguilera-Navarro, Valdir C. Aguilera-Navarro, Ricardo C. Ferreira, and Neuza Teramon. Rational approximations to power expansions. College Mathematics Journal, 32 (4):276–278, September 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2001.11921890.

Anderson:1980:EDF

[And80] David H. Anderson. Elementary derivation of a formula for approximating n!. Two-Year College Mathematics Journal, 11(3): 201–202, June 1980. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1980.11972436.

Anderson:1981:ONE

[And81] David H. Anderson. Obtaining a numerical estimate for e. Two-Year College Mathematics Journal, 12(1):30–33, January 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972488.

Anderson:1984:SR

[And84] G. Ernest Anderson, Jr. Software reviews. College Mathematics Journal, 15(5):444, November 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1984.
11972833.

Andrews:1996:MEC

[And96] George E. Andrews. Mathematical education: a case for balance. College Mathematics Journal, 27(5):341–348, November 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1996.11973805.

Angel:1977:IAP

[Ang77] Allen R. Angel. An interesting approach to δ, ϵ proofs. Two-Year College Mathematics Journal, 8(5):278–280, November 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1977.11974528.

Anne:1998:EFI

[Ann98] Premchand Anne. Egyptian fractions and the inheritance problem. College Mathematics Journal, 29(4):296–300, September 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1998.11973958.

Anonymous:1970:CEa

[Ano70a] Anonymous. Coming events. Two-Year College Mathematics Journal, 1(1):2–3, Spring 1970. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1970.11973157.

Anonymous:1970:CEb

[Ano70b] Anonymous. Coming events. Two-Year College Mathematics Journal, 1(2):3–4, Autumn 1970. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1970.11973179.

Anonymous:1970:EBa

[Ano70c] Anonymous. Editorial board. Two-Year College Mathematics Journal, 1(1):??, Spring 1970. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1970.11973155.

Anonymous:1970:EBb

[Ano70d] Anonymous. Editorial board. Two-Year College Mathematics Journal, 1(2):??, Autumn 1970. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1970.11973177.

Anonymous:1970:MTW

[Ano70e] Anonymous. For the mathematics teacher who has almost everything. Two-Year College Mathematics Journal, 1(2):108, Autumn 1970. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1970.11973196.

Anonymous:1970:LE

[Ano70f] Anonymous. Letters to the Editor. Two-Year College Mathematics Journal, 1(1):1, Spring 1970. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1970.11973156.

Anonymous:1970:PC

[Ano70g] Anonymous. Publishers’ corner. Two-Year College Mathematics Journal, 1(1):91, Spring 1970. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1970.11973174.

Anonymous:1970:Ra

[Ano70h] Anonymous. Reviews. Two-Year College Mathematics Journal, 1(1):59, Spring 1970. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1970.11973170.

Anonymous:1970:Rb

[Ano70i] Anonymous. Reviews. Two-Year College Mathematics Journal, 1(2):59–75, Autumn 1970. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1970.11973190.

Anonymous:1970:SYV

[Ano70j] Anonymous. Second-year vocational. Two-Year College Mathematics Journal, 1(1):36, Spring 1970. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1970.
11973165.

Anonymous:1970:YM

[Ano70k] Anonymous. So you have a manuscript. Two-Year College Mathematics Journal, 1(2):104–105, Autumn 1970. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1970.
11973193.

Anonymous:1970:SMR

[Ano70l] Anonymous. Statement of the MAA’s role in The Two-Year College Mathematics Journal. Two-Year College Mathematics Journal, 1(1):87, Spring 1970. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1970.11973172.

Anonymous:1970:SI

[Ano70m] Anonymous. Subscription information. Two-Year College Mathematics Journal, 1(1):96, Spring 1970. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1970.
11973176.

Anonymous:1971:CEa

[Ano71a] Anonymous. Coming events. Two-Year College Mathematics Journal, 2(1):2–3, Spring 1971. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1971.11973989.

Anonymous:1971:CEb

[Ano71b] Anonymous. Coming events. Two-Year College Mathematics Journal, 2(2):9–11, Autumn 1971. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1971.11974006.

Anonymous:1971:CCB

[Ano71c] Anonymous. The CUPM Consultants Bureau. Two-Year College Mathematics Journal, 2(1):70, Spring 1971. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1971.
11974002.

Anonymous:1971:EBa

[Ano71d] Anonymous. Editorial board. Two-Year College Mathematics Journal, 2(1):??, Spring 1971. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1971.11973987.

Anonymous:1971:EBb

[Ano71e] Anonymous. Editorial board. Two-Year College Mathematics Journal, 2(2):??, Autumn 1971. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1971.11974004.

Anonymous:1971:MTW

[Ano71f] Anonymous. For the mathematics teacher who has ALMOST everything. Two-Year College Mathematics Journal, 2(2):1–8, Autumn 1971. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1971.11974005.

Anonymous:1971:Ra

[Ano71g] Anonymous. Reviews. Two-Year College Mathematics Journal, 2(1):65–69, Spring 1971. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1971.11974001.

Anonymous:1971:Rb

[Ano71h] Anonymous. Reviews. Two-Year College Mathematics Journal, 2(2):74–86, Autumn 1971. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1971.11974016.

Anonymous:1972:CEa

[Ano72a] Anonymous. Coming events. Two-Year College Mathematics Journal, 3(1):8–9, Spring 1972. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1972.11974188.

Anonymous:1972:CEb

[Ano72b] Anonymous. Coming events. Two-Year College Mathematics Journal, 3(2):70–71, Autumn 1972. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1972.
11974219.

Anonymous:1972:EBa

[Ano72c] Anonymous. Editorial board. Two-Year College Mathematics Journal, 3(1):??, Spring 1972. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1972.11974186.

Anonymous:1972:EBb

[Ano72d] Anonymous. Editorial board. Two-Year College Mathematics Journal, 3(2):??, Autumn 1972. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1972.11974204.

Anonymous:1972:ISa

[Ano72e] Anonymous. Institutional subscribers. Two-Year College Mathematics Journal, 3(1):1–7, Spring 1972. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1972.
11974187.

Anonymous:1972:ISb

[Ano72f] Anonymous. Institutional subscribers. Two-Year College Mathematics Journal, 3(2):1–6, Autumn 1972. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1972.
11974205.

Anonymous:1972:NN

[Ano72g] Anonymous. News and notes. Two-Year College Mathematics Journal, 3(2):53–55, Autumn 1972. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1972.
11974217.

Anonymous:1972:Ra

[Ano72h] Anonymous. Reviews. Two-Year College Mathematics Journal, 3(1):48–54, Spring 1972. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1972.11974197.

Anonymous:1972:Rb

[Ano72i] Anonymous. Reviews. Two-Year College Mathematics Journal, 3(2):42–52, Autumn 1972. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1972.11974216.

Anonymous:1972:SIa

[Ano72j] Anonymous. Subscription information. Two-Year College Mathematics Journal, 3(1):76, Spring 1972. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1972.
11974203.

Anonymous:1972:SIb

[Ano72k] Anonymous. Subscription information. Two-Year College Mathematics Journal, 3(2):82, Autumn 1972. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1972.
11974225.

Anonymous:1973:CEa

[Ano73a] Anonymous. Calendar of events. Two-Year College Mathematics Journal, 4(1):71–72, Winter 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1973.11974240.

Anonymous:1973:CEb

[Ano73b] Anonymous. Calendar of events. Two-Year College Mathematics Journal, 4(2):83–84, Spring 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1973.11974258.

Anonymous:1973:CEc

[Ano73c] Anonymous. Calendar of events. Two-Year College Mathematics Journal, 4(3):79, Autumn 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1973.11974277.

Anonymous:1973:EBa

[Ano73d] Anonymous. Editorial board. Two-Year College Mathematics Journal, 4(1):??, Winter 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1973.11974226.

Anonymous:1973:EBb

[Ano73e] Anonymous. Editorial board. Two-Year College Mathematics Journal, 4(2):??, Spring 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1973.11974245.

Anonymous:1973:EBc

[Ano73f] Anonymous. Editorial board. Two-Year College Mathematics Journal, 4(3):??, Autumn 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1973.11974264.

Anonymous:1973:IS

[Ano73g] Anonymous. Institutional subscribers. Two-Year College Mathematics Journal, 4(3):1–6, Autumn 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1973.
11974266.

Anonymous:1973:NNa

[Ano73h] Anonymous. News and notes. Two-Year College Mathematics Journal, 4(1):70, Winter 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1973.11974239.

Anonymous:1973:NNb

[Ano73i] Anonymous. News and notes. Two-Year College Mathematics Journal, 4(2):82, Spring 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1973.11974257.

Anonymous:1973:NNc

[Ano73j] Anonymous. News and notes. Two-Year College Mathematics Journal, 4(3):80, Autumn 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1973.11974278.

Anonymous:1973:PSa

[Ano73k] Anonymous. Problems and solutions. Two-Year College Mathematics Journal, 4(1):73, Winter 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1973.
11974241.

Anonymous:1973:PSb

[Ano73l] Anonymous. Problems and solutions. Two-Year College Mathematics Journal, 4(2):85, Spring 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1973.
11974259.

Anonymous:1973:R

[Ano73m] Anonymous. Reviews. Two-Year College Mathematics Journal, 4(2):75–81, Spring 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1973.11974256.

Anonymous:1973:SIa

[Ano73n] Anonymous. Subscription information. Two-Year College Mathematics Journal, 4(1):??, Winter 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1973.
11974227.

Anonymous:1973:SIb

[Ano73o] Anonymous. Subscription information. Two-Year College Mathematics Journal, 4(2):94, Spring 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1973.
11974263.

Anonymous:1973:SIc

[Ano73p] Anonymous. Subscription information. Two-Year College Mathematics Journal, 4(3):??, Autumn 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1973.
11974265.

Anonymous:1974:CEa

[Ano74a] Anonymous. Calendar of events. Two-Year College Mathematics Journal, 5(1):67, Winter 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1974.11974295.

Anonymous:1974:CEb

[Ano74b] Anonymous. Calendar of events. Two-Year College Mathematics Journal, 5(2):63–64, Spring 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1974.11974317.

Anonymous:1974:EBa

[Ano74c] Anonymous. Editorial board. Two-Year College Mathematics Journal, 5(1):??, Winter 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1974.11974283.

Anonymous:1974:EBb

[Ano74d] Anonymous. Editorial board. Two-Year College Mathematics Journal, 5(2):??, Spring 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1974.11974303.

Anonymous:1974:EBc

[Ano74e] Anonymous. Editorial board. Two-Year College Mathematics Journal, 5(3):??, September 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1974.11974322.

Anonymous:1974:EBd

[Ano74f] Anonymous. Editorial board. Two-Year College Mathematics Journal, 5(4):??, December 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1974.11974341.

Anonymous:1974:E

[Ano74g] Anonymous. Editorials. Two-Year College Mathematics Journal, 5(3):8, September 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1974.11974325.

Anonymous:1974:ISa

[Ano74h] Anonymous. Institutional subscribers. Two-Year College Mathematics Journal, 5(1):1–6, Winter 1974. CODEN ????

ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1974.
11974284.

Anonymous:1974:ISb

[Ano74i] Anonymous. Institutional subscribers. Two-Year College Mathematics Journal, 5(3):2–7, September 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1974.
11974324.

Anonymous:1974:NNa

[Ano74j] Anonymous. News and notes. Two-Year College Mathematics Journal, 5(1):68, Winter 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1974.11974296.

Anonymous:1974:NNb

[Ano74k] Anonymous. News and notes. Two-Year College Mathematics Journal, 5(2):60–62, Spring 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1974.11974316.

Anonymous:1974:NNc

[Ano74l] Anonymous. News and notices. Two-Year College Mathematics Journal, 5(3):46–49, September 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1974.
11974337.

Anonymous:1974:NNd

[Ano74m] Anonymous. News and notices. Two-Year College Mathematics Journal, 5(4):35–37, December 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1974.
11974350.

Anonymous:1974:PSa

[Ano74n] Anonymous. Problems and solutions. Two-Year College Mathematics Journal, 5(1):69–72, Winter 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1974.
11974297.

Anonymous:1974:PSb

[Ano74o] Anonymous. Problems and solutions. Two-Year College Mathematics Journal, 5(4):38–41, December 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1974.
11974351.

Anonymous:1974:SIa

[Ano74p] Anonymous. Subscription information. Two-Year College Mathematics Journal, 5(1):83, Winter 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1974.
11974302.

Anonymous:1974:SIb

[Ano74q] Anonymous. Subscription information. Two-Year College Mathematics Journal, 5(2):??, Spring 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1974.
11974304.

Anonymous:1974:SIc

[Ano74r] Anonymous. Subscription information. Two-Year College Mathematics Journal, 5(3):1, September 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1974.
11974323.

Anonymous:1975:EBa

[Ano75a] Anonymous. Editorial board. Two-Year College Mathematics Journal, 6(1):??, February 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1975.11974355.

Anonymous:1975:EBb

[Ano75b] Anonymous. Editorial board. Two-Year College Mathematics Journal, 6(2):??, May 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1975.11974367.

Anonymous:1975:EBc

[Ano75c] Anonymous. Editorial board. Two-Year College Mathematics Journal, 6(3):??, September 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1975.11974381.

Anonymous:1975:EBd

[Ano75d] Anonymous. Editorial board. Two-Year College Mathematics Journal, 6(4):??, December 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1975.11974395.

Anonymous:1975:E

[Ano75e] Anonymous. Editorials. Two-Year College Mathematics Journal, 6(1):1, February 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1975.11974356.

Anonymous:1975:NNb

[Ano75f] Anonymous. News & notices. Two-Year College Mathematics Journal, 6(2):29–30, May 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1975.11974376.

Anonymous:1975:NNa

[Ano75g] Anonymous. News and notices. Two-Year College Mathematics Journal, 6(1):30–31, February 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1975.
11974364.

Anonymous:1975:NNc

[Ano75h] Anonymous. News and notices. Two-Year College Mathematics Journal, 6(3):31–33, September 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1975.
11974390.

Anonymous:1975:PSa

[Ano75i] Anonymous. Problems and solutions. Two-Year College Mathematics Journal, 6(1):32–34, February 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1975.
11974365.

Anonymous:1975:PSb

[Ano75j] Anonymous. Problems and solutions. Two-Year College Mathematics Journal, 6(2):31–35, May 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1975.
11974377.

Anonymous:1975:PSc

[Ano75k] Anonymous. Problems and solutions. Two-Year College Mathematics Journal, 6(3):34–37, September 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1975.
11974391.

Anonymous:1975:PSd

[Ano75l] Anonymous. Problems and solutions. Two-Year College Mathematics Journal, 6(4):24–28, December 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1975.
11974403.

Anonymous:1976:EBa

[Ano76a] Anonymous. Editorial board. Two-Year College Mathematics Journal, 7(1):??, February 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1976.11974407.

Anonymous:1976:EBb

[Ano76b] Anonymous. Editorial board. Two-Year College Mathematics Journal, 7(2):??, May 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1976.11974421.

Anonymous:1976:EBc

[Ano76c] Anonymous. Editorial board. Two-Year College Mathematics Journal, 7(3):??, September 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1976.11974434.

Anonymous:1976:EBd

[Ano76d] Anonymous. Editorial board. Two-Year College Mathematics Journal, 7(4):??, December 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1976.11974452.

Anonymous:1976:PSa

[Ano76e] Anonymous. Problems and solutions. Two-Year College Mathematics Journal, 7(3):47–50, September 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1976.
11974447.

Anonymous:1976:PSb

[Ano76f] Anonymous. Problems and solutions. Two-Year College Mathematics Journal, 7(4):33–37, December 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1976.
11974461.

Anonymous:1976:TYC

[Ano76g] Anonymous. Two-year College news. Two-Year College Mathematics Journal, 7(4):29–32, December 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1976.
11974460.

Anonymous:1977:EBa

[Ano77a] Anonymous. Editorial board. Two-Year College Mathematics Journal, 8(1):??, January 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1977.11974468.

Anonymous:1977:EBb

[Ano77b] Anonymous. Editorial board. Two-Year College Mathematics Journal, 8(2):??, March 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1977.11974483.

Anonymous:1977:EBc

[Ano77c] Anonymous. Editorial board. Two-Year College Mathematics Journal, 8(3):??, June 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1977.11974494.

Anonymous:1977:EBd

[Ano77d] Anonymous. Editorial board. Two-Year College Mathematics Journal, 8(4):??, September 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1977.11974508.

Anonymous:1977:EBe

[Ano77e] Anonymous. Editorial board. Two-Year College Mathematics Journal, 8(5):??, November 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1977.11974522.

Anonymous:1977:PSa

[Ano77f] Anonymous. Problems and solutions. Two-Year College Mathematics Journal, 8(1):42–46, January 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1977.
11974480.

Anonymous:1977:PSb

[Ano77g] Anonymous. Problems and solutions. Two-Year College Mathematics Journal, 8(2):95–100, March 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1977.
11974490.

Anonymous:1977:PSc

[Ano77h] Anonymous. Problems and solutions. Two-Year College Mathematics Journal, 8(3):177–181, June 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1977.
11974505.

Anonymous:1977:PSd

[Ano77i] Anonymous. Problems and solutions. Two-Year College Mathematics Journal, 8(4):240–243, September 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1977.
11974518.

Anonymous:1977:PSe

[Ano77j] Anonymous. Problems and solutions. Two-Year College Mathematics Journal, 8(5):292–295, November 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1977.
11974531.

Anonymous:1977:TYCa

[Ano77k] Anonymous. Two-year College news. Two-Year College Mathematics Journal, 8(1):38–41, January 1977. CODEN ????

ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1977.
11974479.

Anonymous:1977:TYCb

[Ano77l] Anonymous. Two-year College news. Two-Year College Mathematics Journal, 8(2):90–94, March 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1977.
11974489.

Anonymous:1977:TYCc

[Ano77m] Anonymous. Two-year College news. Two-Year College Mathematics Journal, 8(3):168–176, June 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1977.
11974504.

Anonymous:1977:TYCd

[Ano77n] Anonymous. Two-year College news. Two-Year College Mathematics Journal, 8(4):236–239, September 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1977.
11974517.

Anonymous:1977:TYCe

[Ano77o] Anonymous. Two-year College news. Two-Year College Mathematics Journal, 8(5):286–291, November 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1977.
11974530.

Anonymous:1978:EBa

[Ano78a] Anonymous. Editorial board. Two-Year College Mathematics Journal, 9(1):??, January 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1978.11974534.

Anonymous:1978:EBb

[Ano78b] Anonymous. Editorial board. Two-Year College Mathematics Journal, 9(2):??, March 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1978.11974548.

Anonymous:1978:EBc

[Ano78c] Anonymous. Editorial board. Two-Year College Mathematics Journal, 9(3):??, June 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1978.11974560.

Anonymous:1978:EBd

[Ano78d] Anonymous. Editorial board. Two-Year College Mathematics Journal, 9(4):??, September 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1978.11974576.

Anonymous:1978:EBe

[Ano78e] Anonymous. Editorial board. Two-Year College Mathematics Journal, 9(5):??, November 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1978.11974588.

Anonymous:1978:PSa

[Ano78f] Anonymous. Problems and solutions. Two-Year College Mathematics Journal, 9(1):40–45, January 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1978.
11974545.

Anonymous:1978:PSb

[Ano78g] Anonymous. Problems and solutions. Two-Year College Mathematics Journal, 9(2):95–100, March 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1978.
11974555.

Anonymous:1978:PSc

[Ano78h] Anonymous. Problems and solutions. Two-Year College Mathematics Journal, 9(3):176–181, June 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1978.
11974572.

Anonymous:1978:PSd

[Ano78i] Anonymous. Problems and solutions. Two-Year College Mathematics Journal, 9(4):236–242, September 1978. CODEN ????

ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1978.
11974585.

Anonymous:1978:PSe

[Ano78j] Anonymous. Problems and solutions. Two-Year College Mathematics Journal, 9(5):297–302, November 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1978.
11974596.

Anonymous:1978:TYCa

[Ano78k] Anonymous. Two-year College news. Two-Year College Mathematics Journal, 9(2):90–94, March 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1978.
11974554.

Anonymous:1978:TYCb

[Ano78l] Anonymous. Two-year College news. Two-Year College Mathematics Journal, 9(3):171–175, June 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1978.
11974571.

Anonymous:1978:TYCc

[Ano78m] Anonymous. Two-year College news. Two-Year College Mathematics Journal, 9(4):228–235, September 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1978.
11974584.

Anonymous:1978:TYCd

[Ano78n] Anonymous. Two-year College news. Two-Year College Mathematics Journal, 9(5):294–296, November 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1978.
11974595.

Anonymous:1979:EBa

[Ano79a] Anonymous. Editorial board. Two-Year College Mathematics Journal, 10(1):??, January 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1979.11972313.

Anonymous:1979:EBb

[Ano79b] Anonymous. Editorial board. Two-Year College Mathematics Journal, 10(2):??, March 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1979.11972328.

Anonymous:1979:EBc

[Ano79c] Anonymous. Editorial board. Two-Year College Mathematics Journal, 10(3):??, June 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1979.11972343.

Anonymous:1979:EBd

[Ano79d] Anonymous. Editorial board. Two-Year College Mathematics Journal, 10(5):??, November 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1979.11972374.

Anonymous:1979:GPT

[Ano79e] Anonymous. Guidelines for prospective TYCMJ authors. Two-Year College Mathematics Journal, 10(1):57–58, January 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1979.11972326.

Anonymous:1979:LLBa

[Ano79f] Anonymous. Letters and late-breaking news. Two-Year College Mathematics Journal, 10(1):59–60, January 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1979.
11972327.

Anonymous:1979:LLBb

[Ano79g] Anonymous. Letters and late-breaking news. Two-Year College Mathematics Journal, 10(2):135, March 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1979.
11972340.

Anonymous:1979:LLBc

[Ano79h] Anonymous. Letters and late-breaking news. Two-Year College Mathematics Journal, 10(3):221–222, June 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1979.
11972356.

Anonymous:1979:P

[Ano79i] Anonymous. Problems. Two-Year College Mathematics Journal, 10(1):52–57, January 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1979.11972325.

Anonymous:1980:CCa

[Ano80a] Anonymous. Computers & calculators. Two-Year College Mathematics Journal, 11(2):110, March 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1980.
11972415.

Anonymous:1980:CCb

[Ano80b] Anonymous. Computers & calculators. Two-Year College Mathematics Journal, 11(3):188, June 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1980.
11972433.

Anonymous:1980:CCc

[Ano80c] Anonymous. Computers and calculators. Two-Year College Mathematics Journal, 11(4):251, September 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1980.
11972450.

Anonymous:1980:EBa

[Ano80d] Anonymous. Editorial board. Two-Year College Mathematics Journal, 11(1):??, January 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1980.11972394.

Anonymous:1980:EBb

[Ano80e] Anonymous. Editorial board. Two-Year College Mathematics Journal, 11(2):??, March 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1980.11972409.

Anonymous:1980:EBc

[Ano80f] Anonymous. Editorial board. Two-Year College Mathematics Journal, 11(3):??, June 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1980.11972430.

Anonymous:1980:EBd

[Ano80g] Anonymous. Editorial board. Two-Year College Mathematics Journal, 11(4):??, September 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1980.
11972446.

Anonymous:1980:EBe

[Ano80h] Anonymous. Editorial board. Two-Year College Mathematics Journal, 11(5):??, November 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1980.11972463.

Anonymous:1980:LLB

[Ano80i] Anonymous. Letters and late-breaking news. Two-Year College Mathematics Journal, 11(2):146–149, March 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1980.
11972429.

Anonymous:1980:TRa

[Ano80j] Anonymous. Telegraphic reviews. Two-Year College Mathematics Journal, 11(1):59–60, January 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1980.
11972406.

Anonymous:1980:TRb

[Ano80k] Anonymous. Telegraphic reviews. Two-Year College Mathematics Journal, 11(2):128–130, March 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1980.
11972425.

Anonymous:1980:TRc

[Ano80l] Anonymous. Telegraphic reviews. Two-Year College Mathematics Journal, 11(4):281–282, September 1980. CODEN ????

ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1980.
11972461.

Anonymous:1981:CFMa

[Ano81a] Anonymous. Calendar of future meetings. Two-Year College Mathematics Journal, 12(4):273–274, September 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972546.

Anonymous:1981:CFMb

[Ano81b] Anonymous. Calendar of future meetings. Two-Year College Mathematics Journal, 12(5):347–348, November 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972568.

Anonymous:1981:CCa

[Ano81c] Anonymous. Computers & calculators. Two-Year College Mathematics Journal, 12(1):23, January 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1981.
11972486.

Anonymous:1981:CCb

[Ano81d] Anonymous. Computers & calculators. Two-Year College Mathematics Journal, 12(5):301, November 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1981.
11972554.

Anonymous:1981:EBa

[Ano81e] Anonymous. Editorial board. Two-Year College Mathematics Journal, 12(1):??, January 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1981.11972480.

Anonymous:1981:EBb

[Ano81f] Anonymous. Editorial board. Two-Year College Mathematics Journal, 12(2):??, March 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1981.11972502.

Anonymous:1981:EBc

[Ano81g] Anonymous. Editorial board. Two-Year College Mathematics Journal, 12(3):??, June 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1981.11972519.

Anonymous:1981:EBd

[Ano81h] Anonymous. Editorial board. Two-Year College Mathematics Journal, 12(4):??, September 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1981.
11972534.

Anonymous:1981:EBe

[Ano81i] Anonymous. Editorial board. Two-Year College Mathematics Journal, 12(5):??, November 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1981.11972552.

Anonymous:1981:EE

[Ano81j] Anonymous. Euclid’s Elements. Two-Year College Mathematics Journal, 12(2):117–133, March 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1981.
11972507.

Anonymous:1981:SPM

[Ano81k] Anonymous. Sidney Penner — in memoriam. Two-Year College Mathematics Journal, 12(1):69, January 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/00494925.
1981.11972501.

Anonymous:1981:TRa

[Ano81l] Anonymous. Telegraphic reviews. Two-Year College Mathematics Journal, 12(1):61–62, January 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1981.
11972498.

Anonymous:1981:TRb

[Ano81m] Anonymous. Telegraphic reviews. Two-Year College Mathematics Journal, 12(2):142–144, March 1981. CODEN ????

ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1981.
11972512.

Anonymous:1982:EBa

[Ano82a] Anonymous. Editorial board. Two-Year College Mathematics Journal, 13(1):??, January 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1982.11972570.

Anonymous:1982:EBb

[Ano82b] Anonymous. Editorial board. Two-Year College Mathematics Journal, 13(2):??, March 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1982.11972587.

Anonymous:1982:EBc

[Ano82c] Anonymous. Editorial board. Two-Year College Mathematics Journal, 13(3):??, June 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1982.11972602.

Anonymous:1982:EBd

[Ano82d] Anonymous. Editorial board. Two-Year College Mathematics Journal, 13(4):??, September 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1982.
11972620.

Anonymous:1982:EBe

[Ano82e] Anonymous. Editorial board. Two-Year College Mathematics Journal, 13(5):??, November 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1982.11972635.

Anonymous:1983:EBa

[Ano83a] Anonymous. Editorial board. Two-Year College Mathematics Journal, 14(1):??, January 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1983.11972645.

Anonymous:1983:EBb

[Ano83b] Anonymous. Editorial board. Two-Year College Mathematics Journal, 14(2):??, March 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1983.11972663.

Anonymous:1983:EBc

[Ano83c] Anonymous. Editorial board. Two-Year College Mathematics Journal, 14(3):??, June 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1983.11972680.

Anonymous:1983:EBd

[Ano83d] Anonymous. Editorial board. Two-Year College Mathematics Journal, 14(4):??, September 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1983.
11972699.

Anonymous:1983:EBe

[Ano83e] Anonymous. Editorial board. Two-Year College Mathematics Journal, 14(5):??, November 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1983.11972721.

Anonymous:1983:Pa

[Ano83f] Anonymous. Problems. Two-Year College Mathematics Journal, 14(2):172–178, March 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1983.11972678.

Anonymous:1983:Pb

[Ano83g] Anonymous. Problems. Two-Year College Mathematics Journal, 14(3):260–265, June 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1983.11972695.

Anonymous:1983:Pc

[Ano83h] Anonymous. Problems. Two-Year College Mathematics Journal, 14(4):352–359, September 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1983.11972716.

Anonymous:1983:Pd

[Ano83i] Anonymous. Problems. Two-Year College Mathematics Journal, 14(5):438, November 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1983.11972737.

Anonymous:1983:SNCa

[Ano83j] Anonymous. The signpost: News & commentary. Two-Year College Mathematics Journal, 14(3):268–269, June 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1983.11972697.

Anonymous:1983:SNCb

[Ano83k] Anonymous. The signpost: News & commentary. Two-Year College Mathematics Journal, 14(4):366–367, September 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1983.11972720.

Anonymous:1984:A

[Ano84a] Anonymous. Acknowledgements. College Mathematics Journal, 15(5):460, November 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1984.11972836.

Anonymous:1984:ARR

[Ano84b] Anonymous. Anthony Ralston’s response. College Mathematics Journal, 15(5):380–382, November 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1984.
11972810.

Anonymous:1984:EBa

[Ano84c] Anonymous. Editorial board. College Mathematics Journal, 15 (1):??, January 1984. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1984.11972738.

Anonymous:1984:EBb

[Ano84d] Anonymous. Editorial board. College Mathematics Journal, 15 (2):??, March 1984. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1984.11972757.

Anonymous:1984:EBc

[Ano84e] Anonymous. Editorial board. College Mathematics Journal, 15(3):??, June 1984. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1984.11972772.

Anonymous:1984:EBd

[Ano84f] Anonymous. Editorial board. College Mathematics Journal, 15 (4):??, September 1984. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1984.11972786.

Anonymous:1984:EBe

[Ano84g] Anonymous. Editorial board. College Mathematics Journal, 15 (5):??, November 1984. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1984.11972801.

Anonymous:1984:FRR

[Ano84h] Anonymous. Fred Roberts’ response. College Mathematics Journal, 15(5):397–399, November 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1984.
11972820.

Anonymous:1984:GA

[Ano84i] Anonymous. Guidelines for authors. College Mathematics Journal, 15(1):75, January 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1984.11972755.

Anonymous:1984:Pa

[Ano84j] Anonymous. Problems. College Mathematics Journal, 15(1): 68–74, January 1984. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1984.11972754.

Anonymous:1984:Pb

[Ano84k] Anonymous. Problems. College Mathematics Journal, 15(2): 163–169, March 1984. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1984.11972770.

Anonymous:1984:Pc

[Ano84l] Anonymous. Problems. College Mathematics Journal, 15(3):266– 273, June 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972784.

Anonymous:1984:Pd

[Ano84m] Anonymous. Problems. College Mathematics Journal, 15(4):346– 353, September 1984. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1984.11972799.

Anonymous:1984:Pe

[Ano84n] Anonymous. Problems. College Mathematics Journal, 15(5):445– 451, November 1984. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1984.11972834.

Anonymous:1984:SRa

[Ano84o] Anonymous. Software reviews. College Mathematics Journal, 15(1):65–67, January 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1984.11972753.

Anonymous:1984:SRb

[Ano84p] Anonymous. Software reviews. College Mathematics Journal, 15 (3):263–265, June 1984. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1984.11972783.

Anonymous:1985:A

[Ano85a] Anonymous. Acknowledgements. College Mathematics Journal, 16(5):434, November 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1985.11972923.

Anonymous:1985:EBa

[Ano85b] Anonymous. Editorial board. College Mathematics Journal, 16 (1):??, January 1985. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1985.11972837.

Anonymous:1985:EBb

[Ano85c] Anonymous. Editorial board. College Mathematics Journal, 16 (2):??, March 1985. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1985.11972860.

Anonymous:1985:EBc

[Ano85d] Anonymous. Editorial board. College Mathematics Journal, 16(3):??, June 1985. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1985.11972876.

Anonymous:1985:EBd

[Ano85e] Anonymous. Editorial board. College Mathematics Journal, 16 (4):??, September 1985. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1985.11972889.

Anonymous:1985:EBe

[Ano85f] Anonymous. Editorial board. College Mathematics Journal, 16 (5):??, November 1985. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1985.11972905.

Anonymous:1985:GA

[Ano85g] Anonymous. Guidelines for authors. College Mathematics Journal, 16(1):61, January 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1985.11972855.

Anonymous:1985:Pa

[Ano85h] Anonymous. Problems. College Mathematics Journal, 16(1): 74–80, January 1985. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1985.11972858.

Anonymous:1985:Pb

[Ano85i] Anonymous. Problems. College Mathematics Journal, 16(2): 153–161, March 1985. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1985.11972874.

Anonymous:1985:Pc

[Ano85j] Anonymous. Problems. College Mathematics Journal, 16(3):223– 231, June 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1985.11972887.

Anonymous:1985:Pd

[Ano85k] Anonymous. Problems. College Mathematics Journal, 16(4):304– 311, September 1985. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1985.11972903.

Anonymous:1985:Pe

[Ano85l] Anonymous. Problems. College Mathematics Journal, 16(5):416– 424, November 1985. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1985.11972920.

Anonymous:1986:EBa

[Ano86a] Anonymous. Editorial board. College Mathematics Journal, 17 (1):??, January 1986. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1986.11972924.

Anonymous:1986:EBb

[Ano86b] Anonymous. Editorial board. College Mathematics Journal, 17 (2):??, March 1986. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1986.11972942.

Anonymous:1986:EBc

[Ano86c] Anonymous. Editorial board. College Mathematics Journal, 17(3):??, May 1986. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1986.11972957.

Anonymous:1986:EBd

[Ano86d] Anonymous. Editorial board. College Mathematics Journal, 17 (4):??, September 1986. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1986.11972970.

Anonymous:1986:EBe

[Ano86e] Anonymous. Editorial board. College Mathematics Journal, 17 (5):??, November 1986. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1986.11972984.

Anonymous:1986:GA

[Ano86f] Anonymous. Guidelines for authors. College Mathematics Journal, 17(1):110, January 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1986.11972941.

Anonymous:1986:MH

[Ano86g] Anonymous. Media highlights. College Mathematics Journal, 17(5):450–456, November 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1986.11972998.

Anonymous:1986:Pa

[Ano86h] Anonymous. Problems. College Mathematics Journal, 17(1): 92–99, January 1986. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1986.11972939.

Anonymous:1986:Pb

[Ano86i] Anonymous. Problems. College Mathematics Journal, 17(2): 184–189, March 1986. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1986.11972955.

Anonymous:1986:Pc

[Ano86j] Anonymous. Problems. College Mathematics Journal, 17(3):249– 254, May 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1986.11972967.

Anonymous:1986:Pd

[Ano86k] Anonymous. Problems. College Mathematics Journal, 17(4):360– 365, September 1986. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1986.11972982.

Anonymous:1986:Pe

[Ano86l] Anonymous. Problems. College Mathematics Journal, 17(5):442– 449, November 1986. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1986.11972997.

Anonymous:1987:EBa

[Ano87a] Anonymous. Editorial board. College Mathematics Journal, 18 (1):??, January 1987. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1987.11972999.

Anonymous:1987:EBb

[Ano87b] Anonymous. Editorial board. College Mathematics Journal, 18 (2):??, March 1987. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1987.11973016.

Anonymous:1987:EBc

[Ano87c] Anonymous. Editorial board. College Mathematics Journal, 18(3):??, May 1987. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1987.11973031.

Anonymous:1987:EBd

[Ano87d] Anonymous. Editorial board. College Mathematics Journal, 18 (4):??, September 1987. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1987.11973044.

Anonymous:1987:EBe

[Ano87e] Anonymous. Editorial board. College Mathematics Journal, 18 (5):??, November 1987. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1987.11973059.

Anonymous:1987:GA

[Ano87f] Anonymous. Guidelines for authors. College Mathematics Journal, 18(1):33, January 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1987.11973003.

Anonymous:1987:Pa

[Ano87g] Anonymous. Problems. College Mathematics Journal, 18(1): 69–77, January 1987. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1987.11973014.

Anonymous:1987:Pb

[Ano87h] Anonymous. Problems. College Mathematics Journal, 18(2): 160–166, March 1987. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1987.11973029.

Anonymous:1987:Pc

[Ano87i] Anonymous. Problems. College Mathematics Journal, 18(3):248– 252, May 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1987.11973042.

Anonymous:1987:Pd

[Ano87j] Anonymous. Problems. College Mathematics Journal, 18(4):340– 345, September 1987. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1987.11973057.

Anonymous:1987:Pe

[Ano87k] Anonymous. Problems. College Mathematics Journal, 18(5):425– 432, November 1987. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1987.11973070.

Anonymous:1988:DMR

[Ano88a] Anonymous. David Moore’s response …. College Mathematics Journal, 19(1):32–34, January 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1988.
11973086.

Anonymous:1988:EBa

[Ano88b] Anonymous. Editorial board. College Mathematics Journal, 19 (1):??, January 1988. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1988.11973072.

Anonymous:1988:EBb

[Ano88c] Anonymous. Editorial board. College Mathematics Journal, 19 (2):??, March 1988. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1988.11973100.

Anonymous:1988:EBc

[Ano88d] Anonymous. Editorial board. College Mathematics Journal, 19(3):??, May 1988. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1988.11973116.

Anonymous:1988:EBd

[Ano88e] Anonymous. Editorial board. College Mathematics Journal, 19 (4):??, September 1988. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1988.11973130.

Anonymous:1988:EBe

[Ano88f] Anonymous. Editorial board. College Mathematics Journal, 19 (5):??, November 1988. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1988.11973144.

Anonymous:1988:ECR

[Ano88g] Anonymous. The Editor’s concluding remarks. College Mathematics Journal, 19(1):34–35, January 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1988.
11973087.

Anonymous:1988:MHa

[Ano88h] Anonymous. Media highlights. College Mathematics Journal, 19 (3):295–300, May 1988. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1988.11973129.

Anonymous:1988:MHb

[Ano88i] Anonymous. Media highlights. College Mathematics Journal, 19(4):374–378, September 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1988.11973143.

Anonymous:1988:MHc

[Ano88j] Anonymous. Media highlights. College Mathematics Journal, 19(5):451–454, November 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1988.11973154.

Anonymous:1988:Pa

[Ano88k] Anonymous. Problems. College Mathematics Journal, 19(1): 81–87, January 1988. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1988.11973098.

Anonymous:1988:Pb

[Ano88l] Anonymous. Problems. College Mathematics Journal, 19(2): 186–193, March 1988. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1988.11973114.

Anonymous:1988:Pc

[Ano88m] Anonymous. Problems. College Mathematics Journal, 19(3):290– 294, May 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1988.11973128.

Anonymous:1988:Pd

[Ano88n] Anonymous. Problems. College Mathematics Journal, 19(4):370– 373, September 1988. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1988.11973142.

Anonymous:1988:Pe

[Ano88o] Anonymous. Problems. College Mathematics Journal, 19(5):448– 450, November 1988. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1988.11973153.

Anonymous:1989:A

[Ano89a] Anonymous. Acknowledgement. College Mathematics Journal, 20(5):402, November 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1989.11973263.

Anonymous:1989:EBa

[Ano89b] Anonymous. Editorial board. College Mathematics Journal, 20 (1):??, January 1989. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1989.11973197.

Anonymous:1989:EBb

[Ano89c] Anonymous. Editorial board. College Mathematics Journal, 20 (2):??, March 1989. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1989.11973214.

Anonymous:1989:EBc

[Ano89d] Anonymous. Editorial board. College Mathematics Journal, 20(3):??, May 1989. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1989.11973229.

Anonymous:1989:EBd

[Ano89e] Anonymous. Editorial board. College Mathematics Journal, 20 (4):??, September 1989. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1989.11973244.

Anonymous:1989:EBe

[Ano89f] Anonymous. Editorial board. College Mathematics Journal, 20 (5):??, November 1989. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1989.11973259.

Anonymous:1989:FFFa

[Ano89g] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 20(1):49–50, January 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1989.
11973202.

Anonymous:1989:FFFb

[Ano89h] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 20(2):132–133, March 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1989.
11973219.

Anonymous:1989:FFFc

[Ano89i] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 20(3):226–227, May 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1989.
11973233.

Anonymous:1989:FFFd

[Ano89j] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 20(4):317–319, September 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1989.
11973248.

Anonymous:1989:FFFe

[Ano89k] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 20(5):403–404, November 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1989.
11973264.

Anonymous:1989:GA

[Ano89l] Anonymous. Guidelines for authors. College Mathematics Journal, 20(1):80, January 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1989.11973212.

Anonymous:1989:MH

[Ano89m] Anonymous. Media highlights. College Mathematics Journal, 20(1):81–86, January 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1989.11973213.

Anonymous:1989:Pa

[Ano89n] Anonymous. Problems. College Mathematics Journal, 20(1): 68–79, January 1989. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1989.11973211.

Anonymous:1989:Pb

[Ano89o] Anonymous. Problems. College Mathematics Journal, 20(2): 163–173, March 1989. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1989.11973227.

Anonymous:1989:Pc

[Ano89p] Anonymous. Problems. College Mathematics Journal, 20(3):256– 265, May 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1989.11973242.

Anonymous:1989:Pd

[Ano89q] Anonymous. Problems. College Mathematics Journal, 20(4):343– 354, September 1989. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1989.11973257.

Anonymous:1989:Pe

[Ano89r] Anonymous. Problems. College Mathematics Journal, 20(5):442– 451, November 1989. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1989.11973274.

Anonymous:1990:EBa

[Ano90a] Anonymous. Editorial board. College Mathematics Journal, 21 (1):??, January 1990. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1990.11973276.

Anonymous:1990:EBb

[Ano90b] Anonymous. Editorial board. College Mathematics Journal, 21 (2):??, March 1990. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1990.11973289.

Anonymous:1990:EBc

[Ano90c] Anonymous. Editorial board. College Mathematics Journal, 21(3):??, May 1990. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1990.11973304.

Anonymous:1990:EBd

[Ano90d] Anonymous. Editorial board. College Mathematics Journal, 21 (4):??, September 1990. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1990.11973318.

Anonymous:1990:EBe

[Ano90e] Anonymous. Editorial board. College Mathematics Journal, 21 (5):??, November 1990. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1990.11973333.

Anonymous:1990:FFFb

[Ano90f] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 21(2):127–128, March 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1990.
11973295.

Anonymous:1990:FFFc

[Ano90g] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 21(3):216–218, May 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1990.
11973308.

Anonymous:1990:FFFd

[Ano90h] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 21(4):302–304, September 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1990.
11973323.

Anonymous:1990:FFFe

[Ano90i] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 21(5):394–396, November 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1990.
11973339.

Anonymous:1990:FFFa

[Ano90j] Anonymous. Flaws, fallacies, and flimflam. College Mathematics Journal, 21(1):35–36, January 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1990.
11973280.

Anonymous:1990:PSa

[Ano90k] Anonymous. Problems and solutions. College Mathematics Journal, 21(1):65–72, January 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1990.11973287.

Anonymous:1990:PSb

[Ano90l] Anonymous. Problems and solutions. College Mathematics Journal, 21(2):150–157, March 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1990.11973302.

Anonymous:1990:PSc

[Ano90m] Anonymous. Problems and solutions. College Mathematics Journal, 21(3):246–254, May 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1990.11973316.

Anonymous:1990:PSd

[Ano90n] Anonymous. Problems and solutions. College Mathematics Journal, 21(4):334–341, September 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1990.
11973331.

Anonymous:1990:PSe

[Ano90o] Anonymous. Problems and solutions. College Mathematics Journal, 21(5):423–430, November 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1990.
11973349.

Anonymous:1991:EBa

[Ano91a] Anonymous. Editorial board. College Mathematics Journal, 22 (1):??, January 1991. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1991.11973351.

Anonymous:1991:EBb

[Ano91b] Anonymous. Editorial board. College Mathematics Journal, 22 (2):??, March 1991. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1991.11973365.

Anonymous:1991:EBc

[Ano91c] Anonymous. Editorial board. College Mathematics Journal, 22(3):??, May 1991. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1991.11973380.

Anonymous:1991:EBd

[Ano91d] Anonymous. Editorial board. College Mathematics Journal, 22 (4):??, September 1991. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1991.11973396.

Anonymous:1991:EBe

[Ano91e] Anonymous. Editorial board. College Mathematics Journal, 22 (5):??, November 1991. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1991.11973413.

Anonymous:1991:FFFa

[Ano91f] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 22(1):39–41, January 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1991.
11973356.

Anonymous:1991:FFFb

[Ano91g] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 22(2):131–133, March 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1991.
11973369.

Anonymous:1991:FFFc

[Ano91h] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 22(3):220–223, May 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1991.
11973384.

Anonymous:1991:FFFd

[Ano91i] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 22(4):307–310, September 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1991.
11973400.

Anonymous:1991:FFFe

[Ano91j] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 22(5):403–406, November 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1991.
11973415.

Anonymous:1991:PSa

[Ano91k] Anonymous. Problems and solutions. College Mathematics Journal, 22(1):70–76, January 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1991.11973363.

Anonymous:1991:PSb

[Ano91l] Anonymous. Problems and solutions. College Mathematics Journal, 22(2):167–172, March 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1991.11973378.

Anonymous:1991:PSc

[Ano91m] Anonymous. Problems and solutions. College Mathematics Journal, 22(3):255–260, May 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1991.11973394.

Anonymous:1991:PSd

[Ano91n] Anonymous. Problems and solutions. College Mathematics Journal, 22(4):343–348, September 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1991.
11973411.

Anonymous:1991:PSe

[Ano91o] Anonymous. Problems and solutions. College Mathematics Journal, 22(5):443–449, November 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1991.
11973425.

Anonymous:1991:RR

[Ano91p] Anonymous. Reward of the rings. College Mathematics Journal, 22(5):418–420, November 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1991.11973421.

Anonymous:1991:SF

[Ano91q] Anonymous. Secrets of the Faro. College Mathematics Journal, 22(2):144–147, March 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1991.11973374.

Anonymous:1992:EBa

[Ano92a] Anonymous. Editorial board. College Mathematics Journal, 23 (1):??, January 1992. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1992.11973427.

Anonymous:1992:EBb

[Ano92b] Anonymous. Editorial board. College Mathematics Journal, 23 (2):??, March 1992. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1992.11973441.

Anonymous:1992:EBc

[Ano92c] Anonymous. Editorial board. College Mathematics Journal, 23(3):??, May 1992. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1992.11973454.

Anonymous:1992:EBd

[Ano92d] Anonymous. Editorial board. College Mathematics Journal, 23 (4):??, September 1992. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1992.11973468.

Anonymous:1992:EBe

[Ano92e] Anonymous. Editorial board. College Mathematics Journal, 23 (5):??, November 1992. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1992.11973485.

Anonymous:1992:ERM

[Ano92f] Anonymous. Erratum: Relative maxima or minima for a function of two variables. College Mathematics Journal, 23(4):314, September 1992. CODEN ???? ISSN 0746-8342 (print), 1931-

1346 (electronic). URL http://www.tandfonline.com/doi/
abs/10.1080/07468342.1992.11973478.

Anonymous:1992:FFFa

[Ano92g] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 23(1):36–38, January 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1992.
11973430.

Anonymous:1992:FFFb

[Ano92h] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 23(2):142–144, March 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1992.
11973445.

Anonymous:1992:FFFc

[Ano92i] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 23(3):203–206, May 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1992.
11973457.

Anonymous:1992:FFFd

[Ano92j] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 23(4):304–307, September 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1992.
11973474.

Anonymous:1992:PSa

[Ano92k] Anonymous. Problems and solutions. College Mathematics Journal, 23(1):69–76, January 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1992.11973439.

Anonymous:1992:PSb

[Ano92l] Anonymous. Problems and solutions. College Mathematics Journal, 23(2):162–167, March 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1992.11973452.

Anonymous:1992:PSc

[Ano92m] Anonymous. Problems and solutions. College Mathematics Journal, 23(3):247–252, May 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1992.11973466.

Anonymous:1992:PSd

[Ano92n] Anonymous. Problems and solutions. College Mathematics Journal, 23(4):340–347, September 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1992.
11973483.

Anonymous:1992:PSe

[Ano92o] Anonymous. Problems and solutions. College Mathematics Journal, 23(5):435–440, November 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1992.
11973497.

Anonymous:1992:SR

[Ano92p] Anonymous. Software reviews. College Mathematics Journal, 23 (3):241–246, May 1992. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1992.11973465.

Anonymous:1993:EBa

[Ano93a] Anonymous. Editorial board. College Mathematics Journal, 24 (1):??, January 1993. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1993.11973499.

Anonymous:1993:EBb

[Ano93b] Anonymous. Editorial board. College Mathematics Journal, 24 (2):??, March 1993. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1993.11973517.

Anonymous:1993:EBc

[Ano93c] Anonymous. Editorial board. College Mathematics Journal, 24(3):??, May 1993. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1993.11973529.

Anonymous:1993:EBd

[Ano93d] Anonymous. Editorial board. College Mathematics Journal, 24 (4):??, September 1993. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1993.11973546.

Anonymous:1993:EBe

[Ano93e] Anonymous. Editorial board. College Mathematics Journal, 24 (5):??, November 1993. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1993.11973561.

Anonymous:1993:FFFa

[Ano93f] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 24(2):149–154, March 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1993.
11973519.

Anonymous:1993:FFFb

[Ano93g] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 24(3):229–232, May 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1993.
11973533.

Anonymous:1993:FFFc

[Ano93h] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 24(4):343–346, September 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1993.
11973551.

Anonymous:1993:LTA

[Ano93i] Anonymous. The linear transformation associated with a graph. College Mathematics Journal, 24(1):76–78, January 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1993.11973512.

Anonymous:1993:MHa

[Ano93j] Anonymous. Media highlights. College Mathematics Journal, 24(4):386–393, September 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1993.11973560.

Anonymous:1993:MHb

[Ano93k] Anonymous. Media highlights. College Mathematics Journal, 24(5):481–488, November 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1993.11973572.

Anonymous:1993:PSa

[Ano93l] Anonymous. Problems and solutions. College Mathematics Journal, 24(1):95–100, January 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1993.11973515.

Anonymous:1993:PSb

[Ano93m] Anonymous. Problems and solutions. College Mathematics Journal, 24(2):183–190, March 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1993.11973527.

Anonymous:1993:PSc

[Ano93n] Anonymous. Problems and solutions. College Mathematics Journal, 24(3):271–276, May 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1993.11973544.

Anonymous:1993:PSd

[Ano93o] Anonymous. Problems and solutions. College Mathematics Journal, 24(4):377–385, September 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1993.
11973559.

Anonymous:1993:PSe

[Ano93p] Anonymous. Problems and solutions. College Mathematics Journal, 24(5):472–480, November 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1993.
11973571.

Anonymous:1994:CCa

[Ano94a] Anonymous. Computer corner. College Mathematics Journal, 25(3):231, May 1994. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1994.11973616.

Anonymous:1994:CCb

[Ano94b] Anonymous. Computer corner. College Mathematics Journal, 25(4):325, September 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1994.11973632.

Anonymous:1994:EBa

[Ano94c] Anonymous. Editorial board. College Mathematics Journal, 25 (1):??, January 1994. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1994.11973573.

Anonymous:1994:EBb

[Ano94d] Anonymous. Editorial board. College Mathematics Journal, 25 (2):??, March 1994. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1994.11973587.

Anonymous:1994:EBc

[Ano94e] Anonymous. Editorial board. College Mathematics Journal, 25(3):??, May 1994. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1994.11973604.

Anonymous:1994:EBd

[Ano94f] Anonymous. Editorial board. College Mathematics Journal, 25 (4):??, September 1994. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1994.11973621.

Anonymous:1994:EBe

[Ano94g] Anonymous. Editorial board. College Mathematics Journal, 25 (5):??, November 1994. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1994.11973638.

Anonymous:1994:FFFa

[Ano94h] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 25(1):35–37, January 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1994.
11973577.

Anonymous:1994:FFFb

[Ano94i] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 25(2):128–131, March 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1994.
11973595.

Anonymous:1994:FFFc

[Ano94j] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 25(3):223–225, May 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1994.
11973611.

Anonymous:1994:FFFd

[Ano94k] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 25(5):432–435, November 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1994.
11973647.

Anonymous:1994:MHa

[Ano94l] Anonymous. Media highlights. College Mathematics Journal, 25(1):71–76, January 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1994.11973586.

Anonymous:1994:MHb

[Ano94m] Anonymous. Media highlights. College Mathematics Journal, 25(5):468–473, November 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1994.11973655.

Anonymous:1994:PSa

[Ano94n] Anonymous. Problems and solutions. College Mathematics Journal, 25(1):64–70, January 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1994.11973585.

Anonymous:1994:PSb

[Ano94o] Anonymous. Problems and solutions. College Mathematics Journal, 25(2):157–163, March 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1994.11973602.

Anonymous:1994:PSc

[Ano94p] Anonymous. Problems and solutions. College Mathematics Journal, 25(5):462–467, November 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1994.
11973654.

Anonymous:1995:Ab

[Ano95a] Anonymous. Acknowledgements. College Mathematics Journal, 26(5):404, November 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1995.11973740.

Anonymous:1995:Aa

[Ano95b] Anonymous. Acknowledgment. College Mathematics Journal, 26 (1):83, January 1995. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1995.11973674.

Anonymous:1995:AYC

[Ano95c] Anonymous. Announcing 25-year CMJ Topic Index. College Mathematics Journal, 26(1):82, January 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1995.
11973673.

Anonymous:1995:CC

[Ano95d] Anonymous. Computer corner. College Mathematics Journal, 26(4):316, September 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1995.11973722.

Anonymous:1995:EBa

[Ano95e] Anonymous. Editorial board. College Mathematics Journal, 26 (1):??, January 1995. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1995.11973656.

Anonymous:1995:EBb

[Ano95f] Anonymous. Editorial board. College Mathematics Journal, 26 (2):??, March 1995. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1995.11973675.

Anonymous:1995:EBc

[Ano95g] Anonymous. Editorial board. College Mathematics Journal, 26(3):??, May 1995. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1995.11973692.

Anonymous:1995:EBd

[Ano95h] Anonymous. Editorial board. College Mathematics Journal, 26 (4):??, September 1995. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1995.11973710.

Anonymous:1995:EBe

[Ano95i] Anonymous. Editorial board. College Mathematics Journal, 26 (5):??, November 1995. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1995.11973727.

Anonymous:1995:FFFa

[Ano95j] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 26(1):35–38, January 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1995.
11973664.

Anonymous:1995:FFFb

[Ano95k] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 26(2):132–134, March 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1995.
11973682.

Anonymous:1995:FFFc

[Ano95l] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 26(3):221–222, May 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1995.
11973700.

Anonymous:1995:FFFd

[Ano95m] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 26(4):302–304, September 1995. CODEN ????

ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1995.
11973717.

Anonymous:1995:FFFe

[Ano95n] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 26(5):382–384, November 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1995.
11973734.

Anonymous:1995:MH

[Ano95o] Anonymous. Media highlights. College Mathematics Journal, 26 (3):251–258, May 1995. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1995.11973709.

Anonymous:1995:PSa

[Ano95p] Anonymous. Problems and solutions. College Mathematics Journal, 26(1):66–72, January 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1995.11973671.

Anonymous:1995:PSb

[Ano95q] Anonymous. Problems and solutions. College Mathematics Journal, 26(2):157–162, March 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1995.11973690.

Anonymous:1995:PSc

[Ano95r] Anonymous. Problems and solutions. College Mathematics Journal, 26(3):244–250, May 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1995.11973708.

Anonymous:1995:PSd

[Ano95s] Anonymous. Problems and solutions. College Mathematics Journal, 26(4):329–334, September 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1995.
11973725.

Anonymous:1995:PSe

[Ano95t] Anonymous. Problems and solutions. College Mathematics Journal, 26(5):405–410, November 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1995.
11973741.

Anonymous:1996:A

[Ano96a] Anonymous. Acknowledgments. College Mathematics Journal, 27(5):418, November 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1996.11973821.

Anonymous:1996:CCa

[Ano96b] Anonymous. Computer corner. College Mathematics Journal, 27(3):220, May 1996. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1996.11973783.

Anonymous:1996:CCb

[Ano96c] Anonymous. Computer corner. College Mathematics Journal, 27(5):394, November 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1996.11973816.

Anonymous:1996:EBa

[Ano96d] Anonymous. Editorial board. College Mathematics Journal, 27 (1):??, January 1996. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1996.11973743.

Anonymous:1996:EBb

[Ano96e] Anonymous. Editorial board. College Mathematics Journal, 27 (2):??, March 1996. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1996.11973756.

Anonymous:1996:EBc

[Ano96f] Anonymous. Editorial board. College Mathematics Journal, 27(3):??, May 1996. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1996.11973771.

Anonymous:1996:EBd

[Ano96g] Anonymous. Editorial board. College Mathematics Journal, 27 (4):??, September 1996. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1996.11973788.

Anonymous:1996:EBe

[Ano96h] Anonymous. Editorial board. College Mathematics Journal, 27 (5):??, November 1996. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1996.11973803.

Anonymous:1996:FFFa

[Ano96i] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 27(1):43–46, January 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1996.
11973748.

Anonymous:1996:FFFb

[Ano96j] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 27(2):116–118, March 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1996.
11973761.

Anonymous:1996:FFFc

[Ano96k] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 27(3):204–205, May 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1996.
11973777.

Anonymous:1996:FFFd

[Ano96l] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 27(4):282–284, September 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1996.
11973793.

Anonymous:1996:FFFe

[Ano96m] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 27(5):370–373, November 1996. CODEN ????

ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1996.
11973810.

Anonymous:1996:GAR

[Ano96n] Anonymous. George Andrews replies. College Mathematics Journal, 27(5):354–355, November 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1996.
11973807.

Anonymous:1996:MHa

[Ano96o] Anonymous. Media highlights. College Mathematics Journal, 27(1):79–85, January 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1996.11973755.

Anonymous:1996:MHb

[Ano96p] Anonymous. Media highlights. College Mathematics Journal, 27 (3):236–243, May 1996. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1996.11973787.

Anonymous:1996:PSa

[Ano96q] Anonymous. Problems and solutions. College Mathematics Journal, 27(1):74–78, January 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1996.11973754.

Anonymous:1996:PSb

[Ano96r] Anonymous. Problems and solutions. College Mathematics Journal, 27(2):149–155, March 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1996.11973769.

Anonymous:1996:PSc

[Ano96s] Anonymous. Problems and solutions. College Mathematics Journal, 27(3):231–235, May 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1996.11973786.

Anonymous:1996:PSd

[Ano96t] Anonymous. Problems and solutions. College Mathematics Journal, 27(4):312–316, September 1996. CODEN ????

ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1996.
11973801.

Anonymous:1996:PSe

[Ano96u] Anonymous. Problems and solutions. College Mathematics Journal, 27(5):404–409, November 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1996.
11973819.

Anonymous:1997:A

[Ano97a] Anonymous. Acknowledgments. College Mathematics Journal, 28(5):380, November 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1997.11973894.

Anonymous:1997:CCa

[Ano97b] Anonymous. Computer corner. College Mathematics Journal, 28 (1):59, January 1997. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1997.11973835.

Anonymous:1997:CCb

[Ano97c] Anonymous. Computer corner. College Mathematics Journal, 28(5):394, November 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1997.11973900.

Anonymous:1997:EBa

[Ano97d] Anonymous. Editorial board. College Mathematics Journal, 28 (1):??, January 1997. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1997.11973822.

Anonymous:1997:EBb

[Ano97e] Anonymous. Editorial board. College Mathematics Journal, 28 (2):??, March 1997. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1997.11973840.

Anonymous:1997:EBc

[Ano97f] Anonymous. Editorial board. College Mathematics Journal, 28(3):??, May 1997. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1997.11973856.

Anonymous:1997:EBd

[Ano97g] Anonymous. Editorial board. College Mathematics Journal, 28 (4):??, September 1997. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1997.11973872.

Anonymous:1997:EBe

[Ano97h] Anonymous. Editorial board. College Mathematics Journal, 28 (5):??, November 1997. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1997.11973886.

Anonymous:1997:FFFa

[Ano97i] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 28(1):44–46, January 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1997.
11973828.

Anonymous:1997:FFFb

[Ano97j] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 28(2):119–121, March 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1997.
11973846.

Anonymous:1997:FFFc

[Ano97k] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 28(3):198–200, May 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1997.
11973861.

Anonymous:1997:FFFd

[Ano97l] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 28(4):285–287, September 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1997.
11973876.

Anonymous:1997:FFFe

[Ano97m] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 28(5):377–380, November 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1997.
11973893.

Anonymous:1997:GA

[Ano97n] Anonymous. Guidelines for authors. College Mathematics Journal, 28(1):43, January 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1997.11973827.

Anonymous:1997:MHa

[Ano97o] Anonymous. Media highlights. College Mathematics Journal, 28(1):73–79, January 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1997.11973839.

Anonymous:1997:MHb

[Ano97p] Anonymous. Media highlights. College Mathematics Journal, 28(2):151–156, March 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1997.11973855.

Anonymous:1997:MHc

[Ano97q] Anonymous. Media highlights. College Mathematics Journal, 28(5):412–419, November 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1997.11973904.

Anonymous:1997:PSa

[Ano97r] Anonymous. Problems and solutions. College Mathematics Journal, 28(1):67–72, January 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1997.11973838.

Anonymous:1997:PSb

[Ano97s] Anonymous. Problems and solutions. College Mathematics Journal, 28(2):145–150, March 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1997.11973854.

Anonymous:1997:PSc

[Ano97t] Anonymous. Problems and solutions. College Mathematics Journal, 28(3):231–236, May 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1997.11973870.

Anonymous:1997:PSd

[Ano97u] Anonymous. Problems and solutions. College Mathematics Journal, 28(4):315–321, September 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1997.
11973884.

Anonymous:1997:PSe

[Ano97v] Anonymous. Problems and solutions. College Mathematics Journal, 28(5):406–411, November 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1997.
11973903.

Anonymous:1998:CC

[Ano98a] Anonymous. Computer corner. College Mathematics Journal, 29 (2):148, March 1998. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1998.11973932.

Anonymous:1998:EBa

[Ano98b] Anonymous. Editorial board. College Mathematics Journal, 29 (1):??, January 1998. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1998.11973905.

Anonymous:1998:EBb

[Ano98c] Anonymous. Editorial board. College Mathematics Journal, 29 (2):??, March 1998. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1998.11973920.

Anonymous:1998:EBc

[Ano98d] Anonymous. Editorial board. College Mathematics Journal, 29(3):??, May 1998. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1998.11973937.

Anonymous:1998:EBd

[Ano98e] Anonymous. Editorial board. College Mathematics Journal, 29 (4):??, September 1998. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1998.11973952.

Anonymous:1998:EBe

[Ano98f] Anonymous. Editorial board. College Mathematics Journal, 29 (5):??, November 1998. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1998.11973970.

Anonymous:1998:FFFa

[Ano98g] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 29(1):34–36, January 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1998.
11973910.

Anonymous:1998:FFFb

[Ano98h] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 29(2):134–136, March 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1998.
11973926.

Anonymous:1998:FFFc

[Ano98i] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 29(4):301–303, September 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1998.
11973959.

Anonymous:1998:FFFd

[Ano98j] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 29(5):405–407, November 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1998.
11973976.

Anonymous:1998:MH

[Ano98k] Anonymous. Media highlights. College Mathematics Journal, 29(2):158–165, March 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1998.11973936.

Anonymous:1998:PSa

[Ano98l] Anonymous. Problems and solutions. College Mathematics Journal, 29(1):66–72, January 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1998.11973918.

Anonymous:1998:PSb

[Ano98m] Anonymous. Problems and solutions. College Mathematics Journal, 29(2):152–157, March 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1998.11973935.

Anonymous:1998:PSc

[Ano98n] Anonymous. Problems and solutions. College Mathematics Journal, 29(3):242–248, May 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1998.11973950.

Anonymous:1998:PSd

[Ano98o] Anonymous. Problems and solutions. College Mathematics Journal, 29(4):330–336, September 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1998.
11973968.

Anonymous:1998:PSe

[Ano98p] Anonymous. Problems and solutions. College Mathematics Journal, 29(5):434–440, November 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1998.
11973985.

Anonymous:1999:EBa

[Ano99a] Anonymous. Editorial board. College Mathematics Journal, 30 (1):??, January 1999. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1999.11974022.

Anonymous:1999:EBb

[Ano99b] Anonymous. Editorial board. College Mathematics Journal, 30 (2):??, March 1999. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1999.11974037.

Anonymous:1999:EBc

[Ano99c] Anonymous. Editorial board. College Mathematics Journal, 30(3):??, May 1999. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1999.11974052.

Anonymous:1999:EBd

[Ano99d] Anonymous. Editorial board. College Mathematics Journal, 30 (4):??, September 1999. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1999.11974068.

Anonymous:1999:EBe

[Ano99e] Anonymous. Editorial board. College Mathematics Journal, 30 (5):??, November 1999. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1999.11974083.

Anonymous:1999:FFFa

[Ano99f] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 30(1):39–43, January 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1999.
11974028.

Anonymous:1999:FFFb

[Ano99g] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 30(2):129–132, March 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1999.
11974044.

Anonymous:1999:FFFc

[Ano99h] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 30(3):209–211, May 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1999.
11974058.

Anonymous:1999:FFFd

[Ano99i] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 30(4):306–308, September 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1999.
11974075.

Anonymous:1999:FFFe

[Ano99j] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 30(5):383–386, November 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1999.
11974090.

Anonymous:1999:MHa

[Ano99k] Anonymous. Media highlights. College Mathematics Journal, 30(1):69–76, January 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1999.11974035.

Anonymous:1999:MHb

[Ano99l] Anonymous. Media highlights. College Mathematics Journal, 30(2):149–158, March 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1999.11974050.

Anonymous:1999:MHc

[Ano99m] Anonymous. Media highlights. College Mathematics Journal, 30 (3):236–243, May 1999. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1999.11974066.

Anonymous:1999:MHd

[Ano99n] Anonymous. Media highlights. College Mathematics Journal, 30(4):329–337, September 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1999.11974082.

Anonymous:1999:MHe

[Ano99o] Anonymous. Media highlights. College Mathematics Journal, 30(5):413–421, November 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1999.11974098.

Anonymous:1999:Ma

[Ano99p] Anonymous. Miscellanea. College Mathematics Journal, 30(4): 324–328, September 1999. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1999.11974081.

Anonymous:1999:Mb

[Ano99q] Anonymous. Miscellanea. College Mathematics Journal, 30(5): 422–424, November 1999. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1999.11974099.

Anonymous:1999:PSa

[Ano99r] Anonymous. Problems and solutions. College Mathematics Journal, 30(1):60–66, January 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1999.11974033.

Anonymous:1999:PSb

[Ano99s] Anonymous. Problems and solutions. College Mathematics Journal, 30(2):143–148, March 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1999.11974049.

Anonymous:1999:PSc

[Ano99t] Anonymous. Problems and solutions. College Mathematics Journal, 30(3):230–235, May 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1999.11974065.

Anonymous:1999:PSd

[Ano99u] Anonymous. Problems and solutions. College Mathematics Journal, 30(4):318–323, September 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1999.
11974080.

Anonymous:1999:PSe

[Ano99v] Anonymous. Problems and solutions. College Mathematics Journal, 30(5):406–412, November 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1999.
11974097.

Anonymous:2000:EBa

[Ano00a] Anonymous. Editorial board. College Mathematics Journal, 31 (1):??, January 2000. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2000.11974101.

Anonymous:2000:EBb

[Ano00b] Anonymous. Editorial board. College Mathematics Journal, 31 (2):??, March 2000. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2000.11974117.

Anonymous:2000:EBc

[Ano00c] Anonymous. Editorial board. College Mathematics Journal, 31(3):??, May 2000. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2000.11974136.

Anonymous:2000:EBd

[Ano00d] Anonymous. Editorial board. College Mathematics Journal, 31 (4):??, September 2000. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2000.11974153.

Anonymous:2000:EBe

[Ano00e] Anonymous. Editorial board. College Mathematics Journal, 31 (5):??, November 2000. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2000.11974170.

Anonymous:2000:FFF

[Ano00f] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 31(5):393–396, November 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2000.
11974179.

Anonymous:2000:MHa

[Ano00g] Anonymous. Media highlights. College Mathematics Journal, 31(1):69–76, January 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2000.11974115.

Anonymous:2000:MHb

[Ano00h] Anonymous. Media highlights. College Mathematics Journal, 31(2):148–156, March 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2000.11974134.

Anonymous:2000:MHc

[Ano00i] Anonymous. Media highlights. College Mathematics Journal, 31 (3):227–235, May 2000. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2000.11974150.

Anonymous:2000:MHd

[Ano00j] Anonymous. Media highlights. College Mathematics Journal, 31(4):321–328, September 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2000.11974168.

Anonymous:2000:MHe

[Ano00k] Anonymous. Media highlights. College Mathematics Journal, 31(5):415–420, November 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2000.11974185.

Anonymous:2000:Ma

[Ano00l] Anonymous. Miscellanea. College Mathematics Journal, 31(1): 68, January 2000. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2000.11974114.

Anonymous:2000:Mb

[Ano00m] Anonymous. Miscellanea. College Mathematics Journal, 31(3): 236–239, May 2000. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2000.11974151.

Anonymous:2000:PSa

[Ano00n] Anonymous. Problems and solutions. College Mathematics Journal, 31(1):62–67, January 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2000.11974113.

Anonymous:2000:PSb

[Ano00o] Anonymous. Problems and solutions. College Mathematics Journal, 31(2):137–143, March 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2000.11974132.

Anonymous:2000:PSc

[Ano00p] Anonymous. Problems and solutions. College Mathematics Journal, 31(3):219–226, May 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2000.11974149.

Anonymous:2000:PSd

[Ano00q] Anonymous. Problems and solutions. College Mathematics Journal, 31(4):314–320, September 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2000.
11974167.

Anonymous:2000:PSe

[Ano00r] Anonymous. Problems and solutions. College Mathematics Journal, 31(5):408–414, November 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2000.
11974184.

Anonymous:2000:SR

[Ano00s] Anonymous. Software reviews. College Mathematics Journal, 31(4):310–313, September 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2000.11974166.

Anonymous:2001:CCa

[Ano01a] Anonymous. Classroom capsules. College Mathematics Journal, 32(1):52–53, January 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2001.11921853.

Anonymous:2001:CCb

[Ano01b] Anonymous. Classroom capsules. College Mathematics Journal, 32(2):119–137, March 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2001.11921864.

Anonymous:2001:CCc

[Ano01c] Anonymous. Classroom capsules. College Mathematics Journal, 32(3):201–210, May 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2001.11921877.

Anonymous:2001:CCd

[Ano01d] Anonymous. Classroom capsules. College Mathematics Journal, 32(4):283–296, September 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2001.11921892.

Anonymous:2001:CCe

[Ano01e] Anonymous. Classroom capsules. College Mathematics Journal, 32(5):369–380, November 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2001.11921906.

Anonymous:2001:EBa

[Ano01f] Anonymous. Editorial board. College Mathematics Journal, 32 (1):??, January 2001. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2001.11921843.

Anonymous:2001:EBb

[Ano01g] Anonymous. Editorial board. College Mathematics Journal, 32 (2):??, March 2001. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2001.11921857.

Anonymous:2001:EBc

[Ano01h] Anonymous. Editorial board. College Mathematics Journal, 32(3):??, May 2001. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2001.11921869.

Anonymous:2001:EBd

[Ano01i] Anonymous. Editorial board. College Mathematics Journal, 32 (4):??, September 2001. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2001.11921883.

Anonymous:2001:EBe

[Ano01j] Anonymous. Editorial board. College Mathematics Journal, 32 (5):??, November 2001. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2001.11921897.

Anonymous:2001:FFFa

[Ano01k] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 32(1):48–51, January 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2001.
11921852.

Anonymous:2001:FFFc

[Ano01l] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 32(3):197–200, May 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2001.
11921876.

Anonymous:2001:FFFd

[Ano01m] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 32(4):279–282, September 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2001.
11921891.

Anonymous:2001:FFFe

[Ano01n] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 32(5):363–368, November 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2001.
11921905.

Anonymous:2001:FFFb

[Ano01o] Anonymous. Fallacies, flaws, and flimplam. College Mathematics Journal, 32(2):115–118, March 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2001.
11921863.

Anonymous:2001:MHa

[Ano01p] Anonymous. Media highlights. College Mathematics Journal, 32(1):72–80, January 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2001.11921856.

Anonymous:2001:MHb

[Ano01q] Anonymous. Media highlights. College Mathematics Journal, 32(2):147–154, March 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2001.11921867.

Anonymous:2001:MHc

[Ano01r] Anonymous. Media highlights. College Mathematics Journal, 32 (3):223–231, May 2001. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2001.11921880.

Anonymous:2001:MHd

[Ano01s] Anonymous. Media highlights. College Mathematics Journal, 32(4):304–313, September 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2001.11921894.

Anonymous:2001:MHe

[Ano01t] Anonymous. Media highlights. College Mathematics Journal, 32(5):394–401, November 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2001.11921909.

Anonymous:2001:Ma

[Ano01u] Anonymous. Miscellanea. College Mathematics Journal, 32(1): 68–71, January 2001. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2001.11921855.

Anonymous:2001:Mb

[Ano01v] Anonymous. Miscellanea. College Mathematics Journal, 32(3): 238–240, May 2001. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2001.11921882.

Anonymous:2001:Mc

[Ano01w] Anonymous. Miscellanea. College Mathematics Journal, 32(4): 317–318, September 2001. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2001.11921896.

Anonymous:2001:PSa

[Ano01x] Anonymous. Problems and solutions. College Mathematics Journal, 32(1):60–67, January 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2001.11921854.

Anonymous:2001:PSb

[Ano01y] Anonymous. Problems and solutions. College Mathematics Journal, 32(2):141–146, March 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2001.11921866.

Anonymous:2001:PSc

[Ano01z] Anonymous. Problems and solutions. College Mathematics Journal, 32(3):211–217, May 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2001.11921878.

Anonymous:2001:PSd

[Ano01-27] Anonymous. Problems and solutions. College Mathematics Journal, 32(4):297–303, September 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2001.
11921893.

Anonymous:2001:PSe

[Ano01-28] Anonymous. Problems and solutions. College Mathematics Journal, 32(5):381–387, November 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2001.
11921907.

Anonymous:2001:SR

[Ano01-29] Anonymous. Software reviews. College Mathematics Journal, 32(2):138–140, March 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2001.11921865.

Anonymous:2001:SRP

[Ano01-30] Anonymous. Student research projects. College Mathematics Journal, 32(5):388–393, November 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2001.
11921908.

Anonymous:2002:CCa

[Ano02a] Anonymous. Classroom capsules. College Mathematics Journal, 33(1):42–53, January 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2002.11921918.

Anonymous:2002:CCb

[Ano02b] Anonymous. Classroom capsules. College Mathematics Journal, 33(2):140–149, March 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2002.11921933.

Anonymous:2002:CCc

[Ano02c] Anonymous. Classroom capsules. College Mathematics Journal, 33(3):230–239, May 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2002.11921947.

Anonymous:2002:CCd

[Ano02d] Anonymous. Classroom capsules. College Mathematics Journal, 33(4):313–324, September 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2002.11921960.

Anonymous:2002:CCe

[Ano02e] Anonymous. Classroom capsules. College Mathematics Journal, 33(5):399–400, November 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2002.11921971.

Anonymous:2002:EBa

[Ano02f] Anonymous. Editorial board. College Mathematics Journal, 33 (1):??, January 2002. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2002.11921910.

Anonymous:2002:EBb

[Ano02g] Anonymous. Editorial board. College Mathematics Journal, 33 (2):??, March 2002. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2002.11921922.

Anonymous:2002:EBc

[Ano02h] Anonymous. Editorial board. College Mathematics Journal, 33(3):??, May 2002. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2002.11921937.

Anonymous:2002:EBd

[Ano02i] Anonymous. Editorial board. College Mathematics Journal, 33 (4):??, September 2002. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2002.11921951.

Anonymous:2002:EBe

[Ano02j] Anonymous. Editorial board. College Mathematics Journal, 33 (5):??, November 2002. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2002.11921965.

Anonymous:2002:FFFa

[Ano02k] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 33(1):38–41, January 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2002.
11921917.

Anonymous:2002:FFFb

[Ano02l] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 33(2):137–139, March 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2002.
11921932.

Anonymous:2002:FFFc

[Ano02m] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 33(3):226–229, May 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2002.
11921946.

Anonymous:2002:FFFd ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2002.
11921959.

Anonymous:2002:FFFe

[Ano02o] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 33(5):395–398, November 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2002.
11921970.

Anonymous:2002:MHa

[Ano02p] Anonymous. Media highlights. College Mathematics Journal, 33(1):60–68, January 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2002.11921920.

Anonymous:2002:MHb

[Ano02q] Anonymous. Media highlights. College Mathematics Journal, 33(2):157–165, March 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2002.11921935.

Anonymous:2002:MHc

[Ano02r] Anonymous. Media highlights. College Mathematics Journal, 33 (3):247–255, May 2002. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2002.11921949.

Anonymous:2002:MHd

[Ano02s] Anonymous. Media highlights. College Mathematics Journal, 33(4):332–340, September 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2002.11921962.

Anonymous:2002:MHe

[Ano02t] Anonymous. Media highlights. College Mathematics Journal, 33(5):417–425, November 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2002.11921973.

Anonymous:2002:Ma

[Ano02u] Anonymous. Miscellanea. College Mathematics Journal, 33(2): 166–171, March 2002. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2002.11921936.

Anonymous:2002:Mb

[Ano02v] Anonymous. Miscellanea. College Mathematics Journal, 33(4): 345–352, September 2002. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2002.11921964.

Anonymous:2002:Mc

[Ano02w] Anonymous. Miscellanea. College Mathematics Journal, 33(5): 431–432, November 2002. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2002.11921975.

Anonymous:2002:PSa

[Ano02x] Anonymous. Problems and solutions. College Mathematics Journal, 33(1):54–59, January 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2002.11921919.

Anonymous:2002:PSb

[Ano02y] Anonymous. Problems and solutions. College Mathematics Journal, 33(2):150–156, March 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2002.11921934.

Anonymous:2002:PSc

[Ano02z] Anonymous. Problems and solutions. College Mathematics Journal, 33(3):240–246, May 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2002.11921948.

Anonymous:2002:PSd

[Ano02-27] Anonymous. Problems and solutions. College Mathematics Journal, 33(4):325–331, September 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2002.
11921961.

Anonymous:2002:PSe

[Ano02-28] Anonymous. Problems and solutions. College Mathematics Journal, 33(5):410–416, November 2002. CODEN ????

ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2002.
11921972.

Anonymous:2003:CCa

[Ano03a] Anonymous. Classroom capsules. College Mathematics Journal, 34(1):55–67, January 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2003.11921986.

Anonymous:2003:CCb

[Ano03b] Anonymous. Classroom capsules. College Mathematics Journal, 34(2):139–152, March 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2003.11921998.

Anonymous:2003:CCc

[Ano03c] Anonymous. Classroom capsules. College Mathematics Journal, 34(3):227–238, May 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2003.11922011.

Anonymous:2003:CCd

[Ano03d] Anonymous. Classroom capsules. College Mathematics Journal, 34(4):315–326, September 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2003.11922024.

Anonymous:2003:CCe

[Ano03e] Anonymous. Classroom capsules. College Mathematics Journal, 34(5):394–404, November 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2003.11922038.

Anonymous:2003:EBa

[Ano03f] Anonymous. Editorial board. College Mathematics Journal, 34 (1):??, January 2003. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2003.11921976.

Anonymous:2003:EBb

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2003.11921989.

Anonymous:2003:EBc

[Ano03h] Anonymous. Editorial board. College Mathematics Journal, 34(3):??, May 2003. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2003.11922002.

Anonymous:2003:EBd

[Ano03i] Anonymous. Editorial board. College Mathematics Journal, 34 (4):??, September 2003. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2003.11922014.

Anonymous:2003:EBe

[Ano03j] Anonymous. Editorial board. College Mathematics Journal, 34 (5):??, November 2003. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2003.11922028.

Anonymous:2003:FFFa

[Ano03k] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 34(1):50–54, January 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2003.
11921985.

Anonymous:2003:FFFb

[Ano03l] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 34(2):134–138, March 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2003.
11921997.

Anonymous:2003:FFFc

[Ano03m] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 34(3):224–226, May 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2003.
11922010.

Anonymous:2003:FFFd

[Ano03n] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 34(4):312–314, September 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2003.
11922023.

Anonymous:2003:FFFe

[Ano03o] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 34(5):390–393, November 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2003.
11922037.

Anonymous:2003:MHa

[Ano03p] Anonymous. Media highlights. College Mathematics Journal, 34(1):75–83, January 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2003.11921988.

Anonymous:2003:MHb

[Ano03q] Anonymous. Media highlights. College Mathematics Journal, 34(2):159–167, March 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2003.11922000.

Anonymous:2003:MHc

[Ano03r] Anonymous. Media highlights. College Mathematics Journal, 34 (3):246–254, May 2003. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2003.11922013.

Anonymous:2003:MHd

[Ano03s] Anonymous. Media highlights. College Mathematics Journal, 34(4):334–342, September 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2003.11922026.

Anonymous:2003:MHe

[Ano03t] Anonymous. Media highlights. College Mathematics Journal, 34(5):412–420, November 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2003.11922040.

Anonymous:2003:Ma

[Ano03u] Anonymous. Miscellanea. College Mathematics Journal, 34(2): 168–172, March 2003. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2003.11922001.

Anonymous:2003:Mb

[Ano03v] Anonymous. Miscellanea. College Mathematics Journal, 34(5): 421–422, November 2003. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2003.11922041.

Anonymous:2003:PSa

[Ano03w] Anonymous. Problems and solutions. College Mathematics Journal, 34(1):68–74, January 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2003.11921987.

Anonymous:2003:PSb

[Ano03x] Anonymous. Problems and solutions. College Mathematics Journal, 34(2):153–158, March 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2003.11921999.

Anonymous:2003:PSc

[Ano03y] Anonymous. Problems and solutions. College Mathematics Journal, 34(3):239–245, May 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2003.11922012.

Anonymous:2003:PSd

[Ano03z] Anonymous. Problems and solutions. College Mathematics Journal, 34(4):327–333, September 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2003.
11922025.

Anonymous:2003:PSe

[Ano03-27] Anonymous. Problems and solutions. College Mathematics Journal, 34(5):405–411, November 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2003.
11922039.

Anonymous:2004:CCa

[Ano04a] Anonymous. Classroom capsules. College Mathematics Journal, 35(1):43–54, January 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2004.11922051.

Anonymous:2004:CCb

[Ano04b] Anonymous. Classroom capsules. College Mathematics Journal, 35(2):125–135, March 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2004.11922064.

Anonymous:2004:CCc

[Ano04c] Anonymous. Classroom capsules. College Mathematics Journal, 35(3):216–226, May 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2004.11922076.

Anonymous:2004:CCd

[Ano04d] Anonymous. Classroom capsules. College Mathematics Journal, 35(4):300–309, September 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2004.11922090.

Anonymous:2004:CCe

[Ano04e] Anonymous. Classroom capsules. College Mathematics Journal, 35(5):385–401, November 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2004.11922101.

Anonymous:2004:EBa

[Ano04f] Anonymous. Editorial board. College Mathematics Journal, 35 (1):??, January 2004. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2004.11922043.

Anonymous:2004:EBb

[Ano04g] Anonymous. Editorial board. College Mathematics Journal, 35 (2):??, March 2004. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2004.11922054.

Anonymous:2004:EBc

[Ano04h] Anonymous. Editorial board. College Mathematics Journal, 35(3):??, May 2004. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2004.11922068.

Anonymous:2004:EBd

[Ano04i] Anonymous. Editorial board. College Mathematics Journal, 35 (4):??, September 2004. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2004.11922079.

Anonymous:2004:EBe

[Ano04j] Anonymous. Editorial board. College Mathematics Journal, 35 (5):??, November 2004. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2004.11922093.

Anonymous:2004:FFFa

[Ano04k] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 35(1):40–42, January 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2004.
11922050.

Anonymous:2004:FFFb

[Ano04l] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 35(2):121–124, March 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2004.
11922063.

Anonymous:2004:FFFc

[Ano04m] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 35(3):213–215, May 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2004.
11922075.

Anonymous:2004:FFFd ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2004.
11922089.

Anonymous:2004:FFFe

[Ano04o] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 35(5):382–384, November 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2004.
11922100.

Anonymous:2004:MHa

[Ano04p] Anonymous. Media highlights. College Mathematics Journal, 35(1):61–71, January 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2004.11922053.

Anonymous:2004:MHb

[Ano04q] Anonymous. Media highlights. College Mathematics Journal, 35(2):152–160, March 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2004.11922067.

Anonymous:2004:MHc

[Ano04r] Anonymous. Media highlights. College Mathematics Journal, 35 (3):233–241, May 2004. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2004.11922078.

Anonymous:2004:MHd

[Ano04s] Anonymous. Media highlights. College Mathematics Journal, 35(4):318–327, September 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2004.11922092.

Anonymous:2004:MHe

[Ano04t] Anonymous. Media highlights. College Mathematics Journal, 35(5):409–417, November 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2004.11922103.

Anonymous:2004:PSa

[Ano04u] Anonymous. Problems and solutions. College Mathematics Journal, 35(1):55–60, January 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2004.11922052.

Anonymous:2004:PSb

[Ano04v] Anonymous. Problems and solutions. College Mathematics Journal, 35(2):136–142, March 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2004.11922065.

Anonymous:2004:PSc

[Ano04w] Anonymous. Problems and solutions. College Mathematics Journal, 35(3):227–232, May 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2004.11922077.

Anonymous:2004:PSd

[Ano04x] Anonymous. Problems and solutions. College Mathematics Journal, 35(4):310–317, September 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2004.
11922091.

Anonymous:2004:PSe

[Ano04y] Anonymous. Problems and solutions. College Mathematics Journal, 35(5):402–408, November 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2004.
11922102.

Anonymous:2004:SRP

[Ano04z] Anonymous. Student research projects. College Mathematics Journal, 35(2):143–151, March 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2004.
11922066.

Anonymous:2005:CCa

[Ano05a] Anonymous. Classroom capsules. College Mathematics Journal, 36(1):52–72, January 2005. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2005.11922111.

Anonymous:2005:CCb

[Ano05b] Anonymous. Classroom capsules. College Mathematics Journal, 36(2):144–159, March 2005. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2005.11922122.

Anonymous:2005:CCc

[Ano05c] Anonymous. Classroom capsules. College Mathematics Journal, 36(3):224–237, May 2005. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2005.11922133.

Anonymous:2005:CCd

[Ano05d] Anonymous. Classroom capsules. College Mathematics Journal, 36(4):318–329, September 2005. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2005.11922145.

Anonymous:2005:CCe

[Ano05e] Anonymous. Classroom capsules. College Mathematics Journal, 36(5):397–412, November 2005. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2005.11922156.

Anonymous:2005:EBa

[Ano05f] Anonymous. Editorial board. College Mathematics Journal, 36 (1):??, January 2005. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2005.11922104.

Anonymous:2005:EBb

[Ano05g] Anonymous. Editorial board. College Mathematics Journal, 36 (2):??, March 2005. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2005.11922114.

Anonymous:2005:EBc

[Ano05h] Anonymous. Editorial board. College Mathematics Journal, 36(3):??, May 2005. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2005.11922125.

Anonymous:2005:EBd

[Ano05i] Anonymous. Editorial board. College Mathematics Journal, 36 (4):??, September 2005. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2005.11922137.

Anonymous:2005:EBe

[Ano05j] Anonymous. Editorial board. College Mathematics Journal, 36 (5):??, November 2005. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2005.11922148.

Anonymous:2005:FFFa

[Ano05k] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 36(1):49–51, January 2005. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2005.
11922110.

Anonymous:2005:FFFb

[Ano05l] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 36(2):141–143, March 2005. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2005.
11922121.

Anonymous:2005:FFFc

[Ano05m] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 36(3):222–223, May 2005. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2005.
11922132.

Anonymous:2005:FFFd

[Ano05n] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 36(4):315–317, September 2005. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2005.
11922144.

Anonymous:2005:FFFe

[Ano05o] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 36(5):394–396, November 2005. CODEN ????

ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2005.
11922155.

Anonymous:2005:MHa

[Ano05p] Anonymous. Media highlights. College Mathematics Journal, 36(1):80–88, January 2005. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2005.11922113.

Anonymous:2005:MHb

[Ano05q] Anonymous. Media highlights. College Mathematics Journal, 36(2):165–173, March 2005. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2005.11922124.

Anonymous:2005:MHc

[Ano05r] Anonymous. Media highlights. College Mathematics Journal, 36 (3):246–254, May 2005. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2005.11922135.

Anonymous:2005:MHd

[Ano05s] Anonymous. Media highlights. College Mathematics Journal, 36(4):341–349, September 2005. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2005.11922147.

Anonymous:2005:MHe

[Ano05t] Anonymous. Media highlights. College Mathematics Journal, 36(5):419–427, November 2005. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2005.11922158.

Anonymous:2005:PSa

[Ano05u] Anonymous. Problems and solutions. College Mathematics Journal, 36(1):73–79, January 2005. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2005.11922112.

Anonymous:2005:PSb

[Ano05v] Anonymous. Problems and solutions. College Mathematics Journal, 36(2):160–164, March 2005. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2005.11922123.

Anonymous:2005:PSc

[Ano05w] Anonymous. Problems and solutions. College Mathematics Journal, 36(3):238–245, May 2005. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2005.11922134.

Anonymous:2005:PSd

[Ano05x] Anonymous. Problems and solutions. College Mathematics Journal, 36(4):330–340, September 2005. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2005.
11922146.

Anonymous:2005:PSe

[Ano05y] Anonymous. Problems and solutions. College Mathematics Journal, 36(5):413–418, November 2005. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2005.
11922157.

Anonymous:2006:CCa

[Ano06a] Anonymous. Classroom capsules. College Mathematics Journal, 37(1):43–53, January 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2006.11922166.

Anonymous:2006:CCb

[Ano06b] Anonymous. Classroom capsules. College Mathematics Journal, 37(2):125–142, March 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2006.11922176.

Anonymous:2006:CCc

[Ano06c] Anonymous. Classroom capsules. College Mathematics Journal, 37(3):218–228, May 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2006.11922188.

Anonymous:2006:CCd

[Ano06d] Anonymous. Classroom capsules. College Mathematics Journal, 37(4):293–307, September 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2006.11922198.

Anonymous:2006:CCe

[Ano06e] Anonymous. Classroom capsules. College Mathematics Journal, 37(5):385–391, November 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2006.11922209.

Anonymous:2006:EBa

[Ano06f] Anonymous. Editorial board. College Mathematics Journal, 37 (1):??, January 2006. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2006.11922159.

Anonymous:2006:EBb

[Ano06g] Anonymous. Editorial board. College Mathematics Journal, 37(3):??, May 2006. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2006.11922181.

Anonymous:2006:EBc

[Ano06h] Anonymous. Editorial board. College Mathematics Journal, 37 (4):??, September 2006. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2006.11922191.

Anonymous:2006:EBd

[Ano06i] Anonymous. Editorial board. College Mathematics Journal, 37 (5):??, November 2006. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2006.11922202.

Anonymous:2006:FFFa

[Ano06j] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 37(1):39–42, January 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2006.
11922165.

Anonymous:2006:FFFb

[Ano06k] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 37(2):121–124, March 2006. CODEN ????

ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2006.
11922175.

Anonymous:2006:FFFc

[Ano06l] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 37(3):214–217, May 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2006.
11922187.

Anonymous:2006:FFFd

[Ano06m] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 37(4):290–292, September 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2006.
11922197.

Anonymous:2006:FFFe

[Ano06n] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 37(5):380–384, November 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2006.
11922208.

Anonymous:2006:GPW

[Ano06o] Anonymous. George pólya winners. College Mathematics Journal, 37(2):165–167, March 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2006.11922180.

Anonymous:2006:MHa

[Ano06p] Anonymous. Media highlights. College Mathematics Journal, 37(1):69–77, January 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2006.11922169.

Anonymous:2006:MHb

[Ano06q] Anonymous. Media highlights. College Mathematics Journal, 37(2):152–160, March 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2006.11922178.

Anonymous:2006:MHc

[Ano06r] Anonymous. Media highlights. College Mathematics Journal, 37 (3):237–245, May 2006. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2006.11922190.

Anonymous:2006:MHd

[Ano06s] Anonymous. Media highlights. College Mathematics Journal, 37(4):319–327, September 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2006.11922200.

Anonymous:2006:MHe

[Ano06t] Anonymous. Media highlights. College Mathematics Journal, 37(5):398–408, November 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2006.11922211.

Anonymous:2006:PSa

[Ano06u] Anonymous. Problems and solutions. College Mathematics Journal, 37(1):59–68, January 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2006.11922168.

Anonymous:2006:PSb

[Ano06v] Anonymous. Problems and solutions. College Mathematics Journal, 37(2):143–151, March 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2006.11922177.

Anonymous:2006:PSc

[Ano06w] Anonymous. Problems and solutions. College Mathematics Journal, 37(3):229–236, May 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2006.11922189.

Anonymous:2006:PSd

[Ano06x] Anonymous. Problems and solutions. College Mathematics Journal, 37(4):308–318, September 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2006.
11922199.

Anonymous:2006:PSe

[Ano06y] Anonymous. Problems and solutions. College Mathematics Journal, 37(5):392–397, November 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2006.
11922210.

Anonymous:2006:SRP

[Ano06z] Anonymous. Student research projects. College Mathematics Journal, 37(1):54–58, January 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2006.
11922167.

Anonymous:2007:CCa

[Ano07a] Anonymous. Classroom capsules. College Mathematics Journal, 38(2):135–147, March 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2007.11922229.

Anonymous:2007:CCb

[Ano07b] Anonymous. Classroom capsules. College Mathematics Journal, 38(3):221–226, May 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2007.11922240.

Anonymous:2007:CCc

[Ano07c] Anonymous. Classroom capsules. College Mathematics Journal, 38(4):300–303, September 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2007.11922252.

Anonymous:2007:CCd

[Ano07d] Anonymous. Classroom capsules. College Mathematics Journal, 38(5):378–387, November 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2007.11922263.

Anonymous:2007:EBa

[Ano07e] Anonymous. Editorial board. College Mathematics Journal, 38 (1):??, January 2007. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2007.11922212.

Anonymous:2007:EBb

[Ano07f] Anonymous. Editorial board. College Mathematics Journal, 38 (2):??, March 2007. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2007.11922220.

Anonymous:2007:EBc

[Ano07g] Anonymous. Editorial board. College Mathematics Journal, 38(3):??, May 2007. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2007.11922232.

Anonymous:2007:EBd

[Ano07h] Anonymous. Editorial board. College Mathematics Journal, 38 (4):??, September 2007. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2007.11922244.

Anonymous:2007:EBe

[Ano07i] Anonymous. Editorial board. College Mathematics Journal, 38 (5):??, November 2007. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2007.11922256.

Anonymous:2007:FFFa

[Ano07j] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 38(1):43–46, January 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2007.
11922217.

Anonymous:2007:FFFb

[Ano07k] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 38(2):131–134, March 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2007.
11922228.

Anonymous:2007:FFFc

[Ano07l] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 38(3):218–220, May 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2007.
11922239.

Anonymous:2007:FFFd

[Ano07m] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 38(4):297–299, September 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2007.
11922251.

Anonymous:2007:FFFe

[Ano07n] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 38(5):375–377, November 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2007.
11922262.

Anonymous:2007:MHa

[Ano07o] Anonymous. Media highlights. College Mathematics Journal, 38(1):68–76, January 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2007.11922219.

Anonymous:2007:MHb

[Ano07p] Anonymous. Media highlights. College Mathematics Journal, 38(2):157–164, March 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2007.11922231.

Anonymous:2007:MHc

[Ano07q] Anonymous. Media highlights. College Mathematics Journal, 38 (3):233–240, May 2007. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2007.11922242.

Anonymous:2007:MHd

[Ano07r] Anonymous. Media highlights. College Mathematics Journal, 38(4):319–327, September 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2007.11922255.

Anonymous:2007:MHe

[Ano07s] Anonymous. Media highlights. College Mathematics Journal, 38(5):397–404, November 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2007.11922265.

Anonymous:2007:PSa

[Ano07t] Anonymous. Problems and solutions. College Mathematics Journal, 38(1):60–67, January 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2007.11922218.

Anonymous:2007:PSb

[Ano07u] Anonymous. Problems and solutions. College Mathematics Journal, 38(2):148–156, March 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2007.11922230.

Anonymous:2007:PSc

[Ano07v] Anonymous. Problems and solutions. College Mathematics Journal, 38(3):227–232, May 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2007.11922241.

Anonymous:2007:PSd

[Ano07w] Anonymous. Problems and solutions. College Mathematics Journal, 38(4):309–318, September 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2007.
11922254.

Anonymous:2007:PSe

[Ano07x] Anonymous. Problems and solutions. College Mathematics Journal, 38(5):388–396, November 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2007.
11922264.

Anonymous:2007:R

[Ano07y] Anonymous. Referees in 2006. College Mathematics Journal, 38(5):409–410, November 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2007.11922267.

Anonymous:2007:SRP

[Ano07z] Anonymous. Student research projects. College Mathematics Journal, 38(4):304–308, September 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2007.
11922253.

Anonymous:2008:CCa

[Ano08a] Anonymous. Classroom capsules. College Mathematics Journal, 39(1):51–64, January 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2008.11922275.

Anonymous:2008:CCb

[Ano08b] Anonymous. Classroom capsules. College Mathematics Journal, 39(2):139–152, March 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2008.11922288.

Anonymous:2008:CCc

[Ano08c] Anonymous. Classroom capsules. College Mathematics Journal, 39(3):230–239, May 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2008.11922298.

Anonymous:2008:CCd

[Ano08d] Anonymous. Classroom capsules. College Mathematics Journal, 39(4):301–306, September 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2008.11922308.

Anonymous:2008:CCe

[Ano08e] Anonymous. Classroom capsules. College Mathematics Journal, 39(5):387–400, November 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2008.11922320.

Anonymous:2008:EBa

[Ano08f] Anonymous. Editorial board. College Mathematics Journal, 39 (1):??, January 2008. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2008.11922268.

Anonymous:2008:EBb

[Ano08g] Anonymous. Editorial board. College Mathematics Journal, 39 (2):??, March 2008. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2008.11922279.

Anonymous:2008:EBc

[Ano08h] Anonymous. Editorial board. College Mathematics Journal, 39(3):??, May 2008. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2008.11922291.

Anonymous:2008:EBd

[Ano08i] Anonymous. Editorial board. College Mathematics Journal, 39 (4):??, September 2008. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2008.11922301.

Anonymous:2008:EBe

[Ano08j] Anonymous. Editorial board. College Mathematics Journal, 39 (5):??, November 2008. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2008.11922312.

Anonymous:2008:FFFa

[Ano08k] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 39(1):49–50, January 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2008.
11922274.

Anonymous:2008:FFFb

[Ano08l] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 39(2):136–138, March 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2008.
11922287.

Anonymous:2008:FFFc

[Ano08m] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 39(3):227–229, May 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2008.
11922297.

Anonymous:2008:FFFd ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2008.
11922307.

Anonymous:2008:FFFe

[Ano08o] Anonymous. Fallacies, flaws, and flimflam. College Mathematics Journal, 39(5):380–386, November 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2008.
11922319.

Anonymous:2008:GPAa

[Ano08p] Anonymous. George pólya awards. College Mathematics Journal, 39(1):83–86, January 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2008.11922278.

Anonymous:2008:GPAb

[Ano08q] Anonymous. George pólya awards. College Mathematics Journal, 39(5):422–424, November 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2008.11922324.

Anonymous:2008:MHa

[Ano08r] Anonymous. Media highlights. College Mathematics Journal, 39(1):74–82, January 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2008.11922277.

Anonymous:2008:MHb

[Ano08s] Anonymous. Media highlights. College Mathematics Journal, 39(2):162–170, March 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2008.11922290.

Anonymous:2008:MHc

[Ano08t] Anonymous. Media highlights. College Mathematics Journal, 39 (3):248–256, May 2008. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2008.11922300.

Anonymous:2008:MHd

[Ano08u] Anonymous. Media highlights. College Mathematics Journal, 39(4):315–323, September 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2008.11922310.

Anonymous:2008:MHe

[Ano08v] Anonymous. Media highlights. College Mathematics Journal, 39(5):410–418, November 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2008.11922322.

Anonymous:2008:PSa

[Ano08w] Anonymous. Problems and solutions. College Mathematics Journal, 39(1):65–73, January 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2008.11922276.

Anonymous:2008:PSb

[Ano08x] Anonymous. Problems and solutions. College Mathematics Journal, 39(2):153–161, March 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2008.11922289.

Anonymous:2008:PSc

[Ano08y] Anonymous. Problems and solutions. College Mathematics Journal, 39(3):240–247, May 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2008.11922299.

Anonymous:2008:PSd

[Ano08z] Anonymous. Problems and solutions. College Mathematics Journal, 39(4):307–314, September 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2008.
11922309.

Anonymous:2008:PSe

[Ano08-27] Anonymous. Problems and solutions. College Mathematics Journal, 39(5):401–409, November 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2008.
11922321.

Anonymous:2008:R

[Ano08-28] Anonymous. Referees in 2007. College Mathematics Journal, 39(5):425–427, November 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2008.11922325.

Anonymous:2009:ACE

[Ano09a] Anonymous. Additions, corrections, emendations, and revisions. College Mathematics Journal, 40(5):399, November 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2009.11922393.

Anonymous:2009:CCa

[Ano09b] Anonymous. Classroom capsules. College Mathematics Journal, 40(1):40–46, January 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2009.11922335.

Anonymous:2009:CCb

[Ano09c] Anonymous. Classroom capsules. College Mathematics Journal, 40(2):125–130, March 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2009.11922348.

Anonymous:2009:CCc

[Ano09d] Anonymous. Classroom capsules. College Mathematics Journal, 40(3):213–214, May 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2009.11922364.

Anonymous:2009:CCd

[Ano09e] Anonymous. Classroom capsules. College Mathematics Journal, 40(4):284–292, September 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2009.11922377.

Anonymous:2009:EBa

[Ano09f] Anonymous. Editorial board. College Mathematics Journal, 40 (1):??, January 2009. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2009.11922326.

Anonymous:2009:EBb

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2009.11922340.

Anonymous:2009:EBc

[Ano09h] Anonymous. Editorial board. College Mathematics Journal, 40(3):??, May 2009. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2009.11922352.

Anonymous:2009:EBd

[Ano09i] Anonymous. Editorial board. College Mathematics Journal, 40 (4):??, September 2009. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2009.11922369.

Anonymous:2009:EBe

[Ano09j] Anonymous. Editorial board. College Mathematics Journal, 40 (5):??, November 2009. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2009.11922381.

Anonymous:2009:IMG

[Ano09k] Anonymous. Interview with Martin Gardner. College Mathematics Journal, 40(3):158–161, May 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2009.
11922353.

Anonymous:2009:MHa

[Ano09l] Anonymous. Media highlights. College Mathematics Journal, 40(1):68–76, January 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2009.11922339.

Anonymous:2009:MHb

[Ano09m] Anonymous. Media highlights. College Mathematics Journal, 40(2):146–155, March 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2009.11922351.

Anonymous:2009:MHc

[Ano09n] Anonymous. Media highlights. College Mathematics Journal, 40 (3):228–236, May 2009. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2009.11922368.

Anonymous:2009:MHd

[Ano09o] Anonymous. Media highlights. College Mathematics Journal, 40(4):313–320, September 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2009.11922380.

Anonymous:2009:MHe

[Ano09p] Anonymous. Media highlights. College Mathematics Journal, 40(5):387–396, November 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/074683409X475733.

Anonymous:2009:PSa

[Ano09q] Anonymous. Problems and solutions. College Mathematics Journal, 40(1):55–64, January 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2009.11922337.

Anonymous:2009:PSb

[Ano09r] Anonymous. Problems and solutions. College Mathematics Journal, 40(2):131–141, March 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2009.11922349.

Anonymous:2009:PSc

[Ano09s] Anonymous. Problems and solutions. College Mathematics Journal, 40(3):215–221, May 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2009.11922365.

Anonymous:2009:PSd

[Ano09t] Anonymous. Problems and solutions. College Mathematics Journal, 40(4):293–305, September 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2009.
11922378.

Anonymous:2009:PSe

[Ano09u] Anonymous. Problems and solutions. College Mathematics Journal, 40(5):376–383, November 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.

tandfonline.com/doi/abs/10.4169/074683409X475715.

Anonymous:2009:R

[Ano09v] Anonymous. Referees in 2008. College Mathematics Journal, 40(5):397–398, November 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2009.11922392.

Anonymous:2009:SPB

[Ano09w] Anonymous. Solutions to puzzles and brainteasers. College Mathematics Journal, 40(3):226–227, May 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2009.
11922367.

Anonymous:2009:SRP

[Ano09x] Anonymous. Student research projects. College Mathematics Journal, 40(2):119–124, March 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2009.
11922347.

Anonymous:2010:ACE

[Ano10a] Anonymous. Additions, corrections, emendations, and revisions. College Mathematics Journal, 41(5):430–431, November 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2010.11922464.

Anonymous:2010:EBa

[Ano10b] Anonymous. Editorial board. College Mathematics Journal, 41 (1):??, January 2010. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2010.11922394.

Anonymous:2010:EBb

[Ano10c] Anonymous. Editorial board. College Mathematics Journal, 41 (2):??, March 2010. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468384.2010.11922407.

Anonymous:2010:EBc

[Ano10d] Anonymous. Editorial board. College Mathematics Journal, 41(3):??, May 2010. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2010.11922420.

Anonymous:2010:EBd

[Ano10e] Anonymous. Editorial board. College Mathematics Journal, 41 (4):??, September 2010. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2010.11922435.

Anonymous:2010:EBe

[Ano10f] Anonymous. Editorial board. College Mathematics Journal, 41 (5):??, November 2010. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2010.11922447.

Anonymous:2010:MHa

[Ano10g] Anonymous. Media highlights. College Mathematics Journal, 41(1):77–85, January 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/074683410X475173.

Anonymous:2010:MHb

[Ano10h] Anonymous. Media highlights. College Mathematics Journal, 41(2):173–180, March 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/074683410X480302.

Anonymous:2010:MHc

[Ano10i] Anonymous. Media highlights. College Mathematics Journal, 41 (3):252–260, May 2010. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/074683410X488764.

Anonymous:2010:MHd

[Ano10j] Anonymous. Media highlights. College Mathematics Journal, 41(4):341–348, September 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/074683410X510353.

Anonymous:2010:MHe

[Ano10k] Anonymous. Media highlights. College Mathematics Journal, 41(5):419–427, November 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/074683410X522062.

Anonymous:2010:PSa

[Ano10l] Anonymous. Problems and solutions. College Mathematics Journal, 41(1):64–71, January 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/074683410X475146.

Anonymous:2010:PSb

[Ano10m] Anonymous. Problems and solutions. College Mathematics Journal, 41(2):160–169, March 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/074683410X480285.

Anonymous:2010:PSc

[Ano10n] Anonymous. Problems and solutions. College Mathematics Journal, 41(3):242–247, May 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/074683410X488746.

Anonymous:2010:PSd

[Ano10o] Anonymous. Problems and solutions. College Mathematics Journal, 41(4):329–338, September 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.

tandfonline.com/doi/abs/10.4169/074683410X510335.

Anonymous:2010:PSe

[Ano10p] Anonymous. Problems and solutions. College Mathematics Journal, 41(5):409–415, November 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.

tandfonline.com/doi/abs/10.4169/074683410X522044.

Anonymous:2010:R

[Ano10q] Anonymous. Referees in 2010. College Mathematics Journal, 41(5):428–429, November 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2010.11922463.

Anonymous:2011:ACE

[Ano11a] Anonymous. Additions, corrections, emendations, and revisions. College Mathematics Journal, 42(5):429–431, November 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.42.5.429.

Anonymous:2011:ABE

[Ano11b] Anonymous. Anne Burns explains how Iterated Möbius was constructed. College Mathematics Journal, 42(1):14, January 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2011.11922468.

Anonymous:2011:EBa

[Ano11c] Anonymous. Editorial board. College Mathematics Journal, 42 (1):??, January 2011. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2011.11922465.

Anonymous:2011:EBb

[Ano11d] Anonymous. Editorial board. College Mathematics Journal, 42 (2):??, March 2011. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2011.11922481.

Anonymous:2011:EBc

[Ano11e] Anonymous. Editorial board. College Mathematics Journal, 42(3):??, May 2011. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2011.11922497.

Anonymous:2011:EBd

[Ano11f] Anonymous. Editorial board. College Mathematics Journal, 42 (4):??, September 2011. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2011.11922511.

Anonymous:2011:EBe

[Ano11g] Anonymous. Editorial board. College Mathematics Journal, 42 (5):??, November 2011. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.42.5.fm.

Anonymous:2011:GPA

[Ano11h] Anonymous. George pólya awards for 2010. College Mathematics Journal, 42(5):426, November 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
42.5.426.

Anonymous:2011:MHa

[Ano11i] Anonymous. Media highlights. College Mathematics Journal, 42(1):76–84, January 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.42.1.076.

Anonymous:2011:MHb

[Ano11j] Anonymous. Media highlights. College Mathematics Journal, 42(2):163–172, March 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.42.2.163.

Anonymous:2011:MHc

[Ano11k] Anonymous. Media highlights. College Mathematics Journal, 42 (3):244–252, May 2011. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.42.3.244.

Anonymous:2011:MHd

[Ano11l] Anonymous. Media highlights. College Mathematics Journal, 42(4):339–346, September 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.42.4.339.

Anonymous:2011:MHe

[Ano11m] Anonymous. Media highlights. College Mathematics Journal, 42(5):416–425, November 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.42.5.416.

Anonymous:2011:PSa

[Ano11n] Anonymous. Problems and solutions. College Mathematics Journal, 42(1):62–70, January 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.42.1.062.

Anonymous:2011:PSb

[Ano11o] Anonymous. Problems and solutions. College Mathematics Journal, 42(2):151–158, March 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.42.2.151.

Anonymous:2011:PSc

[Ano11p] Anonymous. Problems and solutions. College Mathematics Journal, 42(3):232–240, May 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.42.3.232.

Anonymous:2011:PSd

[Ano11q] Anonymous. Problems and solutions. College Mathematics Journal, 42(4):329–335, September 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
42.4.329.

Anonymous:2011:PSe

[Ano11r] Anonymous. Problems and solutions. College Mathematics Journal, 42(5):407–412, November 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
42.5.407.

Anonymous:2011:R

[Ano11s] Anonymous. Referees in 2011. College Mathematics Journal, 42(5):427–429, November 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.42.5.427.

Anonymous:2012:ACE

[Ano12a] Anonymous. Additions, corrections, emendations, and revisions. College Mathematics Journal, 43(5):429–430, November 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.43.5.429.

Anonymous:2012:EBa

[Ano12b] Anonymous. Editorial board. College Mathematics Journal, 43 (1):??, January 2012. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2012.11922549.

Anonymous:2012:EBb

[Ano12c] Anonymous. Editorial board. College Mathematics Journal, 43 (2):??, March 2012. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2012.11922568.

Anonymous:2012:EBc

[Ano12d] Anonymous. Editorial board. College Mathematics Journal, 43(3):??, May 2012. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2012.11922584.

Anonymous:2012:EBd

[Ano12e] Anonymous. Editorial board. College Mathematics Journal, 43 (4):??, September 2012. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2012.11922600.

Anonymous:2012:EBe

[Ano12f] Anonymous. Editorial board. College Mathematics Journal, 43 (5):??, November 2012. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2012.11922617.

Anonymous:2012:GPA

[Ano12g] Anonymous. George Pólya Awards for 2011. College Mathematics Journal, 43(5):431, November 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2012.
11922636.

Anonymous:2012:MHa

[Ano12h] Anonymous. Media highlights. College Mathematics Journal, 43(1):105–112, January 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.43.1.105.

Anonymous:2012:MHb

[Ano12i] Anonymous. Media highlights. College Mathematics Journal, 43(2):184–192, March 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.43.2.184.

Anonymous:2012:MHc

[Ano12j] Anonymous. Media highlights. College Mathematics Journal, 43 (3):265–272, May 2012. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.43.3.265.

Anonymous:2012:MHd

[Ano12k] Anonymous. Media highlights. College Mathematics Journal, 43(4):345–352, September 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.43.4.345.

Anonymous:2012:MHe

[Ano12l] Anonymous. Media highlights. College Mathematics Journal, 43(5):420–427, November 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.43.5.420.

Anonymous:2012:PSa

[Ano12m] Anonymous. Problems and solutions. College Mathematics Journal, 43(1):95–104, January 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.43.1.095.

Anonymous:2012:PSb

[Ano12n] Anonymous. Problems and solutions. College Mathematics Journal, 43(2):174–180, March 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.43.2.174.

Anonymous:2012:PSc

[Ano12o] Anonymous. Problems and solutions. College Mathematics Journal, 43(3):257–264, May 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.43.3.257.

Anonymous:2012:PSd

[Ano12p] Anonymous. Problems and solutions. College Mathematics Journal, 43(4):337–344, September 2012. CODEN ????

ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
43.4.337.

Anonymous:2012:PSe

[Ano12q] Anonymous. Problems and solutions. College Mathematics Journal, 43(5):410–419, November 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
43.5.410.

Anonymous:2012:R

[Ano12r] Anonymous. Referees in 2012. College Mathematics Journal, 43(5):428–429, November 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.43.5.428.

Anonymous:2013:ACE

[Ano13a] Anonymous. Additions, corrections, emendations, and revisions. College Mathematics Journal, 44(5):458–460, November 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.44.5.458.

Anonymous:2013:EBa

[Ano13b] Anonymous. Editorial board. College Mathematics Journal, 44 (1):??, January 2013. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2013.11922637.

Anonymous:2013:EBb

[Ano13c] Anonymous. Editorial board. College Mathematics Journal, 44 (2):??, March 2013. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2013.11922655.

Anonymous:2013:EBc

[Ano13d] Anonymous. Editorial board. College Mathematics Journal, 44(3):??, May 2013. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2013.11922671.

Anonymous:2013:EBd

[Ano13e] Anonymous. Editorial board. College Mathematics Journal, 44 (4):??, September 2013. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2013.11922689.

Anonymous:2013:EBe

[Ano13f] Anonymous. Editorial board. College Mathematics Journal, 44 (5):??, November 2013. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2013.11922704.

Anonymous:2013:MHa

[Ano13g] Anonymous. Media highlights. College Mathematics Journal, 44(1):73–80, January 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.44.1.073.

Anonymous:2013:MHb

[Ano13h] Anonymous. Media highlights. College Mathematics Journal, 44(2):156–164, March 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.44.2.156.

Anonymous:2013:MHc

[Ano13i] Anonymous. Media highlights. College Mathematics Journal, 44 (3):248–255, May 2013. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.44.3.248.

Anonymous:2013:MHd

[Ano13j] Anonymous. Media highlights. College Mathematics Journal, 44(4):336–344, September 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.44.4.336.

Anonymous:2013:MHe

[Ano13k] Anonymous. Media highlights. College Mathematics Journal, 44(5):450–456, November 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.44.5.450.

Anonymous:2013:PSa

[Ano13l] Anonymous. Problems and solutions. College Mathematics Journal, 44(1):65–72, January 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.44.1.65.

Anonymous:2013:PSb

[Ano13m] Anonymous. Problems and solutions. College Mathematics Journal, 44(2):142–148, March 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.44.2.142.

Anonymous:2013:PSc

[Ano13n] Anonymous. Problems and solutions. College Mathematics Journal, 44(3):233–240, May 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.44.3.233.

Anonymous:2013:PSd

[Ano13o] Anonymous. Problems and solutions. College Mathematics Journal, 44(4):325–331, September 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
44.4.325.

Anonymous:2013:PSe

[Ano13p] Anonymous. Problems and solutions. College Mathematics Journal, 44(5):437–445, November 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
44.5.437.

Anonymous:2013:R

[Ano13q] Anonymous. Referees in 2013. College Mathematics Journal, 44(5):457–458, November 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.44.5.457.

Anonymous:2014:CPS

[Ano14a] Anonymous. Crossword puzzle solution. College Mathematics Journal, 45(4):332–333, September 2014. CODEN ????

ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
45.4.332.

Anonymous:2014:EBa

[Ano14b] Anonymous. Editorial board. College Mathematics Journal, 45 (1):??, January 2014. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2014.11922723.

Anonymous:2014:EBb

[Ano14c] Anonymous. Editorial board. College Mathematics Journal, 45 (2):??, March 2014. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2014.11922741.

Anonymous:2014:EBc

[Ano14d] Anonymous. Editorial board. College Mathematics Journal, 45(3):??, May 2014. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2014.11922758.

Anonymous:2014:EBd

[Ano14e] Anonymous. Editorial board. College Mathematics Journal, 45 (4):??, September 2014. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2014.11922774.

Anonymous:2014:EBe

[Ano14f] Anonymous. Editorial board. College Mathematics Journal, 45 (5):??, November 2014. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2014.11922792.

Anonymous:2014:FR

[Ano14g] Anonymous. Film review. College Mathematics Journal, 45(1): 65–70, January 2014. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.45.1.065.

Anonymous:2014:GPAa

[Ano14h] Anonymous. George pólya awards for 2013. College Mathematics Journal, 45(1):79, January 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.45.1.079.

Anonymous:2014:GPAb

[Ano14i] Anonymous. George pólya awards for 2014. College Mathematics Journal, 45(5):414–415, November 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
45.5.414.

Anonymous:2014:MHa

[Ano14j] Anonymous. Media highlights. College Mathematics Journal, 45(1):71–78, January 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.45.1.071.

Anonymous:2014:MHb

[Ano14k] Anonymous. Media highlights. College Mathematics Journal, 45(2):152–159, March 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.45.2.152.

Anonymous:2014:MHc

[Ano14l] Anonymous. Media highlights. College Mathematics Journal, 45 (3):232–239, May 2014. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.45.3.232.

Anonymous:2014:MHd

[Ano14m] Anonymous. Media highlights. College Mathematics Journal, 45(4):324–331, September 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.45.4.324.

Anonymous:2014:MHe

[Ano14n] Anonymous. Media highlights. College Mathematics Journal, 45(5):406–413, November 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.45.5.406.

Anonymous:2014:PSa

[Ano14o] Anonymous. Problems and solutions. College Mathematics Journal, 45(1):57–64, January 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.45.1.057.

Anonymous:2014:PSb

[Ano14p] Anonymous. Problems and solutions. College Mathematics Journal, 45(2):145–151, March 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.45.2.145.

Anonymous:2014:PSc

[Ano14q] Anonymous. Problems and solutions. College Mathematics Journal, 45(3):222–229, May 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.45.3.222.

Anonymous:2014:PSd

[Ano14r] Anonymous. Problems and solutions. College Mathematics Journal, 45(4):318–323, September 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
45.4.318.

Anonymous:2014:PSe

[Ano14s] Anonymous. Problems and solutions. College Mathematics Journal, 45(5):393–399, November 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
45.5.393.

Anonymous:2014:R

[Ano14t] Anonymous. Referees in 2014. College Mathematics Journal, 45(5):415–416, November 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.45.5.415.

Anonymous:2015:BRE

[Ano15a] Anonymous. Book review: Excerpts from MAA Interviews. College Mathematics Journal, 46(5):377–384, November 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.46.5.377.

Anonymous:2015:EBa

[Ano15b] Anonymous. Editorial board. College Mathematics Journal, 46 (1):??, January 2015. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2015.11922812.

Anonymous:2015:EBb

[Ano15c] Anonymous. Editorial board. College Mathematics Journal, 46 (2):??, March 2015. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2015.11922830.

Anonymous:2015:EBc

[Ano15d] Anonymous. Editorial board. College Mathematics Journal, 46(3):??, May 2015. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2015.11922848.

Anonymous:2015:EBd

[Ano15e] Anonymous. Editorial board. College Mathematics Journal, 46 (4):??, September 2015. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2015.11922863.

Anonymous:2015:EBe

[Ano15f] Anonymous. Editorial board. College Mathematics Journal, 46 (5):??, November 2015. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.46.5.fm.

Anonymous:2015:GPA

[Ano15g] Anonymous. George pólya awards for 2015. College Mathematics Journal, 46(5):393–394, November 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
46.5.393.

Anonymous:2015:MAC

[Ano15h] Anonymous. MAA 100th anniversary CMJ puzzles solutions. College Mathematics Journal, 46(4):323, September 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.46.4.323.

Anonymous:2015:MHa

[Ano15i] Anonymous. Media highlights. College Mathematics Journal, 46(1):73–80, January 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.46.1.73.

Anonymous:2015:MHb

[Ano15j] Anonymous. Media highlights. College Mathematics Journal, 46(2):153–160, March 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.46.2.153.

Anonymous:2015:MHc

[Ano15k] Anonymous. Media highlights. College Mathematics Journal, 46 (3):233–240, May 2015. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.46.3.233.

Anonymous:2015:MHd

[Ano15l] Anonymous. Media highlights. College Mathematics Journal, 46(4):315–322, September 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.46.4.315.

Anonymous:2015:MHe

[Ano15m] Anonymous. Media highlights. College Mathematics Journal, 46(5):385–392, November 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.46.5.385.

Anonymous:2015:NV

[Ano15n] Anonymous. Notes on volume 45. College Mathematics Journal, 46(5):395–397, November 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.46.5.395.

Anonymous:2015:PSa

[Ano15o] Anonymous. Problems and solutions. College Mathematics Journal, 46(1):61–66, January 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.46.1.61.

Anonymous:2015:PSb

[Ano15p] Anonymous. Problems and solutions. College Mathematics Journal, 46(2):142–148, March 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.46.2.142.

Anonymous:2015:PSc

[Ano15q] Anonymous. Problems and solutions. College Mathematics Journal, 46(3):220–227, May 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.46.3.220.

Anonymous:2015:PSd

[Ano15r] Anonymous. Problems and solutions. College Mathematics Journal, 46(4):300–308, September 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
46.4.300.

Anonymous:2015:PSe

[Ano15s] Anonymous. Problems and solutions. College Mathematics Journal, 46(5):369–376, November 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
46.5.369.

Anonymous:2015:R

[Ano15t] Anonymous. Referees in 2015. College Mathematics Journal, 46(5):398–399, November 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.46.5.398.

Anonymous:2016:EBa

[Ano16a] Anonymous. Editorial board. College Mathematics Journal, 47 (1):??, January 2016. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2016.11922896.

Anonymous:2016:EBb

[Ano16b] Anonymous. Editorial board. College Mathematics Journal, 47 (2):??, March 2016. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2016.11922910.

Anonymous:2016:EBc

[Ano16c] Anonymous. Editorial board. College Mathematics Journal, 47(3):??, May 2016. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2016.11922929.

Anonymous:2016:EBd

[Ano16d] Anonymous. Editorial board. College Mathematics Journal, 47 (4):??, September 2016. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2016.11922947.

Anonymous:2016:EBe

[Ano16e] Anonymous. Editorial board. College Mathematics Journal, 47 (5):??, November 2016. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2016.11922965.

Anonymous:2016:GPA

[Ano16f] Anonymous. George pólya awards for 2016. College Mathematics Journal, 47(5):394–395, November 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
47.5.394.

Anonymous:2016:MHa

[Ano16g] Anonymous. Media highlights. College Mathematics Journal, 47(1):73–80, January 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.47.1.73.

Anonymous:2016:MHb

[Ano16h] Anonymous. Media highlights. College Mathematics Journal, 47(2):153–160, March 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.47.2.153.

Anonymous:2016:MHc

[Ano16i] Anonymous. Media highlights. College Mathematics Journal, 47 (3):232–239, May 2016. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.47.3.232.

Anonymous:2016:MHd

[Ano16j] Anonymous. Media highlights. College Mathematics Journal, 47(4):312–319, September 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.47.4.312.

Anonymous:2016:MHe

[Ano16k] Anonymous. Media highlights. College Mathematics Journal, 47(5):386–393, November 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.47.5.386.

Anonymous:2016:NV

[Ano16l] Anonymous. Notes on volume 46. College Mathematics Journal, 47(5):396–397, November 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.47.5.396.

Anonymous:2016:PSa

[Ano16m] Anonymous. Problems and solutions. College Mathematics Journal, 47(1):61–66, January 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.47.1.61.

Anonymous:2016:PSb

[Ano16n] Anonymous. Problems and solutions. College Mathematics Journal, 47(2):138–145, March 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.47.2.138.

Anonymous:2016:PSc

[Ano16o] Anonymous. Problems and solutions. College Mathematics Journal, 47(3):221–227, May 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.47.3.221.

Anonymous:2016:PSd

[Ano16p] Anonymous. Problems and solutions. College Mathematics Journal, 47(4):300–306, September 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
47.4.300.

Anonymous:2016:PSe

[Ano16q] Anonymous. Problems and solutions. College Mathematics Journal, 47(5):369–374, November 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
47.5.369.

Anonymous:2016:R

[Ano16r] Anonymous. Referees in 2016. College Mathematics Journal, 47(5):398–399, November 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.47.5.398.

Anonymous:2017:EBa

[Ano17a] Anonymous. Editorial board. College Mathematics Journal, 48 (1):??, January 2017. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2017.11922984.

Anonymous:2017:EBb

[Ano17b] Anonymous. Editorial board. College Mathematics Journal, 48 (2):??, March 2017. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2017.11923002.

Anonymous:2017:EBc

[Ano17c] Anonymous. Editorial board. College Mathematics Journal, 48(3):??, May 2017. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2017.11923017.

Anonymous:2017:EBd

[Ano17d] Anonymous. Editorial board. College Mathematics Journal, 48 (4):??, September 2017. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2017.11919677.

Anonymous:2017:EBe

[Ano17e] Anonymous. Editorial board. College Mathematics Journal, 48 (5):??, November 2017. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.48.5.fm.

Anonymous:2017:GPA

[Ano17f] Anonymous. George pólya awards for 2017. College Mathematics Journal, 48(5):391–392, November 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
48.5.391.

Anonymous:2017:MHa

[Ano17g] Anonymous. Media highlights. College Mathematics Journal, 48(1):73–80, January 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.48.1.73.

Anonymous:2017:MHb

[Ano17h] Anonymous. Media highlights. College Mathematics Journal, 48(2):153–160, March 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.48.2.153.

Anonymous:2017:MHc

[Ano17i] Anonymous. Media highlights. College Mathematics Journal, 48 (3):233–240, May 2017. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.48.3.233.

Anonymous:2017:MHd

[Ano17j] Anonymous. Media highlights. College Mathematics Journal, 48(4):313–320, September 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.48.4.313.

Anonymous:2017:MHe

[Ano17k] Anonymous. Media highlights. College Mathematics Journal, 48(5):383–390, November 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.48.5.383.

Anonymous:2017:MHf

[Ano17l] Anonymous. Media highlights. College Mathematics Journal, 49(1):73–80, January 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2017.1397976.

Anonymous:2017:MHg

[Ano17m] Anonymous. Media highlights. College Mathematics Journal, 49(2):153–160, March 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2017.1419010.

Anonymous:2017:NV

[Ano17n] Anonymous. Notes on volume 47. College Mathematics Journal, 48(5):393–396, November 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.48.5.393.

Anonymous:2017:PSa

[Ano17o] Anonymous. Problems and solutions. College Mathematics Journal, 48(1):58–63, January 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.48.1.58.

Anonymous:2017:PSb

[Ano17p] Anonymous. Problems and solutions. College Mathematics Journal, 48(2):138–145, March 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.48.2.138.

Anonymous:2017:PSc

[Ano17q] Anonymous. Problems and solutions. College Mathematics Journal, 48(3):219–225, May 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.48.3.219.

Anonymous:2017:PSd

[Ano17r] Anonymous. Problems and solutions. College Mathematics Journal, 48(4):292–301, September 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
48.4.292.

Anonymous:2017:PSe

[Ano17s] Anonymous. Problems and solutions. College Mathematics Journal, 48(5):370–377, November 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
48.5.370.

Anonymous:2017:PSf

[Ano17t] Anonymous. Problems and solutions. College Mathematics Journal, 49(1):60–67, January 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2017.1403145.

Anonymous:2017:R

[Ano17u] Anonymous. Referees in 2017. College Mathematics Journal, 48(5):397–398, November 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.48.5.397.

Anonymous:2018:ACa

[Ano18a] Anonymous. About the cover. College Mathematics Journal, 49 (1):??, January 2018. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2018.1411080.

Anonymous:2018:ACb

[Ano18b] Anonymous. About the cover. College Mathematics Journal, 49(2):??, March 2018. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2018.1432208.

Anonymous:2018:ACc

[Ano18c] Anonymous. About the cover. College Mathematics Journal, 49(3):??, May 2018. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2018.1460541.

Anonymous:2018:ACd

[Ano18d] Anonymous. About the cover. College Mathematics Journal, 49 (4):??, September 2018. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2018.1516962.

Anonymous:2018:ACe

[Ano18e] Anonymous. About the cover. College Mathematics Journal, 49 (5):??, November 2018. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2018.1540687.

Anonymous:2018:GPA

[Ano18f] Anonymous. George pólya awards for 2018. College Mathematics Journal, 49(5):392–394, November 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2018.
1483670.

Anonymous:2018:MHa

[Ano18g] Anonymous. Media highlights. College Mathematics Journal, 49 (3):233–240, May 2018. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2018.1436840.

Anonymous:2018:MHb

[Ano18h] Anonymous. Media highlights. College Mathematics Journal, 49(4):313–320, September 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2018.1483665.

Anonymous:2018:NV

[Ano18i] Anonymous. Notes on volume 48. College Mathematics Journal, 49(5):395–398, November 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2018.1528806.

Anonymous:2018:PSa

[Ano18j] Anonymous. Problems and solutions. College Mathematics Journal, 49(2):140–146, March 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2018.1432203.

Anonymous:2018:PSb

[Ano18k] Anonymous. Problems and solutions. College Mathematics Journal, 49(3):222–228, May 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2018.1445359.

Anonymous:2018:PSc

[Ano18l] Anonymous. Problems and solutions. College Mathematics Journal, 49(4):295–301, September 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2018.
1503483.

Anonymous:2018:R

[Ano18m] Anonymous. Referees in 2018. College Mathematics Journal, 49(5):399–400, November 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2018.1512307.

Anonymous:2019:ACa

[Ano19a] Anonymous. About the cover. College Mathematics Journal, 50 (1):??, January 2019. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2019.1570746.

Anonymous:2019:ACb

[Ano19b] Anonymous. About the cover. College Mathematics Journal, 50 (2):??, 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2019.1584504.

Anonymous:2019:ACc

[Ano19c] Anonymous. About the cover. College Mathematics Journal, 50 (3):??, 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2019.1610310.

Anonymous:2019:ACd

[Ano19d] Anonymous. About the cover. College Mathematics Journal, 50 (4):??, 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2019.1666605.

Anonymous:2019:ACe

[Ano19e] Anonymous. About the cover. College Mathematics Journal, 50 (5):??, 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2019.1685274.

Anonymous:2019:E

[Ano19f] Anonymous. Erratum. College Mathematics Journal, 50(5): 399, 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2018.1426944.

Anonymous:2019:GPA

[Ano19g] Anonymous. George Pólya Awards for 2019. College Mathematics Journal, 50(5):397–398, 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2019.1667196.

Anonymous:2019:MHa

[Ano19h] Anonymous. Media highlights. College Mathematics Journal, 50(1):73–80, January 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2019.1537454.

Anonymous:2019:MHb

[Ano19i] Anonymous. Media highlights. College Mathematics Journal, 50(2):153–160, 2019. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2019.1560204.

Anonymous:2019:R

[Ano19j] Anonymous. Referees in 2019. College Mathematics Journal, 50 (5):??, 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2019.1664861.

Anonymous:2020:ACa

[Ano20a] Anonymous. About the cover. College Mathematics Journal, 51 (1):??, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2020.1697093.

Anonymous:2020:ACb

[Ano20b] Anonymous. About the cover. College Mathematics Journal, 51 (2):??, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2020.1727720.

Anonymous:2020:ACc

[Ano20c] Anonymous. About the cover. College Mathematics Journal, 51 (3):??, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2020.1753466.

Anonymous:2020:ACd

[Ano20d] Anonymous. About the cover. College Mathematics Journal, 51 (4):??, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2020.1823773.

Anonymous:2020:ACe

[Ano20e] Anonymous. About the cover. College Mathematics Journal, 51 (5):??, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2020.1841532.

Anonymous:2020:GPA

[Ano20f] Anonymous. George Pólya Awards for 2020. College Mathematics Journal, 51(5):369–370, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2020.1820775.

Anonymous:2020:LE

[Ano20g] Anonymous. Letter from the Editor. College Mathematics Journal, 51(1):2, 2020. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2020.1681794.

Anonymous:2020:R

[Ano20h] Anonymous. Referees in 2020. College Mathematics Journal, 51 (5):??, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2020.1847922.

Anonymous:2021:ACa

[Ano21a] Anonymous. About the cover. College Mathematics Journal, 52 (1):??, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2021.1875771.

Anonymous:2021:ACb

[Ano21b] Anonymous. About the cover. College Mathematics Journal, 52 (2):??, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2021.1899502.

Anonymous:2021:ACc

[Ano21c] Anonymous. About the cover. College Mathematics Journal, 52 (3):??, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2021.1927440.

Anonymous:2021:ACd

[Ano21d] Anonymous. About the cover. College Mathematics Journal, 52 (4):??, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2021.1965436.

Anonymous:2021:ACe

[Ano21e] Anonymous. About the cover. College Mathematics Journal, 52 (5):??, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2021.1988483.

Anonymous:2021:Ca

[Ano21f] Anonymous. Correction. College Mathematics Journal, 52(3): 241, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2021.1886555.

Anonymous:2021:Cb

[Ano21g] Anonymous. Correction. College Mathematics Journal, 52(4): ??, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2021.1924519.

Anonymous:2021:GPA

[Ano21h] Anonymous. George Pólya awards for 2021. College Mathematics Journal, 52(5):385–387, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2021.1974747.

Anonymous:2021:R

[Ano21i] Anonymous. Referees in 2021. College Mathematics Journal, 52 (5):??, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2021.1988719.

Anonymous:2022:ACa

[Ano22a] Anonymous. About the cover. College Mathematics Journal, 53 (2):??, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2022.2049107.

Anonymous:2022:ACb

[Ano22b] Anonymous. About the cover. College Mathematics Journal, 53 (5):??, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2022.2143743.

Anonymous:2022:C

[Ano22c] Anonymous. Correction. College Mathematics Journal, 53(4): 334, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2022.2065166.

Anonymous:2022:GPA

[Ano22d] Anonymous. George Pólya Awards for 2022. College Mathematics Journal, 53(5):347–350, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2022.2122670.

Anonymous:2022:MH

[Ano22e] Anonymous. Media highlights. College Mathematics Journal, 53(4):326–333, 2022. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2022.2100156.

Anonymous:2022:R

[Ano22f] Anonymous. Referees in 2022. College Mathematics Journal, 53(5):416–417, 2022. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2022.2142017.

Anonymous:2023:ACa

[Ano23a] Anonymous. About the cover. College Mathematics Journal, 54 (1):??, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2023.2180929.

Anonymous:2023:ACb

[Ano23b] Anonymous. About the cover. College Mathematics Journal, 54 (2):??, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2023.2201123.

Anonymous:2023:ACc

[Ano23c] Anonymous. About the cover. College Mathematics Journal, 54 (3):??, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2023.2209495.

Anonymous:2023:ACd

[Ano23d] Anonymous. About the cover. College Mathematics Journal, 54 (4):??, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2023.2259285.

Anonymous:2023:AC

[Ano23e] Anonymous. About the cover. College Mathematics Journal, 54 (5):??, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2023.2281232.

Anonymous:2023:C

[Ano23f] Anonymous. Correction. College Mathematics Journal, 54(4): 414, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2023.2224688.

Anonymous:2023:R

[Ano23g] Anonymous. Referees in 2023. College Mathematics Journal, 54(5):508–509, 2023. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2023.2264744.

Anonymous:2024:ACa

[Ano24a] Anonymous. About the cover. College Mathematics Journal, 55 (1):??, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2024.2302297.

Anonymous:2024:ACb

[Ano24b] Anonymous. About the cover. College Mathematics Journal, 55 (2):??, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2024.2324653.

Anonymous:2024:ACc

[Ano24c] Anonymous. About the cover. College Mathematics Journal, 55 (3):??, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2024.2345528.

Anonymous:2024:ACd

[Ano24d] Anonymous. About the cover. College Mathematics Journal, 55 (4):??, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2024.2374222.

Anonymous:2024:AC

[Ano24e] Anonymous. About the cover. College Mathematics Journal, 55 (5):??, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic).

Anonymous:2024:CP

[Ano24f] Anonymous. Crossword puzzle. College Mathematics Journal, 55(4):350–351, 2024. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2024.2354106.

Anonymous:2024:GPL

[Ano24g] Anonymous. George Pólya awards for 2023 and 2024. College Mathematics Journal, 55(5):454–459, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic).

Anonymous:2024:R

[Ano24h] Anonymous. Referees in 2024. College Mathematics Journal, 55 (5):462–463, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic).

Anonymous:2025:AC

[Ano25] Anonymous. About the cover:. College Mathematics Journal, 56 (1):??, 2025. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic).

Alsina:2014:PWD

[ANU14] Claudi Alsina, Roger Nelsen, and Hasan Unal. Proof without words: The difference of consecutive integer cubes is congruent to 1 modulo 6. College Mathematics Journal, 45(2):135, March 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.45.2.135.

Aboufadel:2005:BHI

[AOW05] Edward Aboufadel, Julia Olsen, and Jesse Windle. Breaking the Holiday Inn Priority Club CAPTCHA. College Mathematics Journal, 36(2):101–108, March 2005. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2005.
11922116.

Armbruster:1974:B

[AP74] Frank O. Armbruster and Jean J. Pedersen. Bubbles. Two-Year College Mathematics Journal, 5(3):34–38, September 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1974.11974334.

Abrams:1992:PB

[AP92] Aaron D. Abrams and Matteo J. Paris. The probability that (a, b ) = 1. College Mathematics Journal, 23(1):47, January 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1992.11973434.

Appelbaum:2001:MG

[App01] Elizabeth B. Appelbaum. Models for growth. College Mathematics Journal, 32(4):258–259, September 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2001.
11921886.

Albers:1986:IGB

[AR86] Donald J. Albers and Constance Reid. An interview with George B. Dantzig: The father of linear programming. College Mathematics Journal, 17(4):292–314, September 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1986.11972971.

Albers:1987:ILB

[AR87] Donald J. Albers and Constance Reid. An interview with Lipman Bers. College Mathematics Journal, 18(4):266–290, September 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1987.11973045.

Albers:1988:IME

[AR88] Donald J. Albers and Constance Reid. An interview with Mary Ellen Rudin. College Mathematics Journal, 19(2):114–137, March 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1988.11973101.

Abbott:1997:TWB

[AR97] Stephen D. Abbott and Matt Richey. Take a walk on the boardwalk. College Mathematics Journal, 28(3):162–171, May 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1997.11973857.

Aravind:2011:HSA

[Ara11] P. K. Aravind. How spherical are the Archimedean solids and their duals? College Mathematics Journal, 42(2):98–107, March 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.42.2.098.

Arganbright:1984:ESM

[Arg84] Deane E. Arganbright. The electronic spreadsheet and mathematical algorithms. College Mathematics Journal, 15(2):148–157, March 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972767.

Arganbright:1993:SR

[Arg93] Deane Arganbright. Software reviews. College Mathematics Journal, 24(1):89–94, January 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1993.11973514.

Armstrong:1974:CEI

[Arm74] Robert Owen Armstrong. Construction of an exercise involving minimum time. Two-Year College Mathematics Journal, 5(2): 12–14, Spring 1974. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1974.11974307.

Arnow:1994:LEB

[Arn94] Barry J. Arnow. On Laplace’s extension of the Buffon Needle Problem. College Mathematics Journal, 25(1):40–43, January 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1994.11973580.

Arpaia:1974:RSI

[Arp74] Pasquale J. Arpaia. Rings, subrings, identities and homomorphisms. Two-Year College Mathematics Journal, 5(1):25–28, Winter 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1974.11974288.

Albers:1982:CDKa

[AS82a] Donald J. Albers and Lynn Arthur Steen. A conversation with Don Knuth: Part 1. Two-Year College Mathematics Journal, 13 (1):2–18, January 1982. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1982.11972571. Reprinted in An-

nals of the History of Computing 4 (1982), 257–274. Reprinted in Mathematical People, ed. by Donald J. Albers and G. L. Alexanderson (Boston: Birkhäuser Boston, 1985), 182–203. Japanese translation by Nobuko Kishi in bit 16 (1985), 370–377, 506–512, 902–906, 1020–1025; reprinted in [Knu91, pp. 130–167].

Albers:1982:CDKb

[AS82b] Donald J. Albers and Lynn Arthur Steen. A conversation with Don Knuth: Part 2. Two-Year College Mathematics Journal, 13 (2):128–141, March 1982. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1982.11972592. Reprinted in An-

nals of the History of Computing 4 (1982), 257–274. Reprinted in Mathematical People, ed. by Donald J. Albers and G. L. Alexanderson (Boston: Birkhäuser Boston, 1985), 182–203. Japanese translation by Nobuko Kishi in bit 16 (1985), 370–377, 506–512, 902–906, 1020–1025; reprinted in [Knu91, pp. 130–167].

Avery:1988:ATN

[AS88] Chris W. Avery and Frank P. Soler. Applications of transformations to numerical integration. College Mathematics Journal, 19(2):166–173, March 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1988.11973106.

Anschuetz:1996:WFI

[AS96] Robert Anschuetz II and H. Sherwood. When is a function’s inverse equal to its reciprocal? College Mathematics Journal, 27(5):388–393, November 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1996.11973815.

Agard:2002:NLP

[AS02] David B. Agard and Michael W. Shackleford. A new look at the probabilities in bingo. College Mathematics Journal, 33(4):301– 305, September 2002. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2002.11921957.

Axtell:2010:GZD

[AS10] M. Axtell and J. Stickles. Graphs and zero-divisors. College Mathematics Journal, 41(5):396–399, November 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
4169/074683410X522008.

Ahstrom:2022:FBS

[AS22] Genevieve Ahstrom and Thomas Q. Sibley. Fantasy on a baseball seam. College Mathematics Journal, 53(4):265–272, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2022.2047559.

Asadulla:1984:DRM

[Asa84] Syed Asadulla. Digital roots of Mersenne primes and even perfect numbers. College Mathematics Journal, 15(1):53–54, January 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972747.

Ash:1997:NWC

[Ash97] J. Marshall Ash. Neither a worst convergent series nor a best divergent series exists. College Mathematics Journal, 28(4):296– 297, September 1997. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1997.11973879.

Ash:2020:DFL

[Ash20] J. Marshall Ash. Discontinuous functions as limits of compactly supported formulas. College Mathematics Journal, 51(5):337– 344, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2020.1820284.

Askey:1992:MH

[Ask92] Richard Askey. Media highlights. College Mathematics Journal, 23(5):441–448, November 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1992.11973498.

Abboud:2019:CC

[ASR19] Elias Abboud, Fathi Saleh, and Amal-Sharif Rasslan. Coloring a 1-by-n chessboard. College Mathematics Journal, 50(5):322– 330, 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2019.1662707.

Althoen:1994:CCF

[ASW94] S. C. Althoen, K. E. Schilling, and M. F. Wyneken. Cutting corners: a four-gon conclusion. College Mathematics Journal, 25(4):266–279, September 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1994.11973622.

Albers:1984:CHP

[AT84] Donald J. Albers and Michael J. Thibodeaux. A conversation with Henry Pollak. College Mathematics Journal, 15(3):194–217, June 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972773.

Adler:2000:MVP

[AT00] C. L. Adler and James Tanton. π is the minimum value for pi. College Mathematics Journal, 31(2):102–106, March 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2000.11974122.

Arledge:2008:NPR

[AT08] Jane Arledge and Sarah Tekansik. A new property of repeating decimals. College Mathematics Journal, 39(2):107–111, March 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2008.11922283.

Atkinson:2002:WDF

[Atk02] Leigh Atkinson. Where do functions come from? College Mathematics Journal, 33(2):107–112, March 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2002.
11921927.

Atkinson:2005:WPW

[Atk05] Leigh Atkinson. When the Pope was a mathematician. College Mathematics Journal, 36(5):354–362, November 2005. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2005.11922149.

Augros:2012:RPM

[Aug12] Michael Augros. Rediscovering Pascal’s mystic hexagon. College Mathematics Journal, 43(3):194–202, May 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
43.3.194.

Austin:1976:CC

[Aus76] Joe Dan Austin. Computer corner. Two-Year College Mathematics Journal, 7(4):39–42, December 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1976.
11974463.

Austin:1982:PSL

[Aus82] Homer W. Austin. Probability solution to a limit problem. Two-Year College Mathematics Journal, 13(4):272, September 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1982.11972628.

Alexanderson:1983:CGB

[AW83] G. L. Alexanderson and Carroll Wilde. A conversation with Garrett Birkhoff. Two-Year College Mathematics Journal, 14(2): 126–145, March 1983. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1983.11972669.

Althoen:1985:RRS

[AW85] S. C. Althoen and J. F. Weidner. Related rates and the speed of light. College Mathematics Journal, 16(3):186–189, June 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1985.11972878.

Althoen:1990:MFD

[AW90] S. C. Althoen and M. F. Wyneken. A mathematical field day. College Mathematics Journal, 21(5):379–383, November 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1990.11973336.

Ayoub:1987:BGR

[Ayo87] Ayoub B. Ayoub. Behold! The graphs of f and f^{−1}, are reflections about the line y = x. College Mathematics Journal, 18 (1):52, January 1987. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1987.11973010.

Ayoub:2003:ECS

[Ayo03] Ayoub B. Ayoub. The eccentricity of a conic section. College Mathematics Journal, 34(2):116–121, March 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2003.
11921994.

Bacharach:2009:BR

[Bac09] Deborah Bacharach. Book review. College Mathematics Journal, 40(5):384–386, November 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/074683409X475724.

Badger:1982:NP

[Bad82] Lee Badger. A nonlogarithmic proof that (1 + 1/n)^{n} → e. Two-Year College Mathematics Journal, 13(5):331–332, November 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1982.11972642.

Baildon:1985:APC

[Bai85] John D. Baildon. Arithmetic progressions and the consumer. College Mathematics Journal, 16(5):395–397, November 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1985.11972913.

Bailey:1986:HBY

[Bai86] Herbert R. Bailey. How to balance a yardstick on an Apple. College Mathematics Journal, 17(3):220–225, May 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1986.11972959.

Bailey:1987:BS

[Bai87] Herbert R. Bailey. Bat and superbat. College Mathematics Journal, 18(4):307–314, September 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1987.
11973048.

Bailar:1988:MSS

[Bai88] Barbara A. Bailar. The mathematicians’ statistics has a subsidiary role …. College Mathematics Journal, 19(1):14–15, January 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1988.11973077.

Bailey:1989:DEC

[Bai89] D. F. Bailey. Differentials and elementary calculus. College Mathematics Journal, 20(1):52–53, January 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1989.
11973204.

Bailey:1991:HBC

[Bai91] Herbert R. Bailey. ‘hidden’ boundaries in constrained maxmin problems. College Mathematics Journal, 22(3):227–229, May 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1991.11973386.

Bain:1998:CCC

[Bai98] Van Bain. Classroom Computer Capsule: an algorithm for drawing the n-cube. College Mathematics Journal, 29(4):320– 322, September 1998. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1998.11973966.

Bailey:2002:RR

[Bai02] Herb Bailey. On running in the rain. College Mathematics Journal, 33(2):88–92, March 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2002.11921924.

Bailey:2009:JPH

[Bai09] Herb Bailey. Jeeps penetrating a hostile desert. College Mathematics Journal, 40(3):182–187, May 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2009.
11922357.

Bailey:2013:SR

[Bai13] Herb Bailey. Slouching in the rain. College Mathematics Journal, 44(2):136–138, March 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.44.2.136.

Bajnok:2023:RAU

[Baj23] Béla Bajnok. Report on the 13th Annual USA Junior Mathematical Olympiad. College Mathematics Journal, 54(1):15– 21, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2023.2165373.

Baldino:1978:GPC

[Bal78] R. R. Baldino. A geometrical proof of Cramer’s Rule. Two-Year College Mathematics Journal, 9(2):106–107, March 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1978.11974559.

Baldwin:1982:NAa

[Bal82a] James Baldwin. News from amatyc. Two-Year College Mathematics Journal, 13(3):222, June 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1982.
11972619.

Baldwin:1982:NAb

[Bal82b] James Baldwin. News from amatyc. Two-Year College Mathematics Journal, 13(4):285, September 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1982.
11972633.

Balsam:2022:PWL

[Bal22] Jonathan Balsam. Proof without words: Lagrange’s trigonometric identity. College Mathematics Journal, 53(5):399, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2022.2118996.

Balsam:2023:PWL

[Bal23] Jonathan Balsam. Proof without words: Lagrange’s trigonometric identity (Part II). College Mathematics Journal, 54(3): 235, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2023.2206782.

Banks:1990:RDE

[Ban90] Bernard W. Banks. A rich differential equation for computer demonstrations. College Mathematics Journal, 21(1):45–50, January 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1990.11973284.

Banerjee:2004:RST

[Ban04] Sudipto Banerjee. Revisiting spherical trigonometry with orthogonal projectors. College Mathematics Journal, 35(5):375– 381, November 2004. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2004.11922099.

Banchoff:2006:CCC

[Ban06] Thomas Banchoff. Complementary coffee cups. College Mathematics Journal, 37(3):170–175, May 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2006.
11922182.

Bannon:2015:OQ

[Ban15] Thomas Bannon. The origin of quaternions. College Mathematics Journal, 46(1):43–50, January 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
46.1.43.

Barcellos:1979:CMG

[Bar79] Anthony Barcellos. A conversation with Martin Gardner. Two-Year College Mathematics Journal, 10(4):233–244, September 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1979.11972358.

Barcellos:1981:ISU

[Bar81a] Anthony Barcellos. An interview with Stan Ulam. Two-Year College Mathematics Journal, 12(3):182–189, June 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972521.

Barlow:1981:APD

[Bar81b] Penelope Barlow. An alternative proof to Dirac’s Theorem. Two-Year College Mathematics Journal, 12(1):57–58, January 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972496.

Barcellos:1984:FGM

[Bar84] Anthony Barcellos. The fractal geometry of Mandelbrot. College Mathematics Journal, 15(2):98–114, March 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1984.
11972758.

Barbeau:1991:MH

[Bar91] Ed Barbeau. Media highlights. College Mathematics Journal, 22(2):173–180, March 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1991.11973379.

Barry:1997:AL

[Bar97] P. D. Barry. On arc length. College Mathematics Journal, 28 (5):338–347, November 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1997.11973887.

Barker:2009:EST

[Bar09] Andrew T. Barker. Evolutionary stability in the Traveler’s Dilemma. College Mathematics Journal, 40(1):33–38, January 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2009.11922333.

Barbanel:2010:GAF

[Bar10] Julius Barbanel. A geometric approach to fair division. College Mathematics Journal, 41(4):268–280, September 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
4169/074683410X510263.

Barbeau:2011:BR

[Bar11] Edward J. Barbeau. Book review. College Mathematics Journal, 42(1):71–75, January 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.42.1.071.

Barrett:2022:UCL

[Bar22] Johnner Barrett. Unlawful calculations: A look into lie’s notebook. College Mathematics Journal, 53(2):104–115, 2022.

CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2021.2019550.

Bartroff:2023:RYP

[Bar23] Jay Bartroff. Rebalance your portfolio without selling. College Mathematics Journal, 54(4):290–298, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2023.
2228671.

Basu:2016:NRL

[Bas16] Kaushik Basu. A new and rather long proof of the Pythagorean Theorem by way of a proposition on isosceles triangles. College Mathematics Journal, 47(5):356–360, November 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
47.5.356.

Bautista:1993:SR

[Bau93] Maurino P. Bautista. Software reviews. College Mathematics Journal, 24(5):462–471, November 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1993.
11973570.

Bauldry:1998:SR

[Bau98] William C. Bauldry. Software reviews. College Mathematics Journal, 29(3):237–241, May 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1998.
11973949.

Bailey:1987:SMM

[BAW87] Herbert Bailey, Steven C. Althoen, and John Weidner. A surprising max-min result. College Mathematics Journal, 18(3):225– 229, May 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1987.11973038.

Brase:1976:BAB

[BB76] Corrinne J. Brase and Charles H. Brase. Basic algebra in a balanced lecture-program format. Two-Year College Mathematics Journal, 7(4):13–17, December 1976. CODEN ????

ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1976.
11974456.

Bevis:1982:CC

[BB82] Jean H. Bevis and Jan L. Boal. Computers and calculators. Two-Year College Mathematics Journal, 13(2):122–127, March 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1982.11972591.

Brown:1986:SR

[BB86] Judith R. Brown and Robert A. Beezer. Software reviews. College Mathematics Journal, 17(4):358–359, September 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1986.11972981.

Banchoff:1988:MH

[BB88] Thomas F. Banchoff and Thomas Banchoff. Media highlights. College Mathematics Journal, 19(2):194–202, March 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1988.11973115.

Bailey:1997:STI

[BB97] Herbert R. Bailey and Robert Bannister. A stronger triangle inequality. College Mathematics Journal, 28(3):182–186, May 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1997.11973859.

Beslin:2001:ALF

[BB01] Scott J. Beslin and Douglas J. Baney. Another look at factoring polynomials. College Mathematics Journal, 32(4):273– 275, September 2001. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2001.11921889.

Bannon:2008:MG

[BB08a] Thomas J. Bannon and Robert E. Bradley. Mind the gap. College Mathematics Journal, 39(2):95–101, March 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2008.
11922281.

Brown:2008:FFN

[BB08b] Ezra Brown and Jason C. Brunson. Fibonacci’s forgotten number. College Mathematics Journal, 39(2):112–120, March 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2008.11922284.

Barbanel:2011:TPP

[BB11] Julius B. Barbanel and Steven J. Brams. Two-person pie-cutting: The fairest cuts. College Mathematics Journal, 42(1):25–32, January 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.42.1.025.

Benjamin:2014:CMS

[BB14] Arthur T. Benjamin and Ethan J. Brown. Challenging magic squares for magicians. College Mathematics Journal, 45(2):92– 100, March 2014. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.45.2.092.

Boualem:2015:C

[BB15] Hassan Boualem and Robert Brouzet. To be (a circle) or not to be? College Mathematics Journal, 46(3):197–206, May 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.46.3.197.

Bagdasaryan:2016:ACT

[BB16] Armen Bagdasaryan and Josep Batle. Algebraic characterization of two independent events. College Mathematics Journal, 47(5):367–368, November 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.47.5.367.

Barrs:2003:RRT

[BBB03] Sharon Barrs, James Braselton, and Lorraine Braselton. A rational root theorem for imaginary roots. College Mathematics Journal, 34(5):380–382, November 2003. CODEN ????

ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2003.
11922035.

Baker:2013:SPC

[BBB^{+}13] Mark Baker, Jane Beltran, Jason Buell, Brian Conrey, Tom Davis, Brianna Donaldson, Jeanne Detorre-Ozeki, Leila Dibble, Tom Freeman, Robert Hammie, Julie Montgomery, Avery Pickford, and Justine Wong. Sets, planets, and comets. College Mathematics Journal, 44(4):258–264, September 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
44.4.258.

Bisgaier:2002:CHA

[BBHS02] Steven P. Bisgaier, Benjamin S. Bradley, Peter D. Harwood, and Paul M. Sommers. Chasing Hank Aaron’s home run record. College Mathematics Journal, 33(4):293–295, September 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2002.11921955.

Beezer:1987:SR

[BBM87] Robert A. Beezer, Gary Bitter, and Andrew Matchett. Software reviews. College Mathematics Journal, 18(4):337–339, September 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1987.11973056.

Beineke:2024:MHa

[BBS24a] Lowell Beineke, Jennifer Beineke, and Philip Straffin. Media highlights. College Mathematics Journal, 55(3):269–276, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2024.2316551.

Beineke:2024:MHb

[BBS24b] Lowell Beineke, Jennifer Beineke, and Philip Straffin. Media highlights. College Mathematics Journal, 55(5):446–453, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic).

Beineke:2025:MH

[BBS25] Lowell Beineke, Jennifer Beineke, and Philip Straffin. Media highlights. College Mathematics Journal, 56(1):76–83, 2025. CO-DEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic).

Brenner:1979:LLB

[BBW79] Joel Brenner, Ralph P. Boas, and Richard K. Williams. Letters and late-breaking news. Two-Year College Mathematics Journal, 10(5):372–374, November 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1979.11972393.

Bolton:1975:LIT

[BC75] Walter W. Bolton and Sterling G. Crim. A linear integral transform with a simple kernel. Two-Year College Mathematics Journal, 6(1):5–7, February 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1975.11974358.

Bushaw:1985:MH

[BC85] D. Bushaw and Philip Cheifetz. Media highlights. College Mathematics Journal, 16(3):233–247, June 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1985.
11972888.

Borrelli:1994:CLF

[BC94] Robert L. Borrelli and Courtney S. Coleman. Computers, lies, and the fishing season. College Mathematics Journal, 25(5):401– 412, November 1994. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1994.11973644.

Bonomo:2004:HDY

[BC04] John P. Bonomo and Carolyn K. Cuff. How do you stack up? College Mathematics Journal, 35(5):351–361, November 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2004.11922096.

Benjamin:2016:M

[BC16a] Arthur Benjamin and Rohan Chandra. Multiplying by 9. College Mathematics Journal, 47(4):281, September 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
47.4.281.

Berele:2016:APB

[BC16b] Allan Berele and Stefan Catoiu. Area and perimeter bisecting lines of a triangle. College Mathematics Journal, 47(1):19–28, January 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.47.1.19.

Berele:2017:CNA

[BC17] Allan Berele and Stefan Catoiu. The centroid as a nontrivial area bisecting center of a triangle. College Mathematics Journal, 49(1):27–34, January 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2017.1395659.

Benko:2018:NAF

[BC18a] David Benko and Dan Coroian. A new angle on the Fermat– Torricelli Point. College Mathematics Journal, 49(3):195–199, May 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2018.1440865.

Brown:2018:FFA

[BC18b] Ezra Brown and Matthew Crawford. Five families around a well: a new look at an old problem. College Mathematics Journal, 49(3):162–168, May 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2018.1447203.

Bajnok:2022:RAU

[BC22] Béla Bajnok and Evan Chen. Report on the 12th Annual USA Junior Mathematical Olympiad. College Mathematics Journal, 53(1):13–20, 2022. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2021.1991168.

Borodin:2024:SMP

[BCD^{+}24] Matvey Borodin, Eric Chen, Aidan Duncan, Tanya Khovanova, Boyan Litchev, Jiahe Liu, Veronika Moroz, Matthew Qian, Rohith Raghavan, Garima Rastogi, and Michael Voigt. The stable marriage problem and Sudoku. College Mathematics Journal, 55(4):303–316, 2024. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2023.2261183.

Brown:1985:RAH

[BCEW85] Fon Brown, L. O. Cannon, Joe Elich, and David G. Wright. On rearrangements of the alternating harmonic series. College Mathematics Journal, 16(2):135–138, March 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1985.
11972869.

Beres:2021:LNK

[BCK^{+}21] Steven Beres, Vesta Coufal, Kate Kearney, Ryan Lattanzi, and Hayley Olson. Linking numbers of Klein links. College Mathematics Journal, 52(2):106–114, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2021.1877524.

Birky:2011:OPN

[BCR^{+}11] Geoffrey Birky, Connie Campbell, Manya Raman, James Sandefur, and Kay Somers. One problem, nine student-produced proofs. College Mathematics Journal, 42(5):355–360, November 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.42.5.355.

Bajnok:2025:RAU

[BCT25] Béla Bajnok, Evan Chen, and Enrique Treviño. Report on the 15th Annual USA Junior Mathematical Olympiad. College Mathematics Journal, 56(1):84–91, 2025. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic).

Babcock:1978:NAL

[BD78] Bruce S. Babcock and John W. Dawson, Jr. A neglected approach to the logarithm. Two-Year College Mathematics Journal, 9(3):136–140, June 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1978.11974563.

Bivens:1996:SGWa

[BD96a] Irl Bivens and Stephen L. Davis. Some graphs whose vertices pair off by degree: Part I. College Mathematics Journal, 27(2):

127–135, March 1996. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1996.11973765.

Bivens:1996:SGWb

[BD96b] Irl Bivens and Stephen L. Davis. Some graphs whose vertices pair off by degree: Part II. College Mathematics Journal, 27(3):213– 219, May 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1996.11973782.

Brigham:1985:RPV

[BDCH85] Robert C. Brigham, Ronald D. Dutton, Phyllis Z. Chinn, and Frank Harary. Realization of parity visits in walking a graph. College Mathematics Journal, 16(4):280–282, September 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1985.11972897.

Bruckman:2013:SRT

[BDDY13] Paul Bruckman, Joseph B. Dence, Thomas P. Dence, and Justin Young. Series of reciprocal triangular numbers. College Mathematics Journal, 44(3):177–184, May 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
44.3.177.

Britton:2018:PQH

[BDLS18] Sam Britton, Joseph DiMuro, Daniel Lundstrom, and Joshua Sansonetti. Probabilities of Qwirkle hand values. College Mathematics Journal, 49(4):270–276, September 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2018.1490592.

Boyce:1995:COC

[BE95] William E. Boyce and Joseph G. Ecker. The computer-oriented calculus course at Rensselaer Polytechnic Institute. College Mathematics Journal, 26(1):45–50, January 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1995.
11973668.

Brenton:2003:SSC

[BE03] Lawrence Brenton and Thomas G. Edwards. Sets of sets: a cognitive obstacle. College Mathematics Journal, 34(1):31–38, January 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2003.11921982.

Beauregard:2006:NAC

[Bea06] Raymond A. Beauregard. No arithmetic cyclic quadrilaterals. College Mathematics Journal, 37(2):110–113, March 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2006.11922173.

Beauregard:2009:DQT

[Bea09] Raymond A. Beauregard. Diametric quadrilaterals with two equal sides. College Mathematics Journal, 40(1):17–21, January 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2009.11922331.

Beauregard:2011:TTS

[Bea11] R. A. Beauregard. Teaching tip: a slippery slope. College Mathematics Journal, 42(3):206, May 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2011.
11922502.

Beasley:2012:MKT

[Bea12] John D. Beasley. Magic knight’s tours. College Mathematics Journal, 43(1):72–75, January 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
43.1.072.

Beam:2017:PMN

[Bea17] John Beam. A powerful method of non-proof. College Mathematics Journal, 48(1):52–54, January 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
48.1.52.

Behr:1973:AAA

[Beh73] Anthony N. Behr. Achievement, aptitude and attitude in mathematics. Two-Year College Mathematics Journal, 4(2):72–74, Spring 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1973.11974255.

Beineke:1989:GFS

[Bei89] Lowell W. Beineke. Are graphs finally surfacing? College Mathematics Journal, 20(3):206–225, May 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1989.
11973232.

Beineke:2020:CMJ

[Bei20] Lowell Beineke. The College Mathematics Journal during the decade of the 2000s. College Mathematics Journal, 51(4):302– 303, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2020.1787756.

Birnbaum:1975:LE

[BEL75] Saul Birnbaum, David W. Erbach, and Peter A. Lindstrom. Letters to the Editor. Two-Year College Mathematics Journal, 6(4): 2–3, December 1975. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1975.11974397.

Belfi:1984:CEC

[Bel84] Victor A. Belfi. Convexity in elementary calculus: Some geometric equivalences. College Mathematics Journal, 15(1):37–41, January 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972742.

Bell:2015:ACS

[Bel15] Denis Bell. Area of a circle via the second fundamental theorem of calculus. College Mathematics Journal, 46(4):299, September 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.46.4.299.

Bender:1975:HF

[Ben75] David Bender. Hyperbolic functions. Two-Year College Mathematics Journal, 6(3):42–45, September 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1975.
11974394.

Bennett:1976:ANC

[Ben76a] Larry F. Bennett. Another note on “closing the loopholes”. Two-Year College Mathematics Journal, 7(3):56–58, September 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1976.11974451.

Bennett:1976:ANE

[Ben76b] Larry F. Bennett. Another note on epsilons and deltas. Two-Year College Mathematics Journal, 7(3):18, September 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1976.11974440.

Burrell:1978:ECM

[BEN78] Benjamin Burrell, Jessie Ann Engle, and Henry C. Nixt. Experiment and conjecture in mathematics: a discovery course for College freshmen and sophomores. Two-Year College Mathematics Journal, 9(4):210–215, September 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1978.
11974580.

Benson:1994:EGA

[Ben94] Steve Benson. Euclid’s (Gaussian) algorithm: a lattice approach. College Mathematics Journal, 25(2):118–124, March 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1994.11973593.

Benyi:2005:FSH

[Bén05] Árpád Bényi. Finding the sums of harmonic series of even order. College Mathematics Journal, 36(1):44–48, January 2005. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2005.11922109.

Benjamin:2012:SCC

[Ben12a] Arthur T. Benjamin. Squaring, cubing, and cube rooting. College Mathematics Journal, 43(1):58–64, January 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
43.1.058.

Benko:2012:BPT

[Ben12b] David Benko. The Basel Problem as a telescoping series. College Mathematics Journal, 43(3):244–250, May 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
43.3.244.

Benjamin:2017:FNC

[Ben17] Arthur T. Benjamin. Factoring numbers with Conway’s 150 method. College Mathematics Journal, 49(2):122–125, March 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2017.1411141.

Beran:1975:NTI

[Ber75] David Beran. Note on teaching the implication. Two-Year College Mathematics Journal, 6(3):18–19, September 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1975.11974387.

Berman:1980:AAI

[Ber80] Martin Berman. Altitudes ad infinitum. Two-Year College Mathematics Journal, 11(5):300–304, November 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1980.
11972466.

Berry:1987:MH

[Ber87] John W. Berry. Media highlights. College Mathematics Journal, 18(3):253–261, May 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1987.11973043.

Berman:1994:LDO

[Ber94] David M. Berman. Lottery drawings often have consecutive numbers. College Mathematics Journal, 25(1):45–47, January 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1994.11973582.

Berndt:1998:MH

[Ber98] Bruce Berndt. Media highlights. College Mathematics Journal, 29(5):441–448, November 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1998.11973986.

Berresford:2002:RCT

[Ber02] Geoffrey C. Berresford. Runs in coin tossing: Randomness revealed. College Mathematics Journal, 33(5):391–394, November 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2002.11921969.

Berman:2006:FB

[Ber06] Leah Wrenn Berman. Folding beauties. College Mathematics Journal, 37(3):176–186, May 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2006.
11922183.

Berendonk:2014:SCC

[Ber14a] Stephan Berendonk. Stretched circles are conic sections, a geometric proof. College Mathematics Journal, 45(4):316–317, September 2014. CODEN ???? ISSN 0746-8342 (print), 1931-

1346 (electronic). URL http://www.tandfonline.com/doi/
abs/10.4169/college.math.j.45.4.316.

Berger:2014:GJW

[Ber14b] Ruth I. Berger. Green jello world and Escher’s world. College Mathematics Journal, 45(1):50–53, January 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
45.1.050.

Berendonk:2018:PWS

[Ber18] Stephan Berendonk. Proof without words: Sums of squares in a thin rectangle. College Mathematics Journal, 49(3):180, May 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2018.1424424.

Berendonk:2020:GSE

[Ber20] Stephan Berendonk. Geometric series in an equilateral triangle — three proofs without words. College Mathematics Journal, 51(5):385, 2020. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2020.1830649.

Besenyei:2014:TWT

[Bes14] Ádám Besenyei. Traveling waves and Taylor series: Do they have something in common? College Mathematics Journal, 45 (1):29–32, January 2014. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.45.1.029.

Betounes:2012:BTO

[Bet12] David Betounes. Better than optimal by taking a limit? College Mathematics Journal, 43(5):379–386, November 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
43.5.379.

Birnbaum:1977:ERD

[BF77] David A. Birnbaum and Northrup Fowler III. An elementary result on derivatives. Two-Year College Mathematics Journal, 8 (1):10–11, January 1977. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1977.11974471.

Benjamin:2000:RSC

[BF00] Arthur T. Benjamin and Matthew T. Fluet. A rational solution to cootie. College Mathematics Journal, 31(2):124–125, March 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2000.11974127.

Bravo:2013:MMR

[BF13] Daniel Bravo and Joseph Fera. Mathematical minute: Rotating a function graph. College Mathematics Journal, 44(2):124–125, March 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.44.2.124.

Bonomo:2023:SF

[BF23] John P. Bonomo and Montana Ferita. A small fib. College Mathematics Journal, 54(4):274–289, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2023.
2225397.

Burgette:2004:DNC

[BG04] Lane F. Burgette and Russell A. Gordon. On determining the non-circularity of a plane curve. College Mathematics Journal, 35(2):74–83, March 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2004.11922055.

Bedient:2015:GMG

[BG15] Richard Bedient and Courtney Gibbons. Grandma makes granola. College Mathematics Journal, 46(1):58–60, January 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.46.1.58.

Burk:1986:URS

[BGR86] Frank Burk, Sudhir K. Goel, and Dennis M. Rodriguez. Using Riemann sums in evaluating a familiar limit. College Mathematics Journal, 17(2):170–171, March 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1986.
11972952.

Bange:1993:DPA

[BH93] David Bange and Linda Host. Does a parabola have an asymptote? College Mathematics Journal, 24(4):331–342, September 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1993.11973550.

Bell:2022:IMC

[BH22] Lindsey Bell and Paul Hill. An investigation of metropolitan crime distribution. College Mathematics Journal, 53(5):364– 371, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2022.2125263.

Bailey:2004:RMH

[BHB04] Mark Bailey, Mike Hilgert, and Herb Bailey. Recirculation models, homogenized milk, and biotech applications. College Mathematics Journal, 35(4):283–288, September 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2004.11922086.

Becsi:2021:AST

[BHS^{+}21] Brian Becsi, Solomon Huang, Verenalei Schoenfeld, Bogdan D. Suceava and Ashley Thune-Aguayo. Applications of squeeze theorem to limiting processes involving Riemann integration. College Mathematics Journal, 52(3):224–226, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2021.1909980.

Bolt:2010:DDN

[BI10] Michael Bolt and Daniel C. Isaksen. Dogs don’t need calculus. College Mathematics Journal, 41(1):10–16, January 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/074683410X475074.

Bialostocki:1998:AEG

[Bia98] Arie Bialostocki. An application of elementary group theory to central solitaire. College Mathematics Journal, 29(3):208–212, May 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1998.11973941.

Bickle:2017:MG

[Bic17] Allan Bickle. MegaMenger graphs. College Mathematics Journal, 49(1):20–26, January 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2017.1394688.

Bidwell:1986:BGA

[Bid86] James K. Bidwell. A Babylonian geometrical algebra. College Mathematics Journal, 17(1):22–31, January 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1986.
11972926.

Bilodeau:1993:ER

[Bil93] G. E. Bilodeau. An exponential rule. College Mathematics Journal, 24(4):350–351, September 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1993.
11973553.

Bilodeau:2018:PWR

[Bil18] Gerald E. Bilodeau. Proof without words: Riemann meets Euler– Mascheroni. College Mathematics Journal, 49(5):341, November 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2018.1526575.

Birgen:2010:VEU

[Bir10] Mariah Birgen. Visualizing elections using Saari triangles. College Mathematics Journal, 41(4):325–328, September 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
4169/074683410X510326.

Birgen:2015:URP

[Bir15] Brian J. Birgen. The uniqueness of rock–paper–scissors–lizard– Spock. College Mathematics Journal, 46(4):270–273, September 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.46.4.270.

Bissinger:1980:SIV

[Bis80] B. H. Bissinger. Stochastic independent versus intuitive independence. Two-Year College Mathematics Journal, 11(2):122–123, March 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1980.11972421.

Bitter:1970:CCE

[Bit70] Gary G. Bitter. Calculus and the computer: an evaluation by participants. Two-Year College Mathematics Journal, 1(2):41–49, Autumn 1970. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1970.11973187.

Bivens:1986:WTL

[Biv86] Irl C. Bivens. What a tangent line is when it isn’t a limit. College Mathematics Journal, 17(2):133–143, March 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1986.11972944.

Bivens:1997:MDC

[Biv97] Irl C. Bivens. Multiple derivatives of compositions: Investigating some special cases. College Mathematics Journal, 28(4):299– 300, September 1997. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1997.11973881.

Bixler:1995:MH

[Bix95] Harry Bixler. Media highlights. College Mathematics Journal, 26(4):335–342, September 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1995.11973726.

Brown:2024:PBC

[BJ24] Tova Brown and Brody Johnson. Pull-back cars: Vehicles for the instruction of differential equations. College Mathematics Journal, 55(3):192–204, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2024.2302300.

Burke:2017:FLP

[BJK17] Brittany A. Burke, Zach Jackson, and Steven J. Kifowit. The falling ladder paradox revisited. College Mathematics Journal, 49(1):36–40, January 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2017.1388690.

Bayne:1997:EF

[BJKL97] Richard E. Bayne, James E. Joseph, Myung H. Kwack, and Thomas H. Lawson. Exploiting a factorization of x^{n} − y^{n}. College Mathematics Journal, 28(3):206–209, May 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1997.11973865.

Bivens:1985:RDU

[BK85] Irl C. Bivens and L. R. King. Relating differentiability and uniform continuity. College Mathematics Journal, 16(4):283, September 1985. CODEN ???? ISSN 0746-8342 (print), 1931-

1346 (electronic). URL http://www.tandfonline.com/doi/
abs/10.1080/07468342.1985.11972899.

Bailey:2011:WSF

[BK11] Herb Bailey and Dan Kalman. Walking with a slower friend. College Mathematics Journal, 42(5):384–387, November 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.42.5.384.

Braswell:2014:CMD

[BK14] Leigh Marie Braswell and Tanya Khovanova. Cookie Monster devours Naccis. College Mathematics Journal, 45(2):129–135, March 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.45.2.129.

Basu:2019:FPL

[BKSU19] R. Basu, V. Kannan, K. Sannyasi, and N. Unnikrishnan. Functions preserving limit superior. College Mathematics Journal, 50(1):58–60, January 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2019.1534489.

Bloomfield:1975:CL

[BL75] Morton Bloomfield and Frank Lasak. Closing the loopholes. Two-Year College Mathematics Journal, 6(2):42–44, May 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1975.11974379.

Brieske:1978:MGF

[BL78] Tom Brieske and Johnny Lott. The motion geometry of a finite plane. Two-Year College Mathematics Journal, 9(5):259– 266, November 1978. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1978.11974589.

Blackburn-Lynch:2022:WEO

[BL22a] James Blackburn-Lynch. Are we ever our best possible selves? An application of Bézout’s identity to find coincident peaks of multiple sine curves. College Mathematics Journal, 53(3):183– 189, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2022.2040264.

Brawner:2022:ISA

[BL22b] James N. Brawner and Nadou Lawson. Integer solutions to angle optimization problems. College Mathematics Journal, 53(3): 197–208, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2022.2045158.

Blaie:1991:TVC

[Bla91] Larry Blaie. Theory vs. computation in some very simple dynamical systems. College Mathematics Journal, 22(1):42–44, January 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1991.11973357.

Blanchard:1994:TDE

[Bla94] Paul Blanchard. Teaching differential equations with a dynamical systems viewpoint. College Mathematics Journal, 25(5):385– 393, November 1994. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1994.11973642.

Blank:2001:BR

[Bla01] Brian Blank. Book review. College Mathematics Journal, 32(2): 155–160, March 2001. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2001.11921868.

Blake:2016:F

[Bla16a] Sarah Blake. Form. College Mathematics Journal, 47(5):333, November 2016. CODEN ???? ISSN 0746-8342 (print), 1931-

1346 (electronic). URL http://www.tandfonline.com/doi/
abs/10.4169/college.math.j.47.5.333.

Blaasjo:2016:HFL

[Blå16b] Viktor Blåsjö. How to find the logarithm of any number using nothing but a piece of string. College Mathematics Journal, 47(2):95–100, March 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.47.2.95.

Benyi:2009:PTR

[BlC09] Árpád Bényi and loan Casu. Pompeiu’s Theorem revisited. College Mathematics Journal, 40(4):252–258, September 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2009.11922372.

Bleck:1975:CC

[Ble75] Daniel T. Bleck. Computer corner. Two-Year College Mathematics Journal, 6(3):38–40, September 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1975.
11974392.

Bloom:1998:SSC

[Blo98] David M. Bloom. Singles in a sequence of coin tosses. College Mathematics Journal, 29(2):120–127, March 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1998.
11973924.

Bloch:2001:TGG

[Blo01] Bill Goldbloom Bloch. Tension in generalized geometric sequences. College Mathematics Journal, 32(1):44–47, January 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2001.11921851.

Beresin:1989:CCP

[BLW89] May Beresin, Eugene Levine, and John Winn. A chessboard coloring problem. College Mathematics Journal, 20(2):106–114, March 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1989.11973216.

Burton:1977:CUC

[BM77] Charles R. Burton and Wanda A. Marosz. The construction and uses of CATIA, a computerized mathematics testbank. Two-Year College Mathematics Journal, 8(4):212–216, September 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1977.11974512.

Brechner:1988:ANH

[BM88] Beverly L. Brechner and John C. Mayer. Antoine’s necklace or how to keep a necklace from falling apart. College Mathematics Journal, 19(4):306–320, September 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1988.11973131.

Brualdi:1993:SAE

[BM93] Richard A. Brualdi and Jennifer J. Q. Massey. Some applications of elementary linear algebra in combinatorics. College Mathematics Journal, 24(1):10–19, January 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1993.
11973501.

Bauman:1995:AQS

[BM95] Steven F. Bauman and William O. Martin. Assessing the quantitative skills of College juniors. College Mathematics Journal, 26 (3):214–220, May 1995. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1995.11973699.

Barnier:1997:UFE

[BM97] William Barnier and Douglas Martin. Unifying a family of extrema problems. College Mathematics Journal, 28(5):388–391, November 1997. CODEN ???? ISSN 0746-8342 (print), 1931-

1346 (electronic). URL http://www.tandfonline.com/doi/
abs/10.1080/07468342.1997.11973898.

Berkove:1998:LAC

[BM98a] Ethan Berkove and Rich Marchand. The long arm of calculus. College Mathematics Journal, 29(5):376–386, November 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1998.11973973.

Boman:1998:DPM

[BM98b] Eugene C. Boman and Margaret A. Misconish. A diagonal perspective on matrices. College Mathematics Journal, 29(1):37–38, January 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1998.11973911.

Benko:2013:BPR

[BM13] David Benko and John Molokach. The Basel Problem as a rearrangement of series. College Mathematics Journal, 44(3):171– 176, May 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.44.3.171.

Benashaski:2000:IHP

[BMO^{+}00] Jessica Benashaski, John Meier, Kevin O’Brien, Paige Reinheimer, and Margaret Skarbek. Introducing hyperbolicity via piecewise Euclidean complexes. College Mathematics Journal, 31 (3):213–217, May 2000. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2000.11974147.

Bermudez:2013:SDN

[BMRS13] Frank Bermudez, Anthony Medina, Amber Rosin, and Eren Scott. Are stupid dice necessary? College Mathematics Journal, 44(4):315–322, September 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.44.4.315.

Boelkins:2006:CBT

[BMV06] Matthew Boelkins, Jennifer Miller, and Benjamin Vugteveen. From Chebyshev to Bernstein: a tour of polynomials small and large. College Mathematics Journal, 37(3):194–204, May 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2006.11922185.

Bolt:2014:FWR

[BMV14] Michael Bolt, Anthony Meyer, and Nicholas Visser. The fastest way not to run a four-minute mile. College Mathematics Journal, 45(2):101–107, March 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.45.2.101.

Bosse:2003:WEE

[BN03] Michael J. Bossé and N. R. Nandakumar. When equalities are not equal: Missing mathematical precision in teaching, texts, and technology. College Mathematics Journal, 34(5):383–389, November 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2003.11922036.

Barrows:2023:TCE

[BNUW23] Sara Barrows, Emily Noye, Sarah Uttormark, and Matthew Wright. Three’s a crowd: an exploration of subprime Tribonacci sequences. College Mathematics Journal, 54(5):464– 475, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2023.2263109.

Boneh:1995:CQM

[BO95] Shahar Boneh and David C. Ogden. A combinatorial queueing model. College Mathematics Journal, 26(5):346–357, November 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1995.11973728.

Boas:1971:MUP

[Boa71a] Mary L. Boas. Mathematics for the undergraduate physics major. Two-Year College Mathematics Journal, 2(1):49–52, Spring 1971. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1971.11973998.

Boas:1971:CES

[Boa71b] R. P. Boas, Jr. Calculus as an experimental science. Two-Year College Mathematics Journal, 2(1):36–39, Spring 1971. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1971.11973995.

Boas:1972:AC

[Boa72] R. P. Boas, Jr. Anomalous cancellation. Two-Year College Mathematics Journal, 3(2):21–24, Autumn 1972. CODEN ????

ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1972.
11974211.

Boas:1979:TS

[Boa79a] R. P. Boas. Travelers’ surprises. Two-Year College Mathematics Journal, 10(2):82–88, March 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1979.11972329.

Boas:1979:JLT

[Boa79b] Ralph P. Boas. Jazz, literature, and the teaching of mathematics. Two-Year College Mathematics Journal, 10(4):264–265, September 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1979.11972362.

Boas:1980:SEP

[Boa80] Ralph Boas. Snowfalls and elephants, pop bottles and π. Two-Year College Mathematics Journal, 11(2):82–89, March 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1980.11972410.

Boas:1981:NLG

[Boa81a] R. P. Boas. Nevertheless, let’s get on with the job. Two-Year College Mathematics Journal, 12(2):141–142, March 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972511.

Boas:1981:WNM

[Boa81b] Ralph P. Boas. Who needs those mean-value theorems, anyway? Two-Year College Mathematics Journal, 12(3):178–181, June 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972520.

Boas:1982:PSP

[Boa82] Ralph Boas. Power series for practical purposes. Two-Year College Mathematics Journal, 13(3):191–195, June 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1982.11972606.

Boas:1983:VB

[Boa83] Ralph P. Boas. The versed of Boas. Two-Year College Mathematics Journal, 14(4):342–344, September 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1983.
11972710.

Boas:1985:IF

[Boa85] Ralph P. Boas. Inverse functions. College Mathematics Journal, 16(1):42–47, January 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1985.11972848.

Boas:1986:DHW

[Boa86] R. P. Boas. Does “holds water” hold water? College Mathematics Journal, 17(4):341, September 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1986.
11972978.

Boas:2019:OTB

[Boa19] Harold P. Boas. The oldest trig in the book. College Mathematics Journal, 50(1):9–20, January 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2019.
1535158.

Bobo:1995:SRH

[Bob95] E. Ray Bobo. A sequence related to the harmonic series. College Mathematics Journal, 26(4):308–310, September 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1995.11973719.

Bodenrader:1971:FF

[Bod71] J. C. Bodenrader. Factoring functions. Two-Year College Mathematics Journal, 2(1):23–26, Spring 1971. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1971.
11973992.

Bodenrader:1974:MC

[Bod74] Joseph C. Bodenrader. Mini-calculus. Two-Year College Mathematics Journal, 5(2):74, Spring 1974. CODEN ????

ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1974.
11974321.

Bollinger:1983:HMB

[Bol83] Richard C. Bollinger. How to make a bank shot. Two-Year College Mathematics Journal, 14(2):169–170, March 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1983.11972676.

Bompart:1982:AMS

[Bom82] Bill Bompart. An alternate method for solving radical equations. Two-Year College Mathematics Journal, 13(3):198–199, June 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1982.11972610.

Boman:2009:FPD

[Bom09] Eugene C. Boman. False position, double false position and Cramer’s Rule. College Mathematics Journal, 40(4):279–283, September 2009. CODEN ???? ISSN 0746-8342 (print), 1931-

1346 (electronic). URL http://www.tandfonline.com/doi/
abs/10.1080/07468342.2009.11922376.

Bonomo:2008:YCT

[Bon08] John P. Bonomo. You can Teach an Old Magician New Tricks. College Mathematics Journal, 39(5):346–356, November 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2008.11922315.

Bonomo:2014:ANC

[Bon14] John P. Bonomo. Not all numbers can be created equally. College Mathematics Journal, 45(1):3–10, January 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
45.1.003.

Bonomo:2016:WPH

[Bon16] John P. Bonomo. Winning a pool is harder than you thought. College Mathematics Journal, 47(5):347–355, November 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.47.5.347.

Bonomo:2018:SHI

[Bon18] John P. Bonomo. The solution to a Hanoi-ing little problem. College Mathematics Journal, 49(4):288–291, September 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2018.1502983.

Bonomo:2021:BT

[Bon21] John P. Bonomo. Back to the tower. College Mathematics Journal, 52(4):265–273, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2021.1941540.

Book:1979:WFS

[Boo79] Stephen A. Book. Why n − 1 in the formula for the sample standard deviation? Two-Year College Mathematics Journal, 10(5):330–333, November 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1979.11972378.

Boone:1996:MP

[Boo96] James R. Boone. The moise plane. College Mathematics Journal, 27(3):182–185, May 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1996.11973773.

Borges:2011:DVR

[Bor11] Carlos F. Borges. Discretization vs. rounding error in Euler’s Method. College Mathematics Journal, 42(5):396–399, November 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.42.5.396.

Boston:1995:TMP

[Bos95] Nigel Boston. A Taylor-made plug for Wiles’ proof. College Mathematics Journal, 26(2):100–105, March 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1995.
11973677.

Bosch:2000:OCC

[Bos00] Robert A. Bosch. Optimal card-collecting strategies for magic: The gathering. College Mathematics Journal, 31(1):15–21, January 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2000.11974103.

Botts:1978:TYC

[Bot78] Truman Botts. Two-year College news. Two-Year College Mathematics Journal, 9(1):37–39, January 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1978.
11974544.

Bourgeois:1981:BR

[Bou81] Helen D. Bourgeois. Book reviews. Two-Year College Mathematics Journal, 12(3):222–223, June 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1981.
11972532.

Boute:1999:SGS

[Bou99] Raymond Boute. Simple geometric solutions to De l’Hospital’s pulley problem. College Mathematics Journal, 30(4):311–314, September 1999. CODEN ???? ISSN 0746-8342 (print), 1931-

1346 (electronic). URL http://www.tandfonline.com/doi/
abs/10.1080/07468342.1999.11974077.

Boulet:2023:SLS

[Bou23] Romain Boulet. Solving linear systems with column reduction. College Mathematics Journal, 54(2):104–112, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2023.2186082.

Boucher:2024:UTA

[Bou24] Chris Boucher. Using trapezoidal approximations to estimate averages. College Mathematics Journal, 55(5):381–386, 2024. CO-DEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic).

Boyer:1970:HC

[Boy70] Carl B. Boyer. The history of THE CALCULUS. Two-Year College Mathematics Journal, 1(1):60–86, Spring 1970.

CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1970.11973171.

Boyd:1981:CDE

[Boy81] Eddie Boyd, Jr. Constructing “different” examples for beginning abstract algebra students. Two-Year College Mathematics Journal, 12(5):333–334, November 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1981.
11972565.

Boyce:1994:NDE

[Boy94] William E. Boyce. New directions in elementary differential equations. College Mathematics Journal, 25(5):364–371, November 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1994.11973640.

Boyadzhiev:1999:SCF

[Boy99] Khristo N. Boyadzhiev. Spirals and conchospirals in the flight of insects. College Mathematics Journal, 30(1):23–31, January 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1999.11974025.

Boyadzhiev:2007:ESS

[Boy07] Khristo N. Boyadzhiev. Equiangular surfaces, self-similar surfaces, and the geometry of seashells. College Mathematics Journal, 38(4):265–271, September 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2007.
11922247.

Boyles:2010:FRP

[Boy10] Dave Boyles. Finding rational parametric curves of relative degree one or two. College Mathematics Journal, 41(5):371– 382, November 2010. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/074683410X521973.

Boyles:2015:RIE

[Boy15] Dave Boyles. Rational and implicit equations for some polar curves. College Mathematics Journal, 46(3):189–196, May 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.46.3.189.

Brown:2012:WRM

[BP12] Tom Brown and Brian Pasko. Winning a racquetball match. College Mathematics Journal, 43(5):395–400, November 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.43.5.395.

Bacinski:2020:HWT

[BP20] Steve Bacinski and Timothy Pennings. How to win at Tenzi. College Mathematics Journal, 51(4):242–253, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2020.1774296.

Bacinski:2015:ELM

[BPP15] Steve J. Bacinski, Mark J. Panaggio, and Timothy J. Pennings. Elvis lives: Mathematical surprises inspired by Elvis, the Welsh Corgi. College Mathematics Journal, 46(2):82–91, March 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.46.2.82.

Bell:2015:FPB

[BPR15] Kathleen Bell, Shania Polson, and Tom Richmond. The fastest path between two points, with a symmetric obstacle. College Mathematics Journal, 46(2):92–97, March 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
46.2.92.

Bekes:2012:MTP

[BPS12] Robert Bekes, Jean Pedersen, and Bin Shao. Mad tea party cyclic partitions. College Mathematics Journal, 43(1):25–36, January 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.43.1.025.

Benjamin:1999:RFL

[BQ99] Arthur T. Benjamin and Jennifer J. Quinn. Recounting Fibonacci and Lucas identities. College Mathematics Journal, 30 (5):359–366, November 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1999.11974086.

Benjamin:2008:AAA

[BQ08] Arthur T. Benjamin and Jennifer J. Quinn. An alternate approach to alternating sums: a method to DIE for. College Mathematics Journal, 39(3):191–202, May 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2008.
11922293.

Brand:2013:MTV

[BQ13] Neal Brand and John A. Quintanilla. Modeling terminal velocity. College Mathematics Journal, 44(1):57–61, January 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.44.1.057.

Boyd:1983:ACC

[BR83a] J. N. Boyd and P. N. Raychowdhury. An application of convex coordinates. Two-Year College Mathematics Journal, 14 (4):348–349, September 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1983.11972713.

Boyd:1983:CAL

[BR83b] J. N. Boyd and P. N. Raychowdhury. A computational approach to logical statements. Two-Year College Mathematics Journal, 14(4):326–328, September 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1983.11972708.

Boyd:1987:CCP

[BR87] J. N. Boyd and P. N. Raychowdhury. Convex coordinates, probabilities, and the superposition of states. College Mathematics Journal, 18(3):186–194, May 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1987.
11973032.

Boyd:1989:DDP

[BR89] J. N. Boyd and P. N. Raychowdhury. Discrete Dirichlet problems, convex coordinates, and a random walk on a triangle. College Mathematics Journal, 20(5):385–392, November 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1989.11973261.

Boyd:1994:SR

[BR94] James N. Boyd and P. N. Raychowdhury. Software reviews. College Mathematics Journal, 25(2):155–156, March 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1994.11973601.

Bataille:1999:CWF

[BR99] Michel Bataille and Guanshen Ren. Construction without words: Focus and directrix. College Mathematics Journal, 30(3):212, May 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1999.11974059.

Bailey:2000:VTI

[BR00] Herb Bailey and Yanir A. Rubinstein. A variety of triangle inequalities. College Mathematics Journal, 31(5):350–355, November 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2000.11974173.

Betounes:2017:DTS

[BR17] David Betounes and Mylan Redfern. The demise of trig substitutions? College Mathematics Journal, 48(4):284–287, September 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.48.4.284.

Bagyan:2024:CTM

[BR24] Armine Bagyan and Donald Richards. A continuous-time Markov chain model for the spread of COVID-19. College Mathematics Journal, 55(3):210–214, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2023.2291296.

Brazier:1980:WOC

[Bra80] Gerald D. Brazier. What are the odds? — constructing competition probabilities. Two-Year College Mathematics Journal, 11(5):290–295, November 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1980.11972464.

Braden:1985:PFC

[Bra85] Bart Braden. Picturing functions of a complex variable. College Mathematics Journal, 16(1):63–72, January 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1985.
11972856.

Braden:1986:SAF

[Bra86] Bart Braden. The surveyor’s area formula. College Mathematics Journal, 17(4):326–337, September 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1986.
11972974.

Braden:1990:CNC

[Bra90] Bart Braden. China’s 1989 National College Entrance Examination. College Mathematics Journal, 21(5):390–393, November 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1990.11973338.

Braden:1994:NE

[Bra94] Bart Braden. A note from the Editor. College Mathematics Journal, 25(3):178, May 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1994.11973605.

Bradstreet:2000:FDI

[Bra00] Thomas E. Bradstreet. Food and drug interaction: What role does statistics play? College Mathematics Journal, 31(4):268– 273, September 2000. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2000.11974157.

Bradley:2001:VLC

[Bra01a] David M. Bradley. Verhulst’s logistic curve. College Mathematics Journal, 32(2):94–98, March 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2001.11921860.

Bragg:2001:AS

[Bra01b] Louis Bragg. Arctangent sums. College Mathematics Journal, 32(4):255–257, September 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2001.11921885.

Brannen:2001:SMC

[Bra01c] Noah Samuel Brannen. The Sun, the Moon, and convexity. College Mathematics Journal, 32(4):268–272, September 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2001.11921888.

Bradley:2013:BRJ

[Bra13] Robert E. Bradley. Book review: Journey through Mathematics: Creative Episodes in Its History by Enrique A. González-Velasco. College Mathematics Journal, 44(3):241–243, May 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.44.3.241.

Bradie:2016:YMW

[Bra16] Brian Bradie. Yet more ways to skin a definite integral. College Mathematics Journal, 47(1):11–18, January 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
47.1.11.

Bradley:2019:MDC

[Bra19] David M. Bradley. The method of determined coefficients. College Mathematics Journal, 50(1):46–53, January 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2019.
1535106.

Brenner:1983:LG

[Bre83] J. L. Brenner. Labeling of graphs. Two-Year College Mathematics Journal, 14(1):36–41, January 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1983.
11972649.

Bressoud:1992:HSW

[Bre92] David M. Bressoud. How should we introduce integration? College Mathematics Journal, 23(4):296–298, September 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1992.11973472.

Bressoud:2002:WCI

[Bre02] David Bressoud. Was calculus invented in India? College Mathematics Journal, 33(1):2–13, January 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2002.
11921911.

Brenton:2008:RWG

[Bre08] Lawrence Brenton. Remainder wheels and group theory. College Mathematics Journal, 39(2):129–135, March 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2008.
11922286.

Bridger:1975:LHR

[Bri75a] Clyde A. Bridger. Leonardo, his rabbits, and other curiosa. Two-Year College Mathematics Journal, 6(1):14–20, February 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1975.11974361.

Brieske:1975:FFR

[Bri75b] Thomas J. Brieske. Factoring functions and relations. Two-Year College Mathematics Journal, 6(3):8–12, September 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1975.11974384.

Brieske:1978:MDC

[Bri78] Thomas J. Brieske. Mapping, diagrams, continuous functions and derivatives. Two-Year College Mathematics Journal, 9(2):67–72, March 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1978.11974549.

Brieske:1984:VTA

[Bri84] Tom Brieske. Visual thinking about rotations and reflections. College Mathematics Journal, 15(5):406–410, November 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972822.

Bridger:1988:LMS

[Bri88] Mark Bridger. Looking at the Mandelbrot set. College Mathematics Journal, 19(4):353–363, September 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1988.
11973139.

Bridger:1991:IEI

[Bri91] Mark Bridger. Image expansion in integer arithmetic. College Mathematics Journal, 22(5):429–435, November 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1991.11973423.

Bridger:1992:CCV

[Bri92] Mark Bridger. Calculus and computer vision. College Mathematics Journal, 23(2):132–141, March 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1992.
11973444.

Bridger:1996:DFV

[Bri96] Mark Bridger. Dynamic function visualization. College Mathematics Journal, 27(5):361–369, November 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1996.
11973809.

Brilleslyper:2020:URQ

[Bri20] Michael A. Brilleslyper. Unimodular roots of quadrinomials. College Mathematics Journal, 51(3):219–221, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/full/10.1080/07468342.
2020.1701384.

Broman:1974:CCI

[Bro74] Arne Broman. Computer computation of integrals. Two-Year College Mathematics Journal, 5(4):4–11, December 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1974.11974343.

Brooks:1980:CSA

[Bro80] Sarah Brooks. Computer solution of alphametics. Two-Year College Mathematics Journal, 11(2):111–114, March 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1980.11972416.

Bronson:1982:CC

[Bro82] Richard Bronson. Computers & calculators. Two-Year College Mathematics Journal, 13(5):300–310, November 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1982.11972637.

Brozinsky:1984:RPE

[Bro84] Michael K. Brozinsky. Reflection property of the ellipse and the hyperbola. College Mathematics Journal, 15(2):140–142, March 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972763.

Brodie:1999:RRC

[Bro99a] Scott E. Brodie. On “rethinking rigor in calculus …,” or why we don’t do calculus on the rational numbers. College Mathematics Journal, 30(2):135–138, March 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1999.
11974046.

Brown:1999:SRD

[Bro99b] Ezra Brown. Square roots from 1; 24, 51, 10 to Dan Shanks. College Mathematics Journal, 30(2):82–95, March 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1999.11974038.

Brown:2000:TFT

[Bro00] Ezra Brown. Three Fermat trails to elliptic curves. College Mathematics Journal, 31(3):162–172, May 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2000.
11974137.

Brown:2001:MSF

[Bro01] Ezra Brown. Magic squares, finite planes, and points of inflection on elliptic curves. College Mathematics Journal, 32(4):260– 267, September 2001. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2001.11921887.

Brothers:2004:ICN

[Bro04] Harlan J. Brothers. Improving the convergence of Newton’s series approximation for e. College Mathematics Journal, 35(1):34–39, January 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2004.11922049.

Brown:2005:PF

[Bro05] Ezra Brown. Phoebe floats! College Mathematics Journal, 36(2):114–122, March 2005. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2005.11922118.

Brown:2006:IHW

[Bro06] Scott H. Brown. An interview with H. W. Gould. College Mathematics Journal, 37(5):370–379, November 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2006.11922207.

Brooks:2007:NMT

[Bro07] David Alan Brooks. A new method of trisection. College Mathematics Journal, 38(2):78–81, March 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2007.
11922221.

Brown:2015:SST

[Bro15] Ezra Brown. Saints and scoundrels and two theorems that are really the same. College Mathematics Journal, 46(5):326– 334, November 2015. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.46.5.326.

Brown:2023:CDR

[Bro23] Jacob Brown. Counting divisions of a 2 × n rectangular grid. College Mathematics Journal, 54(3):212–221, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2023.2201166.

Brunson:1983:CEI

[Bru83] Pansy Waycaster Brunson. A classroom experiment involving basic mathematics and women. Two-Year College Mathematics Journal, 14(4):318–321, September 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1983.
11972705.

Burris:1972:DMS

[BS72] Joanna S. Burris and Lee Schroeder. Developmental mathematics: Self-instruction with mathematics laboratory. Two-Year College Mathematics Journal, 3(1):16–22, Spring 1972. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1972.11974191.

Bundrick:1978:NPM

[BS78] Charles M. Bundrick and David L. Sherry. A note on the principle of mathematical induction. Two-Year College Mathematics Journal, 9(1):17–18, January 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1978.11974539.

Beauregard:1996:PTH

[BS96] Raymond A. Beauregard and E. R. Suryanarayan. Pythagorean triples: The hyperbolic view. College Mathematics Journal, 27 (3):170–181, May 1996. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1996.11973772.

Beauregard:1998:BT

[BS98] Raymond A. Beauregard and E. R. Suryanarayan. The Brahmagupta triangles. College Mathematics Journal, 29(1):13–17, January 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1998.11973907.

Beauregard:2000:GAT

[BS00] Raymond A. Beauregard and E. R. Suryanarayan. General arithmetic triangles and Bhaskara’s equation. College Mathematics Journal, 31(2):111–115, March 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2000.
11974124.

Beauregard:2002:RRV

[BS02] Raymond A. Beauregard and E. R. Suryanarayan. A Ramanujan result viewed from matrix algebra. College Mathematics Journal, 33(3):212–214, May 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2002.11921943.

Boskoff:2004:WEL

[BS04] Wladimir G. Boskoff and Bogdan D. Suceava. When is Euler’s line parallel to a side of a triangle? College Mathematics Journal, 35(4):292–296, September 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2004.11922088.

Bohorquez:2014:ARG

[BS14a] Louis Bohorquez and Jennifer Switkes. American roulette: a gambler’s ruin. College Mathematics Journal, 45(1):33–40, January 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.45.1.033.

Brilleslyper:2014:LUR

[BS14b] Michael Brilleslyper and Beth Schaubroeck. Locating unimodular roots. College Mathematics Journal, 45(3):162–168, May 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.45.3.162.

Beasley:2015:JPS

[BS15] Brian D. Beasley and David R. Stone. Journal problems sections: Modern challenges and teaching tools. College Mathematics Journal, 46(5):336–346, November 2015. CODEN ????

ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
46.5.336.

Beineke:2023:MH

[BS23] Lowell Beineke and Philip Straffin. Media highlights. College Mathematics Journal, 54(4):406–413, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2023.
2237382.

Blum-Smith:2017:FTS

[BSC17] Ben Blum-Smith and Samuel Coskey. The Fundamental Theorem on Symmetric Polynomials: History’s first whiff of Galois theory. College Mathematics Journal, 48(1):18–29, January 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.48.1.18.

Bradley:1999:AH

[BSF99] Michael J. Bradley, Michael St. Vincent, and David L. Finn. Additivity ⊕ homogeneity. College Mathematics Journal, 30(2): 133–135, March 1999. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1999.11974045.

Bary-Soroker:2009:RTT

[BSL09] Lior Bary-Soroker and Eli Leher. On the remainder in the Taylor theorem. College Mathematics Journal, 40(5):373–375, November 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/074683409X475706.

Bookman:1985:SR

[BSR85] Jack Bookman, David Smith, and Donald E. Rossi. Software reviews. College Mathematics Journal, 16(3):218–222, June 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1985.11972886.

Bains:2007:NPR

[BT07] Manjinder S. Bains and J. B. Thoo. The normals to a parabola and the real roots of a cubic. College Mathematics Journal, 38 (4):272–277, September 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2007.11922248.

Bajnok:2024:RAU

[BT24] Béla Bajnok and Enrique Treviño. Report on the 14th Annual USA Junior Mathematical Olympiad. College Mathematics Journal, 55(1):74–82, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2023.2269804.

Boman:2002:WA

[BU02] Eugene Boman and Frank Uhlig. When 1/(a + b) = 1/a + 1/b, anyway? College Mathematics Journal, 33(4):296–300, September 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2002.11921956.

Buckhiester:1983:BD

[Buc83] Philip G. Buckhiester. Blackjack with n decks. Two-Year College Mathematics Journal, 14(4):345–346, September 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1983.11972711.

Buchthal:1985:PD

[Buc85a] David C. Buchthal. Predetermined determinants. College Mathematics Journal, 16(4):277–279, September 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1985.
11972894.

Buck:1985:BMR

[Buc85b] Frank Buck. Behold! The midpoint rule is better than the trapezoidal rule for concave functions. College Mathematics Journal, 16(1):56, January 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1985.11972852.

Buckley:1987:FLP

[Buc87] Fred Buckley. Facility location problems. College Mathematics Journal, 18(1):24–32, January 1987. CODEN ????

ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1987.
11973002.

Bukowski:2008:CHP

[Buk08] John Bukowski. Christiaan Huygens and the problem of the hanging chain. College Mathematics Journal, 39(1):2–11, January 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2008.11922269.

Bumcrot:1983:ETD

[Bum83] Robert Bumcrot. Exactly n-times differentiable functions. Two-Year College Mathematics Journal, 14(3):258–259, June 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1983.11972694.

Bumcrot:1984:SSL

[Bum84] Robert J. Bumcrot. Some subtleties in L’Hôpital’s Rule. College Mathematics Journal, 15(1):51–52, January 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1984.
11972745.

Bumcrot:1986:IME

[Bum86] Robert J. Bumcrot. Irrationality made easy. College Mathematics Journal, 17(3):243–244, May 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1986.
11972964.

Burditt:1972:TPW

[Bur72] Chris Burditt. Two-Pan weighings. Two-Year College Mathematics Journal, 3(2):80–81, Autumn 1972. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1972.
11974224.

Burns:1977:SER

[Bur77] James A. Burns. Some effects of rationing. Two-Year College Mathematics Journal, 8(4):203–206, September 1977.

CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1977.11974510.

Burgstahler:1983:GCR

[Bur83a] Sylvan Burgstahler. A generalization of Cramer’s Rule. Two-Year College Mathematics Journal, 14(3):203–205, June 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1983.11972683.

Burk:1983:MI

[Bur83b] Frank Burk. Mean inequalities. Two-Year College Mathematics Journal, 14(5):431–434, November 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1983.
11972734.

Burgstahler:1984:ACP

[Bur84] Sylvan Burgstahler. An alternative for certain partial fractions. College Mathematics Journal, 15(1):57–58, January 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972749.

Burk:1985:EC

[Bur85] Frank Burk. Euler’s constant. College Mathematics Journal, 16(4):279, September 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1985.11972895.

Burk:1986:NII

[Bur86a] Frank Burk. Numerical integration via integration by parts. College Mathematics Journal, 17(5):418–422, November 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1986.11972993.

Burk:1986:SIC

[Bur86b] Frank Burk. Some interesting consequences of a hyperbolic inequality. College Mathematics Journal, 17(1):75–76, January 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1986.11972935.

Burk:1987:R

[Bur87a] Frank Burk. π/4 and ln 2 recursively. College Mathematics Journal, 18(1):51, January 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1987.11973009.

Burk:1987:AQS

[Bur87b] Frank Burk. Archimedes’ quadrature and Simpson’s Rule. College Mathematics Journal, 18(3):222–223, May 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/07468342.
1987.11973036.

Burns:1992:FSR

[Bur92] Anne M. Burns. The flowering of string rewriting systems. College Mathematics Journal, 23(3):225–236, May 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1992.11973463.

Burk:1994:FFF

[Bur94] Frank Burk. Fallacies, flaws, and flimflam. College Mathematics Journal, 25(4):309–311, September 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1994.
11973626.

Burns:1995:MMS

[Bur95] Anne M. Burns. Making mountains from a sum of molehills. College Mathematics Journal, 26(1):51–57, January 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1995.11973669.

Burns:1998:MTS

[Bur98] Anne M. Burns. Modeling trees with a stochastic matrix. College Mathematics Journal, 29(3):230–236, May 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1998.11973948.

Burk:1999:NLL

[Bur99] Frank Burk. Natural logarithms via long division. College Mathematics Journal, 30(4):309–311, September 1999.

CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1999.11974076.

Burk:2000:SSI

[Bur00] Frank Burk. Summing series via integrals. College Mathematics Journal, 31(3):178–181, May 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2000.
11974139.

Burgiel:2016:BR

[Bur16] Heidi Burgiel. Book review. College Mathematics Journal, 47 (3):228–231, May 2016. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.47.3.228.

Bushaw:1971:FDD

[Bus71] D. Bushaw. Functions defined by differential equations: a short course in trigonometry. Two-Year College Mathematics Journal, 2(1):32–35, Spring 1971. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1971.11973994.

Bushaw:1996:MH

[Bus96] Donald W. Bushaw. Media highlights. College Mathematics Journal, 27(5):410–417, November 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1996.
11973820.

Butts:1981:FPI

[But81] Thomas Butts. Fixed point iteration — an interesting way to begin a calculus course. Two-Year College Mathematics Journal, 12(1):2–7, January 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1981.11972481.

Butler:2003:TLT

[But03] Steven Butler. Tangent line transformations. College Mathematics Journal, 34(2):105–106, March 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2003.
11921991.

Byrd:1991:CCC

[BW91] Stan Byrd and Terry Walters. Classroom computer capsule: Graphs of rational functions for computer assisted calculus. College Mathematics Journal, 22(4):332–334, September 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1991.11973409.

Bleecker:1998:WBT

[BW98] David D. Bleecker and Lawrence J. Wallen. The world’s biggest taco. College Mathematics Journal, 29(1):2–12, January 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1998.11973906.

Boudreaux:2013:DCS

[BW13] Gregory Mark Boudreaux and Jess E. Walls. Descartes’ calculus of subnormals: What might have been. College Mathematics Journal, 44(5):409–420, November 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
44.5.409.

Bergner:2015:GCE

[BW15] Julia E. Bergner and Christopher D. Walker. Groupoid cardinality and Egyptian fractions. College Mathematics Journal, 46(2):122–129, March 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.46.2.122.

Bell:2019:MER

[BW19] Lindsey Bell and Rachel Wagner. Modeling emergency room arrivals using the Poisson process. College Mathematics Journal, 50(5):343–350, 2019. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2019.1662710.

Bernhardt:1997:PPD

[BY97] Chris Bernhardt and Thomas Yuster. Periodic points of the difference operator. College Mathematics Journal, 28(1):20–26, January 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1997.11973824.

Byrne:2011:TTA

[Byr11] Jesse Byrne. Teaching tip: Actuarial science and Gompertz’s Law of Mortality. College Mathematics Journal, 42(1):40–42, January 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2011.11922472.

Beauregard:2004:TDG

[BZ04] Raymond A. Beauregard and Konstantine D. Zelator. On a three-dimensional generalization of Fermat’s Area Theorem. College Mathematics Journal, 35(4):289–291, September 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2004.11922087.

Bridger:2005:VR

[BZ05] Mark Bridger and Andrei Zelevinsky. Visibles revisited. College Mathematics Journal, 36(4):289–300, September 2005. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2005.11922142.

Caglayan:2015:PWS

[Cag15] Gunhan Caglayan. Proof without words: Series of reciprocals of tetrahedral numbers. College Mathematics Journal, 46(2):130, March 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.46.2.130.

Caglayan:2020:PWC

[Cag20a] Günhan Caglayan. Proof without words: Centered nonagonal numbers are triangular. College Mathematics Journal, 51(5): 371, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2020.1820287.

Caglayan:2020:PWS

[Cag20b] Günhan Caglayan. Proof without words: Sums of polygonal numbers. College Mathematics Journal, 51(4):304, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2020.1807239.

Caglayan:2023:CSN

[Cag23] Günhan Caglayan. Covering a star number with pentagonal numbers. College Mathematics Journal, 54(1):63, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2022.2153554.

Caglayan:2024:PWRb

[Cag24a] Günhan Caglayan. Proof without words: a relationship of heptagonal and nonagonal numbers. College Mathematics Journal, 55(1):58, 2024. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2023.2225384.

Caglayan:2024:PWRa

[Cag24b] Günhan Caglayan. Proof without words: a relationship of pentagonal and hexagonal numbers. College Mathematics Journal, 55(1):56–57, 2024. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2023.2224731.

Caglayan:2024:RTS

[Cag24c] Günhan Caglayan. A relationship of triangular, square and pentagonal numbers. College Mathematics Journal, 55(2):159– 160, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2023.2234798.

Cairns:1973:TRS

[Cai73] Stewart S. Cairns. Topological regular solids. Two-Year College Mathematics Journal, 4(1):74–76, Winter 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/00494925.
1973.11974242.

Callan:1995:PTV

[Cal95] David Callan. Pair them up! A visual approach to the Chung– Feller Theorem. College Mathematics Journal, 26(3):196–198, May 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1995.11973695.

Callan:1998:WRA

[Cal98]  David Callan. When is “rank” additive?  College Math-

ematics Journal, 29(2):145–147, March 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1998.
11973931.

Calvis:2000:PDY

[Cal00] David Calvis. Partially differentiable, yes; continuous, no. College Mathematics Journal, 31(1):42–47, January 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2000.
11974106.

Campbell:1982:BR

[Cam82] Paul J. Campbell. Book reviews. Two-Year College Mathematics Journal, 13(3):217–218, June 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1982.
11972616.

Campbell:1991:SR

[Cam91] Douglas Campbell. Software reviews. College Mathematics Journal, 22(2):163–166, March 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1991.11973377.

Campuzano:2019:CEF

[Cam19] Juan Carlos Ponce Campuzano. Conics as envelopes of families of plane curves. College Mathematics Journal, 50(2):115– 122, 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2019.1560207.

Cannon:1986:GRP

[Can86] Larry Cannon. Geometry of the rational plane. College Mathematics Journal, 17(5):392–402, November 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1986.
11972986.

Cannon:1989:LMP

[Can89] Lawrence O. Cannon. Locating multiples of primes in Pascal’s Triangle. College Mathematics Journal, 20(4):324–328, September 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1989.11973253.

Capitan:2017:HPE

[Cap17] Francisco Javier García Capitán. Homographic pencils for the ellipse and the hyperbola. College Mathematics Journal, 48(2): 134–136, March 2017. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.48.2.134.

Carter:1970:MMP

[Car70] Larry D. Carter. Mini-math: a program of short courses. Two-Year College Mathematics Journal, 1(2):36–38, Autumn 1970. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1970.11973185.

Carleton:1972:UPD

[Car72] D. L. Carleton. Using polyhedrons to define maximum volumes. Two-Year College Mathematics Journal, 3(1):30–32, Spring 1972. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1972.11974193.

Carter:1975:SMP

[Car75] Nancy F. Carter. A survey of mathematics programs. Two-Year College Mathematics Journal, 6(4):14–16, December 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1975.11974400.

Carter:1977:WIG

[Car77] Nancy F. Carter. What’s it good for? Two-Year College Mathematics Journal, 8(2):79–80, March 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1977.
11974486.

Carpenter:1978:NLR

[Car78a] James E. Carpenter. Note on L’Hôpital’s Rule for the indeterminate form ∞/∞. Two-Year College Mathematics Journal, 9 (2):73–74, March 1978. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1978.11974550.

Carson:1978:FNP

[Car78b] Judithlynne Carson. Fibonacci numbers and pineapple phyllotaxy. Two-Year College Mathematics Journal, 9(3):132–136, June 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1978.11974562.

Carchidi:1987:TSR

[Car87] Michael Carchidi. Two simple recursive formulas for summing 1 k + 2 ^{k} + 3 ^{k} + … + n^{k} . College Mathematics Journal, 18(5):406–409, November 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1987.11973065.

Carlson:1993:TLA

[Car93] David Carlson. Teaching linear algebra: Must the fog always roll in? College Mathematics Journal, 24(1):29–40, January 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1993.11973503.

Carchidi:1998:GEL

[Car98] Michael A. Carchidi. Generating exotic-looking vector spaces. College Mathematics Journal, 29(4):304–308, September 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1998.11973960.

Carpenter:2017:BR

[Car17] Jenna P. Carpenter. Book review. College Mathematics Journal, 48(1):64–68, January 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.48.1.64.

Cashing:1987:DPR

[Cas87] Douglas L. Cashing. Is the distributive property redundant? College Mathematics Journal, 18(5):402–403, November 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1987.11973063.

Case:1999:IAA

[Cas99] James Case. Interval arithmetic and analysis. College Mathematics Journal, 30(2):106–111, March 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1999.
11974040.

CastineiraMerino:2003:LFC

[Cas03] Julio Castiñeira Merino. Lissajous figures and Chebyshev polynomials. College Mathematics Journal, 34(2):122–127, March 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2003.11921995.

Conway:2020:CBD

[CAS20] John Conway, Gabrielle Agus, and David Slusky. Converting between dates in the Hebrew and Roman calendars. College Mathematics Journal, 51(5):322–329, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2020.
1814651.

Catone:2019:BCD

[Cat19] Christopher J. Catone. Bringing calculus into discrete math via the discrete derivative. College Mathematics Journal, 50(1): 21–27, January 2019. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2019.1530553.

Cavallini:1993:FLC

[Cav93] Fabio Cavallini. Fitting a logistic curve to data. College Mathematics Journal, 24(3):247–253, May 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1993.
11973540.

Cavior:1996:MKT

[Cav96] Stephen R. Cavior. A master key for ten locks. College Mathematics Journal, 27(1):33–36, January 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1996.
11973746.

Coxeter:1985:MH

[CB85] H. S. M. Coxeter and R. P. Boas. Media highlights. College Mathematics Journal, 16(2):162–171, March 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1985.
11972875.

Colley:2009:BR

[CB09] Susan Jane Colley and Mark Bollman. Book reviews. College Mathematics Journal, 40(3):222–225, May 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/07468342.
2009.11922366.

Corey:2019:STT

[CBL19] Douglas Lyman Corey, Jacob Badger, and Steven Lauzon. Spirals, triangles, and tie-dyed T-shirts. College Mathematics Journal, 50(4):250–259, 2019. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2019.1655374.

Chee:1980:GPF

[CC80] P. S. Chee and S. T. Chin. On the general power function. Two-Year College Mathematics Journal, 11(1):51, January 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1980.11972401.

Cohen:1995:CSD

[CC95] Steve M. Cohen and Paul R. Coe. Card shuffling in discrete mathematics. College Mathematics Journal, 26(3):224–227, May 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1995.11973702.

Chen:2009:CT

[CC09] Hang Chen and Curtis Cooper. n-card tricks. College Mathematics Journal, 40(3):196–203, May 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2009.
11922360.

Chan:2019:DOS

[CC19] Kevin L. T. Chan and Wai-Sum Chan. Developing an optimal strategy for a maximization dice game. College Mathematics Journal, 50(4):272–279, 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2019.1632662.

Chakraborty:2023:GEL

[CC23] Bikash Chakraborty and Sagar Chakraborty. A generalization of Euler’s limit. College Mathematics Journal, 54(2):140– 141, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2023.2183543.

Conrad:1976:NAC

[CCH76] Clifford L. Conrad, Nancy J. Conrad, and Harry B. Higley. A new approach for computer mathematics. Two-Year College Mathematics Journal, 7(2):34–39, May 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1976.
11974427.

Carlson:2021:PFC

[CCH^{+}21] Joshua Carlson, Alex Christensen, Pamela E. Harris, Zakiya Jones and Andrés Ramos Rodríguez. Parking functions: Choose your own adventure. College Mathematics Journal, 52 (4):254–264, 2021. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2021.1943115.

Cericola:2024:AUG

[CCHR24] Christopher Cericola, Justus Curry, Allison Henrich, and Mitchell Rask. The arc unknotting game. College Mathematics Journal, 55(4):280–293, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2023.2285866.

Carrillo:2025:GTP

[CCKL25] Angel Carrillo, Jonathan Cervantes, Mike Krebs, and Francisco Leon. A graph theorist plants a tree. College Mathematics Journal, 56(1):42–48, 2025. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic).

Clark:1993:AMA

[CD93] Dean Clark and Dilip K. Datta. Arithmetic matrices and the amazing nine-card Monte. College Mathematics Journal, 24(1): 52–56, January 1993. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1993.11973506.

Coll:2014:ISB

[CDH^{+}14] Vincent Coll, Jeremy Davis, Martin Hall, Colton Magnant, James Stankewicz, and Hua Wang. Integer solutions to Box optimization problems. College Mathematics Journal, 45(3):180–190, May 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.45.3.180.

Cooney:1990:RRS

[CDK^{+}90] Miriam P. Cooney, Jacqueline M. Dewar, Patricia Clark Kenschaft, Vivian Kraines, Brenda Latka, and Barbara LiSanti. Recruitment and retention of students in undergraduate mathematics. College Mathematics Journal, 21(4):294–301, September 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1990.11973322.

Czarnocha:2001:CAS

[CDL^{+}01] Bronislaw Czarnocha, Ed Dubinsky, Sergio Loch, Vrunda Prabhu, and Draga Vidakovic. Conceptions of area: In students and in history. College Mathematics Journal, 32(2):99–109, March 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2001.11921861.

Cahill:2002:FD

[CDNN02] Nathan D. Cahill, John R. D’Errico, Darren A. Narayan, and Jack Y. Narayan. Fibonacci determinants. College Mathematics Journal, 33(3):221–225, May 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2002.
11921945.

Carvalho:2017:GDW

[CdSS17] Alda Carvalho, Carlos Pereira dos Santos, and Jorge Nuno Silva. The geometer Dog who did not know calculus. College Mathematics Journal, 48(5):339–345, November 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
48.5.339.

Chakerian:2023:HA

[CE23] Don Chakerian and Stephen Erfle. Up the hill and down again. College Mathematics Journal, 54(4):256–268, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2023.2223509.

Ceasar:1981:SPP

[Cea81] Allan J. Ceasar. The Saint Petersburg Paradox and some related series. Two-Year College Mathematics Journal, 12(5):306– 308, November 1981. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1981.11972556.

Ceasar:1984:MCS

[Cea84] Allan J. Ceasar. A Monte Carlo simulation related to the St. Petersburg Paradox. College Mathematics Journal, 15(4):339– 342, September 1984. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1984.11972797.

Cella:1999:RC

[Cel99] Paul Cella. Reexamining the catenary. College Mathematics Journal, 30(5):391–393, November 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1999.
11974093.

Chartrand:1988:HDI

[CEO88] Gary Chartrand, Paul Erdős, and Ortrud R. Oellermann. How to define an irregular graph. College Mathematics Journal, 19(1): 36–42, January 1988. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1988.11973088.

Cereceda:2015:EFF

[Cer15] José Luis Cereceda. Explicit form of the Faulhaber polynomials. College Mathematics Journal, 46(5):359–363, November 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.46.5.359.

Ciaurri:2017:VAG

[CF17] Óscar Ciaurri and Emilio Fernández. Variations on an Archimedean ground: The generalized Salinon. College Mathematics Journal, 48(5):355–365, November 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
48.5.355.

Ciaurri:2012:VPP

[CFLR12] Óscar Ciaurri, Emilio Fernández, Rodolfo Larrea, and Luz Roncal. Viète’s product proved in the finest ancient style. College Mathematics Journal, 43(4):291–296, September 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
43.4.291.

Connors:2014:CLL

[CG14] Darcie Connors and Darren Glass. Chutes and ladders with large spinners. College Mathematics Journal, 45(4):289–295, September 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.45.4.289.

Carlisle:2019:SSD

[CG19] Sylvia Carlisle and William R. Green. Solving systems of differential equations in the case of a defective coefficient matrix. College Mathematics Journal, 50(5):372–374, 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2019.1667200.

Constans:2024:HFT

[CG24] Eric Constans and Nicola Golfari. How to fold a triangle: The connection between the Grashof Fourbar linkage and Pythagorean triangles. College Mathematics Journal, 55(1):12– 23, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2023.2213156.

Chapman:2021:DBF

[CGSO21] Scott Chapman, Pedro García-Sánchez, and Christopher O’Neill. Distances between factorizations in the Chicken McNugget monoid. College Mathematics Journal, 52(3):158–176, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2021.1909953.

Ho:1994:SGP

[cH94] Jeffrey Li chieh Ho. A simple geometric proof of the addition formula for the sine. College Mathematics Journal, 25(3):229– 230, May 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1994.11973614.

Clark:1999:ESS

[CH99] K. E. Clark and S. Hill. The effects of a stiffening spring. College Mathematics Journal, 30(5):379–382, November 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1999.11974089.

Carchidi:2017:CBT

[CH17] Michael A. Carchidi and Robert L. Higgins. Covariances between transient states in finite absorbing Markov chains. College Mathematics Journal, 48(1):42–50, January 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
48.1.42.

Charpentier:1977:WM

[Cha77] Debra Charpentier. Women mathematicians. Two-Year College Mathematics Journal, 8(2):73–79, March 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/00494925.
1977.11974485.

Chamberlain:1979:CTP

[Cha79] Michael W. Chamberlain. Coin-tossing problem revisited. Two-Year College Mathematics Journal, 10(5):349–350, November 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1979.11972384.

Chakerian:1980:CS

[Cha80] G. D. Chakerian. Circles and spheres. Two-Year College Mathematics Journal, 11(1):26–41, January 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1980.
11972397.

Chamberlain:1983:CC

[Cha83] Michael W. Chamberlain. Computers & calculators. Two-Year College Mathematics Journal, 14(2):159–164, March 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1983.11972672.

Chakerian:1984:MH

[Cha84] Don Chakerian. Media highlights. College Mathematics Journal, 15(2):170–177, March 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1984.11972771.

Chakerian:1991:IQA

[Cha91] G. D. Chakerian. The isoperimetric quotient: Another look at an old favorite. College Mathematics Journal, 22(4):313– 315, September 1991. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1991.11973403.

Chacon:1992:RMM

[Cha92] Paul Chacon. Relative maxima or minima for a function of two variables: a neglected approach. College Mathematics Journal, 23(2):145–146, March 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1992.11973446.

Chamberland:1999:SI

[Cha99] Marc Chamberland. The series for e via integration. College Mathematics Journal, 30(5):397, November 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1999.
11974095.

Cha:2007:TFI

[Cha07] Byungchul Cha. Transcendental functions and initial value problems: a different approach to calculus II. College Mathematics Journal, 38(4):288–296, September 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2007.
11922250.

Chatham:2009:RQP

[Cha09] R. Douglas Chatham. Reflections on the N + k Queens problem. College Mathematics Journal, 40(3):204–210, May 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2009.11922361.

Chamberland:2013:PWA

[Cha13] Marc Chamberland. Proof without words: The area of an inner square. College Mathematics Journal, 44(4):322, September 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.44.4.322.

Chew:1981:AAV

[Che81] James Chew. An alternative approach to the vibrating string problem. Two-Year College Mathematics Journal, 12(2):147–149, March 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972514.

Chen:1995:RF

[Che95] Xuming Chen. Recursive formulas for ξ(2k) and L(2k − 1). College Mathematics Journal, 26(5):372–376, November 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1995.11973732.

Chen:1996:SCP

[Che96] Zhibo Chen. Some comments on “parametric equations and plane curves”. College Mathematics Journal, 27(3):210–211, May 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1996.11973780.

Chen:2005:PDM

[Che05] Yung-Pin Chen. On primes, density measures, and statistical independence. College Mathematics Journal, 36(4):284–288, September 2005. CODEN ???? ISSN 0746-8342 (print), 1931-

1346 (electronic). URL http://www.tandfonline.com/doi/
abs/10.1080/07468342.2005.11922141.

Chen:2009:FIR

[Che09] Hongwei Chen. The Fresnel integrals revisited. College Mathematics Journal, 40(4):259–262, September 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2009.
11922373.

Chen:2013:NRT

[Che13] Hongwei Chen. A new ratio test for positive monotone series. College Mathematics Journal, 44(2):139–141, March 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.44.2.139.

Chen:2015:BR

[Che15] Jiang-Ping Jeff Chen. Book review. College Mathematics Journal, 46(2):149–152, March 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.46.2.149.

Chen:2018:STL

[Che18] Jiang-Ping Jeff Chen. A systematic treatment of “linear algebra” in 17th-century China. College Mathematics Journal, 49 (3):169–179, May 2018. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2018.1448670.

Chen:2022:CET

[Che22] William Y. C. Chen. Cubic equations through the looking glass of Sylvester. College Mathematics Journal, 53(5):396– 398, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2022.2118991.

Chimenti:1973:ALR

[Chi73] F. A. Chimenti. Arc length revisited. Two-Year College Mathematics Journal, 4(3):88–89, Autumn 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1973.
11974282.

Chinn:1975:NN

[Chi75] William G. Chinn. News and notices. Two-Year College Mathematics Journal, 6(4):21–23, December 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1975.
11974402.

Chinn:1976:NN

[Chi76] William G. Chinn. News and notices. Two-Year College Mathematics Journal, 7(1):27, February 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1976.
11974417.

Chinn:1982:TNM

[Chi82] William G. Chinn. Two new mathematics research institutes. Two-Year College Mathematics Journal, 13(2):162–163, March 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1982.11972601.

Cheteyan:2011:CLI

[CHJ11] Leslie A. Cheteyan, Stewart Hengeveld, and Michael A. Jones. Chutes and ladders for the impatient. College Mathematics Journal, 42(1):2–8, January 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.42.1.002.

Chosid:1979:Q

[Cho79] Leo Chosid. Queries. Two-Year College Mathematics Journal, 10(5):354–355, November 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1979.11972388.

Choike:1980:PDI

[Cho80a] James R. Choike. The pentagram and the discovery of an irrational number. Two-Year College Mathematics Journal, 11 (5):312–316, November 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1980.11972468.

Chosid:1980:QTR

[Cho80b] Leo Chosid. A quick test for rational roots of a polynomial. Two-Year College Mathematics Journal, 11(3):205–206, June 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1980.11972439.

Chosid:1980:TSQ

[Cho80c] Leo Chosid. Trigonometric solutions to the quadratic equation. Two-Year College Mathematics Journal, 11(5):330–331, November 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1980.11972474.

Chosid:1981:CES

[Cho81] Leo Chosid. Compounding energy savings. Two-Year College Mathematics Journal, 12(1):56–57, January 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1981.
11972495.

Chong:1982:DOA

[Cho82] Kong-Ming Chong. Differential operators applied to integration. Two-Year College Mathematics Journal, 13(2):155–157, March 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1982.11972597.

Choate:1991:SR

[Cho91] Jonathan Choate. Software reviews. College Mathematics Journal, 22(1):65–69, January 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1991.11973362.

Christian:1983:AWI

[Chr83] Robert R. Christian. Another way to introduce natural logarithms and e. Two-Year College Mathematics Journal, 14(5): 424–426, November 1983. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1983.11972732.

Christensen:1996:NMR

[Chr96] Chris Christensen. Newton’s Method for resolving affected equations. College Mathematics Journal, 27(5):330–340, November 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1996.11973804.

Cole:1986:PMS

[CHS86] David Cole, Loren Haarsma, and Jack Snoeyink. The problem of managing a strategic reserve. College Mathematics Journal, 17(1):48–60, January 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1986.11972929.

Cronk:1987:PEF

[CHS87] Jeff Cronk, Duff Howell, and Keith Saints. Positioning of emergency facilities in an obstructed traffic grid. College Mathematics Journal, 18(1):34–43, January 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1987.
11973004.

Chuang:2009:MSB

[Chu09] Jer-Chin (Luke) Chuang. Minimal solutions to the Box problem. College Mathematics Journal, 40(5):354–360, November 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/074683409X475670.

Ciaurri:2017:PWA

[Cia17] Óscar Ciaurri. Proof without words: an arithmetic–geometric series. College Mathematics Journal, 48(1):41, January 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.48.1.41.

Cibes:1993:OCC

[Cib93] Margaret Cibes. Old calculus chestnuts: Roast, or light a fire? College Mathematics Journal, 24(3):241–243, May 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1993.11973537.

Cicenia:1986:PIV

[Cic86] Robert A. Cicenia. Picturing infinite values. College Mathematics Journal, 17(4):322–325, September 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1986.
11972973.

Cid:2021:ETD

[Cid21] José Ángel Cid. An excursion through the double sidedness of the matrix inverse. College Mathematics Journal, 52(1):54– 56, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2021.1848177.

Cipra:1987:DSC

[Cip87] Barry A. Cipra. The derivatives of the sine and cosine functions. College Mathematics Journal, 18(2):139–141, March 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1987.11973024.

Cigler:2023:MJB

[CJ23] Gregor Cigler and Marjan Jerman. Modifying Jordan bases. College Mathematics Journal, 54(3):222–231, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/full/10.1080/07468342.
2023.2201146.

Carlson:1992:GEE

[CJLP92] David Carlson, Charles R. Johnson, David Lay, and A. Duane Porter. Gems of exposition in elementary linear algebra. College Mathematics Journal, 23(4):299–303, September 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1992.11973473.

Carlson:1993:LAC

[CJLP93] David Carlson, Charles R. Johnson, David C. Lay, and A. Duane Porter. The Linear Algebra Curriculum Study Group recommendations for the first course in linear algebra. College Mathematics Journal, 24(1):41–46, January 1993. CODEN ????

ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1993.
11973504.

Chen:2012:HSM

[CK12] Hongwei Chen and Chris Kennedy. Harmonic series meets Fibonacci sequence. College Mathematics Journal, 43(3):237–243, May 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.43.3.237.

Clancy:2014:CLB

[CK14] Daniel T. Clancy and Steven J. Kifowit. A closer look at Bobo’s sequence. College Mathematics Journal, 45(3):199–206, May 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.45.3.199.

Capaldi:2017:CC

[CK17] Mindy Capaldi and Tiffany Kolba. Carcassonne in the classroom. College Mathematics Journal, 48(4):265–273, September 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.48.4.265.

Caulfield:1986:HGB

[CKM86] Michael Caulfield, John Kent, and Daniel McCaffrey. Harvesting a grizzly bear population. College Mathematics Journal, 17(1): 34–46, January 1986. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1986.11972928.

Chinn:1983:SSC

[CL83] William G. Chinn and John B. Lewis. Shiing-shen Chern a man and his times. Two-Year College Mathematics Journal, 14 (5):370–376, November 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1983.11972722.

Cheng:1985:BAL

[CL85] Sui-Sun Cheng and Tzon-Tzer Lu. The bisection algorithm is not linearly convergent. College Mathematics Journal, 16(1):56–57, January 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1985.11972853.

Clark:1990:HHM

[CL90] Dean S. Clark and James T. Lewis. Herbert and the Hungarian mathematician: Avoiding certain subsequence sums. College Mathematics Journal, 21(2):100–104, March 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1990.
11973291.

Cadeddu:2015:MMD

[CL15] Lucio Cadeddu and Giampaolo Lai. Maxima and minima without derivatives? College Mathematics Journal, 46(1):15–22, January 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.46.1.15.

Clark:1973:SSR

[Cla73] Colin Clark. Some socially relevant applications of elementary calculus. Two-Year College Mathematics Journal, 4(2):1–15, Spring 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1973.11974246.

Clark:1992:FFF

[Cla92] Dean Clark. Fallacies, flaws, and flimflam. College Mathematics Journal, 23(5):402–406, November 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1992.
11973490.

Clark:2011:DSP

[Cla11] Jeffrey Clark. Derivative sign patterns. College Mathematics Journal, 42(5):379–382, November 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
42.5.379.

Clark:2021:CCI

[Cla21] Thomas J. Clark. Classroom and computational investigations of Camel Up. College Mathematics Journal, 52(4):289– 296, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2021.1941538.

Cook:1983:CC

[CM83] Lyle Cook and James McWilliam. Computers & calculators. Two-Year College Mathematics Journal, 14(1):52–54, January 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1983.11972652.

Carlson:1990:RCL

[CM90] Stephan C. Carlson and Jerry M. Metzger. A recursively computed limit. College Mathematics Journal, 21(3):222–224, May 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1990.11973311.

Chen:1998:FDI

[CM98a] Yong-Zhuo Chen and Richard F. Melka. Finding a determinant and inverse matrix by bordering. College Mathematics Journal, 29(1):38–39, January 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1998.11973912.

Claus-McGahan:1998:SR

[CM98b] Elly Claus-McGahan. Software reviews. College Mathematics Journal, 29(5):427–433, November 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1998.
11973984.

Cherniavsky:2014:ZDC

[CM14a] Yonah Cherniavsky and Artour Mouftakhov. Zbikowski’s divisibility criterion. College Mathematics Journal, 45(1):17–21, January 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.45.1.017.

Cooper:2014:TDL

[CM14b] Charles L. Cooper and Michael S. McClendon. A topological definition of limits for use in elementary calculus. College Mathematics Journal, 45(4):313–315, September 2014. CODEN ????

ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
45.4.313.

Chatwin:2015:HWO

[CM15] Richard E. Chatwin and Dana Mackenzie. How to win at (one-round) war. College Mathematics Journal, 46(4):242–253, September 2015. CODEN ???? ISSN 0746-8342 (print), 1931-

1346 (electronic). URL http://www.tandfonline.com/doi/
abs/10.4169/college.math.j.46.4.242.

Ciesielski:2017:GPE

[CM17] Krzysztof Chris Ciesielski and David Miller. On a Genocchi-Peano example. College Mathematics Journal, 48(3):205–213, May 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.48.3.205.

Ciraulo:2020:SKK

[CM20] Francesco Ciraulo and Samuele Maschio. Solving knights-and-knaves with one equation. College Mathematics Journal, 51(2): 82–89, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2020.1698931.

Cotton:2015:CCF

[CMP15] Nicholas Cotton, Cam McLeman, and Daniel Pinchock. On combining and convolving fractals. College Mathematics Journal, 46(2):99–108, March 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.46.2.99.

Christian:2020:TRC

[CMS20] J. M. Christian and H. A. J. Middleton-Spencer. On the N th roots of −1 and complex basin boundaries: Fractals from Newton–Raphson. College Mathematics Journal, 51(2):95–104, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2020.1703452.

Crowley:1992:SR

[CO92] Lillie R. F. Crowley and J. Stephen Ott. Software review. College Mathematics Journal, 23(1):64–68, January 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1992.11973438.

Chappell:1999:TT

[CO99] James Chappell and Thomas J. Osler. The trinomial triangle. College Mathematics Journal, 30(2):141–142, March 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1999.11974048.

Coates:2011:IPS

[Coa11] Keith J. Coates. An intuitive proof of the singular value decomposition of a matrix. College Mathematics Journal, 42(5):394– 396, November 2011. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.42.5.394.

Cocks:1976:WOP

[Coc76] Orrin G. Cocks. The well-ordering principle as an alternative to mathematical induction in our lower division recursive formula proofs. Two-Year College Mathematics Journal, 7(1):13–14, February 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1976.11974413.

Coffin:2009:PPC

[Cof09] Stewart Coffin. Polyomino problems to confuse computers. College Mathematics Journal, 40(3):169–173, May 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2009.11922355.

Cogswell:1999:CAM

[Cog99] Kurt Cogswell. Computers and advanced mathematics in the calculus classroom. College Mathematics Journal, 30(3):213–216, May 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1999.11974060.

Cohen:1977:LE

[Coh77a] Donald Cohen. Letters to the Editor. Two-Year College Mathematics Journal, 8(5):259, November 1977. CODEN ????

ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1977.
11974523.

Cohen:1977:SIA

[Coh77b] Donald Cohen. The sequencing of instructional activities in written materials. Two-Year College Mathematics Journal, 8(2): 81–87, March 1977. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1977.11974487.

Colvin:1971:FPP

[Col71] Charles R. Colvin. The Fredonia Plan for preparing two-year college teachers. Two-Year College Mathematics Journal, 2(2): 69–73, Autumn 1971. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1971.11974015.

Collister:1975:CC

[Col75] Larry Collister. Computer corner. Two-Year College Mathematics Journal, 6(2):36–41, May 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1975.11974378.

Collins:1992:AS

[Col92] Raymond J. Collins. Approximating series. College Mathematics Journal, 23(2):153–157, March 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1992.
11973450.

Colley:1994:CB

[Col94] Susan Jane Colley. Calculus in the brewery. College Mathematics Journal, 25(3):226–227, May 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1994.
11973612.

Colley:2003:CMA

[Col03] Susan Jane Colley. Calculus, π, and the Machine Age. College Mathematics Journal, 34(4):264–269, September 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2003.11922016.

Colley:2010:BR

[Col10] Susan Jane Colley. Book reviews. College Mathematics Journal, 41(2):170–172, March 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/074683410X480294.

Colley:2013:BRM

[Col13] Susan Jane Colley. Book review: The Manga Guide to Linear Algebra, by Shin Takahashi, illustrated by Iroha Inoue and Math Girls, by Hiroshi Yuki (translated by Tony Gonzalez). College Mathematics Journal, 44(3):244–247, May 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.44.3.244.

Connolly:1999:CGU

[Con99] Dennis Connolly. Casino gambling, the ultimate strategy. College Mathematics Journal, 30(4):276–278, September 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1999.11974071.

Constans:2017:LSF

[Con17] Eric Constans. A Lagrangian simulation of the floating-arm trebuchet. College Mathematics Journal, 48(3):179–187, May 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.48.3.179.

Coon:1970:GWD

[Coo70] Lewis H. Coon. Gleanings from the Washington, D.C. NCTM Meeting. Two-Year College Mathematics Journal, 1(2):100–103, Autumn 1970. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1970.11973192.

Coon:1974:DDC

[Coo74] Lewis H. Coon. A doctor’s degree for community college teachers — why? Two-Year College Mathematics Journal, 5(4):22– 26, December 1974. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1974.11974347.

Cook:1982:SEF

[Coo82] F. Lee Cook. A simple explicit formula for the Bernoulli numbers. Two-Year College Mathematics Journal, 13(4):273–274, September 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1982.11972629.

Cooper:1984:AGF

[Coo84] Curtis Cooper. Application of a generalized Fibonacci sequence. College Mathematics Journal, 15(2):145–147, March 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972766.

Cooper:1986:TTR

[Coo86] Curtis Cooper. Trees and tennis rankings. College Mathematics Journal, 17(1):76–78, January 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1986.
11972936.

Cook:2013:DPT

[Coo13] William J. Cook. An n-dimensional Pythagorean Theorem. College Mathematics Journal, 44(2):98–101, March 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
44.2.098.

Cooper:2024:PWB

[Coo24] Thomas E. Cooper. Proof without words: Bisecting a quadrilateral’s perimeter. College Mathematics Journal, 55(2):161, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2023.2274250.

Costello:1994:CJC

[Cos94] Patrick Costello. Computing Jordan canonical forms. College Mathematics Journal, 25(3):231–234, May 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1994.
11973617.

Costello:2017:TSB

[Cos17] Patrick J. Costello. Tiling squares with big holes with Ltrominoes. College Mathematics Journal, 48(4):259–263, September 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.48.4.259.

Cowen:1976:SPR

[Cow76] R. H. Cowen. A simple proof of the reflection property for parabolas. Two-Year College Mathematics Journal, 7(2):59–60, May 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1976.11974433.

Cox:2012:BR

[Cox12] David A. Cox. Book review. College Mathematics Journal, 43(2):181–183, March 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.43.2.181.

Cox:2025:UC

[Cox25] Barry J. Cox. Unchained catenaries. College Mathematics Journal, 56(1):15–24, 2025. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic).

Caveny:1989:FPS

[CP89] James Caveny and Warren Page. The fundamental periods of sums of periodic functions. College Mathematics Journal, 20(1): 32–41, January 1989. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1989.11973200.

Craswell:1973:HFN

[Cra73a] Keith J. Craswell. How to find a needle in a haystack. Two-Year College Mathematics Journal, 4(3):18–21, Autumn 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1973.11974269.

Craswell:1973:IPG

[Cra73b] Keith J. Craswell. An interesting Penny game. Two-Year College Mathematics Journal, 4(1):18–25, Winter 1973.

CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1973.11974230.

Craswell:1974:WIP

[Cra74] Keith J. Craswell. Why isn’t Penny flipping fairer? Two-Year College Mathematics Journal, 5(3):18–19, September 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1974.11974329.

Crawford:1977:UIF

[Cra77] Robert C. Crawford. Using inverse functions in integration. Two-Year College Mathematics Journal, 8(2):107–109, March 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1977.11974493.

Crannell:2011:BR

[Cra11a] Annalisa Crannell. Book review. College Mathematics Journal, 42(3):241–243, May 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.42.3.241.

Crannell:2011:SFT

[Cra11b] Annalisa Crannell. The Shad–Fack transom. College Mathematics Journal, 42(4):309–316, September 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
42.4.309.

Crachiola:2012:PWP

[Cra12] Anthony J. Crachiola. Proof without words: Partial sums of an arithmetic sequence. College Mathematics Journal, 43(4): 321, September 2012. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.43.4.321.

Crisman:2011:SGP

[Cri11] Karl-Dieter Crisman. The symmetry group of the permutahedron. College Mathematics Journal, 42(2):135–139, March 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.42.2.135.

Crofoot:1995:CS

[Cro95] R. Bruce Crofoot. Continuity on a set. College Mathematics Journal, 26(1):29–30, January 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1995.11973662.

Carroll:2010:WTL

[CRS10] Maureen T. Carroll, Elyn K. Rykken, and Jody M. Sorensen. Waiting to turn left? College Mathematics Journal, 41(1):60–63, January 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/074683410X475137.

Carter:1974:BLM

[CS74] Nancy F. Carter and Marc Swadener. A bibliography of literature: Mathematics education in the junior and community colleges. Two-Year College Mathematics Journal, 5(1):53–59, Winter 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1974.11974293.

Chu:1986:HFC

[CS86a] Sydney C. K. Chu and Man-Keung Siu. How far can you stick out your neck? College Mathematics Journal, 17(2):122–132, March 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1986.11972943.

Cunningham:1986:MSD

[CS86b] R. S. Cunningham and David A. Smith. A mathematics software database. College Mathematics Journal, 17(3):255–266, May 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1986.11972968.

Cunningham:1987:MSD

[CS87] R. S. Cunningham and David A. Smith. A mathematics software database update. College Mathematics Journal, 18(3):242–247, May 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1987.11973041.

Cunningham:1988:CMS

[CS88] R. S. Cunningham and David A. Smith. The compleat mathematics software database. College Mathematics Journal, 19(3):268– 289, May 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1988.11973127.

Curjel:1993:CLM

[CS93] Caspar R. Curjel and John Sylvester. A computer lab for multivariable calculus. College Mathematics Journal, 24(2):175–177, March 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1993.11973525.

Clark:1995:ITP

[CS95] Dean Clark and E. R. Suryanarayan. Inductive tiling of the plane by Penrose aperiodic rhombi. College Mathematics Journal, 26 (4):266–267, September 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1995.11973711.

Czyz:2003:RCE

[CS03] Jerzy Czyz and William Self. The rationals are countable — Euclid’s proof. College Mathematics Journal, 34(5):367–369, November 2003. CODEN ???? ISSN 0746-8342 (print), 1931-

1346 (electronic). URL http://www.tandfonline.com/doi/
abs/10.1080/07468342.2003.11922032.

Chan:2006:MDD

[CS06] O-Yeat Chan and James Smoak. More designer decimals: The integers and their geometric extensions. College Mathematics Journal, 37(5):355–363, November 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2006.
11922205.

Carter:2017:RIT

[CS17] Paul Carter and Yitzchak Elchanan Solomon. Relaxing the integral test: a challenge for the advanced calculus student. College Mathematics Journal, 48(4):290–291, September 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.48.4.290.

Clar:1974:R

[CSL^{+}74] Lawrence Clar, Gerald M. Smith, Peter A. Lindstrom, Arthur P. Dull, Olene C. Zacher, and Philip F. Rcichmeider. Reviews. Two-Year College Mathematics Journal, 5(2):55–59, Spring 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1974.11974315.

Cruz-Sampedro:2010:PNA

[CSTM10] Jaime Cruz-Sampedro and Margarita Tetlalmatzi-Montiel. POEM’s and Newton’s aerodynamic frustum. College Mathematics Journal, 41(2):145–153, March 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.

tandfonline.com/doi/abs/10.4169/074683410X480258.

Clevenson:2001:ASH

[CSWW01] Larry Clevenson, Mark F. Schilling, Ann E. Watkins, and William Watkins. The average speed on the highway. College Mathematics Journal, 32(3):169–171, May 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/07468342.
2001.11921871.

Cullen:1990:MFC

[Cul90] M. R. Cullen. Moiré fringes and the conic sections. College Mathematics Journal, 21(5):370–378, November 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1990.11973335.

Cullen:1997:CCC

[Cul97] Michael R. Cullen. Cylinder and cone cutting. College Mathematics Journal, 28(2):122–123, March 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1997.
11973847.

Cunningham:2008:RAC

[Cun08] Clifton Cunningham. Report from the Ambassador to Cida-2. College Mathematics Journal, 39(5):337–345, November 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2008.11922314.

Cuoco:2000:MPM

[Cuo00] Al Cuoco. Meta-problems in mathematics. College Mathematics Journal, 31(5):373–378, November 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2000.
11974176.

Cupillari:2000:M

[Cup00] Antonella Cupillari. Miscellanea. College Mathematics Journal, 31(2):144–147, March 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2000.11974133.

Curtin:2005:JHQ

[Cur05a] Daniel J. Curtin. Jan hudde and the quotient rule before Newton and Leibniz. College Mathematics Journal, 36(4):262–272, September 2005. CODEN ???? ISSN 0746-8342 (print), 1931-

1346 (electronic). URL http://www.tandfonline.com/doi/
abs/10.1080/07468342.2005.11922138.

Curtis:2005:TWF

[Cur05b] Dan Curtis. Taking a whipper-the fall-factor concept in rock climbing. College Mathematics Journal, 36(2):135–140, March 2005. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2005.11922120.

Curgus:2006:EEF

[Ćur06] Branko Ćurgus. An exceptional exponential function. College Mathematics Journal, 37(5):344–354, November 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2006.11922204.

Curtis:2010:WMD

[Cur10] Dan Curtis. What’s my domain? College Mathematics Journal, 41(2):113–121, March 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/074683410X480212.

Curgus:2014:STC

[Ćur14] Branko Ćurgus. A squeeze for two common sequences that converge to e. College Mathematics Journal, 45(5):391–392, November 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.45.5.391.

Cusick:1983:CCL

[Cus83] David Cusick. Computers & calculators: a logarithm algorithm for four-function calculators. Two-Year College Mathematics Journal, 14(4):322–324, September 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1983.
11972706;   https://www.jstor.org/stable/3027283.

Crisman:2014:RH

[CV14] Karl-Dieter Crisman and Mike Veatch. Reinventing Heron. College Mathematics Journal, 45(3):191–197, May 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
45.3.191.

Cass:1990:RBS

[CW90] Daniel Cass and Gerald Wildenberg. Relations between surface area and volume in lakes. College Mathematics Journal, 21(5):384–389, November 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1990.11973337.

Cox:2009:MCS

[CW09] Barry Cox and Stan Wagon. Mechanical circle-squaring. College Mathematics Journal, 40(4):238–247, September 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2009.11922370.

Cordier:2023:SPL

[CW23] Michelle Cordier and Meaghan Wheeler. Searching for point locations using lines. College Mathematics Journal, 54(5):423– 433, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2023.2263074.

deAlwis:1995:PMA

[dA95] Tilak de Alwis. Projectile motion with arbitrary resistance. College Mathematics Journal, 26(5):361–367, November 1995.

CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1995.11973730.

DAlessio:2023:SVP

[D’A23] Serge D’Alessio. Solutions of the Van der Pol equation. College Mathematics Journal, 54(2):90–98, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2023.
2191376.

Dalthorp:2019:WFM

[Dal19] Mark Dalthorp. When fractions make cycles. College Mathematics Journal, 50(1):3–8, January 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2019.
1534491.

Dantzig:1985:SMR

[Dan85] George B. Dantzig. Stephen Maurer’s response. College Mathematics Journal, 16(1):17–21, January 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1985.
11972846.

Danby:1992:CCC

[Dan92] J. M. A. Danby. Classroom computer capsules: Fireworks. College Mathematics Journal, 23(3):237–240, May 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/07468342.
1992.11973464.

Daniel:1998:HMM

[Dan98] James W. Daniel. How much money do you (or your parents) need for retirement? College Mathematics Journal, 29(4):278– 283, September 1998. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1998.11973955.

Dashti:2011:FT

[Das11] Mohammad Torabi Dashti. Faulhaber’s triangle. College Mathematics Journal, 42(2):96–97, March 2011. CODEN ????

ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
42.2.096.

Davis:1980:MF

[Dav80]  Philip J. Davis. Mathematics by fiat?  Two-Year Col-

lege Mathematics Journal, 11(4):255–263, September 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1980.11972452.

Davis:1982:SNCa

[Dav82a] Ronald M. Davis. The signpost: News & commentary. Two-Year College Mathematics Journal, 13(3):221–222, June 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1982.11972618.

Davis:1982:SNCb

[Dav82b] Ronald M. Davis. The signpost: News and commentary. Two-Year College Mathematics Journal, 13(4):283–284, September 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1982.11972632.

Davis:1982:SNCc

[Dav82c] Ronald M. Davis. The signpost: News commentary. Two-Year College Mathematics Journal, 13(5):343–346, November 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1982.11972644.

Davis:1982:TMR

[Dav82d] Tom Davis. Teaching mathematics with Rubik’s Cube. Two-Year College Mathematics Journal, 13(3):178–185, June 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1982.11972603.

Davis:1983:T

[Dav83] Philip J. Davis. The thread. Two-Year College Mathematics Journal, 14(2):98–104, March 1983. CODEN ????

ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1983.
11972664.

David:1984:OIP

[Dav84a] Karl David. An odd induction proof. College Mathematics Journal, 15(3):251, June 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1984.11972779.

Davis:1984:WYD

[Dav84b] Robert B. Davis. It’s not what you do, it’s how you do it …. College Mathematics Journal, 15(5):391–393, November 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972815.

Davis:1985:WDK

[Dav85] Philip J. Davis. What do I know? A study of mathematical self-awareness. College Mathematics Journal, 16(1):22–41, January 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1985.11972847.

Davis:1986:MH

[Dav86] Philip J. Davis. Media highlights. College Mathematics Journal, 17(1):101–109, January 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1986.11972940.

David:1988:RR

[Dav88] Karl David. Rencontres reencountered. College Mathematics Journal, 19(2):139–148, March 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1988.
11973102.

Davis:1992:SFS

[Dav92] Stephen L. Davis. Square-free sets on square grids. College Mathematics Journal, 23(3):214–224, May 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1992.
11973462.

Davis:1994:AGQ

[Dav94] Paul Davis. Asking good questions about differential equations. College Mathematics Journal, 25(5):394–400, November 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1994.11973643.

Dawson:1976:SRL

[Daw76] John W. Dawson, Jr. Some ridge-length problems. Two-Year College Mathematics Journal, 7(4):43–45, December 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1976.11974464.

Dawson:1980:WBD

[Daw80] John W. Dawson, Jr. Wavefronts, Box diagrams, and the product rule: a discovery approach. Two-Year College Mathematics Journal, 11(2):102–106, March 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1980.11972413.

Dawson:1990:HBF

[Daw90] John W. Dawson, Jr. Hanging a bird feeder: Food for thought. College Mathematics Journal, 21(2):129–130, March 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1990.11973296.

Dawson:1994:TPM

[Daw94] Robert J. MacG. Dawson. Towers of powers modulo m. College Mathematics Journal, 25(1):22–28, January 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1994.
11973575.

Dawson:1997:BDF

[Daw97] Bryan Dawson. A better draft: Fair division of the talent pool. College Mathematics Journal, 28(2):82–88, March 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1997.11973841.

Dawson:2019:CLU

[Daw19] C. Bryan Dawson. Calculus limits unified and simplified. College Mathematics Journal, 50(5):331–342, 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2019.
1662706.

Day:2013:PRP

[Day13] Colin Day. A power rule proof without limits. College Mathematics Journal, 44(4):323–324, September 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
44.4.323.

Dennis:1995:FCL

[DC95] David Dennis and Jere Confrey. Functions of a curve: Leibniz’s original notion of functions and its meaning for the parabola. College Mathematics Journal, 26(2):124–131, March 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1995.11973681.

Dence:1995:SRE

[DD95] Joseph B. Dence and Thomas P. Dence. A surprise regarding the equation ϕ(x) = 2(6n + 1). College Mathematics Journal, 26(4):297–301, September 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1995.11973716.

Dence:2010:FWS

[DD10] Joseph B. Dence and Thomas P. Dence. Four ways to skin a definite integral. College Mathematics Journal, 41(2):134–144, March 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/074683410X480249.

DeLeon:2010:ECU

[De 10] Doreen De Leon. Euler–Cauchy using undetermined coefficients. College Mathematics Journal, 41(3):235–237, May 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/074683410X488728.

Duncan:1977:FEG

[DE77] Hilda Duncan and David Emery. Finite Euclidean geometries of order p. Two-Year College Mathematics Journal, 8(1):4–10, January 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1977.11974470.

Dean:1971:VPT

[Dea71] Richard A. Dean. Vectors point toward Pisa. Two-Year College Mathematics Journal, 2(2):28–39, Autumn 1971. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1971.
11974009.

Deever:1999:ADE

[Dee99] David L. Deever. The average distance of the Earth from the Sun. College Mathematics Journal, 30(3):218–220, May 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1999.11974062.

DeMaio:2012:CTS

[DeM12] Joe DeMaio. Counting triangles to sum squares. College Mathematics Journal, 43(4):297–303, September 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
43.4.297.

DeMaio:2013:PWA

[DeM13] Joe DeMaio. Proof without words: an alternating sum of squares. College Mathematics Journal, 44(3):170, May 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.44.3.170.

Dence:1977:CC

[Den77] Thomas P. Dence. Computer corner. Two-Year College Mathematics Journal, 8(4):244–246, September 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1977.
11974519.

Dence:2007:SHR

[Den07] Thomas P. Dence. Some half-row sums from Pascal’s Triangle via Laplace transforms. College Mathematics Journal, 38(3):205– 209, May 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2007.11922237.

Denny:2012:REC

[Den12] J. K. Denny. Rational exponentials and continued fractions. College Mathematics Journal, 43(5):405–407, November 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.43.5.405.

Denny:2013:SOS

[Den13] J. K. Denny. SAGE: Open source mathematics software system (http://sagemath.org). College Mathematics Journal, 44(2): 149–155, March 2013. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.44.2.149.

DeTemple:1984:GPG

[DeT84] Duane W. DeTemple. Geometric procedures for graphing the general quadratic equation. College Mathematics Journal, 15 (4):313–323, September 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1984.11972790.

DeTemple:1988:GPC

[DeT88] Duane W. DeTemple. The generalized polygonal cycloid. College Mathematics Journal, 19(5):417–419, November 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1988.11973148.

DeTurck:1993:SR

[DeT93] Dennis DeTurck. Software reviews. College Mathematics Journal, 24(4):370–376, September 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1993.11973558.

DeTemple:1994:APB

[DeT94] Duane W. DeTemple. An Archimedean property of the bicylinder. College Mathematics Journal, 25(4):312–314, September 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1994.11973627.

DeTemple:1998:CPT

[DeT98] Duane W. DeTemple. Colored polygon triangulations. College Mathematics Journal, 29(1):43–47, January 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1998.
11973915.

DeTemple:2010:CSS

[DeT10] Duane W. DeTemple. Counting squares to sum squares. College Mathematics Journal, 41(3):214–221, May 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://

www.tandfonline.com/doi/abs/10.4169/074683410X488700.

Devaney:1989:MH

[Dev89] Robert L. Devaney. Media highlights. College Mathematics Journal, 20(5):452–459, November 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1989.
11973275.

Devaney:1991:ODM

[Dev91a] Robert L. Devaney. The orbit diagram and the Mandelbrot set. College Mathematics Journal, 22(1):23–38, January 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1991.11973355.

Devaney:1991:SID

[Dev91b] Robert L. Devaney. Special issue on dynamical systems. College Mathematics Journal, 22(1):2, January 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1991.
11973352.

Devlin:1995:MH

[Dev95] Keith Devlin. Media highlights. College Mathematics Journal, 26(2):163–170, March 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1995.11973691.

Devlin:2005:BR

[Dev05] Keith Devlin. Book review. College Mathematics Journal, 36(3): 255–256, May 2005. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2005.11922136.

Deddens:1985:SR

[DF85] James A. Deddens and Marjorie A. Fitting. Software reviews. College Mathematics Journal, 16(2):151–152, March 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1985.11972873.

diFranco:1989:SR

[dF89] Roland di Franco. Software review. College Mathematics Journal, 20(1):67, January 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1989.11973210.

Driscoll:1996:PKT

[DF96] Patrick J. Driscoll and William P. Fox. Presenting the Kuhn– Tucker conditions using a geometric approach. College Mathematics Journal, 27(2):101–108, March 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1996.
11973759.

diFranco:1990:SR

[dFMN^{+}90] Roland di Franco, T. J. Mahar, Tyre A. Newton, Tyre A. Newton, and Leonard G. Henscheid. Software reviews. College Mathematics Journal, 21(3):242–245, May 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/07468342.
1990.11973315.

Donaghey:1980:IAG

[DG80] Robert Donaghey and Warren Gordon. An investment approach to geometric series. Two-Year College Mathematics Journal, 11(2):120–121, March 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1980.11972419.

Dorff:2003:SWA

[DH03] Michael Dorff and Leon Hall. Solids in R^{n} whose area is the derivative of the volume. College Mathematics Journal, 34(5): 350–358, November 2003. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2003.11922029.

Derrick:2012:PWP

[DH12a] William Derrick and James Hirstein. Proof without words: Ptolemy’s Theorem. College Mathematics Journal, 43(5):386, November 2012. CODEN ???? ISSN 0746-8342 (print), 1931-

1346 (electronic). URL http://www.tandfonline.com/doi/
abs/10.4169/college.math.j.43.5.386.

Deveau:2012:QRI

[DH12b] Michael Deveau and Robie Hennigar. Quotient-rule-integration-by-parts. College Mathematics Journal, 43(3):254–256, May 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.43.3.254.

Denman:2010:TGM

[DHR10] Richard T. Denman, David Hailey, and Michael Rothenberg. The tower and glass marbles problem. College Mathematics Journal, 41(5):350–356, November 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/074683410X521946.

Dziabenko:2008:CHT

[DI08] Dimitri Dziabenko and Oleg Ivrii. Centaurs: Here, there, everywhere! College Mathematics Journal, 39(4):267–272, September 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2008.11922303.

Diamond:2017:RAS

[Dia17] Harvey Diamond. The rational approximation of small angles. College Mathematics Journal, 49(1):57–59, January 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2017.1397994.

Dickey:2006:DDK

[Dic06] Leonid A. Dickey. Do dogs know calculus of variations? College Mathematics Journal, 37(1):20–23, January 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2006.
11922162.

Dick:2017:BR

[Dic17] Thomas Dick. Book review. College Mathematics Journal, 48 (2):146–152, March 2017. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.48.2.146.

Dieudonne:1981:UDG

[Die81] J. Dieudonné. The universal domination of geometry. Two-Year College Mathematics Journal, 12(4):227–231, September 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972536.

Dierckx:2023:EPP

[Die23] Stijn Dierckx. An extension of the perplexing polynomial puzzle. College Mathematics Journal, 54(4):299–307, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2023.2225391.

Dillon:2014:PGA

[Dil14] Meighan Dillon. Projective geometry for all. College Mathematics Journal, 45(3):169–178, May 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
45.3.169.

Diniz:2013:APF

[Din13] Marcio Diniz. Asset pricing, financial markets, and linear algebra. College Mathematics Journal, 44(1):2–8, January 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.44.1.002.

Ding:2022:LMA

[Din22] Pisheng Ding. Less mundane applications of the most mundane functions. College Mathematics Journal, 53(3):230–232, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2022.2040266.

Dasef:1997:SSS

[DK97] Martha E. Dasef and Steven M. Kautz. Some sums of some significance. College Mathematics Journal, 28(1):52–55, January 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1997.11973831.

Day:2001:TLA

[DK01] Jane M. Day and Dan Kalman. Teaching linear algebra: Issues and resources. College Mathematics Journal, 32(3):162–168, May 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2001.11921870.

Douthett:2008:DTC

[DK08] Jack Douthett and Richard J. Krantz. Dinner tables and concentric circles: a harmony of mathematics, music, and physics. College Mathematics Journal, 39(3):203–211, May 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2008.11922294.

Duncan:1973:CSC

[DL73] David R. Duncan and Bonnie H. Litwiller. Checkerboards and sugar cubes: Geometric counting patterns. Two-Year College Mathematics Journal, 4(2):41–47, Spring 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1973.
11974250.

Duncan:1975:ILR

[DL75] David R. Duncan and Bonnie H. Litwiller. Investigations of linear and reciprocal functions by the line-to-line technique. Two-Year College Mathematics Journal, 6(2):2–9, May 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1975.11974369.

Detmer:1984:SR

[DL84] Richard C. Detmer and L. Carl Leinbach. Software reviews. College Mathematics Journal, 15(2):159–162, March 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972769.

Duncan:2004:VPN

[DLM04] J. Duncan, Daniel H. Luecking, and C. M. McGregor. On the values of pi for norms on R^{2}. College Mathematics Journal, 35 (2):84–92, March 2004. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2004.11922056.

Diniz:2016:SCP

[DLPS16] M. A. Diniz, D. Lopes, A. Polpo, and L. E. B. Salasar. The sticker collector’s problem. College Mathematics Journal, 47 (4):255–263, September 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.47.4.255.

Dray:2003:UDB

[DM03] Tevian Dray and Corinne A. Manogue. Using differentials to bridge the vector calculus gap. College Mathematics Journal, 34(4):283–290, September 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2003.11922019.

Dray:2010:PDB

[DM10] Tevian Dray and Corinne A. Manogue. Putting differentials back into calculus. College Mathematics Journal, 41(2):90–100, March 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/074683410X480195.

Davis:1970:JCC

[DN70] James C. Davis and Ralph H. Niemann. Junior College cooperative program in Colorado. Two-Year College Mathematics Journal, 1(1):41–43, Spring 1970. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1970.11973167.

Doane:1989:UMS

[Doa89] David P. Doane. Using median splits to motivate learning. College Mathematics Journal, 20(3):228–229, May 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1989.11973234.

Dobyns:1972:CEA

[Dob72] Roy Dobyns. A condition equivalent to associativity for finite groups. Two-Year College Mathematics Journal, 3(1):10–13, Spring 1972. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1972.11974189.

Dobbs:1984:PHF

[Dob84] David E. Dobbs. Proving Heron’s formula tangentially. College Mathematics Journal, 15(3):252–253, June 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1984.
11972780.

Dobbs:1985:AVL

[Dob85] David E. Dobbs. Average values and linear functions. College Mathematics Journal, 16(2):132–135, March 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1985.
11972868.

Dominici:2011:AM

[Dom11] Diego Dominici. An arithmetic metric. College Mathematics Journal, 42(3):207–214, May 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
42.3.207.

Dorn:1977:CC

[Dor77a] Willjam S. Dorn. Computer corner. Two-Year College Mathematics Journal, 8(2):101–105, March 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1977.
11974491.

Dorwart:1977:CPF

[Dor77b] Harold L. Dorwart. Can this polynomial be factored? Two-Year College Mathematics Journal, 8(2):67–72, March 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1977.11974484.

Dorner:2005:DRA

[Dor05] Bryan Dorner. Discovering roots: Ancient, medieval, and serendipitous. College Mathematics Journal, 36(1):35–43, January 2005. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2005.11922108.

Dossey:1996:MH

[Dos96] John A. Dossey. Media highlights. College Mathematics Journal, 27(4):317–324, September 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1996.11973802.

Douthitt:1973:ELA

[Dou73] Cameron Douthitt. The effects of a laboratory on achievement in College freshman mathematics. Two-Year College Mathematics Journal, 4(1):55–59, Winter 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1973.11974237.

Douglas:1985:ONM

[Dou85] R. G. Douglas. One needs more than the algorithmic approach. College Mathematics Journal, 16(1):5–6, January 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1985.11972839.

Davis:1971:TYC

[DP71] John B. Davis, Jr. and T. J. Pignani. Two-year College faculty participation in professional mathematics. Two-Year College Mathematics Journal, 2(1):53–57, Spring 1971. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/00494925.
1971.11973999.

Dunkels:1980:CTM

[DP80] Andrejs Dunkels and Lars-Erik Persson. Creative teaching by mistakes. Two-Year College Mathematics Journal, 11(5):296– 300, November 1980. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1980.11972465.

Davis:2021:BRA

[DPV21] Matt Davis, Adam E. Parker, and Daniel A. N. Vargas. Being rational about algebraic numbers. College Mathematics Journal, 52(5):327–337, 2021. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2021.1967651.

Dawson:2004:HR

[DR04] C. Bryan Dawson and Troy D. Riggs. Highway relativity. College Mathematics Journal, 35(4):246–250, September 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2004.11922080.

DeTemple:2006:MCP

[DR06] Duane De Temple and H. David Reynolds II. More combinatorial proofs via flagpole arrangements. College Mathematics Journal, 37(4):279–285, September 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2006.11922195.

Ding:2013:TTW

[DR13] J. Ding and N. H. Rhee. Teaching tip: When a matrix and its inverse are stochastic. College Mathematics Journal, 44(2): 108–109, March 2013. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.44.2.108.

Durham:2020:CSR

[DR20] Samuel Durham and Tom Richmond. Connected subsets of an n × 2 rectangle. College Mathematics Journal, 51(1):32– 42, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2020.1674597.

Dr:2024:PWP

[Dr24a] Günhan Caglayan Dr. Proof without words: a pattern block dissection to square the regular dodecagon. College Mathematics Journal, 55(4):344–345, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2024.2302793.

Dr:2024:TOD

[Dr24b] Ricardo E. Rojas Dr. A two-for-one deal on rational triangles. College Mathematics Journal, 55(3):250–252, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2023.2282345.

Draayer:2000:SI

[Dra00] Bret Draayer. N-site insights. College Mathematics Journal, 31(4):250–258, September 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2000.11974154.

Dray:2013:UDD

[Dra13] Tevian Dray. Using differentials to differentiate trigonometric and exponential functions. College Mathematics Journal, 44(1): 17–23, January 2013. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.44.1.017.

Dray:2023:HL

[Dra23] Tevian Dray. Hyperbolic lunes. College Mathematics Journal, 54(4):308–315, 2023. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2023.2231314.

Derderian:1998:USW

[DRC98] Jean-Claude Derderian and Enriqueta Rodríguez-Carrington. Undersampled sine waves. College Mathematics Journal, 29(3): 213–218, May 1998. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1998.11973942.

Drucker:1992:MRD

[Dru92] Daniel Drucker. A mathematical roller derby. College Mathematics Journal, 23(5):396–401, November 1992. CODEN ????

ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1992.
11973489.

Drucker:1995:MTD

[Dru95] Jack Drucker. Minimal time of descent. College Mathematics Journal, 26(3):232–235, May 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1995.
11973705.

Deeba:1996:CPP

[DRW96] Elias Y. Deeba, Dennis M. Rodriguez, and Ibrahim Wazir. Critical points of polynomial families. College Mathematics Journal, 27(4):291–295, September 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1996.11973797.

Dubbs:1987:CD

[DS87] Clyde Dubbs and David Siegel. Computing determinants. College Mathematics Journal, 18(1):48–50, January 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1987.
11973007.

Dou:1999:MAC

[DS99] Ze-Li Dou and Susan G. Staples. Maximizing the arclength in the cannonball problem. College Mathematics Journal, 30(1):44–45, January 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1999.11974029.

Dickinson:2008:RRT

[DS08] William Dickinson and Mohammad Salmassi. The right Right Triangle on the Sphere. College Mathematics Journal, 39(1): 24–33, January 2008. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2008.11922271.

Dodge:2012:BHS

[DS12] John Dodge and Andrew Simoson. Ben-hur staircase climbs. College Mathematics Journal, 43(4):274–284, September 2012.

CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.43.4.274.

DeMaio:2013:DIT

[DT13] Joe DeMaio and Hong Lien Tran. Domination and independence on a triangular honeycomb chessboard. College Mathematics Journal, 44(4):307–314, September 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
44.4.307.

Du:2020:FGS

[Du20] Kevin Y. Du. A fast-growing sequence inspired by TREE(k). College Mathematics Journal, 51(1):43–50, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2020.
1680227.

Dubeau:1992:WMO

[Dub92] François Dubeau. Weighted means of order r and related inequalities: an elementary approach. College Mathematics Journal, 23(3):211–213, May 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1992.11973461.

Dudley:1994:MH

[Dud94] Underwood Dudley. Media highlights. College Mathematics Journal, 25(3):247–254, May 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1994.11973620.

Dudley:1997:MN

[Dud97] Underwood Dudley. Is mathematics necessary? College Mathematics Journal, 28(5):360–364, November 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1997.
11973890.

Duemmel:1993:C

[Due93] James E. Duemmel. The curious 1/3. College Mathematics Journal, 24(3):236–237, May 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1993.11973535.

Duemmel:1996:CER

[Due96] James Duemmel. Complex eigenvalues and rotations: Are your students going in circles? College Mathematics Journal, 27 (5):378–381, November 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1996.11973812.

Dundas:1977:QPT

[Dun77] Kay Dundas. Quasi-Pythagorean triples for an oblique triangle. Two-Year College Mathematics Journal, 8(3):152–155, June 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1977.11974500.

Dundas:1984:BBB

[Dun84] Kay Dundas. To build a better Box. College Mathematics Journal, 15(1):30–36, January 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1984.11972741.

Dunham:1987:BHS

[Dun87] William Dunham. The Bernoullis and the harmonic series. College Mathematics Journal, 18(1):18–23, January 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1987.11973001.

Dunham:1991:EFT

[Dun91] William Dunham. Euler and the fundamental theorem of algebra. College Mathematics Journal, 22(4):282–293, September 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1991.11973397.

Dunbar:1997:ADB

[Dun97a] Steven R. Dunbar. The average distance between points in geometric figures. College Mathematics Journal, 28(3):187–197, May 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1997.11973860.

Dunham:1997:MH

[Dun97b] William Dunham. Media highlights. College Mathematics Journal, 28(3):237–246, May 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1997.11973871.

Dunham:2020:ME

[Dun20] William Dunham. A morsel from Euler. College Mathematics Journal, 51(1):3–8, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2020.1674577.

Dunham:2023:BMC

[Dun23] William Dunham. Bryn Mawr College matriculation exams from a century ago. College Mathematics Journal, 54(2):83– 89, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2023.2189820.

Dyson:1974:PMI

[DW74] Louise Dyson and Edward B. Wright. Pills: Mathematics instructional model. Two-Year College Mathematics Journal, 5 (3):31–33, September 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1974.11974333.

Demana:1987:PSU

[DW87] Franklin Demana and Bert Waits. Problem solving using microcomputers. College Mathematics Journal, 18(3):236–241, May 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1987.11973040.

Demana:1988:PGC

[DW88] Franklin Demana and Bert K. Waits. Pitfalls in graphical computation, or why a single graph isn’t enough. College Mathematics Journal, 19(2):177–183, March 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1988.
11973112.

Deng:1998:MC

[DW98] Min Deng and Mary T. Whalen. The mathematics of cootie. College Mathematics Journal, 29(3):222–224, May 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1998.11973944.

Ding:2009:DEF

[DW09] Jiu Ding and Zizhong Wang. Dynamics of exponential functions. College Mathematics Journal, 40(5):361–369, November 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/074683409X475689.

Dlab:2019:MSP

[DW19] Vlastimil Dlab and Kenneth S. Williams. The many sides of the Pythagorean theorem. College Mathematics Journal, 50(3):162– 172, 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2019.1580527.

Denny:2004:TMW

[DY04] J. K. Denny and C. A. Yackel. Temperature models for Ware Hall. College Mathematics Journal, 35(3):162–170, May 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2004.11922069.

Edwards:1999:AFS

[EB99] Thomas G. Edwards and Lawrence Brenton. An attempt to Foster students’ construction of knowledge during a semester course in abstract algebra. College Mathematics Journal, 30(2):120–128, March 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1999.11974043.

Editor:2024:MHa

[EBS24a] Lowell Beineke Section Editor, Jennifer Beineke, and Philip Straffin. Media highlights. College Mathematics Journal, 55(1): 66–73, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2024.2285198.

Editor:2024:MHb

[EBS24b] Lowell Beineke Section Editor, Jennifer Beineke, and Philip Straffin. Media highlights. College Mathematics Journal, 55 (2):180–187, 2024. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2024.2299196.

Editor:2024:MHc

[EBS24c] Lowell Beineke Section Editor, Jennifer Beineke, and Philip Straffin. Media highlights. College Mathematics Journal, 55 (4):363–370, 2024. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2024.2374218.

Ecker:1981:MDN

[Eck81] Michael W. Ecker. Must a “dud” necessarily be an inflection point? Two-Year College Mathematics Journal, 12(5):332– 333, November 1981. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1981.11972564.

Ecker:1983:ALC

[Eck83] Michael W. Ecker. The alluring lore of cyclic numbers. Two-Year College Mathematics Journal, 14(2):105–109, March 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1983.11972665.

Eckert:1992:PPT

[Eck92] Ernest J. Eckert. Primitive Pythagorean triples. College Mathematics Journal, 23(5):413–417, November 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1992.
11973493.

Ecker:1997:DHS

[Eck97] Michael W. Ecker. Divergence of the harmonic series by rearrangement. College Mathematics Journal, 28(3):209–210, May 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1997.11973866.

Ecker:1998:NAG

[Eck98] Michael W. Ecker. A novel approach to geometric series. College Mathematics Journal, 29(5):419–420, November 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1998.11973982.

Ecker:2007:MPB

[Eck07] Michael W. Ecker. Maximizing the probability of a big sweepstakes win. College Mathematics Journal, 38(1):32–36, January 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2007.11922215.

Ecker:2011:LIL

[Eck11] Michael W. Ecker. Limit interchange and L’Hôpital’s Rule. College Mathematics Journal, 42(5):382–383, November 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.42.5.382.

Lakins:2024:LEb

[(Ed24] Tamara Lakins (Editor). Letter from the Editor. College Mathematics Journal, 55(2):85, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2024.2303271.

Eddy:1985:BAG

[Edd85] Roland H. Eddy. Behold! The Arithmetic–Geometric Mean Inequality. College Mathematics Journal, 16(3):208, June 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1985.11972881.

Edgar:2016:PWM

[Edg16] Tom Edgar. Proof without words: Matchstick triangles. College Mathematics Journal, 47(3):207, May 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
47.3.207.

Edgar:2018:RNB

[Edg18] Tom Edgar. Roger Nelsen’s books, so far. College Mathematics Journal, 49(4):302–312, September 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2018.
1498263.

Edgar:2020:VDP

[Edg20] Tom Edgar. Visual decompositions of polygonal numbers. College Mathematics Journal, 51(1):9–12, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2020.
1680238.

Edgar:2021:VTN

[Edg21] Tom Edgar. Visual triangular number identities from positional number systems. College Mathematics Journal, 52(2):133– 136, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2021.1877488.

Edgar:2022:VSI

[Edg22] Tom Edgar. Visual sums of integers in polygonal arrays. College Mathematics Journal, 53(1):45–51, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2022.
1996186.

Edgar:2023:BSR

[Edg23] Tom Edgar. Balancing the square root of two. College Mathematics Journal, 54(3):176–178, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2023.2201569.

Edwards:2014:PWI

[Edw14] Steven Edwards. Proof without words: an infinite series using Golden triangles. College Mathematics Journal, 45(2):120, March 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.45.2.120.

Eenigenburg:1997:NRA

[Een97] Paul Eenigenburg. A note on the ratio of arc length to chordal length. College Mathematics Journal, 28(5):391–393, November 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1997.11973899.

Eddy:1994:OO

[EF94] R. H. Eddy and R. Fritsch. An optimization oddity. College Mathematics Journal, 25(3):227–229, May 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1994.
11973613.

Engdahl:1975:RCS

[EG75] Roland Engdahl and Karl Greger. Random charity. a stochastic sieving problem and its connection with the Euclidean algorithm. Two-Year College Mathematics Journal, 6(4):4–9, December 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1975.11974398.

Eggers:2020:CLC

[Egg20] John Eggers. A closer look at the compensating polar planimeter. College Mathematics Journal, 51(2):105–116, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2020.1702852.

Engbers:2016:DPI

[EH16] John Engbers and Adam Hammett. A direct proof of the integral formulae for the inverse hyperbolic functions. College Mathematics Journal, 47(4):297–299, September 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
47.4.297.

Engbers:2019:CP

[EH19] John Engbers and Adam Hammett. The n-children problem. College Mathematics Journal, 50(4):242–249, 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/full/10.1080/07468342.
2019.1632663.

Engbers:2020:RRQ

[EH20] John Engbers and Adam Hammett. To replace or not to replace — that is the question. College Mathematics Journal, 51(2):117–123, 2020. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2020.1698932.

Ennis:2024:DPP

[EH24] Christopher Ennis and Inge Helland. Disjoint placement probability of line segments via induction. College Mathematics Journal, 55(1):32–39, 2024. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2023.2255087.

Ehrismann:1994:QTD

[Ehr94] Martin Ehrismann. Quenching a thirst with differential equations. College Mathematics Journal, 25(5):413–418, November 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1994.11973645.

Ehrenborg:2021:FGH

[Ehr21] Richard Ehrenborg. Fano, Galois, Hamming and a card trick. College Mathematics Journal, 52(4):274–280, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2021.1942678.

Ehrenborg:2024:PWG

[Ehr24] Richard Ehrenborg. Proofs without words: Game of parts. College Mathematics Journal, 55(5):428, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic).

Ervynck:1989:PSE

[EI89] G. Ervynck and P. Igodt. Power series and exponential generating functions. College Mathematics Journal, 20(5):411–415, November 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1989.11973267.

Eidswick:1993:TTP

[Eid93] Jack Eidswick. Two trisectrices for the price of one rolling coin. College Mathematics Journal, 24(5):422–430, November 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1993.11973563.

Eisenberg:1997:WME

[Eis97] Bennett Eisenberg. What is the margin of error of a poll? College Mathematics Journal, 28(3):201–203, May 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1997.11973862.

Eisenberg:1998:HMS

[Eis98a] Bennett Eisenberg. How much should you pay for a derivative? College Mathematics Journal, 29(5):412–414, November 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1998.11973978.

Eisner:1998:PPM

[Eis98b] Milton P. Eisner. The probability of passing a multiple-choice test. College Mathematics Journal, 29(5):421–426, November 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1998.11973983.

Eisenberg:2017:VSP

[Eis17] Bennett Eisenberg. A very short proof that the sum of independent normal random variables is normal. College Mathematics Journal, 48(2):137, March 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.48.2.137.

Eisenberg:1992:OL

[EK92] Bennett Eisenberg and Samir Khabbaz. Optimal locations. College Mathematics Journal, 23(4):282–289, September 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1992.11973470.

Eggleton:2011:PQR

[EK11] Roger Eggleton and Vladimir Kustov. The product and quotient rules revisited. College Mathematics Journal, 42(4):323– 326, September 2011. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.42.4.323.

Edwards:2022:TIE

[EK22] Abe Edwards and Brielle Komosinski. Tetration: Iterative enjoyment. College Mathematics Journal, 53(3):209–219, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2022.2045145.

Eklund:2000:PFS

[Ekl00] Neil Eklund. t-probabilities as finite sums. College Mathematics Journal, 31(3):217–218, May 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2000.
11974148.

Eklund:2001:CEF

[Ekl01] Neil Eklund. CORDIC: Elementary function computation using recursive sequences. College Mathematics Journal, 32(5):330– 333, November 2001. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2001.11921899.

Early:2015:GP

[EKP15] Edward Early, Patrick Kim, and Michael Proulx. Goldbach’s pigeonhole. College Mathematics Journal, 46(2):131–135, March 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.46.2.131.

Elliott:1983:PGF

[Ell83] Donald D. Elliott. A prime-generating function. Two-Year College Mathematics Journal, 14(1):57, January 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1983.11972654.

Ellis:1984:TMT

[Ell84] Wade Ellis, Jr. Teaching is mathematics, too …. College Mathematics Journal, 15(5):393–394, November 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972816.

Elran:2012:RPN

[Elr12] Yossi Elran. Retrolife and the pawns neighbors. College Mathematics Journal, 43(2):147–151, March 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
43.2.147.

Elsner:1982:TT

[Els82] Thomas E. Elsner. Taxes on taxes. Two-Year College Mathematics Journal, 13(4):266–269, September 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1982.
11972625.

Edgar:2017:VVV

[EM17] Tom Edgar and N. Chris Meyer. A visual validation of Viète’s verification. College Mathematics Journal, 48(2):90–96, March 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.48.2.90.

Emmons:2009:TP

[Emm09] Caleb Emmons. Three poems. College Mathematics Journal, 40(3):188, May 2009. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2009.11922358.

Eidolon:2017:SPB

[EO17] Katrina Eidolon and Greg Oman. A short proof of the Bolzano– Weierstrass Theorem. College Mathematics Journal, 48(4):288– 289, September 2017. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.48.4.288.

Erickson:2022:HMW

[Eri22] William Q. Erickson. Haste makes waste: An optimization problem. College Mathematics Journal, 53(2):122–133, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2021.2022955.

Edwards:1985:ICC

[ES85] Bruce H. Edwards and Patricia L. Sharpe. III-Conditioning: a constant surprise in computational mathematics. College Mathematics Journal, 16(2):141–148, March 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1985.
11972871.

Editor:2023:MH

[ES23a] Lowell Beineke Section Editor and Philip Straffin. Media highlights. College Mathematics Journal, 54(5):500–507, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2023.2271357.

Ennis:2023:DPP

[ES23b] Christopher Ennis and John Shier. Disjoint placement probability of line segments via geometry. College Mathematics Journal, 54(1):44–53, 2023. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2022.2160619.

Essary:1982:RAP

[Ess82] Alice W. Essary. Rectangular aids for polar graphs. Two-Year College Mathematics Journal, 13(3):200–203, June 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1982.11972612.

Edwards:2008:DPS

[ESS08] William F. Edwards, Ray C. Shiflett, and Harris S. Shultz. Dependent probability spaces. College Mathematics Journal, 39(3): 221–226, May 2008. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2008.11922296.

Oman:2024:PSd

[EST24] Greg Oman (Editors), Mátyás A. Sustik, and Katherine Thompson. Problems and solutions. College Mathematics Journal, 55(4):352–362, 2024. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2024.2373015.

Ettl:1974:DMA

[Ett74] Joseph J. Ettl. The delta method approximates the roots of polynomial equations. Two-Year College Mathematics Journal, 5(2): 19–20, Spring 1974. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1974.11974309.

Euler:1986:ND

[Eul86] Russell Euler. A note on differentiation. College Mathematics Journal, 17(2):166–167, March 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1986.
11972948.

Euler:1993:MEL

[Eul93] Russell Euler. Maclaurin expansion of Arctanx via L’Hôspital’s Rule. College Mathematics Journal, 24(4):347–350, September 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1993.11973552.

Euler:1997:NTS

[Eul97] Russell Euler. A note on Taylor’s series for sin(ax + b) and cos(ax + b). College Mathematics Journal, 28(4):297–298, September 1997. CODEN ???? ISSN 0746-8342 (print), 1931-

1346 (electronic). URL http://www.tandfonline.com/doi/
abs/10.1080/07468342.1997.11973880.

Eves:1991:TST

[Eve91] Howard Eves. Two surprising theorems on Cavalieri congruence. College Mathematics Journal, 22(2):118–124, March 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1991.11973367.

Eves:2002:VBH

[Eve02] Howard W. Eves. A very brief history of statistics. College Mathematics Journal, 33(4):306–308, September 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2002.11921958.

Elbrink:1970:SAM

[EW70] Larry C. Elbrink and Bert K. Waits. A statistical analysis of multiple-choice examinations in mathematics. Two-Year College Mathematics Journal, 1(1):25–29, Spring 1970. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1970.11973162.

Ewing:1995:CWS

[Ewi95]  John Ewing. Can we see the Mandelbrot set?  College

Mathematics Journal, 26(2):90–99, March 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1995.
11973676.

Ewing:1999:BR

[Ewi99] John Ewing. Book review. College Mathematics Journal, 30(2): 159–161, March 1999. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1999.11974051.

Forbes:1971:WNT

[FA71] Jack E. Forbes and Josephine P. Andree. What’s new in TYC mathematics? Two-Year College Mathematics Journal, 2(2): 93–94, Autumn 1971. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1971.11974018.

Faase:1985:MH

[FA85] Thomas P. Faase and G. L. Alexanderson. Media highlights. College Mathematics Journal, 16(5):425–432, November 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1985.11972921.

Fagin:2020:IFC

[Fag20] Barry S. Fagin. Idempotent factorizations in the cryptography classroom. College Mathematics Journal, 51(3):195–203, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2020.1724495.

Fakler:1999:DMC

[Fak99] Robert Fakler. Do most cubic graphs have two turning points? College Mathematics Journal, 30(5):367–369, November 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1999.11974087.

Falk:1981:ALM

[Fal81] Ruma Falk. Another look at the mean, median, and standard deviation. Two-Year College Mathematics Journal, 12(3):207– 208, June 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972527.

Fala:1985:IAC

[Fal85] Gino T. Fala. Income averaging can increase your tax liability. College Mathematics Journal, 16(1):53–55, January 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1985.11972850.

Falk:1986:PKK

[Fal86] Ruma Falk. Of probabilistic knights and knaves. College Mathematics Journal, 17(2):156–164, March 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1986.
11972946.

Falbo:2005:GRC

[Fal05] Clement Falbo. The Golden ratio — a contrary viewpoint. College Mathematics Journal, 36(2):123–134, March 2005. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2005.11922119.

Farnsworth:2000:GS

[Far00] David L. Farnsworth. The geometry of statistics. College Mathematics Journal, 31(3):200–204, May 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2000.
11974143.

Farmer:2002:DCM

[Far02] Tom Farmer. Designing a calculus mobile. College Mathematics Journal, 33(2):131–136, March 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2002.
11921931.

Farmer:2008:HSP

[Far08] Tom Farmer. A helical stairway project. College Mathematics Journal, 39(4):291–298, September 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2008.
11922306.

Farrington:2015:PEC

[Far15] Eleanor Farrington. Parametric equations at the circus: Trochoids and poi flowers. College Mathematics Journal, 46(3):173– 177, May 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.46.3.173.

Fay:1997:HOP

[Fay97] Temple H. Fay. Harmonic oscillators with periodic forcing. College Mathematics Journal, 28(2):98–105, March 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1997.11973843.

Fraser:1978:LE

[FBF78] W. Roy Fraser, Dorothy Buerk, and Herta T. Freitag. Letters to the Editor. Two-Year College Mathematics Journal, 9(1):3–5, January 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1978.11974535.

Falk:1983:PDB

[FBH83] Ruma Falk and Maya Bar-Hillel. Probabilistic dependence between events. Two-Year College Mathematics Journal, 14(3):240– 247, June 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1983.11972689.

Fisher:1991:IST

[FC91] William Fisher and Richard Campbell. Investigating spirolaterals through LOGO. College Mathematics Journal, 22(2):148–159, March 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1991.11973375.

Fossa:2001:LBW

[FE01] John A. Fossa and Glenn W. Erickson. The Lord over better and worse births. College Mathematics Journal, 32(3):185–193, May 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2001.11921874.

Feeman:2001:CEF

[Fee01] Timothy G. Feeman. Conformality, the exponential function, and world map projections. College Mathematics Journal, 32 (5):334–342, November 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2001.11921900.

Feinstein:1976:DAD

[Fei76] Irwin K. Feinstein. The Doctor of Arts Degree in Mathematics University of Illinois at Chicago Circle. Two-Year College Mathematics Journal, 7(4):18–20, December 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1976.11974457.

Fekete:1988:APD

[Fek88] Antal E. Fekete. Apropos predetermined determinants. College Mathematics Journal, 19(3):254–257, May 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/07468342.
1988.11973124.

Fernandez:1998:CPP

[Fer98] Federico Fernández. A class of pleasing periodic designs. College Mathematics Journal, 29(1):18–26, January 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1998.
11973908.

Flanders:1987:ABMe

[FF87] Harley Flanders and J. Sutherland Frame. Algorithm of the bimonth. College Mathematics Journal, 18(5):417–421, November 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1987.11973068.

Ferdinands:2024:EBT

[FF24] John Ferdinands and Timothy Ferdinands. An extension of Babbage’s theorem. College Mathematics Journal, 55(2):162– 164, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2023.2231312.

Farmer:1992:PDC

[FG92] Tom Farmer and Fred Gass. Physical demonstrations in the calculus classroom. College Mathematics Journal, 23(2):146–148, March 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1992.11973447.

Francis:1993:FFF

[FG93] Richard L. Francis and Constance J. Gardner. Fallacies, flaws, and flimflam. College Mathematics Journal, 24(1):63–66, January 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1993.11973508.

Fink:2007:NPG

[FG07] Alex Fink and Richard Guy. The number-pad game. College Mathematics Journal, 38(4):260–264, September 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2007.11922246.

Fried:2010:PLI

[FG10] Michael N. Fried and Mayer Goldberg. A pumping lemma for invalid reductions of fractions. College Mathematics Journal, 41 (5):357–364, November 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/074683410X521955.

French:2019:BTA

[FG19] Robert M. French and Patrick Géhant. The barycenter theorem: Averaging possible-paths to produce optimal discrete straight-line segments. College Mathematics Journal, 50(2):103– 114, 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2019.1559680.

Frassetto:2018:SSD

[FGL^{+}18] Erin Frassetto, Michael Gableman, McKenzie Lamb, Tyler Shimek, and Andrea Young. Strange spinners and diversity of dice in chutes and ladders. College Mathematics Journal, 49 (4):251–260, September 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2018.1488500.

French:1994:DNP

[FH94] A. P. French and R. S. Hu. In defense of Newton: a physicist’s view. College Mathematics Journal, 25(3):206–211, May 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1994.11973609.

Frederick:1995:MJF

[FH95] William G. Frederick and James R. Hersberger. The mathematical judge: a fable. College Mathematics Journal, 26(5):377– 381, November 1995. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1995.11973733.

Fjelstad:2000:PRA

[FH00] Paul Fjelstad and Peter Hammer. A picture for real arithmetic. College Mathematics Journal, 31(1):56–60, January 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2000.11974111.

Fontaine:2011:GMF

[FH11] Anne Fontaine and Susan Hurley. Golden matrix families. College Mathematics Journal, 42(2):140–147, March 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.42.2.140.

Frey:2021:IFD

[FH21] Darrin Frey and Adam Hammett. An infinite family of divisibility tests. College Mathematics Journal, 52(1):2–10, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2021.1845530.

Fink:1994:SR

[Fin94] James P. Fink. Software reviews. College Mathematics Journal, 25(4):329–333, September 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1994.11973635.

Fink:1997:SR

[Fin97] James P. Fink. Software reviews. College Mathematics Journal, 28(1):63–66, January 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1997.11973837.

Finn:2002:CBC

[Fin02] David L. Finn. Can a bicycle create a unicycle track? College Mathematics Journal, 33(4):283–292, September 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2002.11921954.

Field:2014:TVH

[FIRW14] Elizabeth Field, Rachael Ivison, Amanda Reyher, and Steven Warner. Truck versus human: Mathematics under pressure. College Mathematics Journal, 45(2):116–120, March 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.45.2.116.

Fisk:1975:SAM

[Fis75] Robert S. Fisk. Some applications of modeling in mathematics for two-year colleges. Two-Year College Mathematics Journal, 6(4):10–13, December 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1975.11974399.

Fischbeck:1996:SR

[Fis96] Sally Fischbeck. Software reviews. College Mathematics Journal, 27(3):224–230, May 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1996.11973785.

Fjelstad:1995:RIP

[Fje95] Paul Fjelstad. The repeating integer paradox. College Mathematics Journal, 26(1):11–15, January 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1995.
11973659.

Ferzola:1994:ED

[FK94] Anthony P. Ferzola and Sydney H. Kung. Euler and differentials. College Mathematics Journal, 25(2):102–111, March 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1994.11973590.

Fengming:2013:FIA

[FKY13] Dong Fengming, Ho Weng Kin, and Lee Tuo Yeong. A family of identities via arbitrary polynomials. College Mathematics Journal, 44(1):43–47, January 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.44.1.043.

Fuller:1975:EDC

[FL75] L. E. Fuller and J. D. Logan. On the evaluation of determinants by Chiò’s Method. Two-Year College Mathematics Journal, 6(1):8–10, February 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1975.11974359.

Flanders:1986:PME

[Fla86] Harley Flanders. On polynomial matrix equations. College Mathematics Journal, 17(5):388–391, November 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1986.11972985.

Flanders:1987:ABMa

[Fla87a] Harley Flanders. Algorithm of the bi-month. College Mathematics Journal, 18(1):53–57, January 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1987.
11973011.

Flanders:1987:ABMb

[Fla87b] Harley Flanders. Algorithm of the bi-month. College Mathematics Journal, 18(2):153–156, March 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1987.
11973027.

Flanders:1987:ABMc

[Fla87c] Harley Flanders. Algorithm of the bi-month. College Mathematics Journal, 18(3):230–235, May 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1987.
11973039.

Flanders:1987:ABMd

[Fla87d] Harley Flanders. Algorithm of the bi-month. College Mathematics Journal, 18(4):323–327, September 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1987.
11973054.

Flanders:1988:ABMa

[Fla88a] Harley Flanders. Algorithm of the bi-month. College Mathematics Journal, 19(1):72–78, January 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1988.
11973096.

Flanders:1988:ABMc

[Fla88b] Harley Flanders. Algorithm of the bi-month. College Mathematics Journal, 19(5):444–447, November 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1988.
11973152.

Flashman:1990:MHa

[Fla90a] Martin E. Flashman. Media highlights. College Mathematics Journal, 21(1):73–80, January 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1990.
11973288.

Flashman:1990:MHb

[Fla90b] Martin E. Flashman. Media highlights. College Mathematics Journal, 21(2):158–165, March 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1990.
11973303.

Flashman:1990:MHc

[Fla90c] Martin E. Flashman. Media highlights. College Mathematics Journal, 21(3):255–262, May 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1990.
11973317.

Flashman:1990:MHd

[Fla90d] Martin E. Flashman. Media highlights. College Mathematics Journal, 21(4):342–349, September 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1990.
11973332.

Flashman:1990:MHe

[Fla90e] Martin E. Flashman. Media highlights. College Mathematics Journal, 21(5):431–437, November 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1990.
11973350.

Flapan:2011:HGT

[Fla11] Erica Flapan. How to be a good teacher is an undecidable problem. College Mathematics Journal, 42(5):350–354, November 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.42.5.350.

Flath:2020:DM

[Fla20] Dan Flath. Derivatives are multipliers. College Mathematics Journal, 51(4):260–270, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2020.1776041.

Fleming:1979:AST

[Fle79] Wendell H. Fleming. 1978 AMS survey: Two-year College report. Two-Year College Mathematics Journal, 10(2):143–144, March 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1979.11972342.

Fleming:1980:TYC

[Fle80] Wendell H. Fleming. Two-year College survey. Two-Year College Mathematics Journal, 11(3):222, June 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1980.
11972445.

Flesner:1993:MH

[Fle93] David E. Flesner. Media highlights. College Mathematics Journal, 24(1):101–108, January 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1993.11973516.

Fleron:1999:GWC

[Fle99] Julian F. Fleron. Gabriel’s wedding cake. College Mathematics Journal, 30(1):35–38, January 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1999.
11974027.

Fiore:2018:TTT

[FLP18] Thomas M. Fiore, Alexander Lang, and Antonella Perucca. Tactile tools for teaching: Implementing Knuth’s algorithm for mastering Mastermind. College Mathematics Journal, 49(4):278– 286, September 2018. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2018.1501989.

Fillers:2014:MN

[FLS14] Whitney Rhianna Fillers, Bill Linderman, and Andrew Simoson. Mancala as Nim. College Mathematics Journal, 45(5):350– 356, November 2014. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.45.5.350.

Flusser:1988:TSR

[Flu88] Peter Flusser. Theory, simulation and reality. College Mathematics Journal, 19(3):210–222, May 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1988.
11973117.

Flynn:1995:NMI

[Fly95] Gerald Flynn. Numerical methods for improper integrals. College Mathematics Journal, 26(4):284–291, September 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1995.11973714.

Ford:1977:FCM

[FM77] Janet E. Ford and Douglas B. McLeod. Flow charts in mathematics classes for elementary school teachers. Two-Year College Mathematics Journal, 8(1):15–19, January 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1977.11974473.

Furi:1994:SID

[FM94] Massimo Furi and Mario Martelli. Strictly increasing differentiable functions. College Mathematics Journal, 25(2):125–127, March 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1994.11973594.

Feeman:2000:SCP

[FM00] Timothy G. Feeman and Osvaldo Marrero. Sequences of chords and of parabolic segments enclosing proportional areas. College Mathematics Journal, 31(5):379–382, November 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2000.11974177.

Feeman:2005:ITL

[FM05] Timothy G. Feeman and Osvaldo Marrero. Intersections of tangent lines of exponential functions. College Mathematics Journal, 36(3):205–208, May 2005. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2005.11922129.

Forbes:1972:WNT

[FMA72] Jack E. Forbes, Lawrence L. Mithchell, and Allen Angel. What’s new in TYC mathematics? Two-Year College Mathematics Journal, 3(1):63–65, Spring 1972. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1972.11974199.

Fredricks:1992:SP

[FN92] Gregory Fredricks and Roger B. Nelsen. Summation by parts. College Mathematics Journal, 23(1):39–42, January 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1992.11973431.

Feist:2004:AAH

[FN04] Curtis Feist and Ramin Naimi. Almost alternating harmonic series. College Mathematics Journal, 35(3):183–191, May 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2004.11922071.

Frame:2007:NMW

[FN07] Michael Frame and Nial Neger. Newton’s Method and the Wada Property: a graphical approach. College Mathematics Journal, 38(3):192–204, May 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2007.11922236.

Feist:2009:TEW

[FN09] Curtis Feist and Ramin Naimi. Topology explains why automobile sunshades fold oddly. College Mathematics Journal, 40 (2):93–98, March 2009. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2009.11922343.

Foote:2016:HAV

[FN16] Robert L. Foote and Han Nie. How to approximate the volume of a lake. College Mathematics Journal, 47(3):162–170, May 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.47.3.162.

Fontenot:1996:TAF

[Fon96] Robert A. Fontenot. Three applications of a familiar formula. College Mathematics Journal, 27(5):356–360, November 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1996.11973808.

Fontenot:2009:CUT

[Fon09] Robert A. Fontenot. Correlation of the union of two bivariate data sets. College Mathematics Journal, 40(5):370–373, November 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/074683409X475698.

Forman:1992:SAP

[For92] Susan Forman. A serendipitous application of the Pythagorean triplets. College Mathematics Journal, 23(4):312–314, September 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1992.11973477.

Forringer:2017:PTE

[For17] Edward R. Forringer. Pythagorean triples for easy solutions of certain quadratics and a newly generated tree. College Mathematics Journal, 48(2):112–114, March 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
48.2.112.

Fossa:1999:LE

[Fos99] John Fossa. Letter to the Editor. College Mathematics Journal, 30(1):67–68, January 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1999.11974034.

Foster:2012:TTH

[Fos12] Colin Foster. Teaching tip: How to manipulate test scores. College Mathematics Journal, 43(2):121–122, March 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2012.11922570.

Foster:2016:PWI

[Fos16] Colin Foster. Proof without words: Integer right triangle hypotenuses without Pythagoras. College Mathematics Journal, 47 (2):101, March 2016. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.47.2.101.

Faiziev:2013:WCO

[FPS13] Valerii Faiziev, Robert Powers, and Prasanna Sahoo. When can one expect a stronger triangle inequality? College Mathematics Journal, 44(1):24–31, January 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
44.1.024.

Francis:1975:PBP

[Fra75] Richard L. Francis. Polygons, both perfect and regular. Two-Year College Mathematics Journal, 6(2):20–21, May 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1975.11974373.

Francis:1978:NAC

[Fra78] Richard L. Francis. A note on angle construction. Two-Year College Mathematics Journal, 9(2):75–80, March 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1978.11974551.

Fraser:1984:GGD

[Fra84] Marshall Fraser. The grazing goat in n dimensions. College Mathematics Journal, 15(2):126–134, March 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1984.
11972761.

Francis:1986:NIC

[Fra86a] Richard L. Francis. From none to infinity: Challenging problems in cardinality classification. College Mathematics Journal, 17(3):226–230, May 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1986.11972960.

Frantz:1986:IGM

[Fra86b] Michael E. Frantz. Interactive graphics for multivariable calculus. College Mathematics Journal, 17(2):172–181, March 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1986.11972953.

Franklin:1987:BTB

[Fra87] LeRoy A. Franklin. Bayes’ Theorem, binominal probabilities, and fair numbers of peremptory challenges in jury trials. College Mathematics Journal, 18(4):291–299, September 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1987.11973046.

Francis:1988:MHA

[Fra88] Richard L. Francis. Mathematical haystacks: Another look at repunit numbers. College Mathematics Journal, 19(3):240–246, May 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1988.11973120.

Franklin:1992:USS

[Fra92] LeRoy A. Franklin. Using simulation to study linear regression. College Mathematics Journal, 23(4):290–295, September 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1992.11973471.

Frame:1997:SR

[Fra97] Michael Frame. Software reviews. College Mathematics Journal, 28(5):398–405, November 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1997.11973902.

Fraenkel:2012:R

[Fra12a] Aviezri S. Fraenkel. Ratwyt. College Mathematics Journal, 43 (2):160–164, March 2012. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.43.2.160.

Frantz:2012:DAP

[Fra12b] Marc Frantz. A different angle on perspective. College Mathematics Journal, 43(5):354–360, November 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
43.5.354.

Frantz:2014:HUC

[Fra14] Marc Frantz. Hyperbola: Under construction! College Mathematics Journal, 45(5):388–390, November 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
45.5.388.

Frantz:2024:NNL

[Fra24] Marc Frantz. The natural natural logarithm. College Mathematics Journal, 55(5):432–434, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic).

Freiling:1986:CBF

[Fre86] Chris Freiling. The change of base formula for logarithms. College Mathematics Journal, 17(5):413, November 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1986.
11972989.

Freund:1988:WRH

[Fre88] John E. Freund. It’s not “by whom,” but rather “how” …. College Mathematics Journal, 19(1):18–20, January 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1988.11973080.

Freund:1995:GAS

[Fre95] Dwight D. Freund. A genuine application of synthetic division, Descartes’ Rule of Signs, and all that stuff. College Mathematics Journal, 26(2):106–110, March 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1995.
11973678.

Frederickson:2003:NWO

[Fre03] Greg N. Frederickson. A new wrinkle on an old folding problem. College Mathematics Journal, 34(4):258–263, September 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2003.11922015.

Frederickson:2008:DTB

[Fre08] Greg N. Frederickson. Designing a table both swinging and stable. College Mathematics Journal, 39(4):258–266, September 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2008.11922302.

Frederickson:2011:FPO

[Fre11] Greg N. Frederickson. Folding polyominoes from one level to two. College Mathematics Journal, 42(4):265–274, September 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.42.4.265.

Frederickson:2012:BRP

[Fre12] Greg N. Frederickson. Bracing regular polygons as we race into the future. College Mathematics Journal, 43(1):51–57, January 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.43.1.051.

Freedman:2013:SW

[Fre13] Walden Freedman. Sharing the work. College Mathematics Journal, 44(3):229–232, May 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.44.3.229.

Friesen:1978:LGE

[Fri78] Charles D. Friesen. A look at general education mathematics programs. Two-Year College Mathematics Journal, 9(4):218– 221, September 1978. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1978.11974582.

Friedlander:1981:FF

[Fri81] Richard J. Friedlander. Factoring factorials. Two-Year College Mathematics Journal, 12(1):12–20, January 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1981.
11972483.

Friedberg:1984:DFD

[Fri84] Stephen H. Friedberg. The distribution of first digits. College Mathematics Journal, 15(2):120–125, March 1984. CO-DEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.jstor.org/stable/2686516;
http://www.tandfonline.com/doi/abs/10.1080/00494925.
1984.11972760.

Friedberg:1988:APC

[Fri88] Stephen H. Friedberg. An alternate proof of Cramer’s Rule. College Mathematics Journal, 19(2):171, March 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1988.
11973108.

Friedberg:1989:PRB

[Fri89] Stephen H. Friedberg. The power rule and the binomial formula. College Mathematics Journal, 20(4):322, September 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1989.11973251.

Fried:2021:OAF

[Fri21] Michael N. Fried. An observation about first-order linear ODEs. College Mathematics Journal, 52(2):137–139, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2021.1868927.

Fried:2024:PWP

[Fri24] Michael N. Fried. Proof without words: a property of tangent circles. College Mathematics Journal, 55(4):343, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2024.2301898.

Frauenthal:1979:FIH

[FS79] James C. Frauenthal and Thomas L. Saaty. Foresight–insight– hindsight. Two-Year College Mathematics Journal, 10(4):245– 254, September 1979. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1979.11972359.

Fowler:1999:RD

[FS99] Brian C. Fowler and Randall J. Swift. Relabeling dice. College Mathematics Journal, 30(3):204–208, May 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/07468342.
1999.11974057.

Fink:2013:OIF

[FS13] Kurt Fink and Jawad Sadek. Other indeterminate forms. College Mathematics Journal, 44(1):55–57, January 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
44.1.055.

Foote:2015:IFC

[FS15] Robert L. Foote and Xidian Sun. An intrinsic formula for the cross ratio in spherical and hyperbolic geometries. College Mathematics Journal, 46(3):182–188, May 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
46.3.182.

Fulghesu:2023:IFI

[FST23] Damiano Fulghesu, James A. Sellers, and Courtney K. Taylor. Infinite families of infinite series with integer sums. College Mathematics Journal, 54(1):33–43, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2022.
2160611.

Flath:2004:RM

[FSW04] Daniel Flath, Cliff Stoll, and Stan Wagon. Rocket math. College Mathematics Journal, 35(4):262–273, September 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2004.11922084.

Flanders:1988:ABMb

[FT88] Harley Flanders and Alan F. Tomala. Algorithm of the bimonth. College Mathematics Journal, 19(4):364–367, September 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1988.11973140.

Freitas:2018:DC

[FT18] Pedro J. Freitas and Hugo Tavares. Dividing the circle. College Mathematics Journal, 49(3):187–194, May 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/full/10.1080/07468342.
2018.1440871.

Fulling:2000:RRC

[Ful00] S. A. Fulling. Related rates collide with vectors. College Mathematics Journal, 31(2):116–119, March 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2000.
11974125.

Fulling:2005:HAI

[Ful05] S. A. Fulling. How to avoid the inverse secant (and even the secant itself). College Mathematics Journal, 36(5):381–387, November 2005. CODEN ???? ISSN 0746-8342 (print), 1931-

1346 (electronic). URL http://www.tandfonline.com/doi/
abs/10.1080/07468342.2005.11922153.

Fusaro:2013:BRC

[Fus13] Ben Fusaro. Book review and commentary. College Mathematics Journal, 44(5):446–449, November 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
44.5.446.

Flint:1994:SR

[FW94] Andrew Flint and Ron Wood. Software reviews. College Mathematics Journal, 25(5):458–461, November 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1994.
11973653.

Fan:2012:PWS

[FZ12] Xingya Fan and Yixin Zhu. Proof without words: The sine is subadditive on [0, π]. College Mathematics Journal, 43(5): 376, November 2012. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.43.5.376.

Galda:1981:IHF

[Gal81a] Klaus Galda. An informal history of formal proofs: From vigor to rigor? Two-Year College Mathematics Journal, 12(2):126–140, March 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972510.

Gallian:1981:OFS

[Gal81b] Joseph A. Gallian. An optimal football strategy. Two-Year College Mathematics Journal, 12(5):330–331, November 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972562.

Gallian:1991:MIN

[Gal91] Joseph A. Gallian. The mathematics of identification numbers. College Mathematics Journal, 22(3):194–202, May 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1991.11973381.

Gannon:1981:VA

[Gan81] Gerald E. Gannon. Visual application of Sin(θ_{1} + θ_{2}) = Sinθ_{1}Cosθ_{2} + Cosθ_{1}Sinθ_{2}. Two-Year College Mathematics Journal, 12(3):206, June 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1981.11972526.

Gantner:1984:CRD

[Gan84] T. E. Gantner. The computation of repeating decimals. College Mathematics Journal, 15(5):436–440, November 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972831.

Ganzell:2017:DTO

[Gan17] Sandy Ganzell. Divisibility tests, old and new. College Mathematics Journal, 48(1):36–40, January 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
48.1.36.

Gao:1997:AEG

[Gao97] Ji Gao. An application of elementary geometry in functional analysis. College Mathematics Journal, 28(1):39–43, January 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1997.11973826.

Gao:2013:SPR

[Gao13] Fuchang Gao. A simple proof of the right-hand rule. College Mathematics Journal, 44(3):227–229, May 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
44.3.227.

Gardner:1982:WRS

[Gar82] Martin Gardner. Who is Raymond smullyan? Two-Year College Mathematics Journal, 13(2):146, March 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1982.
11972594.

Gardner:1993:UNR

[Gar93] Robert B. Gardner. A useful notation for rules of differentiation. College Mathematics Journal, 24(4):351–352, September 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1993.11973554.

Gardner:1999:AP

[Gar99] Martin Gardner. The asymmetric propeller. College Mathematics Journal, 30(1):18–22, January 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1999.
11974024.

Gardner:2000:MMP

[Gar00a] Martin Gardner. Modeling mathematics with playing cards. College Mathematics Journal, 31(3):173–177, May 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/07468342.
2000.11974138.

Gardner:2000:SNR

[Gar00b] Martin Gardner. Some new results on magic hexagrams. College Mathematics Journal, 31(4):274–280, September 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2000.11974158.

Gardner:2001:AO

[Gar01] Martin Gardner. Against the odds. College Mathematics Journal, 32(1):39–43, January 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2001.11921850.

Gardner:2007:BR

[Gar07] Martin Gardner. Book review. College Mathematics Journal, 38 (3):241–242, May 2007. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2007.11922243.

Gardner:2009:TTM

[Gar09] Martin Gardner. L-tromino tiling of mutilated chessboards. College Mathematics Journal, 40(3):162–168, May 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/07468342.
2009.11922354.

Gardner:2012:H

[Gar12] Martin Gardner. Hexaflexagons. College Mathematics Journal, 43(1):2–5, January 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.43.1.002.

Garrett:2015:BR

[Gar15] Christina C. Garrett. Book review. College Mathematics Journal, 46(3):228–232, May 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.46.3.228.

Garcia:2022:VET

[Gar22] Emmanuel Antonio José García. A variant of the eyeball theorem. College Mathematics Journal, 53(2):147–148, 2022.

CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2022.2022905.

Garibaldi:2024:LSE

[Gar24] Skip Garibaldi. The luckiest strategy on Earth? A better way to buy lottery scratch-off tickets. College Mathematics Journal, 55(2):117–125, 2024. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2023.2266315.

Gass:1994:TCE

[Gas94] Frederick Gass. Tangents to conics, eccentrically. College Mathematics Journal, 25(1):43–45, January 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1994.
11973581.

Gasull:2023:PWT

[Gas23] Armengol Gasull. Proof without words: Tangent of π/8. College Mathematics Journal, 54(4):385, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2023.
2238581.

Gauld:1984:AAT

[Gau84] David Gauld. Approximate angle trisection. College Mathematics Journal, 15(5):420–422, November 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1984.
11972824.

Gazdar:1978:CC

[Gaz78] Abdus Sattar Gazdar. Computer corner. Two-Year College Mathematics Journal, 9(3):182–185, June 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1978.
11974573.

Gazdar:1984:DPL

[Gaz84] Abdus Sattar Gazdar. Distance from a point to a line. College Mathematics Journal, 15(4):328–329, September 1984.

CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972793.

Gazdar:1992:DPP

[Gaz92] Abdus Sattar Gazdar. Distance from a point to a plane with a variation on the Pythagorean Theorem. College Mathematics Journal, 23(5):410–412, November 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1992.
11973492.

Garcia-Caballero:2018:DSM

[GCM18] Esther M. García-Caballero and Samuel G. Moreno. The doublesidedness of matrix inverses; yet another proof. College Mathematics Journal, 49(2):136–137, March 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2018.
1426315.

Gomez-Calderon:1996:WPR

[GCW96] Javier Gomez-Calderon and David M. Wells. Why polynomials have roots. College Mathematics Journal, 27(2):90–94, March 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1996.11973757.

Graham-Eagle:2009:DC

[GE09] James Graham-Eagle. The draining cylinder. College Mathematics Journal, 40(5):337–344, November 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://

www.tandfonline.com/doi/abs/10.4169/074683409X475652.

Greer:2013:CUC

[GECW13] Meredith L. Greer, Holly A. Ewing, Kathryn L. Cottingham, and Kathleen C. Weathers. Collaborative understanding of cyanobacteria in lake ecosystems. College Mathematics Journal, 44(5): 376–385, November 2013. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.44.5.376.

Geist:1984:HL

[Gei84] Robert Geist. How to live to be 100. College Mathematics Journal, 15(3):256–263, June 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1984.11972782.

Guggenheimer:1995:SEC

[GEJ95] Heinrich W. Guggenheimer, Alan S. Edelman, and Charles R. Johnson. A simple estimate of the condition number of a linear system. College Mathematics Journal, 26(1):2–5, January 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1995.11973657.

Gemignani:1989:CLI

[Gem89] Michael Gemignani. Copyright law as it applies to computer software. College Mathematics Journal, 20(4):332–338, September 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1989.11973255.

Gentry:1973:GRT

[Gen73] Rodney D. Gentry. Generalizing Rolle’s Theorem in elementary calculus. Two-Year College Mathematics Journal, 4(3):11–17, Autumn 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1973.11974268.

Gentile:1989:PT

[Gen89] Enzo R. Gentile. Pythagorean Theorem: a · a′ + b · b′ = c · c′. College Mathematics Journal, 20(1):58, January 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1989.11973208.

Gentile:1991:API

[Gen91] Enzo R. Gentile. Another proof of the irrationality of \sqrt{2}. College Mathematics Journal, 22(2):143, March 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1991.
11973373.

Gersting:1975:MII

[Ger75] Judith L. Gersting. Mathematical induction: If student k understands it, will student k + 1? Two-Year College Mathematics Journal, 6(2):18–19, May 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1975.11974372.

Gerdes:2000:LDC

[Ger00] Paulus Gerdes. On Lunda-designs and the construction of associated magic squares of order 4p. College Mathematics Journal, 31(3):182–188, May 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2000.11974140.

Grossman:1982:DND

[GG82] Jane T. Grossman and Michael P. Grossman. Dimple or no dimple. Two-Year College Mathematics Journal, 13(1):52–55, January 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1982.11972578.

Gordon:1984:MDC

[GG84] Florence S. Gordon and Sheldon P. Gordon. Mathematical discovery via computer graphics: Hypocycloids and epicycloids. College Mathematics Journal, 15(5):440–443, November 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972832.

Gallant:2022:AB

[GG22a] Robert Gallant and Georg Gunther. Arranging beetles. College Mathematics Journal, 53(1):3–12, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2022.
1991770.

Gothen:2022:CRM

[GG22b] Peter B. Gothen and António Guedes de Oliveira. On the classification of the rigid motions of the plane. College Mathematics Journal, 53(4):315–317, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2022.2099718.

Gothen:2023:ERT

[GG23] Peter Gothen and António Guedes de Oliveira. On Euler’s rotation theorem. College Mathematics Journal, 54(3):171– 175, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2022.2015214.

Guy:1978:LE

[GGF78] Richard Guy, Leonard Gillman, and Marjorie A. Fitting. Letters to the Editor. Two-Year College Mathematics Journal, 9 (4):203–204, September 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1978.11974578.

Grygorian:2018:CPT

[GI18] Artur Grygorian and Ionut E. Iacob. A concise proof of the triangle inequality for the Jaccard distance. College Mathematics Journal, 49(5):363–365, November 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2018.
1526020.

Giangrasso:1973:FRO

[Gia73] Anthony Giangrasso. A flexible response to open admissions. Two-Year College Mathematics Journal, 4(2):55–58, Spring 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1973.11974252.

Gilpin:1982:CC

[Gil82] Michael Gilpin. Computer and calculators. Two-Year College Mathematics Journal, 13(4):252–259, September 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1982.11972623.

Gillman:1988:MS

[Gil88] Leonard Gillman. More on the series for in 2. College Mathematics Journal, 19(3):252–253, May 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1988.
11973123.

Gillman:1991:MTI

[Gil91] Leonard Gillman. More on tabular integration by parts. College Mathematics Journal, 22(5):407–410, November 1991.

CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1991.11973416.

Gilligan:1995:SRa

[Gil95a] Lawrence G. Gilligan. Software reviews. College Mathematics Journal, 26(1):58–65, January 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1995.
11973670.

Gilligan:1995:SRb

[Gil95b] Lawrence G. Gilligan. Software reviews. College Mathematics Journal, 26(3):238–243, May 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1995.
11973707.

Gillman:1997:ORP

[Gil97] Leonard Gillman. Order relations and a proof of l’Hôpital’s rule. College Mathematics Journal, 28(4):288–292, September 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1997.11973877.

Gillman:1998:RAL

[Gil98] Leonard Gillman. Revisiting arc length. College Mathematics Journal, 29(2):137–138, March 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1998.
11973927.

Ginsberg:2004:MNS

[Gin04] Brian D. Ginsberg. Midy’s (nearly) secret theorem — an extension after 165 years. College Mathematics Journal, 35(1):26–30, January 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2004.11922047.

Graff:2023:IIS

[GJJ23] Robert Graff, Kelly Jabbusch, and David James. Intuitive interpretations of SVD vectors. College Mathematics Journal, 54 (3):200–211, 2023. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2023.2201567.

Gersting:1977:WHU

[GK77] Judith L. Gersting and Joseph E. Kuczkowski. Why and how to use small groups in the mathematics classroom. Two-Year College Mathematics Journal, 8(5):270–274, November 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1977.11974526.

Gnanadesikan:1988:STN

[GK88] R. Gnanadesikan and J. R. Kettenring. Statistics teachers need experience with data …. College Mathematics Journal, 19(1): 12–14, January 1988. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1988.11973076.

Goehle:2013:PGS

[GK13] Geoff Goehle and Mitsuo Kobayashi. Polynomial graphs and symmetry. College Mathematics Journal, 44(1):37–42, January 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.44.1.037.

Gross:1975:LE

[GKB^{+}75] Herbert I. Gross, Lawrence L. Krajewski, Ralph P. Boas, R. P. Boas, and D. H. Erkiletian. Letters to the Editor. Two-Year College Mathematics Journal, 6(3):1–3, September 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1975.11974382.

Gasparini:2019:GAN

[GKR19] Andrea Gasparini, Eric Key, and David Radcliffe. A geometric approach to the natural exponential function. College Mathematics Journal, 50(5):357–363, 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2019.1664846.

Gilbert:1992:SBP

[GL92] George T. Gilbert and Loren C. Larson. A sliding block problem. College Mathematics Journal, 23(4):315–319, September 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1992.11973479.

Graver:2000:LFR

[GL00] Jack E. Graver and Lawrence J. Lardy. Linear functions and rounding. College Mathematics Journal, 31(2):132–136, March 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2000.11974131.

Glaister:1993:DWG

[Gla93] P. Glaister. Does what goes up take the same time to come down? College Mathematics Journal, 24(2):155–158, March 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1993.11973520.

Glass:2012:SPA

[Gla12] Darren Glass. The Secretary Problem from the applicant’s point of view. College Mathematics Journal, 43(1):76–81, January 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.43.1.076.

Gray:1999:WAL

[GM99] S. I. B. Gray and Tagui Malakyan. The Witch of Agnesi: a lasting contribution from the first surviving mathematical work written by a woman. College Mathematics Journal, 30(4):258– 268, September 1999. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1999.11974069.

Gilligan:1974:WD

[GN74] Lawrence G. Gilligan and Robert B. Nenno. Who dunnit? Two-Year College Mathematics Journal, 5(1):78–79, Winter 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1974.11974300.

Glomski:2012:EDL

[GO12] Matthew Glomski and Edward Ohanian. Eradicating a disease: Lessons from mathematical epidemiology. College Mathematics Journal, 43(2):123–132, March 2012. CODEN ????

ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
43.2.123.

Goberstein:1988:EUF

[Gob88] Simon M. Goberstein. Evaluating “uniformly filled” determinants. College Mathematics Journal, 19(4):343–345, September 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1988.11973134.

Godfrey:1988:SMT

[God88] A. Blanton Godfrey. Should mathematicians teach statistics? …. College Mathematics Journal, 19(1):8–10, January 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1988.11973074.

Goebeler:2011:CDF

[Goe11] Thomas E. Goebeler, Jr. The Cobb–Douglas function and Hölder’s Inequality. College Mathematics Journal, 42(5):387– 389, November 2011. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.42.5.387.

Goldberg:1975:PUA

[Gol75] Samuel Goldberg. A precalculus unit on area under curves. Two-Year College Mathematics Journal, 6(4):29–35, December 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1975.11974404.

Goldberg:1983:IWM

[Gol83a] Dorothy Goldberg. Integrating writing into the mathematics curriculum. Two-Year College Mathematics Journal, 14(5):421– 424, November 1983. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1983.11972731.

Goldstine:1983:FTN

[Gol83b]  Herman Goldstine. Faster than von Neumann?  Two-

Year College Mathematics Journal, 14(5):389, November 1983.

CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1983.11972725.

Gollwitzer:1994:MPU

[Gol94] Herman Gollwitzer. Matrix patterns and undetermined coefficients. College Mathematics Journal, 25(5):444–448, November 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1994.11973650.

Goldberg:2010:BR

[Gol10] Samuel Goldberg. Book review. College Mathematics Journal, 41(4):339–340, September 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/074683410X510344.

Gold:2011:HYP

[Gol11a] Bonnie Gold. How your philosophy of mathematics impacts your teaching! College Mathematics Journal, 42(3):174–182, May 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.42.3.174.

Goldberg:2011:BR

[Gol11b] Samuel Goldberg. Book review. College Mathematics Journal, 42(5):413–415, November 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.42.5.413.

Goldberg:2013:BRP

[Gol13] Samuel Goldberg. Book review: Probability Tales by Charles M. Grinstead, William P. Peterson, and J. Laurie Snell. College Mathematics Journal, 44(1):64, January 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
44.1.064.

Goldberg:2016:AGC

[Gol16] Timothy E. Goldberg. Algebra from geometry in the card game SET. College Mathematics Journal, 47(4):265–273, September 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.47.4.265.

Gold:2017:BR

[Gol17] Bonnie Gold. Book review. College Mathematics Journal, 48(3): 226–232, May 2017. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.48.3.226.

Goodner:1970:CTY

[Goo70] Dwight B. Goodner. The CUPM and the two-year College. Two-Year College Mathematics Journal, 1(1):88–90, Spring 1970. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1970.11973173.

Goodhart:2008:R

[Goo08] Gregory S. Goodhart. Review. College Mathematics Journal, 39(4):324–326, September 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2008.11922311.

Gordon:1982:CC

[Gor82] Sheldon P. Gordon. Computers and calculators. Two-Year College Mathematics Journal, 13(1):22–27, January 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1982.11972574.

Gordon:1984:EUL

[Gor84] Sheldon P. Gordon. Evaluating e^{x} using limits. College Mathematics Journal, 15(1):63–65, January 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1984.
11972752.

Gordon:1986:NSS

[Gor86] Florence S. Gordon. A nonstandard solution to a standard problem. College Mathematics Journal, 17(1):74, January 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1986.11972934.

Gorfin:1989:ESS

[Gor89] Alan Gorfin. Evaluating the sum of the series ∑_{k=1}^{∞} k^{j}/M^{k}. College Mathematics Journal, 20(4):329–331, September 1989.

CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1989.11973254.

Gordon:1992:CCC

[Gor92] Sheldon P. Gordon. Classroom computer capsules: Individualized computer investigations for calculus. College Mathematics Journal, 23(5):426–428, November 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1992.
11973495.

Gordon:1993:TPA

[Gor93] Sheldon P. Gordon. Taylor polynomial approximations in polar coordinates. College Mathematics Journal, 24(4):325–330, September 1993. CODEN ???? ISSN 0746-8342 (print), 1931-

1346 (electronic). URL http://www.tandfonline.com/doi/
abs/10.1080/07468342.1993.11973549.

Gordon:1994:RSE

[Gor94] Sheldon P. Gordon. Riemann sums and the exponential function. College Mathematics Journal, 25(1):39–40, January 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1994.11973579.

Gordon:2003:GMA

[Gor03] Russell A. Gordon. A generalization of a minimum area problem. College Mathematics Journal, 34(1):21–23, January 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2003.11921980.

Gordon:2008:SCS

[Gor08] Russell A. Gordon. Squaring a circular segment. College Mathematics Journal, 39(3):212–220, May 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2008.
11922295.

Gordon:2024:CKH

[Gor24] Marshall Gordon. Chen’s and Kennedy’s harmonic series proof as a teaching tool. College Mathematics Journal, 55(3):255– 256, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2023.2291293.

Gant:1975:CRR

[GP75] Student Keith Gant and Dean B. Priest. The countability of the rationals revisited. Two-Year College Mathematics Journal, 6(3):41–42, September 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1975.11974393.

Galit:1976:IRC

[GP76] Mark D. Galit and John P. Pace. Inter-related concepts: an example. Two-Year College Mathematics Journal, 7(1):7–10, February 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1976.11974411.

Guy:2002:NCS

[GP02] Richard K. Guy and Marc M. Paulhus. Nine cubits or simple Soma. College Mathematics Journal, 33(3):188–195, May 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2002.11921939.

Grotto:2024:RSP

[GPV24] Francesco Grotto, Antonella Perucca, and Tatjana Van Steenbergen Bergeron. Rubik’s snakes on a plane. College Mathematics Journal, 55(4):317–327, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2023.2293416.

Gilligan:1974:CC

[GR74] Lawrence G. Gilligan and Nelson G. Rich. Computer corner. Two-Year College Mathematics Journal, 5(4):42–48, December 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1974.11974352.

Grant:1990:MRP

[Gra90] Douglass L. Grant. Moments on a Rose petal. College Mathematics Journal, 21(3):225–227, May 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1990.
11973312.

Grant:1994:LSC

[Gra94] Hardy Grant. Leibniz and the spell of the continuous. College Mathematics Journal, 25(4):291–294, September 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1994.11973624.

Graver:1995:GAL

[Gra95] Jack E. Graver. A geometric approach to linear functions. College Mathematics Journal, 26(5):389–394, November 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1995.11973737.

Grant:1999:MLAa

[Gra99a] Hardy Grant. Mathematics and the liberal arts. College Mathematics Journal, 30(2):96–105, March 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1999.
11974039.

Grant:1999:MLAb

[Gra99b] Hardy Grant. Mathematics and the liberal arts-II. College Mathematics Journal, 30(3):197–203, May 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1999.
11974056.

Grant:2003:BR

[Gra03] Hardy Grant. Book review. College Mathematics Journal, 34 (4):343–348, September 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2003.11922027.

Graver:2011:ETG

[Gra11a] Jack E. Graver. An elementary treatment of general inner products. College Mathematics Journal, 42(1):57–59, January 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.42.1.057.

Graver:2011:MCE

[Gra11b] Jack E. Graver. Making change efficiently. College Mathematics Journal, 42(4):317–322, September 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
42.4.317.

Gray:2018:BRW

[Gra18] Shirley B. Gray. Book review: The World of Maria Gaetana Agnesi, Mathematician of God by Massimo Mazzotti, The Johns Hopkins University Press, Baltimore, MD, 2007. xx + 217 pp., ISBN 978-0-801-88709-3, $56.95 (hardback); 2018, ISBN 978-1-421-42515-3, $24.95 (paperback). College Mathematics Journal, 49(3):229–232, May 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2018.1440864.

Graf:2023:WGS

[Gra23] Otis F. Graf Jr. Why the golden spiral is golden. College Mathematics Journal, 54(5):434–445, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2023.
2273194.

Greenberg:1973:DI

[Gre73] Benjamin Greenberg. A “doodling” inequality. Two-Year College Mathematics Journal, 4(1):78–79, Winter 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/00494925.
1973.11974244.

Greenberg:1974:SIC

[Gre74a] Benjamin Greenberg. Some insight into the convex quadrilateral. Two-Year College Mathematics Journal, 5(3):14–17, September 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1974.11974328.

Greger:1974:RSP

[Gre74b] Karl Greger. Random sieving and the prime number theorem. Two-Year College Mathematics Journal, 5(1):41–46, Winter 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1974.11974291.

Greene:1975:DRA

[Gre75] Myles Greene. Distributivity with respect to all four rational operations. Two-Year College Mathematics Journal, 6(2):10–12, May 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1975.11974370.

Green:1976:IIE

[Gre76] Thomas M. Green. Identities, inequalities and equations: a computer-graphical approach. Two-Year College Mathematics Journal, 7(1):33–37, February 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1976.
11974419.

Greenwood:1977:TFI

[Gre77] Michael E. Greenwood. Two factors involved in successful individualized mathematics programs. Two-Year College Mathematics Journal, 8(4):219–222, September 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1977.
11974514.

Greenwell:1985:UML

[Gre85] Raymond N. Greenwell. Using Minitab in linear algebra. College Mathematics Journal, 16(3):216–218, June 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1985.
11972885.

Green:2000:PDE

[Gre00] Bo Green. A project for discovery, extension, and generalization in abstract algebra. College Mathematics Journal, 31(4):329– 332, September 2000. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2000.11974169.

Green:2011:ATS

[Gre11] William R. Green. Abel’s Theorem simplifies reduction of order. College Mathematics Journal, 42(5):399–402, November 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.42.5.399.

Greenwell:2016:BR

[Gre16] Raymond N. Greenwell. Book review. College Mathematics Journal, 47(4):307–311, September 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
47.4.307.

Green:2020:NFM

[Gre20] William R. Green. The natural frequency: More natural and more frequent than expected. College Mathematics Journal, 51(5):372–374, 2020. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2020.1811045.

Grinstein:1971:IUC

[Gri71] Louise S. Grinstein. Integration by undetermined coefficients. Two-Year College Mathematics Journal, 2(2):98–101, Autumn 1971. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1971.11974021.

Grinstein:1974:CM

[Gri74] Louise S. Grinstein. Calculus by mistake. Two-Year College Mathematics Journal, 5(4):49–53, December 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1974.
11974353.

Grinstein:1978:NIS

[Gri78] Louise S. Grinstein. A note on infinite series. Two-Year College Mathematics Journal, 9(1):46–47, January 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/00494925.
1978.11974546.

Grime:2017:BWN

[Gri17] James Grime. The bizarre world of nontransitive dice: Games for two or more players. College Mathematics Journal, 48(1):2–9, January 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.48.1.2.

Groetsch:1993:IPT

[Gro93] C. W. Groetsch. Inverse problems and Torricelli’s Law. College Mathematics Journal, 24(3):210–217, May 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/07468342.
1993.11973530.

Growney:1996:GC

[Gro96] JoAnne Growney. Geometry class. College Mathematics Journal, 27(2):143, March 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1996.11973767.

Groetsch:1997:HGR

[Gro97] C. W. Groetsch. Halley’s gunnery rule. College Mathematics Journal, 28(1):47–50, January 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1997.
11973829.

Groetsch:2005:ABS

[Gro05a] C. W. Groetsch. Another broken symmetry. College Mathematics Journal, 36(2):109–113, March 2005. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2005.
11922117.

Grove:2005:PST

[Gro05b] Larry Grove. The Platonic solids from their rotation groups. College Mathematics Journal, 36(4):278–283, September 2005. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2005.11922140.

Groetsch:2008:DR

[Gro08] C. W. Groetsch. The depletion ratio. College Mathematics Journal, 39(1):43–48, January 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2008.11922273.

Groetsch:2010:SBP

[Gro10] C. W. Groetsch. Suspension bridge profiles. College Mathematics Journal, 41(3):237–241, May 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.

tandfonline.com/doi/abs/10.4169/074683410X488737.

Groetsch:2016:MST

[Gro16] C. W. Groetsch. De Morgan’s series test. College Mathematics Journal, 47(2):136–137, March 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
47.2.136.

Grobler:2019:VRS

[Gro19] Trienko Grobler. Visualization of the Riemann–Stieltjes integral. College Mathematics Journal, 50(3):198–209, 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2019.1580109.

Glesser:2019:EIA

[GRSS19] Adam Glesser, Matt Rathbun, Isabel M. Serrano, and Bogdan D. Suceava. Eclectic illuminism: Applications of affine geometry. College Mathematics Journal, 50(2):82–92, 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2019.1565588.

Grunbaum:1981:SWT

[Grü81] Branko Grünbaum. Shouldn’t we teach GEOMETRY? Two-Year College Mathematics Journal, 12(4):232–238, September 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972537.

Grunbaum:1984:CVD

[Grü84a] Branko Grünbaum. The construction of Venn diagrams. College Mathematics Journal, 15(3):238–247, June 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1984.
11972776.

Grunbaum:1984:VDC

[Grü84b] Branko Grünbaum. On Venn diagrams and the counting of regions. College Mathematics Journal, 15(5):433–435, November 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972830.

Gruszka:1994:BEC

[Gru94] Thomas Gruszka. A balloon experiment in the classroom. College Mathematics Journal, 25(5):442–444, November 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1994.11973649.

Goldstone:2021:SPC

[GRV21] Richard Goldstone, Rachel Roca, and Robert Suzzi Valli. Shortest paths on cubes. College Mathematics Journal, 52(2):121– 132, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2021.1866944.

Grzesik:2024:NDG

[Grz24] J. A. Grzesik. A novel derivation of the Gregory–Leibniz series for pi. College Mathematics Journal, 55(4):346–348, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2023.2286167.

Gearhart:1990:F

[GS90a] William B. Gearhart and Harris S. Shultz. The function sin x/x. College Mathematics Journal, 21(2):90–99, March 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1990.11973290.

Gopen:1990:WAL

[GS90b] George D. Gopen and David A. Smith. What’s an assignment like you doing in a course like this?: Writing to learn mathematics. College Mathematics Journal, 21(1):2–19, January 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1990.11973277.

Grimaldi:1996:RRC

[GS96] Ralph P. Grimaldi and Douglas R. Shier. Redundancy and reliability of communication networks. College Mathematics Journal, 27(1):59–67, January 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1996.11973752.

Gearhart:2003:TBH

[GS03] W. B. Gearhart and H. S. Shultz. Tugging a barge with hyperbolic functions. College Mathematics Journal, 34(1):42–49, January 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2003.11921984.

Gupta:2010:CPR

[GS10] Shiv K. Gupta and Waclaw Szymanski. Cubic polynomials with rational roots and critical points. College Mathematics Journal, 41(5):365–369, November 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/074683410X521964.

Gani:2011:RBR

[GS11] Joseph Gani and Randall Swift. Random breakage of a rod into unit lengths. College Mathematics Journal, 42(3):201–205, May 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.42.3.201.

Greenwell:1987:SR

[GSH87] Raymond Greenwell, Carol M. Smith, and Eugene A. Herman. Software reviews. College Mathematics Journal, 18(5):422–424, November 1987. CODEN ???? ISSN 0746-8342 (print), 1931-

1346 (electronic). URL http://www.tandfonline.com/doi/
abs/10.1080/07468342.1987.11973069.

Glesser:2015:ICP

[GSS15] Adam Glesser, James Shade, and Bogdan D. Suceava. On the inverse curvature problem. College Mathematics Journal, 46(3): 207–214, May 2015. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.46.3.207.

Gueron:2000:FFFa

[Gue00a] Shay Gueron. Fallacies, flaws, and flimflam. College Mathematics Journal, 31(2):120–123, March 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2000.
11974126.

Gueron:2000:FFFb

[Gue00b] Shay Gueron. Fallacies, flaws, and flimflam. College Mathematics Journal, 31(3):205–207, May 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2000.
11974144.

Gueron:2001:SNA

[Gue01] Shay Gueron. Speeding up a numerical algorithm. College Mathematics Journal, 32(1):33–38, January 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2001.
11921849.

Guggenheimer:1992:SLD

[Gug92] H. Guggenheimer. Systems of linear differential equations by Laplace transform. College Mathematics Journal, 23(3):196–202, May 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1992.11973456.

Guggenheimer:1996:FLG

[Gug96] Heinrich W. Guggenheimer. Formulas of linear geometry. College Mathematics Journal, 27(1):24–32, January 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1996.
11973745.

Guichard:1995:SSB

[Gui95] David R. Guichard. Sums of selected binomial coefficients. College Mathematics Journal, 26(3):209–213, May 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/07468342.
1995.11973698.

Gulick:1993:SR

[Gul93] Frances Gulick. Software reviews. College Mathematics Journal, 24(2):178–182, March 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1993.11973526.

Gurtas:2019:AIP

[Gur19] Yusuf Z. Gurtas. An alternative to integration by partial fractions technique. College Mathematics Journal, 50(2):140– 142, 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2019.1561125.

Gurtas:2022:UAS

[Gür22] Yusuf Z. Gürtas. An unorthodox approach to skinning a definite integral. College Mathematics Journal, 53(2):134–139, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2022.2011543.

Guthery:2019:BRL

[Gut19] Scott Guthery. Book review: Lost in Math: How Beauty Leads Physics Astray by Sabine Hossenfelder: Basic Books, New York, 2018. 304 pp., ISBN 978-0-465-09425-7, $20.40. College Mathematics Journal, 50(2):150–152, 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2019.
1563387.

Guy:1982:JHC

[Guy82] Richard K. Guy. John Horton Conway: Mathematical magus. Two-Year College Mathematics Journal, 13(5):290–299, November 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1982.11972636.

Guy:1984:HCW

[Guy84] Richard K. Guy. How can we lead in an up-to-date and fair fashion? …. College Mathematics Journal, 15(5):396–397, November 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972819.

Goldstone:2019:UC

[GV19] Richard Goldstone and Robert Suzzi Valli. Unfoldings of the cube. College Mathematics Journal, 50(3):173–184, 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2019.1580108.

Gaeta:2023:EPO

[GV23] Michele Gaeta and Giovanni Vincenzi. Examples of pairs of ordered congruent-like n-gons with different areas. College Mathematics Journal, 54(2):99–103, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2023.
2190299.

Ganzell:2024:KCC

[GV24] Sandy Ganzell and Caroline VanBlargan. Knots with composite colors. College Mathematics Journal, 55(3):205–209, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2023.2294670.

Gustafson:1970:PRA

[GW70] R. David Gustafson and Arnold Wendt. Progress report on articulation in Illinois. Two-Year College Mathematics Journal, 1 (1):37–40, Spring 1970. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1970.11973166.

Gray:2018:ARC

[GW18] Shirley Gray and Cye H. Waldman. Archimedes redux: Center of mass applications from the method. College Mathematics Journal, 49(5):346–352, November 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2018.
1524647.

Grove:2000:CS

[GY00] L. Grove and O. Yiparaki. Contumacious spheres. College Mathematics Journal, 31(1):35–41, January 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2000.
11974105.

Heft:1970:SRC

[HA70] Phil Heft and Charles Ainley. Spring retreat for community College mathematics teachers in Washington. Two-Year College Mathematics Journal, 1(2):56–58, Autumn 1970. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1970.11973189.

Haack:1991:CMC

[Haa91] Joel K. Haack. Clapping music — a combinatorial problem. College Mathematics Journal, 22(3):224–227, May 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1991.11973385.

Huehn:1988:SR

[HAC88] K. L. Huehn, Thomas J. Abram, and Stephen Chang. Software reviews. College Mathematics Journal, 19(1):78–80, January 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1988.11973097.

Hadlock:2013:UM

[Had13] Charles R. Hadlock. Underground mathematics. College Mathematics Journal, 44(5):364–375, November 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
44.5.364.

Hahn:1988:WSI

[Hah88] Gerald J. Hahn. What should the introductory statistics course contain? …. College Mathematics Journal, 19(1):26–29, January 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1988.11973084.

Hahn:1998:THA

[Hah98] Alexander J. Hahn. Two historical applications of calculus. College Mathematics Journal, 29(2):93–103, March 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1998.11973922.

Hall:1980:EVP

[Hal80] W. J. Hall. An expected value problem revisited. Two-Year College Mathematics Journal, 11(3):204–205, June 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1980.11972438.

Halmos:1982:TA

[Hal82] P. R. Halmos. The thrills of abstraction. Two-Year College Mathematics Journal, 13(4):243–251, September 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1982.11972622.

Halmos:1985:PTB

[Hal85] P. R. Halmos. Pure thought is better yet. College Mathematics Journal, 16(1):14–16, January 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1985.
11972845.

Hall:1992:TRT

[Hal92] Leon M. Hall. Trochoids, roses, and thorns-beyond the spirograph. College Mathematics Journal, 23(1):20–35, January 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1992.11973429.

Hall:1993:CAD

[Hal93] Richard L. Hall. Computer-aided delusions. College Mathematics Journal, 24(4):366–369, September 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1993.
11973557.

Halpern:2010:TP

[Hal10] Nicole Yunger Halpern. Three poems. College Mathematics Journal, 41(3):233–234, May 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/074683410X488980.

Halleck:2012:TTC

[Hal12] Ezra Halleck. Teaching tip: Consider a circular cow. College Mathematics Journal, 43(2):133, March 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2012.
11922572.

Halevy:2013:OTN

[Hal13] Avner Halevy. Old tails and new trails in high dimensions. College Mathematics Journal, 44(1):48–52, January 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.44.1.048.

Hamming:1984:CDM

[Ham84] R. W. Hamming. Calculus and discrete mathematics …. College Mathematics Journal, 15(5):388–389, November 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972813.

Hammersley:1985:TAE

[Ham85] J. M. Hammersley. Three algorithmic exercises. College Mathematics Journal, 16(1):12–14, January 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1985.
11972844.

Hammack:2007:GVC

[Ham07] Richard Hammack. A geometric view of complex trigonometric functions. College Mathematics Journal, 38(3):210–217, May 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2007.11922238.

Hammett:2020:ELS

[Ham20] Adam Hammett. Euler’s limit and Stirling’s estimate. College Mathematics Journal, 51(5):330–336, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2020.
1811058.

Hanner:1970:MSG

[Han70] Olof Hanner. Mathematics, a solitary game. Two-Year College Mathematics Journal, 1(2):5–16, Autumn 1970. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/00494925.
1970.11973180.

Hanson:1982:IMW

[Han82] Robert Hanson. Integer matrices whose inverses contain only integers. Two-Year College Mathematics Journal, 13(1):18–21, January 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1982.11972572.

Hanson:1985:SII

[Han85] Robert Hanson. Self-inverse integer matrices. College Mathematics Journal, 16(3):190–198, June 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1985.
11972879.

Hanson:1988:LET

[Han88] David L. Hanson. Let the experts teach and judge …. College Mathematics Journal, 19(1):21–23, January 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1988.
11973082.

Hanson:1990:MDS

[Han90] Robert Hanson. Minimum dimension for a square matrix of order n. College Mathematics Journal, 21(1):28–34, January 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1990.11973279.

Hanin:2020:TDP

[Han20] Leonid Hanin. A tour of discrete probability guided by a problem in genomics. College Mathematics Journal, 51(4):284– 294, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2020.1782100.

Hartnett:1972:UF

[Har72] William E. Hartnett. On the use of functions. Two-Year College Mathematics Journal, 3(2):25–28, Autumn 1972. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1972.11974212.

Hardy:1986:BPT

[Har86] Michael Hardy. Behold! The Pythagorean Theorem via mean proportions. College Mathematics Journal, 17(5):422, November 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1986.11972994.

Harris:1987:BS

[Har87] Edward M. Harris. Behold! Sums of Arctan. College Mathematics Journal, 18(2):141, March 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1987.
11973025.

Harper:1988:ESA

[Har88] James D. Harper. Estimating the sum of alternating series. College Mathematics Journal, 19(2):149–153, March 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1988.11973103.

Hartzler:1995:DRC

[Har95] J. S. Hartzler. Designing a Rose cutter. College Mathematics Journal, 26(1):41–43, January 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1995.
11973666.

Harkleroad:1996:HMK

[Har96a] Leon Harkleroad. How mathematicians know what computers can’t do. College Mathematics Journal, 27(1):37–42, January 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1996.11973747.

Hartman:1996:TDF

[Har96b] James L. Hartman. A terminally discontinuous function. College Mathematics Journal, 27(3):211–212, May 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/07468342.
1996.11973781.

Hare:1997:SR

[Har97] Angela C. Hare. Software reviews. College Mathematics Journal, 28(2):140–144, March 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1997.11973853.

Harkleroad:2008:TT

[Har08] Leon Harkleroad. Tuning with triangles. College Mathematics Journal, 39(5):367–373, November 2008. CODEN ????

ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2008.
11922317.

Hartman:2010:CDI

[Har10] Jim Hartman. Computing definite inegrals using the definition. College Mathematics Journal, 41(1):58–60, January 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/074683410X475128.

Hashisaki:1971:E

[Has71a] Joseph Hashisaki. From the Editor. Two-Year College Mathematics Journal, 2(2):12–13, Autumn 1971. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1971.
11974007.

Hashisaki:1971:MMT

[Has71b] Joseph Hashisaki. The MAA and the mathematics teacher in the two-year College. Two-Year College Mathematics Journal, 2(2): 63–68, Autumn 1971. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1971.11974014.

Hativa:1983:WMM

[Hat83] Nira Hativa. What makes mathematics lessons easy to follow, understand, and remember? Two-Year College Mathematics Journal, 14(5):398–406, November 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1983.
11972727.

Hathout:2004:OHP

[Hat04] Heba Hathout. The old hats problem revisited. College Mathematics Journal, 35(2):97–102, March 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2004.
11922058.

Hathout:2007:JHA

[Hat07] Leith Hathout. Not just hats anymore: Binomial inversion and the problem of multiple coincidences. College Mathematics Journal, 38(3):179–184, May 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2007.11922234.

Hathaway:2011:UCI

[Hat11] Dan Hathaway. Using continuity induction. College Mathematics Journal, 42(3):229–231, May 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
42.3.229.

Haunsperger:2005:SNA

[Hau05] Deanna Haunsperger. Saari, with no apologies. College Mathematics Journal, 36(2):90–100, March 2005. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2005.
11922115.

Haunsperger:2008:BMI

[Hau08] Deanna Haunsperger. A break for mathematics: an interview with Joe Gallian. College Mathematics Journal, 39(3):174–190, May 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2008.11922292.

Haver:1978:DSC

[Hav78] William E. Haver. Developing skills in college algebra — a mastery approach. Two-Year College Mathematics Journal, 9 (5):282–287, November 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1978.11974593.

Hayashi:1992:FNR

[Hay92] Elmer K. Hayashi. Fibonacci numbers, recursion, complexity, and induction proofs. College Mathematics Journal, 23(5):407– 410, November 1992. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1992.11973491.

Hill:1992:CCCa

[HB92] Robert J. Hill and Robert D. Bechtel. Classroom computer capsules: Eigenvectors and Jordan bases using symbolic programs. College Mathematics Journal, 23(1):59–63, January 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1992.11973437.

Hathaway:1997:FPF

[HB97] Dale K. Hathaway and Stephen L. Brown. Fibonacci powers and a fascinating triangle. College Mathematics Journal, 28(2): 124–128, March 1997. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1997.11973848.

Hector:1975:MAM

[Hec75] Judith Harle Hector. A mastery approach to mathematical literacy. Two-Year College Mathematics Journal, 6(2):22–27, May 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1975.11974374.

Hecker:2003:CPL

[Hec03] David Hecker. Constructing a Poincaré line with straightedge and compass. College Mathematics Journal, 34(5):362–366, November 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2003.11922031.

Hedman:2000:CMQ

[Hed00] Bruce Hedman. Colin Maclaurin’s quaint word problems. College Mathematics Journal, 31(4):286–289, September 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2000.11974160.

Hegarty:1987:DMC

[Heg87] John C. Hegarty. A depreciation model for calculus classes. College Mathematics Journal, 18(3):219–221, May 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/07468342.
1987.11973035.

Hegarty:1996:BDA

[Heg96] John C. Hegarty. Bond duration: an application of calculus. College Mathematics Journal, 27(1):47–49, January 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1996.11973749.

Heiny:1978:USD

[Hei78] Robert L. Heiny. The use of sports data for integrating topics in introductory statistics. Two-Year College Mathematics Journal, 9(1):28–33, January 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1978.11974542.

Hellman:1973:SAL

[Hel73] Morton J. Hellman. Some applications of the law of the contrapositive. Two-Year College Mathematics Journal, 4(3):86–88, Autumn 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1973.11974281.

Hendel:1990:SP

[Hen90] Russell Jay Hendel. (x^{n})′ = nx^{n−1}: Six proofs. College Mathematics Journal, 21(4):312–313, September 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1990.11973326.

Hendel:1991:DPF

[Hen91] Russell Jay Hendel. Differentiation via partial fractions: a case against CAS. College Mathematics Journal, 22(5):415– 417, November 1991. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1991.11973419.

Hendel:1994:AFT

[Hen94] Russell Jay Hendel. Approaches to the formula for the n th Fibonacci number. College Mathematics Journal, 25(2):139–142, March 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1994.11973598.

Henle:1995:SR

[Hen95] Michael G. Henle. Software reviews. College Mathematics Journal, 26(4):320–328, September 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1995.
11973724.

Henle:2014:BR

[Hen14] Michael Henle. Book review. College Mathematics Journal, 45 (4):308–312, September 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.45.4.308.

Henrich:2022:HGY

[Hen22] Allison Henrich. How to Get Your PhD: A Handbook for the Journey by Gavin Brown. College Mathematics Journal, 53(3): 235–236, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2022.2045150.

Heo:2015:PWP

[Heo15] Nam Gu Heo. Proof without words: The Pythagorean Theorem. College Mathematics Journal, 46(1):51, January 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.46.1.51.

Hersh:1990:LTP

[Her90] Reuben Hersh. Let’s teach philosophy of mathematics! College Mathematics Journal, 21(2):105–111, March 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1990.
11973292.

Hern:1993:GEI

[Her93] Thomas Hern. Gaussian elimination in integer arithmetic: an application of the l-u factorization. College Mathematics Journal, 24(1):67–71, January 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1993.11973509.

Hersh:1998:MH

[Her98] Reuben Hersh. Media highlights. College Mathematics Journal, 29(1):73–80, January 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1998.11973919.

Hersh:2002:OG

[Her02] Reuben Hersh. The “origin” of geometry. College Mathematics Journal, 33(3):207–211, May 2002. CODEN ????

ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2002.
11921942.

Hersh:2003:IT

[Her03] Reuben Hersh. Independent thinking. College Mathematics Journal, 34(2):112–115, March 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2003.11921993.

Hersh:2004:HAF

[Her04] Reuben Hersh. Heron’s area formula: What about a tetrahedron? College Mathematics Journal, 35(2):112–114, March 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2004.11922061.

Hersh:2009:BR

[Her09] Reuben Hersh. Book reviews. College Mathematics Journal, 40(2):142–145, March 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2009.11922350.

Hersh:2010:URT

[Her10] Reuben Hersh. Under-represented then over-represented: a memoir of Jews in American mathematics. College Mathematics Journal, 41(1):2–9, January 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/074683410X475065.

Hersh:2012:WFP

[Her12] Reuben Hersh. Why the Faulhaber polynomials are sums of even or odd powers of (n + 1/2). College Mathematics Journal, 43 (4):322–324, September 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.43.4.322.

Hesterberg:1995:SSD

[Hes95] Tim Hesterberg. The spider’s spacewalk derivation of sin′ and cos′. College Mathematics Journal, 26(2):144–145, March 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1995.11973687.

Hetzler:1997:CVN

[Het97] Steven M. Hetzler. A continuous version of Newton’s Method. College Mathematics Journal, 28(5):348–351, November 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1997.11973888.

Herz-Fischler:1993:VPT

[HF93] Roger Herz-Fischler. A “very pleasant theorem”. College Mathematics Journal, 24(4):318–324, September 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1993.
11973548.

Hess:2008:HPB

[HGGB08] Richard Hess, Charles Grinstead, Marshall Grinstead, and Deborah Bergstrand. Hermit points on a Box. College Mathematics Journal, 39(1):12–23, January 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2008.
11922270.

Howard:1988:EI

[HH88] Jim Howard and Joe Howard. Equivalent inequalities. College Mathematics Journal, 19(4):350–352, September 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1988.11973138.

Hickey:1970:UCM

[Hic70] Albert E. Hickey. The use of the computer in mathematics instruction. Two-Year College Mathematics Journal, 1(1):44–54, Spring 1970. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1970.11973168.

Higgins:2023:LBP

[Hig23] Peter M. Higgins. Limiting behaviour of pairs with equal images for polynomial functions. College Mathematics Journal, 54(4):391–393, 2023. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2023.2237854.

Hilton:1980:MASa

[Hil80a] Peter Hilton. Math anxiety: Some suggested causes and cures. Two-Year College Mathematics Journal, 11(3):174–188, June 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1980.11972432.

Hilton:1980:MASb

[Hil80b] Peter Hilton. Math anxiety: Some suggested causes and cures: Part 2. Two-Year College Mathematics Journal, 11(4):246–251, September 1980. CODEN ???? ISSN 0049-4925 (print), 2325-

9116 (electronic). URL http://www.tandfonline.com/doi/
abs/10.1080/00494925.1980.11972449.

Hilton:1984:MH

[Hil84a] Peter J. Hilton. Media highlights. College Mathematics Journal, 15(3):274–283, June 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1984.11972785.

Hilton:1984:NCU

[Hil84b] Peter J. Hilton. The need for coherence and unity …. College Mathematics Journal, 15(5):375, November 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1984.
11972806.

Hildebrand:1985:GPT

[Hil85a] W. J. Hildebrand. Generalized Pythagorean triples. College Mathematics Journal, 16(1):48–52, January 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1985.
11972849.

Hilton:1985:AE

[Hil85b] Peter Hilton. Algorithms are not enough. College Mathematics Journal, 16(1):8–9, January 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1985.11972841.

Hillel:1989:NC

[Hil89] J. Hillel. Notational collisions. College Mathematics Journal, 20(5):418–422, November 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1989.11973270.

Hildebrand:1990:CDP

[Hil90] Wilbur J. Hildebrand. Connecting the dots parametrically: an alternative to cubic splines. College Mathematics Journal, 21(3): 208–215, May 1990. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1990.11973307.

Hill:1992:CCCb

[Hil92a] David R. Hill. Classroom computer capsules: Interpolating polynomials and their coordinates relative to a basis. College Mathematics Journal, 23(4):329–333, September 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1992.11973481.

Hilton:1992:JMM

[Hil92b] Peter Hilton. The joy of mathematics: a Mary P. Dolciani lecture. College Mathematics Journal, 23(4):274–281, September 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1992.11973469.

Hilton:1996:MH

[Hil96] Peter Hilton. Media highlights. College Mathematics Journal, 27(2):156–163, March 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1996.11973770.

Hirschhorn:1998:FTA

[Hir98] Michael D. Hirschhorn. The fundamental theorem of algebra. College Mathematics Journal, 29(4):276–277, September 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1998.11973954.

Hirschhorn:2000:MC

[Hir00] Michael D. Hirschhorn. More on cootie. College Mathematics Journal, 31(2):126–128, March 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2000.
11974128.

Hirose:2024:CSP

[Hir24] Kenichi Hirose. Computing sums of powers of integers with the hockey-stick identity and the Stirling numbers. College Mathematics Journal, 55(3):247–249, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2023.
2264719.

Huddy:2017:CBG

[HJ17] Stanley R. Huddy and Michael A. Jones. The calculus behind generic drug equivalence. College Mathematics Journal, 49(1): 2–9, January 2017. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2017.1391502.

Hwang:2019:OCP

[HK19] Suk-Geun Hwang and Ik-Pyo Kim. Orientation of the cross product of 3-vectors. College Mathematics Journal, 50(4):298– 299, 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2019.1656509.

Hale:2022:MHN

[HKM22] Cameron G. Hale, Jonathan R. Kelleher, and John C. Mayer. Multiplicity of hexagon numbers. College Mathematics Journal, 53(5):335–346, 2022. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2022.2120326.

Hamilton:2015:SSP

[HKP15] Gordon Hamilton, Kiran S. Kedlaya, and Henri Picciotto. Square-sum pair partitions. College Mathematics Journal, 46 (4):264–269, September 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.46.4.264.

Hohn:2012:PD

[HL12] Tiina Hohn and Andy Liu. Polyomino dissections. College Mathematics Journal, 43(1):88–94, January 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
43.1.088.

Hern:1998:LOI

[HLL98] Thomas Hern, Cliff Long, and Andy Long. Looking at order of integration and a minimal surface. College Mathematics Journal, 29(2):128–133, March 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1998.11973925.

Hill:1995:MEF

[HM95] Robert J. Hill and Mark S. Mazur. The matrix exponential function and systems of differential equations using Derive. College Mathematics Journal, 26(2):146–151, March 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1995.11973688.

Hirst:1997:BRP

[HM97] Holly P. Hirst and Wade T. Macey. Bounding the roots of polynomials. College Mathematics Journal, 28(4):292–295, September 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1997.11973878.

Humphreys:2005:UGV

[HM05] L. D. Humphreys and P. J. McKenna. Using a gradient vector to find multiple periodic oscillations in suspension bridge models. College Mathematics Journal, 36(1):16–26, January 2005. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2005.11922106.

Hill:2010:CCC

[HM10] Theodore P. Hill and Kent E. Morrison. Cutting cakes carefully. College Mathematics Journal, 41(4):281–288, September 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/074683410X510272.

Holley:2023:LAS

[HM23] Alexander Holley and Anastasiia Minenkova. A linear algebra story: How we reconstructed a matrix from its eigenvalues. College Mathematics Journal, 54(5):454–463, 2023.

CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2023.2266319.

Ho:2020:SPC

[HMF20] Chungwu Ho, Gregory Mellblom, and Marc Frodyma. On the sum of powers of consecutive integers. College Mathematics Journal, 51(4):295–301, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2020.1776051.

Henrich:2011:MKD

[HMN^{+}11] A. Henrich, N. MacNaughton, S. Narayan, O. Pechenik, R. Silversmith, and J. Townsend. A midsummer knot’s dream. College Mathematics Journal, 42(2):126–134, March 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
42.2.126.

Hubbard:1994:NLA

[HMNW94] John H. Hubbard, Jean Marie McDill, Anne Noonburg, and Beverly H. West. A new look at the Airy equation with fences and funnels. College Mathematics Journal, 25(5):419–431, November 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1994.11973646.

Hodge:2010:GC

[HMP10] Jonathan K. Hodge, Emily Marshall, and Geoff Patterson. Gerrymandering and convexity. College Mathematics Journal, 41 (4):312–324, September 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/074683410X510317.

Horvath:2011:ODC

[HMT11] Eszter K. Horváth, Attila Máder, and Andreja Tepavcević. Onedimensional Czédli-type islands. College Mathematics Journal, 42(5):374–378, November 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.42.5.374.

Hummel:1991:SR

[HN91] J. A. Hummel and W. P. Nanry. Software review. College Mathematics Journal, 22(4):335–342, September 1991.

CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1991.11973410.

Hoaglin:1988:ULI

[Hoa88] David C. Hoaglin. Using leverage and influence to introduce regression diagnostics. College Mathematics Journal, 19(5):387– 416, November 1988. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1988.11973146.

Hoban:2013:ESH

[Hob13] Ryan Hoban. Euclidean, spherical, and hyperbolic shadows. College Mathematics Journal, 44(2):126–134, March 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.44.2.126.

Hoehn:1984:GIW

[Hoe84] Larry Hoehn. A geometrical interpretation of the weighted mean. College Mathematics Journal, 15(2):135–139, March 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972762.

Hoehn:1987:SPP

[Hoe87] Larry Hoehn. Some properties of polygons inside a circle. College Mathematics Journal, 18(5):397–401, November 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1987.11973062.

Hoenig:1990:SR

[Hoe90] Alan Hoenig. Software reviews. College Mathematics Journal, 21(2):146–149, March 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1990.11973301.

Hoehn:1997:CTG

[Hoe97] Larry Hoehn. A concurrency theorem and Geometer’s Sketchpad. College Mathematics Journal, 28(2):129–132, March 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1997.11973850.

Hoensch:2009:PAC

[Hoe09] Ulrich A. Hoensch. A “paperclip” approach to curvature, torsion, and the Frénet–Serret formulas. College Mathematics Journal, 40(2):113–118, March 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2009.11922346.

Hoffman:1974:PP

[Hof74] Nathan Hoffman. Partitions of the plane. Two-Year College Mathematics Journal, 5(2):71–73, Spring 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1974.
11974320.

Hoffman:1978:BGR

[Hof78a] Nathan Hoffman. Binary grids and a related counting problem. Two-Year College Mathematics Journal, 9(5):267–272, November 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1978.11974590.

Hoffman:1978:SNW

[Hof78b] Nathan Hoffman. Some new ways of solving a coin tossing problem. Two-Year College Mathematics Journal, 9(1):6–10, January 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1978.11974536.

Hofstadter:1982:AME

[Hof82] Douglas R. Hofstadter. Analogies and metaphors to explain Gödel’s Theorem. Two-Year College Mathematics Journal, 13 (2):98–114, March 1982. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1982.11972588.

Hoffmann:2016:WGI

[Hof16] Heiko Hoffmann. Weighted AM-GM inequality via elementary multivariable calculus. College Mathematics Journal, 47(1):56– 58, January 2016. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.47.1.56.

Hogg:1984:MH

[Hog84] Robert V. Hogg. Media highlights. College Mathematics Journal, 15(1):76–84, January 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1984.11972756.

Hollev:1976:FNI

[Hol76] Ann D. Hollev. Functional notation — an intuitive approach. Two-Year College Mathematics Journal, 7(3):14–15, September 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1976.11974438.

Holley:1978:QI

[Hol78] Ann D. Holley. A question of interest. Two-Year College Mathematics Journal, 9(2):81–83, March 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1978.
11974552.

Holley:1979:MST

[Hol79] Ann D. Holley. Mathematics in seventeen three-hour lessons: a challenge. Two-Year College Mathematics Journal, 10(3):191– 192, June 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1979.11972349.

Holden:1981:ACA

[Hol81a] Herbert L. Holden. Applying complex arithmetic. Two-Year College Mathematics Journal, 12(3):190–194, June 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972522.

Holley:1981:IGI

[Hol81b] Ann D. Holley. Integration by geometric insight — a student’s approach. Two-Year College Mathematics Journal, 12(4):268– 270, September 1981. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1981.11972543.

Holden:1982:CP

[Hol82] Herb Holden. Chords of the parabola. Two-Year College Mathematics Journal, 13(3):186–190, June 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1982.
11972605.

Holly:2023:LIS

[Hol23] Melissa Holly. Large independent set vs. large clique. College Mathematics Journal, 54(4):374–377, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2023.
2237851.

Honsberger:1979:MGb

[Hon79a] Ross Honsberger. Mathematical gems. Two-Year College Mathematics Journal, 10(2):108–111, March 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1979.
11972334.

Honsberger:1979:MGc

[Hon79b] Ross Honsberger. Mathematical gems. Two-Year College Mathematics Journal, 10(3):195–197, June 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1979.
11972351.

Honsberger:1979:MGd

[Hon79c] Ross Honsberger. Mathematical gems. Two-Year College Mathematics Journal, 10(4):271–276, September 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1979.
11972364.

Honsberger:1979:MGe

[Hon79d] Ross Honsberger. Mathematical gems. Two-Year College Mathematics Journal, 10(5):344–347, November 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1979.
11972382.

Honsberger:1979:MGa

[Hon79e] Ross A. Honsberger. Mathematical gems. Two-Year College Mathematics Journal, 10(1):30–32, January 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1979.
11972319.

Honsberger:1980:MGa

[Hon80a] Ross Honsberger. Mathematical gems. Two-Year College Mathematics Journal, 11(2):116–119, March 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1980.
11972418.

Honsberger:1980:MGb

[Hon80b] Ross Honsberger. Mathematical gems. Two-Year College Mathematics Journal, 11(3):196–200, June 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1980.
11972435.

Honsberger:1980:MGc

[Hon80c] Ross Honsberger. Mathematical gems. Two-Year College Mathematics Journal, 11(5):305–312, November 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1980.
11972467.

Honsberger:1981:MGa

[Hon81a] Ross Honsberger. Mathematical gems. Two-Year College Mathematics Journal, 12(1):34–40, January 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1981.
11972490.

Honsberger:1981:MGb

[Hon81b] Ross Honsberger. Mathematical gems. Two-Year College Mathematics Journal, 12(4):280–282, September 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1981.
11972548.

Honsberger:1981:MGc

[Hon81c] Ross Honsberger. Mathematical gems. Two-Year College Mathematics Journal, 12(5):325–329, November 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1981.
11972560.

Honsberger:1982:MGa

[Hon82a] Ross Honsberger. Mathematical gems. Two-Year College Mathematics Journal, 13(1):36–44, January 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1982.
11972576.

Honsberger:1982:MGb

[Hon82b] Ross Honsberger. Mathematical gems. Two-Year College Mathematics Journal, 13(4):260–265, September 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1982.
11972624.

Honsberger:1983:MGa

[Hon83a] Ross Honsberger. Mathematical gems. Two-Year College Mathematics Journal, 14(1):2–8, January 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1983.
11972646.

Honsberger:1983:MGb

[Hon83b] Ross Honsberger. Mathematical gems. Two-Year College Mathematics Journal, 14(2):154–158, March 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1983.
11972671.

Honsberger:1987:MH

[Hon87] Ross Honsberger. Media highlights. College Mathematics Journal, 18(1):78–85, January 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1987.11973015.

Honsberger:1995:MMG

[Hon95] Ross A. Honsberger. More mathematical gems. College Mathematics Journal, 26(4):281–283, September 1995. CODEN ????

ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1995.
11973713.

Hong:2015:WGF

[Hon15] Dae S. Hong. When is the generating function of the Fibonacci numbers an integer? College Mathematics Journal, 46(2):110– 112, March 2015. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.46.2.110.

Hooley:1997:UCM

[Hoo97] Donald E. Hooley. Using the College mathematics journal topic index in undergraduate courses. College Mathematics Journal, 28(2):106–109, March 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1997.11973844.

Hooley:2000:CMA

[Hoo00] Donald E. Hooley. Collapsed matrices with (almost) the same eigenstuff. College Mathematics Journal, 31(4):297–299, September 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2000.11974162.

Hopkins:2010:TT

[Hop10] Brian Hopkins. Taking turns. College Mathematics Journal, 41 (4):289–297, September 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/074683410X510281.

Hopkins:2012:YBS

[Hop12] Brian Hopkins. 30 years of Bulgarian solitaire. College Mathematics Journal, 43(2):135–140, March 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
43.2.135.

Hopkins:2014:BR

[Hop14a] Brian Hopkins. Book review. College Mathematics Journal, 45 (5):400–405, November 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.45.5.400.

Hopkins:2014:EAY

[Hop14b] Brian Hopkins. From the Editor: Anonymity and youth. College Mathematics Journal, 45(2):82, March 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
45.2.082.

Hopkins:2014:ECC

[Hop14c] Brian Hopkins. From the Editor: Constancy and change. College Mathematics Journal, 45(1):2, January 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
45.1.002.

Hopkins:2015:ECD

[Hop15] Brian Hopkins. From the Editor: Celebration and delay. College Mathematics Journal, 46(1):2, January 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
46.1.2.

Hopkins:2016:BR

[Hop16] Brian Hopkins. Book review. College Mathematics Journal, 47 (5):375–380, November 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.47.5.375.

Hopkins:2018:BRC

[Hop18a] Brian Hopkins. Book review: Combinatorics: a Very Short Introduction by Robin Wilson, Oxford University Press, Oxford, 2016, xviii + 158 pp., ISBN 978-0-19-872349-3, $11.95 paperback, $7.99 electronic. Infinity: a Very Short Introduction by Ian Stewart, Oxford University Press, Oxford, 2017, xvi + 144 pp., ISBN 978-0-19-875523-4, $11.95 paperback, $7.99 electronic. College Mathematics Journal, 49(2):147–152, March 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2018.1432202.

Hopkins:2018:EAT

[Hop18b] Brian Hopkins. From the Editor: All things move. College Mathematics Journal, 49(5):322, November 2018. CODEN ????

ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2018.
1528807.

Hopkins:2019:BRP

[Hop19] Brian Hopkins. Book review: Pythagorean Triangles by Waclaw Sierpiński, trans. Ambikeshwar Sharma, Yeshiva University, New York, 1962. Reprint, Dover, Mineola, NY, 2013. viii + 72 pp., ISBN 978-0-486-43278-6, $12.95 (paperback). College Mathematics Journal, 50(1):68–72, January 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2019.
1547955.

Hopkins:2020:FDC

[Hop20] Brian Hopkins. The fifth decade of the CMJ. College Mathematics Journal, 51(5):358–359, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2020.1822068.

Horner:1974:NNN

[Hor74] Donald R. Horner. Nonlab, nonprogrammed, and nonlecture: Any chance? Two-Year College Mathematics Journal, 5(2): 39–44, Spring 1974. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1974.11974312.

Hornsby:1990:GGS

[Hor90a] E. John Hornsby, Jr. Geometrical and graphical solutions of quadratic equations. College Mathematics Journal, 21(5):362– 378, November 1990. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1990.11973334.

Horowitz:1990:TIP

[Hor90b] David Horowitz. Tabular integration by parts. College Mathematics Journal, 21(4):307–311, September 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1990.
11973325.

Hosack:1986:SR

[Hos86] John M. Hosack. Software reviews. College Mathematics Journal, 17(5):434–441, November 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1986.11972996.

Hoseana:2018:ZZF

[Hos18a] Jonathan Hoseana. On zero-over-zero form limits of a special type. College Mathematics Journal, 49(3):219–221, May 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2018.1445916.

Hoseana:2018:VBM

[Hos18b] Jonathan Hoseana. Variations on the binary mind-reading trick. College Mathematics Journal, 49(4):262–268, September 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2018.1500061.

Hotchkiss:2002:PR

[Hot02] Philip K. Hotchkiss. It’s perfectly rational. College Mathematics Journal, 33(2):113–117, March 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2002.
11921928.

Howard:2002:GCR

[How02] Fredric T. Howard. A generalized Chinese Remainder Theorem. College Mathematics Journal, 33(4):279–282, September 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2002.11921953.

Howard:2018:P

[How18] C. Douglas Howard. It’s puzzling. College Mathematics Journal, 49(4):242–249, September 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2018.1500062.

Hilton:1993:GGU

[HP93] Peter Hilton and Jean Pedersen. Geometry: a gateway to understanding. College Mathematics Journal, 24(4):298–317, September 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1993.11973547.

Hilton:1999:RGAa

[HP99a] Peter Hilton and Jean Pedersen. Relating geometry and algebra in the Pascal triangle, hexagon, tetrahedron, and cuboctahedron Part I: Binomial coefficients, extended binomial coefficients and preparation for further work. College Mathematics Journal, 30(3):170–186, May 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1999.11974053.

Hilton:1999:RGAb

[HP99b] Peter Hilton and Jean Pedersen. Relating geometry and algebra in the Pascal triangle, hexagon, tetrahedron, and cuboctahedron part II: Geometry and algebra in higher dimensions: Identifying the Pascal cuboctahedron. College Mathematics Journal, 30 (4):279–292, September 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1999.11974072.

Heneghan:2014:PSM

[HP14] Fiacha Heneghan and Kyle Petersen. Power series for up-down min-max permutations. College Mathematics Journal, 45(2): 83–91, March 2014. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.45.2.083.

Harary:1984:CIG

[HR84] Frank Harary and Robert W. Robinson. Connect-it games. College Mathematics Journal, 15(5):411–419, November 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972823.

Hammett:2024:SLS

[HR24] Adam Hammett and Kevin Roper. A sterling look at Stirling’s constant. College Mathematics Journal, 55(1):24–31, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2023.2216126.

Hocking:1986:EBP

[HS86] Robert L. Hocking and Neil C. Schwertman. An extension of the birthday problem to exactly k matches. College Mathematics Journal, 17(4):315–321, September 1986.

CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1986.11972972.

Hardin:1990:TPL

[HS90] Doug Hardin and Gilbert Strang. A thousand points of light. College Mathematics Journal, 21(5):406–409, November 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1990.11973345.

Howard:1996:COR

[HS96] Ralph Howard and Paul Sisson. Capturing the origin with random points: Generalizations of a Putnam problem. College Mathematics Journal, 27(3):186–192, May 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/07468342.
1996.11973774.

Humphreys:2000:FUB

[HS00] L. D. Humphreys and R. Shammas. Finding unpredictable behavior in a simple ordinary differential equation. College Mathematics Journal, 31(5):338–346, November 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2000.
11974171.

Hsiao:1989:SML

[Hsi89] Frank S. T. Hsiao. The simplex method of linear programming on microcomputer spreadsheets. College Mathematics Journal, 20(2):153–160, March 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1989.11973225.

Holmes:1977:SIG

[HSL77] Allen H. Holmes, Walter Sanders, and John LeDuc. Statistical inference for the general education student-it can be done. Two-Year College Mathematics Journal, 8(4):223–230, September 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1977.11974515.

Hsu:1980:AWL

[Hsu80] David Hsu. Another way of looking at n!. Two-Year College Mathematics Journal, 11(5):333–335, November 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1980.11972476. See note [Mos84].

Hsu:1982:SPF

[Hsu82] David Y. Hsu. Sums of powers of the first n integers. Two-Year College Mathematics Journal, 13(3):196–197, June 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1982.11972608.

He:2005:ICM

[HSZ05] Tian-Xiao He, Zachariah Sinkala, and Xiaoya Zha. If F (x) = ∫_{x}^{2x} f(t) dt is constant, must f(t) = c/t? College Mathematics Journal, 36(3):199–204, May 2005. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2005.
11922128.

Henry:1984:ITA

[HT84] Michael Henry and G. E. Trapp, Jr. Income tax averaging and convexity. College Mathematics Journal, 15(3):253–255, June 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972781.

Hu:2021:PWC

[Hu21] Xiaoyan Hu. Proof without words: Center of mass. College Mathematics Journal, 52(4):297, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2021.1959208.

Huang:1988:DLR

[Hua88] Xun-Cheng Huang. A discrete L’Hôpital’s Rule. College Mathematics Journal, 19(4):321–329, September 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1988.
11973132.

Huang:1991:SPF

[Hua91] Xun-Cheng Huang. A shortcut in partial fractions. College Mathematics Journal, 22(5):413–415, November 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1991.11973418.

Huang:2025:NIC

[Hua25] Shaowu Huang. A new insight into Cramer’s rule via the determinants of block matrices. College Mathematics Journal, 56(1): 63–64, 2025. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic).

Huber:1981:PSC

[Hub81] John Huber. Partial and semipartial correlation — a vector approach. Two-Year College Mathematics Journal, 12(2):151–153, March 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972516.

Hubbard:1994:WIM

[Hub94] John H. Hubbard. What it means to understand a differential equation. College Mathematics Journal, 25(5):372–384, November 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1994.11973641.

Huddleston:2011:TTP

[Hud11] Meg B. Huddleston. Teaching tip: Practice integration on problem triplets. College Mathematics Journal, 42(3):214, May 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2011.11922504.

Huerta:1992:WNS

[Hue92] Carlos C. Huerta. Who needs the sine anyway? College Mathematics Journal, 23(1):43–44, January 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1992.
11973432.

Huff:1989:CEC

[Huf89] Barthel W. Huff. Conditional expectations and the correlation function. College Mathematics Journal, 20(1):55–57, January 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1989.11973206.

Hughes:1979:PR

[Hug79] Barnabas B. Hughes. Polygonal roots. Two-Year College Mathematics Journal, 10(5):313–318, November 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1979.
11972376.

Hughes:1989:HAD

[Hug89] Barnabas B. Hughes. Hippocrates and Archytas double the cube: a heuristic interpretation. College Mathematics Journal, 20(1): 42–48, January 1989. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1989.11973201.

Hughes:2000:AFS

[Hug00] James R. Hughes. Applications of Fourier series in classical guitar technique. College Mathematics Journal, 31(4):300–303, September 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2000.11974163.

Hughes:2006:BR

[Hug06] Christopher Hughes. Book review. College Mathematics Journal, 37(2):161–164, March 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2006.11922179.

Hundhausen:1998:ZDF

[Hun98] Joan R. Hundhausen. Zeroing in on the delta function. College Mathematics Journal, 29(1):27–33, January 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1998.
11973909.

Hutchins:1983:LND

[Hut83] Harry Hutchins. License numbers and divisibility rules. Two-Year College Mathematics Journal, 14(2):122–125, March 1983.

CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1983.11972668.

Huber:1984:DEE

[HW84] John C. Huber and Joseph Wiener. Deriving the equations of the ellipse and hyperbola. College Mathematics Journal, 15(1):58–59, January 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972750.

Hopkins:2004:TAK

[HW04] Brian Hopkins and Robin J. Wilson. The truth about Königsberg. College Mathematics Journal, 35(3):198–207, May 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2004.11922073.

Hwang:2012:RPP

[Hwa12] Suk-Geun Hwang. A real proof of the principal axis theorem. College Mathematics Journal, 43(2):172–173, March 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.43.2.172.

Hyde:1988:PTC

[Hyd88] Kendell Hyde. Pascal triangles and combinations where repetitions are allowed. College Mathematics Journal, 19(1):60–62, January 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1988.11973094.

Isihara:2013:MPT

[IB13] Paul Isihara and Doeke Buursma. Multi-peg Tower of Hanoi. College Mathematics Journal, 44(2):110–116, March 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
44.2.110.

Imbrogno:2022:CNW

[IB22] Jason Imbrogno and David Bergman. Computing the number of winning NFL survivor pool entries. College Mathematics Journal, 53(4):282–291, 2022. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2022.2099704.

Iga:2008:TDS

[Iga08] Kevin Iga. The Truck Driver’s Straw Problem and Cantor sets. College Mathematics Journal, 39(4):280–290, September 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2008.11922305.

Iga:2006:TDS

[IK06] Kevin Iga and Kendra Killpatrick. Truck drivers, a straw, and two glasses of water. College Mathematics Journal, 37(2):82–92, March 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2006.11922170.

Indika:2024:EET

[IL24] S. H. Sathish Indika and Lawrence M. Leemis. Exact expressions for trigonometric functions. College Mathematics Journal, 55(1):40–45, 2024. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2023.2241316.

Illich:2017:WM

[Ill17] Donald Illich. Water mathematics. College Mathematics Journal, 48(3):170, May 2017. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.48.3.170.

Iltis:2000:RNW

[Ilt00] Richard Iltis. Runs with no winner in a lottery. College Mathematics Journal, 31(5):356–361, November 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2000.
11974174.

Intermont:2011:AGT

[IM11] Michele Intermont and Aileen Murphy. An application of group theory to change ringing. College Mathematics Journal, 42(3): 223–228, May 2011. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.42.3.223.

Iacob:2012:VFT

[IMW12] Ionut E. Iacob, T. Bruce McLean, and Hua Wang. The v-flex, triangle orientation, and Catalan numbers in hexaflexagons. College Mathematics Journal, 43(1):6–10, January 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.43.1.006.

Insel:1989:DPI

[Ins89] Arnold J. Insel. A direct proof of the integral formula for arctangent. College Mathematics Journal, 20(3):235–237, May 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1989.11973238.

Insel:1993:RMP

[Ins93] Arnold J. Insel. Rotation matrices in the plane without trigonometry. College Mathematics Journal, 24(1):71–73, January 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1993.11973510.

Isaksen:1995:LAG

[Isa95] Daniel C. Isaksen. Linear algebra on the gridiron. College Mathematics Journal, 26(5):358–360, November 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1995.11973729.

Isaksen:1996:HKF

[Isa96] Daniel C. Isaksen. How to Kick a field goal. College Mathematics Journal, 27(4):267–271, September 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1996.
11973790.

Iversen:1988:SET

[Ive88] Gudmund R. Iversen. Statistician, examine thyself. College Mathematics Journal, 19(1):16–18, January 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1988.
11973079.

Jacobson:1997:PML

[Jac97] Richard A. Jacobson. Paths of minimum length in a regular tetrahedron. College Mathematics Journal, 28(5):394–397, November 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1997.11973901.

Janardan:1980:FON

[Jan80] Konanur G. Janardan. Further observations on “a neglected probability formula”. Two-Year College Mathematics Journal, 11(1):52–54, January 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1980.11972403. See [SY78].

Javier:2017:PWT

[Jav17] García Capitán Francisco Javier. Proof without words: Tangents of 15 and 75 degrees. College Mathematics Journal, 48(1):35, January 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.48.1.35.

Jensen:1992:NTA

[JB92] David W. Jensen and Eric R. Bussian. A number-theoretic approach to counting subgroups of dihedral groups. College Mathematics Journal, 23(2):150–152, March 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1992.
11973449.

James:2013:USV

[JB13] David James and Cynthia Botteron. Understanding singular vectors. College Mathematics Journal, 44(3):220–226, May 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.44.3.220.

Jones:1989:SR

[JBC89] Albyn Jones, Amberse Banks, and Judith L. Clarke. Software reviews. College Mathematics Journal, 20(3):252–255, May 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1989.11973241.

Johnson:1990:WL

[JD90] Diane Johnson and Roy Dowling. Ways of looking at n! College Mathematics Journal, 21(3):219–220, May 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/07468342.
1990.11973309.

Johnsonbaugh:1998:DIV

[JD98] Richard Johnsonbaugh and Duane W. DeTemple. A discrete intermediate value theorem. College Mathematics Journal, 29 (1):42, January 1998. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1998.11973914.

Jepsen:1979:BBN

[Jep79] Charles H. Jepsen. Billiard balls and a number theory result. Two-Year College Mathematics Journal, 10(5):306–312, November 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1979.11972375.

Jepsen:1986:CPU

[Jep86] Charles H. Jepsen. Coloring points in the unit square. College Mathematics Journal, 17(3):231–237, May 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/07468342.
1986.11972961.

Jungreis:1988:HMB

[JF88] Douglas S. Jungreis and Erich Friedman. How many bridge auctions? College Mathematics Journal, 19(2):171–172, March 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1988.11973109.

Joseph:2011:GP

[JHG11] Dan Joseph, Gregory Hartman, and Caleb Gibson. Generalized parabolas. College Mathematics Journal, 42(4):275–282, September 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.42.4.275.

Janardan:1978:CC

[JJ78] Satish K. Janardan and Konanur G. Janardan. Computer corner. Two-Year College Mathematics Journal, 9(4):243–248, September 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1978.11974586.

Johnson:1998:CHP

[JK98] Charles R. Johnson and Brenda K. Kroschel. Clock hands pictures for 2 × 2 real matrices. College Mathematics Journal, 29 (2):148–150, March 1998. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1998.11973933.

James:2011:SVS

[JLR11] David James, Michael Lachance, and Joan Remski. Singular vectors’ subtle secrets. College Mathematics Journal, 42(2):86–95, March 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.42.2.086.

Johnson:1997:PPT

[JM97] Aimee Johnson and Kathleen Madden. Putting the pieces together: Understanding Robinson’s nonperiodic tilings. College Mathematics Journal, 28(3):172–181, May 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1997.
11973858.

Jones:2024:WRP

[JMW24] Michael A. Jones, Allison Mocny, and Jennifer Wilson. The Will Rogers phenomenon and sequential migration. College Mathematics Journal, 55(3):235–244, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2024.
2305210.

Jobbings:2011:DFS

[Job11] Andrew Jobbings. From the dance of the foci to a strophoid. College Mathematics Journal, 42(4):289–298, September 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.42.4.289.

Johnson:1972:AQN

[Joh72a] Phillip E. Johnson. Academic qualifications of North Carolina’s community college professors. Two-Year College Mathematics Journal, 3(2):33–36, Autumn 1972. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1972.
11974214.

Johnson:1972:GDS

[Joh72b] Phillip E. Johnson. The Genesis and development of set theory. Two-Year College Mathematics Journal, 3(1):55–62, Spring 1972. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1972.11974198.

Johnson:1976:ECB

[Joh76] Phillip E. Johnson. Eighteenth Century British mathematics. Two-Year College Mathematics Journal, 7(2):22–27, May 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1976.11974425.

Johnston:1989:ACO

[Joh89] Elgin H. Johnston. An alternative to changing the order of integration. College Mathematics Journal, 20(5):405–408, November 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1989.11973265.

Johnson:1993:TVO

[Joh93a] Jerry Johnson. Tangent vectors and orthogonal projections. College Mathematics Journal, 24(3):259–262, May 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/07468342.
1993.11973542.

Johnson:1993:HDN

[Joh93b] Roger W. Johnson. How does the NFL rate the passing ability of quarterbacks? College Mathematics Journal, 24(5):451– 453, November 1993. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1993.11973566.

Johnson:1994:PS

[Joh94] Roger W. Johnson. Problems and solutions. College Mathematics Journal, 25(4):334–340, September 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1994.
11973636.

Johnson:2000:GOT

[Joh00] Colonel Johnson, Jr. Group operation tables and normalizers. College Mathematics Journal, 31(1):50–51, January 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2000.11974108.

Johnson:2004:BP

[Joh04] Nils P. Johnson. The brachistochrone problem. College Mathematics Journal, 35(3):192–197, May 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2004.
11922072.

Johnson:2013:NGH

[Joh13] Jeremiah William Johnson. The number of group homomorphisms from D_{m} into D_{n}. College Mathematics Journal, 44(3): 190–192, May 2013. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.44.3.190.

Johnson:2017:BCM

[Joh17] Roger W. Johnson. Bet(ch)a my team wins the playoffs. College Mathematics Journal, 48(5):347–353, November 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
48.5.347.

Johnson:2022:DEE

[Joh22] Jeremiah W. Johnson. A Diophantine equation with an elementary solution. College Mathematics Journal, 53(5):361– 363, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2022.2118993.

Jones:1971:BMC

[Jon71] J. A. Jones. Basic mathematics for colleges — the CUPM recommendations. Two-Year College Mathematics Journal, 2(2):87–92, Autumn 1971. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1971.11974017.

Jones:1974:SCJ

[Jon74] David L. Jones. A study of the coefficient j(n, i). Two-Year College Mathematics Journal, 5(4):12–15, December 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1974.11974344.

Jones:1983:SSP

[Jon83] Lenny K. Jones. Sequences, series, and Pascal’s Triangle. Two-Year College Mathematics Journal, 14(3):253–256, June 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1983.11972691.

Jones:1985:MH

[Jon85] Charles V. Jones. Media highlights. College Mathematics Journal, 16(1):81–86, January 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1985.11972859.

Jongsma:1993:SER

[Jon93] Calvin Jongsma. Stacking ellipses-revisited. College Mathematics Journal, 24(5):453, November 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1993.
11973567.

Jones:2004:WLD

[Jon04] Michael A. Jones. Win, lose, or draw: a Markov chain analysis of overtime in the national football league. College Mathematics Journal, 35(5):330–336, November 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2004.
11922094.

Jones:2010:ISJ

[Jon10] Michael A. Jones. An interview with Steven J. Brams. College Mathematics Journal, 41(4):262–267, September 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
4169/074683410X510254.

Jones:2011:MMR

[Jon11] Michael A. Jones. Mathematical minute: Removing a dot. College Mathematics Journal, 42(2):139, March 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2011.
11922490.

Jones:2012:PWS

[Jon12a] Michael A. Jones. Proof without words: The square of a balancing number is a triangular number. College Mathematics Journal, 43(3):212, May 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.43.3.212.

Jones:2012:EIL

[Jon12b] Timothy W. Jones. Euler’s identity, Leibniz tables, and the irrationality of pi. College Mathematics Journal, 43(5):361– 364, November 2012. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.43.5.361.

Jones:2020:CEL

[Jon20a] Brian D. Jones. Chance encounters with large polynomials. College Mathematics Journal, 51(3):174–181, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2020.
1718966.

Jones:2020:DEA

[Jon20b] Michael A. Jones. A difference equation approach to finite differences of polynomials. College Mathematics Journal, 51(5):375– 377, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2020.1760065.

Jones:2024:TMD

[Jon24] Brian D. Jones. Taxicab metric of a different kind. College Mathematics Journal, 55(4):328–335, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2023.
2266314.

Jordan:1977:ECC

[Jor77] James H. Jordan. An elementary construction of the common log tables. Two-Year College Mathematics Journal, 8(5):274– 278, November 1977. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1977.11974527.

Joshi:1983:ETP

[Jos83] Padmini T. Joshi. Efficient techniques for partial fractions. Two-Year College Mathematics Journal, 14(2):110–118, March 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1983.11972666.

Josevich:2016:TDM

[Jos16] Piotr Josevich. Trigonometric derivatives made easy. College Mathematics Journal, 47(5):365–366, November 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
47.5.365.

Joyner:2014:MWF

[Joy14] David Joyner. The man who found God’s number. College Mathematics Journal, 45(4):258–266, September 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
45.4.258.

Jewett:1988:RWZ

[JR88] Robert I. Jewett and Kenneth A. Ross. Random walks on z. College Mathematics Journal, 19(4):330–342, September 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1988.11973133.

Just:1975:TMP

[JS75] Erwin Just and Norman Schaumberger. Two more proofs of a familiar inequality. Two-Year College Mathematics Journal, 6(2): 45, May 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1975.11974380.

Jerrard:2006:SB

[JSSW06] Richard Jerrard, Joel Schneider, Ralph Smallberg, and John Wetzel. Straw in a Box. College Mathematics Journal, 37(2): 93–102, March 2006. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2006.11922171.

Jones:2014:GNR

[JSW14] Michael A. Jones, Brittany C. Shelton, and Miriam E. Weaverdyck. On God’s number(s) for Rubik’s slide. College Mathematics Journal, 45(4):267–275, September 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
45.4.267.

Jones:2022:ASR

[JSW22] Michael A. Jones, Aidan Shor, and Jennifer Wilson. Allocating scarce resources using a weighted lottery. College Mathematics Journal, 53(4):298–305, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2022.2099706.

Junghans:1974:SF

[Jun74] Helmer Junghans. Square functions. Two-Year College Mathematics Journal, 5(2):15–18, Spring 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1974.
11974308.

Juraschek:1994:ICP

[Jur94] Bill Juraschek. Investigating circles in the Poincaré disk using Geometer’s Sketchpad. College Mathematics Journal, 25(2): 145–154, March 1994. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1994.11973600.

Just:1973:PS

[Jus73] Erwin Just. Problems and solutions. Two-Year College Mathematics Journal, 4(3):81–84, Autumn 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1973.
11974279.

Just:1974:PSa

[Jus74a] Erwin Just. Problems and solutions. Two-Year College Mathematics Journal, 5(2):65–67, Spring 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1974.
11974318.

Just:1974:PSb

[Jus74b] Erwin Just. Problems and solutions. Two-Year College Mathematics Journal, 5(3):50–54, September 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1974.
11974338.

Just:1976:PSa

[Jus76a] Erwin Just. Problems and solutions. Two-Year College Mathematics Journal, 7(1):28–32, February 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1976.
11974418.

Just:1976:PSb

[Jus76b] Erwin Just. Problems and solutions. Two-Year College Mathematics Journal, 7(2):49–53, May 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1976.
11974430.

Just:1979:Pa

[Jus79a] Erwin Just. Problems. Two-Year College Mathematics Journal, 10(2):127–131, March 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1979.11972338.

Just:1979:Pb

[Jus79b] Erwin Just. Problems. Two-Year College Mathematics Journal, 10(3):210–217, June 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1979.11972354.

Just:1979:Pc

[Jus79c] Erwin Just. Problems. Two-Year College Mathematics Journal, 10(4):293–299, September 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1979.11972372.

Just:1979:Pd

[Jus79d] Erwin Just. Problems. Two-Year College Mathematics Journal, 10(5):359–367, November 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1979.11972390.

Just:1980:Pa

[Jus80a] Erwin Just. Problems. Two-Year College Mathematics Journal, 11(1):61–65, January 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1980.11972407.

Just:1980:Pb

[Jus80b] Erwin Just. Problems. Two-Year College Mathematics Journal, 11(2):131–141, March 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1980.11972426.

Just:1980:Pc

[Jus80c] Erwin Just. Problems. Two-Year College Mathematics Journal, 11(3):208–214, June 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1980.11972441.

Just:1980:Pd

[Jus80d] Erwin Just. Problems. Two-Year College Mathematics Journal, 11(4):275–280, September 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1980.11972460.

Just:1980:Pe

[Jus80e] Erwin Just. Problems. Two-Year College Mathematics Journal, 11(5):335–341, November 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1980.11972477.

Just:1981:Pa

[Jus81a] Erwin Just. Problems. Two-Year College Mathematics Journal, 12(1):63–68, January 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1981.11972499.

Just:1981:Pb

[Jus81b] Erwin Just. Problems. Two-Year College Mathematics Journal, 12(2):154–162, March 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1981.11972517.

Just:1981:Pc

[Jus81c] Erwin Just. Problems. Two-Year College Mathematics Journal, 12(3):212–218, June 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1981.11972530.

Just:1981:Pd

[Jus81d] Erwin Just. Problems. Two-Year College Mathematics Journal, 12(4):274–279, September 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1981.11972547.

Just:1981:Pe

[Jus81e] Erwin Just. Problems. Two-Year College Mathematics Journal, 12(5):336–347, November 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1981.11972567.

Just:1982:Pa

[Jus82a] Erwin Just. Problems. Two-Year College Mathematics Journal, 13(1):64–71, January 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1982.11972584.

Just:1982:Pb

[Jus82b] Erwin Just. Problems. Two-Year College Mathematics Journal, 13(2):147–154, March 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1982.11972595.

Just:1982:Pc

[Jus82c] Erwin Just. Problems. Two-Year College Mathematics Journal, 13(3):206–212, June 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1982.11972614.

Just:1982:Pd

[Jus82d] Erwin Just. Problems. Two-Year College Mathematics Journal, 13(4):276–283, September 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1982.11972631.

Just:1982:Pe

[Jus82e] Erwin Just. Problems. Two-Year College Mathematics Journal, 13(5):333–342, November 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1982.11972643.

Justice:1989:EOG

[Jus89] Howard K. Justice. Equality in overlapping gravitational fields. College Mathematics Journal, 20(1):27–31, January 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1989.11973199.

Jakubowski:2015:PWV

[JV15] Matthew Jakubowski and Raymond Viglione. Proof without words: The vertex angle sum of a regular star polygon. College Mathematics Journal, 46(2):109, March 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
46.2.109.

Jerrard:2008:USR

[JW08] Richard P. Jerrard and John E. Wetzel. Universal stoppers are Rupert. College Mathematics Journal, 39(2):90–94, March 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2008.11922280.

Jones:2015:ACS

[JW15] Michael A. Jones and Jennifer Wilson. Adjusting child support payments in Michigan. College Mathematics Journal, 46(1):3–9, January 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.46.1.3.

Jones:2019:TDP

[JW19] Michael A. Jones and Jennifer Wilson. A two-dimensional perspective on Simpson’s Paradox and its likelihood. College Mathematics Journal, 50(4):295–297, 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2019.
1634425.

Johnson:2022:UIV

[JW22] Brian Johnson and Lorna Wenzel. Using the intermediate value theorem to circumscribe hyperbolic triangles. College Mathematics Journal, 53(2):116–121, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2022.2013056.

Jones:2016:RT

[JWW16] Michael A. Jones, Alexander Webb, and Jennifer Wilson. Rankings over time. College Mathematics Journal, 47(4):242–248, September 2016. CODEN ???? ISSN 0746-8342 (print), 1931-

1346 (electronic). URL http://www.tandfonline.com/doi/
abs/10.4169/college.math.j.47.4.242.

Jepsen:1998:MSP

[JY98] Charles Jepsen and Roc Yang. Making squares from Pythagorean triangles. College Mathematics Journal, 29(4):284–288, September 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1998.11973956.

Kaczkowski:2017:MMG

[Kac17] Stephen Kaczkowski. Mathematical models for global mean sea level rise. College Mathematics Journal, 48(3):162–169, May 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.48.3.162.

Kaczkowski:2019:RSA

[Kac19] Stephen Kaczkowski. A Riemann sum approach to Buffon’s needle. College Mathematics Journal, 50(2):93–102, 2019.

CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2019.1560206.

Kaigh:2001:CML

[Kai01] W. D. Kaigh. The case of the missing lottery number. College Mathematics Journal, 32(1):15–19, January 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2001.
11921845.

Kalman:1984:BNS

[Kal84a] Dan Kalman. Basic null space calculations. College Mathematics Journal, 15(1):42–47, January 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1984.
11972743.

Kalman:1984:MMT

[Kal84b] Dan Kalman. The maximum and minimum of two numbers using the quadratic formula. College Mathematics Journal, 15 (4):329–330, September 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1984.11972794.

Kalman:1985:GAC

[Kal85a] Dan Kalman. Geometrically asymptotic curves. College Mathematics Journal, 16(3):199–206, June 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1985.
11972880.

Kalman:1985:RTT

[Kal85b] Dan Kalman. Rediscovering Taylor’s Theorem. College Mathematics Journal, 16(2):103–107, March 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1985.
11972862.

Kalman:1986:APT

[Kal86] Dan Kalman. Angling for Pythagorean triples. College Mathematics Journal, 17(2):167–168, March 1986. CODEN ????

ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1986.
11972949.

Kalman:1987:CRS

[Kal87] Dan Kalman. Cramer’s Rule via selective annihilation. College Mathematics Journal, 18(2):136–137, March 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1987.
11973022.

Kalman:1993:SWS

[Kal93] Dan Kalman. Six ways to sum a series. College Mathematics Journal, 24(5):402–421, November 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1993.
11973562.

Kalman:1996:FCD

[Kal96a] Dan Kalman. Fractions with cycling digit patterns. College Mathematics Journal, 27(2):109–115, March 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1996.
11973760.

Kalman:1996:SVD

[Kal96b] Dan Kalman. A singularly valuable decomposition: The SVD of a matrix. College Mathematics Journal, 27(1):2–23, January 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1996.11973744.

Kalman:1999:SR

[Kal99] Dan Kalman. Software reviews. College Mathematics Journal, 30(5):398–405, November 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1999.11974096.

Kalman:2001:GLE

[Kal01] Dan Kalman. A generalized logarithm for exponential–linear equations. College Mathematics Journal, 32(1):2–14, January 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2001.11921844.

Kalman:2002:ULS

[Kal02] Dan Kalman. An underdetermined linear system for GPS. College Mathematics Journal, 33(5):384–390, November 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2002.11921968.

Kalman:2006:MDE

[Kal06] Dan Kalman. The maximal deflection on an ellipse. College Mathematics Journal, 37(4):250–260, September 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2006.11922192.

Kaput:1985:MH

[Kap85] James J. Kaput. Media highlights. College Mathematics Journal, 16(4):312–322, September 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1985.11972904.

Kaput:1992:MH

[Kap92] James J. Kaput. Media highlights. College Mathematics Journal, 23(3):253–260, May 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1992.11973467.

Kaplan:2017:BR

[Kap17] Craig S. Kaplan. Book review. College Mathematics Journal, 48(1):69–72, January 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.48.1.69.

Karaali:2013:BR

[Kar13] Gizem Karaali. Book review. College Mathematics Journal, 44 (4):332–335, September 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.44.4.332.

Karamzadeh:2014:JTC

[Kar14] O. A. S. Karamzadeh. Johnson’s three circles theorem revisited. College Mathematics Journal, 45(3):217–218, May 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.45.3.217.

Kardos:2022:CCF

[Kar22] Judit Kardos. Constructing continuous functions. College Mathematics Journal, 53(1):21–32, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2022.
1995302.

Katz:1992:MH

[Kat92] Victor J. Katz. Media highlights. College Mathematics Journal, 23(1):77–84, January 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1992.11973440.

Katsuura:1999:EF

[Kat99] Hidefumi Katsuura. From Euler to Fermat. College Mathematics Journal, 30(2):118–119, March 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1999.
11974042.

Katsuura:2003:GAG

[Kat03] Hidefumi Katsuura. Generalizations of the arithmetic–geometric mean inequality and a three dimensional puzzle. College Mathematics Journal, 34(4):280–282, September 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2003.
11922018.

Katsuura:2009:SIB

[Kat09] Hidefumi Katsuura. Summations involving binomial coefficients. College Mathematics Journal, 40(4):275–278, September 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2009.11922375.

Katsuura:2012:EAS

[Kat12] Hidefumi Katsuura. Extending the alternating series test. College Mathematics Journal, 43(4):325–330, September 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
43.4.325.

Katz:2018:BRT

[Kat18] Brian P. Katz. Book review: A TEXas-Style Introduction to Proof by Ron Taylor and Patrick X. Rault, MAA Press, Washington, DC, 2017, xiv + 161 pp., ISBN 978-1-93951-213-0, $50.00 paperback, $50.00 electronic. Discovering Discrete Dynamical Systems by Aimee S. A. Johnson, Kathleen M. Madden, and Ayse A. Sahin, MAA Press, Washington, DC, 2017, xiii + 116 pp., ISBN 978-0-88385-793-9, $50.00 paperback, $50.00 electronic. College Mathematics Journal, 49(5):378–383, November 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2018.1512305.

Kaufman:2020:LLP

[Kau20] Richard Kaufman. Limits on legs of Pythagorean triples and Fermat’s Last Theorem. College Mathematics Journal, 51(1):53– 56, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2020.1674620.

Kelly-Boode:1999:BR

[KB99a] Stan Kelly-Boode. Book review. College Mathematics Journal, 30(3):244–247, May 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1999.11974067.

Knox:1999:NSB

[KB99b] John A. Knox and Harlan J. Brothers. Novel series-based approximations to e. College Mathematics Journal, 30(4):269– 275, September 1999. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1999.11974070.

Kolba:2021:CCT

[KB21] Tiffany N. Kolba and Jonathan Beagley. Counting Christmas trees. College Mathematics Journal, 52(5):338–344, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2021.1967674.

Kennedy:1984:NDN

[KC84] Robert E. Kennedy and Curtis N. Cooper. On the natural density of the Niven numbers. College Mathematics Journal, 15 (4):309–312, September 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1984.11972789.

Kennedy:1987:BUN

[KC87] Robert E. Kennedy and Curtis N. Cooper. Bach, 5465, and upside-down numbers. College Mathematics Journal, 18(2):111– 115, March 1987. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1987.11973018.

Kung:1990:ZRR

[KC90] Sidney H. Kung and Yap S. Chua. A zero-row reduction algorithm for obtaining the gcd of polynomials. College Mathematics Journal, 21(2):138–141, March 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1990.
11973299.

Kerley:1992:UFA

[KD92] Lyndell M. Kerley and William P. Dotson. Using Fourier analysis in digital signal processing. College Mathematics Journal, 23 (4):320–328, September 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1992.11973480.

Klima:1999:AAC

[KD99] Richard E. Klima and Robert G. Donnelly. Amortization: an application of calculus. College Mathematics Journal, 30(5):388– 391, November 1999. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1999.11974092.

Kemp:1982:ISF

[Kem82] Franklin Kemp. Infinite series flow chart for ∑ a_{n}^{†}. Two-Year College Mathematics Journal, 13(3):199, June 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1982.11972611.

Kenschaft:1987:CAS

[Ken87] Patricia C. Kenschaft. Charlotte Angas Scott, 1858–1931. College Mathematics Journal, 18(2):98–110, March 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1987.
11973017.

Kendig:1991:SEC

[Ken91] Keith M. Kendig. Single equations can draw pictures. College Mathematics Journal, 22(2):134–139, March 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1991.
11973370.

Kendig:1999:PSH

[Ken99a] Keith Kendig. Pictures suggest how to improve elementary numerical integration. College Mathematics Journal, 30(1):45–50, January 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1999.11974030.

Kennedy:1999:TLA

[Ken99b] Dan Kennedy. Things I have learned at the AP reading. College Mathematics Journal, 30(5):346–355, November 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1999.11974084.

Kenton:1999:NPC

[Ken99c] Stephen Kenton. A natural proof of the chain rule. College Mathematics Journal, 30(3):216–218, May 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1999.
11974061.

Kenschaft:2000:MMW

[Ken00] Patricia Clark Kenschaft. 455 mathematics majors: What have they done since? College Mathematics Journal, 31(3):193–199, May 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2000.11974142.

Kenschaft:2001:BR

[Ken01] Patricia Clark Kenschaft. Book review. College Mathematics Journal, 32(4):314–316, September 2001. CODEN ????

ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2001.
11921895.

Kennedy:2020:FC

[Ken20] Gary Kennedy. Flattening the curve. College Mathematics Journal, 51(4):254–259, 2020. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2020.1785250.

Kepner:1985:PBT

[Kep85] James L. Kepner. On the probability that the better team wins the world series. College Mathematics Journal, 16(4):250– 256, September 1985. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1985.11972890.

Kersten:1983:CSC

[Ker83] Thomas Kersten. Computer simulations to clarify key ideas of statistics. Two-Year College Mathematics Journal, 14(5):416– 421, November 1983. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1983.11972730.

Krebs:2009:HDI

[KES09] Mike Krebs, Caleb Emmons, and Anthony Shaheen. How to differentiate an integer modulo n. College Mathematics Journal, 40(5):345–353, November 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/074683409X475661.

Key:1994:DSI

[Key94] Eric Key. Disks, shells, and integrals of inverse functions. College Mathematics Journal, 25(2):136–138, March 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1994.11973597.

Koshy:2012:CCS

[KG12] Thomas Koshy and Zhenguang Gao. Convergence of a Catalan series. College Mathematics Journal, 43(2):141–146, March 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.43.2.141.

Kadane:1988:RRJ

[KGSP88] Joseph B. Kadane, William A. Golomski, Daniel A. Sankowsky, and Benjamin M. Perles. Reader’s responses to the January ’88 Forum: Should mathematicians teach statistics? College Mathematics Journal, 19(2):164–165, March 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1988.
11973105.

Kung:1984:VDG

[KH84] Mou-Liang Kung and George C. Harrison. Is the Venn diagram good enough? College Mathematics Journal, 15(1):48–50, January 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972744.

Karabegov:2008:FAS

[KH08] Alexander Karabegov and Jason Holland. Finding all solutions to the magic hexagram. College Mathematics Journal, 39(2): 102–106, March 2008. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2008.11922282.

Khan:2003:CEM

[Kha03] M. A. Khan. A class of exponential matrices. College Mathematics Journal, 34(3):194–195, May 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2003.
11922005.

Khadjavi:2013:APC

[Kha13] Lily Khadjavi. About the pace of climate change: Write a report to the President. College Mathematics Journal, 44(5):428– 432, November 2013. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.44.5.428.

Khovanova:2012:MGM

[Kho12] Tanya Khovanova. Martin Gardner’s mistake. College Mathematics Journal, 43(1):20–24, January 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
43.1.020.

Khovanova:2014:BR

[Kho14] Tanya Khovanova. Book review. College Mathematics Journal, 45(3):230–231, May 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.45.3.230.

Kidwell:1993:MH

[Kid93] Mark Kidwell. Media highlights. College Mathematics Journal, 24(2):191–198, March 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1993.11973528.

Kiernan:1985:CLF

[Kie85] Gerard Kiernan. Computing large factorials. College Mathematics Journal, 16(5):403–412, November 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1985.
11972918.

Kiers:2020:DGE

[Kie20] Claire Kiers. The dynamics of the greenhouse effect. College Mathematics Journal, 51(3):182–194, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2020.
1726703.

Kingston:1974:CPC

[Kin74] J. Maurice Kingston. Cyclically permuted code: a variation on binary arithmetic. Two-Year College Mathematics Journal, 5(1): 29–36, Winter 1974. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1974.11974289.

King:1979:SNCa

[Kin79a] Bruce King. The signpost: News & commentary. Two-Year College Mathematics Journal, 10(1):49–50, January 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1979.11972323.

King:1979:SNCb

[Kin79b] Bruce King. The signpost: News & commentary. Two-Year College Mathematics Journal, 10(2):132–135, March 1979.

CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1979.11972339.

King:1979:SNCc

[Kin79c] Bruce King. The signpost: News & commentary. Two-Year College Mathematics Journal, 10(3):218–220, June 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1979.11972355.

King:1979:SNCd

[Kin79d] Bruce King. The signpost: News & commentary. Two-Year College Mathematics Journal, 10(4):299–301, September 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1979.11972373.

King:1979:SNCe

[Kin79e] Bruce King. The signpost: News & commentary. Two-Year College Mathematics Journal, 10(5):367–370, November 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1979.11972391.

King:1980:SNCa

[Kin80a] Bruce King. The signpost: News & commentary. Two-Year College Mathematics Journal, 11(1):66–67, January 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1980.11972408.

King:1980:SNCd

[Kin80b] Bruce King. The signpost: News & commentary. Two-Year College Mathematics Journal, 11(4):283–285, September 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1980.11972462.

King:1980:SNCe

[Kin80c] Bruce King. The signpost: News & commentary. Two-Year College Mathematics Journal, 11(5):346–347, November 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1980.11972479.

King:1980:SNCb

[Kin80d] Bruce King. The signpost: News commentary. Two-Year College Mathematics Journal, 11(2):142–144, March 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1980.11972427.

King:1980:SNCc

[Kin80e] Bruce King. The signpost: News commentary. Two-Year College Mathematics Journal, 11(3):214–216, June 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1980.
11972442.

King:1981:SNCa

[Kin81a] Bruce King. The signpost: News & commentary. Two-Year College Mathematics Journal, 12(1):68–69, January 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972500.

King:1981:SNCb

[Kin81b] Bruce King. The signpost: News & commentary. Two-Year College Mathematics Journal, 12(2):163–164, March 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972518.

King:1981:SNCc

[Kin81c] Bruce King. The signpost: News & commentary. Two-Year College Mathematics Journal, 12(3):219–221, June 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972531.

King:1983:C

[Kin83] Larry King. The ϵ–δ connection. Two-Year College Mathematics Journal, 14(1):42–47, January 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1983.
11972650.

King:2010:BR

[Kin10] Dan King. Book review. College Mathematics Journal, 41(5): 416–418, November 2010. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/074683410X522053.

King:2013:EMC

[Kin13] L. R. King. An ellipse morphs to a cosine graph! College Mathematics Journal, 44(2):117–123, March 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
44.2.117.

Kipps:1970:WCW

[Kip70] Carol Kipps. Who’s committed? Who’s involved? Two-Year College Mathematics Journal, 1(2):32–35, Autumn 1970. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1970.11973184.

Kirby:1988:ELA

[Kir88] James C. Kirby. An efficient logarithm algorithm for calculators. College Mathematics Journal, 19(3):257–260, May 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1988.11973125.

Kirby:1992:RH

[Kir92] James C. Kirby. Rubberbanding and holding out. College Mathematics Journal, 23(2):148–149, March 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1992.
11973448.

Kirby:1996:BTZ

[Kir96] James C. Kirby. Bargaining theory, or Zeno’s used cars. College Mathematics Journal, 27(4):285–286, September 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1996.11973794.

Kirby:1997:MNR

[Kir97] Kevin G. Kirby. Of memories, neurons, and rank-one corrections. College Mathematics Journal, 28(1):2–19, January 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1997.11973823. This paper contains an exten-

sive tutorial on the Dirac bra and ket notation.

Kittipoom:2019:RVP

[Kit19] Pisamai Kittipoom. A remark on variation of parameters. College Mathematics Journal, 50(3):221–223, 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2019.
1603057.

Koo:2011:P

[KJ11] Reginald Koo and Martin Jones. Probability 1/e. College Mathematics Journal, 42(1):9–13, January 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
42.1.009.

Kraines:1990:SR

[KK90] David P. Kraines and Vivian Y. Kraines. Software reviews. College Mathematics Journal, 21(1):57–64, January 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1990.11973286.

Kerley:1993:CVI

[KK93a] Lyndell Kerley and Jeff Knisley. Complex vectors and image identification. College Mathematics Journal, 24(2):166–174, March 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1993.11973524.

King:1993:D

[KK93b] L. R. King and Benjamin G. Klein. The doors. College Mathematics Journal, 24(3):245–246, May 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1993.
11973539.

Kapranidis:2008:VSL

[KK08] Stelios Kapranidis and Reginald Koo. Variations of the sliding ladder problem. College Mathematics Journal, 39(5):374– 379, November 2008. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2008.11922318.

Kennedy:2009:CP

[KK09] Gary Kennedy and Stephen Kennedy. Crossword puzzle: π_{1} ≈ Z ⊕ Z. College Mathematics Journal, 40(3):212, May 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2009.11922363.

Kennedy:2012:PT

[KK12] Gary Kennedy and Stephen Kennedy. The play’s the thing! College Mathematics Journal, 43(1):70–71, January 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
43.1.70.

Kell:2013:SBB

[KK13a] Nat Kell and Matt Kretchmar. Suspense at the ballot Box. College Mathematics Journal, 44(1):9–16, January 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
44.1.009.

Kose:2013:CMC

[KK13b] Emek Köse and Jennifer Kunze. Climate modeling in the calculus and differential equations classroom. College Mathematics Journal, 44(5):424–427, November 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
44.5.424.

Kungozhin:2013:PWA

[KK13c] Madeubek Kungozhin and Sidney Kung. Proof without words: an algebraic inequality. College Mathematics Journal, 44(1):16, January 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.44.1.016.

Kraines:1989:CCC

[KKS^{+}89] David P. Kraines, Vivian Y. Kraines, David A. Smith, David P. Kraines, Vivian Y. Kraines, David A. Smith, David P. Kraines, Vivian Y. Kraines, and David A. Smith. Classroom computer capsules. College Mathematics Journal, 20(5):433–438, November 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1989.11973272.

Kraines:1990:CCCc

[KKS90a] David P. Kraines, Vivian Y. Kraines, and David A. Smith. Classroom computer capsule: The Cauchy integral formula. College Mathematics Journal, 21(4):327–329, September 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1990.11973329.

Kraines:1990:CCCb

[KKS90b] David P. Kraines, Vivian Y. Kraines, and David A. Smith. Classroom computer capsules: Binary operations. College Mathematics Journal, 21(3):240–241, May 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1990.
11973314.

Kraines:1990:CCCa

[KKS^{+}90c] David P. Kraines, Vivian Y. Kraines, David A. Smith, David P. Kraines, Vivian Y. Kraines, and David A. Smith. Classroom computer capsules. College Mathematics Journal, 21(2):142–145, March 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1990.11973300.

Kraines:1990:CCCd

[KKS^{+}90d] David P. Kraines, Vivian Y. Kraines, David A. Smith, David P. Kraines, Vivian Y. Kraines, David A. Smith, David P. Kraines, Vivian Y. Kraines, and David A. Smith. Classroom computer capsules. College Mathematics Journal, 21(5):415–419, November 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1990.11973347.

Kraines:1991:CCCa

[KKS91] David P. Kraines, Vivian Y. Kraines, and David A. Smith. Classroom computer capsule: Predator-prey model. College Mathematics Journal, 22(2):160–162, March 1991. CODEN ????

ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1991.
11973376.

Klyve:2011:EAS

[KL11] Dominic Klyve and Anna Lauren. An empirical approach to the St. Petersburg Paradox. College Mathematics Journal, 42 (4):260–264, September 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.42.4.260.

Kunkle:1987:SR

[KLB^{+}87] Dan Kunkle, L. Carl Leinbach, Judith R. Brown, Thomas J. Abram, and Carol M. Smith. Software reviews. College Mathematics Journal, 18(1):66–68, January 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1987.
11973013.

Klee:1971:SF

[Kle71] Victor Klee. Shapes of the future. Two-Year College Mathematics Journal, 2(2):14–27, Autumn 1971. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1971.
11974008.

Kleiman:1972:PTG

[Kle72] Steven L. Kleiman. Pythagorean triples by geometry. Two-Year College Mathematics Journal, 3(1):39–41, Spring 1972. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1972.11974195.

Kleitman:1984:CDL

[Kle84] Daniel J. Kleitman. Calculus defended …. College Mathematics Journal, 15(5):377–378, November 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1984.
11972808.

Kleiner:1989:EFC

[Kle89] Israel Kleiner. Evolution of the function concept: a brief survey. College Mathematics Journal, 20(4):282–300, September 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1989.11973245.

Kletzing:1991:CPR

[Kle91] Dennis Kletzing. Coset products in rings. College Mathematics Journal, 22(4):323–326, September 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1991.
11973407.

Klyve:2015:BR

[Kly15] Dominic Klyve. Book review. College Mathematics Journal, 46(1):67–72, January 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.46.1.67.

Klyve:2019:LE

[Kly19] Dominic Klyve. Letter from the Editor. College Mathematics Journal, 50(1):2, January 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2019.1557492.

Klyve:2020:YCF

[Kly20] Dominic Klyve. 50 years of CMJ: The first decade. College Mathematics Journal, 51(1):51–52, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2020.
1681788.

Klamkin:1989:GCN

[KM89] M. S. Klamkin and V. N. Murty. Generalizations of a complex number identity. College Mathematics Journal, 20(5):415– 416, November 1989. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1989.11973268.

Krebs:2013:CTM

[KM13] Mike Krebs and Natalie Martinez. The combinatorial trace method in action. College Mathematics Journal, 44(1):32–36, January 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.44.1.032.

Knapper:2015:PWL

[KM15] Elizabeth Knapper and Hope McIlwain. Predicting wins and losses: a volleyball case study. College Mathematics Journal, 46(5):352–358, November 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.46.5.352.

Kantrowitz:2015:AFA

[KN15] Robert Kantrowitz and Michael Neumann. Another face of the Archimedean property. College Mathematics Journal, 46(2): 139–141, March 2015. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.46.2.139.

Knapp:1985:ISP

[Kna85] Thomas R. Knapp. Instances of Simpson’s Paradox. College Mathematics Journal, 16(3):209–211, June 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1985.
11972882.

Knop:1985:CDT

[Kno85] Larry E. Knop. Cantor’s disappearing table. College Mathematics Journal, 16(5):398–399, November 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1985.
11972915.

Knuth:1991:KSN

[Knu91] Donald E. Knuth. Kunusu Sensei no Program-Ron [Professor Knuth’s Programming Discipline]. Kyoritsu-Shuppan, Tokyo, Japan, 1991. ISBN ???? v + 199 pp. LCCN ???? Anthology edited by Makoto Arisawa.

Knudson:2018:CAG

[Knu18] Kevin P. Knudson. Cutting against the grain: Volumes of solids of revolution via cross-sections parallel to the rotation axis. College Mathematics Journal, 49(2):114–120, March 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2018.1411656.

Kleinz:2000:CGF

[KO00] Marcia Kleinz and Thomas J. Osler. A child’s garden of fractional derivatives. College Mathematics Journal, 31(2):82–88, March 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2000.11974118.

Kobayashi:1998:GVV

[Kob98] Yukio Kobayashi. A geometric view of a vector identity. College Mathematics Journal, 29(4):309–310, September 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1998.11973962.

Kobayashi:2013:PWTa

[Kob13a] Yukio Kobayashi. Proof without words: Tangent double angle identity. College Mathematics Journal, 44(1):47, January 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.44.1.047.

Kobayashi:2013:PWTb

[Kob13b] Yukio Kobayashi. Proof without words: Triangular sums. College Mathematics Journal, 44(3):189, May 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
44.3.189.

Kobayashi:2014:PWC

[Kob14] Yukio Kobayashi. Proof without words: Componendo et dividendo, a theorem on proportions. College Mathematics Journal, 45(2):115, March 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.45.2.115.

Kobayashi:2017:PWS

[Kob17] Yukio Kobayashi. Proof without words: a sum computed by self-similarity. College Mathematics Journal, 49(1):10, January 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2017.1389563.

Kohl:2012:WAH

[Koh12] Timothy Kohl. When Abelian = Hausdorff. College Mathematics Journal, 43(3):213–215, May 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
43.3.213.

Kolata:1981:PRG

[Kol81] Gina Bari Kolata. A profile of Ronald L. Graham. Two-Year College Mathematics Journal, 12(5):290–301, November 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972553.

Korte:1985:AMV

[Kor85] Bernhard Korte. Algorithmic mathematics versus dialectic mathematics. College Mathematics Journal, 16(1):6–8, January 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1985.11972840.

Koshy:2009:LGC

[Kos09] Thomas Koshy. Lobb’s generalization of Catalan’s parenthesization problem. College Mathematics Journal, 40(2):99–107, March 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2009.11922344.

Kossek:2017:TSG

[Kos17] Wojciech Kossek. Is a Taylor series also a generalized Fourier series? College Mathematics Journal, 49(1):54–56, January 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2017.1397449.

Koshkin:2020:BHA

[Kos20] Sergiy Koshkin. Bisecting Horn angles. College Mathematics Journal, 51(2):124–131, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2020.1700745.

Kouba:1995:CWU

[Kou95] Duane Kouba. Can we use the first derivative to determine inflection points? College Mathematics Journal, 26(1):31–34, January 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1995.11973663.

Kowalski:2008:HMA

[Kow08] Travis Kowalski. How to measure angles with a ruler. College Mathematics Journal, 39(4):273–279, September 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2008.11922304.

Kowalski:2016:SSD

[Kow16] Travis Kowalski. The sine of a single degree. College Mathematics Journal, 47(5):322–332, November 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
47.5.322.

Kowalski:2024:PPS

[Kow24] R. Travis Kowalski. On a particular particular solution. College Mathematics Journal, 55(3):215–225, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2024.
2301900.

Kalman:1985:NPE

[KP85] Dan Kalman and Warren Page. Nested polynomials and efficient exponentiation algorithms for calculators. College Mathematics Journal, 16(1):57–60, January 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1985.
11972854.

Kunoff:1986:TEP

[KP86] Sharon Kunoff and Sylvia Pines. Teaching elementary probability through its history. College Mathematics Journal, 17(3):210–219, May 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1986.11972958.

King:1991:GGS

[KR91] L. R. King and Harold B. Reiter. Graceful graphs and sparsely marked rulers. College Mathematics Journal, 22(3):232–234, May 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1991.11973389.

Krikorian:1996:CSS

[KR96] Nishan Krikorian and Mark Ramras. Cubic splines from Simpson’s Rule. College Mathematics Journal, 27(2):124–126, March 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1996.11973764.

Klebanoff:1998:SCD

[KR98] Aaron Klebanoff and John Rickert. Studying the Cantor dust at the edge of Feigenbaum diagrams. College Mathematics Journal, 29(3):189–198, May 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1998.11973939.

Krylov:2011:DFP

[KR11] Nikolai Krylov and Edwin L. Rogers. Dynamics of folds in the plane. College Mathematics Journal, 42(2):108–114, March 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.42.2.108.

Kraines:1991:CCCb

[Kra91] David P. Kraines. Classroom computer capsules: Rotations in space and orthogonal matrices. College Mathematics Journal, 22 (3):245–247, May 1991. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1991.11973392.

Krantz:2001:BR

[Kra01] Steven G. Krantz. Book review. College Mathematics Journal, 32(3):232–237, May 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2001.11921881.

Krasopoulos:2003:TPQ

[Kra03] Panagiotis T. Krasopoulos. Tangent planes of a quadratic function. College Mathematics Journal, 34(3):205–206, May 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2003.11922007.

Kreminski:1997:USR

[Kre97] Rick Kreminski. Using Simpson’s Rule to approximate sums of infinite series. College Mathematics Journal, 28(5):368–376, November 1997. CODEN ???? ISSN 0746-8342 (print), 1931-

1346 (electronic). URL http://www.tandfonline.com/doi/
abs/10.1080/07468342.1997.11973892.

Kreminski:2010:TTE

[Kre10] Rick Kreminski. Taylor’s Theorem: The elusive c is not so elusive. College Mathematics Journal, 41(3):186–192, May 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/074683410X488656.

Krebs:2017:CST

[Kre17] Mike Krebs. Can a subset’s topology detect continuous extensions? College Mathematics Journal, 49(2):138–139, March 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2017.1397450.

Krishnapriyan:1995:EPF

[Kri95] H. K. Krishnapriyan. Eulerian polynomials and Faulhaber’s result on sums of powers of integers. College Mathematics Journal, 26(2):118–123, March 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1995.11973680.

Kristensen:2025:MIA

[Kri25] Kai Forsberg Kristensen. A multiple integration approach to the broken stick problem with predetermined minimum piece lengths. College Mathematics Journal, 56(1):25–35, 2025. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic).

Kroopnick:1979:HGR

[Kro79] Allen Kroopnick. How good is the “rule of 72”? Two-Year College Mathematics Journal, 10(4):279–280, September 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1979.11972368.

Kroopnick:1982:NPC

[Kro82] Allan J. Kroopnick. A note on parallel curves. Two-Year College Mathematics Journal, 13(1):59–61, January 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1982.11972581.

Krom:1983:SPN

[Kro83] Myren Krom. On sums of powers of natural numbers. Two-Year College Mathematics Journal, 14(4):349–351, September 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1983.11972714.

Kronholm:2016:IWP

[Kro16] William Kronholm. Integration by the wrong parts. College Mathematics Journal, 47(2):102–105, March 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
47.2.102.

Khotmongkon:2018:PWR

[KRSW18] Poontarika Khotmongkon, Nahathai Rerkruthairat, Sujittra Suriwong, and Kornkanok Watcharakarn. The probability of winning a racquetball game with Deuce. College Mathematics Journal, 49(5):353–358, November 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2018.
1526018.

Krusemeyer:2000:ECT

[Kru00] Mark Krusemeyer. An “extremely” cautionary tale. College Mathematics Journal, 31(2):128–130, March 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2000.
11974129.

Kraines:1988:CCT

[KS88] David P. Kraines and David A. Smith. A computer in the classroom: The time is right. College Mathematics Journal, 19(3): 261–267, May 1988. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1988.11973126.

Klerlein:1996:GMO

[KS96] Joe Klerlein and Scott Sportsman. Generalizations of a mathematical Olympiad problem. College Mathematics Journal, 27 (4):296–297, September 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1996.11973798.

Klein:2003:TSW

[KS03] Jennifer C. Klein and Thomas Q. Sibley. Taking the sting out of wasp nests: a dialogue on modeling in mathematical biology. College Mathematics Journal, 34(3):207–215, May 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2003.11922008.

Klyve:2006:GLS

[KS06] Dominic Klyve and Lee Stemkoski. Graeco-Latin squares and a mistaken conjecture of Euler. College Mathematics Journal, 37 (1):2–15, January 2006. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2006.11922160.

Kantrowitz:2012:SCA

[KS12] Robert Kantrowitz and Michael Schramm. Series that converge absolutely but don’t converge. College Mathematics Journal, 43 (4):331–333, September 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.43.4.331.

Komornik:2024:SIE

[KS24] Vilmos Komornik and Reinhard Schäfke. A simple introduction to the exponential function. College Mathematics Journal, 55(2):165–168, 2024. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2023.2234256.

Kuczmarski:2016:RBC

[Kuc16] Fred Kuczmarski. Rubber band calculus. College Mathematics Journal, 47(2):82–93, March 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.47.2.82.

Kung:1970:X

[Kun70] H. L. Kung. ∫ f(x)e^{ax} dx. Two-Year College Mathematics Journal, 1(2):106–107, Autumn 1970. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1970.11973195.

Kung:1972:TAP

[Kun72] Sidney H. L. Kung. Talking about particular solutions. Two-Year College Mathematics Journal, 3(1):67–71, Spring 1972. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1972.11974201.

Kung:1973:PS

[Kun73] H. L. Kung. On particular solutions of P_{n}(D)Y = 0. Two-Year College Mathematics Journal, 4(1):14–17, Winter 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1973.11974229.

Kung:1974:AEA

[Kun74] Sidney H. L. Kung. An alternative to Euclidean algorithm. Two-Year College Mathematics Journal, 5(2):8–11, Spring 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1974.11974306.

Kundert:1980:CVA

[Kun80a] Kenneth R. Kundert. Correlation — a vector approach. Two-Year College Mathematics Journal, 11(1):52, January 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1980.11972402.

Kundert:1980:WTS

[Kun80b] Kenneth R. Kundert. Why not teach synthetic multiplication? Two-Year College Mathematics Journal, 11(2):121–122, March 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1980.11972420.

Kung:1988:NAC

[Kun88] Sidney H. Kung. A nonstandard approach to Cramer’s Rule. College Mathematics Journal, 19(1):59–60, January 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1988.11973093.

Kundert:1989:SQR

[Kun89] Kenneth R. Kundert. Sensitive questions and randomized response techniques. College Mathematics Journal, 20(5):409– 411, November 1989. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1989.11973266.

Kung:1998:MC

[Kun98] Sidney H. Kung. More coconuts. College Mathematics Journal, 29(4):312–313, September 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1998.11973964.

Kung:2011:ASR

[Kun11] Sidney H. Kung. An application of Sylvester’s rank inequality. College Mathematics Journal, 42(2):148–150, March 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.42.2.148.

Kutz:1981:TUS

[Kut81] Ronald E. Kutz. Two unusual sequences. Two-Year College Mathematics Journal, 12(5):316–319, November 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972558.

Kalman:1998:SSC

[KW98] Dan Kalman and James White. A simple solution of the cubic. College Mathematics Journal, 29(5):415–418, November 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1998.11973980.

Kwon:2016:PWL

[Kwo16] Miyeon Kwon. Proof without words: The lateral surface area of a conical frustum. College Mathematics Journal, 47(5):346, November 2016. CODEN ???? ISSN 0746-8342 (print), 1931-

1346 (electronic). URL http://www.tandfonline.com/doi/
abs/10.4169/college.math.j.47.5.346.

Khovanova:2015:CMP

[KX15] Tanya Khovanova and Joshua Xiong. Cookie Monster plays games. College Mathematics Journal, 46(4):283–293, September 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.46.4.283.

Laible:1970:NGD

[Lai70] Jon M. Laible. A new graduate degree for mathematics teachers in two-year colleges. Two-Year College Mathematics Journal, 1(1):55–58, Spring 1970. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1970.11973169.

Lakins:2024:EE

[Lak24a] Tamara Lakins. Editor’s endnotes. College Mathematics Journal, 55(5):460–461, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic).

Lakins:2024:LEa

[Lak24b] Tamara Lakins. Letter from the Editor. College Mathematics Journal, 55(1):3, 2024. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2024.2285197.

Lakins:2024:LEd

[Lak24c] Tamara Lakins. Letter from the Editor. College Mathematics Journal, 55(4):279, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2024.2374220.

Lakins:2024:LE

[Lak24d] Tamara Lakins. Letter from the Editor. College Mathematics Journal, 55(5):375, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic).

Lakins:2024:LEc

[Lak24e] Tami Lakins. Letter from the Editor. College Mathematics Journal, 55(3):191, 2024. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2024.2334663.

Lakins:2025:EE

[Lak25a] Tamara Lakins. Editor’s endnotes. College Mathematics Journal, 56(1):92, 2025. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic).

Lakins:2025:LE

[Lak25b] Tamara Lakins. Letter from the Editor. College Mathematics Journal, 56(1):3, 2025. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic).

Lamberson:1977:EP

[Lam77] Roland H. Lamberson. An environmental problem. Two-Year College Mathematics Journal, 8(4):252–253, September 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1977.11974521.

Lamphere:1996:RRS

[Lam96] Robert L. Lamphere. A recurrence relation in the spinout puzzle. College Mathematics Journal, 27(4):286–289, September 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1996.11973795.

Lamphere:2024:HCV

[Lam24] Robert L. Lamphere. How to compute the volumes of hyperbolic solids of revolution. College Mathematics Journal, 55(2):140– 152, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2023.2276651.

Lange:1972:LMI

[Lan72] Lester H. Lange. A look at that 1971 MAA information services survey. Two-Year College Mathematics Journal, 3(2):56–69, Autumn 1972. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1972.11974218.

Lange:1974:M

[Lan74] L. H. Lange. Maximize x(a − x). Two-Year College Mathematics Journal, 5(1):22–24, Winter 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1974.11974287.

Lange:1976:SHE

[Lan76] L. H. Lange. Several hyperbolic encounters. Two-Year College Mathematics Journal, 7(1):2–6, February 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/00494925.
1976.11974410.

Landauer:1977:CCM

[Lan77] Edwin G. Landauer. Computer corner: Methods of random number generation. Two-Year College Mathematics Journal, 8 (5):296–303, November 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.jstor.org/

stable/3026786;   http://www.tandfonline.com/doi/abs/
10.1080/00494925.1977.11974532.

Lange:1980:CMH

[Lan80] Lester H. Lange. I.C.M. Helsinki — some personal notes. Two-Year College Mathematics Journal, 11(2):144–146, March 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1980.11972428.

Lange:1983:BR

[Lan83] Lester H. Lange. Book reviews. Two-Year College Mathematics Journal, 14(5):415, November 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1983.
11972729.

Lange:1993:MH

[Lan93a] Les Lange. Media highlights. College Mathematics Journal, 24 (3):277–284, May 1993. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1993.11973545.

Lange:1993:WBR

[Lan93b] Lester H. Lange. What is the biggest rectangle you can put inside a given triangle? College Mathematics Journal, 24(3):237–240, May 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1993.11973536.

Lange:1995:DPG

[Lan95] Lester H. Lange. Did Plutarch get Archimedes’ wishes right? College Mathematics Journal, 26(3):199–204, May 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1995.11973696.

Larson:1974:CDP

[Lar74] Roland E. Larson. Continuous deformation of a polynomial into its derivatives. Two-Year College Mathematics Journal, 5(2): 68–70, Spring 1974. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1974.11974319.

Larson:1975:GCU

[Lar75] Roland E. Larson. Graphing a cubic using calculus and a computer. Two-Year College Mathematics Journal, 6(1):35–40, February 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1975.11974366.

Larson:1985:DLL

[Lar85] Loren C. Larson. A discrete look at 1 + 2 + … + n. College Mathematics Journal, 16(5):369–382, November 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1985.11972910.

Larsen:1989:ETH

[Lar89] Mogens Esrom Larsen. The eternal triangle — a history of a counting problem. College Mathematics Journal, 20(5):370– 384, November 1989. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1989.11973260.

Lathan:1973:R

[Lat73] Calvin A. Lathan. Reviews. Two-Year College Mathematics Journal, 4(1):60–69, Winter 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1973.11974238.

Lathan:1975:Ea

[Lat75a] Calvin Lathan. Editorial. Two-Year College Mathematics Journal, 6(2):1, May 1975. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1975.11974368.

Lathan:1975:Eb

[Lat75b] Calvin A. Lathan. Editorial. Two-Year College Mathematics Journal, 6(4):1, December 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1975.11974396.

Lathan:1976:E

[Lat76] Calvin Lathan. Editorial. Two-Year College Mathematics Journal, 7(1):1, February 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1976.11974408.

Lathan:1977:E

[Lat77] Calvin A. Lathan. Editorial. Two-Year College Mathematics Journal, 8(1):3, January 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1977.11974469.

Latina:1983:SMR

[Lat83] Michael R. Latina. Some maximal rectangles and the realities of applied mathematics. Two-Year College Mathematics Journal, 14(3):248–252, June 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1983.11972690.

LaTorre:1993:IMI

[LaT93] Donald R. LaTorre. Iterative methods in introductory linear algebra. College Mathematics Journal, 24(1):79–88, January 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1993.11973513.

Laudano:2020:PWM

[Lau20a] Francesco Laudano. Proof without words: Magic of tangential polygons. College Mathematics Journal, 51(3):218, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2020.1704131.

Laumakis:2020:APC

[Lau20b] Paul Laumakis. Analyzing proportionality coefficients in differential equation models. College Mathematics Journal, 51(5):360– 368, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2020.1820788.

Laudano:2025:NRR

[Lau25] Francesco Laudano. On the number of rational roots of a rational polynomial. College Mathematics Journal, 56(1):36–41, 2025. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic).

Lawrisuk:1971:CCH

[Law71] Paul Lawrisuk. Committing curricular heresy. Two-Year College Mathematics Journal, 2(1):58–64, Spring 1971. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/00494925.
1971.11974000.

Lawlor:2021:EIR

[Law21] Gary R. Lawlor. An eigenargument for irrational roots. College Mathematics Journal, 52(2):140–141, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2021.
1868926.

Lawlor:2024:PTN

[Law24] Gary R. Lawlor. The product of two natural numbers equals the product of their GCD and LCM. College Mathematics Journal, 55(1):62–63, 2024. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2023.2231311.

Lax:1976:LAP

[Lax76] Anneli Lax. Linear algebra, a potent tool. Two-Year College Mathematics Journal, 7(2):3–15, May 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1976.
11974423.

Lax:1984:PCL

[Lax84] Peter D. Lax. In praise of calculus …. College Mathematics Journal, 15(5):378–380, November 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1984.
11972809.

Lax:1998:FFF

[Lax98] Peter D. Lax. Fallacies, flaws, and flimflam. College Mathematics Journal, 29(3):219–221, May 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1998.
11973943.

Lay:1993:SEF

[Lay93] David C. Lay. Subspaces and echelon forms. College Mathematics Journal, 24(1):57–62, January 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1993.
11973507.

Lindstrom:1980:BRb

[LB80] Peter A. Lindstrom and Leon Bankoff. Book reviews. Two-Year College Mathematics Journal, 11(2):127–128, March 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1980.11972424.

Lenard:1994:KOM

[LB94] Andrew Lenard and Michael K. Brozinsky. Kepler orbits more geometrico. College Mathematics Journal, 25(2):90–98, March 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1994.11973588.

Loe:1997:CCC

[LB97] Brian J. Loe and Nathaniel Beagley. The coffee cup caustic for calculus students. College Mathematics Journal, 28(4):277– 284, September 1997. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1997.11973875.

Lindstrom:1980:BRc

[LC80] Peter A. Lindstrom and Sarah Christiansen. Book reviews. Two-Year College Mathematics Journal, 11(3):218–221, June 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1980.11972444.

Lehet:2023:NMC

[LC23] Ellen Lehet and Kuo-Liang Chang. A novel method for completing the square. College Mathematics Journal, 54(5):476– 480, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2023.2273195.

Lowry-Duda:2015:UCA

[LD15] David Lowry-Duda. Unexpected conjectures about −5 modulo primes. College Mathematics Journal, 46(1):56–57, January 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.46.1.56.

Leise:2022:MH

[LE22] Tanya Leise and Philip Straffin (Section Editor). Media highlights. College Mathematics Journal, 53(5):408–415, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2022.2127286.

Leavitt:1973:FN

[Lea73] W. G. Leavitt. Fermat numbers. Two-Year College Mathematics Journal, 4(3):7–10, Autumn 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1973.11974267.

Leavitt:1976:ACA

[Lea76] W. G. Leavitt. An algorithm and its connection with Abelian groups. Two-Year College Mathematics Journal, 7(2):16–21, May 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1976.11974424.

Leavitt:1984:RD

[Lea84] W. G. Leavitt. Repeating decimals. College Mathematics Journal, 15(4):299–308, September 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1984.
11972788.

Leach:2016:PWP

[Lea16a] C. David Leach. Proof without words: Powers of three and triangular numbers. College Mathematics Journal, 47(2):120, March 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.47.2.120.

Leahy:2016:WNC

[Lea16b] Andrew Leahy. William Neile’s contribution to calculus. College Mathematics Journal, 47(1):42–49, January 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
47.1.42.

Leahy:2018:FCT

[Lea18] Andrew Leahy. A fourth century theorem for twenty-first century calculus. College Mathematics Journal, 49(2):103–108, March 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2018.1421817.

LeCuyer:1986:TMU

[LeC86] Edward J. LeCuyer, Jr. Teaching mathematics using APL. College Mathematics Journal, 17(4):344–358, September 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1986.11972980.

LeDuc:1973:FSC

[LeD73] John W. LeDuc. δ as a function of ϵ, a suggestion for the calculus teacher. Two-Year College Mathematics Journal, 4(3):85–86, Autumn 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1973.11974280.

Lederman:1977:CCB

[Led77] Barbara J. Lederman. The community College basic mathematics course. Two-Year College Mathematics Journal, 8(1):29–34, January 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1977.11974477.

Ledder:2016:EMC

[Led16] Glenn Ledder. Empirical modeling: Choosing models and fitting them to data. College Mathematics Journal, 47(2):109–119, March 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.47.2.109.

Lee:1998:PS

[Lee98] Cheng-Shyong Lee. Polishing the star. College Mathematics Journal, 29(2):144–145, March 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1998.
11973930.

Lee:1999:FRF

[Lee99] Kil S. Lee. A far-reaching formula. College Mathematics Journal, 30(2):138–140, March 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1999.11974047.

Lee:2024:SPM

[Lee24] Kang-Ju Lee. Squares 1^{2}, 2^{2}, …, ((p − 1)/2)^{2} modulo a prime p. College Mathematics Journal, 55(5):430–431, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic).

Lefton:1983:CCP

[Lef83] Phyllis Lefton. Congruences of cyclotomic polynomials. Two-Year College Mathematics Journal, 14(3):257–258, June 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1983.11972693.

Lehning:1990:CAA

[Leh90] Hervé Lehning. Computer-aided or analytic proof? College Mathematics Journal, 21(3):228–239, May 1990. CODEN ????

ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1990.
11973313.

Leher:2022:TDT

[Leh22] Eli Leher. On the two “Dimension theorems”. College Mathematics Journal, 53(2):151, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2022.2020589.

Leinbach:1987:SR

[LEHE87] L. Carl Leinbach, Michael W. Ecker, Rosemary Hirschfelder, and Sheryl E. Edwards. Software reviews. College Mathematics Journal, 18(2):157–159, March 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1987.11973028.

Leitzel:1983:TCU

[Lei83] Joan R. Leitzel. Toward a common understanding of the content of College preparatory mathematics. Two-Year College Mathematics Journal, 14(3):206–209, June 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1983.
11972684.

Leinbach:1990:SR

[Lei90] L. Carl Leinbach. Software reviews. College Mathematics Journal, 21(5):420–422, November 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1990.
11973348.

Leinbach:1993:SR

[Lei93] L. Carl Leinbach. Software reviews. College Mathematics Journal, 24(3):263–270, May 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1993.11973543.

Leinbach:1995:SR

[Lei95] L. Carl Leinbach. Software reviews. College Mathematics Journal, 26(2):152–156, March 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1995.11973689.

Leinbach:1996:SR

[Lei96] L. Carl Leinbach. Software reviews. College Mathematics Journal, 27(5):398–403, November 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1996.
11973818.

Leinbach:1997:SR

[Lei97] Carl Leinbach. Software reviews. College Mathematics Journal, 28(3):224–230, May 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1997.11973869.

Leise:2007:PWT

[Lei07] Tanya Leise. As the planimeter’s wheel turns: Planimeter proofs for calculus class. College Mathematics Journal, 38(1):24–31, January 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2007.11922214.

Leise:2021:MHc

[Lei21] Tanya Leise. Media highlights. College Mathematics Journal, 52(3):233–240, 2021. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2021.1909995.

Lemmermeyer:2016:LVP

[Lem16] Franz Lemmermeyer. Leonardo da Vinci’s proof of the Pythagorean Theorem. College Mathematics Journal, 47(5):361– 364, November 2016. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.47.5.361.

Lenz:1974:NAM

[Len74] G. Lenz. The neglected aspect of math-ed: The liberal arts course. Two-Year College Mathematics Journal, 5(1):7–8, Winter 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1974.11974285.

Lenz:1976:HMC

[Len76] Gerald E. Lenz. History in the mathematics curriculum. Two-Year College Mathematics Journal, 7(3):27–28, September 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1976.11974442.

Lenz:1981:SSS

[Len81] Gerald Lenz. Some sum of sums. Two-Year College Mathematics Journal, 12(3):208–209, June 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1981.
11972528.

Lengvarszky:2015:PPT

[Len15] Zsolt Lengvárszky. Proving the Pythagorean Theorem by letting the sides vary. College Mathematics Journal, 46(1):52–55, January 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.46.1.52.

Lerma:2021:TVH

[Ler21] Miguel A. Lerma. Truck versus human 2.0: Mathematical follow-up under increasing pressure, and how Kepler’s Laws come to the rescue. College Mathematics Journal, 52(1):22– 30, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2021.1847590.

Lesko:2004:SHT

[Les04] James Lesko. Sums of harmonic-type series. College Mathematics Journal, 35(3):171–182, May 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2004.
11922070.

Levin:1981:CC

[Lev81] Eugene M. Levin. Computers & calculators. Two-Year College Mathematics Journal, 12(4):260–266, September 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972541.

Levin:2014:SP

[Lev14] Oscar Levin. Story puzzles. College Mathematics Journal, 45(4): 296, September 2014. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.45.4.296.

Lewis:1973:TII

[Lew73] Frances W. Lewis. Teaching inequalities involving absolute values. Two-Year College Mathematics Journal, 4(2):87–90, Spring 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1973.11974261.

Lewin:1984:AAE

[Lew84] Jonathan W. Lewin. An analytic approach to the Euler line. College Mathematics Journal, 15(1):52–53, January 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972746.

Lewis:1986:DAP

[Lew86] Frances W. Lewis. Defining area in polar coordinates. College Mathematics Journal, 17(5):414–416, November 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1986.11972990.

Lewis:1995:MPS

[Lew95] Jerome L. Lewis. The meeting of the plows: a simulation. College Mathematics Journal, 26(5):395–400, November 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1995.11973738.

Lewin:2002:SR

[Lew02] Jonathan Lewin. Software reviews. College Mathematics Journal, 33(5):426–430, November 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2002.11921974.

Lindstrom:1979:BRb

[LF79] Peter A. Lindstrom and Donald C. Fuller. Book reviews. Two-Year College Mathematics Journal, 10(2):123–126, March 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1979.11972337.

Lindstrom:1980:BRa

[LG80] Peter A. Lindstrom and Klaus Galda. Book reviews. Two-Year College Mathematics Journal, 11(1):56–58, January 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1980.11972405.

Lindstrom:1981:BR

[LGA81] Peter A. Lindstrom, Marilyn Mays Gilchrist, and Henry Africk. Book reviews. Two-Year College Mathematics Journal, 12(1): 58–61, January 1981. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1981.11972497.

Lindstrom:1980:BRd

[LH80] Peter A. Lindstrom and S. S. Holland, Jr. Book reviews. Two-Year College Mathematics Journal, 11(5):341–345, November 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1980.11972478.

Leinbach:1986:SR

[LH86] L. Carl Leinbach and Rosemary Hirschfelder. Software reviews. College Mathematics Journal, 17(1):90–91, January 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1986.11972938.

Long:1989:GCZ

[LH89] Cliff Long and Thomas Hern. Graphing the complex zeros of polynomials using modulus surfaces. College Mathematics Journal, 20(2):98–105, March 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1989.11973215.

Leinbach:1992:SR

[LH92] L. Carl Leinbach and Edward A. Huff. Software reviews. College Mathematics Journal, 23(5):429–434, November 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1992.11973496.

Leinbach:1988:SRb

[LHL88] L. Carl Leinbach, Thomas Hannula, and Daniel G. Lamet. Software reviews. College Mathematics Journal, 19(4):367–369, September 1988. CODEN ???? ISSN 0746-8342 (print), 1931-

1346 (electronic). URL http://www.tandfonline.com/doi/
abs/10.1080/07468342.1988.11973141.

Li:2013:PWMa

[Li13a] Xiaoxue H. Li. Proof without words: Monotonicity of sin x/x on (0, π/2). College Mathematics Journal, 44(5):408, November 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.44.5.408.

Li:2013:PWMb

[Li13b] Xiaoxue H. Li. Proof without words: Monotonicity of tan x/x on (0, π/2). College Mathematics Journal, 44(5):420, November 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.44.5.420.

Lial:1970:PIE

[Lia70] Margaret L. Lial. Programmed instruction in elementary algebra: an experiment. Two-Year College Mathematics Journal, 1(2):17–21, Autumn 1970. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1970.11973181.

Libeskind:1981:SFS

[Lib81] Shlomo Libeskind. Summation of finite series — a unified approach. Two-Year College Mathematics Journal, 12(1):41–50, January 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972491.

Lichtman:1981:CC

[Lic81] Denis R. Lichtman. Computers & calculators. Two-Year College Mathematics Journal, 12(3):199–203, June 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1981.
11972524.

Liebars:1997:TLE

[Lie97] Cathy Liebars. Tying up loose ends with probability. College Mathematics Journal, 28(5):386–388, November 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1997.11973897.

Lindstrom:1970:ICN

[Lin70] Peter A. Lindstrom. The irrationality of certain numbers. Two-Year College Mathematics Journal, 1(1):30–31, Spring 1970. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1970.11973163.

Lindley:1974:IP

[Lin74a] Roger G. Lindley. The interpolating polynomial. Two-Year College Mathematics Journal, 5(2):21–31, Spring 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1974.11974310.

Lindstrom:1974:NED

[Lin74b] Peter A. Lindstrom. A note on epsilons and deltas. Two-Year College Mathematics Journal, 5(3):12–13, September 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1974.11974327.

Lindstrom:1975:STY

[Lin75] Peter A. Lindstrom. Survival of the two-year College mathematics teacher. Two-Year College Mathematics Journal, 6(1):11–13, February 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1975.11974360.

Lindstrom:1976:TYCa

[Lin76a] Peter A. Lindstrom. Two-year College news. Two-Year College Mathematics Journal, 7(2):43–48, May 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1976.
11974429.

Lindstrom:1976:TYCb

[Lin76b] Peter A. Lindstrom. Two-year College news. Two-Year College Mathematics Journal, 7(3):41–46, September 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1976.11974446.

Lindstrom:1978:NIT

[Lin78] Peter A. Lindstrom. A note on the integral test. Two-Year College Mathematics Journal, 9(2):105–106, March 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1978.11974558.

Lindstrom:1981:TRa

[Lin81a] Peter A. Lindstrom. Telegraphic reviews. Two-Year College Mathematics Journal, 12(3):223–224, June 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1981.
11972533.

Lindstrom:1981:TRb

[Lin81b] Peter A. Lindstrom. Telegraphic reviews. Two-Year College Mathematics Journal, 12(4):286–287, September 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972551.

Lindstrom:1981:TRc

[Lin81c] Peter A. Lindstrom. Telegraphic reviews. Two-Year College Mathematics Journal, 12(5):348–349, November 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972569.

Lindstrom:1982:TRa

[Lin82a] Peter A. Lindstrom. Telegraphic reviews. Two-Year College Mathematics Journal, 13(1):74–76, January 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1982.
11972586.

Lindstrom:1982:TRb

[Lin82b] Peter A. Lindstrom. Telegraphic reviews. Two-Year College Mathematics Journal, 13(2):114–164, March 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1982.
11972589.

Lindstrom:1982:TRc

[Lin82c] Peter A. Lindstrom. Telegraphic reviews. Two-Year College Mathematics Journal, 13(3):219–220, June 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1982.
11972617.

Lindstrom:1983:TRa

[Lin83a] Peter A. Lindstrom. Telegraphic reviews. Two-Year College Mathematics Journal, 14(2):179–181, March 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1983.
11972679.

Lindstrom:1983:TRb

[Lin83b] Peter A. Lindstrom. Telegraphic reviews. Two-Year College Mathematics Journal, 14(3):269–271, June 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1983.
11972698.

Lindstrom:1983:TRc

[Lin83c] Peter A. Lindstrom. Telegraphic reviews. Two-Year College Mathematics Journal, 14(4):363–364, September 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1983.11972718.

Lindstrom:1985:SCD

[Lin85] Peter A. Lindstrom. A self-contained derivation of the formula d/dx(x^{r}) = rx^{r−1} for rational r. College Mathematics Journal, 16(2):131–132, March 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1985.11972867.

Linderholm:2000:BR

[Lin00a] Carl Linderholm. Book review. College Mathematics Journal, 31(1):77–79, January 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2000.11974116.

Linwood:2000:CGT

[Lin00b] D. A. Linwood. The calculus and Gamow’s theory of gravitation. College Mathematics Journal, 31(4):281–285, September 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2000.11974159.

Linhart:2019:ALI

[Lin19] Jean Marie Linhart. The art of logic in an illogical world. College Mathematics Journal, 50(5):385–388, 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2019.
1673639.

Linhart:2021:MHF

[Lin21] Jean Marie Linhart. Mathematics for human flourishing. College Mathematics Journal, 52(1):59–63, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2020.
1848202.

Lipkin:2003:TFC

[Lip03] Leonard Lipkin. Tossing a fair coin. College Mathematics Journal, 34(2):128–133, March 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2003.11921996.

Liu:2001:SML

[Liu01] Andy Liu. In search of a missing link: a case study in errorcorrecting codes. College Mathematics Journal, 32(5):343–347, November 2001. CODEN ???? ISSN 0746-8342 (print), 1931-

1346 (electronic). URL http://www.tandfonline.com/doi/
abs/10.1080/07468342.2001.11921901.

Liu:2009:TAH

[Liu09] Andy Liu. Two applications of a Hamming code. College Mathematics Journal, 40(1):2–5, January 2009. CODEN ????

ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2009.
11922327.

Liu:2011:PER

[Liu11] Andy Liu. Is parallelism an equivalence relation? College Mathematics Journal, 42(5):372, November 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
42.5.372.

Landau:1973:IQR

[LJ73] Martin D. Landau and William R. Jones. An interpolation question resolved by calculus. Two-Year College Mathematics Journal, 4(1):36–39, Winter 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1973.11974233.

Lima:2022:WIR

[LJ22] F. M. S. Lima and P. G. F. Jordão. Why is it that the ratio of any circle’s circumference to its diameter is a constant? College Mathematics Journal, 53(3):171–182, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2022.2040326.

Lindquist:1971:HCB

[LL71] Norman F. Lindquist and Norman F. Lindquist. How are computers being used in two-year colleges? Two-Year College Mathematics Journal, 2(1):71, Spring 1971. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1971.11974003.

Long:2001:SAI

[LL01] Andrew E. Long and Clifford A. Long. Surface approximation and interpolation via matrix SVD. College Mathematics Journal, 32(1):20–25, January 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2001.11921846.

Li:2021:CDC

[LLH21] Hou-Biao Li, Hong Li, and Ting-Zhu Huang. Chaos in determinant condensation calculations. College Mathematics Journal, 52(5):345–354, 2021. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2021.1967695.

Lindstrom:1979:BRe

[LM79] Peter A. Lindstrom and Kenneth E. Martin. Book reviews. Two-Year College Mathematics Journal, 10(5):355–358, November 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1979.11972389.

Lopez:1990:UCA

[LM90] Robert Lopez and John Mathews. Using a computer algebra system to solve for maxima and minima. College Mathematics Journal, 21(5):410–414, November 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1990.
11973346.

Lange:1992:RLG

[LM92] Lester H. Lange and James W. Miller. A random ladder game: Permutations, eigenvalues, and convergence of Markov chains. College Mathematics Journal, 23(5):373–385, November 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1992.11973487.

Lutzer:2006:DBH

[LM06] Carl V. Lutzer and James E. Marengo. The divergence of balanced harmonic-like series. College Mathematics Journal, 37 (5):364–369, November 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2006.11922206.

Lopez:2010:UBE

[LM10] Manuel Lopez and James Marengo. An upper bound for the expected range of a random sample. College Mathematics Journal, 41(1):42–48, January 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/074683410X475100.

Littleton:2016:SEB

[LM16] Leonard Littleton and Russell May. Simplified expectations in the birthday problem. College Mathematics Journal, 47(1):50–55, January 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.47.1.50.

Litvinov:2020:SPC

[LM20] Semyon Litvinov and Frantisek Marko. Sums of powers of consecutive integers and Pascal’s triangle. College Mathematics Journal, 51(1):25–31, 2020. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2020.1675428.

Levin:2021:PC

[LM21] Oscar Levin and Tyler Markkanen. Puzzles of cardinality. College Mathematics Journal, 52(4):243–253, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2021.
1941539.

Litvinov:2023:TCB

[LM23] Semyon Litvinov and Frantisek Marko. Two comments on the binomial theorem. College Mathematics Journal, 54(2):142– 144, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2023.2186084.

Lynch:2024:PPR

[LM24] Peter Lynch and Michael Mackey. Parity and partition of the rational numbers. College Mathematics Journal, 55(5):387–399, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic).

Laible:1970:LE

[LMC^{+}70] Jon M. Laible, John W. Milsom, Robert J. Cohen, Darrell J. Peterson, G. Thomas Riggle II, Ronald Johnston, C. M. Adams, Martha C. Jordan, and J. I. Cone. Letters to the Editor. Two-Year College Mathematics Journal, 1(2):1–2, Autumn 1970. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1970.11973178.

Lindstrom:1979:BRa

[LMK79] Peter A. Lindstrom, Shelba Jean Morman, and James C. Kropa. Book reviews. Two-Year College Mathematics Journal, 10(1):

43–48, January 1979. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1979.11972322.

Long:1980:CC

[LN80] Cliff Long and Vic Norton. Computers & calculators. Two-Year College Mathematics Journal, 11(5):320–325, November 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1980.11972470.

Leigh:1999:MAL

[LN99] Robert Barrington Leigh and Richard Travis Ng. Minimizing aroma loss. College Mathematics Journal, 30(5):356–358, November 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1999.11974085.

LeCrone:2014:MG

[LN14] Jeremy LeCrone and Nancy Ann Neudauer. Matroids on groups? College Mathematics Journal, 45(2):121–128, March 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.45.2.121.

LoBello:1991:VCS

[Lo 91] Anthony Lo Bello. The volume and centroid of the Step Pyramid of Zoser. College Mathematics Journal, 22(4):318–321, September 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1991.11973405.

LoBello:1997:NBR

[Lo 97] Anthony Lo Bello. A nod to Bertrand Russell. College Mathematics Journal, 28(2):133, March 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1997.
11973851.

Loehr:2014:CNM

[Loe14] Nick Loehr. Classifying nilpotent maps via partition diagrams. College Mathematics Journal, 45(2):108–115, March 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.45.2.108.

Loeb:2021:CRN

[Loe21] Peter A. Loeb. Chain rule note. College Mathematics Journal, 52(1):57–58, 2021. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2020.1854007.

Loftus:1978:LE

[Lof78] Marie Loftus. Letters to the Editor. Two-Year College Mathematics Journal, 9(3):131, June 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1978.11974561.

Logothetti:1979:SBM

[Log79] Dave Logothetti. Super bat meets the word problem. Two-Year College Mathematics Journal, 10(5):371, November 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1979.11972392.

Logothetti:1980:IHM

[Log80] Dave Logothetti. An interview with H. S. M. Coxeter, the king of geometry. Two-Year College Mathematics Journal, 11(1):2–19, January 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1980.11972395.

Lomen:1994:EPD

[Lom94] David O. Lomen. Experiments with probes in the differential equations classroom. College Mathematics Journal, 25(5):453– 457, November 1994. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1994.11973652.

Long:1972:QPZ

[Lon72] C. A. Long. The quadratic polynomial and its zeros. Two-Year College Mathematics Journal, 3(1):23–29, Spring 1972. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1972.11974192.

Long:1976:FMP

[Lon76] Cliff Long. A flexible model for peak, ridge and pass. Two-Year College Mathematics Journal, 7(3):16–17, September 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1976.11974439.

Long:1981:HMP

[Lon81] Calvin T. Long. How many positive integers have nines in their decimal representations? Two-Year College Mathematics Journal, 12(5):320–324, November 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1981.
11972559.

Long:1983:RIE

[Lon83] Calvin T. Long. On the radii of the inscribed and escribed circles of right triangles — a second look. Two-Year College Mathematics Journal, 14(5):382–389, November 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1983.11972724.

Long:1984:PTD

[Lon84] Calvin Long. Pascal’s Triangle, difference tables and arithmetic sequences of order N. College Mathematics Journal, 15(4):290– 298, September 1984. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1984.11972787.

Lopez:1993:UPD

[Lop93] Antonio M. Lopez, Jr. Using PROLOG in discrete mathematics. College Mathematics Journal, 24(4):357–365, September 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1993.11973556.

Lorch:2000:MGR

[Lor00] John D. Lorch. Modeling the gaitpath of a running animal. College Mathematics Journal, 31(2):93–97, March 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2000.11974120.

Lotto:2023:EPR

[Lot23] Benjamin A. Lotto. An elementary proof that reduced row echelon form of a matrix is unique. College Mathematics Journal, 54(2):145–146, 2023. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2023.2184168.

Luciano:1987:CCC

[LP87] Dennis Luciano and Gordon Prichett. Cryptology: From Caesar ciphers to public-key cryptosystems. College Mathematics Journal, 18(1):2–17, January 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1987.11973000.

Leinbach:1988:SRa

[LP88] L. Carl Leinbach and Bill Parker. Software reviews. College Mathematics Journal, 19(2):184–185, March 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1988.
11973113.

Laubenbacher:1994:EMG

[LP94] Reinhard C. Laubenbacher and David J. Pengelley. Eisenstein’s misunderstood geometric proof of the Quadratic Reciprocity Theorem. College Mathematics Journal, 25(1):29–34, January 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1994.11973576.

Lee:2014:RRM

[LP14] Christopher Lee and Valerie Peterson. The rank of recurrence matrices. College Mathematics Journal, 45(3):207–215, May 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.45.3.207.

Leake:2016:ACT

[LP16] Jacqueline Leake and Nicholas Pritchard. The advantage of the coin toss for the new overtime system in the national football league. College Mathematics Journal, 47(1):2–9, January 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.47.1.2.

Loomis:2008:SEF

[LPP08] Paul Loomis, Michael Plytage, and John Polhill. Summing up the Euler ϕ function. College Mathematics Journal, 39(1):34–42, January 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2008.11922272.

Leggett:2011:CDD

[LPT11] Deanna Leggett, John Perry, and Eve Torrence. Computing determinants by double-crossing. College Mathematics Journal, 42(1):43–54, January 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.42.1.043.

Levin:2013:CKK

[LR13] Oscar Levin and Gerri M. Roberts. Counting knights and knaves. College Mathematics Journal, 44(4):300–306, September 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.44.4.300.

Lawson:2020:SGC

[LR20] Jeffrey Lawson and Matthew Rave. Sweeping gestures: A control theory model for curling. College Mathematics Journal, 51(2):132–140, 2020. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2020.1707044.

Lawson:2024:ZZP

[LR24] Dan Lawson and David Richeson. Zig-zag paths and neusis constructions of a heptagon and a nonagon. College Mathematics Journal, 55(1):46–55, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2023.2261347.

Lindstrom:1979:BRd

[LRM79] Peter A. Lindstrom, Henry J. Ricardo, and Barnabas Hughes O. F. M. Book reviews. Two-Year College Mathematics Journal, 10(4):287–292, September 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1979.11972371.

Leung:2014:CSP

[LRSZ14] Kimberly Leung, Chris Rasmussen, Samuel S. P. Shen, and Dov Zazkis. Calculus from a statistics perspective. College Mathematics Journal, 45(5):377–386, November 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
45.5.377.

Leonard:1979:TS

[LS79] Bill Leonard and Harris S. Shultz. Triangular squares. Two-Year College Mathematics Journal, 10(3):169–171, June 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1979.11972345.

Laubenbacher:1994:GPM

[LS94] Reinhard C. Laubenbacher and Michael Siddoway. Great problems of mathematics: a Summer workshop for high school students. College Mathematics Journal, 25(2):112–114, March 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1994.11973591.

Littleton:2001:DCS

[LS01] Pam Littleton and David A. Sánchez. Dipsticks for cylindrical storage tanks-exact and approximate. College Mathematics Journal, 32(5):352–358, November 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2001.
11921903.

Latulippe:2012:SNM

[LS12] Joe Latulippe and Jennifer Switkes. Sometimes Newton’s Method always Cycles. College Mathematics Journal, 43(5):365–370, November 2012. CODEN ???? ISSN 0746-8342 (print), 1931-

1346 (electronic). URL http://www.tandfonline.com/doi/
abs/10.4169/college.math.j.43.5.365.

Leise:2019:MHa

[LS19a] Tanya Leise and Philip Straffin. Media highlights. College Mathematics Journal, 50(3):233–240, 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2019.
1603712.

Leise:2019:MHb

[LS19b] Tanya Leise and Philip Straffin. Media highlights. College Mathematics Journal, 50(4):313–320, 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2019.
1650598.

Leise:2019:MHc

[LS19c] Tanya Leise and Philip Straffin. Media highlights. College Mathematics Journal, 50(5):389–396, 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2019.
1664859.

Leise:2020:MHa

[LS20a] Tanya Leise and Philip Straffin. Media highlights. College Mathematics Journal, 51(1):73–80, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2020.
1680225.

Leise:2020:MHb

[LS20b] Tanya Leise and Philip Straffin. Media highlights. College Mathematics Journal, 51(2):153–160, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2019.
1701916.

Leise:2020:MHc

[LS20c] Tanya Leise and Philip Straffin. Media highlights. College Mathematics Journal, 51(3):233–240, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2020.
1724528.

Leise:2020:MHd

[LS20d] Tanya Leise and Philip Straffin. Media highlights. College Mathematics Journal, 51(4):313–320, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2020.
1791622.

Leise:2020:MHe

[LS20e] Tanya Leise and Philip Straffin. Media highlights. College Mathematics Journal, 51(5):393–400, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2020.
1826770.

Leise:2021:MHa

[LS21a] Tanya Leise and Philip Straffin. Media highlights. College Mathematics Journal, 52(1):73–80, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2020.
1853982.

Leise:2021:MHb

[LS21b] Tanya Leise and Philip Straffin. Media highlights. College Mathematics Journal, 52(2):149–156, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2021.
1872293.

Leise:2021:MHd

[LS21c] Tanya Leise and Philip Straffin. Media highlights. College Mathematics Journal, 52(4):316–323, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2021.
1949222.

Leise:2021:MHe

[LS21d] Tanya Leise and Philip Straffin. Media highlights. College Mathematics Journal, 52(5):397–404, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2021.
1969182.

Leise:2022:MHa

[LS22a] Tanya Leise and Philip Straffin. Media highlights. College Mathematics Journal, 53(1):77–84, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2021.
1999727.

Leise:2022:MHb

[LS22b] Tanya Leise and Philip Straffin. Media highlights. College Mathematics Journal, 53(2):161–168, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2022.
2026089.

Leise:2022:MHc

[LS22c] Tanya Leise and Philip Straffin. Media highlights. College Mathematics Journal, 53(3):245–252, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2022.
2045155.

Lengvarszky:2022:LCD

[LS22d] Zsolt Lengvárszky and Tibor Szarvas. The law of cosines with differential calculus and without. College Mathematics Journal, 53(2):98–103, 2022. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2022.2011542.

Leise:2023:MHa

[LS23a] Tanya Leise and Philip Straffin. Media highlights. College Mathematics Journal, 54(1):72–79, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2023.
2165374.

Leise:2023:MHb

[LS23b] Tanya Leise and Philip Straffin. Media highlights. College Mathematics Journal, 54(2):161–168, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2023.
2186092.

Lindstrom:1979:BRc

[LTM79] Peter A. Lindstrom, Elaine Johnson Tatham, and Charles Miller. Book reviews. Two-Year College Mathematics Journal, 10(3): 205–209, June 1979. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1979.11972353.

Lu:2022:ULI

[Lu22] Yuanting Lu. Using linear interpolation to implement the change of variables in double integrals. College Mathematics Journal, 53(1):64–66, 2022. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2021.1971032.

Lucas:1983:PPN

[Luc83] John F. Lucas. Paths and Pascal numbers. Two-Year College Mathematics Journal, 14(4):329–341, September 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1983.11972709.

Lucas:1984:DMC

[Luc84] William F. Lucas. Discrete mathematics courses have constituents besides computer scientists …. College Mathematics Journal, 15(5):386–388, November 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1984.
11972812.

Lucarelli:1998:FBA

[Luc98] Vincent Lucarelli. On factoring n with the b-algorithm. College Mathematics Journal, 29(4):289–295, September 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1998.11973957.

Luthar:1980:LMS

[Lut80] R. S. Luthar. Luddhar’s Method of solving a cubic equation with a rational root. Two-Year College Mathematics Journal, 11(2):107–110, March 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1980.11972414.

Luthar:1984:IST

[Lut84] R. S. Luthar. Integer-sided triangles with one angle twice another. College Mathematics Journal, 15(1):55–56, January 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972748.

Luther:2011:TTI

[Lut11] Ken Luther. Teaching tip: Is this integral zero? College Mathematics Journal, 42(5):373, November 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
42.5.373.

Lutzer:2017:CFR

[Lut17] Carl V. Lutzer. A curious feature of regression. College Mathematics Journal, 48(3):189–198, May 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
48.3.189.

Liu:2010:CML

[LW10] Yingfan Liu and Youguo Wang. A class of multivariable limits. College Mathematics Journal, 41(2):154–156, March 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/074683410X480267.

Lorenz:2018:SRM

[LW18] Dirk Lorenz and Karl-Joachim Wirths. Sarrus rules for matrix determinants and dihedral groups. College Mathematics Journal, 49(5):333–340, November 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2018.1526019.

Lycan:2023:TQE

[Lyc23] Ron Lycan. Two quasigroup elements can commute with any positive rational probability. College Mathematics Journal, 54 (4):388–390, 2023. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2023.2237853.

LaHaye:2018:GIG

[LZ18] Roberta La Haye and Petr Zizler. The Gini index and grayscale images. College Mathematics Journal, 49(3):205–211, May 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2018.1426330.

Mabry:2010:HSP

[Mab10] Rick Mabry. The hardest straight-in pool shot. College Mathematics Journal, 41(1):49–57, January 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.

tandfonline.com/doi/abs/10.4169/074683410X475119.

MacDonald:1978:IMA

[Mac78] Carolyn T. MacDonald. Introductory mathematics and the adult woman student. Two-Year College Mathematics Journal, 9(3): 158–161, June 1978. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1978.11974568.

MacLane:1984:CDL

[Mac84] Saunders MacLane. Calculus is a discipline …. College Mathematics Journal, 15(5):373, November 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1984.
11972804.

Mackenzie:1992:TP

[Mac92a] Dana N. Mackenzie. Triquetras and porisms. College Mathematics Journal, 23(2):118–131, March 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1992.
11973443.

Mackiw:1992:NTL

[Mac92b] George Mackiw. Number theory and linear algebra: Exact solutions of integer systems. College Mathematics Journal, 23(1): 52–58, January 1992. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1992.11973436.

Macula:1995:PSF

[Mac95] Anthony J. Macula. The point-slope formula leads to the fundamental theorem of calculus. College Mathematics Journal, 26(2):135–139, March 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1995.11973683.

Mackiw:1996:CAA

[Mac96] George Mackiw. Computing in abstract algebra. College Mathematics Journal, 27(2):136–142, March 1996. CODEN ????

ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1996.
11973766.

MacKendrick:2003:WF

[Mac03] Sharon MacKendrick. For what functions is f^{−1}(x) = 1/f(x)? College Mathematics Journal, 34(4):304–311, September 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2003.11922022.

Maesumi:2019:TIT

[Mae19] Arman Maesumi. Triangle inscribed-triangle picking. College Mathematics Journal, 50(5):364–371, 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2019.
1669383.

Mahmoudi:2020:GIG

[Mah20] M. G. Mahmoudi. The AM-GM inequality via gradient. College Mathematics Journal, 51(2):141–143, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2020.
1697605.

Main:1970:EMI

[Mai70] Carl L. Main. Encouraging mathematical inquisitiveness. Two-Year College Mathematics Journal, 1(1):32–36, Spring 1970. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1970.11973164.

Malraison:1981:BR

[Mal81] Pierre J. Malraison. Book reviews. Two-Year College Mathematics Journal, 12(4):285–286, September 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1981.
11972550.

Maltenfort:2015:STR

[Mal15] Michael Maltenfort. Secants, tangents, rotations, and reflections. College Mathematics Journal, 46(1):24–34, January 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.46.1.24.

Maltenfort:2016:CCW

[Mal16] Michael Maltenfort. A canine conundrum, or what would Elvis do? College Mathematics Journal, 47(2):106–107, March 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.47.2.106.

Maltenfort:2017:FWS

[Mal17] Michael Maltenfort. A function worth a second look. College Mathematics Journal, 48(1):55–57, January 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
48.1.55.

Malinovsky:2022:SPP

[Mal22a] Yaakov Malinovsky. A short probabilistic proof of a binomial identity. College Mathematics Journal, 53(5):394–395, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2022.2072093.

Maltenfort:2022:AIM

[Mal22b] Michael Maltenfort. The associativity of infinite matrix multiplication, revisited. College Mathematics Journal, 53(1):39– 44, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2022.1990648.

Mansfield:1970:TYC

[Man70] Ralph Mansfield. Two-year colleges and post-secondary education in Western Europe. Two-Year College Mathematics Journal, 1(2):50–55, Autumn 1970. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1970.11973188.

Mansfield:1974:MMW

[Man74] Ralph Mansfield. The man-made world: Cultural vs remedial mathematics. Two-Year College Mathematics Journal, 5(1):9–21, Winter 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1974.11974286.

Mansfield:1975:MIY

[Man75] Ralph Mansfield. Mathematics — is it any of your business? Two-Year College Mathematics Journal, 6(3):20–26, September 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1975.11974388.

Mansfield:1978:AMT

[Man78] Ralph Mansfield. Applicable mathematics in two year colleges. Two-Year College Mathematics Journal, 9(3):148–153, June 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1978.11974566.

Manin:1981:DP

[Man81] Yu. I. Manin. A digression on proof. Two-Year College Mathematics Journal, 12(2):104–107, March 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1981.
11972505.

Mandelbrot:1984:APF

[Man84] Benoît B. Mandelbrot. Additional perspectives on fractals. College Mathematics Journal, 15(2):115–119, March 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972759.

Mann:1998:BRT

[Man98] James E. Mann, Jr. The buckled rail: Three formulations. College Mathematics Journal, 29(2):138–141, March 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1998.11973928.

Martic:1982:REG

[Mar82] Branislav Martić. Remark on an elementary gem concerning Π(n). Two-Year College Mathematics Journal, 13(2):158–159, March 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1982.11972598.

Marcus:1990:DS

[Mar90a] Marvin Marcus. Determinants of sums. College Mathematics Journal, 21(2):130–135, March 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1990.
11973297.

Marty:1990:NPR

[Mar90b] Roger H. Marty. The number of paths in a rooted binary tree of infinite height. College Mathematics Journal, 21(4):305– 307, September 1990. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1990.11973324.

Maruszewski:1991:PLC

[Mar91] Richard F. Maruszewski, Jr. Programs for a logic course. College Mathematics Journal, 22(3):235–241, May 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/07468342.
1991.11973390.

Marcus:1992:DL

[Mar92a] Marvin Marcus. Determinantal loci. College Mathematics Journal, 23(1):44–47, January 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1992.11973433.

Markowsky:1992:MAG

[Mar92b] George Markowsky. Misconceptions about the Golden ratio. College Mathematics Journal, 23(1):2–19, January 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1992.11973428.

Markelov:1998:GCS

[Mar98a] Sergey Markelov. Geometric characterization of the shortest path in a tetrahedron. College Mathematics Journal, 29(2):150–151, March 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1998.11973934.

Martins:1998:SDR

[Mar98b] Luiz Felipe Martins. A simple decision rule for a guessing game. College Mathematics Journal, 29(5):371–375, November 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1998.11973972.

Martin:2008:SCT

[Mar08] Greg Martin. Sets that contain their circle centers. College Mathematics Journal, 39(5):357–366, November 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2008.11922316.

Maruszewski:2009:FFN

[Mar09] Richard Maruszewski. Fibonacci’s forgotten number revisited. College Mathematics Journal, 40(4):248–251, September 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2009.11922371.

Martin:2010:HG

[Mar10] John Martin. The Helen of geometry. College Mathematics Journal, 41(1):17–28, January 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/074683410X475083.

Marrero:2013:SVE

[Mar13] Osvaldo Marrero. Seasonal variation in epidemiology. College Mathematics Journal, 44(5):386–398, November 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
44.5.386.

Maritz:2024:EPC

[Mar24] Milton F. Maritz. Extracting pi from chaos. College Mathematics Journal, 55(2):86–99, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2023.2265282.

Mason:1984:NBR

[Mas84] John Mason. A need for better role models and attitudes …. College Mathematics Journal, 15(5):395–396, November 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972818.

Massell:1991:UCG

[Mas91] Paul B. Massell. Using computer graphics to help analyze complicated functions. College Mathematics Journal, 22(4):327– 331, September 1991. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1991.11973408.

Mason:2011:GPC

[Mas11] John Mason. A generalization of the parabolic chord property. College Mathematics Journal, 42(4):326–328, September 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.42.4.326.

Matthews:1988:ABM

[Mat88] Keith Matthews. Algorithm of the bi-month. College Mathematics Journal, 19(2):174–176, March 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1988.
11973111.

Mathews:1991:SO

[Mat91] John H. Mathews. The sum is one. College Mathematics Journal, 22(4):322, September 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1991.11973406.

Mathews:1992:GP

[Mat92] John H. Mathews. A geometric proof of lim_{d→0^{+}}(−d ln d) = 0. College Mathematics Journal, 23(3):209, May 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/07468342.
1992.11973459.

Mathews:1994:IE

[Mat94] John H. Mathews. The integral of x^{1/2}, etc. College Mathematics Journal, 25(2):142–144, March 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1994.
11973599.

Mathews:1996:MER

[Mat96] David M. Mathews. Mathematics education: a response to Andrews. College Mathematics Journal, 27(5):349–353, November 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1996.11973806.

Matsko:2014:SOR

[Mat14] Vincent J. Matsko. Second-order recurrences with nonconsecutive initial conditions. College Mathematics Journal, 45(1):41–49, January 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.45.1.041.

Matsko:2017:DKL

[Mat17] Vincent J. Matsko. Designing Koch-like curves. College Mathematics Journal, 49(1):11–19, January 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2017.
1395264.

Mateer:2025:CPB

[Mat25a] Todd Mateer. A combinatorial proof of a binomial coefficient identity. College Mathematics Journal, 56(1):65–67, 2025. CO-DEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic).

Matsko:2025:RGS

[Mat25b] Vincent J. Matsko. Recursively generated spirals. College Mathematics Journal, 56(1):4–14, 2025. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic).

Maurer:1983:IAW

[Mau83a] Stephen B. Maurer. An interview with Albert W. Tucker. Two-Year College Mathematics Journal, 14(3):210–224, June 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1983.11972685.

Maurer:1983:SUL

[Mau83b] Stephen B. Maurer. Some unusual locus problems. Two-Year College Mathematics Journal, 14(2):146–153, March 1983.

CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1983.11972670.

Maurer:1984:MH

[Mau84] Stephen B. Maurer. Media highlights. College Mathematics Journal, 15(5):452–460, November 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1984.
11972835.

Maurer:1985:AWL

[Mau85] Stephen B. Maurer. The algorithmic way of life is best. College Mathematics Journal, 16(1):2–5, January 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/07468342.
1985.11972838.

Maurer:2002:D

[Mau02] Stephen B. Maurer. Hat derivatives. College Mathematics Journal, 33(1):32–37, January 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2002.11921916.

Maxfield:1974:IE

[Max74] Margaret Maxfield. Innovative evaluation. Two-Year College Mathematics Journal, 5(1):47–52, Winter 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1974.
11974292.

Mayer:1997:PPR

[May97a] Uwe F. Mayer. A proof that polynomials have roots. College Mathematics Journal, 28(1):58, January 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1997.
11973834.

Mayes:1997:SR

[May97b] Robert Mayes. Software reviews. College Mathematics Journal, 28(4):310–314, September 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1997.11973883.

Mayfield:2013:WMT

[May13] Betty Mayfield. Women and mathematics in the time of Euler. College Mathematics Journal, 44(2):82–88, March 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.44.2.082.

May:2024:WTP

[May24] Rus May. The weight of a ten-pound bag of potatoes. College Mathematics Journal, 55(2):109–116, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2023.
2270891.

Murty:1982:LSB

[MB82] V. N. Murty and B. H. Bissinger. The law of succession and Bayes’ Rule. Two-Year College Mathematics Journal, 13(1): 44–51, January 1982. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1982.11972577.

Mounteer:1978:RII

[MC78] Gail B. Mounteer and Robert J. Cermele. The role of the instructor in the individualized classroom. Two-Year College Mathematics Journal, 9(5):276–281, November 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1978.11974592.

Maritz:2020:FRN

[MC20] Milton F. Maritz and Marèt Cloete. Fold-over regions in nonlinear first order PDEs. College Mathematics Journal, 51(3):204– 215, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2020.1735866.

Maritz:2024:AVF

[MC24] Milton F. Maritz and Marèt Cloete. Analysis and visualization of fractional reflections. College Mathematics Journal, 55(1):4– 11, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2023.2253129.

McAllister:1983:PSC

[McA83] Byron L. McAllister. Products of sets of complex numbers. Two-Year College Mathematics Journal, 14(5):390–397, November 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1983.11972726.

McCarty:1980:RA

[McC80a] Carl P. McCarty. A rational approximation to \sqrt{n}. Two-Year College Mathematics Journal, 11(2):123–124, March 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1980.11972422.

McCarty:1980:CC

[McC80b] George McCarty. Computers & calculators. Two-Year College Mathematics Journal, 11(1):42–48, January 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1980.
11972398.

McCullough:1980:NVT

[McC80c] Thomas A. McCullough. A note on the vector triple product. Two-Year College Mathematics Journal, 11(3):206–207, June 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1980.11972440.

McCarty:1990:PS

[McC90] Loretta McCarty. Polar summation. College Mathematics Journal, 21(5):397–398, November 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1990.
11973340.

McCartin:1998:PMG

[McC98a] Brian J. McCartin. Prelude to musical geometry. College Mathematics Journal, 29(5):354–370, November 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1998.
11973971.

McClendon:1998:MP

[McC98b] Michael Scott McClendon. Minimal pyramids. College Mathematics Journal, 29(3):224–226, May 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1998.
11973945.

McCutcheon:2001:AGI

[McC01] Randall McCutcheon. Analyzing games of information. College Mathematics Journal, 32(2):82–90, March 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/07468342.
2001.11921858.

McCoart:2002:LCA

[McC02] Richard F. McCoart. Lewis Carroll’s amazing number-guessing game. College Mathematics Journal, 33(5):378–383, November 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2002.11921967.

McCune:2023:SAD

[McC23] David McCune. A simple algorithm for drawing apportionment diagrams. College Mathematics Journal, 54(1):22–32, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2022.2160618.

McDonald:2000:BR

[McD00] Michael A. McDonald. Book review. College Mathematics Journal, 31(2):157–159, March 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2000.11974135.

McGivney:1990:FHM

[McG90] Raymond A. McGivney, Jr. Forward homework — a motivational tool. College Mathematics Journal, 21(5):400–402, November 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1990.11973342.

McGrath:2018:NST

[McG18] Peter McGrath. Newton’s Shell Theorem via Archimedes’s hat box and single-variable calculus. College Mathematics Journal, 49(2):109–113, March 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2018.1411655.

McKenna:1976:CRD

[McK76a] James E. McKenna. Computation of repeating decimals. Two-Year College Mathematics Journal, 7(2):55–58, May 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1976.11974432.

McKenna:1976:ISR

[McK76b] James E. McKenna. Interpolation and square roots. Two-Year College Mathematics Journal, 7(4):49–50, December 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1976.11974467.

McKelvey:1979:TYC

[McK79] Robert McKelvey. The two-year colleges and the graduate schools: The teachers’ perspective. Two-Year College Mathematics Journal, 10(2):136–142, March 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1979.
11972341.

McLeod:1987:MH

[McL87] Douglas B. McLeod. Media highlights. College Mathematics Journal, 18(4):346–355, September 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1987.
11973058.

McMillan:2003:ECP

[McM03] Thomas C. McMillan. Examining continuity, partial derivatives and differentiability with cylindrical coordinates. College Mathematics Journal, 34(1):11–14, January 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2003.
11921978.

McCune:1980:CMI

[MDC80] E. D. McCune, R. G. Dean, and W. D. Clark. Calculators to motivate infinite composition of functions. Two-Year College Mathematics Journal, 11(3):189–195, June 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1980.11972434.

Mathes:2011:CG

[MDL11] Ben Mathes, Chris Dow, and Leo Livshits. Cantor groups. College Mathematics Journal, 42(1):60–61, January 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
42.1.060.

Melo:2019:UCA

[MdMH19] Cesar Adolfo Hernandez Melo and Fernanda Diniz de Melo Hernandez. Uniform continuity: Another way to approach this concept in the classroom. College Mathematics Journal, 50(1): 54–57, January 2019. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2019.1535208.

Mehta:2022:SGP

[Meh22] Jay Mehta. A short generalized proof of infinitude of primes. College Mathematics Journal, 53(1):52–53, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2022.
1987114.

Melman:2022:ESZ

[Mel22] Aaron Melman. An exclusion sector for zeros of polynomials with nonnegative coefficients. College Mathematics Journal, 53 (4):312–314, 2022. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2022.2099717.

Mendelsohn:1988:AFS

[Men88] N. S. Mendelsohn. Amalgamation of formulae for sequences. College Mathematics Journal, 19(5):421–423, November 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1988.11973150.

Mendelsohn:2004:TD

[Men04] N. S. Mendelsohn. Tiling with dominoes. College Mathematics Journal, 35(2):115–120, March 2004. CODEN ????

ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2004.
11922062.

Merritt:1974:SRS

[Mer74] Paul W. Merritt. Some research support for a second chance for beginning algebra students. Two-Year College Mathematics Journal, 5(2):50–54, Spring 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1974.11974314.

Mercier:1997:HCT

[Mer97] William C. Mercier. How chaotic things work. College Mathematics Journal, 28(2):110–118, March 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1997.
11973845.

Mercer:2005:EEN

[Mer05] Peter R. Mercer. Error estimates for numerical integration rules. College Mathematics Journal, 36(1):27–34, January 2005. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2005.11922107.

Merca:2017:AMS

[Mer17] Mircea Merca. On the arithmetic mean of the square roots of the first n positive integers. College Mathematics Journal, 48(2): 129–133, March 2017. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.48.2.129.

Mercier:2018:WCC

[Mer18] William C. Mercier. Why the centroid is the centroid: Modern variations on a theme of Archimedes. College Mathematics Journal, 49(2):93–102, March 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2018.1420982.

Merca:2022:CBP

[Mer22a] Mircea Merca. Connections between partitions and divisors related to the parity of the partition function. College Mathematics Journal, 53(1):33–37, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2022.1990637.

Mercer:2022:EEB

[Mer22b] Peter R. Mercer. Extending extending bernoulli’s inequality. College Mathematics Journal, 53(2):149–150, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/full/10.1080/07468342.
2022.2020587.

Meyerson:1984:RGG

[Mey84] Mark D. Meyerson. Return of the grazing goat in n dimensions. College Mathematics Journal, 15(5):430–432, November 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972829.

Meyers:1986:TWM

[Mey86] Leroy F. Meyers. Three ways to maximize the area of an inscribed quadrilateral. College Mathematics Journal, 17(3):238– 239, May 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1986.11972962.

Meyers:1991:FCE

[Mey91] Leroy F. Meyers. Four crotchets on elementary integration. College Mathematics Journal, 22(5):410–413, November 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1991.11973417.

Meyer:2006:DGR

[Mey06] Walter Meyer. David gaie: Restless pioneer. College Mathematics Journal, 37(1):29–38, January 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2006.
11922164.

Meyer:2007:OFM

[Mey07] Walter Meyer. The origins of finite mathematics: The social science connection. College Mathematics Journal, 38(2):106–118, March 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2007.11922225.

McDill:1994:LSD

[MF94] J. M. McDill and Bjørn Felsager. The lighter side of differential equations. College Mathematics Journal, 25(5):448–452, November 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1994.11973651.

Martelli:1997:WCD

[MG97] Mario Martelli and Gerald Gannon. Weighing coins: Divide and conquer to detect a counterfeit. College Mathematics Journal, 28(5):365–367, November 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1997.11973891.

Moreno:2019:PWS

[MGC19] Samuel G. Moreno and Esther M. García-Caballero. Proof without words: Sophie Germain’s identity. College Mathematics Journal, 50(3):197, 2019. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2019.1603533.

Murray:2009:R

[MH09] Margaret A. M. Murray and Reuben Hersh. Reviews. College Mathematics Journal, 40(4):306–312, September 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2009.11922379.

Movshovitz-Hadar:1992:IMD

[MHS92] Nitsa Movshovitz-Hadar and Alla Shmukler. Infinitely many different quartic polynomial curves. College Mathematics Journal, 23(3):186–195, May 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1992.11973455.

Michael:2011:GGF

[Mic11] T. S. Michael. Guards, galleries, fortresses, and the octoplex. College Mathematics Journal, 42(3):191–200, May 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.42.3.191.

Mickens:2012:ESM

[Mic12] Ronald E. Mickens. An exactly solvable model for the spread of disease. College Mathematics Journal, 43(2):114–121, March 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.43.2.114.

Miceli:2014:CPT

[Mic14] Brian K. Miceli. A combinatorial proof of a theorem of Katsuura. College Mathematics Journal, 45(5):365–369, November 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.45.5.365.

Mielke:1983:TPT

[Mie83] Paul T. Mielke. A tiling of the plane with triangles. Two-Year College Mathematics Journal, 14(5):377–381, November 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1983.11972723.

Mihaila:2002:CFF

[Mih02] Ioana Mihaila. Cantor, 1 /4, and its family and friends. College Mathematics Journal, 33(1):21–23, January 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2002.
11921914.

Mihaila:2004:RCS

[Mih04] Ioana Mihaila. The rationals of the Cantor set. College Mathematics Journal, 35(4):251–255, September 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2004.
11922081.

Miller:1971:RT

[Mil71] William I. Miller. F (1) and F (d) rejection theorems. Two-Year College Mathematics Journal, 2(2):95–96, Autumn 1971. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1971.11974019.

Miller:1974:CMV

[Mil74] Wayne L. Miller. The contract method vs the traditional method of teaching developmental mathematics to underachieves: a comparative analysis. Two-Year College Mathematics Journal, 5(2): 45–49, Spring 1974. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1974.11974313.

Miller:1978:DAV

[Mil78] Jay I. Miller. Differentiating area and volume. Two-Year College Mathematics Journal, 9(1):47–48, January 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/00494925.
1978.11974547.

Miller:1985:BBL

[Mil85] William Miller. Both a borrower and a lender be. College Mathematics Journal, 16(4):284, September 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1985.
11972900.

Mills:1993:FFF

[Mil93] Stuart E. Mills. Fallacies, flaws, and flimflam. College Mathematics Journal, 24(5):446–450, November 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1993.
11973565.

Minor:1991:PET

[Min91] Lee H. Minor. A productive error in a trigonometry text. College Mathematics Journal, 22(4):315–318, September 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1991.11973404.

Minor:1992:VLL

[Min92] Lee H. Minor. Visualization of limits and limits of visualization. College Mathematics Journal, 23(1):48–51, January 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1992.11973435.

Minton:1994:PPE

[Min94] Roland Minton. A progression of projectiles: Examples from sports. College Mathematics Journal, 25(5):436–442, November 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1994.11973648.

Minor:2003:PPH

[Min03] Darrell P. Minor. Parrondo’s Paradox — hope for losers! College Mathematics Journal, 34(1):15–20, January 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2003.
11921979.

Misiurewicz:2013:ISR

[Mis13] Michał Misiurewicz. Irrational square roots. College Mathematics Journal, 44(1):53–55, January 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
44.1.053.

Missa:2023:IMS

[Mis23] Abdel Missa. An intuitive method to solve quadratic equations. College Mathematics Journal, 54(1):54–61, 2023. CO-DEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/full/10.1080/07468342.
2022.2119068.

Mitchell:1972:CPT

[Mit72] Lawrence L. Mitchell. Conference proceedings: Teaching mathematics to occupational and developmental students. Two-Year College Mathematics Journal, 3(1):42–47, Spring 1972. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1972.11974196.

Mitchell:1974:IGE

[Mit74] William Mitchell. Improving general education mathematics. Two-Year College Mathematics Journal, 5(2):32–38, Spring 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1974.11974311.

Mitchem:1981:HSF

[Mit81] John Mitchem. On the history and solution of the four-color map problem. Two-Year College Mathematics Journal, 12(2):108–116, March 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972506.

Maier:1973:CRN

[MM73] E. A. Maier and David Maier. A construction of the real numbers. Two-Year College Mathematics Journal, 4(1):31–35, Winter 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1973.11974232.

Murty:1985:GMD

[MM85] Vedula N. Murty and James F. McCrory. A general method of deriving the auxiliary equation for Cauchy-Euler equations. College Mathematics Journal, 16(3):212–215, June 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1985.11972884.

Maron:1993:GIC

[MM93] Melvin J. Maron and Ghansham M. Manwani. A geometric interpretation of the columns of the (pseudo)inverse of A. College Mathematics Journal, 24(1):73–75, January 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1993.11973511.

MacLaughlin:2013:CD

[MM13] Andrew MacLaughlin and Alex Meadows. Chomp in disguise. College Mathematics Journal, 44(4):284–292, September 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.44.4.284.

Mahmood:2014:PWC

[MM14] Munir Mahmood and Ibtihal Mahmood. Proof without words: Completing the square via the difference of squares. College Mathematics Journal, 45(1):21, January 2014. CODEN ????

ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
45.1.021.

Martinjak:2021:PWV

[MM21] Ivica Martinjak and Ana Mimica. Proofs without words: a visual proof for an infinite alternating sign series. College Mathematics Journal, 52(3):204, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2021.1909963.

Martin-Morales:2014:CVS

[MMOM14] Jorge Martín-Morales and Antonio M. Oller-Marcén. Computing volumes of solids of revolution with double integrals. College Mathematics Journal, 45(3):219–221, May 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
45.3.219.

Molokach:2016:PWP

[Mol16] John Molokach. Proof without words: The product-to-sum identities. College Mathematics Journal, 47(3):199, May 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.47.3.199.

Molokach:2017:PWP

[Mol17] John Molokach. Proof without words: The Pythagorean Theorem. College Mathematics Journal, 48(5):334, November 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.48.5.334.

Monzingo:1982:VII

[Mon82] M. G. Monzingo. A visual interpretation of independent events. Two-Year College Mathematics Journal, 13(3):197–198, June 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1982.11972609.

Montgomery:2021:WB

[Mon21] Aaron Montgomery. What’s in the bag? College Mathematics Journal, 52(3):177–184, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2021.1909952.

Moore:1978:ASE

[Moo78a] Charles G. Moore. The anatomy of the stupid error. Two-Year College Mathematics Journal, 9(5):309–310, November 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1978.11974598.

Moore:1978:CSL

[Moo78b] Charles G. Moore. Completing the square — a laboratory approach. Two-Year College Mathematics Journal, 9(4):215–218, September 1978. CODEN ???? ISSN 0049-4925 (print), 2325-

9116 (electronic). URL http://www.tandfonline.com/doi/
abs/10.1080/00494925.1978.11974581.

Moore:1980:EDC

[Moo80] Charles G. Moore. An error-detecting check by substitution. Two-Year College Mathematics Journal, 11(5):326–327, November 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1980.11972471.

Moore:1983:CIC

[Moo83] Charles G. Moore. The construction of integral cevians. Two-Year College Mathematics Journal, 14(4):301–308, September 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1983.11972702.

Moore:1984:SR

[Moo84] Tom Moore. Software reviews. College Mathematics Journal, 15(4):343–344, September 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1984.11972798.

Moore:1985:CFD

[Moo85] Charles G. Moore. Constructing the foci and directrices of a given ellipse. College Mathematics Journal, 16(2):122–128, March 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1985.11972865.

Moore:1986:DRC

[Moo86] Robert C. Moore. Doubling: Real, complex, quaternion, and beyond … well, Maybe. College Mathematics Journal, 17(4):342– 343, September 1986. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1986.11972979.

Moore:1989:VEP

[Moo89] Charles G. Moore. To view an ellipse in perspective. College Mathematics Journal, 20(2):134–136, March 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1989.
11973220.

Moore:1991:UEF

[Moo91] Thomas L. Moore. Using Euler’s formula to solve plane separation problems. College Mathematics Journal, 22(2):125–130, March 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1991.11973368.

Moore:1995:VGH

[Moo95] Robert C. Moore. Visualizing the group homomorphism theorem. College Mathematics Journal, 26(2):143, March 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1995.11973686.

Moodley:2020:MSG

[Moo20] Mervlyn Moodley. Matrix solution to Gergonne’s pile problem. College Mathematics Journal, 51(5):351–357, 2020. CO-DEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/full/10.1080/07468342.
2020.1748951.

Moore:2023:ETL

[Moo23] Matthew N. Moore. An elementary treatment of the Lambert-W relation. College Mathematics Journal, 54(2):130–138, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2023.2188040.

Morman:1973:ATM

[Mor73] Shelba Jean Morman. An audio-tutorial method of instruction vs the traditional lecture-discussion method. Two-Year College Mathematics Journal, 4(3):56–61, Autumn 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1973.
11974274.

Morris:1986:CCT

[Mor86] J. Richard Morris. Counterexamples to a comparison test for alternating series. College Mathematics Journal, 17(2):165–166, March 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1986.11972947.

Moran:1991:GDI

[Mor91] Daniel A. Moran. Graphs and derivatives of the inverse trig functions. College Mathematics Journal, 22(5):417, November 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1991.11973420.

Moretti:2002:MCA

[Mor02] Christopher Moretti. Moving a couch around a corner. College Mathematics Journal, 33(3):196–200, May 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/07468342.
2002.11921940.

Morrill:2004:FMB

[Mor04] John E. Morrill. Five mathematicians, a bunch of coconuts, a monkey, and a coin. College Mathematics Journal, 35(4):256– 257, September 2004. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2004.11922082.

Morgan:2010:FLT

[Mor10a] Frank Morgan. Fermat’s Last Theorem for fractional and irrational exponents. College Mathematics Journal, 41(3):182–185, May 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/074683410X488647.

Morrison:2010:FP

[Mor10b] Kent E. Morrison. The FedEx problem. College Mathematics Journal, 41(3):222–232, May 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.

tandfonline.com/doi/abs/10.4169/074683410X488719.

Morrison:2011:BR

[Mor11] Kent E. Morrison. Book review. College Mathematics Journal, 42(4):336–338, September 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.42.4.336.

Mortici:2015:VEP

[Mor15] Cristinel Mortici. A very elementary proof of Bernoulli’s Inequality. College Mathematics Journal, 46(2):136–137, March 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.46.2.136.

Moreno:2016:SEP

[Mor16] Samuel G. Moreno. A short and elementary proof of the Basel Problem. College Mathematics Journal, 47(2):134–135, March 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.47.2.134.

Moser:1984:SAW

[Mos84] William Moser. A sequel to “another way of looking at n!”. College Mathematics Journal, 15(2):142–143, March 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972764. See [Hsu80].

Mosteller:1988:GAS

[Mos88] Frederick Mosteller. Growth and advances in statistics. College Mathematics Journal, 19(1):15–16, January 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1988.
11973078.

Motto:1975:MA

[Mot75] Vincent J. Motto. Mathematical astronomy. Two-Year College Mathematics Journal, 6(1):21–26, February 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1975.
11974362.

Montuchi:1988:BTE

[MP88] Paolo Montuchi and Warren Page. Behold! Two extremum problems (and the Arithmetic–Geometic Mean Inequality). College Mathematics Journal, 19(4):347, September 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1988.11973136.

Marrero:2003:CTC

[MP03] Osvaldo Marrero and Paul C. Pasles. Coin ToGa: a cointossing game. College Mathematics Journal, 34(3):183–193, May 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2003.11922004.

Minton:2007:DDK

[MP07] Roland Minton and Timothy J. Pennings. Do dogs know bifurcations? College Mathematics Journal, 38(5):356–361, November 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2007.11922260.

Mazzone:2010:ANT

[MP10] Eric F. Mazzone and Bruce R. Piper. Animating nested Taylor polynomials to approximate a function. College Mathematics Journal, 41(5):405–408, November 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.

tandfonline.com/doi/abs/10.4169/074683410X522035.

Martin:2012:OLI

[MP12] Paul Martin and Kirthi Premadasa. The optimal level of insulation in a home attic. College Mathematics Journal, 43(2): 165–168, March 2012. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.43.2.165.

McQuillan:2014:DFS

[MP14] Daniel McQuillan and Rob Poodiack. On the differentiation formulae for sine, tangent, and inverse tangent. College Mathematics Journal, 45(2):140–142, March 2014. CODEN ????

ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
45.2.140.

Marrero:2023:MPJ

[MP23] Osvaldo Marrero and Paul C. Pasles. The multivariate probabilistic Josephus problem. College Mathematics Journal, 54 (5):446–453, 2023. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2023.2266316.

Matrix:1979:MGD

[MR79] Irving Joshua Matrix and Peter Renz. Martin Gardner: Defending the honor of the human mind. Two-Year College Mathematics Journal, 10(4):227–232, September 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1979.
11972357.

Martins:1995:ATF

[MR95] L. F. Martins and I. W. Rodrigues. Angle trisection by fixed point iteration. College Mathematics Journal, 26(3):205–208, May 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1995.11973697.

Martins:1996:SEC

[MR96] L. F. Martins and I. W. Rodrigues. A serendipitous encounter with the Cantor ternary function. College Mathematics Journal, 27(3):193–198, May 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1996.11973775.

Mulcahy:1998:FAS

[MR98] Colm Mulcahy and John Rossi. A fresh approach to the singular value decomposition. College Mathematics Journal, 29(3):199– 207, May 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1998.11973940.

McShane:2003:DDU

[MR03] Janet M. McShane and Michael I. Ratliff. Dice distributions using combinatorics, recursion, and generating functions. College Mathematics Journal, 34(5):370–376, November 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2003.11922033.

Maharam:1976:WD

[MS76] Lewis G. Maharam and Edward P. Shaughnessy. When does (fg)′ = f′g′? Two-Year College Mathematics Journal, 7(1): 38–39, February 1976. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1976.11974420.

Milles:1978:SFE

[MS78] Stephen J. Milles and Henry J. Schultz. Some functional equations for the calculus student. Two-Year College Mathematics Journal, 9(4):205–209, September 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1978.
11974579.

Mureika:1984:ACS

[MS84] R. A. Mureika and R. D. Small. An almost correct series. College Mathematics Journal, 15(4):334–338, September 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972796.

Moore:1987:MH

[MS87] L. C. Moore and David A. Smith. Media highlights. College Mathematics Journal, 18(5):433–442, November 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1987.11973071.

Manuel:1988:UAG

[MS88] George Manuel and Amalia Santiago. An unexpected appearance of the Golden ratio. College Mathematics Journal, 19(2):168–170, March 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1988.11973107.

Mathews:1989:RI

[MS89] John H. Mathews and Harris S. Shultz. Riemann integral of cos x. College Mathematics Journal, 20(3):237, May 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1989.11973239.

Matthews:1991:CMP

[MS91] Paul C. Matthews and Steven H. Strogatz. Chaotic mappings and probability distributions. College Mathematics Journal, 22 (1):45–47, January 1991. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1991.11973358.

Muller:1994:SR

[MS94] Eric R. Muller and K. J. Srivastava. Software reviews. College Mathematics Journal, 25(1):56–63, January 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1994.
11973584.

Moraleda:2012:LWD

[MS12] Jorge Moraleda and David G. Stork. Lake Wobegon dice. College Mathematics Journal, 43(2):152–159, March 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
43.2.152.

Mark:2016:WYW

[MS16] Melissa Mark and Michael Schramm. When you wander off on a tangent, where do you end up? College Mathematics Journal, 47(5):334–339, November 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.47.5.334.

Miller:2017:IPD

[MS17a] Steven J. Miller and Cesar E. Silva. If a prime divides a product …. College Mathematics Journal, 48(2):123–128, March 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.48.2.123.

Millett:2017:ELC

[MS17b] Julie Millett and Xingping Sun. Existence of limits and continuity. College Mathematics Journal, 48(1):31–35, January 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.48.1.31.

Mixon:2018:SIO

[MS18] Dustin G. Mixon and James Solazzo. A short introduction to optimal line packings. College Mathematics Journal, 49(2): 82–91, March 2018. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2018.1421364.

Miles:2019:CIA

[MS19] William W. Miles and Sammi E. Smith. The choking index: An analysis of performance under pressure on the PGA Tour. College Mathematics Journal, 50(4):260–271, 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2019.1644102.

May:2020:PCC

[MS20] Dan May and Dan Swenson. The proportion of comets in the card game SET. College Mathematics Journal, 51(3):162– 172, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2020.1733893.

Muleshkov:2022:EF

[MS22] Angel S. Muleshkov and Kurt R. Sweat. Elementary functions. College Mathematics Journal, 53(1):54–63, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2022.
1990675.

Moran:2023:KTL

[MS23] Molly Moran and Jessica Schenkman. Knot theory in linear algebra: an example with trip matrices. College Mathematics Journal, 54(3):186–194, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2023.2201163.

Myers:1985:SR

[MSFA85] Roy E. Myers, Carol M. Smith, Marjorie A. Fitting, and Henry Africk. Software reviews. College Mathematics Journal, 16(4): 301–303, September 1985. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1985.11972902.

Miller:2021:WRM

[MST21] Steven J. Miller, Haoyu Sheng, and Daniel Turek. When rooks miss: Probability through chess. College Mathematics Journal, 52(2):82–93, 2021. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2021.1886774.

Mueller:1997:DDE

[MT97] William Mueller and Richard Thompson. Discovering differential equations in optics. College Mathematics Journal, 28(3):217–223, May 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1997.11973868.

Millis:2019:ATE

[MT19] John P. Millis and Courtney K. Taylor. Areas of a total eclipse. College Mathematics Journal, 50(5):375–377, 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2019.1667201.

Miller:2022:STP

[MT22] Steven J. Miller and Enrique Treviño. On the sum of k -th powers in terms of earlier sums. College Mathematics Journal, 53(3):220–225, 2022. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2022.2040236.

Markowsky:2023:TAS

[MT23] Greg Markowsky and David Treeby. Trigonometric angle-sum formulas and conformal maps. College Mathematics Journal, 54(3):195–199, 2023. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2023.2201563.

Muench:2007:IAT

[Mue07] Donald L. Muench. An iterative angle trisection. College Mathematics Journal, 38(2):82–84, March 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2007.
11922222.

Muench:2020:LTD

[Mue20] Donald L. Muench. The last two days in elementary linear algebra. College Mathematics Journal, 51(3):222–224, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2020.1740540.

Mukherjee:2024:PWG

[Muk24] Rajib Mukherjee. Proof without words: Geometric evaluation of another limit. College Mathematics Journal, 55(5):429, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic).

Mulder:2019:GT

[Mul19] Kenneth Mulder. Geek tragedy. College Mathematics Journal, 50(2):133, 2019. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2019.1559549.

Munasinghe:2004:SLD

[Mun04] Ranjith Munasinghe. Some linear differential equations forget that they have variable coefficients. College Mathematics Journal, 35(1):22–25, January 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2004.11922046.

Murty:1977:II

[Mur77] V. N. Murty. Independence and intuition. Two-Year College Mathematics Journal, 8(2):106–107, March 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1977.
11974492.

Murty:1980:HCR

[Mur80a] V. N. Murty. How close are the Riemann sums to the integral they approximate? Two-Year College Mathematics Journal, 11 (4):268–270, September 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1980.11972455.

Murty:1980:IP

[Mur80b] V. N. Murty. Integration by parts. Two-Year College Mathematics Journal, 11(2):90–94, March 1980. CODEN ????

ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1980.
11972411.

Murty:1980:NDC

[Mur80c] V. N. Murty. A note on the derivative of a composite function. Two-Year College Mathematics Journal, 11(1):50, January 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1980.11972400.

Murty:1981:DSF

[Mur81] V. N. Murty. On dice-sum frequencies. Two-Year College Mathematics Journal, 12(3):209–211, June 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1981.
11972529.

Murty:1984:MSD

[Mur84] Vedula N. Murty. On the mean and standard deviation of a random sample. College Mathematics Journal, 15(1):60–63, January 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972751.

Murdoch:1995:SCC

[Mur95] Timothy Murdoch. Spherical coordinates from cylindrical coordinates on a torus. College Mathematics Journal, 26(5):385– 387, November 1995. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1995.11973735.

Murray:2010:BR

[Mur10] Margaret A. M. Murray. Book review. College Mathematics Journal, 41(3):248–251, May 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/074683410X488755.

Muses:1979:GI

[Mus79] C. Musès. Geometric interpretations of Sinφ_{1} + Sinφ_{2} = 1. Two-Year College Mathematics Journal, 10(5):350–351, November 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1979.11972385.

Mellinger:2012:SF

[MV12] Keith E. Mellinger and Raymond Viglione. The spider and the fly. College Mathematics Journal, 43(2):169–172, March 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.43.2.169.

McAllister:1981:OS

[MW81] Byron L. McAllister and J. Eldon Whitesitt. Is Ln the other shoe? Two-Year College Mathematics Journal, 12(1):20–23, January 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972484.

Muench:1983:LAF

[MW83] Donald L. Muench and Gerald Wildenberg. A logarithm algorithm for a five-function calculator. Two-Year College Mathematics Journal, 14(4):324–326, September 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1983.
11972707.

Meier:1997:VGL

[MW97] John Meier and Jessica Wolfson. Visualizing the geometry of Lissajous knots. College Mathematics Journal, 28(3):211–216, May 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1997.11973867.

McCullough:2003:REP

[MW03] Darryl McCullough and Elizabeth Wade. Recursive enumeration of Pythagorean triples. College Mathematics Journal, 34(2): 107–111, March 2003. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2003.11921992.

McDevitt:2018:SPP

[MW18] Timothy McDevitt and Angela Wesneski. A simple probability paradox. College Mathematics Journal, 49(4):292–294, September 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2018.1483661.

Myers:1977:SER

[Mye77] Roy E. Myers. Some elementary results related to the mean value theorem. Two-Year College Mathematics Journal, 8(1):51–53, January 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1977.11974482.

Marchisotto:1994:IIF

[MZ94] Elena Anne Marchisotto and Gholam-Ali Zakeri. An invitation to integration in finite terms. College Mathematics Journal, 25 (4):295–308, September 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1994.11973625.

Martinez:2024:GRA

[MZ24] Lucy Martinez and Doron Zeilberger. A guide to the riskaverse gambler and resolving the St. Petersburg Paradox once and for all. College Mathematics Journal, 55(3):226–234, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2024.2305103.

Nacin:2014:PP

[Nac14] David Nacin. Permudoku puzzle. College Mathematics Journal, 45(4):307, September 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.45.4.307.

Nacin:2015:MACa

[Nac15a] David Nacin. MAA 100th anniversary CMJ Puzzle A. College Mathematics Journal, 46(4):254, September 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
46.4.254.

Nacin:2015:MACb

[Nac15b] David Nacin. MAA 100th anniversary CMJ Puzzle C. College Mathematics Journal, 46(4):263, September 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
46.4.263.

Nacin:2015:MACc

[Nac15c] David Nacin. MAA 100th anniversary CMJ Puzzle J. College Mathematics Journal, 46(4):274, September 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
46.4.274.

Nacin:2015:MACd

[Nac15d] David Nacin. MAA 100th anniversary CMJ Puzzle M. College Mathematics Journal, 46(4):294, September 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
46.4.294.

Nacin:2016:MACa

[Nac16a] David Nacin. MAA 101st anniversary CMJ Puzzle A. College Mathematics Journal, 47(4):249, September 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
47.4.249.

Nacin:2016:MACb

[Nac16b] David Nacin. MAA 101st anniversary CMJ Puzzle C. College Mathematics Journal, 47(4):254, September 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
47.4.254.

Nacin:2016:MACc

[Nac16c] David Nacin. MAA 101st anniversary CMJ Puzzle J. College Mathematics Journal, 47(4):264, September 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
47.4.264.

Nacin:2016:MACd

[Nac16d] David Nacin. MAA 101st anniversary CMJ Puzzle M. College Mathematics Journal, 47(4):274, September 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
47.4.274.

Nacin:2016:MACe

[Nac16e] David Nacin. MAA 101st anniversary CMJ puzzles solutions. College Mathematics Journal, 47(4):293, September 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.47.4.293.

Nacin:2017:DPa

[Nac17a] David Nacin. Dihedoku puzzle 1. College Mathematics Journal, 48(4):248, September 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.48.4.248.

Nacin:2017:DPb

[Nac17b] David Nacin. Dihedoku puzzle 2. College Mathematics Journal, 48(4):258, September 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.48.4.258.

Nacin:2017:DPc

[Nac17c] David Nacin. Dihedoku puzzle 3. College Mathematics Journal, 48(4):264, September 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.48.4.264.

Nacin:2017:DPS

[Nac17d] David Nacin. Dihedoku puzzles solutions. College Mathematics Journal, 48(4):283, September 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
48.4.283.

Nacin:2017:CKP

[Nac17e] David Nacin. On a complex KenKen problem. College Mathematics Journal, 48(4):274–282, September 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
48.4.274.

Nacin:2018:KPa

[Nac18a] David Nacin. Knightdoku puzzle 1. College Mathematics Journal, 49(4):250, September 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2018.1502549.

Nacin:2018:KPb

[Nac18b] David Nacin. Knightdoku puzzle 2. College Mathematics Journal, 49(4):261, September 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2018.1502550.

Nacin:2018:KPc

[Nac18c] David Nacin. Knightdoku puzzle 3. College Mathematics Journal, 49(4):269, September 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2018.1502551.

Nacin:2018:KPd

[Nac18d] David Nacin. Knightdoku puzzle 4. College Mathematics Journal, 49(4):277, September 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2018.1502552.

Nacin:2018:KPS

[Nac18e] David Nacin. Knightdoku puzzles solutions. College Mathematics Journal, 49(4):287, September 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2018.
1502553.

Nacin:2022:KK

[Nac22] David Nacin. Kleinian KenKen. College Mathematics Journal, 53(4):253–264, 2022. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2022.2099702.

Nagahashi:2021:MCT

[Nag21] Hideo Nagahashi. Magic card tricks on Hamming codes over finite fields. College Mathematics Journal, 52(4):281–288, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2021.1942676.

Nakhli:1986:BVA

[Nak86] Fouad Nakhli. Behold! The vertex angles of a star sum to 180^{∘}. College Mathematics Journal, 17(4):338, September 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1986.11972975.

Nam:2024:APU

[Nam24] Kyumin Nam. Approximation of pi using infinite series of arcsine. College Mathematics Journal, 55(2):172–174, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2023.2242213.

Nanyes:2014:CDC

[Nan14] Ollie Nanyes. Continuously differentiable curves detect limits of functions of two variables. College Mathematics Journal, 45(1): 54–56, January 2014. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.45.1.054.

Nash:1985:DEB

[Nas85] David H. Nash. Differential equations and the Battle of Trafalgar. College Mathematics Journal, 16(2):98–102, March 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1985.11972861.

Nash:2023:NTC

[Nas23] David A. Nash. A new take on classic ‘pen problems’. College Mathematics Journal, 54(4):345–354, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2023.
2237843.

Nathan:1991:RFC

[Nat91] Kathirgama Nathan. Review: Fractals for the Classroom. Part One: Introduction to Fractals and Chaos by H. O. Peitgen, H. Jürgens, D. Saupe; Fractals for the Classroom. Part Two: Complex Systems and Mandelbrot Set by H. O. Peitgen, H. Jürgens, D. Saupe. College Mathematics Journal, 22(5):455–456, November 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.jstor.org/stable/2686611.

Natian:2019:PFF

[Nat19] Albert Natian. Pinpoint the flitting fly. College Mathematics Journal, 50(5):351–356, 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2019.1664248.

Nathan:2024:SSP

[Nat24] Joseph Amal Nathan. A simple and short proof of det(AB) = det(A) det(B). College Mathematics Journal, 55(3):253–254, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2024.2302291.

Nauenberg:1994:NPI

[Nau94] M. Nauenberg. Newton’s Principia and inverse-square orbits. College Mathematics Journal, 25(3):212–222, May 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1994.11973610.

Neal:1993:DSS

[Nea93] David Neal. Determining sample sizes for Monte Carlo integration. College Mathematics Journal, 24(3):254–259, May 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1993.11973541.

Neal:1994:SPL

[Nea94] David Neal. The series ∑_{n=1}^{∞} n^{m}x^{n} and a Pascal-like triangle. College Mathematics Journal, 25(2):99–101, March 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1994.11973589.

Needleman:2013:BLP

[Nee13] Jonathan Needleman. Boggle logic puzzles: Minimal solutions. College Mathematics Journal, 44(4):293–299, September 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.44.4.293.

Neff:1982:DTP

[Nef82] John D. Neff. Dice tossing and Pascal’s Triangle. Two-Year College Mathematics Journal, 13(5):311–314, November 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1982.11972638.

Neff:1983:IM

[Nef83] John D. Neff. Imbedding the metric. Two-Year College Mathematics Journal, 14(3):197–202, June 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1983.
11972682.

Neidinger:1989:ADA

[Nei89] Richard D. Neidinger. Automatic differentiation and APL. College Mathematics Journal, 20(3):238–251, May 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/07468342.
1989.11973240.

Nelson:1977:MOT

[Nel77] Eric J. Nelson. A machine-oriented technique for the complete solution of linear systems. Two-Year College Mathematics Journal, 8(3):161–164, June 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1977.11974502.

Nelson:1979:PPP

[Nel79] Robert Nelson. Pictures, probability, and paradox. Two-Year College Mathematics Journal, 10(3):182–190, June 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1979.11972348.

Nelsen:1988:RBH

[Nel88] Roger B. Nelsen. The relationship between hyperbolic and exponential functions. College Mathematics Journal, 19(1):54–56, January 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1988.11973091. See also [Nel92].

Nelsen:1989:DAF

[Nel89a] Roger B. Nelsen. The double-angle formulas. College Mathematics Journal, 20(1):51, January 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1989.
11973203.

Nelsen:1989:RMS

[Nel89b] Roger B. Nelsen. The root mean square-arithmetic mean– geometric mean-harmonic mean inequality. College Mathematics Journal, 20(3):231, May 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1989.11973236.

Nelsen:1992:RBH

[Nel92] Roger B. Nelsen. The relationship between hyperbolic and exponential functions — revisited. College Mathematics Journal, 23 (3):207–208, May 1992. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1992.11973458. See [Nel88].

Nelsen:1994:PS

[Nel94] Roger B. Nelsen. Problems and solutions. College Mathematics Journal, 25(3):240–246, May 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1994.
11973619.

Nelsen:1995:SI

[Nel95] Roger Nelsen. Symmetry and integration. College Mathematics Journal, 26(1):39–41, January 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1995.
11973665.

Nelsen:2012:PWA

[Nel12] Roger B. Nelsen. Proof without words: an alternating series. College Mathematics Journal, 43(5):370, November 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.43.5.370.

Nelsen:2013:PWS

[Nel13] Roger B. Nelsen. Proof without words: Squares modulo 3. College Mathematics Journal, 44(4):283, September 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
44.4.283.

Nelsen:2014:PWA

[Nel14a] Roger Nelsen. Proof without words: Alternating sums of consecutive squares. College Mathematics Journal, 45(1):16, January 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.45.1.016.

Nelsen:2014:PWP

[Nel14b] Roger Nelsen. Proof without words: Pythagorean quadruples. College Mathematics Journal, 45(3):179, May 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
45.3.179.

Nelsen:2014:PWSb

[Nel14c] Roger B. Nelsen. Proof without words: a sine identity for triangles. College Mathematics Journal, 45(5):376, November 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.45.5.376.

Nelsen:2014:PWSa

[Nel14d] Roger B. Nelsen. Proof without words: Summing squares by counting triangles. College Mathematics Journal, 45(5):349, November 2014. CODEN ???? ISSN 0746-8342 (print), 1931-

1346 (electronic). URL http://www.tandfonline.com/doi/
abs/10.4169/college.math.j.45.5.349.

Nelsen:2015:PWT

[Nel15a] Roger B. Nelsen. Proof without words: a trigonometric proof of the arithmetic mean–geometric mean inequality. College Mathematics Journal, 46(1):42, January 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
46.1.42.

Nelsen:2015:PWC

[Nel15b] Roger B. Nelsen. Proof without words: Centered triangular numbers. College Mathematics Journal, 46(5):335, November 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.46.5.335.

Nelsen:2015:PWE

[Nel15c] Roger B. Nelsen. Proof without words: Each but two triangular numbers is a sum of three triangular numbers. College Mathematics Journal, 46(3):172, May 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.46.3.172.

Nelsen:2015:PWS

[Nel15d] Roger B. Nelsen. Proof without words: Sums of every third triangular number. College Mathematics Journal, 46(2):98, March 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.46.2.98.

Nelsen:2015:PWA

[Nel15e] Roger B. Nelsen. Proof without words: The area of a regular dodecagon. College Mathematics Journal, 46(1):10, January 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.46.1.10.

Nelsen:2016:PWR

[Nel16a] Roger B. Nelsen. Proof without words: a right triangle identity. College Mathematics Journal, 47(5):355, November 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.47.5.355.

Nelsen:2016:PWSa

[Nel16b] Roger B. Nelsen. Proof without words: a surprising integer result. College Mathematics Journal, 47(2):94, March 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.47.2.94.

Nelsen:2016:PWN

[Nel16c] Roger B. Nelsen. Proof without words: Nearly cubic Pythagorean boxes. College Mathematics Journal, 47(3):190, May 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.47.3.190.

Nelsen:2016:PWPb

[Nel16d] Roger B. Nelsen. Proof without words: Perfect numbers and triangular numbers. College Mathematics Journal, 47(3):171, May 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.47.3.171.

Nelsen:2016:PWSb

[Nel16e] Roger B. Nelsen. Proof without words: Square triangular numbers and almost isosceles Pythagorean triples. College Mathematics Journal, 47(3):179, May 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.47.3.179.

Nelsen:2016:PWG

[Nel16f] Roger B. Nelsen. Proof without words: The Golden ratio. College Mathematics Journal, 47(2):108, March 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
47.2.108.

Nelsen:2016:PWPa

[Nel16g] Roger B. Nelsen. Proof without words: The pentagon–hexagon– decagon identity. College Mathematics Journal, 47(1):10, January 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.47.1.10.

Nelsen:2017:PWN

[Nel17a] Roger B. Nelsen. Proof without words: Nested square roots. College Mathematics Journal, 48(3):204, May 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
48.3.204.

Nelsen:2017:PWP

[Nel17b] Roger B. Nelsen. Proof without words: Perfect numbers modulo 7. College Mathematics Journal, 48(1):17, January 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.48.1.17.

Nelsen:2018:FIS

[Nel18] Roger B. Nelsen. Further integral skinning with applications. College Mathematics Journal, 49(5):327–332, November 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2018.1507383.

Nelsen:2019:PWP

[Nel19] Roger B. Nelsen. Proof without words: Pythagorean primes and triangular numbers. College Mathematics Journal, 50(5): 378, 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2019.1664854.

Nelsen:2021:MFP

[Nel21] Roger Nelsen. Machin’s formula via a proof without words. College Mathematics Journal, 52(5):355, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2021.
1969860.

Nelsen:2022:PWC

[Nel22a] Roger Nelsen. Proof without words: Consecutive powerful numbers. College Mathematics Journal, 53(4):318, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2022.2099716.

Nelsen:2022:PWT

[Nel22b] Roger Nelsen. Proof without words: Tangents of multiples of π/24.CollegeMathematicsJournal, 53(3) : 226, 2022.CODEN????ISSN0746−

8342(print), 1931−1346(electronic).URLhttpwwwtandfonlinecomdoifull1010800746834220222022904.

Nelsen:2023:HHI

[Nel23] Roger Nelsen. The Hermite–Hadamard inequalities and applications. College Mathematics Journal, 54(5):482–484, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2023.2273181.

Nelsen:2024:PWC

[Nel24a] Roger Nelsen. Proof without words: Cousin prime products modulo 72. College Mathematics Journal, 55(3):245, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2023.2280420.

Nelsen:2024:SSC

[Nel24b] Roger Nelsen. Summing squares and cubes using triangular numbers. College Mathematics Journal, 55(2):169–171, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2023.2264720.

Nelsen:2024:IJN

[Nel24c] Roger B. Nelsen. An introduction to the Jacobsthal numbers. College Mathematics Journal, 55(2):153–158, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2023.2282201.

Nelsen:2024:MSL

[Nel24d] Roger B. Nelsen. Morrie’s Second Law and its consequences. College Mathematics Journal, 55(5):376–380, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic).

Nelsen:2025:PWP

[Nel25] Roger B. Nelsen. Proof without words: a product inequality for pi and the golden ratio. College Mathematics Journal, 56 (1):62, 2025. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic).

Neger:2003:CCC

[NF03] Nial Neger and Michael Frame. Clarifying compositions with cobwebs. College Mathematics Journal, 34(3):196–204, May 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2003.11922006.

Ni:2016:ILF

[Ni16] Xianglong Ni. Iterating the logarithmic function. College Mathematics Journal, 47(3):172–178, May 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
47.3.172.

Nicol:1993:IPS

[Nic93] Sherrie J. Nicol. Integrals of products of sine and cosine with different arguments. College Mathematics Journal, 24(2):158–160, March 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1993.11973521.

Niedra:1978:GSG

[Nie78] Andris Niedra. Geometric series on the gridiron. Two-Year College Mathematics Journal, 9(1):18–20, January 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1978.11974540.

Nievergelt:1991:CCC

[Nie91] Yves Nievergelt. Classroom computer capsule: Fractals illustrate the mathematical way of thinking. College Mathematics Journal, 22(1):60–64, January 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1991.11973361.

Nievergelt:1994:CMT

[Nie94] Yves Nievergelt. Calculus measures tank capacity and avoids oil spills. College Mathematics Journal, 25(2):132–136, March 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1994.11973596.

Nievergelt:2001:WRF

[Nie01] Yves Nievergelt. Well-rounded figures. College Mathematics Journal, 32(1):30–32, January 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2001.
11921848.

Nievergelt:2003:HSQ

[Nie03] Yves Nievergelt. How (not) to solve quadratic equations. College Mathematics Journal, 34(2):90–104, March 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2003.
11921990.

Nievergelt:2021:DIS

[Nie21] Yves Nievergelt. Is doom the inescapable solution of initial value problems? College Mathematics Journal, 52(4):302– 305, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2021.1942677.

Niven:1972:WL

[Niv72] Ivan Niven. Which is larger, e^{π} or π^{e} ? Two-Year College Mathematics Journal, 3(2):13–15, Autumn 1972. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1972.
11974209.

Niven:1979:EPB

[Niv79] Ivan Niven. An ellipse problem beyond the reach of calculus. Two-Year College Mathematics Journal, 10(3):162–168, June 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1979.11972344.

Niven:1986:MH

[Niv86] Ivan Niven. Media highlights. College Mathematics Journal, 17 (3):267–275, May 1986. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1986.11972969.

Nirode:2022:DP

[NK22] Wayne Nirode and Norm Krumpe. Donuts with Pythagoras. College Mathematics Journal, 53(4):306–311, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/full/10.1080/07468342.
2022.2099705.

Norton:1984:CRM

[NL84] Alec Norton and Benjamin Lotto. Complex roots made visible. College Mathematics Journal, 15(3):248–249, June 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972777.

Nord:1996:MRL

[NM96] John Nord and Edward Miller. Mercator’s rhumb lines: a multivariable application of arc length. College Mathematics Journal, 27(5):384–387, November 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1996.11973814.

Nathan:1991:MH

[NN91] Kathirgama Nathan and Kathirgama Nathan. Media highlights. College Mathematics Journal, 22(5):450–457, November 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1991.11973426.

Ngo:1992:TLT

[NO92] V. Ngo and S. Ouzomgi. Teaching the Laplace transform using diagrams. College Mathematics Journal, 23(4):309–312, September 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1992.11973476.

Noether:1988:MH

[Noe88] Gottfried E. Noether. Media highlights. College Mathematics Journal, 19(1):88–95, January 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1988.
11973099.

Nogin:2017:OPA

[Nog17] Maria Nogin. Optimizing prisms of all shapes and dimensions. College Mathematics Journal, 48(3):199–203, May 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
48.3.199.

Norman:1988:PGL

[Nor88] F. Alexander Norman. Parameter-generated loci of critical points of polynomials. College Mathematics Journal, 19(3):223–229, May 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1988.11973118.

Norton:2009:EET

[Nor09] Anderson Norton. Eighty-eight thousand, four hundred and eighteen (more) ways to fill space. College Mathematics Journal, 40(2):108–112, March 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2009.11922345.

Northshield:2017:TSP

[Nor17] Sam Northshield. Two short proofs of the infinitude of primes. College Mathematics Journal, 48(3):214–216, May 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.48.3.214.

Nott:1971:NRR

[Not71] Maurice E. Nott, Jr. New results of research comparing programmed and lecture-text instruction. Two-Year College Mathematics Journal, 2(1):19–22, Spring 1971. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1971.
11973991.

Norman:2022:VTV

[NP22] Sandy Norman and Priya V. Prasad. Variations on a theorem of Varignon. College Mathematics Journal, 53(5):351– 360, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2022.2119792.

Nesky:2016:DSB

[NR16] Amy Nesky and Clara Redwood. Discrete and smooth bicycle “unicycle” paths. College Mathematics Journal, 47(3):180–189, May 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.47.3.180.

Nogueira:2020:SPC

[NRT20] Joaquim Nogueira, Fátima Rodrigues, and Luís Trabucho. Some probability calculations concerning the Egyptian game Senet. College Mathematics Journal, 51(4):271–283, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2020.1776569.

Nelsen:1987:PSR

[NS87] Roger B. Nelsen and James E. Schultz. The probability that the “sum of the rounds” equals the “round of the sum”. College Mathematics Journal, 18(5):390–396, November 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.jstor.org/stable/2686963;
 http://www.tandfonline.com/doi/abs/10.1080/07468342.
1987.11973061.

Needleman:2017:UGS

[NS17] Jonathan Needleman and Felicia Sciortino. UFOs in the game SET: Looking for airplanes and spaceships. College Mathematics Journal, 48(4):249–257, September 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
48.4.249.

Nunemacher:2010:BRC

[Nun10] Jeffrey Nunemacher. Book review: The Calculus of Friendship: What a Teacher and a Student Learned about Life While Corresponding about Math, by Steven Strogatz, Princeton University Press, 2009. College Mathematics Journal, 41(1):74–76, January 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/074683410X475164.

Ngo:1998:WCI

[NW98] Viet Ngo and Saleem Watson. Who cares if X^{2} + 1 = 0 has a solution? College Mathematics Journal, 29(2):141–144, March 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1998.11973929.

Naylor:2010:BCR

[NW10] Michael Naylor and Brian Winkel. Biangular coordinates redux: Discovering a new kind of geometry. College Mathematics Journal, 41(1):29–41, January 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/074683410X475092.

Nymann:1983:APF

[Nym83] J. E. Nymann. An alternative for partial fractions (part of the time). Two-Year College Mathematics Journal, 14(1):60–61, January 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1983.11972656.

Neubauer:2003:AE

[NZ03] Michael G. Neubauer and Joel Zeitlin. Apportionment and the 2000 election. College Mathematics Journal, 34(1):2–10, January 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2003.11921977.

Orosi:2021:EDD

[OAHS21] Greg Orosi, Ricardo Alfaro, Lixing Han, and Kenneth Schilling. An elementary derivation of the duration of play in the gambler’s ruin problem. College Mathematics Journal, 52(4):299– 301, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2021.1942679.

Ouellette:1977:STF

[OB77] Hugh Ouellette and Gordon Bennett. Stirling’s triangle of the first kind-absolute value style. Two-Year College Mathematics Journal, 8(4):195–202, September 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1977.
11974509.

Ouellette:1979:DGE

[OB79] Hugh Ouellette and Gordon Bennett. The discovery of a generalization: an example in problem solving. Two-Year College Mathematics Journal, 10(2):100–106, March 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1979.
11972332.

OBrien:1982:A

[O’B82] Katharine O’Brien. Aftermath. Two-Year College Mathematics Journal, 13(3):185, June 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1982.11972604.

Oman:2018:PS

[OC18] Greg Oman and Charles N. Curtis. Problems and solutions. College Mathematics Journal, 49(5):371–377, November 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2018.1531656.

Oman:2019:PSa

[OC19a] Greg Oman and Charles N. Curtis. Problems and solutions. College Mathematics Journal, 50(1):61–67, January 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2019.1537453.

Oman:2019:PSb

[OC19b] Greg Oman and Charles N. Curtis. Problems and solutions. College Mathematics Journal, 50(2):143–149, 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2019.1561123.

Oman:2019:PSc

[OC19c] Greg Oman and Charles N. Curtis. Problems and solutions. College Mathematics Journal, 50(3):224–232, 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2019.1601453.

Oman:2019:PSd

[OC19d] Greg Oman and Charles N. Curtis. Problems and solutions. College Mathematics Journal, 50(4):301–306, 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2019.1656482.

Oman:2019:PSe

[OC19e] Greg Oman and Charles N. Curtis. Problems and solutions. College Mathematics Journal, 50(5):379–384, 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2019.1667205.

Oman:2020:PSa

[OC20a] Greg Oman and Charles N. Curtis. Problems and solutions. College Mathematics Journal, 51(1):66–72, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2020.
1680235.

Oman:2020:PSb

[OC20b] Greg Oman and Charles N. Curtis. Problems and solutions. College Mathematics Journal, 51(2):146–152, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2020.1700677.

Oman:2020:PSc

[OC20c] Greg Oman and Charles N. Curtis. Problems and solutions. College Mathematics Journal, 51(3):225–232, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2020.1735882.

Oman:2020:PSd

[OC20d] Greg Oman and Charles N. Curtis. Problems and solutions. College Mathematics Journal, 51(4):305–312, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2020.1802179.

Oman:2020:PSe

[OC20e] Greg Oman and Charles N. Curtis. Problems and solutions. College Mathematics Journal, 51(5):386–392, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2020.1826771.

Oman:2021:PSa

[OC21a] Greg Oman and Charles N. Curtis. Problems and solutions. College Mathematics Journal, 52(1):64–72, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2020.
1853981.

Oman:2021:PSb

[OC21b] Greg Oman and Charles N. Curtis. Problems and solutions. College Mathematics Journal, 52(2):142–148, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2021.1872292.

Oman:2021:PSc

[OC21c] Greg Oman and Charles N. Curtis. Problems and solutions. College Mathematics Journal, 52(3):227–232, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2021.1909994.

Oman:2021:PSd

[OC21d] Greg Oman and Charles N. Curtis. Problems and solutions. College Mathematics Journal, 52(4):306–315, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2021.1949221.

Oman:2021:PSe

[OC21e] Greg Oman and Charles N. Curtis. Problems and solutions. College Mathematics Journal, 52(5):388–396, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2021.1969181.

Oman:2022:PSa

[OC22a] Greg Oman and Charles N. Curtis. Problems and solutions. College Mathematics Journal, 53(1):69–76, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2021.
1991163.

Oman:2022:PSb

[OC22b] Greg Oman and Charles N. Curtis. Problems and solutions. College Mathematics Journal, 53(2):152–160, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2022.2026088.

Oman:2022:PSc

[OC22c] Greg Oman and Charles N. Curtis. Problems and solutions. College Mathematics Journal, 53(3):237–244, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2022.2040871.

Oman:2022:PSd

[OC22d] Greg Oman and Charles N. Curtis. Problems and solutions. College Mathematics Journal, 53(4):319–325, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2022.2100155.

Oman:2022:PSe

[OC22e] Greg Oman and Charles N. Curtis. Problems and solutions. College Mathematics Journal, 53(5):400–407, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2022.2127287.

Oman:2023:PSa

[OC23a] Editors Greg Oman and Charles N. Curtis. Problems and solutions. College Mathematics Journal, 54(1):70–71, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2023.2172918.

Oman:2023:PSb

[OC23b] Editors Greg Oman and Charles N. Curtis. Problems and solutions. College Mathematics Journal, 54(2):147–160, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2023.2186094.

Oman:2023:PSc

[OC23c] Editors Greg Oman and Charles N. Curtis. Problems and solutions. College Mathematics Journal, 54(4):394–399, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2023.2229224.

Oman:2023:PSd

[OC23d] Editors Greg Oman and Charles N. Curtis. Problems and solutions. College Mathematics Journal, 54(4):400–405, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2023.2237385.

Oman:2023:PS

[OC23e] Greg Oman and Charles N. Curtis. Problems and solutions. College Mathematics Journal, 54(5):492–499, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2023.2271821.

OHalloran:1985:JDE

[O’H85] P. J. O’Halloran. Josephine Dianne Edwards. College Mathematics Journal, 16(5):433, November 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1985.
11972922.

Osler:2007:SCB

[OHC07] Thomas J. Osler, Abdulkadir Hassen, and Tirupathi R. Chandrupatla. Surprising connections between partitions and divisors. College Mathematics Journal, 38(4):278–287, September 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2007.11922249.

Okumura:2022:SPI

[Oku22] Hiroshi Okumura. A sangaku problem involving two circles and an equilateral triangle. College Mathematics Journal, 53(4):292– 297, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2022.2099703.

OLeary:1998:TPR

[O’L98] Mike O’Leary. Taylor polynomials for rational functions. College Mathematics Journal, 29(3):226–228, May 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/07468342.
1998.11973946.

Olinick:1973:SSL

[Oli73] Michael Olinick. Solving systems of linear differential equations. Two-Year College Mathematics Journal, 4(1):26–30, Winter 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1973.11974231.

Olmsted:1973:RIO

[Olm73] John M. H. Olmsted. Riemann integration in ordered fields. Two-Year College Mathematics Journal, 4(2):34–40, Spring 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1973.11974249.

Olson:1989:WTP

[Ols89] Lloyd Olson. Where there is pattern, there is significance. College Mathematics Journal, 20(4):321, September 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1989.11973250.

Olson:1996:AWG

[Ols96] David Olson. Another way to graph a sequence. College Mathematics Journal, 27(3):208–209, May 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1996.
11973779.

Olson:2013:BRL

[Ols13] Richard Olson. Book review: The Lost Millennium: History’s Timetables under Siege (2nd edition), by Florin Diacu. College Mathematics Journal, 44(1):62–63, January 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
44.1.062.

Oman:2009:IAS

[Oma09] Greg Oman. An independent axiom system for the real numbers. College Mathematics Journal, 40(2):78–86, March 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2009.11922341.

Olinick:1976:DTP

[OP76] Michael Olinick and Bruce B. Peterson. Darboux’s Theorem and points of inflection. Two-Year College Mathematics Journal, 7(3):5–9, September 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1976.11974436.

ORourke:2015:BR

[O’R15] Joseph O’Rourke. Book review. College Mathematics Journal, 46(4):309–314, September 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.46.4.309.

Osserman:1981:GD

[OS81] Robert Osserman and Gideon Schwartz. Geometry is dead. Two-Year College Mathematics Journal, 12(1):33, January 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972489.

Oman:2016:CCG

[OS16] Greg Oman and Victoria Slattum. A characterization of the cyclic groups by subgroup indices. College Mathematics Journal, 47(1):29–33, January 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.47.1.29.

Osler:2000:BB

[Osl00] Thomas J. Osler. Binomials to binomials. College Mathematics Journal, 31(3):211–212, May 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2000.
11974146.

Osler:2003:VTP

[Osl03] Thomas J. Osler. Variations on a theme from Pascal’s Triangle. College Mathematics Journal, 34(3):216–223, May 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2003.11922009.

Osler:2006:AQP

[Osl06] Thomas J. Osler. Archimedes’ quadrature of the parabola: a mechanical view. College Mathematics Journal, 37(1):24–28, January 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2006.11922163.

Osserman:1981:SVS

[Oss81] Robert Osserman. Structure vs. substance: The fall and rise of geometry. Two-Year College Mathematics Journal, 12(4):239– 246, September 1981. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1981.11972538.

Oman:2024:PSa

[OST24a] Greg Oman, Matyas Sustik, and Katherine Thompson. Problems and solutions. College Mathematics Journal, 55(1):64– 65, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2024.2293458.

Oman:2024:PSb

[OST24b] Greg Oman, Mátyás A. Sustik, and Katherine Thompson. Problems and solutions. College Mathematics Journal, 55(2):175– 179, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2024.2316549.

Oman:2024:PSc

[OST24c] Greg Oman, Mátyás A. Sustik, and Katherine Thompson. Problems and solutions. College Mathematics Journal, 55(3):260– 268, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2024.2327980.

Oman:2024:PS

[OST24d] Greg Oman, Mátyás A. Sustik, and Katherine Thompson. Problems and solutions. College Mathematics Journal, 55(5):435–445, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic).

Oman:2025:PS

[OST25] Greg Oman, Mátyás A. Sustik, and Katherine Thompson. Problems and solutions. College Mathematics Journal, 56(1):68–75, 2025. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic).

Olmsted:1972:STP

[OT72] John M. H. Olmsted and Carl G. Townsend. On the sum of two periodic functions. Two-Year College Mathematics Journal, 3(1):33–38, Spring 1972. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1972.11974194.

Otero:1998:MH

[Ote98] Daniel E. Otero. Media highlights. College Mathematics Journal, 29(3):249–256, May 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1998.11973951.

Ovchinnikov:2011:SIF

[Ovc11] Sergei Ovchinnikov. Series with inverse function terms. College Mathematics Journal, 42(4):283–288, September 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
42.4.283.

Osler:2002:TTS

[OW02] Thomas J. Osler and Marcus Wright. A tale of two series. College Mathematics Journal, 33(2):99–106, March 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2002.11921926.

Pace:1978:RAF

[Pac78] John Pace. A rational approach to fractions. Two-Year College Mathematics Journal, 9(3):154–158, June 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/00494925.
1978.11974567.

Page:1972:HSH

[Pag72] Robert L. Page. How steep is a Hill? Two-Year College Mathematics Journal, 3(1):66–67, Spring 1972. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1972.
11974200.

Page:1979:CCa

[Pag79a] Warren Page. Classroom capsules. Two-Year College Mathematics Journal, 10(1):33–39, January 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1979.
11972320.

Page:1979:QRE

[Pag79b] Warren Page. Questions in the round — an effective barometer of understanding. Two-Year College Mathematics Journal, 10 (4):278–279, September 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1979.11972367.

Page:1983:IR

[Pag83] Warren Page. Interfractile ranges. Two-Year College Mathematics Journal, 14(2):170–172, March 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1983.
11972677.

Page:1984:F

[Pag84a] Warren Page. Forum. College Mathematics Journal, 15(5): 370, November 1984. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1984.11972802.

Page:1984:IHR

[Pag84b] Warren Page. An interview with Herbert Robbins. College Mathematics Journal, 15(1):2–24, January 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1984.
11972739.

Page:1985:IUT

[Pag85] Warren Page. An interview with the 1985 USA team to the international mathematical Olympiad. College Mathematics Journal, 16(5):336–360, November 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1985.11972907.

Page:1986:MCM

[Pag86] Warren Page. Mathematical competition in modeling. College Mathematics Journal, 17(1):32–33, January 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1986.
11972927.

Page:1987:MH

[Pag87a] Edward Page. Media highlights. College Mathematics Journal, 18(2):167–175, March 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1987.11973030.

Page:1987:RAL

[Pag87b] Warren Page. A rational approach to lattice polygons. College Mathematics Journal, 18(4):316–318, September 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1987.11973050.

Page:1988:FE

[Pag88] Warren Page. A farewell editorial. College Mathematics Journal, 19(5):386, November 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1988.11973145.

Page:1989:SR

[Pag89] Warren Page. Software reviews. College Mathematics Journal, 20(4):339–342, September 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1989.11973256.

Page:2009:I

[Pag09] Warren Page. Introduction. College Mathematics Journal, 40(1): 47–54, January 2009. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2009.11922336.

Page:2020:TYC

[Pag20] Warren Page. The (two-year) College mathematics journal during the 1980s. College Mathematics Journal, 51(2):144– 145, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2019.1700678.

Palmiter:1992:SR

[Pal92] Jeanette R. Palmiter. Software reviews. College Mathematics Journal, 23(2):158–161, March 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1992.
11973451.

Palais:1998:EDF

[Pal98] Bob Palais. An example demonstrating the fundamental theorem of calculus. College Mathematics Journal, 29(4):311–312, September 1998. CODEN ???? ISSN 0746-8342 (print), 1931-

1346 (electronic). URL http://www.tandfonline.com/doi/
abs/10.1080/07468342.1998.11973963.

Palatnik:2017:PWV

[Pal17] Alik Palatnik. Proof without words: Varignon’s Theorem. College Mathematics Journal, 48(5):354, November 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
48.5.354.

Paparella:2017:SEP

[Pap17] Pietro Paparella. A short and elementary proof of the twosidedness of the matrix inverse. College Mathematics Journal, 48(5):366–367, November 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.48.5.366.

Partridge:1983:ECS

[Par83] Alden R. Partridge. Ellipses from a circular and spherical point of view. Two-Year College Mathematics Journal, 14(5):436– 438, November 1983. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1983.11972736.

Parsons:1984:RTP

[Par84] William Parsons. Right triangles with perimeter and area equal. College Mathematics Journal, 15(5):429, November 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972828.

Parris:1991:RFR

[Par91] Richard Parris. The root-finding route to chaos. College Mathematics Journal, 22(1):48–55, January 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1991.
11973359.

Pare:1995:VPE

[Par95] Robert Paré. A visual proof of Eddy and Fritsch’s minimal area property. College Mathematics Journal, 26(1):43–44, January 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1995.11973667.

Park:2005:RCF

[Par05] Poo-Sung Park. Ramanujan’s continued fraction for a puzzle. College Mathematics Journal, 36(5):363–365, November 2005. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2005.11922150.

Parris:2007:CT

[Par07] Richard Parris. Commensurable triangles. College Mathematics Journal, 38(5):345–355, November 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2007.
11922259.

Parris:2011:LC

[Par11] Richard Parris. Lattice cubes. College Mathematics Journal, 42(2):118–125, March 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.42.2.118.

Parker:2013:WSB

[Par13] Adam E. Parker. Who solved the Bernoulli differential equation and how did they do it? College Mathematics Journal, 44(2): 89–97, March 2013. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.44.2.089.

Parker:2016:WNW

[Par16] Adam E. Parker. What’s in a name: Why Cauchy and Euler share the Cauchy–Euler equation. College Mathematics Journal, 47(3):191–198, May 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.47.3.191.

Paul:1985:HNP

[Pau85] Joey Paul. The house number problem and its variations. College Mathematics Journal, 16(2):108–117, March 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1985.11972863.

Paulsen:2019:GQI

[Pau19] William Paulsen. Greedy queens on an infinite chessboard. College Mathematics Journal, 50(4):288–294, 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2019.
1632671.

Paulsen:2021:SPR

[Pau21] William Paulsen. The sock problem revisited. College Mathematics Journal, 52(3):193–203, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2021.1897371.

Page:1981:SDS

[PC81] Warren Page and Leo Chosid. Synthetic division shortened. Two-Year College Mathematics Journal, 12(5):334–336, November 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972566.

Pezzimenti:2021:EEI

[PDR21] Samantha Pezzimenti, Giovanni DiCicco, Aditya Kommoju and Dhanush Rajesh. Exploring and extending the impossible card location trick. College Mathematics Journal, 52(5):356– 363, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2021.1967696.

Page:1982:SZP

[PE82] Warren Page and Michael W. Ecker. The sums of zeros of polynomial derivatives. Two-Year College Mathematics Journal, 13 (5):328–329, November 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1982.11972640.

Peckham:1970:IPC

[Pec70] Herbert D. Peckham. An integrated physics-calculus course. Two-Year College Mathematics Journal, 1(1):17–24, Spring 1970. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1970.11973161.

Pedersen:1972:SG

[Ped72] Jean J. Pedersen. Sneaking up on a group. Two-Year College Mathematics Journal, 3(2):9–12, Autumn 1972. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/00494925.
1972.11974208.

Pedersen:1973:PPP

[Ped73] Jean J. Pedersen. Plaited Platonic puzzles. Two-Year College Mathematics Journal, 4(3):22–37, Autumn 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1973.
11974270.

Pedersen:1980:CCP

[Ped80a] Jean J. Pedersen. The challenge of classifying polyhedra. Two-Year College Mathematics Journal, 11(3):162–173, June 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1980.11972431.

Pedersen:1980:GAW

[Ped80b] Jean J. Pedersen. Geometry is alive and well: The Coxeter symposium in Toronto. Two-Year College Mathematics Journal, 11(1):19–25, January 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1980.11972396.

Pedersen:1998:MH

[Ped98a] Jean Pedersen. Media highlights. College Mathematics Journal, 29(4):337–344, September 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1998.11973969.

Pedoe:1998:LG

[Ped98b] Daniel Pedoe. In love with geometry. College Mathematics Journal, 29(3):170–188, May 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1998.11973938.

Penner:1971:ICC

[Pen71] Sidney Penner. An interesting correspondence and its consequences. Two-Year College Mathematics Journal, 2(1):40–44, Spring 1971. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1971.11973996.

Penner:1973:A

[Pen73] Sidney Penner. ab=c. Two-Year College Mathematics Journal, 4(2):86–87, Spring 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1973.11974260.

Penn:1986:CGV

[Pen86] Howard Lewis Penn. Computer graphics for the vibrating string. College Mathematics Journal, 17(1):79–89, January 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1986.11972937.

Penn:1988:TR

[Pen88] Howard Lewis Penn. Ties at rotation. College Mathematics Journal, 19(3):230–239, May 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1988.11973119.

Penn:1989:SR

[Pen89] Howard Lewis Penn. Software reviews. College Mathematics Journal, 20(5):439–441, November 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1989.
11973273.

Pennings:2003:DDK

[Pen03]  Timothy J. Pennings. Do dogs know calculus?  College

Mathematics Journal, 34(3):178–182, May 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2003.
11922003.

Penner:2019:FCB

[Pen19] Alvin Penner. Fitting a cubic Bézier to a parametric function. College Mathematics Journal, 50(3):185–196, 2019. CO-DEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/full/10.1080/07468342.
2019.1583038.

Pengelley:2020:LAL

[Pen20] David Pengelley. From lecture to active learning: Rewards for all, and is it really so difficult? College Mathematics Journal, 51(1):13–24, 2020. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2020.1680228.

Perry:1971:ETE

[Per71] Donald Perry. An experiment in teaching elementary algebra. Two-Year College Mathematics Journal, 2(2):40–46, Autumn 1971. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1971.11974010.

Perry:1973:SUC

[Per73] Donald Perry. A study, using CUPM recommendations as criteria, of the academic preparation of two-year College teachers. Two-Year College Mathematics Journal, 4(2):67–71, Spring 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1973.11974254.

Perry:1983:BPF

[Per83] W. L. Perry. A bifurcation problem in first semester calculus. Two-Year College Mathematics Journal, 14(1):57–60, January 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1983.11972655.

Persky:1994:EBI

[Per94] Ronald L. Persky. Extending Bernoulli’s Inequality. College Mathematics Journal, 25(3):230, May 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1994.
11973615.

Perkins:2003:SP

[Per03] David Perkins. A serendipitous proof. College Mathematics Journal, 34(5):359–361, November 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2003.
11922030.

Perucca:2017:CRC

[Per17] Antonella Perucca. The Chinese remainder clock. College Mathematics Journal, 48(2):82–89, March 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
48.2.82.

Peterson:1970:COM

[Pet70] John Peterson. A computer-oriented multiplication algorithm. Two-Year College Mathematics Journal, 1(2):106, Autumn 1970. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1970.11973194.

Peterson:1983:AAT

[Pet83] Glen Peterson. Approximation to an angle trisection. Two-Year College Mathematics Journal, 14(2):166–167, March 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1983.11972674.

Petrillo:2011:CSD

[Pet11] Joseph Petrillo. Counting subgroups in a direct product of finite cyclic groups. College Mathematics Journal, 42(3):215–222, May 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.42.3.215.

Petrache:2014:CBG

[Pet14] Horia I. Petrache. Correspondence between geometric and differential definitions of the sine and cosine. College Mathematics Journal, 45(1):11–15, January 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
45.1.011.

Peters:2015:SVH

[Pet15] Michael H. Peters. On the shrinking volume of the hypersphere. College Mathematics Journal, 46(3):178–180, May 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.46.3.178.

Pfaff:2011:ASP

[Pfa11] Thomas J. Pfaff. Averaging sums of powers of integers. College Mathematics Journal, 42(5):402–404, November 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
42.5.402.

Pfiefer:1988:BPE

[Pfi88] Richard E. Pfiefer. Bounds on the perimeter of an ellipse via Minkowski sums. College Mathematics Journal, 19(4):348–350, September 1988. CODEN ???? ISSN 0746-8342 (print), 1931-

1346 (electronic). URL http://www.tandfonline.com/doi/
abs/10.1080/07468342.1988.11973137.

Pfiefer:1991:SE

[Pfi91] Richard E. Pfiefer. Stacking ellipses. College Mathematics Journal, 22(4):312–313, September 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1991.
11973402.

Page:1979:DPL

[PG79] Warren Page and Warren B. Gordon. Distance from a point to a line. Two-Year College Mathematics Journal, 10(5):348– 349, November 1979. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1979.11972383.

Perruchet:2006:DDK

[PG06] Pierre Perruchet and Jorge Gallego. Do dogs know related rates rather than optimization? College Mathematics Journal, 37(1): 16–19, January 2006. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2006.11922161.

Page:1990:SR

[PGNH90] Warren Page, Raymond N. Greenwell, Tyre A. Newton, and Leonard G. Henscheid. Software reviews. College Mathematics Journal, 21(4):330–333, September 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1990.
11973330.

Partee:1977:DMA

[PH77] Square Partee and Eric Halsey. Discovery method algebra at the University of Washington. Two-Year College Mathematics Journal, 8(1):27–29, January 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1977.11974476.

Page:1983:MBG

[PH83] Warren Page and David Y. Hsu. Minimization based on the greatest common divisor. Two-Year College Mathematics Journal, 14(2):164–166, March 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1983.11972673.

Phrobis:1990:ASS

[Phr90] Winston Phrobis. The age of the solar system. College Mathematics Journal, 21(5):399–400, November 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1990.
11973341.

Pierce:1971:SCG

[Pie71] R. S. Pierce. A strategy for a class of games. Two-Year College Mathematics Journal, 2(2):55–62, Autumn 1971. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1971.11974013.

Pierce:1977:BHL

[Pie77] R. C. Pierce, Jr. A brief history of logarithms. Two-Year College Mathematics Journal, 8(1):22–26, January 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1977.11974475.

Pinker:1972:NDP

[Pin72] Aron Pinker. A note on derivatives of polynomials. Two-Year College Mathematics Journal, 3(2):77–78, Autumn 1972. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1972.11974222.

Pinker:1974:EWO

[Pin74a] Aron Pinker. The equivalence of the well-ordering principle and Dirichlet’s Box Principle. Two-Year College Mathematics Journal, 5(1):76–77, Winter 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1974.11974299.

Pinker:1974:GTI

[Pin74b] Aron Pinker. A generator of trigonometric identities. Two-Year College Mathematics Journal, 5(4):54–55, December 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1974.11974354.

Pinker:1975:IUG

[Pin75] Aron Pinker. An interesting use of generating functions. Two-Year College Mathematics Journal, 6(4):39–45, December 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1975.11974406.

Pinkham:2001:SRC

[Pin01] R. S. Pinkham. Simpson’s Rule with constant weights. College Mathematics Journal, 32(2):91–93, March 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/07468342.
2001.11921859.

Piotrowski:2021:ASR

[Pio21] Andrzej Piotrowski. Atypical series representations of Riemannintegrable functions. College Mathematics Journal, 52(1):31– 38, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2021.1847589.

Page:1981:CEC

[PJS81] Warren Page, Richard Johnsonbaugh, and E. F. Schuster. Corrections to an earlier capsule. Two-Year College Mathematics Journal, 12(3):204–206, June 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1981.
11972525.

Page:1983:SFS

[PK83] Warren Page and Dan Kalman. Sum formulas for sine and cosine. Two-Year College Mathematics Journal, 14(1):55–56, January 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1983.11972653.

Page:1979:CCc

[PKL^{+}79] Warren Page, J. P. King, W. G. Leavitt, David Beran, R. S. Luthar, Stephen A. Book, and Lawrence Sher. Classroom capsules. Two-Year College Mathematics Journal, 10(3):197–204, June 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1979.11972352.

Page:1981:SIF

[PL81] Warren Page and R. S. Luthar. A standard integral formula. Two-Year College Mathematics Journal, 12(5):329–330, November 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972561.

Page:1982:IS

[PL82] Warren Page and William A. Leonard. Intuition out to sea. Two-Year College Mathematics Journal, 13(3):195–196, June 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1982.11972607.

Pulskamp:1994:USF

[PL94] Richard J. Pulskamp and William J. Larkin III. Using the sign function to analyze graphs. College Mathematics Journal, 25 (4):327–328, September 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1994.11973634.

Poodiack:2009:FTA

[PL09] Robert D. Poodiack and Kevin J. LeClair. Fundamental theorems of algebra for the perplexes. College Mathematics Journal, 40 (5):322–336, November 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/074683409X475643.

Plagge:1978:FQA

[Pla78] Richard Plagge. Fractions without quotients: Arithmetic of repeating decimals. Two-Year College Mathematics Journal, 9(1): 11–15, January 1978. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1978.11974537.

Plaza:2014:PWL

[Pla14] Ángel Plaza. Proof without words: Limit of a recursive arithmetic mean. College Mathematics Journal, 45(5):364, November 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.45.5.364.

Plaza:2016:PWA

[Pla16] Ángel Plaza. Proof without words: Arithmetic mean of two means. College Mathematics Journal, 47(2):125, March 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.47.2.125.

Plaza:2017:PWP

[Pla17a] Ángel Plaza. Proof without words: a Pascal-like triangle with Pell number row sums. College Mathematics Journal, 48(5): 346, November 2017. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.48.5.346.

Plaza:2017:PWS

[Pla17b] Ángel Plaza. Proof without words: Sum of a row in Pascal’s Triangle. College Mathematics Journal, 48(3):188, May 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.48.3.188.

Plaza:2017:PWT

[Pla17c] Ángel Plaza. Proof without words: The triangle with maximum area for a given base and perimeter. College Mathematics Journal, 48(1):51, January 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.48.1.51.

Plaza:2018:PWA

[Pla18] Ángel Plaza. Proof without words: an alternating geometric series. College Mathematics Journal, 49(3):200, May 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2018.1437302.

Page:2018:MH

[PLS18] Warren Page, Tanya Leise, and Philip Straffin. Media highlights. College Mathematics Journal, 49(5):384–391, November 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2018.1505134.

Perry:1974:SRP

[PM74] Donald Perry and Wayne L. Miller. A suggested recruiting project: Math contests. Two-Year College Mathematics Journal, 5(4):19–21, December 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1974.11974346.

Page:1982:NRA

[PM82a] Warren Page and V. N. Murty. Nearness relations among measures of central tendency and dispersion: Part 1. Two-Year College Mathematics Journal, 13(5):315–327, November 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1982.11972639.

Perisastry:1982:BRA

[PM82b] M. Perisastry and V. N. Murty. Bounds for the ratio of the arithmetic mean to the geometric mean. Two-Year College Mathematics Journal, 13(2):160–161, March 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1982.
11972600.

Page:1983:NRA

[PM83] Warren Page and V. N. Murty. Nearness relations among measures of central tendency and dispersion: Part 2. Two-Year College Mathematics Journal, 14(1):8–17, January 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1983.11972647.

Page:1988:SMT

[PM88] Warren Page and David S. Moore. Should mathematicians teach statistics? College Mathematics Journal, 19(1):2–7, January 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1988.11973073.

Polya:1976:TSSa

[Pól76a] G. Pólya. As their students see them. Two-Year College Mathematics Journal, 7(2):54, May 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1976.
11974431.

Polya:1976:TSSb

[Pól76b] G. Pólya. As their students see them. Two-Year College Mathematics Journal, 7(3):51, September 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1976.
11974448.

Polya:1976:TSSc

[Pól76c] G. Pólya. As their students see them. Two-Year College Mathematics Journal, 7(4):38, December 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1976.
11974462.

Polya:1978:GP

[Pól78] George Pólya. Guessing and proving. Two-Year College Mathematics Journal, 9(1):21–27, January 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1978.
11974541.

Polya:1979:MGP

[Pól79] George Pólya. More on guessing and proving. Two-Year College Mathematics Journal, 10(4):255–258, September 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1979.11972360.

Pomeranz:1977:CAP

[Pom77] Janet B. Pomeranz. Counterexamples from the algebra of polynomials over a nonfield. Two-Year College Mathematics Journal, 8(1):11–14, January 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1977.11974472.

Pomeranz:1984:DPT

[Pom84] Janet Bellcourt Pomeranz. The dice problem-then and now. College Mathematics Journal, 15(3):229–237, June 1984. CO-DEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/00494925.
1984.11972775.

Pong:2007:SCI

[Pon07] Wai Yan Pong. Sums of consecutive integers. College Mathematics Journal, 38(2):119–123, March 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2007.
11922226.

Pongsriiam:2017:IVG

[Pon17] Prapanpong Pongsriiam. Integral values of the generating functions of Fibonacci and Lucas numbers. College Mathematics Journal, 48(2):97–101, March 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
48.2.97.

Pook:2012:HEF

[Poo12] Les Pook. From hexaflexagons to edge flexagons to point flexagons. College Mathematics Journal, 43(1):11–14, January 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.43.1.011.

Porter:1993:LAA

[Por93] Gerald J. Porter. Linear algebra and affine planar transformations. College Mathematics Journal, 24(1):47–51, January 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1993.11973505.

Powderly:1980:GPC

[Pow80] Mary Powderly. A geometric proof of Cauchy’s Generalized Law of the Mean. Two-Year College Mathematics Journal, 11 (5):329–330, November 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1980.11972473.

Pedersen:1984:PSA

[PP84] Jean Pedersen and George Pólya. On problems with solutions attainable in more than one way. College Mathematics Journal, 15(3):218–228, June 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1984.11972774.

Previte:2021:BCD

[PP21] Joseph Previte and Michelle Previte. The beautiful chaotic dynamics of i^{z}. College Mathematics Journal, 52(5):364–372, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2021.1973820.

Pedersen:1985:TUU

[PR85] Jean Pedersen and Peter Ross. Testing understanding and understanding testing. College Mathematics Journal, 16(3):178–185, June 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1985.11972877.

Polster:2018:MS

[PR18] Burkard Polster and Marty Ross. Marching in squares. College Mathematics Journal, 49(3):181–186, May 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/full/10.1080/07468342.
2018.1440054.

Pratt:2009:FH

[Pra09] Vaughan Pratt. Factoring Heron. College Mathematics Journal, 40(1):15–16, January 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2009.11922330.

Praton:2017:MTU

[Pra17] Iwan Praton. Minimal tilings of a unit square. College Mathematics Journal, 48(4):242–247, September 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
48.4.242.

Prc:1982:SDA

[Prc82] David Y. Hsu Prc. A simple divisibility algorithm. Two-Year College Mathematics Journal, 13(1):58–59, January 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1982.11972580.

Prindle:1972:PL

[Pri72] Paul E. Prindle. Publisher’s letter. Two-Year College Mathematics Journal, 3(2):7, Autumn 1972. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1972.11974206.

Price:1973:TSS

[Pri73a] G. Baley Price. Telescoping sums and the summation of sequences. Two-Year College Mathematics Journal, 4(2):16–29, Spring 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1973.11974247.

Price:1973:DTY

[Pri73b] H. Vernon Price. A doctorate for the two-year College instructor? Two-Year College Mathematics Journal, 4(1):48–50, Winter 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1973.11974235.

Prichett:1976:GS

[Pri76] Gordon D. Prichett. The game of sprouts. Two-Year College Mathematics Journal, 7(4):21–25, December 1976. CODEN ????

ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1976.
11974458.

Price:1989:LMT

[Pri89] J. J. Price. Learning mathematics through writing: Some guidelines. College Mathematics Journal, 20(5):393–401, November 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1989.11973262.

Price:1990:AEP

[Pri90] J. J. Price. Algorithms for evaluation of polynomials. College Mathematics Journal, 21(5):404–405, November 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1990.11973344.

Pritikin:1995:RPF

[Pri95] Dan Pritikin. Rectangular-to-polar folding fans. College Mathematics Journal, 26(4):305–308, September 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1995.
11973718.

Priestley:1999:SRR

[Pri99] W. M. Priestley. From square roots to n-th roots: Newton’s Method in disguise. College Mathematics Journal, 30(5):387– 388, November 1999. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1999.11974091.

Price:2018:PWS

[Pri18] Tim Price. Proof without words: Sum of squares of consecutive Fibonacci numbers. College Mathematics Journal, 49(2): 121, March 2018. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2018.1424425.

Protomastro:1976:APE

[Pro76] Gerard P. Protomastro. An alternative proof of the equality of the mixed partial derivatives. Two-Year College Mathematics Journal, 7(4):47–48, December 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1976.11974466.

Proctor:1995:DOF

[Pro95] T. G. Proctor. Distinguished oscillations of a forced harmonic oscillator. College Mathematics Journal, 26(2):111–117, March 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1995.11973679.

Page:1979:AAD

[PS79] Warren Page and Norman Schaumberger. An alternate approach to the derivatives of the trigonometric functions. Two-Year College Mathematics Journal, 10(4):276–277, September 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1979.11972365.

Page:1981:IP

[PS81] Warren Page and Man-Keung Siu. Inventor’s paradox. Two-Year College Mathematics Journal, 12(4):267, September 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972542.

Page:1982:ERP

[PS82a] Warren Page and K. R. S. Sastry. An equal ratio property for convex polygons. Two-Year College Mathematics Journal, 13 (4):269–270, September 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1982.11972626.

Page:1982:SFC

[PS82b] Warren Page and Allen J. Schwenk. Selection of a fair currency exchange rate. Two-Year College Mathematics Journal, 13(2):154–155, March 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1982.11972596.

Putz:2003:GPT

[PS03] John F. Putz and Timothy A. Sipka. On generalizing the Pythagorean Theorem. College Mathematics Journal, 34(4):291– 295, September 2003. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2003.11922020.

Ponomarenko:2012:TSE

[PS12] Vadim Ponomarenko and Natalie Selinski. Two semigroup elements can commute with any positive rational probability. College Mathematics Journal, 43(4):334–336, September 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.43.4.334.

Pohoata:2015:SCS

[PS15] Cosmin Pohoata and Richard Stong. A sufficient condition for subgroups with prime indices to be normal. College Mathematics Journal, 46(5):348–351, November 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
46.5.348.

Palatnik:2022:PWP

[PS22] Alik Palatnik and Moshe Stupel. Proof without words: A property of a cyclic polygon with an even number of vertices. College Mathematics Journal, 53(2):146, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2022.
2020588.

Partin:1976:LE

[PSSP76] Harold Partin, John Staib, Paul Shapiro, and Herbert Peebles. Letters to the Editor. Two-Year College Mathematics Journal, 7(3):1–4, September 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1976.11974435.

Patil:1986:DFD

[PU86] S. A. Patil and V. R. R. Uppuluri. The distribution of first j digits. College Mathematics Journal, 17(3):240–243, May 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.jstor.org/stable/2686982;
 http://www.tandfonline.com/doi/abs/10.1080/07468342.
1986.11972963.

Pugh:1980:ACD

[Pug80] David L. Pugh. An application from combinatorics to dice-sum frequencies. Two-Year College Mathematics Journal, 11(5):331– 333, November 1980. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1980.11972475.

Pulskamp:1988:CMT

[Pul88] Richard J. Pulskamp. Constructing a map from a table of intercity distances. College Mathematics Journal, 19(2):154–163, March 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1988.11973104.

Putz:1986:PPE

[Put86] John F. Putz. The Pascal polytope: an extension of Pascal’s Triangle to N dimensions. College Mathematics Journal, 17(2): 144–155, March 1986. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1986.11972945.

Poet:2005:CCM

[PV05] Jeffrey L. Poet and Donald L. Vestal, Jr. Curious consequences of a miscopied quadratic. College Mathematics Journal, 36(4):273– 277, September 2005. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2005.11922139.

Pedersen:2009:SMO

[PV09] Ryan M. Pedersen and Timothy L. Vis. Sets of mutually orthogonal sudoku latin squares. College Mathematics Journal, 40(3): 174–181, May 2009. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2009.11922356.

Patel:2013:PWV

[PV13] Purna Patel and Raymond Viglione. Proof without words: a variation on Thébault’s first problem. College Mathematics Journal, 44(2):135, March 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.44.2.135.

Pasternack:2016:HRO

[PV16] Joel Pasternack and Stewart Venit. Horse racing odds: Can you beat the track by hedging your bets? College Mathematics Journal, 47(4):275–281, September 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
47.4.275.

Passias:2021:CCP

[PW21] Georgios Passias and Sven-Ake Wegner. On a counterexample in connection with the Picard–Lindelöf Theorem. College Mathematics Journal, 52(3):221–223, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2021.
1909978.

Xu:1991:SPR

[pX91] Xiao peng Xu. The snowplow problem revisited. College Mathematics Journal, 22(2):139, March 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1991.
11973371.

Packel:2004:PMR

[PY04] Edward W. Packel and David S. Yuen. Projectile motion with resistance and the Lambert W function. College Mathematics Journal, 35(5):337–350, November 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2004.
11922095.

Page:1979:CCb

[PZA^{+}79] Warren Page, Elmar Zemgalis, David H. Anderson, Norman Schaumberger, Alban J. Roques, V. N. Murty, Warren B. Gordon, and Harold D. Shane. Classroom capsules. Two-Year College Mathematics Journal, 10(2):112–118, March 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1979.11972335.

Quinn:2016:TRG

[Qui16] Anne Quinn. Technology review: Graph theory Smartphone apps. College Mathematics Journal, 47(1):67–72, January 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.47.1.67.

Quinn:2017:TRS

[Qui17] Anne Quinn. Technology review: Statistics Web apps. College Mathematics Journal, 48(5):378–382, November 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
48.5.378.

Quintanilla:2022:NDS

[Qui22] John A. Quintanilla. A new derivation of snell’s law without calculus. College Mathematics Journal, 53(2):140–145, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2022.2011544.

Ralston:1984:WDM

[Ral84] Anthony Ralston. Will discrete mathematics surpass calculus in importance? College Mathematics Journal, 15(5):371–373, November 1984. CODEN ???? ISSN 0746-8342 (print), 1931-

1346 (electronic). URL http://www.tandfonline.com/doi/
abs/10.1080/00494925.1984.11972803.

Ratliff:1981:GMF

[Rat81] Michael Ratliff. A geometric motivation of Fermat’s factoring method. Two-Year College Mathematics Journal, 12(1):24–27, January 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972485.

Ratliff:2010:LCV

[Rat10] Thomas C. Ratliff. Lewis Carroll, voting, and the taxicab metric. College Mathematics Journal, 41(4):303–311, September 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/074683410X510308.

Razminia:2015:SPS

[Raz15] Kambiz Razminia. A short proof of symmetric inequalities. College Mathematics Journal, 46(5):364–366, November 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.46.5.364.

Rice:2021:WHC

[RB21] Adrian Rice and Ezra Brown. Why Hamilton couldn’t multiply triples. College Mathematics Journal, 52(3):185–192, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2021.1897418.

Rossman:1999:TRS

[RC99] Allan J. Rossman and Beth L. Chance. Teaching the reasoning of statistical inference: a “top ten” list. College Mathematics Journal, 30(4):297–305, September 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1999.
11974074.

Rubis:2020:BRG

[RC20] Oksana Rubis and Kathleen M. Clark. Book review: A Guide to Higher Learning by Julie Chen, Flying Fish Press, Oakland, 2009. College Mathematics Journal, 51(5):378–384, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2020.1830658.

Riegsecker:1986:SR

[RCS86] John Riegsecker, Judith L. Clarke, and Carol M. Smith. Software reviews. College Mathematics Journal, 17(2):182–183, March 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1986.11972954.

Rebman:1979:PPW

[Reb79] Kenneth R. Rebman. The pigeonhole principle (what it is, how it works, and how it applies to map coloring). Two-Year College Mathematics Journal, 10(1):3–13, January 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1979.11972315.

Reber:1989:PNG

[Reb89] James C. Reber. Pseudorandom number generators in a fourbit computer system. College Mathematics Journal, 20(1):54–55, January 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.jstor.org/stable/pdfplus/
2686821.pdf;   http://www.tandfonline.com/doi/abs/10.
1080/07468342.1989.11973205.

Redmond:1989:FRR

[Red89] Don Redmond. Finding rational roots of polynomials. College Mathematics Journal, 20(2):139–141, March 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1989.
11973222.

Reed:1988:WTW

[Ree88] Michael Reed. Who teaches what to whom? College Mathematics Journal, 19(1):24–25, January 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1988.
11973083.

Reinhardt:1973:FTA

[Rei73a] H. E. Reinhardt. Four theorems about Montana. Two-Year College Mathematics Journal, 4(1):76–78, Winter 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1973.11974243.

Reinhardt:1973:DI

[Rei73b] Howard E. Reinhardt. Dividing inheritances. Two-Year College Mathematics Journal, 4(2):30–33, Spring 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/00494925.
1973.11974248.

Reid:1986:AJR

[Rei86] Constance Reid. The autobiography of Julia Robinson. College Mathematics Journal, 17(1):2–21, January 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/07468342.
1986.11972925.

Reimann:1987:AA

[Rei87] Kurt W. Reimann. Annuities as areas. College Mathematics Journal, 18(1):45–47, January 1987. CODEN ????

ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1987.
11973006.

Rekdal:2016:PLK

[Rek16] Paisley Rekdal. Philip Larkin’s Koan. College Mathematics Journal, 47(2):133, March 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.47.2.133.

Renz:1981:MPW

[Ren81] Peter Renz. Mathematical proof: What it is and what it ought to be. Two-Year College Mathematics Journal, 12(2):83–103, March 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972504.

Renz:1982:HGP

[Ren82]  Peter Renz. How good a picture do you need?  Two-

Year College Mathematics Journal, 13(3):204–205, June 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1982.11972613.

Renz:1985:PH

[Ren85] Peter Renz. The path to Hell. College Mathematics Journal, 16 (1):9–11, January 1985. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1985.11972842.

Renz:1989:MH

[Ren89] Peter W. Renz. Media highlights. College Mathematics Journal, 20(3):266–273, May 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1989.11973243.

Ren:2014:PWPb

[Ren14a] Guanshen Ren. Proof without words: Pythagoras for a clipped rectangle. College Mathematics Journal, 45(3):216, May 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.45.3.216.

Ren:2014:PWPa

[Ren14b] Guanshen Ren. Proof without words: Pythagoras for a right trapezoid. College Mathematics Journal, 45(3):198, May 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.45.3.198.

Reynolds:1993:AVA

[Rey93] Barbara E. Reynolds. The algorists vs. the abacists: an ancient controversy on the use of calculators. College Mathematics Journal, 24(3):218–223, May 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1993.11973531.

Riahi:1995:EIM

[Ria95] Farhad Riahi. An early iterative method for the determination of sin 1^{∘}. College Mathematics Journal, 26(1):16–21, January 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1995.11973660.

Ribenboim:1994:PNR

[Rib94] Paulo Ribenboim. Prime number records. College Mathematics Journal, 25(4):280–290, September 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1994.
11973623.

Ribenboim:1997:TFG

[Rib97] Paulo Ribenboim. Are there functions that generate prime numbers? College Mathematics Journal, 28(5):352–359, November 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1997.11973889.

Rickey:1987:INM

[Ric87] V. Frederick Rickey. Isaac Newton: Man, myth, and mathematics. College Mathematics Journal, 18(5):362–389, November 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1987.11973060.

Richman:1993:CA

[Ric93] Fred Richman. A circular argument. College Mathematics Journal, 24(2):160–162, March 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1993.11973522.

Richeson:2005:CUS

[Ric05] David Richeson. Centers of the United States. College Mathematics Journal, 36(5):366–373, November 2005. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2005.
11922151.

Rich:2008:LN

[Ric08] Andrew Rich. Leftist numbers. College Mathematics Journal, 39(5):330–336, November 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2008.11922313.

Richmond:2010:PPP

[Ric10] Bettina Richmond. On a perplexing polynomial puzzle. College Mathematics Journal, 41(5):400–403, November 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
4169/074683410X522017.

Rice:2012:TTY

[Ric12] Theodore Rice. Teaching tip: Are you changing the rules? again? College Mathematics Journal, 43(4):304, September 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.43.4.304.

Richeson:2015:CRW

[Ric15a] David Richeson. Circular reasoning: Who first proved that C divided by d is a constant? College Mathematics Journal, 46(3): 162–171, May 2015. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.46.3.162.

Richeson:2015:PWM

[Ric15b] David Richeson. Proof without words: The maximum sum of inradii. College Mathematics Journal, 46(1):23, January 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.46.1.23.

Richmond:2018:CC

[Ric18] Tom Richmond. Calculus with curtains. College Mathematics Journal, 49(5):369–370, November 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2018.
1512326.

Richeson:2019:BRC

[Ric19a] David Richeson. Book review: Calculus Reordered: A History of the Big Ideas by David Bressoud, Princeton University Press, Princeton, 2019. Change Is the Only Constant: The Wisdom of Calculus in a Madcap World by Ben Orlin, Black Dog & Leventhal, New York, 2019. Infinite Powers: How Calculus Reveals the Secrets of the Universe by Steven Strogatz, Houghton Mifflin Harcourt, Boston, 2019. College Mathematics Journal, 50(4):307–312, 2019. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2019.1656942.

Richeson:2019:N

[Ric19b] David Richeson. By the numbers. College Mathematics Journal, 50(4):286–287, 2019. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2019.1657340.

Richeson:2019:NS

[Ric19c] David Richeson. By the numbers solution. College Mathematics Journal, 50(4):300, 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2019.1657341.

Ricardo:2022:EDN

[Ric22] Henry Ricardo. The equivalence of definitions of the natural logarithm function. College Mathematics Journal, 53(3):190– 196, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2022.2039553.

Rice:2024:PPR

[Ric24] Alex Rice. A precise probability related to Simpson’s Paradox. College Mathematics Journal, 55(5):400–405, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic).

Ridenhour:1990:SCM

[Rid90] Jim R. Ridenhour. Spiders, computers, and Markov chains. College Mathematics Journal, 21(4):323–326, September 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1990.11973328.

Riddle:1995:ICI

[Rid95] Larry Riddle. Individualized computer investigations for multivariable calculus. College Mathematics Journal, 26(3):235–237, May 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1995.11973706.

Riddle:1998:ALC

[Rid98] Larry Riddle. Arc length contest. College Mathematics Journal, 29(4):314–320, September 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1998.11973965.

Roberts:1971:CMP

[RM71] Keith J. Roberts and Leo E. Michels. The crossover mathematics program at Milwaukee area technical College. Two-Year College Mathematics Journal, 2(2):47–50, Autumn 1971. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1971.11974011.

Robertson:1973:ACC

[Rob73] Jack M. Robertson. Another challenge in the classroom. Two-Year College Mathematics Journal, 4(2):48–54, Spring 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1973.11974251.

Robertson:1978:AFM

[Rob78] Jane I. Robertson. Algorithms for finding maturity value in compound interest problems. Two-Year College Mathematics Journal, 9(4):249–251, September 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1978.
11974587.

Robbins:1984:SBS

[Rob84a] Herbert Robbins. Some breakthroughs in statistical methodology. College Mathematics Journal, 15(1):25–29, January 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972740.

Roberts:1984:IMC

[Rob84b] Fred S. Roberts. The introductory mathematics curriculum: Misleading, outdated, and unfair. College Mathematics Journal, 15 (5):383–385, November 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1984.11972811.

Roberts:1985:ACV

[Rob85] Jerry A. Roberts. Accurate computation of variance. College Mathematics Journal, 16(2):149–150, March 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1985.
11972872.

Robertson:1990:PT

[Rob90] John S. Robertson. Proofs by — Tion. College Mathematics Journal, 21(3):220–222, May 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1990.
11973310.

Robertson:1997:GSP

[Rob97] John S. Robertson. Gudermann and the simple pendulum. College Mathematics Journal, 28(4):271–276, September 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1997.11973874.

Rodin:1980:EBI

[Rod80] Ervin Y. Rodin. Extending Bernoulli’s Inequality. Two-Year College Mathematics Journal, 11(2):124–126, March 1980.

CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1980.11972423.

Roe:1993:BST

[Roe93] P. L. Roe. The best shape for a tin can. College Mathematics Journal, 24(3):233–236, May 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1993.
11973534.

Rogers:2007:IFP

[Rog07] Allen D. Rogers. Integrals of fitted polynomials and an application to Simpson’s Rule. College Mathematics Journal, 38(2): 124–130, March 2007. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2007.11922227.

Rojas:2016:WOC

[Roj16] Ricardo E. Rojas. Waiter! One classic calculus problem, hold the calculus. College Mathematics Journal, 47(1):59–60, January 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.47.1.59.

Rojas:2023:WTS

[Roj23] Dr. Ricardo E. Rojas. When triangles are similar. College Mathematics Journal, 54(1):68–69, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2022.
2150506.

Rokicki:2014:TGN

[Rok14] Tomas Rokicki. Towards God’s number for Rubik’s Cube in the quarter-turn metric. College Mathematics Journal, 45(4):242– 253, September 2014. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.45.4.242.

Roques:1979:CC

[Roq79] Alban J. Roques. Computers & calculators. Two-Year College Mathematics Journal, 10(5):340–343, November 1979.

CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1979.11972380.

Ross:1970:RDC

[Ros70] Don Ross. Remedial or developmental? confusion over terms. Two-Year College Mathematics Journal, 1(2):27–31, Autumn 1970. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1970.11973183.

Rosenbaum:1976:IAE

[Ros76] Burt M. Rosenbaum. An integral approximation exact for fifthdegree polynomials. Two-Year College Mathematics Journal, 7 (3):10–14, September 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1976.11974437.

Rosenfeld:1977:HDO

[Ros77] Robert Rosenfeld. A helpful device: or one more use for Pascal’s Triangle. Two-Year College Mathematics Journal, 8(3):188–191, June 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1977.11974507.

Rosendahl:1987:C

[Ros87] John Rosendahl. Comparing B^{A} and A^{B} for A > B. College Mathematics Journal, 18(1):50, January 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1987.
11973008.

Rosenthal:1989:LTI

[Ros89] William E. Rosenthal. Lattices of trigonometric identities. College Mathematics Journal, 20(3):232–234, May 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/07468342.
1989.11973237.

Rose:1991:D

[Ros91] David A. Rose. The differentiability of Sinx. College Mathematics Journal, 22(2):139–142, March 1991. CODEN ????

ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1991.
11973372.

Ross:1992:MH

[Ros92] Peter Ross. Media highlights. College Mathematics Journal, 23(2):168–175, March 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1992.11973453.

Ross:2000:BR

[Ros00] Peter Ross. Book review. College Mathematics Journal, 31(3): 240–244, May 2000. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2000.11974152.

Rosenholtz:2001:ELI

[Ros01] Ira Rosenholtz. Estimating large integrals: The bigger they are, the harder they fall. College Mathematics Journal, 32(5):322– 329, November 2001. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2001.11921898.

Ross:2002:BR

[Ros02] Peter Ross. Book review. College Mathematics Journal, 33(3): 256–259, May 2002. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2002.11921950.

Rosario:2004:SSC

[Ros04] Héctor Rosario. Simmons’ subliminal channel. College Mathematics Journal, 35(3):208–212, May 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2004.
11922074.

Ross:2006:BR

[Ros06] Peter Ross. Book review. College Mathematics Journal, 37 (4):328–331, September 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2006.11922201.

Ross:2012:DMT

[Ros12] Andrew M. Ross. Designing medical tests: The other side of Bayes’ Theorem. College Mathematics Journal, 43(3):251–254, May 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.43.3.251.

Rosenhouse:2014:KKN

[Ros14] Jason Rosenhouse. Knights, knaves, normals, and neutrals. College Mathematics Journal, 45(4):297–306, September 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.45.4.297.

Ross:2016:BR

[Ros16] Peter Ross. Book review. College Mathematics Journal, 47(2): 146–152, March 2016. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.47.2.146.

Rosenhouse:2017:BR

[Ros17] Jason Rosenhouse. Book review. College Mathematics Journal, 48(4):302–312, September 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.48.4.302.

Rosentrater:2018:FF

[Ros18] C. Ray Rosentrater. Fakin’ flips. College Mathematics Journal, 49(3):201–204, May 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2018.1446655.

Rousseau:2009:BR

[Rou09] Cecil Rousseau. Book reviews. College Mathematics Journal, 40(1):65–67, January 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2009.11922338.

Roussos:2012:SMP

[Rou12] Ioannis M. Roussos. On the Steiner minimizing point and the corresponding algebraic system. College Mathematics Journal, 43(4):305–308, September 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.43.4.305.

Rousseau:2013:HIL

[Rou13] Christiane Rousseau. How inge Lehmann discovered the inner core of the Earth. College Mathematics Journal, 44(5):399– 408, November 2013. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.44.5.399.

Rougetet:2014:PN

[Rou14] Lisa Rougetet. A prehistory of nim. College Mathematics Journal, 45(5):358–363, November 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
45.5.358.

Rowland:2015:CCC

[Row15] Dana Rowland. Candy crush combinatorics. College Mathematics Journal, 46(4):255–262, September 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
46.4.255.

Rozema:1988:WSW

[Roz88] Edward Rozema. Why should we pivot in Gaussian elimination? College Mathematics Journal, 19(1):63–72, January 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1988.11973095.

Rozema:1994:RBA

[Roz94] Edward Rozema. Round-off, batting averages, and illconditioning. College Mathematics Journal, 25(4):314–317, September 1994. CODEN ???? ISSN 0746-8342 (print), 1931-

1346 (electronic). URL http://www.tandfonline.com/doi/
abs/10.1080/07468342.1994.11973628.

Rozema:2007:EMS

[Roz07] Edward Rozema. Epidemic models for SARS and measles. College Mathematics Journal, 38(4):246–259, September 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2007.11922245.

Rousseau:1985:IMO

[RP85] Cecil Rousseau and Gregg Patruno. The international mathematical Olympiad training session. College Mathematics Journal, 16(5):362–365, November 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1985.11972908.

Reiter:1989:CSM

[RR89] Harold Reiter and David Ritchie. A complete solution to the magic hexagram problem. College Mathematics Journal, 20(4): 307–316, September 1989. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1989.11973247.

Richmond:1998:CPF

[RR98] Bettina Richmond and Tom Richmond. Characterizing power functions by volumes of revolution. College Mathematics Journal, 29(1):40–41, January 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1998.11973913.

Rozen:1985:APR

[RS85] R. Rozen and A. Sofo. Area of a parabolic region. College Mathematics Journal, 16(5):400–402, November 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1985.11972917.

Reiter:1996:JCI

[RS96] Harold Reiter and Isaac Sonin. The “join the club” interpretation of some graph algorithms. College Mathematics Journal, 27(1):54–58, January 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1996.11973751.

Ross:2000:WRB

[RS00] Chip Ross and Jody Sorensen. Will the real bifurcation diagram please stand up! College Mathematics Journal, 31(1):2–14, January 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2000.11974102.

Ryan:2016:SBR

[RS16] Kathleen Ryan and Brittany Shelton. Statistics on the bonus round of wheel of fortune. College Mathematics Journal, 47 (4):250–253, September 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.47.4.250.

Roualdes:2021:OPC

[RS21] Edward A. Roualdes and Neil C. Schwertman. Optimal pooling of COVID-19 samples. College Mathematics Journal, 52 (5):380–384, 2021. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2021.1971931.

Richmond:2023:ATA

[RS23a] Tom Richmond and Mia Sword. An arithmetic triangle arising from a one-way street grid. College Mathematics Journal, 54(2):113–122, 2023. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2023.2186083.

Ruff:2023:AA

[RS23b] Hayden Ruff and Kenneth Schilling. An attractive attractor. College Mathematics Journal, 54(4):316–325, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/full/10.1080/07468342.
2023.2237849.

Rice:2007:SLT

[RT07] Adrian Rice and Eve Torrence. “Shutting up like a telescope”: Lewis Carroll’s “curious” condensation method for evaluating determinants. College Mathematics Journal, 38(2):85–95, March 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2007.11922223.

Rucker:1982:MSG

[Ruc82] Rudy Rucker. A machine as smart as God. Two-Year College Mathematics Journal, 13(2):115–121, March 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1982.
11972590.

Ruckle:1987:DGH

[Ruc87] William H. Ruckle. A division game: How far can you stretch mathematical induction? College Mathematics Journal, 18(3): 212–218, May 1987. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1987.11973034.

Richmond:2013:III

[RY13] Tom Richmond and Aaron Young. Instant insanity II. College Mathematics Journal, 44(4):265–272, September 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
44.4.265.

Ryder:1996:EFN

[Ryd96] Jack Ryder. Exploring Fibonacci numbers mod M. College Mathematics Journal, 27(2):122–124, March 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1996.
11973763.

Roseman:1999:SOI

[RZ99] Joseph J. Roseman and Gideon Zwas. Second order iterations. College Mathematics Journal, 30(5):393–396, November 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1999.11974094.

Staib:1979:DAI

[SA79] John Staib and Howard Anton. A discovery approach to integration by parts. Two-Year College Mathematics Journal, 10 (5):353–354, November 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1979.11972387.

Schwartz:1999:SPS

[SA99] Mark Schwartz and Darryl Adams. Shortest path solution by epitrochoid machine. College Mathematics Journal, 30(3):221– 225, May 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1999.11974063.

Saar:1970:FRT

[Saa70] Howard Saar. F(1) rejection theorem. Two-Year College Mathematics Journal, 1(2):39–40, Autumn 1970. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1970.
11973186.

Sachs:2022:SFD

[Sac22] Robert L. Sachs. Similarities in a first differential equations course. College Mathematics Journal, 53(2):87–97, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2022.2011646.

Sadhukhan:2017:PNN

[Sad17] Arpan Sadhukhan. Partitioning the natural numbers to prove the infinitude of primes. College Mathematics Journal, 48(3):217– 218, May 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.48.3.217.

Sagan:1992:SR

[Sag92] Bruce E. Sagan. Software reviews. College Mathematics Journal, 23(4):334–339, September 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1992.11973482.

Saler:1985:IHF

[Sal85] B. M. Saler. Inverse hyperbolic functions as areas. College Mathematics Journal, 16(2):129–131, March 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1985.
11972866.

Salwach:1988:CDC

[Sal88] Chester J. Salwach. Codes that detect and correct errors. College Mathematics Journal, 19(5):402–416, November 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1988.11973147.

Salomon:2014:IP

[Sal14] Paul Salomon. Imbalance puzzles. College Mathematics Journal, 45(4):288, September 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.45.4.288.

Salwinski:2018:ESP

[Sal18] David Salwinski. Euler’s sine product formula: an elementary proof. College Mathematics Journal, 49(2):126–135, March 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2018.1419703.

Sanger:1977:TLC

[San77] Robert C. Sanger. A transfer level computer calculus sequence. Two-Year College Mathematics Journal, 8(4):216–218, September 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1977.11974513.

Sandefur:1991:DDM

[San91] James T. Sandefur. Discrete dynamical modeling. College Mathematics Journal, 22(1):13–22, January 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1991.
11973354.

Sandifer:2002:SCM

[San02] Edward Sandifer. Spanish colonial mathematics: a window on the past. College Mathematics Journal, 33(4):266–278, September 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2002.11921952.

Sanchis:2006:CTM

[San06] Gabriela R. Sanchis. A card trick and the mathematics behind it. College Mathematics Journal, 37(2):103–109, March 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2006.11922172.

Sand:2019:ASC

[San19] Mark Sand. Adding and subtracting in the Cantor set. College Mathematics Journal, 50(1):41–45, January 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2019.
1534505.

Sanghi:2021:VCR

[San21] Aniket Sanghi. Visualizing the complex roots of quadratic and cubic polynomial functions in three dimensions. College Mathematics Journal, 52(5):373–379, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2021.1969178.

Sangwin:2024:HFA

[San24] Chris Sangwin. On Heron’s formula for the area of a plane triangle. College Mathematics Journal, 55(5):406–415, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic).

Sarkar:2023:LQC

[Sar23] Jyotirmoy Sarkar. The largest quadrilateral is cyclic: a new geometric proof. College Mathematics Journal, 54(4):378–384, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2023.2239227.

Sastry:1981:DPL

[Sas81] K. R. S. Sastry. Distance from a point to a line. Two-Year College Mathematics Journal, 12(2):146–147, March 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972513.

Sastry:1991:TLP

[Sas91] K. R. S. Sastry. Triangles in a lattice parabola. College Mathematics Journal, 22(4):301–306, September 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1991.11973399.

Saul:1990:SWC

[Sau90] Mark E. Saul. Sin^{2}x: a sheep in wolf’s clothing. College Mathematics Journal, 21(1):43–44, January 1990. CODEN ????

ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1990.
11973283.

Saul:1994:MH

[Sau94] Mark E. Saul. Media highlights. College Mathematics Journal, 25(2):164–171, March 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1994.11973603.

Saul:1999:BR

[Sau99] Mark Saul. Book review. College Mathematics Journal, 30 (5):425–428, November 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1999.11974100.

Sawhney:2017:UPT

[Saw17] Mehtaab Sawhney. An unusual proof of the triangle inequality. College Mathematics Journal, 49(3):218, May 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2017.1397993.

Sayrafiezadeh:1995:GMV

[Say95] M. Sayrafiezadeh. A generalization of the mean value theorem for integrals. College Mathematics Journal, 26(3):223–224, May 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1995.11973701.

Schultz:1984:APS

[SB84] James E. Schultz and William F. Burger. An approach to problem-solving using equivalence classes modulo n. College Mathematics Journal, 15(5):401–405, November 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972821.

Singh:1997:DCL

[SB97] Sahib Singh and Dip Bhattacharya. On dividing coconuts: a linear Diophantine problem. College Mathematics Journal, 28(3): 203–204, May 1997. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1997.11973863.

Stewart:2003:BAN

[SC03] Jack Stewart and Annalisa Crannell. The band around a (non)convex set. College Mathematics Journal, 34(5):377–379, November 2003. CODEN ???? ISSN 0746-8342 (print), 1931-

1346 (electronic). URL http://www.tandfonline.com/doi/
abs/10.1080/07468342.2003.11922034.

Scavo:1991:MH

[Sca91] Thomas Scavo. Media highlights. College Mathematics Journal, 22(1):77–84, January 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1991.11973364.

Schmidt:1970:CST

[Sch70] Stanley Schmidt. A curriculum suggestion for teaching College arithmetic. Two-Year College Mathematics Journal, 1(1):92–95, Spring 1970. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1970.11973175.

Schaumberger:1972:ACP

[Sch72a] Norman Schaumberger. An alternate classroom proof of the familiar limit for e. Two-Year College Mathematics Journal, 3(2): 72–73, Autumn 1972. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1972.11974220.

Schmalz:1972:CNR

[Sch72b] Rosemary Schmalz, S. P. Complex numbers as residue classes of polynomials mod(x^{2} + 1). Two-Year College Mathematics Journal, 3(2):78–80, Autumn 1972. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1972.11974223.

Schoen:1972:GAO

[Sch72c] Harold L. Schoen. A geometric approach to the orders of infinity. Two-Year College Mathematics Journal, 3(2):74–76, Autumn 1972. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1972.11974221.

Schaumberger:1973:EWN

[Sch73] Norman Schaumberger. Evaluating ∫_{a}^{b} x^{k} dx where k is any negative integer other than −1. Two-Year College Mathematics Journal, 4(2):91–93, Spring 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1973.11974262.

Schaumberger:1974:CTT

[Sch74a] Norman Schaumberger. A classroom theorem on trigonometric irrationalities. Two-Year College Mathematics Journal, 5(1): 73–76, Winter 1974. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1974.11974298.

Schaumberger:1974:STI

[Sch74b] Norman Schaumberger. A set of trigonometric inequalities with applications to maxima and minima. Two-Year College Mathematics Journal, 5(3):26–30, September 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1974.
11974332.

Schaumberger:1974:SCE

[Sch74c] Norman Schaumberger. Some comments on the exceptional case in a basic integral formula. Two-Year College Mathematics Journal, 5(3):58, September 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1974.11974340.

Schoen:1974:FNO

[Sch74d] Harold L. Schoen. A further note on the orders of infinity. Two-Year College Mathematics Journal, 5(1):80–82, Winter 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1974.11974301.

Schaumberger:1976:GAB

[Sch76] Norman Schaumberger. A geometrical approach to a basic limit. Two-Year College Mathematics Journal, 7(1):11–12, February 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1976.11974412.

Schaumberger:1977:SCP

[Sch77] Norman Schaumberger. Some consequences of a property of the centroid of a triangle. Two-Year College Mathematics Journal, 8(3):142–144, June 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1977.11974497.

Schaumberger:1978:CPA

[Sch78a] Norman Schaumberger. A calculus proof of the arithmetic– geometric mean inequality. Two-Year College Mathematics Journal, 9(1):16–17, January 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1978.11974538.

Schoenfeld:1978:SAT

[Sch78b] Alan H. Schoenfeld. A simple antidifferentiation technique. Two-Year College Mathematics Journal, 9(2):104–105, March 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1978.11974557.

Schaumberger:1980:ADD

[Sch80a] Norman Schaumberger. Another derivation of a double inequality. Two-Year College Mathematics Journal, 11(4):273, September 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1980.11972458.

Schaumberger:1980:PA

[Sch80b] Norman Schaumberger. A precalculus approximation of n!. Two-Year College Mathematics Journal, 11(3):202–204, June 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1980.11972437.

Schaumberger:1981:CA

[Sch81a] Norman Schaumberger. A classroom approach to x^{2} + y^{2} + z^{2} = w^{2}. Two-Year College Mathematics Journal, 12(5):331– 332, November 1981. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1981.11972563.

Schaumberger:1981:CGE

[Sch81b] Norman Schaumberger. A coordinate geometry evaluation of | tan(A − B)|. Two-Year College Mathematics Journal, 12(1): 52–54, January 1981. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1981.11972493.

Schmidt:1981:UPS

[Sch81c] Garfield C. Schmidt. Uniqueness of power series representations. Two-Year College Mathematics Journal, 12(1):54–56, January 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972494.

Schulz:1981:VIQ

[Sch81d] William C. Schulz. Vector identities from quaternions. Two-Year College Mathematics Journal, 12(4):271–273, September 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972545.

Schaumberger:1982:CAP

[Sch82a] Norman Schaumberger. A classroom approach to Pythagorean triples. Two-Year College Mathematics Journal, 13(1):61–63, January 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1982.11972582.

Schaumberger:1982:D

[Sch82b] Norman Schaumberger. The derivative of Arctanx. Two-Year College Mathematics Journal, 13(4):274–276, September 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1982.11972630.

Schaumberger:1982:SAP

[Sch82c] Norman Schaumberger. Still another proof of the arithmetic– geometric mean inequality. Two-Year College Mathematics Journal, 13(2):159–160, March 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1982.11972599.

Schaumberger:1983:IH

[Sch83a] Norman Schaumberger. Instant hindsight! Two-Year College Mathematics Journal, 14(4):351, September 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1983.
11972715.

Scheding:1983:SPM

[Sch83b] John Scheding. SSD persistence: a mathematical system for student investigation. Two-Year College Mathematics Journal, 14 (4):309–312, September 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1983.11972703.

Schultz:1983:AAP

[Sch83c] Phillip Schultz. An algebraic approach to partial fractions. Two-Year College Mathematics Journal, 14(4):346–348, September 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1983.11972712.

Schaumberger:1984:AAT

[Sch84a] Norman Schaumberger. Alternate approaches to two familiar results. College Mathematics Journal, 15(5):422–423, November 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972825.

Schaumberger:1984:AL

[Sch84b] Norman Schaumberger. Another look at x^{1/x}. College Mathematics Journal, 15(3):249–250, June 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1984.
11972778.

Schaumberger:1984:D

[Sch84c] Norman Schaumberger. The derivatives of Sinx and Cosx. College Mathematics Journal, 15(2):143–145, March 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972765.

Schwartzman:1984:FTP

[Sch84d] Steven Schwartzman. The factorial triangle and polynomial sequences. College Mathematics Journal, 15(5):424–426, November 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972826.

Schaumberger:1985:ASE

[Sch85a] Norman Schaumberger. Approximating solutions for exponential equations. College Mathematics Journal, 16(3):211–212, June 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1985.11972883.

Schaumberger:1985:IP

[Sch85b] Norman Schaumberger. An instant proof of e^{π} > π^{e}. College Mathematics Journal, 16(4):280, September 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1985.
11972896.

Schaumberger:1985:MAM

[Sch85c] Norman Schaumberger. More applications of the mean value theorem. College Mathematics Journal, 16(5):397–398, November 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1985.11972914.

Schoenberg:1985:HTD

[Sch85d] I. J. Schoenberg. On hypocycloids and their diameters. College Mathematics Journal, 16(4):262–267, September 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1985.11972892.

Schaefer:1986:SRS

[Sch86a] Paul Schaefer. Sums of rearranged series. College Mathematics Journal, 17(1):66–70, January 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1986.
11972931.

Schaumberger:1986:AAC

[Sch86b] Norman Schaumberger. Another approach to a class of slowly diverging series. College Mathematics Journal, 17(5):417, November 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1986.11972991.

Schaumberger:1986:D

[Sch86c] Norman Schaumberger. The derivatives of Arcsecx, Arctanx, and Tanx. College Mathematics Journal, 17(3):244–246, May 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1986.11972965.

Schaumberger:1986:TFF

[Sch86d] Norman Schaumberger. Three familiar formulas for π via geometry. College Mathematics Journal, 17(4):339, September 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1986.11972976.

Schoenberg:1986:MH

[Sch86e] I. J. Schoenberg. Media highlights. College Mathematics Journal, 17(2):190–199, March 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1986.11972956.

Schulz:1986:APV

[Sch86f] William C. Schulz. An alternate proof of the vector triple product formula. College Mathematics Journal, 17(1):73–74, January 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1986.11972933.

Schaumberger:1987:ES

[Sch87a] Norman Schaumberger. Extending the series for in 2. College Mathematics Journal, 18(3):223–225, May 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/07468342.
1987.11973037.

Schoenberg:1987:CRP

[Sch87b] Isaac J. Schoenberg. The Chinese Remainder Problem and polynomial interpolation. College Mathematics Journal, 18(4):320– 322, September 1987. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1987.11973053.

Schaumberger:1988:API

[Sch88a] Norman Schaumberger. Another proof of the inequality between power means. College Mathematics Journal, 19(1):56–58, January 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1988.11973092.

Schaumberger:1988:GFA

[Sch88b] Norman Schaumberger. A general form of the arithmetic– geometric mean inequality via the mean value theorem. College Mathematics Journal, 19(2):172–173, March 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1988.
11973110.

Schattschneider:1989:MH

[Sch89a] Doris Schattschneider. Media highlights. College Mathematics Journal, 20(2):174–181, March 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1989.
11973228.

Schaumberger:1989:GI

[Sch89b] Norman Schaumberger. The AM–GM inequality via x^{1/x}. College Mathematics Journal, 20(4):320, September 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1989.
11973249.

Schaumberger:1989:APC

[Sch89c] Norman Schaumberger. Another proof of Chebysheff’s Inequality. College Mathematics Journal, 20(2):141–142, March 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1989.11973223.

Schaumberger:1989:G

[Sch89d] Norman Schaumberger. A generalization of lim_{n→∞} ^{n} \sqrt{n!}/n = e^{−1}. College Mathematics Journal, 20(5):416–418, November 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1989.11973269.

Schoenfeld:1989:CFA

[Sch89e] Alan H. Schoenfeld. The curious fate of an applied problem. College Mathematics Journal, 20(2):115–123, March 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1989.11973217.

Schilling:1990:LRH

[Sch90a] Mark F. Schilling. The longest run of heads. College Mathematics Journal, 21(3):196–207, May 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1990.
11973306.

Schonefeld:1990:SEI

[Sch90b] Steven Schonefeld. Some examples illustrating Richardson’s improvement. College Mathematics Journal, 21(4):314–322, September 1990. CODEN ???? ISSN 0746-8342 (print), 1931-

1346 (electronic). URL http://www.tandfonline.com/doi/
abs/10.1080/07468342.1990.11973327.

Schumer:1990:SSC

[Sch90c] Peter Schumer. Strings of strongly composite integers and invisible lattice points. College Mathematics Journal, 21(1):37–40, January 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1990.11973281.

Schattschneider:1991:CIT

[Sch91a] Doris Schattschneider. Counting it twice. College Mathematics Journal, 22(3):203–211, May 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1991.
11973382.

Schaumberger:1991:APF

[Sch91b] Norman Schaumberger. Another proof of a familiar inequality. College Mathematics Journal, 22(3):229–230, May 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1991.11973387.

Schielack:1992:SGS

[Sch92a] Vincent P. Schielack, Jr. Summing geometric series by holding a tournament. College Mathematics Journal, 23(3):210–211, May 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1992.11973460.

Schonefeld:1992:RAJ

[Sch92b] Steven Schonefeld. Rotation of axes-not just for conics. College Mathematics Journal, 23(5):418–425, November 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1992.11973494.

Schaumberger:1994:APF

[Sch94a] Norman Schaumberger. Another proof of the formula e = ∑_{0}^{∞}(1/n!). College Mathematics Journal, 25(1):38–39, January 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1994.11973578.

Schwartz:1994:DPC

[Sch94b] Mark Schwartz. On the distance from a point to a curve. College Mathematics Journal, 25(4):317–319, September 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1994.11973629.

Schonefeld:1995:EPE

[Sch95a] Steven Schonefeld. Eigenpictures: Picturing the eigenvector problem. College Mathematics Journal, 26(4):316–319, September 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1995.11973723.

Schuster:1995:MPM

[Sch95b] Eugene F. Schuster. MAD property of medians: an induction proof. College Mathematics Journal, 26(5):387–389, November 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1995.11973736.

Schwartz:1995:CAR

[Sch95c] Mark Schwartz. The chair, the area rug, and the astroid. College Mathematics Journal, 26(3):229–231, May 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1995.11973704.

Schwartzman:1995:NWE

[Sch95d] Steven Schwartzman. Number words in English. College Mathematics Journal, 26(3):191–195, May 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1995.
11973694.

Schumer:1997:PB

[Sch97a] Peter Schumer. The pen and the barn. College Mathematics Journal, 28(3):205–206, May 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1997.
11973864.

Schwenk:1997:ILW

[Sch97b] Allen J. Schwenk. Introduction to limits, or why can’t we just trust the table? College Mathematics Journal, 28(1):51, January 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1997.11973830.

Schay:1999:UCC

[Sch99a] Geza Schay. Using consistency conditions to solve linear systems. College Mathematics Journal, 30(3):226–229, May 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1999.11974064.

Schwenk:1999:SSS

[Sch99b] Allen J. Schwenk. Several Sets of n + 1 Shapes, Each the Similitude Union of the Other n. College Mathematics Journal, 30(2): 112–117, March 1999. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1999.11974041.

Schwenk:2000:PRM

[Sch00] Allen J. Schwenk. A polynomial with a root mod m for every m. College Mathematics Journal, 31(5):403–405, November 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2000.11974181.

Schueller:2004:HLW

[Sch04] Albert Schueller. How long was your day? College Mathematics Journal, 35(1):31–33, January 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2004.
11922048.

Schlatter:2006:HVF

[Sch06] Mark Schlatter. How to view a Flatland painting. College Mathematics Journal, 37(2):114–120, March 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2006.
11922174.

Schechter:2007:WWJ

[Sch07] Murray Schechter. Which way is Jerusalem? navigating on a spheroid. College Mathematics Journal, 38(2):96–105, March 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2007.11922224.

Schilling:2011:BRL

[Sch11a] Kenneth Schilling. Book review: Logical Labyrinths, by Raymond M. Smullyan. A K Peters. College Mathematics Journal, 42(2):159–160, March 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.42.2.159.

Schwartz:2011:DDK

[Sch11b] Mark Schwartz. Do dogs know the trammel of Archimedes? College Mathematics Journal, 42(4):299–308, September 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.42.4.299.

Schaffer:2012:PSS

[Sch12] Karl Schaffer. A Platonic sextet for strings. College Mathematics Journal, 43(1):64–69, January 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
43.1.064.

Schilling:2013:DSP

[Sch13] Kenneth Schilling. Derivative sign patterns in two dimensions. College Mathematics Journal, 44(2):102–108, March 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.44.2.102.

Scheinberg:2016:GBP

[Sch16] Stephen Scheinberg. The generalized birthday problem. College Mathematics Journal, 47(4):294–296, September 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
47.4.294.

Schmidt:2021:TFB

[Sch21a] Phillip H. Schmidt. Two friends and a bike. College Mathematics Journal, 52(1):11–21, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2021.1848198.

Schmitz:2021:PFC

[Sch21b] Michael Schmitz. A plea for finite calculus. College Mathematics Journal, 52(2):94–105, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2021.1866937.

Schultz:2023:PWA

[Sch23] James Schultz. Proof without words: The American flag inspires a proof that the sum of the first n positive odd integers is n^{2}. College Mathematics Journal, 54(5):481, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2023.2273197.

Scott:1987:ELP

[Sco87] P. R. Scott. Equiangular lattice polygons and semiregular lattice polyhedra. College Mathematics Journal, 18(4):300–306, September 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1987.11973047.

Schwertman:1993:QCI

[SD93] Neil C. Schwertman and Larry R. Dion. Quadratic confidence intervals. College Mathematics Journal, 24(5):453–457, November 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1993.11973568.

Schmalzried:1985:SR

[SDC85] Cynthia Schmalzried, James A. Deddens, and Thomas J. Carter. Software reviews. College Mathematics Journal, 16(5):413–415, November 1985. CODEN ???? ISSN 0746-8342 (print), 1931-

1346 (electronic). URL http://www.tandfonline.com/doi/
abs/10.1080/07468342.1985.11972919.

Sansgiry:1996:HHM

[SE96] Prashant S. Sansgiry and Constance C. Edwards. A home heating model for calculus students. College Mathematics Journal, 27 (5):394–397, November 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1996.11973817.

Sedinger:1980:DL

[Sed80] Harry Sedinger. Derivatives without limits. Two-Year College Mathematics Journal, 11(1):54–55, January 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1980.
11972404.

Seery:1998:PRG

[See98] Michael J. Seery. Pursuit and regular N-gons. College Mathematics Journal, 29(3):228–229, May 1998. CODEN ????

ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1998.
11973947.

Segal:1982:MC

[Seg82a] Arthur C. Segal. Motivating e by calculator. Two-Year College Mathematics Journal, 13(4):271, September 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1982.
11972627.

Segal:1982:NSS

[Seg82b] Arthur C. Segal. A note on the surface of a sphere. Two-Year College Mathematics Journal, 13(1):63–64, January 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1982.11972583.

Segal:1983:CFF

[Seg83] Arthur C. Segal. Closed-form formulas for quasi-geometric series. Two-Year College Mathematics Journal, 14(2):118–122, March 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1983.11972667.

Segal:1987:LDM

[Seg87] Arthur C. Segal. A linear diet model. College Mathematics Journal, 18(1):44–45, January 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1987.11973005.

Segal:1989:SS

[Seg89] Arthur C. Segal. Subharmonic series. College Mathematics Journal, 20(3):194–200, May 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1989.11973230.

Segal:2000:QCT

[Seg00] Arthur C. Segal. A quick construction of tangents to an ellipse. College Mathematics Journal, 31(2):131, March 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2000.11974130.

Sekino:2001:BAC

[Sek01] Junpei Sekino. The band around a convex set. College Mathematics Journal, 32(2):110–114, March 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2001.
11921862.

Seligman:1970:ILS

[Sel70] Aaron Seligman. If … some suggestions on presenting the connector “if … then”. Two-Year College Mathematics Journal, 1 (2):22–26, Autumn 1970. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1970.11973182.

Selden:1991:MH

[Sel91] Annie Selden. Media highlights. College Mathematics Journal, 22(3):261–268, May 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1991.11973395.

Semmes:1974:GAL

[Sem74] Pat Semmes. A geometric approach to linear programming in the two-year College. Two-Year College Mathematics Journal, 5(1): 37–40, Winter 1974. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1974.11974290.

Senechal:1995:MH

[Sen95] Marjorie Senechal. Media highlights. College Mathematics Journal, 26(1):73–81, January 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1995.11973672.

Serna:2015:PPW

[Ser15] Juan D. Serna. Partial proof without words: Shaping some cases of the Erdős–Straus conjecture. College Mathematics Journal, 46(3):181, May 2015. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.46.3.181.

Servatius:2017:BRT

[Ser17] Brigitte Servatius. Book review: A Treatise of Conic Sections by George Salmon, Hodges and Smith, Dublin, 1848.

Reprint of sixth edition, American Mathematical Society, Providence, RI, 2005. xv + 399 pp., ISBN 978-0-8218-3777-1, $54.00 (hardback). College Mathematics Journal, 49(1):68–72, January 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2017.1389134.

Setek:1974:LLE

[Set74] William M. Setek, Jr. Logic: a logical elective. Two-Year College Mathematics Journal, 5(3):39–40, September 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1974.11974335.

Sewell:2018:DBS

[Sew18] Granville Sewell. Derivation of the Black–Scholes equation from basic principles. College Mathematics Journal, 49(3):212–215, May 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2018.1439634.

Swanson:1981:MIW

[SH81] Leonard G. Swanson and Rodney T. Hansen. Mathematical induction or “what good is all this stuff if we are going to assume it’s true anyway?”. Two-Year College Mathematics Journal, 12(1):8–12, January 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1981.11972482.

Smith:1985:TSL

[SH85] James Smith and Mason Henderson. Tetrahedra, skew lines, and volume. College Mathematics Journal, 16(2):138–140, March 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1985.11972870.

Seppala-Holtzman:2010:DF

[SH10] David Seppala-Holtzman. The dance of the foci. College Mathematics Journal, 41(2):122–128, March 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.

tandfonline.com/doi/abs/10.4169/074683410X480221.

Seppala-Holtzman:2012:OBF

[SH12] David Seppala-Holtzman. An optimal basketball free throw. College Mathematics Journal, 43(5):387–394, November 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.43.5.387.

Seppala-Holtzman:2014:PWE

[SH14] David Seppala-Holtzman. Proof without words: Ensphering three capped prisms. College Mathematics Journal, 45(1):49, January 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.45.1.049.

Seppala-Holtzman:2018:CCF

[SH18] D. N. Seppala-Holtzman. A canonical conical function. College Mathematics Journal, 49(5):359–362, November 2018. CO-DEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/full/10.1080/07468342.
2018.1516979.

Shaughnessy:1980:AIA

[Sha80] J. Michael Shaughnessy. Applications of intermediate algebra: a possible alternative. Two-Year College Mathematics Journal, 11(2):94–101, March 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1980.11972412.

Shaw:1995:HFP

[Sha95] Howard Shaw. Hyperbolic functions and proper time in relativity. College Mathematics Journal, 26(4):312–315, September 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1995.11973721.

Shao:2015:PWB

[Sha15] Meiyue Shao. Proof without words: Bounding the Euler– Mascheroni constant. College Mathematics Journal, 46(5):347, November 2015. CODEN ???? ISSN 0746-8342 (print), 1931-

1346 (electronic). URL http://www.tandfonline.com/doi/
abs/10.4169/college.math.j.46.5.347.

Shaw:2018:HGG

[Sha18] Douglas J. Shaw. Harris graphs — a graph theory activity for students and their instructors. College Mathematics Journal, 49 (5):323–326, November 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2018.1507382.

Shankar:2025:LAB

[Sha25] P. M. Shankar. Linear algebra and Bayes’ Theorem in probability. College Mathematics Journal, 56(1):49–59, 2025. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic).

Sher:1975:SWS

[She75] Lawrence Sher. Solving whodunits by symbolic logic. Two-Year College Mathematics Journal, 6(4):36–38, December 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1975.11974405.

Sher:1976:LE

[She76] Lawrence Sher. Letters to the Editor. Two-Year College Mathematics Journal, 7(1):1, February 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1976.
11974409.

Sher:1977:PCL

[She77] Lawrence Sher. The perfect curve: at least for grades. Two-Year College Mathematics Journal, 8(3):148–152, June 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1977.11974499.

Sher:1978:BAP

[She78] Lawrence Sher. Boolean algebra as a proof paradigm. Two-Year College Mathematics Journal, 9(3):186–190, June 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1978.11974574.

Shepard:1997:RRR

[She97] Melissa Shepard. A Rose is a Rose is a Rose …. College Mathematics Journal, 28(1):55–56, January 1997. CODEN ????

ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1997.
11973832.

Sheard:2009:TT

[She09] Michael Sheard. Trick or technique? College Mathematics Journal, 40(1):10–14, January 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2009.11922329.

Shekhawat:2022:CS

[She22] Hanumant Singh Shekhawat. Correlation and subtraction. College Mathematics Journal, 53(3):233–234, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2022.
2045151.

Shilgalis:1978:FCI

[Shi78] Thomas W. Shilgalis. Flow chart for infinite series. Two-Year College Mathematics Journal, 9(3):191, June 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/00494925.
1978.11974575.

Shiflet:1988:MN

[Shi88] Angela B. Shiflet. Musical notes. College Mathematics Journal, 19(4):345–347, September 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1988.11973135.

Shi:2010:SDN

[Shi10] Yixun Shi. Sum-difference numbers. College Mathematics Journal, 41(5):404–405, November 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.

tandfonline.com/doi/abs/10.4169/074683410X522026.

Small:1986:CAS

[SHL86] Don Small, John Hosack, and Kenneth Lane. Computer algebra systems in undergraduate instruction. College Mathematics Journal, 17(5):423–433, November 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1986.
11972995.

Shpiro:2014:UPF

[Shp14] Asya Shpiro. Unfamiliar properties of familiar shapes. College Mathematics Journal, 45(5):371–375, November 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
45.5.371.

Shultz:1979:EVP

[Shu79] Harris S. Shultz. An expected value problem. Two-Year College Mathematics Journal, 10(4):277–278, September 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1979.11972366.

Shulman:2002:TEP

[Shu02] Bonnie Shulman. Is there enough poison gas to kill the city?: The teaching of ethics in mathematics classes. College Mathematics Journal, 33(2):118–125, March 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2002.
11921929.

Sibley:2012:IM

[Sib12] Thomas Q. Sibley. Idempotents à la mod. College Mathematics Journal, 43(5):401–404, November 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
43.5.401.

Siehler:2012:FLG

[Sie12] Jacob A. Siehler. The finite lamplighter groups: a guided tour. College Mathematics Journal, 43(3):203–211, May 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
43.3.203.

Siehler:2024:NPP

[Sie24] Jacob A. Siehler. New problems in port-and-sweep solitaire. College Mathematics Journal, 55(4):294–302, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2023.2235226.

Silverman:1989:SAF

[Sil89] Herb Silverman. A simple auxiliary function for the mean value theorem. College Mathematics Journal, 20(4):323, September 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1989.11973252.

Silverman:1990:TIT

[Sil90] Herb Silverman. Trigonometric identities through calculus. College Mathematics Journal, 21(5):403, November 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1990.
11973343.

Simmonds:1995:NLO

[Sim95] J. G. Simmonds. A new look at an old function, e^{iθ}. College Mathematics Journal, 26(1):6–10, January 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/07468342.
1995.11973658.

Simoson:1997:ES

[Sim97] Andrew Simoson. An envelope for a spirograph. College Mathematics Journal, 28(2):134–139, March 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1997.
11973852.

Simoson:2000:FFF

[Sim00] Andrew Simoson. Fast-food-frusta and the center of gravity. College Mathematics Journal, 31(4):303–306, September 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2000.11974164.

Simmonds:2002:PAD

[Sim02] J. G. Simmonds. Perfecting the analog of a deck of cards or why evolution can’t be left to chance. College Mathematics Journal, 33(1):17–20, January 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2002.11921913.

Simoson:2006:PBS

[Sim06] Andrew J. Simoson. Playing ball in a space station. College Mathematics Journal, 37(5):334–343, November 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2006.11922203.

Simoson:2007:PCM

[Sim07] Andrew J. Simoson. Pursuit curves for the man in the moone. College Mathematics Journal, 38(5):330–338, November 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2007.11922257.

Simoson:2009:MSW

[Sim09a] Andrew J. Simoson. Maximizing the spectacle of water fountains. College Mathematics Journal, 40(4):263–274, September 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2009.11922374.

Simoson:2009:SS

[Sim09b] Andrew J. Simoson. Solomon’s Sea and π. College Mathematics Journal, 40(1):22–32, January 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2009.
11922332.

Simmons:2011:AMB

[Sim11a] Charlotte Simmons. Augustus De Morgan behind the scenes. College Mathematics Journal, 42(1):33–40, January 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.42.1.033.

Simoson:2011:NRM

[Sim11b] Andrew J. Simoson. Newton’s radii, Maupertuis’ arc length, and Voltaire’s giant. College Mathematics Journal, 42(3):183–190, May 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.42.3.183.

Simoson:2019:EP

[Sim19] Andrew J. Simoson. Extrapolating Plimpton 322. College Mathematics Journal, 50(3):210–220, 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2019.
1582990.

Siner:1976:OU

[Sin76] Helen B. Siner. The Open University. Two-Year College Mathematics Journal, 7(3):28–32, September 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1976.
11974443.

Singmaster:1990:TIS

[Sin90] David Singmaster. Triangles with integer sides and sharing barrels. College Mathematics Journal, 21(4):278–285, September 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1990.11973320.

Singh:2012:FWS

[Sin12] Satyanand Singh. A fifth way to skin a definite integral. College Mathematics Journal, 43(5):377–378, November 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
43.5.377.

Siu:1981:II

[Siu81] Man-Keung Siu. From an inequality to inversion. Two-Year College Mathematics Journal, 12(2):149–151, March 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972515.

Siu:1984:SC

[Siu84] Man-Keung Siu. On the sphere and cylinder. College Mathematics Journal, 15(4):326–328, September 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1984.
11972792.

Serna:2011:CMS

[SJ11] Juan D. Serna and Amitabh Joshi. The center of mass of a soft spring. College Mathematics Journal, 42(5):389–394, November 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.42.5.389.

Saenz:2000:APR

[SJC00] Gillian Saenz, Christopher Jackson, and Ryan Crumley. The asymmetric propeller revisited. College Mathematics Journal, 31(5):347–349, November 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2000.11974172.

Spero:1977:CC

[SK77] Samuel W. Spero and Mary Koehler. Computer corner. Two-Year College Mathematics Journal, 8(3):182–187, June 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1977.11974506.

Schaumberger:1989:APJ

[SK89] Norman Schaumberger and Bert Kabak. Another proof of Jensen’s Inequality. College Mathematics Journal, 20(1):57–58, January 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1989.11973207.

Strizhevsky:2001:GPG

[SK01] Michael Strizhevsky and Dmitry Kreslavskiy. Geometric progressions — a geometric approach. College Mathematics Journal, 32 (5):359–362, November 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2001.11921904.

Skala:1991:CMV

[Ska91] Helen Skala. Contour maps — a visual experience. College Mathematics Journal, 22(3):241–244, May 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1991.
11973391.

Skala:1996:WRB

[Ska96] Helen Skala. Will the real best fit curve please stand up? College Mathematics Journal, 27(3):220–223, May 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1996.11973784.

Skala:1997:D

[Ska97] Helen Skala. A discover-e. College Mathematics Journal, 28(2): 128–129, March 1997. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1997.11973849.

Schmidt:1984:PRA

[SL84] Robert Schmidt and Robert Lacher. Probabilistic repeatability among some irrationals. College Mathematics Journal, 15 (4):330–332, September 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1984.11972795.

Sloyer:1976:WAM

[Slo76] Clifford Sloyer. What is an application of mathematics? Two-Year College Mathematics Journal, 7(3):19–26, September 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1976.11974441.

Slonneger:1977:TR

[Slo77] Kenneth Slonneger. Types of relations. Two-Year College Mathematics Journal, 8(5):267–269, November 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1977.
11974525.

Subramaniam:1991:RFR

[SM91] T. N. Subramaniam and D. E. G. Malm. Reduction formulas revisited. College Mathematics Journal, 22(5):421–429, November 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1991.11973422.

Smith:2000:LD

[SM00] Brian E. Smith and Francis J. Merceret. The lognormal distribution. College Mathematics Journal, 31(4):259–261, September 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2000.11974155.

Sandefur:2013:CI

[SM13] James Sandefur and John Mason. Circular inclusion. College Mathematics Journal, 44(3):193–201, May 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
44.3.193.

sarah-mariebelcastro:2012:CSS

[smb12] sarah-marie belcastro. The continuing saga of snarks. College Mathematics Journal, 43(1):82–87, January 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
43.1.082.

sarah-mariebelcastro:2002:CFP

[smbH02] sarah-marie belcastro and Thomas C. Hull. Classifying Frieze patterns without using groups. College Mathematics Journal, 33 (2):93–98, March 2002. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2002.11921925.

sarah-mariebelcastro:2016:DTP

[smbV16] sarah-marie belcastro and Tamara Veenstra. Do the twist! (on polygon-base boxes). College Mathematics Journal, 47(5):340– 345, November 2016. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.47.5.340.

Smith:1972:PK

[Smi72] Karl J. Smith. A program for keno. Two-Year College Mathematics Journal, 3(2):16–20, Autumn 1972. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1972.
11974210.

Smith:1973:PT

[Smi73] Karl J. Smith. Pascal’s Triangle. Two-Year College Mathematics Journal, 4(1):1–13, Winter 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1973.11974228.

Smith:1978:HPR

[Smi78] David A. Smith. The homicide problem revisited. Two-Year College Mathematics Journal, 9(3):141–145, June 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1978.11974564.

Smith:1980:OBC

[Smi80] Lehi T. Smith. Orthogonal basis: a computational alternative. Two-Year College Mathematics Journal, 11(4):274, September 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1980.11972459.

Smith:1981:MMM

[Smi81] Betsy Darken Smith. Misguided mathematical Maxim-makers. Two-Year College Mathematics Journal, 12(5):309–316, November 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972557.

Smith:1986:RLA

[Smi86] Robert S. Smith. Rolle over Lagrange-another shot at the mean value theorem. College Mathematics Journal, 17(5):403– 406, November 1986. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1986.11972987.

Smith:1987:NRM

[Smi87] David A. Smith. Newton’s n-th root method without derivatives. College Mathematics Journal, 18(5):403–406, November 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1987.11973064.

Smith:1989:WSA

[Smi89] David A. Smith. What’s significant about a digit? College Mathematics Journal, 20(2):136–139, March 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1989.
11973221.

Smith:1991:PES

[Smi91] Henry Mark Smith. Positivity from evaluation of a single point. College Mathematics Journal, 22(3):230–231, May 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1991.11973388.

Smith:2000:NLC

[Smi00] Kirby C. Smith. Normal lines and curvature. College Mathematics Journal, 31(1):54–56, January 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2000.
11974110.

Smither:2011:SPC

[Smi11] Robert K. Smither. The symmedian point: Constructed and applied. College Mathematics Journal, 42(2):115–117, March 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.42.2.115.

Smith:2014:MF

[Smi14] Charlie Smith. Math frenzy. College Mathematics Journal, 45 (4):276–277, September 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.45.4.276.

Smorynski:1985:NA

[Smo85] Craig Smoryński. The need for abstraction. College Mathematics Journal, 16(1):11–12, January 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1985.
11972843.

Smolinsky:2024:RCC

[Smo24] Lawrence Smolinsky. Rearrangements of a conditionally convergent series summing to logarithms of natural numbers. College Mathematics Journal, 55(1):59–61, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2023.
2223513.

Smullyan:1982:ADTa

[Smu82a] Raymond Smullyan. The asylum of Doctor Tarr and Professor Fether. Two-Year College Mathematics Journal, 13(2):142–145, March 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1982.11972593.

Smullyan:1982:ADTb

[Smu82b] Raymond Smullyan. The asylum of Doctor Tarr and Professor Fether: Solutions. Two-Year College Mathematics Journal, 13(3):213–217, June 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1982.11972615.

Smullyan:1985:SCI

[Smu85] Raymond M. Smullyan. Satan, Cantor, and infinity. College Mathematics Journal, 16(2):118–121, March 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1985.
11972864.

Snee:1988:MOO

[Sne88] Ronald D. Snee. Mathematics is only one tool that statisticians use …. College Mathematics Journal, 19(1):30–32, January 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1988.11973085.

Snipes:1977:DIC

[Sni77] Ray F. Snipes. On the definition of implication: Classroom discussion and justification. Two-Year College Mathematics Journal, 8(4):247–252, September 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1977.
11974520.

Snow:1989:SPS

[Sno89] Donald R. Snow. Spreadsheets, power series, generating functions, and integers. College Mathematics Journal, 20(2):143–152, March 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1989.11973224.

Snyder:2024:MSA

[Sny24] Jason Snyder. A modern spin on Archimedes’ quadrature of the parabola. College Mathematics Journal, 55(2):134–139, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2023.2278998.

Sobczyk:1995:HNP

[Sob95] Garret Sobczyk. The hyperbolic number plane. College Mathematics Journal, 26(4):268–280, September 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1995.
11973712.

Sobczyk:1997:GSD

[Sob97] Garret Sobczyk. The generalized spectral decomposition of a linear operator. College Mathematics Journal, 28(1):27–38, January 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1997.11973825.

Solow:1994:SR

[Sol94] Anita E. Solow. Software reviews. College Mathematics Journal, 25(3):235–239, May 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1994.11973618.

Solow:1999:BR

[Sol99] Anita E. Solow. Book review. College Mathematics Journal, 30(1):77–80, January 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1999.11974036.

Solow:2023:HIP

[Sol23] Daniel Solow. Half an induction proof of the Collatz conjecture. College Mathematics Journal, 54(5):419–422, 2023.

CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2023.2263029.

Somanchi:1980:ADD

[Som80] L. N. Somanchi. An arithmetic description of the dihedral group. Two-Year College Mathematics Journal, 11(5):327–329, November 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1980.11972472.

Sommers:1999:SC

[Som99] Paul M. Sommers. Student’s t and crackers. College Mathematics Journal, 30(1):32–34, January 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1999.
11974026.

Sommers:2000:SBT

[Som00] Paul M. Sommers. The super bowl theory: Fourth and long. College Mathematics Journal, 31(3):189–192, May 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2000.11974141.

Sommers:2002:PGR

[Som02] Paul M. Sommers. Is Presidential greatness related to height? College Mathematics Journal, 33(1):14–16, January 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2002.11921912.

Sommers:2003:BAS

[Som03] Paul M. Sommers. Baseball’s all-stars: Birthplace and distribution. College Mathematics Journal, 34(1):24–30, January 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2003.11921981.

Soule:2014:SFS

[Sou14] Michelle Soule. A single family of semigroups with every positive rational commuting probability. College Mathematics Journal, 45(2):136–139, March 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.45.2.136.

Switney:1971:LE

[SP71] David T. Switney and Richard Plagge. Letters to the Editor. Two-Year College Mathematics Journal, 2(1):1, Spring 1971. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1971.11973988.

Scopelliti:1976:CC

[SP76] Pasquale Scopelliti and Herbert Peebles. Computer corner. Two-Year College Mathematics Journal, 7(3):52–54, September 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1976.11974449.

Sher:1979:BSA

[SP79] Lawrence Sher and Warren Page. Bounds for the sum of absolute standard scores. Two-Year College Mathematics Journal, 10(5):351–353, November 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1979.11972386.

Spangler:1971:LCC

[Spa71] Richard Spangler. Lower Columbia College Mathematics Laboratory. Two-Year College Mathematics Journal, 2(1):27–31, Spring 1971. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1971.11973993.

Spencer:1986:CCF

[Spe86] Joel Spencer. Combinatorics by coin flipping. College Mathematics Journal, 17(5):407–412, November 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1986.
11972988.

Spivey:2006:FID

[Spi06] Michael Z. Spivey. Fibonacci identities via the determinant sum property. College Mathematics Journal, 37(4):286–289, September 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2006.11922196.

Spivey:2010:DE

[Spi10] Michael Z. Spivey. Deranged exams. College Mathematics Journal, 41(3):197–202, May 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/074683410X488674.

Spivey:2016:CVI

[Spi16] Michael Spivey. The Chu–Vandermonde identity via Leibniz’s identity for derivatives. College Mathematics Journal, 47(3):219– 220, May 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.47.3.219.

Sporn:2016:ACN

[Spo16] Howard Sporn. Abbott-and-Costello numbers. College Mathematics Journal, 47(2):126–132, March 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
47.2.126.

Sporn:2017:PTC

[Spo17] Howard Sporn. Pythagorean triples, complex numbers, and perplex numbers. College Mathematics Journal, 48(2):115–122, March 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.48.2.115.

Sporn:2022:PRP

[Spo22] Howard Sporn. A proof of the reflective property of the parabola. College Mathematics Journal, 53(1):67–68, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2022.1991198.

Sporn:2023:FMM

[Spo23a]  Howard Sporn.  Forringer’s method and more trees of

Pythagorean triples. College Mathematics Journal, 54(4):269– 273, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2023.2225398.

Sporn:2023:GCA

[Spo23b] Howard Sporn. The golden circle and the aureate circle. College Mathematics Journal, 54(5):485–491, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2023.
2263318.

Sorensen:2010:SBS

[SR10] Jody Sorensen and Elyn Rykken. Sprinkler bifurcations and stability. College Mathematics Journal, 41(5):383–391, November 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/074683410X521982.

Sanchez-Reyes:2012:CR

[SR12a] Javier Sánchez-Reyes. The catenary as roulette. College Mathematics Journal, 43(3):216–219, May 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
43.3.216.

Sanchez-Reyes:2012:HSC

[SR12b] Javier Sánchez-Reyes. The hyperbolic sine cardinal and the catenary. College Mathematics Journal, 43(4):285–290, September 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.43.4.285.

Sharif-Rasslan:2016:VAG

[SR16] Amal Sharif-Rasslan. A visual approach to geometric series with negative ratio. College Mathematics Journal, 47(3):216– 218, May 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.47.3.216.

Sridharma:1999:D

[Sri99] Selvaratnam Sridharma. The derivative of Sinθ. College Mathematics Journal, 30(4):314–315, September 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1999.
11974078.

Srinivasan:2010:BRE

[Sri10] Bhama Srinivasan. Book review: Emmy Noether: The Mother of Modern Algebra, by M. B. W. Tent. College Mathematics Journal, 41(1):72–73, January 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/074683410X475155.

Sandell:1977:FOS

[SS77] Donn C. Sandell and F. Max Stein. Factorization of operators of second order linear homogeneous ordinary differential equations. Two-Year College Mathematics Journal, 8(3):132–141, June 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1977.11974496.

Slay:1979:NCS

[SS79] Jack C. Slay and J. L. Solomon. A note on the Cauchy–Schwarz inequality. Two-Year College Mathematics Journal, 10(4):280– 283, September 1979. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1979.11972369.

Sher:1981:SCF

[SS81a] Lawrence Sher and David Sher. Some conjectures on Fermat’s Last Conjecture. Two-Year College Mathematics Journal, 12(1): 51–52, January 1981. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1981.11972492.

Slay:1981:MGF

[SS81b] Jack C. Slay and J. L. Solomon. A mean generating function. Two-Year College Mathematics Journal, 12(1):27–29, January 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972487.

Schaumberger:1982:ELV

[SS82] Norman Schaumberger and Norman Schaumberger. The Euler line: a vector approach. Two-Year College Mathematics Journal, 13(5):329–331, November 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1982.11972641.

Shiflett:1983:WDS

[SS83a] Ray C. Shiflett and Harris S. Shultz. When does a square give maximum area? Two-Year College Mathematics Journal, 14(3): 194–196, June 1983. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1983.11972681.

Sommer:1983:EFS

[SS83b] Bruce Sommer and Norman Schaumberger. Evaluating f sec x dx and f sec^{3} x dx. Two-Year College Mathematics Journal, 14(3):256–257, June 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1983.11972692.

Selden:1989:GH

[SS89] John Selden and Annie Selden. Graphing with the HP-28S. College Mathematics Journal, 20(5):423–432, November 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1989.11973271.

Selden:1993:CME

[SS93] Annie Selden and John Selden. Collegiate mathematics education research: What would that be like? College Mathematics Journal, 24(5):431–445, November 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1993.
11973564.

Smith:1994:PEP

[SS94] Kirby C. Smith and Vincent P. Schielack. Parametric equations and planar curves. College Mathematics Journal, 25(4):319– 321, September 1994. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1994.11973630.

Scholten:1996:FLP

[SS96] Paul Scholten and Andrew Simoson. The falling ladder paradox. College Mathematics Journal, 27(1):49–54, January 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1996.11973750.

Strzelecki:1999:CVD

[SS99] Pawel Strzelecki and Abe Shenitzer. Continuous versions of the (Dirichlet) drawer principle. College Mathematics Journal, 30 (3):195–196, May 1999. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1999.11974055.

Sanchez:2000:NLE

[SS00] David Sanchez and Kirby C. Smith. Normal lines and the evolute curve. College Mathematics Journal, 31(5):397–403, November 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2000.11974180.

Shultz:2005:MS

[SS05] Harris S. Shultz and Ray C. Shiflett. M&m sequences. College Mathematics Journal, 36(3):191–198, May 2005. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/07468342.
2005.11922127.

Schaefer:2017:BND

[SS17] Alex Schaefer and Jay Schweig. Balanced nontransitive dice. College Mathematics Journal, 48(1):10–16, January 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
48.1.10.

Stevanovi:2021:VPS

[SS21] Sanja Stevanovi and Dragan Stevanović. Visual proofs for the sums of fourth and fifth powers of the first n natural numbers. College Mathematics Journal, 52(2):115–120, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2021.1886773.

Schmitz:2022:ESM

[SS22] Michael Schmitz and André Streicher. Envelopes are solving machines for quadratics and cubics and certain polynomials of arbitrary degree. College Mathematics Journal, 53(5):372– 381, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2022.2133936.

Schiller:1989:MH

[SSG89] Barry Schiller, Helen Salzberg, and Richard K. Guy. Media highlights. College Mathematics Journal, 20(4):355–362, September 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1989.11973258.

Sanders:1998:MOF

[ST98] Marc D. Sanders and Barry A. Tesman. MATH and other four-letter words. College Mathematics Journal, 29(5):418–419, November 1998. CODEN ???? ISSN 0746-8342 (print), 1931-

1346 (electronic). URL http://www.tandfonline.com/doi/
abs/10.1080/07468342.1998.11973981.

Sauerberg:2002:C

[ST02] Jim Sauerberg and Alan Tarr. Centering. College Mathematics Journal, 33(1):24–31, January 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2002.
11921915.

Seidametova:2019:EIS

[ST19] Zarema Seidametova and Valerii Temnenko. Euler’s insignia: Some admirable curves having a simple trigonometric equation in a natural form. College Mathematics Journal, 50(2):134– 139, 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2019.1562825.

Seidametova:2020:SGO

[ST20] Zarema Seidametova and Valerii Temnenko. Some geometric objects related to a classical problem of Galileo. College Mathematics Journal, 51(1):57–65, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2020.1681871.

Stark:1971:CCC

[Sta71] William Stark. Calculus and the computer — CRICISAM. Two-Year College Mathematics Journal, 2(2):51–54, Autumn 1971. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1971.11974012.

Staib:1979:SPC

[Sta79] John Staib. Some pre-calculus algebra. Two-Year College Mathematics Journal, 10(2):89–95, March 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1979.
11972330.

Staton:1988:PRN

[Sta88] William Staton. On partitioning a real number. College Mathematics Journal, 19(1):53–54, January 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1988.
11973090.

Starr:1998:NCS

[Sta98] Norton Starr. Nothing counts for something. College Mathematics Journal, 29(4):308–309, September 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1998.
11973961.

Stein:1973:MCS

[Ste73] S. K. Stein. Mathematics for the captured student. Two-Year College Mathematics Journal, 4(3):62–71, Autumn 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1973.11974275.

Stein:1974:FID

[Ste74] S. K. Stein. Formal integration: Dangers and suggestions. Two-Year College Mathematics Journal, 5(2):1–7, Spring 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1974.11974305.

Stenger:1975:CIL

[Ste75] William Stenger. Cauchy’s Inequality and the least squares line. Two-Year College Mathematics Journal, 6(1):2–4, February 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1975.11974357.

Stein:1976:EGI

[Ste76a] Robert G. Stein. Exploring the Gaussian integers. Two-Year College Mathematics Journal, 7(4):4–10, December 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1976.11974454.

Sterrett:1976:STA

[Ste76b] Andrew Sterrett. A search for trends among mathematics programs in small colleges. Two-Year College Mathematics Journal, 7(1):21–23, February 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1976.11974415.

Stephens:1977:GSI

[Ste77] Larry J. Stephens. Getting the students involved in the elementary statistics course. Two-Year College Mathematics Journal, 8(1):19–21, January 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1977.11974474.

Stein:1980:CPP

[Ste80] F. Max Stein. The curve parallel to a parabola is not a parabola: Parallel curves. Two-Year College Mathematics Journal, 11(4): 239–246, September 1980. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1980.11972448.

Stein:1981:TCS

[Ste81a] Sherman K. Stein. Teachers, clocks, and students. Two-Year College Mathematics Journal, 12(3):195–198, June 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972523.

Stevens:1981:FBE

[Ste81b] Gary E. Stevens. Forward and backward with Euclid. Two-Year College Mathematics Journal, 12(5):302–306, November 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972555.

Steen:1983:JGK

[Ste83] Lynn Arthur Steen. John G. Kemeny: Computing pioneer. Two-Year College Mathematics Journal, 14(1):18–35, January 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1983.11972648.

Stein:1985:RP

[Ste85] Sherman Stein. Routine problems. College Mathematics Journal, 16(5):383–385, November 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1985.11972911.

Stenlund:2001:TLP

[Ste01] Mikko Stenlund. On the tangent lines of a parabola. College Mathematics Journal, 32(3):194–196, May 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2001.
11921875.

Stewart:2012:CD

[Ste12] Ian Stewart. Cups and downs. College Mathematics Journal, 43(1):15–19, January 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.43.1.015.

Stein:2015:WDE

[Ste15] Sherman Stein. What distributes over exponentiation? College Mathematics Journal, 46(1):11–14, January 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
46.1.11.

Stewart:2023:MSI

[Ste23] Seán M. Stewart. On that most over skinned of improper integrals. College Mathematics Journal, 54(2):123–129, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2023.2186085.

Stoutemyer:1983:CC

[Sto83] David R. Stoutemyer. Computers & calculators. Two-Year College Mathematics Journal, 14(3):233–239, June 1983.

CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1983.11972688.

Stockmeyer:2004:WDL

[Sto04] Paul K. Stockmeyer. What did Lincoln really mean? College Mathematics Journal, 35(2):103–104, March 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2004.
11922059.

Stockmeyer:2007:MMB

[Sto07] Paul K. Stockmeyer. More mathematics in the bedroom: a paradoxical probability. College Mathematics Journal, 38(5):339– 344, November 2007. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2007.11922258.

Stretton:1972:SPU

[Str72] William C. Stretton. Some problems of utmost gravity. Two-Year College Mathematics Journal, 3(1):72–75, Spring 1972. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1972.11974202.

Strum:1977:BM

[Str77] Jay E. Strum. Binomial matrices. Two-Year College Mathematics Journal, 8(5):260–266, November 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1977.
11974524.

Straffin:1979:UIE

[Str79] Philip D. Straffin, Jr. Using integrals to evaluate voting power. Two-Year College Mathematics Journal, 10(3):179–181, June 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1979.11972347.

Straffin:1985:TPW

[Str85] Philip D. Straffin, Jr. Three person winner-take-all games with McCarthy’s Revenge Rule. College Mathematics Journal, 16 (5):386–394, November 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1985.11972912.

Strang:1990:SDV

[Str90a] Gilbert Strang. Sums and differences vs. integrals and derivatives. College Mathematics Journal, 21(1):20–27, January 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1990.11973278.

Struss:1990:EVS

[Str90b] Keith A. Struss. Exploring the volume-surface area relationship. College Mathematics Journal, 21(1):40–43, January 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1990.11973282.

Strang:1991:CS

[Str91] Gilbert Strang. A chaotic search for i. College Mathematics Journal, 22(1):3–12, January 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1991.11973353.

Strang:1993:GMS

[Str93] Gilbert Strang. Graphs, matrices, and subspaces. College Mathematics Journal, 24(1):20–28, January 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1993.
11973502.

Straffin:1997:MH

[Str97a] Philip D. Straffin. Media highlights. College Mathematics Journal, 28(4):322–329, September 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1997.
11973885.

Struthers:1997:CWV

[Str97b] Allan A. Struthers. Counterexamples to a weakened version of the two-variable second derivative test. College Mathematics Journal, 28(5):383–385, November 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1997.11973896.

Strohmer:2004:FOD

[Str04] Gerhard Ströhmer. First order differential equations and the atmosphere. College Mathematics Journal, 35(2):93–96, March 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2004.11922057.

Stump:1995:NDP

[Stu95] Robert K. Stump. A normal density project. College Mathematics Journal, 26(4):310–312, September 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1995.
11973720.

Stuffelbeam:2013:HWW

[Stu13] Ryan Stuffelbeam. How weird are weird fractions? College Mathematics Journal, 44(3):202–209, May 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
44.3.202.

Sullivan:2017:ILC

[STW17] Brendan W. Sullivan, Nikolas Townsend, and Mikayla L. Werzanski. An introduction to lazy Cops and robbers on graphs. College Mathematics Journal, 48(5):322–333, November 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.48.5.322.

Subramanian:2004:SDL

[Sub04] P. K. Subramanian. Successive differentiation and Leibniz’s Theorem. College Mathematics Journal, 35(4):274–282, September 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2004.11922085.

Subramaniam:2015:PWC

[Sub15] K. B. Subramaniam. Proof without words: Cotangent double angle identity. College Mathematics Journal, 46(2):121, March 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.46.2.121.

Suen:1994:RCD

[Sue94] Robert Y. Suen. Roots of cubics via determinants. College Mathematics Journal, 25(2):115–117, March 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1994.
11973592.

Sullivan:1996:KAS

[Sul96] Michael C. Sullivan. Knots about Stokes’ Theorem. College Mathematics Journal, 27(2):119–122, March 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1996.
11973762.

Sultan:2009:CHH

[Sul09] Alan Sultan. CORDIC: How hand calculators calculate. College Mathematics Journal, 40(2):87–92, March 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/07468342.
2009.11922342.

Sullivan:2016:FCD

[Sul16] Patrick Sullivan. The FA cup draw and pairing up probabilities. College Mathematics Journal, 47(4):282–292, September 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.47.4.282.

Sullivan:2024:CPS

[Sul24] Brendan Sullivan. Crossword puzzle solution. College Mathematics Journal, 55(4):371, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2024.2359293.

Sun:2021:PWC

[Sun21] Dennis L. Sun. Proof without words: Convex hulls and Jensen’s inequality. College Mathematics Journal, 52(4):298, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2021.1959785.

Suplee:1984:ZRC

[Sup84] Curt Suplee. Zork, RAMS and the curse of Ra: Computo, ergo sum. College Mathematics Journal, 15(2):158–159, March 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972768.

Sury:2011:UGI

[Sur11] B. Sury. Uncountably generated ideals of functions. College Mathematics Journal, 42(5):404–406, November 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
42.5.404.

Suzuki:1999:MDP

[Suz99] Jeff Suzuki. Multiplying and dividing polynomials using geloxia. College Mathematics Journal, 30(1):50–53, January 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1999.11974031.

Suzuki:2014:SLP

[Suz14] Jeff Suzuki. The self-limiting partisan gerrymander: an optimization approach. College Mathematics Journal, 45(5):338– 348, November 2014. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.45.5.338.

Steger:1976:MLA

[SW76] William F. Steger and Gretchen Willging. Modularizing liberal arts mathematics: an experiment. Two-Year College Mathematics Journal, 7(3):33–37, September 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1976.
11974444.

Schultz:1979:CC

[SW79] James E. Schultz and Bert K. Waits. Computers & calculators. Two-Year College Mathematics Journal, 10(1):20–27, January 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1979.11972317.

Staton:1995:ISG

[SW95] William Staton and Clifton Wingard. Independent sets and the Golden ratio. College Mathematics Journal, 26(4):292–296, September 1995. CODEN ???? ISSN 0746-8342 (print), 1931-

1346 (electronic). URL http://www.tandfonline.com/doi/
abs/10.1080/07468342.1995.11973715.

Swadener:1973:TM

[Swa73] Marc Swadener. On transformations and matrices. Two-Year College Mathematics Journal, 4(3):44–51, Autumn 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1973.11974272.

Swadener:1974:FFF

[Swa74] Marc Swadener. A finite field — a finite geometry and triangles. Two-Year College Mathematics Journal, 5(3):22–25, September 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1974.11974331.

Swanson:2005:PAC

[Swa05] Christopher N. Swanson. The probability an amazing card trick is dull. College Mathematics Journal, 36(3):209–212, May 2005. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2005.11922130.

Swanson:2011:BAC

[Swa11] David Swanson. The band around a convex body. College Mathematics Journal, 42(1):15–24, January 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
42.1.015.

Swain:2013:AC

[Swa13] Gordon A. Swain. Archimedes curves. College Mathematics Journal, 44(3):185–189, May 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.44.3.185.

Swetz:1992:MH

[Swe92] Frank Swetz. Media highlights. College Mathematics Journal, 23(4):348–355, September 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1992.11973484.

Swetz:1995:MH

[Swe95] Frank Swetz. Media highlights. College Mathematics Journal, 26(5):411–418, November 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1995.11973742.

Swetz:2002:BR

[Swe02] Frank Swetz. Book review. College Mathematics Journal, 33(1): 69–70, January 2002. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2002.11921921.

Swetz:2003:BR

[Swe03] Frank Swetz. Book review. College Mathematics Journal, 34(5): 423, November 2003. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2003.11922042.

Swenson:2022:SRA

[Swe22] Dan Swenson. A statistics-related approach to the legislative apportionment problem … and also, a second one. College Mathematics Journal, 53(5):382–391, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2022.
2118994.

Sodano:1978:NPF

[SY78] John Sodano and Arthur Yaspan. A neglected probability formula. Two-Year College Mathematics Journal, 9(3):145–147, June 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1978.11974565. See comments [Jan80].

Schwartz:1987:GSI

[SY87] Judah L. Schwartz and Michal Yerushalmy. The geometric supposer: an intellectual prosthesis for making conjectures. College Mathematics Journal, 18(1):58–65, January 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1987.11973012.

Swenson:1993:ADN

[SY93] Carl Swenson and André Yandl. An alternative definition of the number e. College Mathematics Journal, 24(5):458–461, November 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1993.11973569.

Swenson:2012:PPT

[SY12] Carl Swenson and André Yandl. Partitioning Pythagorean triangles using Pythagorean angles. College Mathematics Journal, 43(3):220–225, May 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.43.3.220.

Schlais:1989:SR

[SZ89] Hal Schlais and Walter Zimmermann. Software reviews. College Mathematics Journal, 20(2):161–162, March 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1989.
11973226.

Szabo:2005:PPG

[Sza05] Sándor Szabó. A paper-and-pencil gcd algorithm for Gaussian integers. College Mathematics Journal, 36(5):374–380, November 2005. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2005.11922152.

Topper:1974:WMI

[TA74] Michael A. Topper and James W. Armstrong. A working model for inservice training. Two-Year College Mathematics Journal, 5(4):16–18, December 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1974.11974345.

Tall:1984:CMD

[Tal84] David Tall. Continuous mathematics and discrete computing are complementary, not alternatives …. College Mathematics Journal, 15(5):389–391, November 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1984.
11972814.

Tamaki:1987:ETP

[Tam87] Robert K. Tamaki. Equalizing a two-person alternation game. College Mathematics Journal, 18(2):134–135, March 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1987.11973021.

Tanis:1985:SR

[Tan85] Elliot Tanis. Software reviews. College Mathematics Journal, 16(1):72–73, January 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1985.11972857.

Tan:1986:GIS

[Tan86a] A. Tan. A geometric interpretation of Simpson’s Paradox. College Mathematics Journal, 17(4):340–341, September 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1986.11972977.

Tan:1986:CTI

[Tan86b] Soo Tang Tan. A case of true interest. College Mathematics Journal, 17(3):247–248, May 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1986.11972966.

Tanur:1988:NWS

[Tan88] Judith Tanur. No! But who should teach statistics? …. College Mathematics Journal, 19(1):11–12, January 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1988.
11973075.

Tandra:2015:IPC

[Tan15] Haryono Tandra. An inductive proof of the compactness of the closed unit ball of an arbitrary dimension. College Mathematics Journal, 46(3):218–219, May 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
46.3.218.

Tandra:2017:BTL

[Tan17] Haryono Tandra. Basic theorems in the language of maximal intervals. College Mathematics Journal, 49(1):41–45, January 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2017.1392101.

Tattersall:1991:MH

[Tat91] J. J. Tattersall. Media highlights. College Mathematics Journal, 22(4):349–356, September 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1991.11973412.

Tattersall:1994:MH

[Tat94] Jim Tattersall. Media highlights. College Mathematics Journal, 25(4):341–348, September 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1994.11973637.

Tattersall:2002:IDS

[Tat02] James J. Tattersall. An interview with Dirk Struik on the eve of his one hundredth birthday. College Mathematics Journal, 33 (3):178–187, May 2002. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2002.11921938.

Turner:1991:SR

[TB91] Peter Turner and Craig Bailey. Software review. College Mathematics Journal, 22(5):436–442, November 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1991.
11973424.

Teixeira:2020:FCM

[TB20] João Teixeira and Maria João Borges. Formula for the computation of the matrix exponential. College Mathematics Journal, 51(5):345–350, 2020. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2020.1820285.

Thomson:1977:LE

[TD77] Thomas R. Thomson and Lester Lange Dean. Letters to the Editor. Two-Year College Mathematics Journal, 8(3):131, June 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1977.11974495.

Teague:2006:GTF

[Tea06] Daniel J. Teague. Group testing: Four student solutions to a classic optimization problem. College Mathematics Journal, 37 (4):261–268, September 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2006.11922193.

Teets:1993:LSQ

[Tee93] Donald Teets. Least squares and quadric surfaces. College Mathematics Journal, 24(3):243–244, May 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1993.
11973538.

Teets:2000:FFF

[Tee00] Donald A. Teets. Fallacies, flaws, and flimflam. College Mathematics Journal, 31(4):307–309, September 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2000.
11974165.

Teets:2007:MGT

[Tee07] Donald Teets. The mathematics of “go to” telescopes. College Mathematics Journal, 38(3):170–178, May 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/07468342.
2007.11922233.

Teets:2012:PTM

[Tee12] Donald Teets. Push-to telescope mathematics. College Mathematics Journal, 43(3):227–231, May 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
43.3.227.

Teets:2017:CSO

[Tee17] Donald Teets. Computing a satellite orbit from photographs. College Mathematics Journal, 48(2):102–110, March 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
48.2.102.

Teixeira:2023:IIA

[Tei23] Katiuscia Teixeira. Improper integrals: an alternative criterion. College Mathematics Journal, 54(3):232–234, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2023.2201566.

Tennor:1983:AP

[Ten83] Michael Tennor. The address problem. Two-Year College Mathematics Journal, 14(5):407–415, November 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1983.
11972728.

Terrell:1996:IPA

[Ter96] Robert E. Terrell. Independence of path and all that. College Mathematics Journal, 27(4):272–277, September 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1996.11973791.

Tharp:2001:SR

[Tha01] Marcia Tharp. Software reviews. College Mathematics Journal, 32(3):218–222, May 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2001.11921879.

Thamattoor:2018:NLB

[Tha18] Ajoy Thamattoor. Normal limit of the binomial via the discrete derivative. College Mathematics Journal, 49(3):216–217, May 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2018.1440872.

Theerakarn:2023:CSA

[The23] Thunwa Theerakarn. On the center of surface area of the boundary of a star-shaped region. College Mathematics Journal, 54 (4):326–336, 2023. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2023.2240203.

Thibodeaux:1979:LSa

[Thi79a] Michael J. Thibodeaux. The lighter side. Two-Year College Mathematics Journal, 10(1):28–29, January 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1979.
11972318.

Thibodeaux:1979:LSb

[Thi79b] Michael J. Thibodeaux. The lighter side. Two-Year College Mathematics Journal, 10(2):107, March 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1979.
11972333.

Thibodeaux:1979:LSc

[Thi79c] Michael J. Thibodeaux. The lighter side. Two-Year College Mathematics Journal, 10(3):193–194, June 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1979.
11972350.

Thibodeaux:1979:LSd

[Thi79d] Michael J. Thibodeaux. The lighter side. Two-Year College Mathematics Journal, 10(4):266–269, September 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1979.11972363.

Thibodeaux:1979:LSe

[Thi79e] Michael J. Thibodeaux. The lighter side. Two-Year College Mathematics Journal, 10(5):343–371, November 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1979.11972381.

Thibodeaux:1980:LSa

[Thi80a] Michael J. Thibodeaux. The lighter side. Two-Year College Mathematics Journal, 11(1):48–49, January 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1980.
11972399.

Thibodeaux:1980:LSb

[Thi80b] Michael J. Thibodeaux. The lighter side. Two-Year College Mathematics Journal, 11(2):114–115, March 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1980.
11972417.

Thibodeaux:1980:LSc

[Thi80c] Michael J. Thibodeaux. The lighter side. Two-Year College Mathematics Journal, 11(3):216–217, June 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1980.
11972443.

Thibodeaux:1980:LSd

[Thi80d] Michael J. Thibodeaux. The lighter side. Two-Year College Mathematics Journal, 11(4):267, September 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1980.
11972454.

Tossavainen:2024:WCD

[THMM24] Timo Tossavainen, Pentti Haukkanen, Jorma K. Merikoski, and Mika Mattila. We can differentiate numbers, too. College Mathematics Journal, 55(2):100–108, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2023.
2268494.

Thompson:1984:ICV

[Tho84] Patrick W. Thompson. Issues of content versus method …. College Mathematics Journal, 15(5):394–395, November 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972817.

Thoo:1992:TE

[Tho92] J. Thoo. Timing is everything. College Mathematics Journal, 23(4):308–309, September 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1992.11973475.

Thornhill:1996:SR

[Tho96] Jerry Thornhill. Software reviews. College Mathematics Journal, 27(4):305–311, September 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1996.11973800.

Thompson:1998:DBC

[Tho98a] Richard B. Thompson. Designing a baseball cover. College Mathematics Journal, 29(1):48–61, January 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1998.
11973916.

Thoo:1998:PWT

[Tho98b] J. B. Thoo. A picture is worth a thousand words. College Mathematics Journal, 29(5):408–411, November 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1998.11973977.

Thorsen:2000:EIM

[Tho00] Bobette Hayden Thorsen. The eigenvalues of an infinite matrix. College Mathematics Journal, 31(2):107–110, March 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2000.11974123.

Thomson:2012:SKR

[Tho12] Brian S. Thomson. A strong kind of Riemann integrability. College Mathematics Journal, 43(4):313–321, September 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.43.4.313.

Thomson:2023:NTR

[Tho23] Brian S. Thomson. A note on teaching the Riemann integral. College Mathematics Journal, 54(1):3–10, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2022.
2160588.

Thuente:1982:CLP

[Thu82] David J. Thuente. Chain letters: a poor investment unless …. Two-Year College Mathematics Journal, 13(1):28–35, January 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1982.11972575.

Thurston:2000:CWI

[Thu00] Hugh Thurston. Can we improve the teaching of calculus? College Mathematics Journal, 31(4):262–267, September 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2000.11974156.

Thurschwell:2015:ESP

[Thu15] Eric Thurschwell. An even simpler proof of the right-hand rule. College Mathematics Journal, 46(3):215–217, May 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
46.3.215.

Tirman:1986:GPP

[Tir86] Alvin Tirman. Geometric parametrization of Pythagorean triples. College Mathematics Journal, 17(2):168, March 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1986.11972950.

Thompson:2000:GNP

[TL00] Peter Thompson and Lorrie Lendvoy. Getting normal probability approximations without using normal tables. College Mathematics Journal, 31(1):51–54, January 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2000.
11974109.

Leise:2023:MH

[TLSES23] In Memoriam) Tanya Leise (Section Editor, 2019-2023 and Philip Straffin. Media highlights. College Mathematics Journal, 54 (3):243–250, 2023. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2023.2207435.

Tattersall:2001:IDM

[TM01] James Tattersall and Shawnee McMurran. An interview with Dame Mary L. Cartwright, D.B.E., F.R.S. College Mathematics Journal, 32(4):242–254, September 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2001.
11921884.

Tolosa:2022:TEP

[Tol22] Juan Tolosa. Two eigenvectors for the price of one. College Mathematics Journal, 53(3):227–229, 2022. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2022.
2040265.

Tong:1997:APV

[Ton97] Jingcheng Tong. Area and perimeter, volume and surface area. College Mathematics Journal, 28(1):57, January 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1997.11973833.

Tong:2015:NFT

[Ton15] Zengxiang Tong. A note on the fundamental theorem of calculus. College Mathematics Journal, 46(5):367–368, November 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.46.5.367.

Torrence:2011:ELG

[Tor11] Bruce Torrence. The easiest lights out games. College Mathematics Journal, 42(5):361–372, November 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
42.5.361.

Touhey:1995:CTG

[Tou95] Pat Touhey. Chebyshev’s Theorem: a geometric approach. College Mathematics Journal, 26(2):139–141, March 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1995.11973684.

Touhey:1996:CBO

[Tou96] Pat Touhey. Countability via bases other than 10. College Mathematics Journal, 27(5):382–384, November 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1996.11973813.

Tou:2013:TFG

[Tou13] Erik R. Tou. A typology for finite groups. College Mathematics Journal, 44(5):432–436, November 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
44.5.432.

Toxopeus:1970:SDM

[Tox70] Fred Toxopeus. Summer Developmental Mathematics Program at Kalamazoo Valley Community College. Two-Year College Mathematics Journal, 1(1):14–16, Spring 1970. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1970.
11973160.

Trivieri:1974:LE

[TP74] Lawrence Trivieri and Bill Peters. Letters to the Editor. Two-Year College Mathematics Journal, 5(4):1–3, December 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1974.11974342.

Tabakovic:1999:MCP

[TPB99] Haris Tabakovic, Joseph Paullet, and Richard Bertram. Measuring the curl of paper. College Mathematics Journal, 30(4):315– 317, September 1999. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1999.11974079.

Treeby:2023:CPT

[TR23] David Treeby and Marty Ross. A curious property of tangents. College Mathematics Journal, 54(1):11–14, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2022.
2160602.

Trachtenberg:1996:GDM

[Tra96] Felicia Moss Trachtenberg. The game of dreidel made fair. College Mathematics Journal, 27(4):278–281, September 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1996.11973792.

Treisman:1992:SSS

[Tre92] Uri Treisman. Studying students studying calculus: a look at the lives of minority mathematics students in College. College Mathematics Journal, 23(5):362–372, November 1992. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1992.11973486.

Treuden:1994:FDF

[Tre94] Michael L. Treuden. Frequencies of digits in factorials: an experimental approach. College Mathematics Journal, 25(1):48–55, January 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1994.11973583.

Trevino:2018:IEP

[Tre18] Enrique Treviño. An inclusion-exclusion proof of Wilson’s Theorem. College Mathematics Journal, 49(5):367–368, November 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2018.1512325.

Trevino:2019:BSP

[Tre19] Enrique Treviño. A birthday in St. Petersburg. College Mathematics Journal, 50(1):36–40, January 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2019.
1534492.

Treeby:2020:RGI

[Tre20] David Treeby. Randomly generated identities. College Mathematics Journal, 51(2):90–94, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2020.1692618.

Troutman:1978:CC

[Tro78] James G. Troutman. Computer corner. Two-Year College Mathematics Journal, 9(2):101–103, March 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1978.
11974556.

Tropp:1982:BR

[Tro82] Henry S. Tropp. Book reviews. Two-Year College Mathematics Journal, 13(1):72–73, January 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1982.
11972585.

Trotter:1989:LPP

[Tro89] William T. Trotter. On-line partitioning of partially ordered sets. College Mathematics Journal, 20(2):124–131, March 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1989.11973218.

Trombetta:2004:TLI

[Tro04] Maurizio Trombetta. Tangent lines and the inverse function differentiation rule. College Mathematics Journal, 35(4):258– 261, September 2004. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2004.11922083.

Trudeau:1983:HBP

[Tru83] Richard J. Trudeau. How big is a point? Two-Year College Mathematics Journal, 14(4):295–300, September 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1983.11972701.

Truc:2019:RSG

[Tru19] Jean-Paul Truc. Riemann sums for generalized integrals. College Mathematics Journal, 50(2):123–132, 2019. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2019.
1560119.

Truc:2023:HMK

[Tru23] Jean-Paul Truc. How to marry the kernel and the image. College Mathematics Journal, 54(4):386–387, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2023.
2225396.

Tucker:1983:MTS

[TS83] A. W. Tucker and Philip D. Straffin, Jr. The mathematics of Tucker: a sampler. Two-Year College Mathematics Journal, 14 (3):228–232, June 1983. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1983.11972687.

Trenkler:2003:SR

[TT03] Dietrich Trenkler and Götz Trenkler. On the square root of aa^{T} + bb^{T}. College Mathematics Journal, 34(1):39–41, January 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2003.11921983.

Trenkler:2012:NRL

[TT12] Götz Trenkler and Dietrich Trenkler. The numerical range of the Luoshu is a piece of cake — almost. College Mathematics Journal, 43(5):371–376, November 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.43.5.371.

Tian:2021:URF

[TT21] Melanie Tian and Enrique Treviño. An unusual recursive formula to answer a question regarding fixed points in permutations. College Mathematics Journal, 52(3):219–220, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2021.1909979.

Taalman:2013:MM

[TTW^{+}13] L. Taalman, A. Tongen, B. Warren, F. Wyrick-Flax, and I. Yoon. Mancala matrices. College Mathematics Journal, 44(4):273– 283, September 2013. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.44.4.273.

Tucker:1983:SLR

[Tuc83] Albert W. Tucker. Solomon Lefschetz: a reminiscence. Two-Year College Mathematics Journal, 14(3):225–227, June 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1983.11972686.

Tucker:1993:GIL

[Tuc93] Alan Tucker. The growing importance of linear algebra in undergraduate mathematics. College Mathematics Journal, 24(1):3–9, January 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1993.11973500.

Turner:1981:PLP

[Tur81] Barbara Turner. Point-and-line proof for the sum of the cubes. Two-Year College Mathematics Journal, 12(4):270–271, September 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972544.

Turner:1985:FUW

[Tur85a] Nura D. Turner. A follow-up on the USAMO winners. College Mathematics Journal, 16(5):366–368, November 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1985.11972909.

Turner:1985:HSO

[Tur85b] Nura D. Turner. A historical sketch of Olympiads: U.S.A. and international. College Mathematics Journal, 16(5):330–335, November 1985. CODEN ???? ISSN 0746-8342 (print), 1931-

1346 (electronic). URL http://www.tandfonline.com/doi/
abs/10.1080/07468342.1985.11972906.

Thoro:1994:IRR

[TV94] Dmitri Thoro and Linda Valdés. Investigation of a recurrence relation. College Mathematics Journal, 25(4):322–324, September 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1994.11973631.

Thomassen:1990:UFD

[TW90] Carsten Thomassen and Herbert S. Wilf. On ‘uniformly filled’ determinants. College Mathematics Journal, 21(2):135–137, March 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1990.11973298.

Teets:1998:CPO

[TW98] Donald A. Teets and Karen Whitehead. Computation of planetary orbits. College Mathematics Journal, 29(5):397–404, November 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1998.11973975.

Tymoczko:1981:CUC

[Tym81] Thomas Tymoczko. Computer use to computer proof: a rational reconstruction. Two-Year College Mathematics Journal, 12(2):120–125, March 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1981.11972509.

Turner:1988:MPD

[TYM88] Danny W. Turner, Dean M. Young, and Virgil R. Marco. Maybe the price doesn’t have to be right: Analysis of a popular TV game show. College Mathematics Journal, 19(5):419–421, November 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1988.11973149.

Uetrecht:1988:CSR

[Uet88] James A. Uetrecht. A clamped Simpson’s Rule. College Mathematics Journal, 19(1):43–52, January 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1988.
11973089.

Urdaletova:1996:SLD

[UK96] A. B. Urdaletova and S. K. Kydyraliev. Solving linear differential equations by operator factorization. College Mathematics Journal, 27(3):199–203, May 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1996.11973776.

Unal:2009:PWS

[Una09] Hasan Unal. Proof without words: Sum of an infinite series. College Mathematics Journal, 40(1):39, January 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2009.11922334.

Unal:2010:PWD

[Una10] Hasan Unal. Proof without words: Double sum for sine and cosine. College Mathematics Journal, 41(5):392, November 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2010.11922454.

Unal:2014:PWF

[Una14a] Hasan Unal. Proof without words: a formula of Hermann. College Mathematics Journal, 45(5):387, November 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.45.5.387.

Unal:2014:PWA

[Una14b] Hasan Unal. Proof without words: an arctangent identity. College Mathematics Journal, 45(5):357, November 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
45.5.357.

Underhill:1984:FBD

[Und84] W. Vance Underhill. Finding bounds for definite integrals. College Mathematics Journal, 15(5):426–429, November 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972827.

Usiskin:1979:PS

[Usi79] Zalman Usiskin. Products of sines. Two-Year College Mathematics Journal, 10(5):334–340, November 1979. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1979.
11972379.

Usiskin:1986:PPR

[Usi86] Zalman P. Usiskin. A pretrigonometry proof of the reflection property of the ellipse. College Mathematics Journal, 17(5): 418, November 1986. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1986.11972992.

Vallin:2015:MTL

[Val15] Robert W. Vallin. A magic trick leads to an identity: Some induction fun. College Mathematics Journal, 46(4):295–298, September 2015. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.46.4.295.

Vandervelde:2014:ACD

[Van14] Sam Vandervelde. An acute case of discontinuity. College Mathematics Journal, 45(1):22–28, January 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
45.1.022.

Varner:1976:CBU

[Var76] John T. Varner III. Comparing a^{b} and b^{a} using elementary calculus. Two-Year College Mathematics Journal, 7(4):46, December 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1976.11974465.

Varner:1977:DCT

[Var77] John Taylor Varner III. Discovering a calculus theorem. Two-Year College Mathematics Journal, 8(5):304, November 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1977.11974533.

Vardanyan:2011:TTC

[Var11] Vagarshak Vardanyan. Teaching tip: Correcting Cramer’s Rule. College Mathematics Journal, 42(1):54–55, January 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2011.11922474.

Vautaw:2008:TPT

[Vau08] William R. Vautaw. Two problems with table saws. College Mathematics Journal, 39(2):121–128, March 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2008.
11922285.

Venit:1986:CRR

[VB86] Stewart Venit and Wayne Bishop. Classifying row-reduced echelon matrices. College Mathematics Journal, 17(2):169–170, March 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1986.11972951.

Vierling-Claassen:2010:WDH

[VC10] Angela Vierling-Claassen. Who does the housework? College Mathematics Journal, 41(4):298–302, September 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
4169/074683410X510290.

vanDeventer:2009:PMP

[vD09] M. Oskar van Deventer. Puzzling mechanisms, part 4: Tricky triplet. College Mathematics Journal, 40(3):211, May 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2009.11922362.

Velleman:2023:ISC

[Vel23] Daniel J. Velleman. Integration by substitution: a case study in the relationship between calculus and analysis. College Mathematics Journal, 54(4):367–373, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2023.
2228673.

Venit:1974:IOI

[Ven74] Stewart Venit. Interchanging the order of integration. Two-Year College Mathematics Journal, 5(3):20–21, September 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1974.11974330.

Venit:1976:RCD

[Ven76] Stewart M. Venit. Remarks concerning the delta method for approximating roots. Two-Year College Mathematics Journal, 7 (4):1–3, December 1976. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1976.11974453.

Venit:1980:CSD

[Ven80] Stewart Venit. Conic section or degenerate form — a simple test. Two-Year College Mathematics Journal, 11(5):316–319, November 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1980.11972469.

Vest:1985:GCR

[Ves85] Floyd Vest. Graphing the complex roots of a quadratic equation. College Mathematics Journal, 16(4):257–261, September 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1985.11972891.

Viertel:1978:IPA

[Vie78] W. K. Viertel. An interesting practical application of solid analytic geometry. Two-Year College Mathematics Journal, 9(5): 273–275, November 1978. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1978.11974591.

Venit:1980:AGA

[VK80] Stewart Venit and Richard Katz. An analytic geometry approach to the least squares line of best fit. Two-Year College Mathematics Journal, 11(4):270–272, September 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1980.11972456.

Venit:2000:EML

[VK00] Stewart Venit and Richard Katz. Eigenvalues of matrices of low rank. College Mathematics Journal, 31(3):208–210, May 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2000.11974145.

Vodola:1997:AGP

[Vod97] Paul Vodola. Adventure games, permutations, and spreadsheets. College Mathematics Journal, 28(4):301–309, September 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1997.11973882.

vonRonik:1997:DS

[vR97] Wolf von Rönik. Doughnut slicing. College Mathematics Journal, 28(5):381–383, November 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1997.11973895.

vonRonik:2001:PBU

[vR01] Wolf von Rönik. The profit in being unbalanced. College Mathematics Journal, 32(5):348–351, November 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2001.
11921902.

Verdejo:2021:SSM

[VU21] Antonio J. Moreno Verdejo, Abraham López Viveros and Rafael Ramirez Uclés. Statistical significance of the median of a set of points on the plane. College Mathematics Journal, 52(3):205–218, 2021. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2021.1909969.

Villarino:2024:AR

[VV24] Mark B. Villarino and Joseph C. Várilly. Archimedes’ revenge. College Mathematics Journal, 55(3):257–259, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2023.2285686.

VanDyke:1973:R

[VW73] James A. Van Dyke and Donald C. Willner. Reviews. Two-Year College Mathematics Journal, 4(3):72–78, Autumn 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1973.11974276.

Wagner:1977:NIT

[Wag77] Carl G. Wagner. Newton’s Inequality and a test for imaginary roots. Two-Year College Mathematics Journal, 8(3):145–147, June 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1977.11974498.

Wagner:1983:USS

[Wag83] Clifford H. Wagner. Uncertainty in science and statistics. Two-Year College Mathematics Journal, 14(4):360–363, September 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1983.11972717.

Wagner:1984:CVD

[Wag84] Daniel H. Wagner. Calculus versus discrete mathematics in OR applications …. College Mathematics Journal, 15(5):374– 375, November 1984. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1984.11972805.

Wagner:1987:EVG

[Wag87] Clifford H. Wagner. Enhancing the value of graphics programs. College Mathematics Journal, 18(2):142–152, March 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1987.11973026.

Wagon:1994:MMT

[Wag94] Stan Wagon. A mathematical magic trick. College Mathematics Journal, 25(4):325–326, September 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1994.
11973633.

Wagner:1999:CAJ

[Wag99] Clifford H. Wagner. Coming to America: The journey of an immigrant scholar. College Mathematics Journal, 30(1):2–17, January 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1999.11974023.

Wagon:2011:BRR

[Wag11] Stan Wagon. Book review: Roads to Infinity, The Mathematics of Truth and Proof, by John Stillwell, A K Peters. College Mathematics Journal, 42(2):160–162, March 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.42.2.160.

Wagon:2014:GTP

[Wag14] Stan Wagon. Graph theory problems from hexagonal and traditional chess. College Mathematics Journal, 45(4):278–287, September 2014. CODEN ???? ISSN 0746-8342 (print), 1931-

1346 (electronic). URL http://www.tandfonline.com/doi/
abs/10.4169/college.math.j.45.4.278.

Wagon:2020:CCS

[Wag20] Stan Wagon. Counting connected sets of squares. College Mathematics Journal, 51(3):173, 2020. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2020.1735878.

Waits:1974:IVM

[Wai74] Bert K. Waits. Instructional videocassettes in mathematics. Two-Year College Mathematics Journal, 5(4):27–30, December 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1974.11974348.

Waits:1978:CAI

[Wai78] Bert K. Waits. Classroom applications of the inexpensive handheld calculator. Two-Year College Mathematics Journal, 9(3): 162–166, June 1978. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1978.11974569.

Wallace:1986:CD

[Wal86] Kyle D. Wallace. Cyclic diversions. College Mathematics Journal, 17(1):61–65, January 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1986.11972930.

Wallis:1987:OFG

[Wal87] W. D. Wallis. One-factorizations of graphs: Tournament applications. College Mathematics Journal, 18(2):116–123, March 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1987.11973019.

Wallen:1995:KTM

[Wal95a] Lawrence J. Wallen. Kepler, the taxicab metric, and beyond: an isoperimetric primer. College Mathematics Journal, 26(3):178– 190, May 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1995.11973693.

Walsh:1995:DNM

[Wal95b] James A. Walsh. The dynamics of Newton’s Method for cubic polynomials. College Mathematics Journal, 26(1):22–28, January 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1995.11973661.

Walsh:1996:FLA

[Wal96] James A. Walsh. Fractals in linear algebra. College Mathematics Journal, 27(4):298–304, September 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1996.
11973799.

Walser:2000:PP

[Wal00a] Hans Walser. The Pascal pyramid. College Mathematics Journal, 31(5):383–392, November 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2000.11974178.

Walser:2000:PC

[Wal00b] Hans Walser. The pop-up cuboctahedron. College Mathematics Journal, 31(2):89–92, March 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2000.11974119.

Walk:2004:MY

[Wal04a] Stephen M. Walk. Mind your ∀’s and ∃’s. College Mathematics Journal, 35(5):362–369, November 2004. CODEN ????

ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2004.
11922097.

Walker:2004:ERL

[Wal04b] James S. Walker. An elementary resolution of the liar paradox. College Mathematics Journal, 35(2):105–111, March 2004. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2004.11922060.

Walk:2011:IVT

[Wal11] Stephen M. Walk. The Intermediate Value Theorem is NOT obvious — and I am going to prove it to you. College Mathematics Journal, 42(4):254–259, September 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
42.4.254.

Wang:2010:ALF

[Wan10] Frank Wang. Application of the Lambert W function to the SIR epidemic model. College Mathematics Journal, 41(2):156–159, March 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/074683410X480276.

Wang:2014:PWC

[Wan14a] Long Wang. Proof without words: Cosine difference identity. College Mathematics Journal, 45(5):370, November 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
45.5.370.

Wang:2014:PWS

[Wan14b] Long Wang. Proof without words: Sine sum identity. College Mathematics Journal, 45(3):190, May 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
45.3.190.

Wang:2021:EEA

[Wan21] Wenjie Wang. The explicit expression of axis and angle of a rotation matrix. College Mathematics Journal, 52(1):39– 44, 2021. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2021.1845540.

Watkins:1984:MH

[Wat84] Ann Watkins. Media highlights. College Mathematics Journal, 15(4):354–361, September 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1984.11972800.

Watkins:1985:VLL

[Wat85] Will Watkins. Vectors in a LOGO learning environment. College Mathematics Journal, 16(4):286–300, September 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1985.11972901.

Watkins:1987:GF

[Wat87] William Watkins. Generating functions. College Mathematics Journal, 18(3):195–211, May 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1987.
11973033.

Waterhouse:1989:TEP

[Wat89] William C. Waterhouse. Two elementary proofs of an inequality (and 11/2 better ones). College Mathematics Journal, 20(3):201– 205, May 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1989.11973231.

Watkins:2012:TNG

[Wat12] John Watkins. Triangular numbers, Gaussian integers, and KenKen. College Mathematics Journal, 43(1):37–42, January 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.43.1.037.

Wayment:1980:AQI

[Way80] Stanley G. Wayment. Another question of interest. Two-Year College Mathematics Journal, 11(4):252–254, September 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1980.11972451.

Whitman:1977:OIA

[WC77] Betsey Whitman and Donald Cook. Operational and intuitive algebra. Two-Year College Mathematics Journal, 8(3):155–161, June 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1977.11974501.

Waiveris:1999:FS

[WC99] Charles Waiveris and Yuanqian Chen. Folding stars. College Mathematics Journal, 30(5):370–378, November 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1999.11974088.

Wiener:1988:IPI

[WCP88] Joseph Wiener, Henjin Chi, and Hushang Poorkarimi. Involutions and problems involving perimeter and area. College Mathematics Journal, 19(3):250–252, May 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1988.
11973122.

Waits:1990:HCN

[WD90] Bert K. Waits and Franklin Demana. A hidden case of negative amortization. College Mathematics Journal, 21(2):121–126, March 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1990.11973294.

Waits:1973:SEM

[WE73] Bert K. Waits and Larry C. Elbrink. Student evaluation of mathematics instruction. Two-Year College Mathematics Journal, 4(2):59–66, Spring 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1973.11974253.

Weimer:1976:CCN

[Wei76a] Richard C. Weimer. Can the complex numbers be ordered? Two-Year College Mathematics Journal, 7(4):10–12, December 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1976.11974455.

Weissglass:1976:SGA

[Wei76b] Julian Weissglass. Small groups: an alternative to the lecture method. Two-Year College Mathematics Journal, 7(1):15–20, February 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1976.11974414.

Weiss:1991:SR

[Wei91] Joan Wyzkoski Weiss. Software reviews. College Mathematics Journal, 22(3):248–254, May 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1991.
11973393.

Wejntrob:1983:ASR

[Wej83] Leon Wejntrob. Approximation of square roots. Two-Year College Mathematics Journal, 14(5):427–431, November 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1983.11972733.

Wenger:1974:LE

[Wen74] John C. Wenger. Letters to the Editor. Two-Year College Mathematics Journal, 5(3):9–11, September 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1974.
11974326.

West:1994:NGE

[Wes94] Beverly West. A note from the guest Editor. College Mathematics Journal, 25(5):362–363, November 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1994.
11973639.

Wetzel:2010:WTF

[Wet10] John E. Wetzel. When are two figures congruent? College Mathematics Journal, 41(3):193–196, May 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.

tandfonline.com/doi/abs/10.4169/074683410X488665.

Weyenberg:2023:DDK

[Wey23] Grady Weyenberg. Do dogs know calculus with early transcendentals? College Mathematics Journal, 54(1):64–67, 2023.

CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2022.2160615.

Wood:1975:Rb

[WH75] June P. Wood and Betty J. Hinman. Reviews. Two-Year College Mathematics Journal, 6(2):28, May 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1975.
11974375.

Wood:1976:Rc

[WHB^{+}76] June P. Wood, Betty J. Hinman, Donald E. Brown, Mary Ann De Vincenzo, and Ralph Mansfield. Reviews. Two-Year College Mathematics Journal, 7(3):38–40, September 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1976.11974445.

Wood:1974:Ra

[WHC^{+}74] June P. Wood, Betty J. Hinman, Eugene P. Cooper, H. Eugene Hall, Amogene F. DeVaney, S. C. Tefteller, and Alan C. Tucker. Reviews. Two-Year College Mathematics Journal, 5 (3):41–45, September 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1974.11974336.

Wood:1974:Rb

[WHD^{+}74] June P. Wood, Betty J. Hinman, Cameron Douthitt, Don Gallagher, Shelba Morman, and Peter A. Lindstrom. Reviews. Two-Year College Mathematics Journal, 5(4):31–34, December 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1974.11974349.

Wood:1978:Rd

[WHDS78] June P. Wood, Betty J. Hinman, Mary Ann De Vincenzo, and Marilyn F. Semrau. Reviews. Two-Year College Mathematics Journal, 9(5):288–293, November 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1978.
11974594.

Wheeler:1986:MH

[Whe86] David Wheeler. Media highlights. College Mathematics Journal, 17(4):366–375, September 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1986.11972983.

Wood:1977:Rd

[WHFM77] June P. Wood, Betty J. Hinman, David Farnsworth, and Samiha Mourad. Reviews. Two-Year College Mathematics Journal, 8 (4):231–235, September 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1977.11974516.

Wood:1977:Rb

[WHGR77] June P. Wood, Betty J. Hinman, Leland D. Graber, and Dennis M. Rodriguez. Reviews. Two-Year College Mathematics Journal, 8(2):88–89, March 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1977.11974488.

Wright:1974:R

[WHH^{+}74] Edward B. Wright, Frank Hacker, Harry L. Hancock, Judith Gersting, and Allen R. Angel. Reviews. Two-Year College Mathematics Journal, 5(1):60–66, Winter 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1974.
11974294.

Wood:1976:Ra

[WHH^{+}76] June P. Wood, Betty J. Hinman, Nancy Holder, James G. Troutman, and Ken Seydel. Reviews. Two-Year College Mathematics Journal, 7(1):24–26, February 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1976.
11974416.

Whitton:1971:GFT

[Whi71] Etta Mae Whitton. A general formula for the N th term of a sequence. Two-Year College Mathematics Journal, 2(2):96–98, Autumn 1971. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1971.11974020.

Whitt:1985:MCF

[Whi85a] Lee Whitt. Medical cozenage on Fermat’s Last Theorem. College Mathematics Journal, 16(1):55–56, January 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1985.
11972851.

Whittaker:1985:NDA

[Whi85b] Joseph Whittaker. A new divisibility algorithm. College Mathematics Journal, 16(4):268–276, September 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1985.
11972893.

White:1988:TCM

[Whi88] James E. White. Teaching with CAL: a mathematics teaching and learning environment. College Mathematics Journal, 19 (5):424–443, November 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1988.11973151.

Whitcher:2011:BC

[Whi11] Ursula Whitcher. Boundary conditions. College Mathematics Journal, 42(1):56, January 2011. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.42.1.056.

Wood:1976:Rd

[WHL^{+}76] June P. Wood, Betty J. Hinman, Peter A. Lindstrom, Marilyn F. Semrau, and Suzanne Brown (Donald). Reviews. Two-Year College Mathematics Journal, 7(4):26–28, December 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1976.11974459.

Wood:1978:Rb

[WHLB78] June P. Wood, Betty J. Hinman, Jean Lane, and Donald E. Brown. Reviews. Two-Year College Mathematics Journal, 9(3): 167–170, June 1978. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1978.11974570.

Wood:1977:Ra

[WHLD77] June P. Wood, Betty J. Hinman, Peter Lindstrom, and Mary Ann DeVincenzo. Reviews. Two-Year College Mathematics Journal, 8(1):35–37, January 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1977.11974478.

Wood:1977:Rc

[WHMB77] June P. Wood, Betty J. Hinman, John McGregor, and Suzanne Brown. Reviews. Two-Year College Mathematics Journal, 8(3): 165–167, June 1977. CODEN ???? ISSN 0049-4925 (print),

2325-9116 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/00494925.1977.11974503.

Wood:1977:Re

[WHMS77] June P. Wood, Betty J. Hinman, David E. Moxness, and Jean B. Smith. Reviews. Two-Year College Mathematics Journal, 8 (5):281–285, November 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1977.11974529.

Wood:1978:Ra

[WHN78] June P. Wood, Betty J. Hinman, and John Niman. Reviews. Two-Year College Mathematics Journal, 9(2):84–89, March 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1978.11974553.

Wood:1976:Rb

[WHPT76] June P. Wood, Betty J. Hinman, John P. Pace, and Wesley W. Tom. Reviews. Two-Year College Mathematics Journal, 7(2):40– 42, May 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1976.11974428.

Wood:1975:Rc

[WHS^{+}75a] June P. Wood, Betty J. Hinman, Jean B. Smith, Cherry Mauk, R. DeJean, and Terral McKellips. Reviews. Two-Year College Mathematics Journal, 6(3):27–30, September 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1975.11974389.

Wood:1975:Ra

[WHS^{+}75b] June P. Wood, Betty J. Hinman, Ara B. Sullenberger, Lawrence Gilligan, and Thomas L. Alexander. Reviews. Two-Year College Mathematics Journal, 6(1):27–29, February 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1975.
11974363.

Wood:1978:Rc

[WHS78] June P. Wood, Betty J. Hinman, and Frank Swetz. Reviews. Two-Year College Mathematics Journal, 9(4):222–227, September 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1978.11974583.

Wood:1975:Rd

[WHT^{+}75] June P. Wood, Betty J. Hinman, Lawrence A. Trivieri, Gerald M. Smith, and Bruce King. Reviews. Two-Year College Mathematics Journal, 6(4):17–20, December 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1975.
11974401.

Wiebe:1975:GTPa

[Wie75a] Richard Wiebe. Gödel’s Theorem (Part I). Two-Year College Mathematics Journal, 6(2):13–17, May 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1975.
11974371.

Wiebe:1975:GTPb

[Wie75b] Richard Wiebe. Gödel’s Theorem (Part II). Two-Year College Mathematics Journal, 6(3):4–7, September 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1975.
11974383.

Wiener:1983:EID

[Wie83] Joseph Wiener. Evaluating integrals by differentiation. Two-Year College Mathematics Journal, 14(2):168–169, March 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1983.11972675.

Wiener:1984:PSN

[Wie84] Joseph Wiener. Pythagorean systems of numbers. College Mathematics Journal, 15(4):324–326, September 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972791.

Wiener:1985:BIN

[Wie85] Joseph Wiener. Bernoulli’s Inequality and the number e. College Mathematics Journal, 16(5):399–400, November 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1985.11972916.

Wiener:1986:AAP

[Wie86] Joseph Wiener. An algebraic approach to partial fractions. College Mathematics Journal, 17(1):71–72, January 1986. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1986.11972932.

Wiener:1987:SME

[Wie87] Joseph Wiener. A shorter, more efficient proof of lim_{n→∞} ^{n} \sqrt{n!}/n = e^{−1}. College Mathematics Journal, 18(4):319, September 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1987.11973052.

Wieschenberg:1990:BEC

[Wie90] Agnes Arvai Wieschenberg. The birth of the eötvös competition. College Mathematics Journal, 21(4):286–293, September 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1990.11973321.

Wiener:2000:I

[Wie00] Joseph Wiener. Integrals of cos^{2n} X and sin^{2n} X. College Mathematics Journal, 31(1):60–61, January 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2000.
11974112.

Wiener:2001:DRP

[Wie01] Joseph Wiener. Differentiation with respect to a parameter. College Mathematics Journal, 32(3):180–184, May 2001. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2001.11921873.

Wilson:1972:DSL

[Wil72] Peter M. Wilson. Do students learn from and like an audiotutorial course in freshman mathematics? Two-Year College Mathematics Journal, 3(2):37–41, Autumn 1972. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1972.
11974215.

Williams:1973:MLS

[Wil73] Ralph C. Williams. The mathematics laboratory and the single student. Two-Year College Mathematics Journal, 4(1):40–47, Winter 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1973.11974234.

Wilansky:1982:SN

[Wil82a] A. Wilansky. Smith numbers. Two-Year College Mathematics Journal, 13(1):21, January 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1982.11972573.

Williams:1982:MA

[Wil82b] Gareth Williams. Mathematics in archaeology. Two-Year College Mathematics Journal, 13(1):56–58, January 1982. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1982.
11972579.

Wilf:1990:SEM

[Wil90] Herbert S. Wilf. Self-esteem in mathematicians. College Mathematics Journal, 21(4):274–277, September 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1990.
11973319.

Willson:1991:TAF

[Wil91] Stephen J. Willson. Teaching about fractals. College Mathematics Journal, 22(1):56–59, January 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1991.
11973360.

Wilson:1993:PP

[Wil93] John H. Wilson. Permutation puzzles. College Mathematics Journal, 24(2):163–165, March 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1993.11973523.

Wilson:1994:NOP

[Wil94] Curtis Wilson. Newton’s orbit problem: a historian’s response. College Mathematics Journal, 25(3):193–200, May 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1994.11973607.

Wilkin:1996:SR

[Wil96] Jon Wilkin. Software reviews. College Mathematics Journal, 27(1):68–73, January 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1996.11973753.

Wilkin:1998:SR

[Wil98a] Jon Wilkin. Software reviews. College Mathematics Journal, 29(1):62–65, January 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1998.11973917.

Wilkinson:1998:NPC

[Wil98b] Steven Wilkinson. Numerically parametrizing curves. College Mathematics Journal, 29(2):104–119, March 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1998.
11973923.

Wilkinson:1998:SR

[Wil98c] Steven Wilkinson. Software reviews. College Mathematics Journal, 29(4):323–329, September 1998. CODEN ????

ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1998.
11973967.

Wilkin:1999:SR

[Wil99a] Jon Wilkin. Software reviews. College Mathematics Journal, 30(1):54–59, January 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1999.11974032.

Will:1999:BDS

[Wil99b] Todd G. Will. The bus driver’s sanity problem. College Mathematics Journal, 30(3):187–194, May 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1999.
11974054.

Willoughby:2002:WGC

[Wil02a] Keith A. Willoughby. Winning games in Canadian football: a logistic regression analysis. College Mathematics Journal, 33(3): 215–220, May 2002. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2002.11921944.

Wilson:2002:ANI

[Wil02b] Robin Wilson. Alice in Numberland — an informal dramatic presentation in 8 fits. College Mathematics Journal, 33(5):354– 377, November 2002. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2002.11921966.

Wilkins:2003:TLC

[Wil03] Daniel Wilkins. The tangent lines of a conic section. College Mathematics Journal, 34(4):296–303, September 2003. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2003.11922021.

Wilson:2016:FR

[Wil16] Jennifer Wilson. Film review. College Mathematics Journal, 47(5):381–385, November 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.4169/college.math.j.47.5.381.

Will:2024:PPA

[Wil24a] Todd G. Will. Pentagons with permutable angles. College Mathematics Journal, 55(4):336–342, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2023.
2279459.

Williams:2024:UGN

[Wil24b] Kenneth S. Williams. Using the Girard–Newton–Waring formula to evaluate trigonometric sums: a blast from the past. College Mathematics Journal, 55(5):416–427, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic).

Wimp:2002:BR

[Wim02] Jet Wimp. Book review. College Mathematics Journal, 33 (4):341–344, September 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2002.11921963.

Winston:1991:CE

[Win91] Elliot Winston. A carpenter’s ellipse. College Mathematics Journal, 22(4):311–312, September 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1991.
11973401.

Winter:1995:FVD

[Win95] Mary Jean Winter. Finding volumes with the definite integral: a group project. College Mathematics Journal, 26(3):227–228, May 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1995.11973703.

Wisner:1999:PUS

[Wis99] Robert J. Wisner. Powers as uniform sums of positive squares. College Mathematics Journal, 30(4):293–296, September 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1999.11974073.

Withers:2017:TIS

[Wit17] Wm. Douglas Withers. Triple integrals for the sketchingimpaired. College Mathematics Journal, 49(1):46–53, January 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2017.1394669.

Wagner:1973:GBI

[WJ73] John Wagner and Howard Jones. Group-based instruction: The best chance for success? Two-Year College Mathematics Journal, 4(1):51–54, Winter 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/00494925.1973.11974236.

Woods:1973:PS

[WK73] Dale Woods and Mary Jane Kohlenberg. Pascal’s k-simplex. Two-Year College Mathematics Journal, 4(3):38–43, Autumn 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1973.11974271.

Westfall:1994:DNH

[WK94] Richard S. Westfall and Sidney H. Kung. In defense of Newton: His biographer replies. College Mathematics Journal, 25(3):201– 205, May 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1994.11973608.

Williams:1976:RDS

[WM76] Eddie R. Williams and Harold W. Mick. Research and development of synchronized slide-tape units for a mathematics laboratory. Two-Year College Mathematics Journal, 7(2):28–33, May 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1976.11974426.

Walsh:2013:MCD

[WM13] Jim Walsh and Richard McGehee. Modeling climate dynamically. College Mathematics Journal, 44(5):350–363, November 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.44.5.350.

Weinstock:1994:INC

[WN94] Robert Weinstock and Roger B. Nelsen. Isaac Newton: Credit where credit won’t do. College Mathematics Journal, 25(3):179– 192, May 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1994.11973606.

Wampler:1996:IP

[WN96] Joe F. Wampler and Stephen E. Newman. Integer programming. College Mathematics Journal, 27(2):95–100, March 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1996.11973758.

Woll:1971:OSS

[Wol71] John W. Woll, Jr. One-sided surfaces and orientability. Two-Year College Mathematics Journal, 2(1):5–18, Spring 1971. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1971.11973990.

Wolper:2023:SDE

[Wol23] James S. Wolper. Not solving differential equations. College Mathematics Journal, 54(4):337–344, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2023.
2240204.

Wong:2000:OUA

[Won00] Roman W. Wong. The orbits of a unimodular affine transformation. College Mathematics Journal, 31(4):290–296, September 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2000.11974161.

Wood:1970:FYM

[Woo70] June P. Wood. First-year mathematics — a challenging variable. Two-Year College Mathematics Journal, 1(1):8–13, Spring 1970. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1970.11973159.

Wood:1972:LE

[Woo72a] June P. Wood. Letters to the Editor. Two-Year College Mathematics Journal, 3(2):8, Autumn 1972. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1972.
11974207.

Wood:1972:STS

[Woo72b] June P. Wood. “sample” tests for students. Two-Year College Mathematics Journal, 3(1):14–15, Spring 1972. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL

http://www.tandfonline.com/doi/abs/10.1080/00494925.
1972.11974190.

Woo:1975:GB

[Woo75] Norman Woo. On generalized h-base. Two-Year College Mathematics Journal, 6(3):16–17, September 1975. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1975.
11974386.

Woodriff:1984:DDI

[Woo84] R. L. Woodriff. A distinctly discrete indiscretion …. College Mathematics Journal, 15(5):376–377, November 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972807.

Woo:2000:SCG

[Woo00] Peter Y. Woo. Straightedge constructions, given a parabola. College Mathematics Journal, 31(5):362–372, November 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2000.11974175.

Wood:2025:PWE

[Woo25] William E. Wood. Proof without words: Estimating generalized pi. College Mathematics Journal, 56(1):60–61, 2025. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic).

Weidman:1990:RPC

[WP90] P. D. Weidman and K. Pfendt. On the radial packing of circles in the plane. College Mathematics Journal, 21(2):112–120, March 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1990.11973293.

Widdis:1989:MMT

[WR89] Daniel B. Widdis and R. Bruce Richter. It’s magic! Multiplication theorems for magic squares. College Mathematics Journal, 20(4):301–306, September 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1989.11973246.

Wright:1989:NPT

[Wri89] Tommy Wright. A note on Pascal’s Triangle and simple random sampling. College Mathematics Journal, 20(1):59–66, January 1989. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1989.11973209.

Wright:1995:PTG

[Wri95] Tommy Wright. Pascal’s Triangle gets its genes from Stirling numbers of the first kind. College Mathematics Journal, 26 (5):368–371, November 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1995.11973731.

Wirkus:1998:HPS

[WRR98] Stephen Wirkus, Richard Rand, and Andy Ruina. How to pump a swing. College Mathematics Journal, 29(4):266–275, September 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1998.11973953.

Waits:1973:NLO

[WS73] Bert K. Waits and Jerry L. Silver. A new look at an old work problem. Two-Year College Mathematics Journal, 4(3):52–55, Autumn 1973. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1973.11974273.

Wheeler:1981:SFT

[WS81] Dan Wheeler and David Sklar. A space-filling torus. Two-Year College Mathematics Journal, 12(4):246–248, September 1981. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1981.11972539.

White:1996:RCP

[WS96] Gary D. White and Kirby Smith. The rental car problem. College Mathematics Journal, 27(5):374–378, November 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1996.11973811.

Wexler:2005:MB

[WS05] Anthony Wexler and Sherman Stein. Making a bed. College Mathematics Journal, 36(3):213–221, May 2005. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2005.
11922131.

Wu:2023:PWG

[Wu23a] Rex H. Wu. Proof without words: a geometric series through the dissection of regular hexagons. College Mathematics Journal, 54(1):62, 2023. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2022.2153563.

Wu:2023:PWS

[Wu23b] Rex H. Wu. Proof without words: The sum of even powers of three. College Mathematics Journal, 54(2):139, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2023.2189821.

Wu:2024:PWA

[Wu24] Rex H. Wu. Proof without words: an alternating geometric series. College Mathematics Journal, 55(3):246, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/07468342.2023.2283364.

Wiener:1988:CAI

[WW88] Joseph Wiener and Will Watkins. A classroom approach to involutions. College Mathematics Journal, 19(3):247–250, May 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1988.11973121.

Weber:1995:PC

[WW95] Griffin Weber and Glenn Weber. Pizza combinatorics. College Mathematics Journal, 26(2):141–143, March 1995. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1995.
11973685.

Wade:2000:RPP

[WW00] Peter W. Wade and William R. Wade. Recursions that produce Pythagorean triples. College Mathematics Journal, 31(2):98–101, March 2000. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2000.11974121.

Watkins:2020:CMJ

[WW20] Ann E. Watkins and William Watkins. The College Mathematics Journal during the 1990s. College Mathematics Journal, 51(3):216–217, 2020. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2020.1733327.

Whitt:1983:AWP

[WWW83] L. Whitt, K. Wilk, and Daniel H. Wagner. Antisubmarine warfare: Passive vs. active sonar. Two-Year College Mathematics Journal, 14(5):434–435, November 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1983.
11972735.

Xu:2010:CQF

[Xu10] Conway Xu. A characterization of a quadratic function in R^{n}. College Mathematics Journal, 41(3):212–213, May 2010. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/074683410X488692.

Xu:2014:PQF

[Xu14] Connie Xu. A proof for a quadratic function without using calculus. College Mathematics Journal, 45(2):143–144, March 2014. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.45.2.143.

Yandl:1980:DEP

[Yan80] André L. Yandl. Delta, epsilon, and polynomials. Two-Year College Mathematics Journal, 11(4):263–266, September 1980. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1980.11972453.

Yandl:1985:NIP

[Yan85] André L. Yandl. A note on integration by parts. College Mathematics Journal, 16(4):282–283, September 1985. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1985.11972898.

Yackel:2007:PFC

[YD07] C. A. Yackel and J. K. Denny. Partial fractions in calculus, number theory, and algebra. College Mathematics Journal, 38 (5):362–374, November 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2007.11922261.

Yerion:1997:GED

[Yer97] Kathie Yerion. Gaussian elimination and dynamical systems. College Mathematics Journal, 28(2):89–97, March 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1997.11973842.

Ying:1987:WRB

[Yin87] Wong Ngai Ying. Why is a restaurant’s business worse in the owner’s eyes than in the customers’? College Mathematics Journal, 18(4):315–316, September 1987. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1987.
11973049.

Yiu:2012:GSD

[Yiu12] Paul Yiu. Geometry of sum-difference numbers. College Mathematics Journal, 43(5):408–409, November 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
43.5.408.

Yates:1990:CCC

[YMH90] Daniel S. Yates, John H. Mathews, and David R. Hill. Classroom computer capsules. College Mathematics Journal, 21(1):51–56, January 1990. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1990.11973285.

Young:1970:PBM

[You70] Gail S. Young. Providing a better mathematics education. Two-Year College Mathematics Journal, 1(1):4–7, Spring 1970. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1970.11973158.

Young:1974:CC

[You74] Frederick H. Young. Computer corner. Two-Year College Mathematics Journal, 5(3):55–57, September 1974. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1974.
11974339.

Young:1993:LCC

[You93] Donald Francis Young. Local conditions for convexity and upward concavity. College Mathematics Journal, 24(3):224–228, May 1993. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1993.11973532.

Young:2006:TT

[You06] Donald Francis Young. The tippy trough. College Mathematics Journal, 37(3):205–213, May 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.2006.
11922186.

Young:2024:CMB

[You24] Cooper Young. Cutting the Möbius strip. College Mathematics Journal, 55(2):126–133, 2024. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/07468342.2023.2271827.

Zacharias:2018:PCM

[Zac18] John Zacharias. Proof of a conjecture of Merca on an average of square roots. College Mathematics Journal, 49(5):342– 345, November 2018. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2018.1502997.

Zahn:1972:DCT

[Zah72] Karl G. Zahn. A design for class testing mathematics textbook materials. Two-Year College Mathematics Journal, 3(2):29–32, Autumn 1972. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1972.11974213.

Zames:1977:SAC

[Zam77] Frieda Zames. Surface area and the cylinder area paradox. Two-Year College Mathematics Journal, 8(4):207–211, September 1977. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1977.11974511.

Zaslavsky:1970:MYP

[Zas70] Claudia Zaslavsky. Mathematics of the Yoruba people and of their neighbors in Southern Nigeria. Two-Year College Mathematics Journal, 1(2):76–99, Autumn 1970. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1970.
11973191.

Zeeman:2013:MSB

[Zee13] Mary Lou Zeeman. Mathematics, sustainability, and a bridge to decision support. College Mathematics Journal, 44(5):346– 349, November 2013. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.44.5.346.

Zehnder:1996:SR

[Zeh96] Alan T. Zehnder. Software reviews. College Mathematics Journal, 27(2):144–148, March 1996. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1996.11973768.

Zemgalis:1971:OSI

[Zem71] Elmar Zemgalis. On one-sided inverses of matrices. Two-Year College Mathematics Journal, 2(1):45–48, Spring 1971. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1971.11973997.

Zeng:1996:NBP

[Zen96] Jim Zeng. A note on the brachistochrone problem. College Mathematics Journal, 27(3):206–208, May 1996. CODEN ????

ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1996.
11973778.

Zerger:2002:VS

[Zer02] Monte J. Zerger. A visit with six. College Mathematics Journal, 33(2):74–87, March 2002. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2002.11921923.

Zizler:1997:ESV

[ZF97] Peter Zizler and Holly Frase. Eigenpictures and singular values of a matrix. College Mathematics Journal, 28(1):59–62, January 1997. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1997.11973836.

Zhou:2006:SF

[ZGN06] Ying Zhou, Walter Gall, and Karen Nabb. Synchronizing fireflies. College Mathematics Journal, 37(3):187–193, May 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2006.11922184.

Zhou:2012:TTW

[Zho12a] Li Zhou. Teaching tip: When does f(g(x)) = x? College Mathematics Journal, 43(4):290, September 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
43.4.290.

Zhou:2012:VPF

[Zho12b] Li Zhou. Viviani polytopes and Fermat points. College Mathematics Journal, 43(4):309–312, September 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
43.4.309.

Zhong:2017:EAM

[Zho17a] Wenti Zhong. On the eigenvalues of anticommuting matrices. College Mathematics Journal, 48(5):368–369, November 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.48.5.368.

Zhou:2017:DEK

[Zho17b] Li Zhou. Did Elvis know Cauchy–Schwarz? College Mathematics Journal, 48(5):335–338, November 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
48.5.335.

Zhou:2018:PWE

[Zho18] Li Zhou. Proof without words: Elvis trades running for swimming. College Mathematics Journal, 49(5):366, November 2018. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/07468342.2018.1526576.

Zia:1991:UFD

[Zia91] Lee Zia. Using the finite difference calculus to sum powers of integers. College Mathematics Journal, 22(4):294–300, September 1991. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1991.11973398.

Zirkel:1976:CLC

[Zir76] Gene Zirkel. Closing the loopholes in “closing the loopholes”. Two-Year College Mathematics Journal, 7(3):55–56, September 1976. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1976.11974450.

Zitarelli:2005:BAM

[Zit05] David E. Zitarelli. The bicentennial of American mathematics journals. College Mathematics Journal, 36(1):2–15, January 2005. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2005.11922105.

Zizza:1998:DFC

[Ziz98] Frank Zizza. Differential forms for constrained max-min problems: Eliminating Lagrange multipliers. College Mathematics Journal, 29(5):387–396, November 1998. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/07468342.1998.
11973974.

Zizler:2013:QTS

[Ziz13]  Peter Zizler. Quiz today: Should I skip class?  College

Mathematics Journal, 44(3):166–170, May 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
44.3.166.

Zobitz:2013:FCU

[Zob13] John Zobitz. Forest carbon uptake and the fundamental theorem of calculus. College Mathematics Journal, 44(5):421–424, November 2013. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.44.5.421.

Zucker:2009:FT

[Zuc09] Marc Zucker. Flipping triangles! College Mathematics Journal, 40(3):189–195, May 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.2009.11922359.

Zucker:2016:LPH

[Zuc16] Marc Zucker. Lattice paths and harmonic means. College Mathematics Journal, 47(2):121–124, March 2016. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
47.2.121.

Zager:2007:CRW

[ZV07] Laura A. Zager and George C. Verghese. Caps and robbers: What can you expect? College Mathematics Journal, 38(3):185–191, May 2007. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2007.11922235.